A Complete Bibliography of Publications in the British Journal for the Philosophy of Science: 1950–date

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

25 May 2024 Version 1.51

Title word cross-reference

−kTr(ρ ln ρ) [She99]. $25.00 [CE04]. $33.95 [Lyo06a]. $48/£ [Per07]. $58.00 [Bun86]. $65.00 [Gly04]. $95 [Pro86]. _{∞} [JHB^{+}23]. P (h|b) [Row13]. Σ_{1}^{0} [Ber23].

0 [Bos06, Bri06, CE04, Gia06b, Gil07, Gly04, Hoo86, Lyo06a, She06, Tau06]. 0-09-158010-2 [O’H86]. 0-19-503215-2 [Rub86]. 0-19-515527-0 [Gly04]. 0-19-875254-7 [Bri06]. 0-19-924618-1 [CE04]. 0-19-926850-9 [Gia06b]. 0-262-19215-2 [Hoo86]. 0-300-03174-2 [Bai86]. 0-415-24203-7 [Lyo06a]. 0-415-31128-4 [Bos06]. 0-471-90320-5 [Hog86]. 0-521-24256-8 [Rai86]. 0-521-24803-5 [Roc86]. 0-521-26530-4 [Kör86]. 0-521-26547-9 [Gib86]. 0-521-28973-4 [Roc86]. 0-521-55372-5 [Tau06]. 0-631-18887-8 [She06]. 0-7546-0678-3 [Gil07]. 0-8204-0118-8 [Den86].

1 [CE04, Gia06a, Per07]. 1-4020-0226-2 [Fes06]. 1-4020-2080-5 [Gia06a]. 1-86094-500-7 [Per07]. 10.95 [Hea86]. 1736 [Far00]. 17th [PP78]. 18.99/US [CE04]. 18th [PP78]. 1926 [Bor54]. 1952 [Ecc52]. 1997 [Dre00].

2 [Hoo86]. 20th [Mar00]. 25s [Qui61]. 294 [Coh86a]. 2nd [Lyo06a].

3 [Gil07]. 301 [Fie86].

4 [Bos06]. 418pp [Gly04]. 4A [Rai86].

5 [Gia06a, Tau06].

6 [Gil03].

7 [Bri06, Lyo06a, Per07]. 70th [AGTZ08].

8 [She06]. 88 [Gil03]. 88-7018-386-6 [Gil03].

9 [Gia06b]. 90 [D’A86]. 90-277-1760-5 [Cap86]. 90-6765-020-X [D’A86]. 90/US$143/EUR [Fes06]. 920 [Pro86]. 9O [Bro86]. 9O-277-1648-X [Bro86].

Abduction [Wil03]. Abiding [WS23]. Aboutness [Sob85]. Absolute [Dew24, Dor78, Hoe98, Hug08, Hug81, Lac70, Lis04, Pap52, Sto86, Wei01a, Zar79]. Absolutism [Hin88]. Absolutus [Wes79a, Wes79b]. Absorber [Pri94]. Abstract [Lig14, LS94, Pin15]. Abstraction [Bla20, Coo12, Grü53b, LU09, Rob20, Sch57, Sha03, Stu16, Fin02, CE04]. Abstractions [JS19]. Absurd [Blo70]. Abuses [Mcl90]. Academic [Fes06, Gia06a, PP78]. Accelerated [Kes85]. Acceptability [Dou02, She02]. Acceptable [LU09]. Acceptance [Lev63, Wil69]. ’accessible [CM00]. Accommodation [Har06, HS04, Lan01]. According [Vai15, BL23, Wal12]. Account [Cas16, Cha79c, Dor19, Geb22, Gim04, Kin10, Kos21, Leu12, Leu14, MSM09, Pea12, Puc76, Sma79, SP12, Wil23a, Woo14, Gel59]. Accounting [Yud13]. Accounts [Bet24, VH03]. Accuracy [Hor78a, Kru97, McC23, Odd19, Sch22b, VC24]. Accuracy-Based [VC24]. Accurate [CMS22, FS94, Ngu21]. Achievement [Lev14]. Achilles [Bar62, Bro55, Gru66]. Achinstein [Gim04, Rub86]. Acknowledgements [Ano13]. Acquisition [Sam80]. across [MP22, Sch15, War22]. Action [Ada01, Gil05, Hut51d, Mac69d, Mun89, Wal07]. Action-Related [Gil05]. Activism [Grü56]. Actual [DE76, FG17, Fis24, Ste16b]. Actuality [You86]. ’Ad [Pop59b, Bam93, FS94, Grü76a, Pop59a, Red78]. ’Ad-Hocnes [Pop59b]. Ad-Hocness [Pop59a]. Adam [Her07]. Adams [NR13]. Adaptation [Cla53, Gri96]. Adaptationism [Sar15]. Adaptations [Nea95a]. Added [McC84]. Additivity [Bar04, How08, Wil99]. Adequacy [Ket04, Mil18, Pop56a]. Adequate [Ten85]. Adolf [Cra92]. Advancement [Wat75]. Advantages [Fri97a]. Aesthetic [Cur23, Mar89a]. Aether [Bor54, Pro63]. affirmations [Ueb99]. Afloat [Mul05a]. After [MG17, Pos60a, Ste01]. Again [Bri82, Lee95, Put05]. Against [Bak09b, But06b, Car02b, Cha89, Jos82, Koe72, KS19, Lar00, Loc03, Mol22a, Res13, RH72, Sta09, SP12, Sto86, Urb78, Wri68, Dod13, Fie86, Háj08, PP78, Ram13, Rig91, Sti83]. Agassi [Ale59a, Wil78a, Aga78]. Age [Abr54, Akk94, Dav54b, Dav56, Gra55, Hut55a, Koe00, Öpi54, Poo55, Sch54a, Scr54, Whi54a, Whi55a]. Agency [Pri91a]. Agent [FS21, Kuy24, Mar17]. Agent-Based [FS21]. Aggregation [War22]. Aharonow [Liu94]. Aid [Aga59b, Hue09]. Aim [Max93c, Max93b]. Aim-oriented [Max93c, Max93b]. Aims [Coh63a, Jar61a]. Ain’t [Lau89]. Air [Jen88]. Air-Pump [Jen88]. Akaike [Kie97]. Al [Hog86]. Al-Daffa [Hog86]. al. [Hub08]. Alan [Jon69a]. Albert [Cal15, Cou52a]. Alchemy [Mad86]. Alessandro [Cam00]. Alex [Joh15]. Alexander [Lip99]. Alferov [Jos10]. Alfred [C.56b, C.56d]. Algebraic [Lup18]. Algorithm [CB23]. algorithms [Fal00]. Ali [Hog86, Sla15a]. Alicia [Ada01]. Alignment [Bro20]. Alisa [Ber10]. Alleged [Coh02, Die67, Eel88, Ker87b, Lon53]. Allen [PS00]. Allies [IG95]. Allowed [Wil75b]. Alloys [GBW61b]. already [Dor78].

Alternative [Puc76, Spo06]. Alternatives [Bar19b, DHS15, Egg16, Sta06a, Sta09, Sta06b, Enf08]. Altruism [HL24, Oka05, Ses99, Ste96a]. Alvin [Fel01, Gan11]. Always [Tor87, Irz85]. Alwyn [Pro86]. Alyssa [Cal15]. Ambiguity [Fre11, Gim04]. Amendment [Wat71a]. Among [Bla79]. Amplification [FN21]. Amplified [Led87]. Analogical [Hor94, She02]. Analogies [Kro89, Nyr21]. Analogue [Bec15, Cav10, DTW17, Mal23]. Analogy [Bun68a, Dor71, Hes52, Mar17, She02, Tur55]. Analyses [FP66]. Analysing [Bec19]. Analysis [Bai87, Bub94, Cle79, Dea76, Geb22, Gia94, Gly11b, Gri93, HF85, Hut55d, KLM10, LRP23, Mac93, Mad57, Mul91, Mül05b, Oka04, Oka08, RO55, RKK19, Sch59b, Sch18, Smy54, Ste16b, Urb92, Van08b, VC24, Wri18, Lil99, Pla00, May00]. Analytic [Ezr56, Sch64a, Spi57, Wis56a, Wis63a]. Analytical [Fai56]. Anarchy [Van08b]. Anatomy [Van94a, BER99]. Ancestry [SO03, SS17]. Ancient [Woo00b]. Andersen [Tha09]. André [Sid01, Pag00]. Andreas [Sch08a]. Andrew [Fod01, Tau06]. Angel [Lar03]. Angles [Why52a]. Animal [Cla53, Dac23, Sin00]. Anjan [Cho11]. Anjum [Cha13]. Anniversary [Bir12]. Annual [Den59a, Den60a, HH63, HH64]. Anomalous [Kle92]. Anomaly [Blu89]. Answer [Boh62, Grü76b, Grü86, Urb88]. Anthropic [Cra88, Wil94]. Anthropocene [San19]. Anthropological [Tri94]. Anthropology [Bro64b, Jar61a]. Anti [Dal89, Die99, FA14, Obe98, SE23, Ueb96]. Anti-Foundationalism [Obe98, Ueb96]. Anti-Platonism [Die99]. Anti-Realism [Dal89, FA14]. Anti-reductionist [SE23]. Antimatter [BH10]. Antipositivism [Gie75]. Antirealism [Hel87]. Antirepresentationalism [Gar08]. Any [Car02b, Ern90, Sha03, Urb78]. Anything [HL24]. Anyway [BE21]. Apparent [Gar73a, Lan01, Mil18, Wat51]. Appearances [Ben12, Van94a]. Applicability [DE76, Luc91, Sim01]. Application [Puc74, Vas21a]. Applied [Bil00, Smi07b]. apply [GB99]. Appraisal [Nia05, Red78]. Appraising [Fes06, KKS02]. Apprehension [Pri60]. Apprentice [Lew14, Ste12b]. Approach [Alb94, Bar85a, Bel86, BM18, Bou15, Bra93, Bux78, For95a, Fre88, Gir76, Goo84, HP05a, HP05b, Hea12, JHB^{+}23, Joz86, Kri18, Oka16, Pot10, Rad87, Sac64, Ten94, Vol95, War22]. Approaches [Far16a]. Appropriate [GP14]. Approximate [Mac94b, Pit11, Smi98]. Arabia [Hog86]. Arbitrariness [Gil14]. Arbitrary [Von83b, dC22]. Arbitration [Set90]. Arch [IG95]. Arched [Tho94a]. Architecture [Ber16, GS08a]. Architectures [Wel93]. Aren’t [PR95]. Argument [Ara11, Bai73, Bla66, But06a, Coa74, DHS15, Dod13, Dor71, DH86, Gla84, Har60, Hau94, Hei78, Kle02, Kuk08, Lac70, LM01, Le 05, Lin23, Mas01, Pit08, Pop58b, Rig91, Ryd81, Ryn94, Saa11, Sch08c, Van88, Wea18, Wei05, Wil17, WHB^{+}23b]. Arguments [Bet16, Cra01, Háj08, Hea02, Kot12, Loc03, Pre61, Ram13, Urb78]. Aristotle [Gil00a, Mil72, Con03]. Arithmetic [Dem01, Hog04, Jen05, Kre53a, Kre53b, Luc91, Pot02, GB99, Pot00, Ket01]. Armstrong [Bir01, San80]. Aronson [HM99]. arrow [Fri00, Ang02, HS22]. Art [Hac71b, Rue11]. Artefacts [CDCH^{+}24, Mac51, Mac53a, May52, Pir52a, VH03]. Article [Bla63a, Bra63a, Bra63b, Lev97a, Pap91, Dow00, Fol00, Hau99, Mac99, Ses99]. Articles [Din65a, Red77]. Artificial [Bod73, Hel87]. Arturo [Gia06a]. Ascribing [VH03]. Ashgate [Gil07]. Aspect [Kar12, Sch59b]. Aspects [Ben56, Ben57, Bla50, Dav62, Har62, HL11, Hut54b, Kne57b, Mil78, Pod75, Pod76, Sär68]. Assay [Hei99]. Assembling [BSM19, Bar23, BS17a]. Assembly [LaC22]. Assertion [Job80]. Assessment [Chr10, Lep82, McA86, Rat02, Rey09]. Assimilation [Par17]. Assumption [BZ84, Dew73, GPW05, Sch60, Sch61b]. astronomy [Gin84, Rai86]. Astrophysical [Kle02]. Astrophysics [Gin84, Rai86]. Asymmetric [CBC^{+}22]. asymmetries [Hau98, Hit00]. Asymmetry [BS92, Die67, Ehr87, Hut95b, Nor02, Nun93, Pap85, Rob20, vR13, Pap88]. Asymptotic [Mee04]. Asymptotics [Bat02, Hoo04]. Atemporal [Cus90]. Atomic [Buc59, Nyr21, Sta09]. Atomism [Pos75a]. Atoms [Hud20]. Attempted [Grü76c]. Attention [FN21, Tay18]. Attitude [Kuk94]. Attraction [Dri11]. Attractors [Dri11]. Aufbau [Ryc99]. Authority [Pas05]. Authors [ABB^{+}24, Ano86a]. Autobiographical [HHS02]. Autonomous [ARR15, Bel78, Pop55b]. Autonomous-Statistical [ARR15]. Autonomy [Leb89, Rob20]. Auxiliary [FW05, Grü76a, Lai88, Lep82, Str01, Str05]. Auyang [Col00]. Aversion [SB19]. Award [Red89]. Away [Oka01]. Axiom [Pop55b]. Axiomatic [Coh16, Mue69, San61]. Axiomatization [Zah91]. Axioms [CCT13]. Azande [Tri88].

B [Kyb58, Mel90]. Bachelard [Roc86, Til84]. Back [Bri82, Gar23, Mil72]. Background [Fre88, Wat60a, Wat87, PP78]. Backhouse [Bra99]. Backward [Rig91, Sch98]. Backwards [Dup13, Sob11]. Bacon [PR91, Urb82]. Baconianism [Coh80a]. Bacteriophages [MP08]. Bad [Ale16, BCP93, Hau11, Row23, Sch22b]. Balaguer [Die99]. Balance [MTH^{+}24]. Baldwin [Dup00, Wat99]. Balint [Fai57]. Bananas [Smi97]. Bang [Pit08, Smi94]. Bar [Pop55a, Pop56a]. Bar-Hillel [Pop55a, Pop56a]. Barbara [Koe99].

Bargain [Fey81]. Barker [Tau06, Tha09]. Barnard [Dia64]. Barnes [Har11, Mee99]. Barrett [Den94]. Barrow [Cra88]. Barry [Mos86, Cha89]. Bas [Pin11a]. Base [Bir22]. Based [Ell88, FS21, Irv16, MC12, Ner82, RS23, VC24]. Basement [Fey81]. Bases [Mil74b]. Basic [Bae16, Fai56, God69, MS06a, Zah83a]. Basil [Jen86]. Basis [Bar05b, Car02a, Gär84, Hal55a, Hub05, NS20, Pir52b, Urb83, Wil69, Haa91]. Bassi [CM00]. Batterman [Pin11c]. Baumer [Kah67]. Bay [Wis58f]. Bayes [CS12, CEH21, DW17, DGO07, For95a, Geb22, Gly10a]. Bayesian [How97b, Wil14a, BH15, Bux78, Chi87, CS12, CH17, CFM08, DE76, FW05, FW07, Gil91a, Gly09, Goo84, Hol98, HF85, How85, HF94, NS21, Nun93, Ott94, Par16, Riv94, Ros94, Str01, Str05, Urb92, Vra04, Wei01b, Whe12, Wil80b, Wil22, Wil23a, Hub08]. Bayesianism [Coh80a, Dor72, Hen14, How84, Nii83, Ols03, Rai20, Val18]. Bayesians [Bel17a, Bel17b, DGO07, VC24]. Be [Ale18, BFHS02, Cla66, GP14, Gar18, Goo74b, How73, Ism96, Lan06, Odd19, Row23, Sob98, Swa21, Tho65, Wil75b, Bai69, Cal04, FA14, Grü69b, Irz96, Max68, Wor89]. Beast [Gea65, KFH65]. Beat [WOB21]. Beauty [Bra11, Bra15, Gro08, Lun21, Wil14b]. Because [Leh11, Mcd72, TH20]. Becomes [Lav23, WC17]. Becoming [Bou04, Myr03, Pau16]. Been [Kop01, Dor78]. Beer [Pre03]. Before [KK83, Mac86, Mil87a]. Beginning [Cra93]. Begriffsschrift [May80a, May80b]. Behavior [Ada01, Col07]. Behavioral [Gro09, MC12, JLZ05]. Behaviour [Buc23, Erw82, Fri04, Gil93, Gre81, Gre84, Gun68, Mac51, Mac53a, May52, PT57, Pir52a, Rei23, Rus51, OB12, Bir13]. Behavioural [Mac62, McW16]. behind [Lan76]. Behold [Spu52]. Beiden [Zah83a].

Being [Ale18, CEH21, Her23, HK08, Pri76, Ero12]. Bekoff [PS00]. Belief [Gär84, Jen16, Lip02, NR13, Sch15, Val18, Wei93, Sti83, Wil14a, Koe99]. Beliefs [Bot94, Par16, Pol52b]. Believe [Bec04, Sch00b, Wal06a]. Bell [Ang02, Bha91, But92, Cav10, EPW13, GPW05, KP99, Lau95, May00, Pla00, Shi84]. Bell-type [GPW05]. Belot [Vic08]. Ben [Sch14, Wri86]. Ben-Zeev [Wri86]. Benardete [Sha05]. Benditt [Sch76]. Benefits [Bor16]. Benjamin [Vas21b]. Bennett [Kne57b]. Berkeley [Hug92, Hut53a, Pir53, Pop53, Ras53, Rus54, Whi53a, Wis53a, Wis53b]. Berkson [Wil78a]. Bermúdez [Car99b]. Bernoulli [Hac71b]. Bertrand [Bra70, GR15]. Bertuglia [Str09]. Best [Bol83, Dor19, DW17, Hen14, Kat13, Lyo06a, Mag10, Pra24, TP19, TP21, Web09, Woo14, Lip04]. Bets [Hei78]. Better [CK20, CG09, Hoe98, JT17, Lav23]. Betting [Pos60a, Sch22a]. Between [Aga63, BS92, Boh61, Car02b, Coh02, DiS99, Die67, Dob51a, Dob51b, Fri04, Gal03, Gar73a, Gie73a, Gro90, Ker87b, Kör67, Kro89, Mil57, Pop66b, Sti79, Wil73a, BM18, HH92, Kel10, Len06, Max68, Mor86, Oka16, Rat20]. Beyond [Ber10, Cod75, Del21, Gel75, Haa78, Hun06, Kan84, Kut00, SW19b, Bok08b, D’A20, And10, Den86]. Bias [Buc23, Pet22]. Biased [Das23]. Biases [Kot12]. Bicchieri [Gua07]. Bickle [Bon00]. Biconditionals [TH20]. Big [Bae16, BCP93, Tos16, VM16, Mee99, Pit08, Smi94]. Bigger [Str03, Wer10]. Bild [Hor94]. Bind [Per18]. Binmore [Bir13, Sky12]. Biochemical [Bar16, Tah21]. Biography [Wil67]. Biological [Bec19, Cal09, CC93, DiF19, Gan04, Gir76, HL24, Hau13, Hul74, HZ13, HL11, Man97, Mar89b, MSM09, MP22, O’M14, Pra72, Ros01, Rus87, SMM11, Pra12]. Biologically [PNSBC94]. Biology [Bae19, Byr07, Can64, Can65, Col00, Con03, CK89, Cur16, FP66, Gou57, Hul77, Hut60d, Jar67a, Kuy24, Leh65, Mat17, Mic81, Pro86, PS00, Rad00, Rat20, Ros21, Rus69, Ste95, Ste97, UH19, VK91, Wil51, Woo52a, Woo60a, Dup14, Fag13, Gar07, Šus07, Cla14, Gra01, Ste10]. Biometry [Mor02]. Biophysics [Why55b]. Biospecies [Hul70]. Bird [Lip99]. birth [Fie86]. birthday [AGTZ08]. Bisection [Puc73]. Bivalent [Mol22b]. Black [Buc23, Mul05a, Sim91, Sob98]. Blackwell [Jen86, She06]. Blind [Kuh80]. Blindspots [Sca10]. blood [Fag13, Cur16]. Blooded [Cur17]. Bloor [Tri94]. Blue [Wit58b, Wit58c]. Boden [All01]. Bodies [Low00, Smi07b]. Body [Cla72, Hei99, Hor96, Pir50, Urb77, Wat77, Wat51, Wil75a, Wis52a]. Body-Mind [Pir50]. Bogdan [Sto99]. Bohm [Liu94]. Bohmian [AGTZ08, EHLD14, Lew07, Loe98]. Bohm’s [Fey60]. Bohr [Bro86, Bub73, Gar72b, Hen84, Hol86, Hoo91]. Bokulich [Ber10]. Bold [Grü76c, Pic11]. Boltzmann [Pri02]. Bomb [Pri02]. Bondage [Hit03]. Bondi [Hal55b, Kap55b]. Book [Ada01, Aga78, Ale16, Amu99, And10, And86, Bai86, Bak15, BD86a, Bar01, Bat15, Bee09, Bel01, Ber10, Bir13, Bon01, Bon00, Bor54, BG16b, Bos06, BL14, Bra99, Bri06, BS00, Bro86, Bro09b, BF00, Bun86, Byr00, Cal15, Cam00, Cap86, Car12, Car99b, Cha13, Cho11, Cla14, Coh86a, Col00, CS16, CE04, Cou52a, Cur16, D’A86, Dem01, Den86, DiS99, Dic01a, Dic01b, Die99, Dou13, Dow08, Dre00, Dup13, Eel00, Enf08, Ero12, Far00, Far16b, Fel01, Fes06, Fie86, Fod01, Fre10, Fre11, FB01, Fri16, Gal00a, Gal00b, Gan11, Gar99, Gat01, Gia06a, Gia06b, Gib86, Gil00a, Gil01, Gil07, Gil09b, Gil13, Gla14, Gly04, Gly11a, Gra01, Gri08a, Gro09, Gro11, Gua07, Gut00, Hal15, Hal99, Hal01, Har11, Haw99, Hea86, Hea99, Hei99, Hei78, HM99].

Book [Hey16b, Hit00, Hoe00, Hof01, Hog86, Hoo86, Jar12, Jen86, Joh15, Ket00, Ket01, Kir01, Koe99, Kör86, Koz14, Kra09, Kre60a, Kus01, Kut00, Lad99, Lar00, LM99, Lev01, Lew14, Lil99, Lip99, Low00, Lyo06a, Mar00, May00, McC16, McK16, Mel99a, Mel00, Mor99, Mos86, Mum13, Nad99, Ney11, Ney14, O’H86, O’M14, Oft07, Oka00, Pag00, Pas16, Pau16, Per07, Phi09, Pin11a, Pin16, Pol14, Pre01, Pri99, Pri01, Pro86, PS00, Put01, Qui61, Rad00, Rai86, RW13, Red00, Reu09, Rey15, Ric16, Roc86, Ros99, Ros08, Ros09, Rub86, Rus00, Ryc99, SL13, Saf86, Sch07a, Sch08a, Sch59b, Sch11a, Sch00a, Sch14, Sco08, She06, She01, SF13, Sid01, Sie10a, Sim01, Sin00, Sky12, Sla15a, Sma00, Smi00a, Smi00b, Spr11, Ste12a, Ste10, Ste00, Sto99]. Book [Str09, Tal15, Tau06, Tei99, Tei13, Tha09, Til00, Tom99, Tso10, Ueb01, Vai15, WF13, Wal14, Wei14, Wer10, Wil14a, Woo00b, Yud16, Bai73, Moo83, Ryd81]. Books [Wit58b, Wit58c, dF81]. Boole [Pit94]. Boolean [BF23]. Boorse [Kin10]. Born [Kes61]. Boston [Fes06]. Botterill [Kir01]. Bottled [Cur21]. Boucher [Hea99]. Bovens [Ols02a]. Box [Sob98]. Boxes [Buc23]. Boyle [Jen88]. Bracken [D’A86]. Brad [RW13]. Bradford [Hoo86, Pau16]. Brain [Ara11, Bub94, CS12, Dew75, Mei22, Pir54, Puc73, Puc74, Puc76, Puc78, Puc81, Puc93, RKK19, Rob76, Sha12, Smy60a, Tha70, Tho94b, Wil80a, Wil22, Pri60, Smy62]. Brain-Damaged [Bub94]. Brain-On [CS12]. Brains [Bog77, Puc89, War22]. Braithwaite [Kyb58, Mel90]. Brakel [Har93]. Branches [DF22]. Branching [Bel05a, BM10, Bra11, Mül05b, SW08a, KP99, Pla00, BM10]. Brandl [She01]. Bread [Sob01]. Breadbox [Bur18]. Brett [Sch14]. Brian [Bai86, Ste10, WF13]. Bridge [Mor09]. Bridging [BM18, Fri04]. Bridgman [Sch59b]. Brief [Sac93]. Bright [Row23]. British [Cla00b, Den60a, HH63, HH64, May60, Sob01]. Broad [Wis62]. Broadbent [Joh15]. Broke [Lau89, LE22]. Brown [Wit58b, Wit58c, Din56, Fol99, Hof01]. Brownian [Cha11]. Bruno [Mil97a]. Bub [Sta85]. Buchdahl [Aga59b]. Bug [BCP93]. Building [Dav01]. Built [Gut00]. Burgess [Pas16]. Burniston [Din56]. Business [MG17]. Bust [DGO07, For95a, Sha17]. Butterfield [Dic01a]. Butts [Cap86]. Bynum [Tom99]. Byproduct [PC12].

C [And10, Fre10, Grü69a, Mar61a, Mel00, Pin11a, PS00]. C. [Str09]. Calcott [Sch14]. Calculi [Nor19]. Calculus [Goo61a, Goo61b, Pop55b]. Caldin [Mac60a]. Calibrating [LE22]. Calibration [Lad84, SW13, Whe12, Wil12]. Called [HO79, Leb60, Mac66, Mil66a]. Cambridge [And86, BD86a, Coh86a, Fie86, Gal00a, Gib86, Kör86, Rai86, Roc86, Tau06]. Campbell [SL13, McC86]. Can [Ale16, Ale18, Ash52, Bok08a, Bro93, BF12a, BF12b, Cas16, Cla66, FAB^{+}89, FW21, FA14, Goo74b, Grü69b, Grü76b, Grü86, Irz96, Max68, Ner79, PLKG24, Rob94, Sta09, Sti81, Tho65]. Canfield [Nis70]. Cantorian [Dau88, May77b, May77a]. Cao [Mar00]. Capabilities [Bog77]. Capacity [Pen21]. Capricious [Lai21]. Care [Kop05]. Carel [Gla14]. Cargile [Sch76]. Carla [Cam00]. Carnap [Hei78, IG95, Pop56b, Pot00, Pot02, Ryc99, Sär68, Smo77, Dem01]. Carroll [DF22]. Carruthers [Fod01, Hea99, Kir01]. Carsetti [Gia06a]. Carter [Gil07, Gre64, Wil94]. Cartesian [Hug04]. Cartographic [War22]. Cartwright [Fra88a, Gib84, KT93]. Carving [COS11, Lyo01, SL13]. Cascade [Fra95, Pat98]. Case [Bar19b, Bon92, DE76, Gil07, Har06, Hub11, Joh88, Kuy24, MP08, Mor86, Pot10, Pri84, Pro63, RH72, Ros56, Sch00c, Skl00, Smo21, Sta17, Van93a, Van94c, Wri68, vEDS06, Car03, Sti83, HK68]. Cases [Set77a]. Castellani [Low00]. Cat [Bub79, Hol88]. Catch [Sta09]. Catch-All [Sta09]. Categorial [LW15]. Categorical [Fel23, KZH^{+}22, LT17]. Categories [Mul01, Sla15a, Sma53, Van93a, Kha13a]. Category [Bel81]. Cauchy [Cle71, CKKR88, Fis79]. Causae [Bel05a]. Causal [Ahm14, And18, Arn90, Buc65, Car02b, Car06, Chi18, Eel00, Egg16, Ehr87, ES19, Eva18, FRSW18, Fri09b, Geb22, Gil09a, Gil07, Gly04, Goo61a, Goo61b, Gou58, Hau98, HW04, Hid05, KW97, Leu14, MW15, MW14, Mil06, Oka16, Ots16, Pap85, Pap88, Pin22, Ram13, Ros21, Shr19, Sko14, Smi07a, SS13, Sta17, Ste05, Ste06a, Ste12a, SO19, WS23, WC17, Zha13, ZG21, Car03, HW99, Joy99, Rus08, WC99, Woo03, Far16b, Hit00]. Causal-Structural [Leu14]. Causality [BY06, BM18, Buc65, Cas16, Din70, Dow93, Dow00, Ehr84, Fis67, Fri09a, Gil05, Gly09, Goo59b, Gri77a, Hoo93, Lyo67, Mac69d, Ner82, Nor09, Pri91a, Pur23, Rus08, Sch61c, Sch62, Ste12a, Gil01]. Causally [Set89]. Causantes [Bel05a]. Causation [Bal14, BF23, Bel05a, Cav10, Edg97, Far20, FGK15, Fis24, Fut05, Geb22, GE23, Gil14, Gly11b, HH15, Hub11, Kis06, McD95, MP93, Mül05b, Nat14, Ney09, Per18, Rig91, Sch01a, Sch07a, Sch98, Smi07a, Sob87, SGS97, Spo06, SP12, Wil51, Win21a, Win21b, Hau99, HK08, Nad99, Mum13, Ero12, Hei99]. Cause [FG17, Hoo03, Mor60, Pla60, RS07, Smi94, Sob01, Ste03a, Tor87, Wil50, Wor07, HSRS99, Sch07a, Woo00b]. Caused [Cra93]. Causes [Bel05a, Cha13, Ful19, Gly10a, HP05a, HP05b, Hoo93, Irz96, KS19, Mel80, Mor94, NS20, Pap91, Woo90, MA11]. Cautious [Gre81]. Cell [Cur16, Fag13]. Cellular [Mat17]. central [Šus07]. Centralized [Avi19]. centuries [PP78]. Century [Mar00, Wis63a, Gin84, Rai86, Cle79, Cro53a, Elk71, Hut53a, Pot53]. Certain [Fai56]. Certainties [Wil14a]. Certainty [Why63]. Ceteris [Dre01, Kle92, PR95, Reu09, Rup07, Sch07c, Sch01b]. Chair [Bas92]. Chairman [Pan60]. Chakravartty [Cho11]. Challenge [Ake86, Ake90a, Bub00, Rad91b, Sar15, VH03, Vic21, You61, Yud13, Hal99]. Challenges [Sti94]. Chalmers [Pol14, Cur83, Gow88]. Chamberlin [Fod01]. Chance [AH03, Bla98, Cus22, Edg97, Eme15, Gly10b, How78, Mau00, Mel65, Mel67a, Pri84, Rob13, Sch07a, Sch07b, Súr00, Mil01, Gly11a]. Chances [BCP93, Fis06, Hau13, Hum04, Ism96, Lan06, NV22, Sch03, Suá18, Her07]. Chang [Gil09b]. Change [BP89, Cal09, Cha79c, CK89, Dur02, HR95, KK65, Nee13, Noo97, Par71, Ten94, Ten97, UH19, Wer16]. Changing [Fey77, San93, Ten94, Ten97, PP78, Tom99]. Chaos [Kop01, Str09, Wag12, Wer09, Wer10, BV05, Str03, FB01]. Chaotic [Fri04]. Character [BD86b, Naj24]. Character-Factor [BD86b]. Charity [Wac87]. Charles [Ecc52, Mos85]. Charlie [Bir01]. Chart [Sch00b]. Chasing [Myr09]. Chemical [Hav18, Mun94, Pan62a, Pan62b, Pot53]. Chemistry [Bra55a, Cal59, Goo10, HN07, Kuy24, Le 05, Mac60a]. Chen [Tha09]. Cherniak [Bro90]. Chess [Ash52, Hal52a, Hic53, Puc74]. Chess-Player [Ash52, Hic53]. Chi [Bai83]. Chi-Squared [Bai83]. Chihara [How97b, Mel00]. Child [Spu52]. Childhood [Car02a]. Chimeras [Myr09]. Chinese [Spr07]. Choice [AH14, Bar13, Bir14, Bra17, Cof14, Col15, GSK13, GT16, KZ16, LC15, Mac71b, PR13, PC12, Sam75, Sar15, Sch79a, Sko14, Sme15, SW13, Ste16c, Wer16, Kar12, Pea12, Ros12]. Choices [Mac73a, Sch74c]. Chomskian [Sti79]. Chomsky [Bra84, Col07, D’A86].

Christian [Har11]. Christopher [Gib86, Put01, Bak15]. Chromatography [LE22]. Chronogeometry [MLR21]. Church [Bla89, But09, Cot03, Pic11]. Churchland [Dre00, Dre00, KH10]. Circle [Obe98, Ueb96]. Circularity [Bam93, Shu19, Sho00]. circumstances [Wal06a, Bun86]. Claims [Ach95, Sta17]. Clarendon [CE04, Mos86]. Clarifications [May08]. Clarity [Wal10]. Clark [But93]. Clarke [Wil75a]. Classes [Cha07b, Mul01]. Classical [AG09, BD86c, Bar15, Bar19d, Ber10, Boh62, Bok08a, Bra55a, Buc51, Cal95, Cur14, Dea61, Dob21, EN98, Fra95, Fri02b, Gil14, Hoe69, Hut93, Hut95b, Lar97, Lav77, Low00, May77b, MS55, Mun89, Myr11, Pat98, PB02, Pop50a, Pop50b, Sav94, Urb92, Vic08, Wil23b, Bok08b, Goo13]. Classification [Hon91, Pra72, Sla15a, Kha13a]. Classifying [Coo02, Coo05, Tso10]. Clatterbaugh [Nad99]. Clavelin [Ste74]. Clerk [Smi00a]. Clifford [Bun86]. Clifton [BG01]. Climate [Kat13, SW13, Wer16, Wer19b]. Cliometric [Mee04]. Clipped [Poo55]. Clock [Dor68, Pod79, Ros78]. Clocks [Kno10, MLR21]. Close [IG95, Poo55]. Close-Clipped [Poo55]. Closed [Set89]. Closer [Sor07, Str95]. Closet [Tho94a]. Closure [Nia05]. cloth [CE04, Fes06, Tau06]. Clothes [Fey81]. co [OB12]. co-operation [OB12]. Coalescence [JHB^{+}23]. Coarse [Rob20]. Coarse-Graining [Rob20]. Cockburn [Gal00b]. Codell [Gil07]. Coding [AB16, Ste16b]. Cogency [Odd97]. Cogito [Sor86]. Cognition [Bub94, Chr10, Ech16, Gia06a, HPE92, Kha02, Nad08, Pet22, SW19b, Wei04, Car04b, Ste03b, Fod01]. Cognitive [Bar19a, Ber16, BG16b, Car02a, CH17, Dow08, Far16a, Fra95, GS08a, Gly94, Hoo86, Irv16, Kuk92, Lyn90, Odd97, PS00, Res09, RKK19, Sha84, SD93, Tal13, Tha09, Wal07, ABC06, LM99, Sti83]. Cognitivist [Sca10]. Cohen [Lev79]. Coherence [CHvI15, Cur17, Dan05, ET97, Gla21, How08, KS19, NW22, Sch11b]. Coherentism [Her23]. Cohorts [Akk94]. Cointegration [Hoo03]. Coksobosation [Pop60]. Collapse [Bec04, CT16, Kri88, CM99]. Colleagues [Wis65]. College [PP78, Per07]. Collingwood [Wil64a]. Collins [Ste00]. Collision [Ang05, Lar04]. Colour [Dav18, Van93a, Van94c]. Colouring [AH14]. Colours [Dav18]. Combinatorial [Le 05]. Combining [Mor13, Yud16]. come [Wal06a]. Command [Kle17]. Command-Following [Kle17]. Comment [Bar08b, Bon55a, Bot94, De 72, Den94, Deu96, Goo85, Han92, Loe96, Pea83, Pop62, Pop66a, Raa05, Sau96, Squ89, Ste06a, Von87, Mee99]. Commentators [Mac78c]. Comments [Aga59b, BH55, BH56, Car02b, Car06, Coh80b, Dob61, Dor76a, EN98, Fai57, Fou57, Gra55, Kul77, Lon53, Mac60a, Mil78, New73, Pan60, Sam60, Sta90, Sut57, Sza56, Tiz76, Was55, Wei89]. Commitments [Ber23]. Committee [KK65]. Common [AGTZ08, Bro20, Che91, Hoo03, KS19, Mac72, Max66, Rus53, Sob01, SO03, SS17, Ste03a, Tor87, Yud16, HSRS99, Mor13]. Commonsense [MK56]. Commonwealth [May60]. Communicable [Mar59a, Mar59b]. Communication [BR20, Fri04, Lig14, McW16, Pet23]. Communicational [Mar61b, Mar61c, Mar61d]. communist [Jos10]. Commutativity [Gal14, Wei09]. Companion [Gal00a]. Company [Bro86]. Comparability [Gie73a, Lep69]. Comparative [BH53, Car53, FW07, Liu22, Rus51, Sch74a, Vol95, Buc23]. Comparativism [Mar22]. Comparing [BH94, WHB^{+}23a]. Comparison [Arc67, BD86c, BD86b, Ell57, Mil74b, Zha13]. Compatible [Pra24]. Competence [Cla90, Fra95, Pat98, Fra99]. Competitors [And18]. Compleat [Kri61]. Complementarity [Aga58a, Bun55a, Bun55b, Hol86, Ves20]. Complements [And18]. Completability [MW15]. Complete [Mag04]. Completeness [CK20, Hoo91]. completers [Rup07]. Completion [CL19, MW15]. Complex [Ada01, Col00, GBW61b, Nea95a]. Complex-System [Col00]. complexes [Hos00]. Complexity [Goo74a, LS94, RS23, SF13, Wer10, BV05, Mit09, Str03, WC99, Str09]. Complexity-Based [RS23]. Composing [LaC22]. Composition [Mas17, Tor15]. Comprehensibility [Sma00]. Comprehensive [Rat02, VC24]. Computation [AB10, Goo71a, Mal23, Mil19, Wel08, Wel93]. Computational [Chi18, Lee21, Mol22a, Res13, Siz00, Ten94]. Computationalist [Lyn90]. Computer [Moo78, Par17]. Computers [Bun56a, Bun56b, But09, Hog04, Puc74, Tom99]. Computing [PNSBC94, Pre01]. Concentration [Nat14]. Concept [AA87, Bar12, Car53, Cle79, Coh86b, Coo65, Din65a, Gre51, Hin56, Jar67a, Kes65a, Kes65b, Kes66a, Kes66b, Kes66c, Koz14, Lev65, MS06a, May52, Pan62a, Pan62b, Pap52, Pir52a, San93, Til93b, Irv13]. Conception [Cro53a, Le 52, Pov18, Ram13, Sma00]. Conceptions [Fai56, Mel99b]. Concepts [Bar82, BM18, Boh62, Bri51b, Bri51a, Bri51c, Car52, CH01, DiS99, Fis24, Gil07, Pir52c, RO55, Ros21, Smi59, VM16, Car03, LM99]. Conceptual [Ber68, Gie73a, UH19, Yud16, Lil99, Mor13, Mar00]. Conceptualisation [SS65]. Conceptually [Mor86]. Concerning [Dew73, Goo74a, Goo85, Kul77, Lar05, Ram23, Sni67, Wei89, Sch00b]. Concrete [Pin22]. Condition [Car02b, Car06, HW04, Lar03, Pla60, Sho03, Ste05, Ste06a, Wer19b, Wil79b, HW99].

Conditional

[Hum04, Hun70, Mil97a, Hut99]. Conditionalisation [Wil80b]. Conditionalization [HF94, NR13, Odd97, Rai20]. Conditionalizers [Gal14, Wei09]. Conditionalizing [Wil98]. Conditioning [MG94, TH20]. Conditions [Bel05a, Bra93, Cal04, Gri88, Ner71, Ver05, Pit94]. Conference [A.56b, Den59a, Den60a, HH63, HH64]. Confidence [Joh88]. Configuration [Lew04, Sce91]. Confirmable [Puc62]. Confirmation [Ale58a, Ale59a, Bai69, Bai71, Bai73, BH55, BH56, Bau64, Bau68, Bra11, Bra15, BH15, CV83, Car53, Car56, Chi87, CFM08, DTW17, ES89, FW05, FW07, For95a, Fre88, Gib69, Gil91b, Goo60a, Goo61c, Goo85, Hub08, Kah69, KLS92, Law85, Leb60, Mac63b, Mac69c, McC86, McG03, Mus74, Nel96, Nun93, Ott94, Pet22, Pop54, Pop55a, Pop56b, Pop57b, Pop58a, Red86, Riv94, Ros94, SS20, Sch70, Sch64b, Sea87, Set70, Sho17, Smy56, SW13, Sti85, Sto66, Wat60a, Whe09, Wil64b, Wil14b, Kah67]. Confirmatory [Ack66]. Conflict [Ste16c]. Conflicting [Wag15]. Conflicts [Gar73a]. Confounding [Ful19]. Confused [RD17]. Confusion [Ses07, Sni67]. Confusions [AL04, Puc81]. Conjectures [Grü76c]. Conjecturing [Hac71b]. Conjunction [SS20]. Conjunctive [Tor87]. Connection [Bea89]. Connectionism [But93, Cla90, Dav89, HPE92, Row94, San08, Sti94, Hoe00, Gar99]. Connectionist [Bot94]. Connections [Arn94, Car02a, Max68]. Connectivity [Mel65, Mel67a]. Conscious [Cla66, Tho65]. Consciousness [Cla07, Dew75, Gun68, Irv12, Irv13, Kle17, Lev97a, Lev01, Mac63a, Mar05, Roe04, SB10, Smy60a, Wal07, Wil84, Koz14]. Consensus [BH92, Mee04]. Consequence [Coo03b, Hor78b, Von87]. Consequences [Geb22, GW14, Ryd81]. Conservation [Cla76, Sch00c]. Conservatism [KZ16]. Conservativeness [Coo12]. Conserving [MP22]. Consideration [Cod75]. Considerations [Gil92b, Hum04]. Consilience [Coh68]. Consistency [How08, Jon82a, Leh70, May77b, May77a, Pop56a]. Consistent [Cas98]. Consolations [Gie71]. Conspiracy [Lew06, Sam18]. Constancy [Ben12, BM93]. Constantine [Dic01a]. Constitution [CB23]. Constitutive [BG16a, KK17, Leu12, SH20]. Constraints [Bra17, Mar89a, Win21a, Win21b]. Constructing [Cha12, Pol14]. Construction [Ale58b, Fei22, HR22, Ryc99, UH19, Wit61a]. Constructive [CC93, Dic07, Han04, Lad00, Lad04, Lev97b, MvF03, Nor08, Sob85, Bil00]. Constructivism [Hel98, Koe00, Bil00]. Constructivist [Ste20]. Contact [Hut51d, Smi07b]. Containing [SS93]. Contemporary [Bet51c, Bet51d, CK89, Gie75, Kle02, Mum13, Sch61c, Sch62, Yud16, Koe99, Mor13]. Content [All92, Bam93, BH56, Boh61, Car56, HF85, Lee21, MS06b, MHB^{+}23, Mun86a, Pop56b, SGSC18, Why93, Zah82a, Zah82b, WC99, Pop55a]. Contentious [Elb94]. Contents [BH94, Elb94]. Context [Car52, Law85, Pir52c]. Contexts [Mor94, PNSBC94]. Contextual [Oka04, Shi84]. Contingency [Kum69, WHB^{+}23b]. Continued [VHH86]. Continuity [Abr58, Cle71, Per18]. Continuous [Smi07b, Sta85]. Continuum [Joz86, Smi07b]. Contra [Han04]. Contract [Ern01]. Contraction [Din59, Dor68, Grü59, Grü60, Pop59a, Pop59b, Pop66b, BF00]. Contrary [Dre00]. Contrast [Cha07b]. Contrastive [Hit99, Kuo24, Sta17, SP12]. Contribution [Tro93, Wis62]. Contributions [Pan60]. Control [Cla07, HPE92]. Controversial [Nia05].

Controversy [Bai83, Gal81, Gie72b, Goo13, Gri95, Len06, Llo67, Mor06, Oka01, Puc78, Zah83a, BER99, Koe99]. Convenience [Gie73c]. Convention [BM01]. Conventional [Chr77, Wag15]. Conventionalism [Mor88, Mor89, Sto86, Swi80a, Zar79]. Conventionality [BHS15, Pet89, Pur23, Rox77]. Convergence [Cle71, Cur13, DE76, Har13, Mac94b, Nie21, Pea12]. Convergence-of-Opinion [DE76]. Convergent [Aro89]. convex [Whe12]. Convexity [Wil12]. Cooper [Gla14, Tso10]. Cooperation [Bir13, Ste16c, Sch14]. Copenhagen [Bub68, Loe98]. Copenhagenism [Fre78]. Copernican [Cha81, Cha83, Cur83, Hal70, Hut83, PS86, Tho92]. Cornell [HM99]. Corpuscular [Sab54]. Correction [Bel17a, Din56, Goo74a, Puc77]. Correlated [Gly08, Oka05]. Correlation [Sol75, Whe09]. Correlations [Elb93, Irz96, Pap91]. Correspondence [Mił23, Rad91a, Sch09, Lar00]. Corrigenda [Goo62]. Corrigendum [Coh67a, Coh67b, TP21, Win21a]. Corroborating [O’N82, Ols02a, Ols02b]. Corroboration [Aga59a, Coh82, Goo68a, Lev63, Mic66, Pop58a, Row13, Spr18]. Cosmological [Cra01, Lin23, Pit08]. Cosmology [Aga58b, Dav62, Grü04, Har62, Ker87b, Kle02, KOV06, McC19, Mun54b, Mun62, Pro86, Wal10, WB53, WB54, Grü00]. Cosmotopology [Mag05]. Costs [Rub22b]. Could [DTW17, Gau66, Ism96, Jon69a, Odd19, Sud73, Tho92]. Count [Hoc17]. Countable [Bar04, How08, Wil99]. Counter [Jos82, Ste03a]. Counter-Example [Ste03a]. Counteracting [Gri88]. Counterarguments [She02]. Counterfactual [BS17b]. Counterfactuals [FGK15, JL22, KDM21]. Countering [San08]. Counterintuitive [TH20]. Counterpart [Bel95]. Counting [BG01, CM00, Lew97, Lew03, May08, SW18]. Counts [May08]. Coupling [DiF23]. Covariance [Oka04]. covers [Roc86]. CPT [Gre10]. Crane [Bir01]. Creaking [Cod75]. Creation [Bro86, Cra86, Cra90, Cra92, Mun54b, Hen84]. Creativity [Car02a]. creator [Cra90]. Credence [Bla98, Rob13]. Credences [Dod13, Ede23]. Credibility [Sea87]. Crick [Sch69a, Šus07]. Crisis [Bir22, Lav23]. Cristina [Gua07]. Criteria [Hut95a, Kie01, Lei13, Par71, Sch70]. Criterion [Hor78b, Kie97, Mac69c, Man76, Mil19, Tho92, vFH17]. Critical [Bet50, But02, Chr10, Dre00, Fai56, Gil93, Gil13, Hon91, Ste01, Wat65a, Gel59, Row11]. Criticism [Bal57, Bun86, Bus52b, DB52, Har59b, Mac78c, Pos62a, Ree78, Wis53a, Aga78]. Criticisms [Kap60, Wat91, Wis58a]. Critics [BM01, Lew03, Mac71b, Ten97]. Critique [Bel01, DF22, DH86, Ern75, KW97, Mar89a, Peć89, PC12, Rig91, She73, Ste00, Van88, Van87, Grü00]. Cronin [Gri95]. Cross [Hun81, SMM11]. Cross-Generation [SMM11]. Cross-Theoretic [Hun81]. Crucial [Bae19, Web09]. Crupi [Hub08]. Crux [Web09]. Crystallographers [MG17]. Cult [Rus53]. Cultural [Cha21, Cro08, Dri11, RD17, Hey16b]. Culturally [Tri94]. Culture [Ste16c]. Cumulative [Dri11]. Curie [Cha70a]. Current [New73, Puc81, Vic11, Why52b, WY72]. Curve [Bel17b, DeV97, For95b, Kie97, Kuk95, Tur90, Bel17a]. Curve-Fitting [Bel17b, DeV97, For95b, Kuk95, Kie97, Bel17a]. Cushing [DiS99]. Cybernetic [Tiz57]. Cybernetics [May51, Pol52a, TS53, Wal51, Wis51]. Cyclotron [BF90].

D [Bro86, Cap86, Gal00b, Hit00, May61, Pri99, Sch00a, SF13, Blu89]. D. [Gly11a]. d’Être [Bla66]. Daffa [Hog86]. Damaged [Bub94]. Damned [Wor89, Car99a]. dams [Car99a]. Dangerous [Goo74b]. Daniel [Bar01, CS16]. Danks [ZG21]. Darrell [Gil13]. Darwin [Dup13, Sco08, Dow05, Hel18, Kle85, Sob11, Sta06a, Ste01, Tuc11, Rad00]. Darwinian [Mol22b, Van88, Sin00]. Darwinism [Gay98, Old82, RS15, Amu99]. Daston [Jar12]. Data [Bra60, Lus21, Par17, Puc76, Smy58b, Smy62, Tre62, War22, Wri18]. David [Cal15, Bir01, Gal59, Hor78a, Mor88, Pol14, Sch00b, Urb94, Vai15]. Davies [Cra86]. Davy [Zuc88]. Dawid [Ric16, Tei13]. Dawkins [Elt03]. De-Coupling [DiF23]. Dead [Set70]. Death [Gra01, Mei22, Ses07]. Debate [And98, Bir01, Far00, Goo13, Hoe98, Kov19, Psi00, RS23, Nad99, Dea76, Tiz76, Urb74a, Urb74b]. Debating [Koe04]. Debt [Zah78]. debunking [Gła23]. Deception [McW16]. Deciding [Hog04]. Decision [Ahm05, Ahm14, Bra93, Cav10, Eel00, Sim65, WS23, Joy99, Mor99]. Decision-making [Bra93]. Decisions [Bir13, Wat55c, Bin09, OB12, Coh86a, Sky12]. Decoding [RKK19]. Decompose [Ric19]. Decompositions [Cla16]. Deconstructing [Obe99]. Decoupled [Chr10]. Dedicated [AGTZ08]. ’deduction [Wor00, Dor71, Sau64, Sle83, Sor86, Zah87]. Deductive [Kör64, Fra87]. Deep [Mul05a]. Default [Dac23]. Defaults [HH15]. Defeated [But06a]. Defence [Fie86, Fis67, Hor65, Kuy24, Old78, VM16, Dup00, Lil99]. Defend [Cha89]. Defended [Sam78, Sle88]. Defending [Har13, Kos21, Mey21]. Defense [Hob94, Mac14, Mor12, Lad99]. Defensible [Cla01]. Defining [MP18, Wer16]. Definition [FG17, Hes52, Pop68, Suc69, Sza56, Wal56a, Wil68a, Wil68b, Win67]. Definitions [Har74, Rus69, Sto66, Tic74].

Defusing [GR15]. Degeneration [Tiz76, Urb74a, Urb74b]. Degree [BH55, BH56, Car56, Pop54, Pop55a, Pop56b, Pop57b, Pop58a]. Degrees [HH21, Hay55, Leb60, Nii79, Wil10, Wil14a]. Deity [Stu61]. della [Gil03]. Delusional [BG16b, Par16, Par21]. Delusions [Bor16, BGH^{+}23]. Demarcation [May96]. Demise [Mei22]. Democratic [Dou13, Dur02, Kit11]. demons [Hit99]. ’demonstration [Wor00, Nor19]. Demonstrations [Abo02]. Demopoulos [Hal15]. Denial [IS24]. Dennett [Rat02, Tod09]. Dennis [Cla66]. density [CM00]. Deoccamization [Dur89]. Depaul [Bon01]. Depend [Wol12]. Dependence [Hun23, Wer19b]. Dependency [Cur17, Del21]. Dependent [Gal14, Sho00]. Derivation [BM93, GPW05]. Derivative [HS17]. Derivatives [Eas14, Swa22]. Descartes [D’A86, Hoe00, Adk52, Bra84, Hug04, Pir50, Sle83, Sle88, Sor86, Wig56]. Description [Fin81]. Descriptions [BM90]. Descriptive [Cro53a, Pet23]. Design [CB69, Cra88, Goo09, Kle02, Mil01, Sch08c, Wei05]. Designer [Ash52]. Desirability [BS17b]. Desire [NR13]. Desires [Sch71]. Details [Bra21, CK20]. Detection [Mor24]. Determinacy [Bro09a]. Determinate [Tum07]. Determination [McK20, PS86]. Determining [Nee17]. Determinism [Bel95, Ben63b, Ben63c, Bra62a, Bri97, Coa74, Dea61, Dob21, Grü56, Lat72, Mac74a, MP18, O’C57, Pla63, Sch87, Wil09b, Zim08, Hit99, Mel99b]. Deterministic [GL23, Gly10b, Hut93, Nie21, Sch07b, Sud76, Tav91, Wag12]. Deutsch [Pri99]. Development [Bet50, Cal59, Coh87, DiF23, Far16a, Gil78, Pla16, She13, Szu77, HH92]. Developmental [Kov19]. Developments [Mar00, Wis63a]. Deviant [Gar72b]. Devices [Lai88]. Devo [NV22]. Dewitt [Puc76]. Dhahran [Hog86]. Diachronic [Tor15]. Diagnoses [Coh80a, Fel23].

Diagnostic [Coo05, Tso10, Coo02].

Diagonal

[Kre53a, OK05, Sle83, Sle88, Sor86].

Diagrams [Nor04]. dialogue

[Hen84, Bro86]. Diathesis [Dob57, Kap57b]. Dichotomy [Hin88, Kav23, Sha05, Ang01]. Dickson [Mau00, Súr00]. Dictionaries [Hes58a, Hes58c]. Dictum [Adk52]. Did [Bec04, Dor78, Pro74, Smi94, Sob11, Zah73a, Zah73b, Dup13]. Dieks [Cli95]. Differ [Ros21]. Difference [BF23, CPG20, Mil57, Ney09, Pop66b, Tod09]. Difference-Making [BF23, Ney09]. Different [Wil73a]. Differential [Ell63, ES89, Sch59b]. Differing [Hut95a]. Diffraction [SV11]. Diffusion [BS00]. Digital [LRP23, Tom99]. Dilemma [Max88, Sam78, Squ89, Wei09, You86, For95b, Mar90, D’A78]. Dilemmas [Dur02]. Dimension [Gil09a]. Dimensional [Cha57, Kar12, Mun53, Mun54a, Til94, Bun58a]. Dimensionality [Abr58]. Dimensionless [Why54a]. Dimensions [Bri06, Bro55, Gro09, Kos88, Whi55e, JLZ05, Le 03]. Dingle [Grü60, Grü61, Kes66b, Sch65a, Van61a, Whi81]. Dirac [DEO19]. Directedness [Smo21]. Direction [CBR90, Nar65, Pla60, Sav96]. Dis [Fri00]. Dis- [Fri00]. Disagreement [NS21]. Disappearance [Kap55c]. Disarming [Vic21]. Disassembled [Nor07]. Discernibility [HN14]. Discerning [MS08]. Discontents [Bir14]. Discord [PG13]. Discount [May08]. Discoveries [Bec19, Hud01]. Discovering [GG86, Han61, Han62a, Hut62, Shr19]. Discovery [Bai87, God69, KW97, Lai91, Max93b, Sch08b, Wil03, Wil68a, Wil68b, Yux90, Zah80, Zah83b]. Discrete [Bot94]. Discs [But06a]. Discussion [BG99, BM10, Bil00, Bla78, Bla63a, BV63, Bog77, Bon55b, Bra63a, Bra63b, Car64, Cha00, Cla66, CM99, CM00, Dup00, Erw82, Fol99, Fra99, Gal59, Lar99, Mac60a, Mat99, Mee99, Mul91, Obe99, Peć89, Pot53, Sch00b, Swi00, Tha70, Ueb99, Urb77, Wat99, WB53]. Discussions [Grü59, HK68, Har59b, Har60, Mar59b, Pop59a, Red78, Ros78, Sam78]. Disease [Gla14, Gre79, GM18, Kin10, Kin16, Le 52, CC13, Car03, Gil07]. Disjunctive [Bal14, Coh02]. Disorder [Bur54, Den89b, Sza56, Wal56a, Wal56c]. Disorders [Coo05, Coo02, Tso10]. Disorganization [Den89b]. Disparity [Gie73a]. Dispersion [Fri09a]. Dispositional [Cox75, Kin16, Wer19a]. Dispositions [Bir01, Coh02, Dre01, Tho88, Bir01]. Disproofs [Wil76]. Disputations [PR91]. Disputed [Lev97a]. Dissolving [Mik04]. Distance [Pir53, Tay23]. Distances [Cha81, Cha83, Hut83, PS86]. Distant [Mun89]. Distinction [BZ84, Kuy24]. Distinctively [Lan13]. distinctiveness [Wor00]. Distinguish [Kin10]. Distributed [Lyn90]. Distributions [Lyo14]. Distributive [IK24, Sob95]. Divergent [Hoo91]. Divide [Lyo06b]. Divine [Bat07, Bee09, Cra88, Fos04]. Dna [Ste09, Sch08b]. Do [Bun56a, Bun56b, CT23, Ewe11, GB99, Gin93, Lan65, Lan06, Mcd72, Net08, Tah21, Tor87, Wol12, Woo53b, Dri11, Gil92b, Goo10, Ros12]. Dobbs [Fer70, Grü69a, Kap57b, Mar61a, Mun53, Mun54a]. Doctrine [Mar69, Sch87]. Doctrines [Hoo71]. Dodges [Aga78]. Does [BL23, Cab20, Col08, ER96, Ern90, Haa91, LP18, Mor09, Pet20, Pit08, TH20, Wal07, GB99, Hoc17]. Dog [Cow98, Spu52]. dogma [Šus07]. Doing [Bol83, Gia06b, Lom02, Mat05]. Domains [Van94a]. Dominic [Bro09b]. Dominion [Sin00]. Don [BL14, Dow08]. Don’t [AS19, Hoo93, Lau89, Mac05, Mcd72, Ric19]. Doomsday [Wil17]. Doppler [Din60a, Din60b]. Dorato [Sch11a]. Dordrecht [Fes06, Gia06a]. Dordrecht/Boston/London [Fes06].

Double [Cap86, May08, Puc93, SW18, Tho94b, But84b]. Double-Counting [May08, SW18]. Doubt [Sle88]. Doubts [Sol75]. Dowe [Sch07a, Sco08]. Down [Set70]. Dr [Aga59b, Ale59a, Das58, Din56, Din61a, Din65b, Dob61, Han62b, Mar61a, Pop55a, Pop56a, Wis58a]. Dr. [Gre64]. Drawing [Cha11]. Dreyfus [Wil76]. Drift [Gil09a]. Drive [Hin56, Rus50]. Drop [Nia05]. Drops [Nyr21]. Dual [Mag65]. Dualism [Fre53, Lan59, You61]. Duality [Puc89, Red77]. Dualling [DH86]. Ducks [May96]. Duhem [Aga57, Ari84, FW07, Web09]. Dumb [DTW17]. Dummett [Gil03, She01]. Dupré [Cla14]. during [PP78]. Dust [Bel05b]. Dutch [Hei78, Bai73, dF81, Moo83, Ryd81]. Dynamic [Ale14, Far16a, Hic18, NW22]. Dynamical [Fri04, HH93, KT93, Lar05, Mey21, Smi98, Yos12, BG99, GB99]. Dynamics [Dis94, EN98, Gär84, Gua07, Lar97, Lar04, RS96, Whi50, BV05, Bic06, Koe99, Skl13, BS94, Bat15, Ada01]. Dysfunction [GM18].

E. [Grü53a, Grü54, LM85, Whi54b]. Each [Bar64, Dia63, Hul74]. Early [Plu05]. Earman [Lar99]. Ecology [Cur21]. Economic [Arc59, Bac95, Bra99, Cod75, Col15, Hut53a, KLM10, Mar57, Rad87, Dow08]. Economics [Bic87, Bol83, Coa74, Col00, GJ88, Han92, Kla64, Lat72, Lin93, Mac74a, Mar64, Mee64, Pik54, Pik55, Sto64, Tiz57, Gua05, Sch05, Gro11, Ros08]. Economists [Hoo93]. Economy [Ros09, Mir04]. Ed [Rai86, Gal00a, Gib86, Gut00, Low00, Pro86, Tei99]. edited [Bir01]. Edition [Bir12, Lyo06a]. Editor [AH14, Bar13, Bir14, Cof14, Col15, GSK13, GT16, Kar12, KZ16, LC15, Pea12, PR13, PC12, Ros12, Sar15, Sko14, Sme15, SW13, Ste16c, Wer16]. Editorial [Ano21a, BL05, BL06, Cla00a, FM11, Mas16, Pap94a, Pap96a, Pap98]. Editorship [KK65]. Edmund [Bor54, W56]. Eds [Sch14, Bir13, Bon01, BL14, Cam00, Ero12, Hea99, Kut00, SL13, She01, Tei13, Tom99]. Edward [Gre84]. Edwin [And10]. Effect [Hul65a, Hul65b, Lan01, Liu94, Mor60, Wat99, Wil50, Dup00, Din60a, Din60b]. Effective [Fra21, Wil19]. Effectiveness [Fra21, Mol23]. Effects [Bal14, Nad08, SW19b, Wei19]. Effortless [Ros09, Mir04]. Effusions [Bat07]. Eggs [Dun11]. Ehring [Pap88]. Eighteenth [Elk71, Hut53a]. Eighth [HH64]. Eileen [HM99]. Einstein [Bha86, Bra22, But07, Cou52a, Cul81, Dor71, Dor78, Fey74, Fey80, Gar72a, Hoo91, KK83, Lau95, Mac86, Max93a, Max93c, Max93b, McC84, McC57, Pop66b, Pro74, Sch74a, Wyk76, Zah73a, Zah73b, Zah77, Zah78, Zah80, Zah87]. Einsteinian [Dor68]. Elaborated [Bor16]. Elder [Pap05]. elections [BD99]. Electoral [Sor84]. Electric [Why52b]. Electricity [Bor54]. Electrodynamics [Fri02b, Vic08]. Electromagnetic [Mun89, Pit11]. Electromagnetism [AG09, Bel98]. Electron [Why52b, Why53b]. Electronic [Sce91]. Electrostatics [San61]. Element [Pan62a, Pan62b, Rus50]. Elementary [Gia94]. Elements [Fow83, Le 05, Gil00a, Mue69, Sch00a]. Elga [Her07]. Eliminability [Dor77, Ryn83, SG73]. eliminating [Mil01]. Eliminative [Bar19b, Kah69]. Eliminativism [Col07, Loc03]. Elizabeth [Koz14]. Elkana [Cla76, Kul77]. Elliott [Dup13]. Ellis [Vol91]. Embody [Goo10]. Emergence [Hoo04, Pap52, Ryc99]. Emergent [Vai15, Wal20, Wal12]. Emily [Fre11]. Emotion [Cha02a, Eva02, Han05, Pri04, Gri08a]. Emotions [Loa24, Sca10]. Emperor [Fey81, Har93]. Empirical [Ach95, Ake88, Bac95, Bam93, Bat10, Bra99, Dor68, Fri15a, Fri72, Gil92b, GW14, Haa91, Ket04, Lai88, LRP23, Mar64, Mil18, MHB^{+}23, Miz23, Oka97, Pap76, Pin11c, Pit11, Pop60, Ram23, Sch09, Smy56, Sta20, Sto64, Wat60b, Zah82a, Zah82b, Kos00]. Empiricism [CC93, Den86, Dic07, Han04, Lad00, Lad04, Lak76, Max93c, Max93b, MvF03, Pla62, Ric77, Ruy19, Sob85, Sti79, Kan84, Ryc99, Sky84, Bai86]. Empiricist [Hul77]. Empty [Lew07]. Encyclopedia [Hes69a]. End [BH15, Gro08, How75b]. ends [Sch99]. Enemies [IG95]. Energy [Cla76, Jam63, Leh11, Rea20]. Engagement [Cur23]. Engineering [Irz85]. English [Pri55]. Enhanced [Saa11]. Enigmas [Bri06, Le 03]. Enough [Ber23, Smi97]. Enquiry [Fin81, DTV11, Tei13]. Enskillment [Far16a]. Entanglement [Bub00, FGK15]. Enthusiastic [Bet24]. Entities [Pau71]. Entity [Cla01]. Entropy [BS92, Bur54, Den89b, Den94, Fri04, Hen03, Kle53, Pop57a, Vas21b, Wal10, She99]. enumeration [BG99]. Environment [Mar89b]. Environmental [CS16]. Eötvös [Bon92]. Epicyclic [Har72]. Epidemiology [Bro13, Joh15]. Epigenetic [Gro09, JLZ05]. Epiphenomenalism [Lon53, Peć89, Wal07]. Episodes [Gro85a]. Epistemic [Avi19, BGTW24, Bar98, Bor16, Cur23, Das16, DF22, Dob21, Hun23, IK19, IK24, McC19, Nic87, Par16, Pot10, Tul12, Vol91, WOB21, Wil13a, Wil78b, Wil22, Wri86]. Epistemics [Cod75]. Epistemological [Jab92, Pan62a, Pan62b, Sco61]. Epistemologies [Giu88]. Epistemologists [Mor00]. Epistemology [Aga59b, Bea89, Coh80b, CT16, Dur89, Ell88, Fra84b, Gal03, Gia94, Gou23, Kit87, Mar62, Mar61b, Mar61c, Mar61d, Mor00, O’D90, Pri09, Pry01, SS17, Tal16, Wal06a, Wit59, Sch99, Wra11, RW13]. Epochs [Bet50]. EPR [Elb93, EPW13]. Epsilon [Rai20]. Epsilon-Conditionalization [Rai20]. Equal [Lan06, PR95]. Equally [Mor88]. Equation [Tur91]. Equations [Bor61, Kes60]. Equilibrium [Alb94, ER96].

Equipossibility [Hac71a]. Equivalence [Cof14, Coh02, Dor68, Fri15a, Ker87b, Kno14, Lup18, New73, Oka97, Pit11, Sac76, Tul12, WY73, Bra22, WY72]. Equivalent [Bar19d, BM90, Sor07]. Ergodic [ER96, Rei23]. Eric [Har11]. Erkenntnistheorie [Zah83a]. Ernest [Kör86, Qui61]. Ernst [Bla85, Bla89]. Erotetic [Kle70]. Errata [Bet51c, Bet51a, Bet51b, Goo62]. Erratum [Ano99, Ben63a, Ben63b]. Error [OL23, Sai95]. Erwin [Gre84]. Escaping [Kav23]. ESP [Cha64, Mar60]. Essay [Bir12, BF00, May77b, May77a, Phi09, Rey15, Ueb01, Le 07, Sla13]. Essays [Bun86, Cal15, Cam00, Cla14, Coh86a, D’A86, Dre00, Dup13, Ero12, HHS02, Pro86, She01, Bra84, CC13, Dup14, Fie86, HK08, NA13, Sob11, Gla14, Kus01]. Essential [Pin22]. Essentialism [Fal79, Hul65a, Hul65b, Max93c, Wal06b]. Essentially [Rei23]. Establish [Chr77]. Establishing [Wil23a]. Estimated [Mic66]. Ether [Pro68]. Ethics [Rus00, Lil99]. Ethological [Hin56]. Ethology [PS00]. Etiological [VH03]. Euclid [Fow83, Mue69]. Euclidean [Mcg70]. Euclides [Luc69a]. Eva [Gro09]. Evaluate [Tho92]. Evaluating [vEDS06]. Evaluation [Boe22, Coh80a, Fai56, Sch74a]. Evaluations [Sor84]. Evans [Tri94]. Evans-Pritchard [Tri94]. Even [Hoo93, Tho92]. Evens [Ang05]. Event [Fis67]. Events [Grü69a, Hei78, Max68, Mil97a, Sko14, Wei72]. Ever [Kne68b]. Everett [Vai15, Bra11, Tap00, Tap08, Tap11, Wal06a, Wal12]. Everettian [Bra15, DF22, DB15, Har22, SC18, Wil13b, Wil14b]. Everything [Bar05b, Mag05]. Evidence [Ach95, BGTW24, Bar08b, Bet16, Bin90, CS16, CFM08, Cur13, Dac23, DTV11, Gal14, Gil90, Gim04, Goo67b, GBW61b, How91, Hub08, Kle10, Kov19, Kuo24, MC12, Mah90, Mon06, Net08, OL23, Sch08b, Set69, SS17, Tap11, Til92, Tri94, Whe12, Bar99, Tei13]. Evidence-Based [MC12]. Evidential [Ahm05, Coh66a, Coh66b, Coh67b, Ede23, TP19, TP21, Wil69, Wri18]. Evo [NV22]. Evolution [Bal92, Bar23, Car02a, Chr10, Coh74, DiF23, Dri11, GC19, GM18, IS24, JLZ05, Lew14, Mat99, MP08, OB12, Pea12, Pow12, Pum53, Pum54, Ric77, Sme15, SF17, Van87, Wag12, WF13, Was78, Hoo84, Sky10, Ste03b, Ste12b, Sto99, Bir13, Gib86, Gro09, Hey16b, Sch14, Fod01]. Evolutionarily [Arm93, Mar90]. Evolutionary [BV63, Car64, Coh73, Coh87, Col00, GSK13, Gre58a, Gre58b, Gre63b, Gre64, Hau94, LB18, MC12, Mic81, Old78, Ots16, PC12, Pur23, Rad00, RW13, Ric07, Sam98, Stu63, UH19, Van63, Wal06b, WHB^{+}23b, Wra11, Fre10]. Evolvability [Bro14]. Evolve [LaC22]. Evolved [Lew14, Ste12b]. Evolving [Goo68a, O’C17]. Exact [Hog86, Tav91, ADS84]. Examination [Coo05, PNSBC94, Tso10, Coo02]. Example [Ake91, Jos82, Ste03a]. Examples [AGTZ14]. Exceeding [Sta06b, Enf08]. Excess [HF85]. Excitations [Lar03]. Exclusion [Kra09, Mas05, Mas01]. Exist [Sud73]. Existence [Cur18, Hol98, Lej54, Pet89, Fos04]. Exit [Lat76]. Expanding [Egg16]. Expectation [Aut20]. Expected [Wei86]. Experience [Ste00, Pit94, Cam00]. Experiment [Abo02, Bon92, Din64a, EPW13, Fra81, Fra84b, Gen98, Hor94, Mil85a, Nia05, NS87, Sab54, Sam60, Til93a, NS84]. Experimental [Bin90, CDCH^{+}24, Dan05, De 82, GBW61b, Jen88, KGB^{+}22, MTH^{+}24, Mck91, Rat20, Ros08, Til92, Til94, Van93b, Gua05]. Experimentalism [Ack89]. Experimentation [Ezr56, Spi57]. Experimenter [LE22]. Experimenting [Mck91]. Experiments [Bae19, Bon92, Cha11, Gan86, Lai84, Mor86, SH20, Web09, Boe22]. Expert [Sor84].

Explain [Bha86, Bok08a, BF12a, BF12b, Cha02b, ER96, Hea15, Ner79, Ros12]. Explained [Cal04, Kop01]. Explaining [Cal09, Dou23, Ern01, Gil93, Gou23, Kor91, Mor06, Nea95a, FB01]. Explanation [And10, Bak09a, Ban13, Bar19c, Bar62, Blu89, Cab20, Cam00, Can64, Can65, Cha02b, Chi18, Cla90, Cle11, Col66, CH17, D’A20, Del21, Die67, DW17, Eme15, Ero12, Fei22, FH16, Fra95, GE23, Gla21, Goo68a, Gor70, Gri88, Gri08b, Hay55, Hen14, HL11, Hut56a, Jar64a, Jon18, Kan88, Kat13, Kha13b, Kle92, Koe92, Kri18, Lan13, Lea68, Leh65, Leu14, Mag10, Mat17, Mil86a, Mor06, Nat17, Nel96, Pet96, Pin22, Pov18, Pra24, Res58, RK87, Ros01, San98, Sch57, Sea87, Shu19, Sob87, Sob95, Tre62, TP19, TP21, Wat52a, Wat57a, Web09, WMR^{+}22, Wil51, Wil03, Woo79, Woo00a, Woo00b, uY94, Yol59, Ach83, Dow00, Fra99, Hit99, HK08, Hun06, Lip04, Mat99, Woo03, Gly04, Lyo06a, Rub86, Sha17]. Explanationist [Hue09]. Explanations [And18, ARR15, Bra21, Cal09, Cal04, CK20, ES59, Gou58, Gre53, HP05a, HP05b, Mey21, Par21, Pin15, Saa18, Sch71, Sko14, Sko15, Tha98, WHB^{+}23a, Wit61a]. Explanatory [Bak03, Bat10, Bla20, Cla16, Coh16, D’A20, ET97, ES19, Gly15, Gri08b, JS19, Kar12, KDM21, McG03, Pin11c, Pla83, Plu05, Pot10, RS96, Saa11, Sch18, Tay23, Tre04]. Explicating [Nii82]. Explication [Gil09a, Kui83]. Explications [Bam93]. Exploration [Avi19]. Explorations [Bra99]. Exploratory [Bai87]. Exploring [Lun21]. Expose [Ale16]. Exposing [Gut00]. Expressionism [Lig14]. Expressions [Ern90, Wil73a]. Extended [LB18]. Extensibility [Sha03]. Extension [Car85, EL72, FG17, Lar04, Ryd81]. Extensive [Grü53b]. Extent [Wel08]. Externalism [Fra91]. Extinctions [WHB^{+}23b]. Extrapolation [Bae16, Dav56]. eye [May96].

F [Bar01, Bla63a, Bra63a, Hea86, Pri09, Rus00, Str09, W56]. F. [Pan63]. F.R.S [Ecc52]. Fabric [Pri99]. Faces [Den89a]. Fact [Nar74, Tho92]. facto [Zen77]. Factor [Bai87, BD86b, Gan86, Mul91]. Factorisation [Jam63]. Factors [Tre04]. Facts [BN94, Gar82, How84, Nii83]. Facts-Well-Put [BN94]. Fagan [Cur16]. Fails [Kin10, Nor08]. Failure [Cum92, Fra95]. Failures [Mac78c]. Fairbairn [Bal57, Fai57, Fou57, Sut57]. Faithful [Bar11]. Faithfulness [Eva18]. Fallacies [Hon91]. Fallacy [Bir22, Lar05]. Fallibilism [Wei93]. False [Hin88, Mil74b, Mor88]. Falsehood [Gel75]. Falsifiability [Din59, Grü59, Grü60, Grü61, Ryn83, Sim85, Sto91, Tur91]. Falsificationism [Grü76a, Rad91b]. Falsifying [Gil71, De 72]. Faraday [Wil67]. Fat [BG16a]. Fat-Handedness [BG16a]. Fatal [Dri11]. Fatuus [Van94c]. Feasible [Fri82]. Features [Bet51c, Bet51d, Kho20]. Federica [Ste12a]. Feeling [Puc67]. Feminism [Koe00]. Fermions [MS08]. Fertility [Seg08, Nol99]. Festschrift [Gar73b]. Few [Col15]. Feyerabend [Lar00, Fos69, Har59b, Set69, Tho94a, Zah78, Zah81, Zah82a, Zah82b]. Feynman [Lee94, Lee95, Pri94]. Fiction [Nar74]. fictions [Mac99]. Fidelity [Cha21]. Field [Bak09b, Bla55, CH01, DEO19, Din67, Fra21, Fra20, Hug00, Lup18, Mar00, Mil18, Pik54, Pik55, Pit11, Ros96, Seb22, Mil86a]. Fieldian [Mac99]. Fields [Bin19, HH21]. Fiftieth [Bir12]. Fifty [MG17]. filosofia [Gil03]. Final [CL19]. Find [Ney14]. Finding [Van94a]. Fine [CE04, Kop05, Mon06, Wei05, Goo64]. Fine-Tuning [Kop05, Mon06, Wei05]. Finetti [Bar04, FRZ12, Hei78, How08, Mil97a]. Finite [Grü69b, Tos16]. Finitist [Coo03b]. Firing [Wei05]. First [AGTZ11]. Fisher [Bai83, Joh87, Len06, Mor02, Mor06, Oka08].

Fisher/Pearson [Bai83]. Fission [Sch15]. Fit [Han05]. Fitelson [Ols02a, Str05]. Fitness [AL04, Aut20, BR04, BR20, Mar17, PR13, Ram13, RD17, SS93, Sim83, Wal96]. Fittest [Elt03]. Fitting [Bel17b, DeV97, For95b, Kuk95, Tur90, Kie97, Bel17a]. Fitzgerald [Din59, Grü59, Grü60, Pop66b]. Fix [Lau89, Wor89]. Fixed [Gin93, Wor88]. Fixity [Ahm14]. Flavour [Kno10]. Flavour-Oscillation [Kno10]. Flaws [Sch08c]. Flesh [Cur16, Fag13]. Floating [Fis06]. Floridi [Pre01]. fMRI [MD22]. Focused [Whe09]. Fodor [LM99]. Folk [Hoc17, Hoo86, Sti83]. Follow [VC24]. Followers [Ake86]. Following [Joh87, Kle17, Sie10b]. Footnotes [McC84]. Forces [Dor77, Ell63, Mas17, Sau64, Sni67, Wil07b]. Foreword [Edi50]. Foris [D’A86]. Forks [Tor87]. Form [Fle79, Ker87a, Sha05, Tho52]. Formal [Ern75, McW16, Pri09, Tia85]. Formalism [Jab92]. Formalized [Kre53a]. Formation [AA87]. Formative [Gat01]. Forms [MG65, ST22]. Formulae [Pot53]. Formulas [Goo10]. Formulation [JL22, San61]. Formulations [Bar19d]. Forschung [BK59]. Forster [Kuk95]. Fossils [Rud64]. Foster [Bee09]. Foucault [Sab54]. Foulkes [Fai57]. Fouls [Dur02]. Foundation [LP18, Lyn90, PR13]. Foundational [Bel01]. Foundationalism [Obe98, Ueb96]. Foundations [Alb94, AGTZ14, Bat15, Bel81, Bet52b, Coo03a, Dav05, Din60a, Din60b, Eel00, Hac64, Hug00, Kre53b, Kre58, Mar89b, May77b, May77a, Mor94, Ols03, Ots16, Ric77, Sla93, ST22, Whi50, Joy99, RS23, Skl13, Ueb99, Col00]. Four [Bri06, Gro09, JLZ05, Le 03]. Fourteenth [Cro53a]. Fozzy [Mac63a]. Fraassen [Cha11, Lad00, Lad04, Max93a, Max93c, Max93b, Mcm85, Pin11a]. Fragility [Tav91]. Frame [Sto86]. Framework [Vas21a, Wil14a]. Francesco [Ros08]. Francis [PR91, Urb82]. Frank [C.56a, C.56c, Pot51]. Franks [Pat98]. Fraser [Sch14]. Frechette [Ale16]. Free [Dew73, EL72, Fis06, Goo71a, LE70, Mcd72, Pla60, Sch74c, Tha70, Til00]. Free-Floating [Fis06]. Freedman [KW97, SGS97]. Freedom [HH21, Ste01, Wat71a, Wil10]. Frege [Mil89, Cur76, Cur82a, Cur82b]. Frequency [Daw85]. Frequentism [Str96, Tos16]. Freud [Pet56, Wal56b]. Fried [Dun11]. Friedman [Man76]. Frisch [Far16b, Vic08]. Front [Ano21b, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b]. Frugal [FRSW18]. Fugitive [Bun86]. Fuhrmann [BF00]. Function [Cal15, CT16, Gao22, Gar18, Gil90, KZH^{+}22, Kin10, Kin16, Man97, Pot53, Rud64, Wal96, NA13]. Functional [Ber16, Cha02b, FP66, Gri93, Mor09, Nis70, Pet96, Rea20, Wei11]. Functionalism [Kat84, Loa24]. Functionalist [Noo97]. Functionally [Bot94]. Functioning [TS53]. Functions [Bat07, Cha02b, GP14, Gri93, Lee21, MSM09, MP22, SMM11, VH03, Von83b, dC22]. Fundamental [Bet51d, HS17, Kav23, Lin93, Oka08, Sac64, San08, Sta09, Wal20, Why52a, Bet51c]. Fundamentality [CL19, HL11, McK14]. Funding [Avi19]. Further [BH56, Bog77, Car64, Hun69a]. Fusion [Dan05]. Future [Grü69a, Har68b, Pow12, Rak97, Sch22a, Sha03, Wei72, Gil03, Gal00b]. futuro [Gil03].

G [Byr00, Hal99, Ket00, Ket01, Wil14a]. G. [Dia64, Kap60, Kap61]. Galavotti [Cam00]. Galilean [Bin17, Mag03]. Galileo [Aga57, Bud97, Cha85a, Dow05, Fin75, Gen98, Ste74, Tho94a, Tho54a, Gal00a, Sco08]. Galison [Jar12]. Gallie [Har59a]. Gallois [Sid01]. Gambling [Col08]. Game [HZ13, Ngu21, PG13, Van08b]. Games [BS17a, Bar23, FC18, O’C19, Pur23, Wag15]. Gap [BM18, Fri04]. Garon [Bot94]. Garvey [Ste10]. Gases [dR96a]. Gather [Mah90]. Gauge [BB04, Teh16]. Gauld [Jon69a]. Gay [Rus00]. Gayon [Amu99]. Gedanken [NS87, NS84]. Gellner [Qui61, Fey76, Kör86]. Gene [Ewe11, LB18, Rey09, San08]. General [Cla53, Dob21, Gar08, JL22, Max93b, Rai86, Von50, Dor78, Kno10, Sac75, Sac76, Wei79]. Generalisation [Bal57]. Generalisations [Nii79]. Generalised [Smi10]. Generalism [Gli20]. Generalization [LT17]. Generalize [O’C17]. Generalized [Cha10, Pov18, Rad91a, RS15, Her07]. Generalizing [DW06]. generally [Wit58c]. generate [Lar99]. Generation [SMM11]. Generations [MP22]. Generous [GS08a]. Genes [Kit82, Vel10, Led87]. Genetic [Bab22, BD86c, BD86b, Gro09, Kit87, Pla16, JLZ05, Šus07, Ste16b]. Genetics [BD86c, BD86b, Gil14, Hal55a, Ots16, Was78, Oka00]. Genic [WC99]. Genomes [Bab22]. Genomic [Bae12, Sar15]. Genuine [SS20, Sko15]. Geodesics [Why53a, Why54b]. Geoffrey [Kut00]. Geometric [Mor88]. Geometrical [Gre10]. Geometricity [Kno10]. Geometrization [Sie10b]. Geometry [Bel00, Bin19, Cra69a, Dur89, Hut54b, Mun86a, Ner79, Ric77, Saa18, Sta85, Wal20]. George [C.56b, C.56d, Fes06, Kir01, Pit94]. German [Zah83a, BK59]. Gerrans [BG16b]. Gestalt [GSK13, RO55]. Gestalt-Switching [GSK13]. Get [Hoo93, Tah21]. Gets [Bla20]. Getting [Kho20, MA11, Cha13]. Ghirardi [AGTZ08, CM00]. Ghose [Dic01b]. Ghosts [Rub22a]. GHZ [KP99]. GianCarlo [AGTZ08]. Gibbsian [FW21]. Gibson [Sto70b]. Gilbert [Hes60b, Hes60c]. Gillies [De 72, Spi74]. Gingerich [Rai86]. given [PP78]. Glass [Han92]. Glassen [O’H84]. Glennan [Cas16, Geb22]. Global [Cur84, Val18]. Gluttons [Cur17]. Glymour [Dur89, Wil09a, uY94]. Go [Cuf18, Oka01, SS13]. Goal [Smo21]. Goal-Directedness [Smo21]. Göbel [Hea88]. God [Cra86, Cra90, Fos04, Hol98, KT93, Mon06]. Gödel [Goo69, Han71, Sle82]. Goes [Hoe98]. Going [Mcd72]. Goldilocks [Bla20]. Goldman [Fel01, Gan11]. Goldstein [Wat59d]. Golfer [For95b]. Gollancz [Qui61]. Good [Fra81, Gan86, Irz85, Lai84, Mor86, Sch22b, Gil90]. Goodman [Alt68, Cha00, Fos69, Saf86, Tel69, Wil79b]. Goodstein [Fin64]. Gottesman [Cuf17]. Göttingen [McL07]. Gould [Elt03]. Gouldian [Pea12, WHB^{+}23b]. Governed [Car12, Rob08a]. Government [Cap86, But84b]. Gower [Cha89]. Gradation [Buc16]. Graded [HH15, Mol22b]. Gradual [Bar13]. Grail [Rob23]. Graining [Rob20]. Grammar [Bar12, Gua07, Tim06, Bic06]. Grand [HF96]. Graphs [Oka16]. Grasp [Egg16, Enf08, Sta06b]. Grasping [Bel23]. Gravitation [Din67, Kap55a, New73, WY72, WY73]. Gravitational [Pit11, Rea20, Seb22]. Gravity [CL19, DTW17, Pit11, Teh16, Wal10]. Greek [Woo00b, Llo67]. Green [Fre88]. Greenstein [Hal99]. Gregory [And86, Wil12]. Grene [Car64]. Griffiths [Gra01]. Grosholz [Fre11]. ground [Gho06, Dic01b]. Grounded [Rub22a]. Groundless [Rub22a]. Group [Bir20, Den59a, Gly08, Oka01, Oka05, Ree78, Rob11]. Growth [Ake90b, Akk94]. Grue [Ake91, Hes69b, Hun69a]. Gruesome [DeV97]. Grünbaum [Cra92, Cra01, Din60c, Din61b, Dob70, Fer70, Hor75, Swi00, Swi05, Zen77, Grü05]. Grundprobleme [Zah83a]. Guala [Ros08]. Guicciardini [Far00]. Guidance [GC19]. Guide [How72, Ros96, Thé16]. Gut [Pri04, Gri08a].

H [And10, BD86a, Gil00a, Gly11a, Mar61a, Rey15, SL13, Tom99]. H. [Grü69a]. H.-C [And10]. Haack [Kus01]. Haag [Mil18]. Hacking [Pin16, Coo04, Jos82]. Haecceitism [Mel99b]. Haldane [Woo56]. Hallucinations [Cha58]. Halpern [FG17]. Halsbury [Bon55a]. Halt [Kuh80]. Hamilton [Bir14]. Hamiltonian [Cur14]. Handedness [BG16a]. Hanging [Hol88]. Hankinson [Woo00b]. Hanne [Tha09]. Hans [Hut53b, Le 52]. Hanson [Pau71]. Happen [Gly04, Woo03]. hard [Roc86]. Hardback [Bos06, Bri06, Cap86, Gia06a, Gia06b, Jen86, Per07, Fes06]. HARKing [Rub22b]. Harman [Smi00a]. ‘Harmonics’ [Tau06, Bar00]. Harmony [PG13]. Harold [BL14, Gal03]. Harrah [Gal59]. Harré [HM99, CM73]. Harrod [Pop58b]. Harry [D’A86]. Hartmann [Ols02a]. Hartry [Mil86a]. Harvard [Saf86]. Hasok [Gil09b]. Hattiangadi [Giu88]. Hausman [Car02b, Hit00, Ste06a]. Havi [Gla14]. Hawking [Sco08, BER99, Cra90, Dow05, Sac93]. hbk [Gly04]. Head [Har06]. Healey [Kut00]. Health [Bro19, CC13, Hau11, Kin10, Sch13, Wer19a, Gla14]. Healthier [Sch13]. Healthy [Sch13]. Hearing [Cle55]. Heerden [Din61a]. Heesen [Row23]. Hegel [Buc72, Zah87]. Hegelian [Aga58a]. Heil [Ney14]. Heisenberg [Boh62]. Hellman [Kut00, Bil00]. Helmholtz [Cla76, Ric77]. Help [Gła23, Pit08]. Helps [Hea15]. Hemispheres [Rob76]. Hemispheric [Bog77]. Hempel [Vra04]. Hendrick [Pri09]. Hendry [Bro86]. Henry [Pop81]. Herbert [Whi81]. Heredity [Amu99, MP22, NS20, Gay98]. Heritability [Bou15, Ses05, Oft07]. Hermann [H56, Ric77]. Hermes [Hug92]. Herring [Fre88, Goo67c, Goo68b, Goo69, Hem67]. Herrings [Mag10]. Herrnstein [Koe99].

Heuristic [Cho15, Nun93]. Heuristics [Hey16a, McG03, Mic81, Nic87, Rad91a]. Hidden [Bha86, Bub68, Dew73, Edw79, Fei15, Shi84, Lan98, Byr00]. High [FH16]. High-Level [FH16]. Higher [Fis06, Geb22, GE23, Kar12]. Higher-Level [Geb22, GE23]. Higher-Order [Fis06]. Highlights [Pry01]. Hilbert [Gia83, Kre53b, McL07]. Hillel [Pop55a, Pop56a]. Hintikka [CS98]. Historian [Coa74]. Historians [Bla79, Hes60b, Hes60c]. Historical [Bla85, Cle11, Cur17, DiS99, Goo85, Gou57, Gri96, Grü56, Mus74, Tuc11, Vic21, Wat52a, Wat57a]. Historicism [Urb78]. Histories [Gil07, Car03]. Historiography [McA86, RH72, Sie10b]. History [Aga58b, Bar62, Ber55a, Bor54, Bra76, Enf08, Fri02a, Gar23, Gie73c, Gou58, Gro09, Hac81, HH92, Kuh80, May60, Mcl90, Old82, Rai86, Sac93, Sch11a, Sma72, Van94a, VK91, Wig61, Wil65, Wil75b, Dor05, Fie86, JLZ05, PK70a, PK70b, PP78, Sta06b]. Hobbes [Jen88, PT57]. Hoc [Bam93, FS94, Grü76a]. Hoc-ness [Bam93]. Hocnes [Pop59b]. Hocness [Pop59a, Red78]. Hodgkin [Lev14]. Hogarth [Man21a, Wel08]. Hohwy [Ero12]. Hole [But89, EN87, Ryn94, Wea18]. Holes [DTW17, Dob21, Mel99b]. Holism [Dar12, Gal14, Tel86, Wei09]. Holistic [Ric19]. Holy [Rob23]. Homeostasis [Wal56b]. Homeostat [Adk51]. Homology [DiF23, Jar67a]. Homotopy [LP18]. Honest [BR20]. Honor [Pro86]. Honour [Cam00]. Hookway [Gib86, Put01]. Hope [Sch11b]. Hopelessly [RD17]. Horgan [Gar99]. Horowitz [Jar65]. Hostile [She06, Ste03b]. Hot [Cur17]. Hot-Blooded [Cur17]. Houkes [Pre03]. House [Gut00]. Huber [CFM08]. Hugh [Til00]. Hull [Coh74, PT57, Rus71]. Human [AH14, Car02a, Coo04, Far16a, Fod01, Gan86, GC19, Goo67a, Luc68, Mac73a, Pow12, Puc74, Red00, Sin00, Sla15a, Ste01, Kha13a, Ste03b]. Humanities [Gie75]. Humans [Lew14, Ste12b]. Hume [Gow91, Hol98, Lip02, Mil87a, Sch84b]. Humean [BCP93, Dar12, Gil91a, Hit03, Urb88]. Humeanism [Hic18]. Humphreys [KW97, SGS97]. Hung [And10]. ‘Hunting [Gly10a]. Hutchinson [O’H86]. Hutchison [Cal95]. Hüttemann [Sch08a]. Hutten [Han62b, Pan60]. Huxley [Lev14]. Hybrid [Kat13]. Hybridism [VM16]. Hypercomputation [Cot03, OK05, Wel04]. Hyperproof [Fri97a]. Hyperproof-like [Fri97a]. Hypotheses [Ach90, Col08, De 82, FW05, Grü76a, Kap55c, Lep82, She02, Str01, Str05, Tho94a]. Hypothesis [Ach62, Amu99, Das58, Din59, Eva02, Gla21, Grü59, Grü60, Han05, Joh88, May51, MLR21, Pol52a, Pop59a, Pop59b, Ros78, Sam98, Sta09, Tre62, Wal51, Wis51, Gay98].

Ian [Pin16]. IBE [Cab20]. Icon [Hor94]. Idea [Row23]. Ideal [Pap76, Wat52a]. Idealisation [Kör64]. Idealism [Bla79]. Idealization [Fra95, Pin22]. Idealizations [FS21, Lin93, Fra99]. Idealized [Ric19]. Ideals [Hoo91]. Ideas [Res66]. Identical [HY02]. Identities [Mil89]. Identity [Ach62, Ara11, Bar13, FR88, Fri15a, Lei13, Liz93, Mas01, Mor60, Mun94, Puc73, Tap00, Wil81a, vR13, Haw99, Pra12, O’M14, Sid01]. ideology [Gel59]. Idiom [McK20]. Idiot [Bun86]. If [HL24, Lau89, Lev69]. Ignis [Van94c]. Ignorance [Mel65, Mel67a, Ols02a, Sai95]. Ignore [Lan65]. II [AK95, Ben57, Bir56b, Bri51a, Bun55b, Bun56b, Coh66b, Din60b, Dob51b, Fra50b, Goo61b, Goo61c, Gre58b, Hal79d, HP05b, Han62a, Hes60c, Hul65b, Kes65b, Kes66b, Lak63b, Mar61c, May77a, May80b, Pan62b, Pik55, Pop50b, Sch62, Sch52b, Spe66b, Urb74b, Wil68b, Zah73b]. III [Bri51c, Kes66a, Lak63c, Mar61d, Mee04, Wit61a]. Illegitimate [Pet22]. Illness [Gla14, CC13]. Illusion [Jen16]. Illusions [Lev85]. I’m [Mcd72]. Image [Har94, Wor84, Mur06, Bro09b]. Images [Kle10, Phi09, Smy58a, Le 07]. Imaginary [Goo94, Sam60]. Imitation [FC18]. Immobile [Hug08]. Immunology [O’M14, Pra12]. Impact [BK59, Cha03, Pet23]. Impasse [Tod09]. Impenetrability [Smi07b]. Impera [Lyo06b]. Imperative [Car23]. Imperial [Per07]. Implementation [CT23, Res13]. Implications [Cal50, Coa74, Dew75, KT93, KOV06, Sti94, Wer09]. Implicit [Win67]. Implies [HLWW17, Sch09]. Imply [Pet20, TH20, GB99]. Important [Mor86, Wil84]. Impossibilities [Hal01, Ver99]. Impossibility [Cra79, Eel88, Red85a, Spr18, Whi78b]. Impossible [Van94b]. Imprecise [Dod13, FC18, Val18]. Imprecision [Swi69]. Imre [Fey75, Hac79, Lar00]. Inaccuracy [RS18]. Inadequate [MS07]. Inapplicability [Mar90]. Inborn [Woo53b]. Incapacitation [Tul12]. Incautious [Erw82]. Including [Lar00]. Inclusive [BR20, Mar17]. Incommensurability [And10, Fey87, San93, Sie80, Szu77, Hun06]. Incommensurable [Fra84a]. Incommunicable [Mar59a, Mar59b]. Incompatibility [New73, WY72]. Incompatible [Fei15]. Incomplete [Liu22]. Incompleteness [Fin64, Goo63, Nor19, SS13]. Inconsistency [Mil18, ST22, Vic08, SV11]. Inconsistent [Fri16, MB21]. Indefinite [Sha03]. Independence [Aga90, Ber16, COS23, HW99, Pop60, Pot10, RK87, Wei15]. Independent [Nor09]. Independently [Cli95]. Indeterminacy [Ber23, Bub00, Gar73a]. Indeterminate [Liu22]. Indeterminateness [Mac73a]. Indeterminism [Bra62a, EN98, Hoe69, Hut54a, Lar97, Pop50a, Pop50b, Ros56, Ste05, Lar99]. Indeterministic [Sch07a]. Index [Ano86a, Mel70]. Indexing [Ech16]. Indifference [Cas98, HLWW17, Pet20, Smi17, TH20]. Indigenous [Lud17]. Indirect [KL17]. Indiscernibility [CB12a, Dis94]. Indiscernibles [FR88, Mas01]. Indispensability [Ban13, Gen98, Saa11]. Indistinguishability [Wil23b]. Indistinguishable [RT92]. individual [Mat99]. Individualism [Bon98, Cur84, Gol58, Ree78, Sco61, Wat52b, Wat59c]. Individuality [Mol22b]. Individuals [DiF19, Rey09, Rus87]. Individuation [Dew18]. Induction [Aga59a, Aga90, Bar85a, Ben56, Ben57, Boe22, Che91, DF89, Das58, Dou23, Ell88, Fey68, FA14, Gra74, Har60, Hob94, Hut58b, Juh94, Kne57b, Kyb58, Leh70, Lip02, Mac91, Max93a, Max93c, Max93b, MS06a, Mil68, Mil87a, NW22, O’N89, Pop58b, Riv94, Row93, Sch64a, Sch00c, Set90, Swa88, Tre04, Vol91, Wat91, Wet94, Cum92, Fos04, Sch00b]. Inductions [Mag10]. Inductive [Bai83, BH53, Bla66, CB69, Eel88, Elb94, Gem89, How75b, How75c, How87a, Hue09, Kui83, Kuo24, Lev79, Mil90a, Nor19, Red85a, Sär68, Smi17, Smo77, Vas21a, Vas22, Wil07a]. Industrial [Pog85]. Inequalities [Ang02, Bha91, Shi84]. Inequality [GPW05, TBR18]. Inequivalent [Bar19d, JHB^{+}23, CH01]. Inertia [Sam18, Val18, Zim08, Whi51b, Why51b]. Inexact [Cle70a, Kum67]. Inexactness [Swi69]. Infant [Ses07]. Inference [Bai83, Bar72a, Bar72b, Bux78, CB23, Dor72, Dor75b, DE76, DW17, FRSW18, GH16, Goo84, Hen14, How87a, Jon18, Kat13, Len06, Lip04, Mac14, MW14, MD22, Mil01, Mor90, Nor19, OL23, Pra24, Rud64, Sär68, Sob98, Ste09, Tei13, TP19, Tuc11, Vas21a, Vas22, Web09, ZG21, DTV11, Lyo06a, TP21, Gil01]. Inferring [Sch00c]. Infinite [AK91, AK95, Bel79a, Coo03b, Cra79, Dav01, Gil80, Pop78, Pot96, Smi10, Whi78b]. Infinitely [Fis79]. Infinitesimal [BHW18, CKKR88, May00, Wis53a, Her07]. Infinitude [Grü69b]. Infirmities [Spi74]. Inflated [Bou04]. Inflationary [KOV06, McC19]. Influence [WOB21, Wil07a]. Influences [CBR90, Sta90]. Informal [Tia85]. Informants [Gie63]. Information [AH14, BHC53, BR68, Das23, HO79, Isa19, Kie97, Mac66, Mił23, Mil66b, Mul91, Par11, Pop66a, Pop66c, RS18, Ste09, Van81, VHH86, Wil80b, BER99, Sky10, Šus07, WF13]. Information-Theoretic [Par11]. Information-Transforming [Mul91]. Informative [Mil89]. Informed [Col15]. Inheritance [Mam04, Mun94]. Inherited [She13]. Initial [Cal04, CBC^{+}22, JL22, Lar03, Wer19b]. Initial-Condition [Wer19b]. Innate [Res66]. Inner [Cla07]. Inquiry [Gri08b, Sch00a, Bon01]. Inscriptionalism [Gor70]. Inscriptions [Sch71]. Insight [Bun68a]. Insights [Sca10]. Insisting [Mag04]. Instability [Swa22]. Instantial [Blu70, Das58, God69]. Instinct [Tho54b]. Instrumentalism [Bon58, Mcm85]. Instruments [Bai87, BF90, Mul91]. Insufficient [Vas21b]. Insurance [FRZ12]. Integrated [BM18, Gir76]. Integration [Man21b]. Integrative [Bar19b, Fei54, Tro93]. intellectual [Koe99, PP78]. Intelligence [Bod73, Hel87]. Intelligibility [Hoo91, Lac70]. Intention [Kle17]. Intentional [Ada01]. Intentionality [Nad08]. Intentions [Gil93]. Inter [Skl67]. Inter-Theoretic [Skl67]. Interaction [FS21, Gly08, Oka05, SW19b, Sme15]. Interactionism [Van87]. Interactions [Smi07b]. Interactive [Kha10]. Interdependence [Pot10]. Interdisciplinary [Hea99]. Interesting [Kil91, Lar05]. Interestingly [Cus90]. Interests [Coh63a, Wag15, Sin00]. Interface [Ech16]. Interlevel [SH20]. Internal [Her07]. Interplay [Sco08, Dow05]. Interpretation [Ach63, Bae12, Bar11, Bor53, Bra55a, Bra11, Bub68, Bux78, Cli95, Dic95, Fin73, Fra80, Hol98, Jab92, Kap55a, Llo67, PR13, Pop59c, Pot51, Tiz57, Wal12, Why51a, Ver99, Wal06a, Hal01, Vai15]. Interpretations [Bak09b, Cli96, Fra50a, Fra50b, HP03, Loc96a, Loc96b, Loe98, Ver05, Grü00]. Interpretative [Bro20, Cof14]. Interpreting [Rue11, Skl00, Tha98, Wal14, Low00, Sto99]. Interrelations [HH92]. Intersubjective [Gil91b]. Intertheoretic [Fri82]. Intervals [Joh88]. Intervening [MG94, Sob98]. Intervention [Til93a]. Interventionism [Bla20, SE23]. Interventionist [FH16]. Interventions [Cas16]. Interview [HHS02]. Interwar [Gat01]. Intimate [Gie73c]. Intimations [KK83, Urb83]. Intransitivity [Rox75]. Intrinsic [Mil86a]. Intrinsicality [HL11]. Introduction [Bec15, Dor71, Gow93, Gra01, Hof01, Sch11a, Dor05, Jen86, Pre01]. Introspection [Mac05, Mor24]. Intuition [Bon01]. Intuitions [Dev06, Dev10, Fit10, GC11, Kui83]. Inus [Bel05a]. Invariance [BM93, MS55, Mil83a, HW99, Woo00a]. Invariant [Sav94]. Invariants [Bla50]. Inventing [Cha04, Gil09b]. Invention [Zah83b]. Inventor [Lai91]. Inverted [Bla79]. Investigating [Fow83]. investigation [Row11, Gil13]. investigations [Wit58c]. Ion [Goo13]. Iowa [Col08]. ‘IQ [Dea76, Tiz76, Urb74a, Urb74b]. IR [Bat71]. Irrationalist [Gie71]. Irrationality [Wil22]. Irrealist [Hac88]. Irrefutable [Wis63b]. Irrelevant [Sch70]. Irreversibility [Cla82a, Cla82b, Den89a, Pop57a, Zen77].

Irvine [Koz14]. Isaac [Coh86a]. ISBN [Bai86, Bos06, Bri06, Bro86, Cap86, CE04, D’A86, Den86, Fes06, Gia06a, Gia06b, Gib86, Gil03, Gil07, Gly04, Hog86, Hoo86, Kör86, Lyo06a, O’H86, Per07, Rai86, Roc86, Rub86, She06, Tau06]. Islam [ADS84, Hog86]. Ismael [Str96]. Isn’t [Ber23]. Isochrony [Chr77]. Isometry [De 82]. Isomorphism [Mad57]. Issue [Cha03]. Italian [Gil03]. IV [Lak64]. ix [Fie86, Rub86].

J [Amu99, DiS99, Gar99, Gro09, Haw99, Hea99, Hei99, Hog86, Lev79, May00, Mel99a, Mil66a, Pol14, Sto99, Woo00b]. J. [Bra93, Hon91, O’D90]. Jablonbka [Gro09]. Jacques [Hac71b]. Jakob [Ero12]. James [BL14, Hof01, Smi00a, Tom99, Fol99, Gly04, Eel00]. Jardine [Fie86]. Jarrett [Lad99]. Jaynes [LM85]. Jean [Hud20]. Jeffrey [Den86, NR13]. Jeffreys [Gal03]. Jennings [Kei93, Tri94]. Jeremy [Dic01a]. Jerrold [HM99, Mel99a]. Jerry [LM99]. Jesper [Ero12]. Jesse [Gri08a]. Johannes [She01]. John [Bro86, Hog86, O’H86, Bee09, Car12, Cla14, Fri09a, Hoe00, Ney14, Pas16, Wor99, Bon00]. Johnny [Haa91]. Johnson [Han92]. Joints [Lyo01, SL13, COS11]. Jon [Gly09]. Jones [HK68, Lev69]. José [Car99b]. Joseph [Pop81, SL13]. Jourdain [Hal81]. Journal [Cla00b]. Joyce [Sch14, Eel00]. Jr [BD86a]. Juarrero [Ada01]. Judea [Gil01]. Judgement [Hey16a]. Judgments [Kla64, Mee64, Nad08, Ros55]. Judy [Vas21b]. Jumps [Sch52a, Sch52b, Whi53b]. June [Cla00b]. Junk [Sha97]. Just [Bla20, FS24, Fre88]. Justice [IK24]. Justifiably [Grü76c]. Justification [Bux78, Ham23, Kan88, Lip02, McC19, Mos91, Sch64a, Sho00]. Justifying [Rob20, Vas22].

K. [Gil07, Gra01]. Kalām} [Lin23, Pit08]. Kalam [Cra01]. Kallestrup [Ero12].

Kaluza [Kar12]. Kampis [Fes06]. Kane [Den86]. Kant [But84b, Dem01, Ste00, Buc65, But84b, Cap86, Pot00, Pot02]. Kaplan [Mor99]. Kapp [Bon55b, Dob58b]. Karl [Gat01, Gil81, Hac81, PP78, Wat94]. Katz [Mel99a]. Kauffman [San08]. Keeping [Cal23]. Keim [SL13]. Keith [Bat71]. Ken [Bir13, Sky12]. Kenneth [Nad99]. Kepler [Fie86]. Keswani [Din65a, Pod75]. Key [Jar64b]. Keynes [Bra93, Coh86b, Gil88, O’D90, Ram89]. Khalidi [Sla15a]. Kim [Hei99, Hor96, Lew14, Sch14, She06]. Kin [Bir20, Dem01, Oka16, Pot02, Pot00]. Kincaid [BL14]. Kind [BEL92, Cha02a, HW18]. Kindhood [Buc16]. Kindness [Sla15b]. Kinds [Bar16, Bet24, Bur18, CB12a, Coo04, DiF19, HW18, Hav18, HR22, Kha10, Lai21, Lud17, Mel77, O’C19, Pla83, Rus87, SL13, Sla15a, Wei11, COS11, Kha13a]. Kinematics [BM93, HF94, Lev67, Van81, Why53c]. Kinetic [Nyh88, dR96a]. Kirchhoff [SV11]. Kirk [Llo67]. Kistler [Mum13]. Kit [CE04]. Kitcher [Dou13, Kar12]. Klein [Kar12]. Kluwer [Fes06, Gia06a]. Knew [Lev69]. Knill [Cuf17]. Knot [Puc78]. Know [AH03, Hoo93, Mac05, Mcd72]. Knowing [Avr02, Cho11, Gia06a, Gia06b, Mat05, Car04b, Cha07a]. Knowledge [Ake90b, Bar12, Boh61, Coh86a, Cur16, D’A77, DeR02, Dem03, Dem11, Egg16, Fre88, Gri06, Gro85b, Hor65, Jen05, Kat84, Koe04, Mor00, Pas15, PLKG24, Ram90, Wat60a, Wat87, Wil98, Dav89, Fag13, Sch79a, Zah83a, Fel01, Ueb01]. known [Wit58c]. Kochen [Cli95, Gar72a]. Koertge [Gut00]. Kolmogorov [Fri04, Leb89]. Koyré [Aga58b]. Kristin [Ale16]. Kuhn [And10, Bro02, Gal81, Hal70, Hun06, IG95, Jon86, Kat84, Mus71, Pea82, RW13, San93, Wra11]. Kuhnian [Sto91]. Kuhn’s [May96, HH92]. Kuhn’s-eye [May96]. Kukla [Pag00, For95b]. Kumar [Kör68].

Kvasz [Fes06]. Kyburg [BD86a]. Kyle [Enf08].

L

[HM99, May00, Mil66a, She01, Sin00, Whi73]. L. [Ber78, Lev79, Whi73]. Lab [Cur21]. Labels [RT92]. Laboratory [Bla55, Hac88, Wei92]. Lacey [Til00]. Lacuna [Vra04]. Laden [Bro93, FAB^{+}89]. Ladislav [Fes06]. Ladyman [BL14, Han04]. Lagrangian [Cur14]. Lakatos [Bra99, Fes06, KKS02, Lar00, Ake86, Ake90a, Fey75, Hac79, Kul77, Red89, Tho92, Zah83a, Zuc88]. Lakatosian [Nic87]. Lamb [Gro09]. Landauer [Par11]. Landé [Dor76a, Gar73b]. Landini [Byr00]. Lane [McL07]. Lang [Den86, Gil00a]. Lange [Yud13]. Language [BH87, D’A77, D’A86, Fod01, Fre53, Mac05, Mac62, Qui57, Res09, Row94, Sam80, Tan89, Bra84, For99, Hea99]. Language-Acquisition [Sam80]. Languages [Pos74, Ryn83]. Laraudogoitia [EN98]. Large [Hac88, Red00]. Larry [Gro90]. Laser [Mor94]. Last [McL07, Sma86]. Lasting [Mcg70]. Lastis [Coa74]. Latency [Hol86]. Latent [Isa19]. Later [Cou52a, Ein50a, Ein50b]. Lattice [Goo59a]. Laudan [Ake93, Gro90, Mar89a, Oka97, Wor89]. Laura [Wal14]. laureate [Jos10]. Lavoisier [Zuc88]. Law [Bar19c, BS94, Buc51, Car12, HS22, NS67, Rob08a, Rob23, Smi07a, Urb88, WS23, Whi51b, Why51b]. Law-Abiding [WS23]. Law-Governed [Car12, Rob08a]. lawmaker [Fos04, Bee09]. Lawrence [Bel01, Bat15]. Laws [Bal92, BEL92, Blu70, Buc65, Dor19, Dre01, GL23, Hau13, HS17, How73, Jam23, Lan76, Lan06, Man57, Mor09, Mum04, Mum13, Old78, PR95, Pri76, Pro59, Rob08b, Sch11a, Sch00c, Shu19, Sud73, Sud76, VK91, Why57, Wis63b, Woo90, Woo14, Dor05, Fos04, Her07, Kis06, Rup07, Sch07c, Yud13, Bos06, Reu09, Sch01b]. Lawson [Wat87]. Layered [Gre17]. Leap [HF96]. Learn [Gal14, Sti81]. Learned [PR91]. Learning [Ale14, Cha73, Goo74b, Sme15, Sul22, WF13, Sky10]. Leaves [Bla89]. Lecture [Red89]. Lectures [Lar00, Fos04, PP78]. Leeds [Pri94]. legacy [Dem13, Hal15]. Lehrer [Bat71]. Leibniz [Art94]. Leibnizian [Cra01, Fut05]. Leng [Wei14]. Length [Dor68, Wei79]. Lenin [Jos10]. Leon [C.56b, C.56d]. Leonardo [Pir52b]. Leplin [Lad99, Oka97]. Less [FS94]. Lessons [Cha11, Ryn94, Wor00]. Letters [Cur82a, Cur82b]. Level [Cla16, FH16, Geb22, GE23, Kri18, Mil06, Oka04, Sam75]. Levels [Fei54, GS08b, Sob01]. Leverage [Bar08b]. Levi [Coh86a]. Leviathan [Jen88]. Levinson [Kes66b]. Lewens [Hey16b, Mat03]. Lewis [Her07, Hub11, Pur23, Tap08]. Lewis-Type [Pur23]. Liar [Scr63, Sle88, Wri83, Pos74]. Liberal [Mar62]. Lie [Set70]. Life [All01, Cla14, Con03, ES59, Fle79, Jen88, Lew04, Nea95b, Rus50, Ste95, Vel10, Wei92, Wis53b, Dup14, JLZ05, Jos10, Gro09]. Lifeless [Kap54]. Light [BM93, Ern90, Kno14, Mag65, MP22, NS84, Nor86, Øhr86, Pod79, Sab54, Tha70, Why53c, Zah83a, NS87]. Light-speed [BM93]. Like [BL23, Mac53a, May52, Pir52a, Chr81, Fri97a]. Likelihood [CCT13, Gan15, Gib69, Hac72, Kot12, SS17]. likelihoods [Her07]. Likely [Wat89]. Lill [Cha13]. Limitation [Hal81]. Limitations [Dau88, Lat76]. Limitative [LU09]. Limited [Hob94]. Limits [BGH^{+}23, CE04, DeR02, GE23, Lup18, MW14, O’M14, RKK19, Zem87, Fin02, Pra12, Wor00]. Lineage [Bab22, Cal09]. Linguistic [Cur76, Dev10, Fit10, GC11, Smi59, Gel59]. Linguistics [D’A76, Dev06]. Linguists [CG09]. Link [Car02b]. Lipton [Lyo06a]. Listening [Mac99]. Literacy [Fey76]. Literalism [Luc91]. Literary [Bus52b, DB52]. Literature [LRP23].

Little [Goo74b]. Living [Kap54]. Local [Elb93, Ram23]. Localisation [HF85]. Locality [But07, Fri02b, Mau00, Súr00]. Locating [Wit00, SC18]. Location [Sch90]. Lockwood [Deu96, Loe96, Sau96]. Logic [Bai87, Bal14, BH53, Bar72a, Bea89, Bel86, Bri82, Bub79, Buc51, CB12a, Coh66a, Coh66b, Coh67b, DF89, Dal89, Ell73, Fey76, Fra87, Gär84, God69, Goo67a, Goo52, Gro85a, How75b, How75c, How97a, Hue09, Hut65a, Jen89, Kap65, Kei93, Kel00, Kle70, Kum67, LaC22, Lan56, Lej54, Luc68, Mil97a, Mun62, Oli00, Sch74a, Sch74b, Sim65, Sla93, Smo77, Sta85, Ten81, Ten85, Tri88, Tri94, Yux90, Zah83b, Chi98, Fol00, BK59]. Logical [Bet16, Cle70a, Dew73, Gie73a, Hal55a, How73, Hul77, Ker87a, LaC22, Mac73a, Mad57, Mil90b, Mus74, Nel96, RO55, Sac76, Sto66, Wil73a, Zem87, Ryc99]. Logically [Sor07]. Logicism [Dem13, Gal03, Mac03, Mil89, Mus77, Hal15]. Logicist [Sha03]. Logico [Fre88]. Logico-Probabilistic [Fre88]. Logics [Bri82, Gar72b]. Logik [BK59]. London [Bos06, Fes06, Lyo06a, Per07, PP78]. Long [Jos82]. Look [AS19, BL23, Mar69, Res66, Rob76, Str95]. Looks [Hul77, Put05]. Lord [Bin54, Bon55a, Smy62]. Lorentz [Bor61, Bra22, Din59, Grü59, Grü60, Kes60, Mil74a, Pop66b, Pro74, Ros78, Sch74a, Zah73a, Zah73b, Zah78]. Lorentzian [Dor68]. Lorraine [Jar12]. Losing [CM99]. loss [BER99]. Lotka [KL17]. Lottery [Bar04, Cha10, DW06, Smi10]. Louis [RH72]. Lowe [Haw99]. Luciano [Pre01]. Luck [McC19]. Lucky [Mor12]. Luis [Car99b]. ‘Lying’ [Kri61].

M [Bra93, D’A86, Die99, Dre00, Eel00, Hit00, O’D90, PS00, Rad00, Smi00a, Spr11]. M. [Car64, Wer10]. Mac [McL07]. Macdonald [A.56a]. Mach [New73, Bla85, Bla89, Fey80, Pop53, Sac72, Sac75, WY72, WY75, Zah77, WY73]. Machamer [Gal00a]. Machian [Mar22, Zah81]. Machine [Buc23, Cla66, Gau66, Goo67a, Jon69a, Luc68, LS94, Sul22, Tho75, Tho65]. Machines [Cla72, Dav01, Foz63, Gib86, Kap54, Man21a, Mil57, Peć89, Puc67, Puc74, Sza58, Hoo84]. Machlup [Lat76]. Mackay [Dew73, Foz63, Wat71a]. Mackie [Cox75, Mil66a]. Macro [Aga63]. Macroeconomics [Hoo93]. Macrophysical [WY75]. macroscopic [GB99]. Mad [Cow98]. Maddy [Rol07, Sie10a, Ros99]. Made [Lew14, Wil19, Ste12b]. Madness [BG16b, Coo05, Tso10, Coo02]. Magnitude [Bec15]. Maher [Lan01]. Mainstream [Pri09]. Make [Bec19, CPG20, Lee82]. Makes [Fra81, Gal57, Lan13, Pol58, Tod09]. Making [BJ04, BF23, Lav23, Ney09, Noo97, Ses05, Ste03a, Tal16, Woo03, Bra93, Gly04, Oft07]. Maladapted [Fre10, Ric07]. Malament [Man21a, Wel08]. Man [Gea65, How72, KFH65, Rob76, Stu63, Fes06, KKS02]. Mandelbaum [Wat65a]. Mandevillian [Pet22]. Maneuver [Sko07]. Manifesto [Kus01]. Manipulability [BG16a, Car06, Leu12]. Manual [Coo05, Tso10, Coo02]. Many [AGTZ11, Cho15, Den89a, HP03, Pap96b, Loc96a, Loc96b]. Maps [Res09]. Marbles [BG01, Lew03, CM99, CM00]. Maré [Gil72a]. Margaret [A.56a, Gro11]. Margenau [Hol86, Mas01]. Margin [Lis04, Sai95]. Maria [Cam00]. Marigold [Pre03]. Marinoff [Arm93]. Marion [Gro09]. Mark [Mor99, Sim01]. Marker [Col08]. Markers [LB13]. Markov [Car02b, Car06, HW99, HW04, Ste05, Ste06a]. Markovian [Zha13]. Markus [Reu09]. Marr [Fra91]. Marras [Mor09]. Marriage [Gie73c]. Mars [Cha85a]. Marshall [Dob61]. Martin [Sch00a, Bir01]. Mary [Roc86, Wei14]. Mass [Leh11, Mar22, Seb22, WHB^{+}23b, CM00]. Masses [Dor77]. Massimi [Kra09]. Massive [Pit11, Sam98, Sch08c]. Massively [Lyn90]. Massless [Pit11]. ’mathematical [Bil00, Fol99, And18, Bak09a, Bar19c, Cla16, D’A20, Dav03, Ern90, Far00, Goo52, Gro85b, Hal79c, Hal79d, Hel98, Hol94, Kle02, Lan13, Lom02, Mad86, MHB^{+}23, Mor93b, Mos85, Pik54, Pik55, San98, Sko15, WHB^{+}23a, Yux90, Hau99]. Mathematician [Bat07, McL07]. Mathematics [Akk94, Bat10, Bel81, BF12a, BF12b, CS98, Coo03a, Fes06, Fra87, Fre11, Gia83, Gil78, Gil80, Ham23, Hof01, Kre56, Kre58, LP18, Lak76, Len10, Mac83, Mac93, May77b, May77a, MB21, McL07, Men54, Mil94a, Mol23, Oli00, Par82, Pas05, Pas15, Pin11c, Pin11b, Pin16, Pro77a, Red04, RK87, Rol07, Saa11, Sha82, Sim01, Smi07b, Tal13, Woo52a, Zah80, Zyc89, Bil00, Fol00, KKS02, Bak15, Wei14, Die99, Ros99]. Mathias [Far16b]. Matteo [Yud16]. Matter [Ano21b, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b, Cla82a, Cla82b, Gly11a, Kap55c, Rat04, Sie10b]. Matters [Ros21, Saf86, Lev01]. Matthen [Gia06b, Lew01]. Matthew [McK16, SL13, Rey15]. Maudlin [Ney11]. Maureen [McC16]. Maurice [Ste74]. Mauro [Sch11a]. Max [Mum13]. Maximizing [Mar17]. Maximizing-Agent [Mar17]. Maximum [How75a]. Maxwell [Per07, Sma00, Squ89, Tur55, Wil79b, Smi00a]. May [Mar57, Row23, Bai69]. Mayr [Set89]. McCaffrey [ZG21]. Mean [Woo53b]. Meaning [Ber55a, Ern75, Ern90, GG86, Gia06a, Gie70, Lep69, Mac71a, Par71, Wag12, Wag15, Car04b]. Meaninglessness [Pre61]. Meanings [Cho15, vR13]. Means [Aut20, BH15, Sch99]. Means-ends [Sch99]. Measurability [Rob08b]. Measure [BG16b, Fri04, Gil90, NS84, NS87, Row13, Sch11b]. Measurement [Din50b, FS24, Gil09b, Kut00, Mic94, Mil86a, Mun86b, Par17, Rob94, Sta20, Suá04, Tal16, Ves20, Cha04, BD86a]. Measurements [Gao22]. Measures [Cal04, Coh16, CCT13, Irv12, KS19, Spr18]. Measuring [Ste12a]. Mechanical [Ash52, Elb93, Hal52a, Hic53, Mac62]. Mechanics [Alb94, Alb62, AGTZ08, Ang05, Bar05b, Bar11, Bar15, Bar19d, Bor53, Bro86, Bro09a, Bub89, Bub00, Buc51, But06b, Cal95, Cal23, CBC^{+}22, Cli95, Cli96, Cur14, Dic01b, DB15, Dor77, ER96, EHLD14, Fei15, Fin73, FW21, Gie73a, Hal01, Har22, HP03, HY02, Hol86, Hut93, Hut95b, Jab92, KZH^{+}22, Lan59, Lav77, Lew97, Lew06, Lew07, Loc96a, Loc96b, Mag04, MS55, Mee22, MS07, Par11, PB02, Put05, Red74b, Red74c, RT92, Rob20, Sav94, Sce91, SC18, Skl00, Smi07b, Str96, Tel86, Thé12, Ver05, Von73, WT10, Wei01a, Why58c, Wil13b, Gho06, Hen84, NA13, She99, Sty00, Ver99, Ang02, Bar01, Cal15]. Mechanism [Bae12, Han71, Mac63a, Sha17]. Mechanisms [Bec19, Cal09, CHvI15, Gar23, Lee21, Leu12, Pir53, Ros21]. Mechanistic [Cas16, CK20, EL72, Geb22, KK17, Kri18, LE70, Mac71b, Mat17, Rat20, Win21a, Win21b, Mey21]. Mediated [Sho17]. Medical [Ahm05, Coh80a]. Medicine [MR94, Til93b, Wig56]. Medieval [Hog86, ADS84]. Meets [Lin23, Rad91b, Van08b]. Melinda [Cur16]. Mellor [Gly11a, Edg97, Pri84, Rig91]. Member [Hul74]. Memes [Ste06b]. Memory [Hoe00, Mar60, Naj24]. Men [Mil57, Sza58]. Mendelism [Mor02]. Menger [Gil81]. Meno [Sha84]. Mensuration [Mil90b]. Mentaculus [Gla24, Lin23]. Mental [All92, Coo05, Gou23, Hei99, Puc89, Sil91, TS53, Tso10, Wel93, Coo02]. Mere [AH14]. Merrilee [Cam00]. Merwe [Pro86]. Mescaline [Smy53]. Messy [Hav18, RS96].

Meta

[Col07, Cum92, Hey16a, Yud13, Fod01].

Meta-Cognition [Fod01]. Meta-heuristics [Hey16a]. Meta-induction [Cum92]. Meta-laws [Yud13]. Meta-scientific [Col07]. Metabiology [Wis53g, Wis54a]. Metabolism [All01, Bod99]. Metamathematics [Bea89]. Metamers [KH10]. Metaphor [BJ04, Mac71a]. Metaphorical [PNSBC94]. Metaphysical [Fra50a, Fra50b, Mac72, Puc62]. Metaphysician [McC54]. Metaphysicians [PT57]. Metaphysics [Cal95, Cal15, Car06, CB12a, Cho11, Dic07, ELLH17, GJ88, Gli20, Han92, Har22, KK17, KT93, Lad00, Lad04, Low04, MT13, Ney11, Reu09, Rey15, SL13, Tuo89, Wat75, Yud16, COS11, Cha07a, Haw99, Lil99, Mau07, Mor13, NA13, Sch07c, Sla13, McK16, BL14]. Metatheory [Mee04]. Method [Ada94, Bla09, Car06, Cur17, Gow88, Grü53b, Grü76c, Koe72, Kre53a, Lar00, Mee04, Mue69, Nol87, OK05, Rus51, Tau06, Tei62, Tur55, Von83b, dC22, Bar00, But84b, Mad07, Wor00, Ket00, Ric16, Sie10a]. Methodological [Gin93, Gol58, GBW61a, GBW61b, Hut54b, McC19, Ree78, Sco61, Tho54a, Wat52b, Wat59c, Wit61a]. Methodologies [Cur86]. Methodologist [Gil88]. Methodology [Boh61, Bra99, Cas58, Chi57, Cus90, Fes06, HZ13, Hut55d, Mil74a, Pog85, Ros08, RS23, Slo93, Sma72, Wor88, Yux90, But84b, Gua05, KKS02, Cap86]. Methods [Bun86, Far00, Gly94, Jar61a, Was55, Gly10a]. Methuen [Hea86]. Metric [Hor75]. Meyerson [Zah80, Zah87]. Michael [Dem01, Den86, Fes06, SL13, She01, Bon01, Gil03, Kan84, Tal15, Wil67, Wil14a]. Michela [Kra09]. Micro [Aga63, Pau71, RS23, Sch61a, WY75]. Micro- [WY75]. Micro-Entities [Pau71]. Micro-foundations [RS23]. Micro-Reduction [Sch61a]. Microbiology [McC16]. Microexplanation [Dow08]. Microphysicalism [Sch08a, Hüt04]. Microphysics [Bra62a]. Mid [Wis63a]. Mid-Century [Wis63a]. Mild [Wal07]. Mill [Hon91]. Miller [BR68, DH86, Goo70, Hor78a, HO79, Mac66, Mor88, Pop66a, Roz69, Zah78]. Mills [Pit11]. Milne [Grü53a, Grü54, Whi54b]. Mimi [Tei13]. Mind [Ara11, Bel78, Ben90, Blo70, BG16b, Bra84, Buc16, Cla72, CEH21, Dew75, Fod01, GS08a, Gre51, Hei99, Hor96, KT93, Lev01, Loc96a, Loc96b, LS94, Lyo01, Mac53a, May52, Pir50, Pir52a, Pir54, Puc78, Puc81, Puc93, PS00, Red04, Saf86, Sha88, Sle82, SD93, Urb77, Wat77, Wat51, Wil75a, Wis52a, And86, D’A86, Red00, Hei99]. Mind-Body [Cla72, Hei99, Wat77, Wis52a]. Mind-Brain [Ara11, Pir54]. Mind-Like [Mac53a, May52, Pir52a]. Mind-Machine [LS94]. Minded [Rai20]. Mindful [Bro96]. Mindlike [Mac51]. Mindreading [Gol06, Gan11]. Minds [Avr02, But96, Dac23, Gan11, Gol06, HP03, Hoo84, Pap96b, Puc89, SB93, Sto99, Gib86]. Minerals [Hog86]. Minimal [Bat02, GPW05, Pov18, Ten85]. Minimizers [Van81, VHH86]. Minimum [Sch60, Sch61b, Wil80b]. Minkowski [Mun86a]. Miracles [Hol98, Pos60a]. Miraculous [SV11]. mirage [Kel10]. Mirowski [Ros09]. Mirrors [Buc23]. Mislead [Mar57]. Miss [Mac63a]. Missing [Abo02, Le 05, Wil09b]. Mistakes [Cha73, Mea10]. Mitchell [SF13]. Mixed [Mag10]. Mixtures [MD22]. Mnemic [Naj24]. Modal [Cli95, Cli96, Dic95, Hal01, Lad04, MvF03, Ruy19, Win21a, Win21b, Chi98, Ver99]. Modalism [Man21b]. Modality [Bri97, Dic07, Har22, Lad00, Mul05a]. Model [Ake90b, Ake93, Ang01, Blu89, Bye69, Cab20, For99, Gre53, Gro90, HP05a, HP05b, Irv16, Kie01, KL17, LC15, PC12, Sce91, Sch69a, Sch08b, SW18, Sto91, Wis52a]. Model-Based [Irv16]. Model-Objects [Bye69]. Model-Selection [SW18]. Model-Structures [Bye69]. Modeler [Wei07]. Modeling [Wil14a]. Modelling [CS12, Col15, Del21, Joz86, KL17, KLM10, Mor02, Ste12a, Ste96a, Tav91, TBR18, Rus08]. Models [Ach65, Ack66, Bat02, Bec19, Bub94, CHvI15, Dac23, Eva18, Far67, FS21, Gil01, Hes53a, Hin56, Hor94, Hut53f, Hut54e, Kat13, Kuk92, Lat76, Len06, LC15, McW16, Mor93b, Ngu21, Par16, Pin22, Pov18, Pra24, Rat20, Red80a, RHH18, Ric19, Rue05, Shr19, Spe65, SW13, Sul22, Swa67, Tal15, Wit61a, Zha13, vR13, GB99, Wei13, Rat20]. Moderate [Kus01]. Modern [BF23, Bor54, Din70, Fey80, Low00, Rad00, Ras53, Zah77, Hei58, Hei62, Nad99]. Modes [Pri60]. Modest [Pic11, Tha98]. Modification [Sab54]. Modular [Lyo01]. Modularity [Ben90, Boo24, Car02b, Ech16, Fod01, HW04, Sam98, Sch08c, Sha88, Dav89]. Modus [Hel18, Wag04]. Mohan [Gia06b]. Molecular [BD86b, CK89, DiF23, FS24, Kuy24, Rat20, Šus07]. Molecules [Hud20]. Momentum [Leh11]. Monism [Kle92]. Monothematic [BGH^{+}23]. Monton [BG01, Lad04]. Mood [Siz00]. Moor [Tom99]. Moral [Nad08, SF17]. Morganti [Yud16]. Moritz [Wat77]. Morphological [DiF23]. Morphospecies [Hul70]. Morrison [Von87]. Mosaic [MP08]. Most [Fri09b]. Motion [Bel00, Cha11, Cla82a, Cla82b, Cro53a, Hug04, Saa18, Wat51, Whi53a, Buc51]. Motivated [PNSBC94]. Motivating [Cli95]. Motivation [CB12b]. Motu [Rus54]. Mr [Cra86, Din65a, Ezr57, Gol59, Goo64, Kap57b, Kes66b, Kör68, Mun53, Mun54a, Pop58b, Ram89, Sto70b, Wat59d]. Mr. [Kne57b]. Msrp [GJ88]. Much [Fre88]. Muhammad [Sla15a]. Muller [Vic08].

Multi [Bun58a, Cha57, Cla16, Oka04]. Multi-Dimensional [Cha57, Bun58a]. Multi-Level [Cla16, Oka04]. Multidimensional [Dob58a]. Multilevel [Oka16]. Multiple [Bar13, Bat00, Hol98, Liz93, Wil81a]. Multisensory [Man21b]. Multivariate [RKK19]. Multiverse [Tap11, Vai15, Tap00, Wal12]. Mumford [Bos06, Cha13, McK16, Bir01]. Murder [Ses07]. Murphy [Bro09b, Rus00]. Musing [Aut20]. Must [GP14, How73, Lan06]. Mute [Puc76]. Mutual [BG16a, Leu12]. My [Cou52a, Puc77, Spu52, Ten97, Ein50a, Ein50b]. Mysteries [Roz69]. Mystery [Wil09b]. Myth [HW18]. Myths [Gas75, Gut00, LB13, Moo78].

N [Fie86, Gut00, Saf86, Blu89]. Nadel [Jar61a]. naevo [Luc69a]. Nagel [Lip82, She73]. Naked [Har93]. Names [Rey09]. Nancy [Fra88a, Gib84, KT93]. Narrow [Mil68]. Nativism [Cow98, Far16a]. natura [Gil03]. Natural [Amu99, ARR15, Aut20, Bal92, BEL92, BR04, Bou15, Bur18, Cha02a, Cle11, Far00, Gou58, Gro11, HW18, Kha13a, Kuk94, Mel77, Men54, Mil06, Mos91, NS67, Oka08, Pop68, Rus87, SL13, Sla15a, Sla15b, Smi00a, Sob95, SO03, Ste16a, Ste74, Suc69, VK91, Was78, Zah91, BV05, COS11, Gay98, Sch05, Sla15a]. Naturalism [Gie72a, Hea86, MR94, Pas05, Rat04, Rol07, Ros96, Ros99]. Naturalistic [Dev91, Sie10a, Mad07]. Naturalizing [Sil91]. Nature [Alb62, BEL92, Bon98, Bri51b, Bri51a, Bri51c, Buc72, CBC^{+}22, Dor19, Fey60, Gil00a, Gua07, Kha02, KZ16, Mag65, Mum13, Nia05, Pop52, Red89, Rub86, SL13, San08, Shu19, Urb88, Woo14, Ach83, Bic06, COS11, Dor05, Fos04, Kel10, Kis06, Mum04, Bos06, Gil03, Sch11a, Lev01]. Natures [Lew10]. Neander [Nea95a, Sob95]. Nearest [Dem01, Pot02, Pot00]. Necessarily [Lyo01]. Necessary [Coh82, Dav03, Hau13, Mat03, Max68, Sac70, Bod99]. Necessity [BEL92, CM73, Hun81, Joh89, Kne61, NS67, Pop68, Suc69, Hun06, And10]. Need [CT23, Fer70, Gin93, Gri08b, Mor09, SW18, Dor78]. Negative [BH94]. Neglected [Lun21]. Neither [Sud76]. Neo [Far16a, Mac03, Pea12, Sha03, WHB^{+}23b]. Neo-Gouldian [Pea12, WHB^{+}23b]. Neo-Logicism [Mac03]. Neo-Logicist [Sha03]. Neo-nativism [Far16a]. ness [Bam93]. Net [Geb22, Gly10a]. Netherlands [Gia06a]. Nets [Gly09]. Network [Bec19]. Networks [BSM19, Bot94, Mat17, WOB21]. Neumann [Bec04, Hen03]. Neural [MA23, RKK19]. Neurally [Col15]. Neurath [Wat74]. Neuroconstructivism [Far16a]. Neuroimaging [Kle10, Wri18]. Neurons [Col15]. Neuropsychology [Gly94, GH16, SD93]. Neuroscience [BG16b, Chi18, CS12, Gan11, GS08a, Gil89, Gol06, Irv16, RKK19, Tal13]. Neurotic [Per07, Max04]. Neutrality [Kum69, Lea68]. Neutrino [Pit11]. Neven [Oft07]. Newburgh [WY73]. Newcomb [Ahm05, BHM72, BR76, Car75, Cav10, Sch98]. Newman [Ain09]. Newton [Far00, Abo02, Ang02, Art94, Buc51, Hal82, Hug92, Sab54, Wor00]. Newtonian [Ang05, Dor77, Gie73a, Ker87b, Kno14, Lac70, Lar04, Wal20, Wil07b]. Newtonianism [Elk71]. Newtonus [Wes79a, Wes79b]. Ney [Cal15]. Neyman [Alb02, Gra78, Len06, MS06a, Red74a, Spi73]. Niccolò [Far00]. Niche [HR22, UH19]. Nicholas [Per07, Sma00]. Nicod [Hor78b]. Nightmare [Gro08]. Nineteenth [Cle79, Pot53]. Nineteenth-Century [Pot53]. Ninian [Dob61]. Nissim [Øhr86]. Nissim-Sabat [Øhr86]. No [BH15, Cal98b, Cuf18, DHS15, Dic95, FV11, Hem67, LW15, Pap96b, Sam18, SS13, Sud73, Wor07, DGO07]. No-Go [Cuf18, SS13].

No-when [Cal98b]. Noisy [BGTW24]. Nomic [Hun81]. Nominalism [AB16, MvF03, Mac99]. Non [Bae12, Bha91, Boh62, Bra22, Chi18, Chr77, Coh02, Das58, Ell57, Fra80, Fra87, Fri02b, Goo13, Kui83, Mau00, Mey21, Qui74, Rox75, Sko14, Súr00, Whe12, Wil10, Win21a, Win21b, Wri86, Yos12]. Non-Causal [Sko14, Chi18]. Non-Classical [Boh62, Goo13]. Non-Conventional [Chr77]. Non-convex [Whe12]. Non-deductive [Fra87]. Non-Epistemic [Wri86]. Non-Equivalence [Coh02, Bra22]. Non-Inductive [Kui83]. Non-Instantial [Das58]. Non-Locality [Fri02b, Mau00, Súr00]. Non-mechanistic [Mey21]. Non-perspectival [Win21a, Win21b]. Non-Process [Ell57]. Non-Reductionist [Bae12]. Non-Reductive [Yos12, Wil10]. Non-separability [Bha91]. Non-Standard [Qui74, Fra80]. Nongenetic [Mam04]. Nonlinear [HH93, RS96]. Nonlinearity [BV05, Str09]. Nonlocal [CBR90, Sta90]. Nonlocality [HP03]. Nonlogical [CM73]. Nonpresuppositionist [Los86]. Nonrandom [Bat07]. Nonrelativistic [Thé12]. Nonsense [Bun68a]. Nonseparability [Myr11]. Nonseparable [Lau95]. Nonstandard [Mac93]. Nonstationary [Hoo03]. Nontrivially [Elb93]. Noordhof [Sch07a]. nor [Sud76]. Nordic [Spu52]. Normal [Car85, Lyo14, May96, Til93b]. Normalcy [Cra01]. Normative [GC19, Kar12, MR94]. Normativity [BGTW24, Bet24]. Norms [CK20, PG13, Pet23, Vas22, Bic06, Gua07]. Norton [Fri09a, Lar99]. Not-So-Naked [Har93]. Note [AB10, Alt68, BH53, BM10, Bla85, Car56, Cha57, Coh68, Coh87, Cro53a, Den89b, Din56, Din65a, Fey68, Fin64, Fre88, Gal59, Goo64, Hes58a, Hol88, Hon91, Hor94, Hor78a, Hut55a, Jon56, Jon69a, KT93, Kap61, Kei93, Lin93, Mac71c, Mac86, May61, McC23, Mil05, Mil83a, Mil85a, Poo55, Pop53, Pop57b, Pop58a, Pop66b, Pop76, Pot51, Pro68, Rus54, Sab54, San61, Sch65a, Sch61b, Sma53, Tur91, Vas21b, Why55a, Wig56, Wig61, Wat99]. Notes [Bin54, Bla78, Har59b, Hol94, Leb89, Red78, Ros78, Sam78]. Nothingnes [Cra01]. Notice [But02, KK65]. Notion [Cle70a, CKKR88, Das16, MS55, HSRS99, Šus07]. Notions [RS07]. Novel [CV83, Gar82, Goo85, How84, KLS92, McC86, Nii83, Tho92, Lad99]. Novelty [Nun93, Red86, SW18]. Nuanced [SW18]. Nuclei [Nyr21]. Null [MLR21]. Number [Ber23]. Number-Theoretic [Ber23]. Numbers [Gru66, Kne72, Mic94]. Numerals [Kne72]. Nurture [Kha02, Kel10].

O [Dob58b]. O.M [Ecc52]. Oberdan [Ueb99]. Obeyed [Sud73]. Obituary [A.56a, C.56c, C.56d, Ecc52, Har91, Hut53b, MS50, Wor99]. Object [Bal57]. Object-Relations [Bal57]. Objection [Ain09, CC93]. Objections [Bat71, How85, Nyh88]. Objective [Ale18, Cus22, McC57, Whe12, Wil13b, Pau16]. Objectivist [Cha79c]. Objectivity [DG10, Hoo91, Kos21, Mar71, Mar72, Mos88a, Mul91, Sie80, dC22, Til84, Roc86, Jar12]. Objects [Bye69, Gła23, Gor70, Low00, Pro59, Wol12, GB99]. Observability [Cha03, Kos88, Mul05a]. Observable [BB04]. Observables [Fei15, Fei18, Har94, Tum07]. Observation [Azz04, Bro93, FAB^{+}89, Gil92b, Hes58a, Hes58c, Mos88a, Sch64b, Spe66a, Spe66b, Zah82a, Zah82b]. Observations [Cha85a, Why53a, Why54b]. Obsession [Mcg70]. Occam [Zha13]. occasion [AGTZ08]. Occasions [Sid01]. Ockam [Kap58]. Odd [BFHS02]. Oddie [BH91, Nii82, Pea83]. Odds [Ang05]. Off [SS13]. Official [Blo70]. O’Hear [Gla84, Gow93]. Oil [Nia05]. Okasha [Bir13]. Old [How91, Irv12, Mon06, Roz69, Bar99, Pro86]. Olsson [BFHS02]. O’Malley [McC16]. Omit [Bra21]. omni [Luc69a]. Omniscience [Bet16]. One [BEL92, Han05, Hau11, Kuk08, Max93a, NS84, NS87, Nor86, Øhr86, Pap96b, She85, Til94, Why55b, Grü76b, Grü86]. One-Way [NS87, Øhr86, Why55b]. Only [Lai88, Leh11, Lev69]. Ontic [Gli20, LW15, McK14, Pov18, Thé16]. Ontogeny [San08]. Ontological [Kuk94, Le 05, Mar72]. Ontology [AGTZ14, DEO19, Dis94, EHLD14, Fol00, Pri55, Roh88]. Open [Rai20]. Open-Minded [Rai20]. operation [OB12]. Operational [Bra55a, Hes52, Sch59b]. Operations [Grü69b]. Operator [Sla93]. Operators [LaC22]. Opinion [DE76]. Opinions [Sor84]. Opportunity [Lav23]. Optics [Abo02, Wor00]. Optimal [NW22]. Optimality [Juh94, Pot10]. Option [Mey21]. Orden [Ros10]. Order [BY06, Fis06, Gil00a, Hoo93, Ket01, Kle53]. Ordering [BH87]. Orders [Fis79]. Ordinary [Gła23, GB99]. Organic [Goo10]. Organisation [Gia06a, Kle53]. Organism [Why55a]. Organisms [LC15, Ram13, Why57]. Organization [MP22, SMM11, Car04b]. Organizational [MSM09]. Organs [Mun94]. Orgin [Van94a]. Orientation [Rus00]. oriented [Max93c, Max93b]. Origin [Cra92, ES59, Kap55c, Kes65a, Kes65b, Kes66a, Kes66b, Kes66c, Kra09, Lev65, Poo55, Sac89, SC18, Mas05, Dup13, Din65a, Sob11]. Originating [Bel05a]. Origins [Con03, Gro11, Nem03, Tuc11, Sch05]. O’Rourke [SL13]. Orthodox [Rai20]. Orthodoxy [Red74b, Red74c]. Orthogonality [Lew97]. Orwellian [Tod09]. Oscillation [Kno10]. Oscillations [MA23]. Oscillatory [Ake90b]. Osherson [Sch00a]. Ostrogradski [Swa22]. Other [Hul74, Mat03, PR95, Saf86, SB93, Was55, Gal00b]. Otherwise [Wel04]. Otto [Wat74]. Our [Avr02, Bri51b, Bri51a, Bri51c, Cha73, Egg16, Enf08, Poo55, Dem03, Sta06b]. Outcomes [KP99, Pla00]. Outline [Von50, Wis53b]. Outlook [Din51]. Outplay [Ash52]. Over- [Tuc11]. overdetermination [Wit00]. Overdose [Hau11]. Overfitting [HS04]. Overview [Gre81]. Owen [Rai86]. Own [Avr02, Nol99]. Oxford [Bri06, CE04, Gia06b, Gly04, Rub86, She06].

P [Dre00, Gal00a, Gil13, Gra01, Hea86, Hea99, Hoo86, May61, Pas16, Smi00a]. P. [Enf08, Har59b, Sch59b]. Paap [Ros10]. Packets [Liu94]. Page [Sma86]. Pagnini [Cam00]. Pagonis [Dic01a]. Paneth [Bla63a, Bra63a, Bra63b, Pan63]. Pangenesis [Sta06a]. Paolo [Ueb01]. Paper [Aga59b, Den94, Fou57, Pod75, Pod76, Sch65a, Sut57, VM16, Mil78]. Paperback [Jen86, Lyo06a, She06, Roc86]. Papineau [Ehr87, Urb94, Pap88]. Para [Sor03]. Para-reflections [Sor03]. Paracetamol [Hau11, Kin10]. Paradigm [HPE92]. Paradigms [Fra84a]. Paradise [Ros94]. Paradox [Bha86, Cha10, Dew73, DW06, Fos69, Gar62, Gib69, Gie70, Goo60a, Goo61c, Har11, Hol88, HO79, Law85, Mac63b, Mac66, MG65, Mik04, Mil66a, Mil66b, Pop62, Pop66a, Pop66c, Sch64b, Smi10, Vra04, Wei01b, Wil64b, Bar08a, BER99, Alb02, Alt68, BHM72, BR68, Car69, Car75, Goo70, Gru66, GR15, Hol94, Kne71, Red74a, Roz69, Sch98, Van94b, Wat87, BR76, Kut00]. Paradoxes [Ale58a, Ale59a, AK91, AK95, Arn90, Bau64, Bau68, Dou02, Kah69, Pin11a, Sti85, Van08b, Van08a, Van10, Kah67]. Parallel [Lyn90, Wel93]. Parallelism [Pea12]. Parallelogram [Sau64]. Parallels [JHB^{+}23]. Parameter [Jon82a].

Paraparticles [CB12b]. Parapsychology [Was55]. Parastatistics [BHS15]. Paresis [Die67, Car99a]. Paribus [Dre01, Kle92, PR95, Reu09, Sch01b, Rup07, Sch07c]. Parochial [Fei18]. Parrini [Ueb01]. Parrot [Fre88]. Parsimonious [GS08a]. Parsimony [Bak03, JT17, Nol97, Sob81]. Parsons [Mos85]. Part [Max93a, Max93b, Dob51a, Dob51b, Fra50a, Fra50b, HP05a, HP05b, Max93c, Pop50b, Sch52a, Sch52b]. Partha [Dic01b]. Partial [Ach63, Mac94b, Mor60]. Participant [Hac88]. Particle [CH01, DEO19, Din67, Dor77, EN98, Lar97, Max88, McK14, Red77, RT92, Sch00c, Squ89, Wil23b]. Particles [ET97, HY02, RT92, Why51a]. Particular [Sko14]. Parts [Bab22, Hug08, Kes66b, Win21a, Win21b]. Passage [Rus54]. Passing [Tod09]. Passionate [Kus01]. Passivity [BZ84]. Past [Ahm14, Bel79a, Cal04, Cra79, Gal00b, Grü69a, Pop78, Rak97, Wei72, Wei79, Whi78b]. Pasteur [RH72]. Path [Wei19]. Path-Specific [Wei19]. Pathological [Til93b]. Pathways [Ros21]. Patient [Kuy24]. Pattern [Puc74, RKK19, Tho75]. Patterns [Fey77, Van94a, Win21a, Win21b]. Patterson [Fra99]. Paul [Lar00, Sch07a]. Pauli [Bro86, Hen81, Hen84, Kra09, Mas05]. Pavlov [Ash57a]. PC [CB23]. Pearce [Ber78]. Pearl [FG17, Gil01]. Pearson [Bai83, Gra78, Hac81, Len06, MS06a, Mor02, PP78, Spi73]. Peculiar [Hor78b]. Peer [Row23]. Peirce [Hoo85, Hoo99, Hoo00, Mil75, Set77a, Put01]. Peircean [Mee04]. Penelope [Ros99, Sie10a]. Penrose [Red00]. People [WOB21]. Perceive [Gau66, Jon69a]. Perception [BZ84, Ech16, Gil89, Gre17, Lom02, Mat05, Pir53, Smo21, Smy56, Smy60b, Tuc58, Tuo78, Gia06b]. Perceptual [Gri08a, Pri04]. Percival [BS00]. Perfect [Elb93, Hea09]. Performance [Bub94]. Performed [GP14, Grü69b]. Period [Nii98].

Permanent [Pit11]. Permeate [PLKG24]. permutation [Hug99]. Perrin [Cha11, Hud20]. Persistent [DEO19, Har06, NS21]. Persists [DB15]. Person [Tho94b]. Personal [Liz93, Puc73, Wil81a]. Personality [Liz93, Wil81a]. Persons [Rob76, Woo60b]. Perspectival [Rue05, Win21a, Win21b]. Perspectivalism [Mol22a]. Perspective [And98, Bla09, Cuf18, Jam23, Nee13, Pir52b, Irv13, Van08a, Van10, Pin11a, Koz14]. Perspectives [Gal00b]. Perspectivism [KZH^{+}22]. Perturbative [Fra20]. Pessimistic [FA14, Hob94]. Peter [Den86, Fod01, Jar12, Kir01, Rub86, She01, Tha09, FB01, Fri16, Lyo06a]. Petrinovich [Sin00]. Petroleum [Hog86]. Phenomena [Bok08a, BF12a, BF12b, Dor71, KK17, Mas07, Nel94, Pro68, Sko15, Smy53, Wor00]. Phenomenal [Cra69a]. Phenomenology [Gil80]. Phil [Sch07a, Sco08]. Philip [BG16b, Dou13, Ros09, Smi00b, Tei13]. Philipp [Pot51]. Philosopher [Bas92, Bla09, Bra70, Cur76, Gil81, Hal01, Hen81, McL07, Put05, Ver99, KW97]. Philosophers [PLKG24, Puc78, Wil75b]. Philosophical [Bec15, Bel01, Bon01, Cha11, CB69, Coh86a, Coo05, Dav62, Dup13, Gal00b, Gea65, Gia06b, Gie72a, Gil13, Gla14, GBW61b, Hal15, Har62, HHS02, Hug00, Kie97, Kie01, KFH65, KOV06, Kör67, Lai84, Mat05, Mcg70, Mos86, Nor04, Nyh88, Oka02, Oka08, Pop52, Rat04, Sim01, Tso10, CC13, Coo02, Dem13, Hoo84, Row11, Sob11, Wit58c, Smy58b, DiS99]. Philosophies [Dem01, Koe00, Pot02, Ten81, Pot00]. Philosophising [Fin75]. Philosophy [A.56b, Ale16, Bac95, Bol83, BG16b, Bra99, Bro13, Bro58, Bub89, Buc65, Buc72, Byr07, Cal50, Cla00b, Cla14, Con03, Den59a, Den60a, DiS99, Ern90, Fag13, Far00, Fey60, Gan11, Gar73a, Gar99, Gar07, Gat01, Gia83, Gib84, Gie75, Gie73c, Gol06, Goo84, Gra01, Hac79, HH63, HH64, Hea54, Hes69a, HH92, Kan88, Ket00, Kir01, Kli87, Koz14, Kre60b, Kro02, Kuh80, Lak76, LRP23, Los86, Mac83, MS06a, May60, Mos85, Mur90, Oli00, Pas05, Pin16, PLKG24, Pri91b, Pro86, PS00, Rad87, Rad91b, Rat02, Sch11a, She01, Sie10a, Skl13, Smi00a, Ste74, Ste95, SD93, Tor84, Ueb96, Whi53a, Whu56, Why55a, Wig61, Wis62, dR96a, But84b, Dor05, Dup14, Fie86, Gel59, Gil03, Hei58, Hei62, Irv13, Mad07, Bat15, CS16, Cur16, Hoe00, Jen86]. Philosophy [Joh15, McC16, Pre01, Ste10, Dic01a, Mum13, Ueb01, Mor99, Nad99, Tom99, Boh62, Hea53, Hof01, Kre60a, Lip99]. phoenix [Tom99]. Photons [Dor71]. Photosynthesis [Dob57, Kap57b]. Phrased [BH87]. Phylogenetics [Vas21a]. Physical [Abr58, AB10, And10, Ang01, Bec04, Ber23, Bri51b, Bri51a, Bri51c, BF12a, BF12b, Cot03, Ell57, Fer70, Fle16, Gre53, Grü69a, Hei99, Hes52, Hoo91, Jam23, Joz86, Kap58, Kro89, LT17, Lup18, Mar60, Mil19, Mun86a, Ney09, Pic11, Roh88, Sch01a, Sko15, Tho88, Tur55, Vic11, WB53, WB54, Whi55e, Why53a, Why54b, Why57, Why63, Zah80, Grü00, Hun06]. Physicalism [Wil10, Yos12]. Physicalist [Sma79]. Physicalization [Mil94a]. Physics [Bla52, Bla53, Cas58, CT16, Dav03, Dea61, DiS99, Dic01a, Din60a, Din60b, Din70, Dob51a, Dob51b, Dob53, Dob56, Dob69, Eas14, ELLH17, FR88, Fri09b, Gib84, Gil00a, Hei58, Hei62, Hes53a, Hes62, Hoe69, Hug04, Hut53f, Hut54b, Hut54e, Hut56a, Hut60d, Kar12, Kra78, Kro02, Lav77, LM85, Lin93, Mac60a, Max66, McK14, Mol23, Mor89, Nor09, Pik54, Pik55, Pop50a, Pop50b, Pri91b, Pro86, Red80a, Sau64, Sch61c, Sch62, Sch22a, Sch00c, Sha82, Sto86, Swa22, Tor84, Vic11, Wal20, Why52a, Why52b, Why54a, Why55b, Wil09b, Woo60a, Yud16, Kos00, Mau07, Mor13, Bla89, Ney11, Col00, Far16b, Low00, Boh62, Sac93]. Physiological [Pir53]. Physiology [Pir50, Tha70]. Picture [Bae16, SW18].

Pictures [Bro97, Fol99, Hof01]. Piecemeal [Irz85, MW14]. Pied [Pir52c]. Pieter [Hal01]. Pietroski [Sch01b]. Pincock [Bak15]. Pink [Goo68b]. Pipers [Pir52c]. Place [Gil00a, Gou57, Hut53a, Mac93, Wei04, Cra90, Bir01]. Plain [How72]. Planck [Bas92, Bla78]. Planetary [Cha81, Cha83, Hut83, PS86]. Plasticity [Kov19, Van93a]. Platonism [Die99, Die99]. Plausible [Lar04]. Player [Ash52, Hal52a, Hic53]. Playing [Puc74]. Please [FW21]. Pleasure [Puc69]. Plenum [Pro86]. Pliny [Pap05]. Pluralism [Bar19b, Bar98, KDM21, Lee21, Ste16a, Tay18, War22, Bok08b, Ber10]. Pluralist [Gar18]. Pluralistic [Roh88]. Pluralists [Bar05a]. Plurals [Hos00]. Podolsky [Bha86]. Poidevin [Bri06, Phi09, Tei99, HN07]. Poincaré [BM01]. Point [ABB^{+}24, Bai83, Gou23, Set69, Tor87, Van61a]. Pointillisme [But06b]. Points [Bar72b]. Poison [Kin10]. Polanyi [Den86, Kan84]. Polarization [NS21]. Polemic [Mar62]. Policy [Boe22, Mit09, CS16, SF13]. Policymakers [WOB21]. Polio [Kin10]. Politics [Gat01, Ket00]. Polythetic [Fel23]. Polywater [Mck91, Van93b]. Popper [Gat01, Bam93, BH55, BH56, Car56, Cra79, DH86, Gar72a, Gil90, Gil13, Gla84, Har74, Har60, How87a, Irz85, Kes66c, Kri84, Leb89, Llo67, Mac78c, May96, Mil74c, Mil75, Mil85a, NS67, Nol87, Peć89, Pos62a, Row11, Row13, Sch74b, Set77a, Set89, Set90, Suc69, Swa88, Tic74, Tur91, Urb78, Urb82, Van87, Wat94, Zah83a, Zen77, Ket00]. Popperian [Bea89, Bra76, Gem89, Mil90a, Sam80, Sti79, Ten85]. Popperians [Sti81]. Popperism [Fre78, Har72]. Population [Akk94, Bir20, Chi57, Gil14, GS08b, Mil06, Ram13, Ste16a, Was78]. Population-Level [Mil06]. Populations [Cla53, Mor02]. Portability [Spr07]. Posidonius [Pap05]. Position [Jef55, Poo55, Why51a, You86].

Positive [BH94, Mic81, Set69, Ueb01]. Positivism [Har59b, Zah81]. Positron [Han61, Han62a, Hut62]. Possession [Bar12]. Possibilities [Bro96, Hal01, Par16, Ver99]. Possibility [AB10, Bae19, Ber68, Eme15, Pop78, Wel04, Chi98, Haw99, Mel00]. Possible [But84a, CBR90, Kit87, Nia05, Pit94, Pos74, Ros01, Sta90, Sti85, AHW95, Rue11, HM99, Ste00]. posteriori [Ros19]. Postmodern [Mur90]. Postmodernism [Koe00]. Postmodernist [Gut00]. Postulate [Dic95, Wyk76]. Postulated [Sob98]. Potter [Dem01]. Poverty [Grü04, LM01]. Power [Bak03, BH91, Coh16, ES19, Odd90, RS96, Tho94a]. Powers [Har70, Hid05, Low04, Win21a, Win21b, MA11, Cha13]. Pp [Bai86, Cap86, Coh86a, Den86, Fie86, Gib86, Kör86, O’H86, Pro86, Rai86, Bro86, Bun86, Hea86, Hog86, Hoo86, Jen86, Roc86, Rub86]. pp. [And86, Mos86, Saf86]. Practical [Ake91, Cuf18]. Practice [Bam93, Cur23, Gre11, Kle02, Kre60a, Kre60b, Kuk94, Miz23, Til92, Fol99, Sto99]. Practices [Rat04]. Pradeu [O’M14]. Pragmatic [DF89, How75a]. Pragmatics [Sky84, Bai86]. Pragmatism [Put01, Hoo00]. Pragmatist [Haa76a, Hea12]. Pre [JHB^{+}23]. Pre-QM [JHB^{+}23]. Precautionary [CS16]. Precisely [Wil23b]. Precision [O’C17]. Precursor [Pop53, Urb82]. Predicates [Cle70a, Kum67]. Predication [Mat71]. Predictability [Bic87, O’C57, Wat71a]. Predicting [Gar82]. Prediction [BS92, Cle11, Den94, Die67, Gar62, Har06, HS04, Hor78a, Lan01, Law85, Mar89b, Mat71, MG65, Pop62, Pur23, Res58, Sal81, Sch57, SGS97]. Predictions [AGTZ14, FS94, Grü56, Har08, Kru97, Par21, Wil94]. Predictive [CEH21, Dor19, Gła23]. Predictivism [Bar05a, Bar08a, Har11]. Predilections [Har08]. Preemption [Ehr84]. Prefatory [Jon56]. Preference [Liu22]. Preferred [Bar05b]. Preferred-Basis [Bar05b]. Prehistory [Bet52b, How78]. Prejudice [Sch61a]. Preliminary [Wit58c]. Premises [Le 05]. Present [Dob69, Gal00b, Grü69a, Jef55, Psi00, Rak97, Sac93, Why51a]. Presentism [BJ03]. Press [And86, Bai86, Bri06, Coh86a, CE04, Fie86, Gia06b, Gib86, Hoo86, Kör86, Mos86, Per07, Pro86, Rai86, Roc86, Rub86, Saf86, Tau06, BD86a]. Presume [NS20]. Presuppositions [Cal50]. Pretence [Car02a]. Pretense [Tal13]. Prevent [Dri11]. Price [EN87, Mag04, Per07, Lee94, Lee95, Nor02]. Prices [Sob01]. Primary [Pri89]. Primate [GG86]. Primer [May00]. Primitive [AGTZ14, ELLH17]. Principal [Bla98, HLWW17, Mea10, NR13, Pet20, TH20]. Principia [Far00]. Principle [Bar13, Bla98, Cas98, Cha70a, CS16, Cra88, Fis67, Gan15, Gil91a, Goo67b, HLWW17, Hoo03, Hut51d, Kno14, Kra09, Mea10, New73, NR13, Nor09, Pet20, Rad91a, SB93, Sch60, Sch61b, Smi17, Sob81, Sob01, Ste03a, Str95, Suz85, TH20, Wat52b, Wil80b, Wil94, WY73, BG99, HSRS99, Mas05, Bla78, Lan95, Mas01, Par11, Sac72, Sac75, WY72, WY75, Kra09]. Principled [Sch03]. Principles [Aut20, BR04, CS98, Gin93, LU09, Sac76, Wac87, How97b]. Prinz [Gri08a]. Prior [Bro20, How87a]. Priori [Ach95, Mar64, Sto64]. Priority [McK14]. Priors [Jon82a]. Prisoner [Mar90]. Pritchard [Tri94]. Private [Tre62]. Privileged [You86]. Probabilism [Háj08]. Probabilist [For95a]. Probabilistic [Ehr84, FG17, Fre88, Gly11b, Pop60, Pri91a, Riv94, Sho03, Sla93, Sud76, Gal03]. Probabilities [Bar98, BHW18, Ede23, Hoo93, How87a, Hub05, Kyb74, Lan06, Liu22, Moo83, Pop55b, Ses07, Hut99, Mee99, Mil01]. Probability [Ach90, Bai69, Bai71, BS94, Ben56, Ben57, BR04, Coo65, CCT13, Daw85, De 72, DB15, Edw79, Eel88, Elb94, Ell73, Fin73, Gil71, Gil72b, Gil91b, Gly15, Goo68a, Gow91, Gri88, Hac66, Hac71a, Hei78, HP03, How72, How73, How75c, HF94, How95, How97a, Hut58b, Jef55, Kne57b, Kyb58, LM85, Leb89, Lev67, Lev79, Mül05b, Nor07, Oka02, Ols02b, Per18, Pop59c, Pri76, Ram89, Red85a, Ree88, SC18, Set70, Set72, Set77a, Smi10, Van81, Von73, Whi72, Wil78b, Wil13b, vFH17, BH56, Her07, PK70a, PK70b, Str03, Tap00, Wil99, Wer10]. Probabilized [Wag04]. Probable [Jon73]. Probably [JT17]. Problem [Ake93, Bae16, Bar85a, Bar05b, Bla20, Bro09a, BR20, Cav10, Cha07b, Cla72, DeV97, DB15, Dor78, DGO07, Egg16, Ell88, Enf08, Fey68, FW07, For95a, FH16, FS24, Fra20, Gir76, GBW61a, GBW61b, Haa91, Hei99, Hor96, Hor78a, How91, Kie97, Kie01, Kuk95, Lai91, Lig14, Lip02, LS94, MW15, Max93a, May77b, May77a, Mee22, Mon06, Mor88, MR94, Ner71, Nor02, Nun93, Ott94, Pir50, Pir54, Pol57, Pos75a, Res66, Rob94, Rus71, SB93, Sch00c, She85, Sob85, Sta06a, Suá04, Thé12, Tur90, Van81, VHH86, Vas21b, Vol91, Wat77, Wat91, Wat51, Whe12, Wil50, Wil79b, Wit61a, Bar99, For99, Mee99, Sta06b, Wit00, Sim01]. Problematic [Pet23]. Problems [Ahm05, Chi87, CB69, Ell22, Giu88, Irv12, Jar64b, Koe92, Kul77, Leu12, O’N89, Pop52, Smy60b, Sor84, Tho54a, Web09, Wet94, Zah83a]. Procedural [Naj24]. Procedure [KGB^{+}22]. Process [Dow93, Ell57, Mil06, Nee13]. Processes [Dup14, Ham23, Mac62, Sac64, Why55b, Why57, Wil03, Cla14]. Processing [CEH21, Gła23, Lyn90]. Produce [PLKG24]. Productive [Fre11]. Prof. [Cra01]. Professor [BH55, BH56, Bla63a, Bon55b, Bra63a, Bra63b, Bra62b, Coh74, Din60c, Din61b, Dob58b, Dob70, Fey60, Fey76, Foz63, Grü60, Grü61, Hal55b, Har60, Jar65, Kap55b, Kes61, Kes66b, Kes66c, Mil66a, Pop56b, Sch65a, Set69, Van61a, W56, Wat62, Woo56]. Program [Lew05, Spr07]. Programmability [Van94a]. Programme [Bea89, Kle85, Pro74, Tho92, Zah73a, Zah73b]. Programmed [Gun68]. Programmes [Ake88, Hal79c, Hal79d, Sch74a, Yux90]. Programs [Bae12]. Progress [BF90, Coh73, Han92, Tiz76, Urb74a, Urb74b, AHW95, Cha04, HM99, Gil09b]. Project [Los86, Ste20]. Projectibility [DiF19]. Projection [Dic95, Hoo91, Smy54, Tel69]. Prolegomenon [Sha03]. Promiscuous [Dup96, Wil96b]. Proof [D’A20, Gan15, Gil91a, Lon53, Pas15]. Proofs [Bro97, Ham23, Hof01, How97b, Lak63a, Lak63b, Lak63c, Lak64, Fol99]. Propaganda [WOB21]. Propensities [Kri84, Kyb74, Mil85a, NV22, Suá04, Suá18, Urb80]. Propensiton [Max88]. Propensity [Edw79, Gil00b, Mil75, PR13, Pop59c, Set72, Whi72]. Proper [dF81, Gou23, Gri93, Rey09]. Properties [Cli96, Coh02, Cox75, HS17, Woo51c]. Property [Bro19]. Proportionality [GE23]. Proposal [Tha98]. Proposed [FG17]. Propositional [BH87]. Propositions [Pos74]. Protective [Gao22]. Protein [MG17, Tah21]. Protocols [Ueb99, Obe99]. Provability [LR07, Raa05]. provenance [Fie86]. Provide [FS94, LP18]. Providing [Smy56]. Proximity [Har13]. Pruning [Nea95b]. Psychiatry [Fel23, Gou23, Mur06, Bro09b]. Psycho [Ezr56, Fai56, Hut55d, Spi57, Wis56a, Wis63a]. Psycho-Analysis [Hut55d]. Psycho-Analytic [Ezr56, Spi57, Wis56a, Wis63a]. Psycho-Analytical [Fai56]. Psychoanalysis [Erw84, Kli87, Tei62]. Psychological [Buc16, HS22, Hoc17, Kle92, Mad57, MD22, Sha84, Smi59, Tre62].

Psychology [Bla55, Bla52, Bla53, Bon01, Dob51a, Dob51b, Dob53, Dob56, Ern90, Fre10, Gan11, Gea65, Gol06, Gre53, Hoc17, Hoo86, Hut56a, Hut65a, Kap65, KFH65, Kli87, Lou62, Ras53, Sam98, Sha17, Sma79, Tay18, Zah83b, Ric07, Rup07, Sti83, Fre10, Gar99, Kir01]. Psychometrics [Ves20]. Psychoneural [Bic98, Bon00]. Psychosis [Gre79]. Psychosomatic [Sza56, Wal56a, Wal56c, Wig56]. Psychotherapy [Bin90]. Ptolemy [Tau06, Bar00]. Public [IK19]. Publications [Ano86b, D’A86]. Publishers [Fes06, Gia06a]. Publishing [Bro86]. Puccetti [Tho75, Wat62]. Pump [Jen88]. Pure [Bar11, Ste00]. Purpose [AA87, Why93]. Pursuit [Nyr21]. Put [BN94]. Putnam [Fey87, Lev97b, Urb77, Wil75a]. Putting [Gar23]. Puzzle [Pit94, Ram23, Rob08b, Fri00]. Pythagoras [Kle02].

QM [JHB^{+}23]. Qua [Goo68b]. Qualitative [Mil74c]. Qualities [Pri89]. Quality [MP93]. Quanta [CH01, HN14, Lan56, Pos73]. Quantification [Par82, Sim85]. Quantifying [GE23]. Quantitative [Bak03, GL23, JT17, Mic94, Nol97, Bar99]. Quantization [Sac70, Thé16]. Quantized [Wal06a]. Quantum [Alb94, Alb62, AGTZ11, AGTZ14, Bak09b, Bar01, Bar05b, Bel86, Ber10, Bha91, Boh62, Bok08a, Bor53, Bro86, Bro96, Bro09a, Bub79, Bub89, Bub00, Bun68a, Cal15, Cal23, Cas58, Cav10, CBC^{+}22, Cli95, Cli96, CH01, CT16, CL19, Dav05, DEO19, Dic01b, DB15, Elb93, Eva17, Eva18, Fei15, Fei22, FGK15, Fin73, Fra20, FR88, Gal81, Gar72a, Gar72b, Gri07, Hal99, Hal01, Har22, Hea12, Hea15, HP03, HY02, Hol86, Hug00, Jab92, KZH^{+}22, Kri84, Lan59, Lan65, Lan76, Lew97, Lew06, Lew07, Loc96a, Loc96b, Loe98, Low00, Lup18, Mag04, Mau00, Max88, Mee22, Mil18, MS07, Mun86b, Myr03, Myr11, Nar74, Pit94, Pop50a, Pop50b, Put05, Red74b, Red74c, RT92, Sce91, Sch74b, Sch52a, Sch52b, Sch53, Sch54b, SC18, Shr19, Sta85, ST22, Str96]. Quantum [Suá04, Súr00, Tel86, Tum07, Vai15, Ver05, Von73, WT10, Wal14, Wei01a, Whi53b, Why58c, Wil13b, Wil17, Bok08b, Gho06, Hen84, NA13, Rue11, She99, Sty00, Ver99, Wal12, Kut00, BS00]. Quantum-Classical [Ber10, Bok08b]. Quantum-Theoretical [Gar72a]. Quasi [Boe22]. Quasi-experiments [Boe22]. Quentin [Cra93]. Querist [Hut53a]. Quest [Nor86]. Question [Ful19, Lad84]. Questions [Coh86b, Grü76b, Grü86, Mat03, Mor93b, Pro86, San98, Smo77, Tei99]. Quine [FW07, Gar73a, Gie72a, Tan89]. Quitting [Wil14a].

R [BH55, Bra99, Dob58b, Fre11, HM99, Rad00, Tei99, W56, Woo00b]. R. [Joh87, Kyb58, Mel90]. Rôle [Hut53f, Hut54e]. Rabbits [May96]. Race [And98, Gan04, Pos60a]. Rachel [Gla14, Tso10]. radiation [Fri00]. Radical [LW15]. Radu [Sto99]. Raison [Bla66]. Ramifications [Hes69b, Hun69a]. Ramsey [Bon01, Bot94, SG73, Spr11, Wat89]. Ramseyfication [MS06b]. Ramsification [Ket04]. Random [Joh89]. Randomization [Pap94b]. Randomize [MTH^{+}24, Wor07]. Randomized [Ful19, Fal00]. Randomness [Eag05]. Range [SB93, Sch90]. Rani [Cha13]. Rate [Bir22]. Rates [GP14]. Ratio [CCT13]. Rational [Bin09, Bra17, Dem03, Dou02, Kul77, Sal81, Sau64, Sim65, Wei93, Sky12]. Rationalism [Gil13, Gro90, Sti79, Row11, Mel99a]. Rationalist [Ste01, Tho94a]. Rationality [BGTW24, Bic87, Bir13, BGH^{+}23, Cur86, Dor72, Gro90, Jen84, Put01, Row93, Sie80, Szu77, Tia85, Tro93, Wac87, Wat91, Wet94, Hoo00, OB12]. Raven [Vra04, Wat87].

Ravens [Gib69, HS68, Sch64b, Sim91, Sto70b]. ray [MG17]. Rays [Gas75]. Razor [Goo68a, Kap58]. Razors [Zha13]. Reaching [Cla07]. Reaction [Puc76]. Reactions [Gri08a, Pri04]. Read [Bou15, Pla16, She13]. Reading [Far00]. Real [Cal23, CH01, Fra20, Gia94, Psi04, Rey15, Tho88, Sla13]. Realism [Aro89, AHW95, Azz04, Bla79, Bro90, Cha03, Cho11, Cla01, Cum92, Dal89, Dup96, Eva17, FA14, Gar72a, Gli20, Har13, Hel87, Jab92, Kav23, Kri88, Kuk94, LW15, Lau95, Lew05, Lyo06b, Max93a, Max93c, Max93b, McK14, Mck91, Mor90, Mur90, Pla62, Psi00, RK87, Rob11, Ruy19, Seg08, She85, Sta09, SF17, Tuo78, Van93b, Vic21, WT10, Wat65a, Wil19, Wil96b, Cha07a, Chi98, Lad99, HM99, Pag00]. Realist [CEH21, Swa21]. Realistic [Elb93, Mel99a]. Realities [Mar59a, Mar59b]. Reality [Cam00, Cha03, Gao22, HW18, Hud20, Liu94, Pol67, Pro77a, Red89, Ueb01, Wei72, Len10, Pri99, Wei14]. Realizability [Bar13, Bat00]. Realization [Rup07, Tah21]. Realized [Sch98]. Really [Bro14, Dic95, Han04, Hut93, Lad00, Mat03, Sac70, Swa22, Wri83]. Reappraisal [GM18]. Reappraising [Van08b]. Reason [Din64a, Leh70, Lev01, MTH^{+}24, Pot00, Pot02, Sco08, Vas21b, Wil03, Dow05, Ste00, Dem01]. Reasonable [Job80, Sch84b]. Reasoning [Far16b, Fri97a, Fri09b, Gil01, Hub05, Lom02, McG03, Sch18]. Reassessing [Boe22]. Recasting [May96]. Receive [Lan06]. Recipe [Bou15]. Reciprocal [Ste96a]. Reckoning [Mag05]. Reclaiming [Rat04]. Recognition [Puc74, Tho75]. Reconceived [Stu16]. Reconceptualization [RS23]. Reconcile [Sab54]. Reconciliation [Bub68]. reconciling [WC99]. Reconditioned [Ash57a]. Reconsideration [Col07]. Reconsidered [Dor77, Hel18, SW19a, Ste16b, Wil64a, Kan84, Ram07b, Den86, Spr11, Hei78]. Reconsidering [Cro08, Cuf18]. Reconstruction [Dem03, Fey77, Gri07, Kul77]. Record [Mag10]. Recursive [SS93]. Recycling [Bar19a]. Red [Fre88, Goo67c, Goo69, Hem67, Mag10]. Redefined [Tic76]. Redeflated [Mil76b]. Redhead [Raa05]. Reduce [Rob23]. Reduced [Ero12, Irz96, HK08]. Reduction [BD86b, Bon00, Dow93, Ero12, Fri82, HN07, Hoo04, Kra78, Le 05, Man97, Mar05, Mor09, Roh88, Ros19, Sce91, Sch61a, Skl67, Thé12, Yos81, vR13, BG99, Bic98, GB99, HK08]. Reductionism [Ber10, Sch69a, Wri68, vEDS06, Bok08b, Oka00]. Reductionist [Bae12, RH72, SE23]. Reductive [Fei22, HL11, Yos12, Wil10]. Redundant [McD95]. Reexamining [Bok08b, Ber10]. Refer [Mac62]. Referees [Cla00b]. Reference [BM05, Cum92, Dav56, Lev97b, Sch09, Sto86]. Referential [Mar71]. Referring [AB16]. Refined [Rad91b]. Reflections [Bub79, Sor03]. Reformulating [Hun23]. Refutability [Wis63b]. Refutation [Arc67, Gem89, Spi73]. Refutations [Grü76c, Lak63a, Lak63b, Lak63c, Lak64]. Refuted [Zuc88]. Regained [Bel79a, Ros94]. Regarding [Mea10, Wea18]. Registration [War22]. Regress [LE22]. Regression [Urb92]. Regularist [Kri18]. Regularities [Ros12]. Regularity [BF23, HH93, Nie21, Ste20]. Regulation [San08]. Regulatory [HPE92]. Rehabilitated [PB02]. Reichenbach [Gar72b, HSRS99, Hut53b, Job80, Juh94]. Reidel [Bro86, Cap86]. Reification [Gan04]. Reinterpreting [Pri94]. Rejecting [Dac23]. Rejections [COS23]. Rejoinder [BH56, Din61b, FW07, Grü60, Grü61, Grü05, Luc68, Wal56c, WY73]. Related [Gil05]. Relation [Bea89, Ber10, DiS99, Din64a, Dob51a, Dob51b, Oka16, Rat20, Ros19, Smi07a, Why57, Wil50, Bok08b]. Relational [Bar82, Bro09a, Dar12, Hoe98, Hoo71, Hug04, Tel86]. Relationalism [Bin17, PB02, Ste20]. Relationism [But84a, Gar77, Hin88]. Relations [Bal57, Bir56a, Bir56b, BH91, Dow93, FRSW18, Kör67, Mic94, Odd90, Wil73a]. Relationship [Boh61, Gie73c, Mor86, Wis52a]. Relative [Fal79, Sac72, Tri94, Van81, VHH86]. Relativism [Gel85, Gro90, Kör86]. Relativist [Rad91b]. Relativistic [BM93, Gie73a, Ker87b, MLR21, Myr03, Pro68, Sch53, Swa21, Zah87, Sch54b]. Relativity [Dob21, FGK15, Fle16, JL22, Mil78, Pod75, Arn90, Art94, BJ03, BY06, Bon58, Chr81, Cul81, Din64a, Dor78, Fle51, Gar77, Gri77a, Hal51, Hut51c, Kes65a, Kes65b, Kes66a, Kes66b, Kes66c, KK83, Kil91, Kno10, Lev65, Mac86, Max93b, McC84, Mun86b, Ner82, Nor08, Pod76, Pod79, Rak97, Raw51c, Sac75, Sac76, Sac89, Sch65a, Tor15, Von87, Wei72, Wei79, Wei89, Whi51a, Whi51h, Why51c, Why51d, Din65a]. Relevance [BG16a, Den94, Gib69, HS68, KS19, Lai84, Leu12, Mac69c, SH20, Sto70b, SO19, vR13]. Relevance-Sensitive [KS19]. Relevant [BS92]. Reliability [Gie63, Fal00]. Reliable [Ham23]. Religion [Din53b, O’H93, PC12, Dow05, Sco08]. Religious [Fle79, PP78]. Remains [GT16]. Remark [Bin54, Mor88]. Remarks [Abo02, Bon55a, Car56, Dav05, Far67, Hal55b, KW97, Kre58, Pod75, Pod76, Sha88, Tri94]. Renaissance [Lak76]. Renormalizability [CL19]. Rényi [Leb89]. Reorganization [Bar13]. Replacement [Yos81]. Replication [Bir22]. Replications [Lav23]. Replies [Loc96b]. Reply [Ale59a, All01, Bai71, BG01, Bel79a, Ber78, BFHS02, Bra15, Bra62b, Can65, Cha83, Cha89, Cra92, CFM08, Cur83, Dia64, Din60c, Din61a, Din61b, Din65b, Dob58b, Dob70, Dup96, Erw82, Ezr57, Fai57, FW05, For95b, Fri09a, Gal59, Gil72a, Goo64, Gre63b, Gre64, Grü69a, Grü05, Han62b, HR95, Har60, Hen03, Hub08, Hul70, Jar65, Kap55b, Kap60, Kap61, Kes61, Kes66b, Kes66c, Kne57b, Kör68, Lad04, Lan01, Lat76, Lee95, Lew01, Lew03, Mac63a, Mac71b, Mar61a, May61, Mil66a, Mil78, Mul91, Nea95a, Nii82, Nor02, O’H84, Ols02a, Pat98, Pop55a, Pop56a, Pop56b, Pop78, Pre03, Pri94, Rus71, SW08b, Sch84b, Sch76, Set69, Set72, Smy62, Sob95, SGS97, Str96, Str05, Swi00, Swi05, Ten97, Urb94, Wat59d, Wat62, Wat89, Wil78a, Wis58a, Woo56, Wor89, Zah78]. Reply [Zah81, CM00, Dup00, Fol99, Obe99, Ros10, Sch00b, Ueb99]. Represent [BF11]. Representation [Car23, Chr10, Dev91, Dew18, Fri15b, JHB^{+}23, KL17, Kuk92, Mal23, Nel94, Ngu21, Pin11a, Ram07b, RKK19, Sil91, Le 07, Pin11b, Van08a, Van10, WC99, Fre11, Spr11, Bak15, Phi09]. Representational [Saa11, Ves20]. Representations [Bec15, Coh16, MA23, Pla16, Sha12, Sha97, She13]. Representative [Smy56]. Representing [MB21]. Reproducing [Cha02b]. Require [Cab20]. Requirement [Zah91]. Requirements [SS93, Sim83]. Rescher [Hau94]. Rescue [Cas16]. rescued [AHW95, HM99]. Research [Ake88, Bae16, Bet52b, Bla55, Boh61, BK59, Ell22, Hal79c, Hal79d, Kle85, Pog85, Sch74a, Tho92, Yux90, Rus00]. resistance [Koe99]. Resolution [Goo70, Max88, Squ89]. Resolving [Alb02]. Resonance [Cha64, Dob61]. Response [BH91, Cha11, Cra93, Cum92, Gre84, Mas01, Mor09, Pap88, Row23, Ste03a, Wil09a, Wil12, Fra99, Her07]. Responses [Wat91]. Restated [Ake90a, Pea82]. Restatement [HW04]. Resting [MD22]. Restoration [Mil90a]. Restoring [Gim04]. Restraints [Smi59]. Restricted [Ros12, SO19, Tor15].

Restriction [Cas98]. Result [KGB^{+}22]. Results [Ake88, Dan05, LU09, Spr18, Sta20, Wri18]. Rethinking [Buc16, D’A76, Sch13, Bon01]. Retraction [NS87]. Retrocausality [Eva18]. Retrodiction [BS92, Den94]. Retrospective [Pos60a]. Reveiws [Den58]. Revenge [Kle02, Pos74]. Reversal [AG09, Cal95, Far20, Sav94]. Reverse [GH16, Mac14]. reversible [Hut93]. Review [Act57a, Ada01, Aga78, Ale59b, Ale16, Amu99, Amu98, And10, And86, Apt84a, BJ04, Bai86, Bak15, BD86a, Bar01, Bar60a, Bat15, Bee98, Bee09, Bel01, Bel96, Ber10, Bir13, Bir96, Bir97, Bir12, Bol96, Bon01, Bon00, Bor54, BG16b, Bos06, Bot96, BL14, Bra65a, Bra99, Bri06, BS00, Bro86, Bro09b, BF00, Bun86, Bun03, Byr00, Cal98a, Cal15, Cal82a, Cam00, Cap86, Car67a, Car97, Car12, Car98, Car99b, Car04a, Cas95, Cha13, Cha95, Cha97a, Cha97b, Cha96, Che97, Cho11, Cla82b, Cla14, Coh86a, Col00, Col97, Col07, CS16, Col98, CE04, Cor97, Cou52a, Cou53a, Cou62a, Cra96, Cro52a, Cro52b, Cur82b, Cur16, D’A86, Dal90, Das58, Dav86, Dem01, Den86, DiS99, Dic98, Dic01a, Dic01b, Die99, Din54, Dor98, Dou13, Dow08, Dow00, Dre00, Dun62a, Dup94]. Review [Dup97, Dup13, Ecc54, Ecc55a, Eel00, Ell66, Ell96, Enç94, Enf08, Ero12, Eva98, Far00, Far16b, Fay95, Fel01, Fes06, Fie86, Fin96, Fod01, Fol00, For97, Fra88b, Fra92a, Fra92b, Fre10, Fre84, Fre95, Fre96, Fre98, Fre11, Fri97b, FB01, Fri16, Gal00a, Gal00b, Gan11, GG86, Gar99, Gat01, Gia06a, Gia06b, Gib84, Gib86, Gib89, Gil00a, Gil93, Gil92a, Gil96a, Gil96b, Gil01, Gil03, Gil07, Gil09b, Gil13, Gla14, Gly04, Gly11a, Gol98, Gow92, Gow97, Gra01, Gre97, Gri08a, Gro09, Gro11, Gua07, Gut00, Hal98, Hal15, Hal79a, Hal86, Hal99, Hal01, Har96a, Har96b, Hau99, Haw97, Haw99, Hea86, Hea99, Hei99, Hen97, HM99, Hey96, Hey16b, Hin95, Hit00, Hod80a, Hod97, HR92, Hoe00, Hof01, Hog86, Hoo86, Hoo96, Hor93, Hul98, Hut51a, Hut53c, Jar12, Jar79a, Jef54, Jen86, Jen92]. Review [Joh03, Joh15, Ken94, Ket00, Ket01, Kil93, Kir01, Kit91, Kne65a, Koe99, Kör59a, Kör86, Koz14, Kra09, Kre60a, Kre60b, Kri92, Kuk97, Kus01, Kut98, Kut00, LaP03, Lad96, Lad99, Lar00, Lau66, LM99, Lav95, Law94, Le 98, Lev97a, Lev01, Lew14, Lil99, Lip93, Lip94, Lip99, Llo66a, Los98, Low97, Low00, Low04, Lyo06a, Mac96a, Mac98, Mac94a, Mac96b, Mac96c, Mai92, Mai95, Man96, Mar00, Mas54, Mau95, May00, McC16, McK16, McM97, Mel99a, Mel00, Mer88, Mil86b, Mil94b, Mil96, Mil97b, Mis97, Mon94, Moo96, Mor93a, Mor97, Mor99, Mos86, Mos88b, Mum13, Nad99, Nee66, Ney11, Ney14, Nic95, Nol80a, Nol89, Nol96, Nor04, O’H86, O’M14, Oft07, Oka00, Pag00, Pap80a, Pap91, Par97, Pas16, Pau16, Per07, Phi09, Pin11a, Pin16, Pol14, Pow92, Pre94, Pre97, Pre01, Pri99]. Review [Pri01, Pro86, Psi94, Psi96, PS00, Put01, Qui61, Rad00, Rai86, Rav60a, RW13, Red74c, Red77, Red96, Red98, Red00, Reu09, Rey15, Ric95, Ric52a, Ric16, Roc86, Ros99, Ros08, Ros09, Row23, Rub86, Rue98, Rus00, Rus55a, Ryc99, SL13, Saf86, Sch07a, Sch08a, Sch59b, Sch11a, Sch00a, Sch14, Sco08, Seg97, Ses99, Sha95, She06, She01, Sho98, SF13, Sid01, Sie97, Sie10a, Sim97, Sim01, Sin00, Skl96, Sky96, Sky12, Sla15a, Sma00, Smi00a, Smi00b, Spi60a, Spi60b, SGS97, Spr11, Squ96, Ste12a, Ste10, Ste96b, Ste00, Sti66, Sto99, Str09, Suá96, Sup98, Swo95, Tal15, Tau06, Tei99, Tei13, Tha09, Tho96, Til00, Tom99, Too95, Tso10, Tuz93, Ueb94, Ueb97, Ueb98, Ueb01, Vai15, WF13, Wal14, Wal97, Wat54, Wat95, Way95, Wei14, Wei94, Wer10, Wes79b]. Review [Wes97, Whi51c, Whi51d, Whi58a, Why52c, Why92, Wil96a, Wil14a, Wil91, Wis54a, Wis54b, Wis57a, Woo98, Woo58a, Woo94, Woo00b, Wor95, You66, YHF68, Yud16, Zah82b, Zah97, dR96b, Har11]. Reviewed [Odd81]. Reviews [Ach68, Ack65, Act56, Act57b, Aga60, Aga87, Ale52a, Ale52b, Ale54, Ale57, Ale59c, Ale60, Ale62a, Ale62b, Ale67, And78, Apt84b, Arc58, Arc59, Arm58, Arm59a, Arm59b, Arm70, Ash55, Ash57b, Ash57c, Ash59, Ash63, Aye70, Bal71, Bar69, Bar85b, Bar79, Bar63, Bar70, Bar71, Bar60b, Bas64, Bea66, Bel61, Bel51, Bel72, Bel75, Bel76, Bel77, Bel79b, Bel82, Bel71, Bel97, Ber55a, Ber59, Ber69, Bet51e, Bet52a, Bet59a, Bet59b, Bev52, Bin71, Bir56c, Bir85, Bla58a, Bla60a, Bla69, Bla70, Bla71, Bla58b, Bla59, Bla60b, Bla62, Bla63b, Bla84, Bla82, Bod64, Bod66, Bod68, Bod69, Bod70a, Bod70b, Bod71, Bod74, Bod75, Boh66, Boh68, BK59, Bon65, Boo79, Bra79, Bra80, Bra55b, Bra65b, Bri71, Bro69, Bro87, Bro62a, Bro65, Bro62b, Bro51, Bro52a, Bro52b, Bro52c, Bro54, Bro56, Bro64a, Bro78, Bro79, Bro84, Buc61, Buc63, Buc83]. Reviews [Bun58b, Bun58c, Bun60, Bun63, Bun64, Bun68b, Bun73, Bus52a, Cal52, Cal58, Cal67, Cal97, Cal82b, Car66, Car67b, Car76, Car77, Car83, Cas57, Cas64a, Cas64b, Cas69, Cas50, Cha70b, Cha79a, Cha79b, Cha82, Cha85b, Cha85c, Chi69, Chi70, Cla87, Cle70b, Clo66, Coh63b, Coh67c, Coh70, Coh72, Coh79, Coh81, Col56, Col57, Coo79, Cop80, Cop51, Cou52b, Cou52c, Cou53b, Cou62b, Cow52, Cra69b, Cra69c, Cra70a, Cra70b, Cra71, Cra68, Cra69d, Cro52c, Cro52d, Cro52e, Cro52f, Cro52g, Cro52h, Cro52i, Cro53b, Cro53c, Cro54, Cro59, Cro63, Cro62, Cur75, Cur79, Cur81a, Cur81b, Cur82c, Cur85, Cur87, D’A84, Dal79, Dar83, Dar54, Dav54a, Dav55, Dav57, Dav59, Daw65, Daw70, De 59, Den82, Den85, Den87a, Den87b, Den69, Den59b, Den60b, Dil84, Dil76, Din50a, Din53a, Din53b, Din64b, Dor75a, Dor76b, Dor76c, Dor76d, Dre85, Dun62b]. Reviews [Dun62c, Dun87, Dun67, Ecc53, Ecc55b, Egg80, Ell67, Elz83, Emm68, Eva68, Eva71, Eve87, Fey56, Fey58, Fey59, Fey61, Fey63, Fey64, Fey67, Fin68a, Fin68b, Fin69, Fin71a, Fin71b, Fle52, For62, Fox75, Fra83, Frö69, Gab57, Gab58, Gab66, Gan71, Gar67, Gar74, Gea69, Gel61, Gel83, Geo52, Geo53, Gib82, Gib81, Gie73b, Gie74, Gil68, Gil82a, Gom54, Goo58, Goo71b, Goo76, Goo60b, Goo60c, Gou60, Gow70, Gra83, Gra73, Gre74, Gre62a, Gre62b, Gre63a, Gre70, Gri68, Gri77b, Gri70, Grl57, Gut53, Haa71, Haa76b, Haa77, Haa80, Hac67, Hac65a, Hac65b, Hal52b, Hal62, Hal67, Hal72, Hal75a, Hal79b, Hal87, Hal78, Hal75b, Hal61, Ham65, Ham67, Ham68, Har61a, Har63, Har67, Har68a, Har61b, Har71, Har84, Hei83, Hen82, Hen66, Hes57, Hes53b, Hes53c, Hes56, Hes58b, Hes59a, Hes59b, Hes60a, Hes60d, Hes61, Hes67a, Hes67b, Hes68a, Hes68b, Hes70]. Reviews [Hir66, Hir67, Hir78, Hir62, Hir63, Hir65, Hod80b, Hod78, Hol79, Hor79, Hot63, How58, How80, How69, How71, How87b, Hum83, Hun69b, Hut51b, Hut52a, Hut52b, Hut52c, Hut53e, Hut53d, Hut53g, Hut54c, Hut54d, Hut55b, Hut55e, Hut55c, Hut56b, Hut58a, Hut59a, Hut59b, Hut60a, Hut60f, Hut60e, Hut60b, Hut60c, Hut63b, Hut63c, Hut63a, Hut65b, Jac53, Jar67b, Jar69, Jar61b, Jar61c, Jar63, Jar64c, Jar64d, Jar64e, Jar73, Jar74, Jar79b, Jar80, Jar83, Jef52, Jef53, Jef56, Jon66, Jon69b, Jon71, Jon80, Jon82b, Kam83, Kam85, Kap53, Kap55d, Kap56a, Kap56b, Kap57a, Kap57c, Kee66, Kee70, Kek81, Ken70, Kil62, Kil81a, Kil81b, Kil83a, Kil83b, Kil83c, Kil84, Kla69, Kle87, Kne54, Kne57a, Kne60, Kne62, Kne63a, Kne63b, Kne65b, Kne66, Kne67, Kne68a, Kne65c, Koe73, Koe76, Koe78, Koe80a, Koe80b, Koe83, Kör59b, Kör59c, Kör59d, Kuh66]. Reviews [Kuh67, Kuk87, Lac69, Lam62, Lan57, Lan61, Lan51, Lan83, Lar81, Lau67, Lau68, Lau69, Lau70, Laz58, Lea69, Lej55, Lej56, Lej57, Lej58b, Lej58a, Lej59, Lej63, Lep87, Lev68a, Lev68b, Lie67, Lin84, Llo66b, Llo76, Lon51, Lon52, Low73, Luc69b, Luc78, Luc80, Luc82, Luc85, ML75, Mac68a, Mac69a, Mac69b, Mac78a, Mac53b, Mac60b, Mac60c, Mac65a, Mac65b, Mac67, Mac71d, Mac78b, Mac79, Mac87, Mac65c, Mac68b, Mac68c, Mac71e, Mac73b, Mac74b, Mah70b, Mah70a, Man78, Man74, Mar87, Mas53, Mat81, Mat82, Max69b, Max69a, Max71, May56a, May56b, May57, May59, May64a, May64b, McC50, McC51a, McC51b, McC52a, McC52b, McC53, Mcf75, Med54, Mel67b, Mel69, Mel83, Mel84, Mid85, Mil51, Mil74d, Mil76a, Mil67, Mil70, Mil83b, Mil84, Mil82, Mil83c, Mil85b, Mil87b, Min66, Mol81, Mor85, Mos84, Mot83, Mou85, Mun56, Mun57, Mur74, Mus69]. Reviews [Nag59, Nar68, Nar70, Ner78, NS69, NS78, NM83, Nid71, Nol80b, Nol84, Nor66, Nor67, Nor71, O’B69, O’C54, O’C56a, O’C56b, O’C56c, O’C63, O’C72, O’C75, O’H83, O’H85, Odd87, Ogd51, Ord83, Ord87, Pad64, Pap78, Pap80b, Pap82, Pet67, Pey57, Phi79, Phi80, Pic85, Pir56, Pol71, Pol69, Pop79, Pop83, Pos62b, Pos71, Pos75b, Pos78, Pot69, Pra79, Pra84, Pri59, Pri60, Pri98, Pri68, Pri53, Pri69, Pro77b, Pry60, Prz62, Puc56, Pur71, Qui77a, Qui79, Qui77b, Rai84, Rat67, Rat68, Rav60b, Rav60c, Rav60d, Rav61, Rav62, Rav65, Rav78, Rav87, Raw50, Raw51a, Raw51b, Red80b, Red81a, Red81b, Red85b, Res87, Ric51, Ric52b, Rit62, Rob65, Rob68, RH77, RH79, RH81, Ros58, Ros70, Rub81, Rub82, Rub80, Rus52, Rus55b, Rus55c, Rus55d, Rus56, Rus57, Rus60, Sab61, Sai85, Sam81, Sau87, Sch84a, Sch69b, Sch81]. Reviews [Sch65b, Sch75, Sch80, Sch82, Sch83, Sch79b, Sci68, Sci78, Scr62, Sel52, Set77b, She67, She70, She71, She64, Shr85, Sin70, Sin71, Sko72, Sko82, Slo85, Slo66, Slo68, Slo70, Slu67, Slu80, Slu85, Sma70, Smi70, Smi85b, Smi85a, Smy60b, Soa54, Spi53a, Spi53b, Spi53c, Spi55a, Spi55b, Spi56, Spi60c, Spi60d, Spi60e, Spr53, Sta67, Ste54, Ste62, Ste71, Ste70, Ste78, Ste84, Sto70a, Sto79, SR63, SR64, Suc67, Suc70, Sum53, Swi87a, Swi87b, Swi70, Swi76, Swi78, Swi80b, Swi81, Syn52, Syn62a, Syn62b, Tan67, Tay52, Tel71a, Tel71b, The70, Tho66a, Tho54b, Tho66b, Tic80, Tou53, Tou58, Tou60, Tow56, Tri87, Tur60, Urb81, Van61b, Ves63, Von75, Von83a, W.52, Wal69, Wat84, Wat53, Wat55c, Wat55a, Wat55b, Wat57b, Wat59a, Wat59b, Wat71b, Wat72, Wat64, Wat65b, Wat65c, Wea71, Web70]. Reviews [Wei84, Whi78a, Whi51e, Whi51f, Whi51g, Whi52, Whi54c, Whi55b, Whi55c, Whi55d, Whi57a, Whi57b, Whi58b, Whi58c, Whi58d, Whi58e, Whi59, Whi62, Whi63a, Whi63b, Whi68, Whi69, Whi75, Whi80, Whi83, Wht61a, Wht61b, Why52d, Why58a, Why58b, Why60, Why67, Why69, Why71, Why72, Wig55, Wig70, Wil82a, Wil82b, Wil83, Wil53, Wil58, Wil79a, Wil82c, Wil60a, Wil60b, Wil66, Wil70, Wil73b, Wil73c, Wil81b, Wil62, Wis50, Wis52b, Wis53g, Wis53c, Wis53d, Wis53e, Wis53f, Wis54c, Wis54d, Wis54e, Wis54f, Wis54g, Wis55a, Wis55b, Wis56b, Wis56c, Wis56d, Wis56e, Wis57b, Wis57c, Wis58b, Wis58c, Wis58d, Wis58e, Wit50, Wit51, Wit52, Wit53, Wit55, Wit56, Wit57a, Wit57b, Wit58a, Wit61b, Wit61c, Wit61d, Wit62a, Wit62b, Wój78, Wol59, Woo51a, Woo51b, Woo52b, Woo52c, Woo52d, Woo53a, Woo58b, Woo58c, Woo58d, Woo59, Wor85, Yos78, You68]. Reviews [You58]. Revised [Pop68, Suc69]. Revision [Yud16, Zah91, Mor13]. Revisionism [Gil82b]. Revisions [Coh86a]. Revisited [BHM72, Ben90, But07, CS98, Dev10, FW05, GC11, Han05, HY02, KL17, Lip82, Liz93, Mus77, Ste06b, RS07]. Revitalized [Slo93]. Reviws [Koe74]. Revolution [Bro64b, CS98, Hal70, Hal82, Kle70, Kne68b, Ueb96, Van08b, Hei58, Hei62]. Revolutionary [Kle85]. revolutions [ABC06, Kuh62, Tha09, Bir12]. Rey [Sch01b]. Rheology [GBW61a, GBW61b, Wit61a]. Rhetoric [Gre79]. Richard [And86, Sch14, Grü05, Kut00, Ric16]. Richards [Jen86]. Richardson [Fre10]. riddle [Cha00, Sch00b]. Right [Bla20, Gly10a, Han04, HN14, Nol99]. Rigor [Dav03, Pas16]. Rimini [AGTZ08]. Rindler [CH01]. Rise [Gil07, Zah77, Car03]. Risk [Bra93, CMS22, HS04, Kos21, Kuo24, SB19, Wei86]. Risk-Sensitive [CMS22]. Risks [FRZ12]. Rivals [Grü76b, Grü86]. Road [How75b, HHS02, Rea04]. Robert [Cap86, Fol99, Hof01, Fre10]. Roberts [Car12]. Robin [Bri06, Phi09]. Robinson [Puc77]. Robot [Kri61]. Robots [Gun68]. Robust [Jon18]. Robustness [Boo24, FS21, KGB^{+}22, KLM10, Sch18, Tro93]. Roger [Dod13, Red00]. Role [AH14, Bat02, Bat10, Bon01, CL19, Cur21, Fin81, dF81, Hor94, Kha13b, Lav23, Lav77, Pin11c, Saa11, Sha82, Sme15, Was78, Why52b, Zah80]. Rooijen [Peć89, Set89]. Rooms [Spr07]. Roots [Pop52]. Ros [Van94b]. Rose [Rad00]. Rosen [Bha86]. Rosenberg [Van88]. Rosenberg/Williams [Van88]. Ross [BL14, Dow08, Hol94, Sch76]. Rotating [But06a, MLR21]. Rotation [Hug81]. Roush [Bar08b]. Routledge [Bos06, Lyo06a, Mum13]. Rowbottom [Gil13]. Roy [Pop58b]. Royal [Ses07]. Ruben [Tuo89]. Ruetsche [Wal14]. Rule [Bat71, Bir14, De 72, DW17, dF81, Gil71, How75c, Hun70, Juh94, Mil68, Jos82, Spi74]. Rules [Lep76, Lev63, Rob94]. Run [Jos82]. Ruse [Hul70]. Russell [Pri60, Bin54, Bra70, Byr00, Hal81, Kne71, Lan98, RS07]. Russo [Ste12a].

S [Col00, Dre00, Gil00a, Hon91, Kat84, Mel00, Oka00, Str09]. S. [W56]. Sabat [Øhr86]. Sacred [Fri09b]. Sad [But09, Hog04]. Salience [Bar23]. Salmon [Cam00]. Sameness [Akk94, Nee17]. Samir [Bir13]. Sampling [Joh89]. Sampson [D’A78]. Sand [Gut00]. Sandra [SF13].

Sarkar [Oka00]. Sarton [C.56b, C.56d]. Saudi [Hog86]. Saunders [BM10, McL07, SW08b, Tap08]. Saved [FA14]. Saving [Lus21, Mas07, Mor00, PR95, Ros10]. Say [FW21]. Says [FW21, Wri83]. Scalar [Pit11]. Scale [Bur18, Mil83a, Pin22]. Scales [Grü54, Whi54b, Grü53a]. Scandal [Bro58]. Sceptic [Wis58f]. Scepticism [Buc59, Cha89, Gem89, Mil90a, Sch84b, Smi17, Wat85, Hea86, Mos86, O’H86]. Schabas [Gro11]. Schema [Nis70]. Schemas [Bae12]. Schizophrenia [Bor16]. Schleichert [Din65b]. Schlesinger [Kap60, Kap61]. Schlick [Pla63, Wat77]. Schrenk [Reu09]. Schrödinger [AGTZ11, Bas92, Bub79, Hol88]. Schulte [Cha00]. Science [A.56b, Aga59b, Ale16, Ale18, Arc59, BH92, BJ04, Bak09a, Bar62, Bat10, Bet50, Bet51c, Bet51d, Bir56a, Bir56b, Bla09, Bla79, Bra70, Buc51, Buc23, Bun86, Bus52b, Cal50, Car85, Cha89, Cla00b, Cle11, Coh73, Coh74, CS16, CH17, Den59a, DB52, Din53b, Dow08, Enf08, Fey66, Fey80, Fra50a, Fra50b, Fri02a, Gar73a, Gie75, Gie73c, Gla24, Goo94, Grü76c, Hac79, Hea54, HH92, HL11, Hut65a, Irv12, Irz85, IK19, IK24, Jar64a, Jen84, Jen86, Kan88, Kap58, Kap65, Kit11, Koe00, Koe04, Kor91, Koz14, Kuh80, Kuk08, LRP23, Lel93, Llo67, Los86, Lou62, Lyn90, Mag10, May96, May60, McA86, Men54, Moo78, Nem03, New17, Nyr21, O’H93, Per07, Pet22, Pet23, Pin15, Pir52c, Pol67, Pop52, Pot51, PLKG24, Qui57, Rad87, Rad91b, Red84]. Science [Roc86, Ros09, Rus00, Saa11, Sam75, Sco08, SB10, SF13, Sma72, Szu77, Ten85, Til94, Til00, Tor84, Tuc11, Wal07, Wat52c, Wat75, Wat85, WOB21, Wei11, Wer19b, WB53, WB54, Whu56, Why63, Wig61, Wil13a, Wil03, Woo51c, Woo60b, Yud16, Zah77, Zem87, But84b, COS11, Dow05, Fie86, For99, Hei58, Hei62, Irv13, Jos10, LM99, Max04, Mir04, Mit09, Mor13, MT13, Sta06b, Sti83, Til84, Pin11c, And86, Den60a, Dou13, HH63, HH64, Hoo86, O’H86, Ale16, Bel01, Bra99, Fri16, Gut00, McK16, SL13, Sma00, Lip99, Hea53]. Sciences [Bla55, Con03, Cur17, Dun11, Ell57, Fra95, Grü56, Hog86, MC12, Mos91, Roh88, Ros12, Ros55, Ste12a, Stu61, Tav91, Was55, Wat57a, ADS84, Gel85, Kha13a, Rus08, Woo00a, Kör86, Fre11, Sla15a]. Scientific [AA87, Ake90b, And10, Azz04, BF90, Bak15, Bam93, Bar00, Bir12, Bla09, Boh61, Bro90, Bro09b, Buc65, Cam00, Cha03, Chi57, Cho11, Cum92, Cur23, Cur86, Cus90, Dan05, DiS99, Din51, Dor72, Dor75b, DE76, Fle79, FV11, FS21, FA14, Gal57, Gen98, Gil09b, Giu88, God69, Goo09, Gro90, Har59a, Hey16a, HH92, Hub05, Jon82a, KZH^{+}22, Kav23, Ket00, Kle70, Kne68b, Koe04, Kör67, Kos21, Koz14, Kra09, Kuk94, KZ16, Lac70, Lan13, Lar00, Leu14, Lud17, Lyo06b, Mar71, Mar72, Max93a, Max93c, Max93b, Miz23, MR94, Mur90, Nol87, OL23, Pag00, Pir52b, Pol58, Pos60b, Pov18, PLKG24, Pra24, Pro59, Psi00, Rat04, Ric16, Sch00a, Seg08, SS93, SG73, Sim83, Smy58b, Sta09, Sud73, Sud76, Tei62, Tha09, Tia85, Til93b, Til00, Tuo78]. Scientific [Van08a, Van10, Vic21, Wat91, Wil19, Wit59, Wit61a, Woo79, Yux90, ABC06, AHW95, Cha07a, Cha04, Col07, HM99, Hun06, Irv13, Kuh62, Lad99, Mas05, Mur06, PP78, Pin11b, BL14, Pin11a, Tau06]. scientist [Mee99]. Scientists [Bla79, Mah90, Puc78]. Scope [Qui57, Why58c, Wri18, Wor00]. Scoring [dF81]. Scott [Lev01, Pri01]. Screening [SS13]. Screening-Off [SS13]. Sea [DEO19, Mul05a, Sob01]. Search [Bla50, CL19, Eva02, Han05, KW97, NS20, Rob23, SGS97, Wal10]. Searching [RS15]. Sebens [DF22]. Second [Bla63a, Bra63a, Eas14, Grü61, Ket01, Mad07, Mus71, Pop56a, Pop57b, Sch61b, Swa22, Swi05, Zah81, BS94, Buc51, HS22, Rob23, Sie10a, Wyk76, Grü05].

Second-Order [Ket01]. Secondary [Bro19, MP93, Pri89]. Securing [Sta20]. Seeing [Ben12, Car04b, Cle55, Dav18, Gil92b, Mat05, Smo21, Gia06a, Gia06b]. Selection [ARR15, Aut20, Bir20, BR04, Bou15, Cla16, Cro08, DiF19, FN21, Gla21, Gly08, GS08b, Gre61, Kie01, Kot12, Lew01, Lew10, Mam04, Man97, Mar89a, Mil06, Mol22b, Oka01, Oka04, Oka05, Oka08, Oka16, Pet96, Ree78, Sob95, SO03, SW18, Ste16a, VC24, Was78, For99, Gay98, Mat99, Amu99]. Selections [Suá04]. Selective [Har13]. Self [BS17a, BSM19, Bar23, Cha02b, Dew75, Dob57, Gia06a, Kap57b, LaC22, Lar03, O’M14, Puc76, Sch79a, SC18, Sho00, Car04b, Pra12]. Self-Assembling [BSM19, Bar23, BS17a]. Self-Assembly [LaC22]. Self-dependent [Sho00]. Self-Excitations [Lar03]. Self-knowledge [Sch79a]. Self-locating [SC18]. Self-Organisation [Gia06a]. self-organization [Car04b]. Self-Reproducing [Cha02b]. Self-Winding [Dob57, Kap57b]. Selye [Le 52]. Semantic [BHC53, FA14, Mił23, Ryn83, Shu19, Van94c, Vol95, Wag15]. Semantical [Smo77]. Semantics [BM10, Hut53e, Isa19, Sil91, Chi98]. Semaphore [Cod75]. Semirealism [Nan13]. Sensation [BZ84, Puc69]. Sense [Bra60, Den53, Fri04, Gia06b, Lee82, Mac72, Man21b, Mat05, Max66, Oft07, Smy58b, Smy62, Vie20, Mor13, Ses05, Yud16]. Sense-Data [Bra60, Smy58b, Smy62]. Senses [KGB^{+}22]. Sensitive [CMS22, KS19]. Sensory [Pen21]. Sentences [Nii79]. Separability [DF22, Bha91]. Separable [Bha86]. Sequences [Coo03b]. Serendipity [Van94a]. Series [Hoo03]. Serious [Kho20, Sta09]. Seriously [Lar03, Zyc89, Sch00b]. Sesardic [Oft07]. Session [Ezr56, Spi57]. sessions [PP78]. Set [Dau88, May77b, May77a, Sha03, WC17]. Sets [Coo03a, Lev97b, Mul01]. Setting [DW17, NS21]. Settings [COS23]. Settle [Bai71]. Severe [MS06a]. Severity [Gil90]. Sex [Lew01, Mat03, Gra01]. Sexual [Rus00]. Shackle [Cod75]. Shadow [vF06]. Shadows [Mac03, Rub22a]. Shaky [Urb83]. Shandy [Sma86]. Shannon [Isa19]. Shape [Ben12, Gre17, Mag05]. Shapere [Los86]. Shared [Kho20]. Sharp [Kle02]. Sharpened [Goo68a]. Shed [Ern90]. Shenker [Hen03]. Sherrington [Ecc52]. Sherwood [C.56a, C.56c]. Ship [Bud97]. Shoe [Fre88, Goo67c, Goo68b, Hem67, Sim91]. Shogenji [Sch11b]. Should [Bra21, Cal04, Kop05, Sob98, VC24, Wil75b, Wal06a]. Show [Col08]. Shrader [Ale16]. Shrader-Frechette [Ale16]. Side [Lan01]. Sight [Pir53]. Signal [Ale14, Mor24, Pod79, Why53c]. Signalling [Bro20, FC18, SGSC18, Wag15]. Signals [CPG20, GG86, Sky10, WF13]. Significance [Bir12, Cuf17, Fin64, Fle16, Gie72b, Goo63, GBW61b, Gre11, Joh87, Lis04, Mos86, Ram23, Tro93, Hug99, Fie86]. Similar [Hul74]. Similarities [SS17]. Similarity [BM18, Fle16, Tal15, Urb83, Wei13]. Simple [RS96, Ryd81, SGSC18]. Simpler [FS94, Fri02a]. Simplicity [DGO07, For95a, Fri72, Goo68a, Kie97, Man76, Mil19, Pos60b, Pos62a, Ram07a, Tur91, Woo14]. Simpson [Ket01]. Simulating [Gol06, Gan11]. Simulation [DTW17, Par17, Wei13, Tal15]. Simultaneity [Giu01, Har68b, Pet89]. Sinai [Fri04]. Since [HHS02, Pop57a, Rea04]. Single [Joh88, Lat76, Pri84, Set77a, Sto86, Why53b]. Single-Exit [Lat76]. Singular [Hei78, Nii79, SP12]. Singularist [Nan13]. Singularity [Pit08]. Sir [W56, Ecc52, Pri60]. Situation [Kin16]. Situation-Specific [Kin16]. Situational [Coa74, Lat72, Mac74a]. Situations [GP14, TP19, TP21]. Size [Dau88, Hal81]. Skinner [Col07]. Sklar [Bat15, Bel01, Sko07]. Skow [Pau16]. Skyrms [WF13, Bai86]. Slant [EPW13]. Slater [Rey15, SL13]. Sleeping [Bra11, Bra15, Gro08, Lun21, Wil14b]. Slit [Mil85a]. Small [Fis79, Red00, Mil01]. Smaller [Bur18]. Smith [FB01, Cra93, Koe99]. Smythies [Bra62b]. Snyder [Ols02a]. So-Called [HO79, Leb60, Mac66, Mil66a]. Soames [Pri01]. Sober [Dup13, Wei05, Den94, Kuk95, Nea95a, Ste03a]. Social [Bir56a, Bir56b, Bla55, DW17, Ern01, Fel01, Gie75, GC19, Grü56, Gua07, Jar61a, Jar64a, Kör86, Nem03, PG13, RW13, Ros12, Ros55, SW19b, Sla15a, Sme15, Smo21, Ste12a, Tuo89, Urb80, Wat52c, Wat57a, BV05, Bic06, Gel85, Kha13a, Rus08, Wra11]. Societal [Man57]. Society [Den60a, Gua07, HH63, HH64, Irz85, Bic06, Kit11, Dou13]. Sociologist [Jon86]. Sociology [Coh63a, Hor65, HH92, Jar64b, Jen84, Jen86, Kat84]. Software [Coh87, Sch11a, Dor05]. Solipsism [Pla62]. Solution [DeV97, Dou02, Fos69, Gal14, Gib69, Lun21, Ott94, Tur90, Vra04]. Solve [Haa91, Max93a, Rob94, Dor78]. solved [Dor78]. Solving [Ake93, Cha07b, Ell88, Pol57, Fri00]. Somatic [Col08, LB13]. Some [Abo02, AK91, AK95, Bat71, BM01, Ben56, Ben57, Bla50, Bri51b, Bri51a, Bri51c, Bub79, Buc51, Chi87, CB69, Coh80b, Dav05, Geb22, Gil92b, Gra55, Hea86, How85, Hum04, Kne57b, Kne71, Kör67, Kre56, Kul77, Mac71b, May08, Mil90b, PNSBC94, Pod75, Pod76, Sam60, Sär68, Sut57, Was55, Wei89, Why63, Ern75, Wor00, LU09, Mil78]. Somebody [FW21]. Sons [Hog86]. Sophistication [Dew19]. Sorites [Car69]. Sort [Rob76]. Sortals [DiF19]. Soundness [Ber23]. Sources [Bet24, Buc59]. Space [Abr58, Art94, Bar82, Bel05a, Bra60, Bri06, Cha58, Chr81, Dor78, Goo59a, Hin88, Hoo71, Hor75, Hug04, Hug08, Kap59, Lac70, Lew04, Mcg70, Mül05b, Sac72, Smy58a, Smy62, Swi80a, Tor84, Whi55e, Why53a, Why54b, Zar79, Kel10, KP99, Le 03, Pla00]. Space-like [Chr81]. Spacelike [Arn94]. Spaces [Smi10]. Spacetime [Bin19, Bud97, Cur18, Dis94, EN87, Hel98, Hoe98, Hug04, Kno14, Leh11, Mel80, Swa21, WT10, Bil00]. Spacetimes [MLR21, Wel08]. Spatial [BE21, BD99]. Spatiotemporally [Ros12]. Speaking [Mac03]. Spearman [Lan95]. Special [Bat07, Cal04, Dav56, Dun11, Hub11, Puc74, Wei11, Woo00a, Arn90, BY06, Chr81, Cul81, Din64a, Dor78, Max93b, McC84, Mun86b, Ner82, Rak97, Tor15]. Specialness [Bat07]. Species [Bar19b, Gir76, Gri06, Hul74, MP08, Rey15, Ros96, Rus69, Rus71, Rus87, Vel10, Sla13, Rey15, PS00]. Specific [Kin16, Wei19]. Specker [Gar72a]. Spectral [AH14]. Spectre [Rad00]. Speed [Goo71a, Juh94, NW22, O’C17, BM93]. Speed-Optimal [NW22]. Speed-Optimality [Juh94]. Spell [Bub73]. Sperber [Dri11]. Spilsbury [Ezr57]. Spin [Sac89, Von87, Wei89]. Split [Bog77, Dew75, Har13, Puc76, Rob76]. Split-Brain [Dew75, Puc76, Rob76]. Split-Brains [Bog77]. Spohn [Hub11]. Springer [Bun86]. Springer-Verlag [Bun86]. Squads [Wei05]. Squared [Bai83]. Stability [BH92, Coo12, Pol52b]. Stable [Arm93, Cus22, Mar90]. Stage [Jon82a]. Stalinesque [Tod09]. Stalinesque/Orwellian [Tod09]. Stand [Ste03a]. Standard [Giu88, MS07, PC12, Qui74, Rox75, Vra04, Fra80]. Standing [Erw84]. Stanford [Enf08]. Stapp [CBR90]. Stasis [Hul65a, Hul65b]. State [Arc59, BS00, CBC^{+}22, Kle17, KZ16, MD22, Psi00, Sac93, SB10, Swa21, WT10]. Statements [Blu70, Gil71, God69, Gou57, Puc62, She73, Was55, Wat60b, De 72].

States [Bon98, Bot94, Wil80a]. Statistical [Alb62, ARR15, BGTW24, Bar72a, Bar72b, Bux78, Col00, Coo05, Dow93, ER96, Fin81, FW21, Gri88, Kie01, Lav77, LM85, Len06, Par11, Red74a, Rob20, Skl00, Tso10, Coo02]. Statistics [Col66, Din70, Gre61, Hac64, Hac81, LM85, vFH17, PK70a, PK70b, PP78]. Status [Ben63b, Ben63c, Cha02a, Nol87, Nor04, Pan62a, Pan62b, Sac76, Kos00]. Steel [CS16]. Stegmüller [Pea82]. Steiner [Sim01]. Stem [Cur16, Fag13]. Stephen [Hoo86, Bir01, Bos06, Cha13, Ket01, McK16]. Sterelny [Gra01, Lew14, Sch14, She06]. Stewart [Jen86]. Stich [Bot94, Hoo86]. Still [Bau68, MTH^{+}24, MG17, Ste03a]. Stimulus [LM01]. Stochastic [Aga90, But07, Pla00]. Stokes [Ket00]. Stöltzner [Fes06]. Stone [KW97]. Story [EN87]. Stove [Row93]. Straight [Hun70, Juh94, Mil68]. Strange [Bar01, Sty00]. Stratagema [Lyo06b]. Strategic [Bir13, OB12]. Strategies [Arm93, Rat20]. Strategy [Bol83, Mag03, Mar90, Wit59]. Strawson [Hea86]. Strength [Nel96]. Strevens [Wer10, FW05, FW07]. Strife [Bun55a, Bun55b]. String [Ric16]. Stroboscope [Smy56]. Strong [Lei13, Lew05, Nie21]. Stroud [Mos86]. Stroyls [Hog86]. Structural [Gli20, Goo10, HP05a, HP05b, Kro89, LW15, Leu14, McK14, Rob11, Ruy19, Sch08c, Sch09, Sha12, Tha98, Why55a]. Structural-Model [HP05a, HP05b]. Structuralism [CB12b, LT17, McK20]. Structuralist [Res13, SW19a]. Structure [AGTZ08, And10, BD86c, BD86b, Bar15, Bir20, Bir12, CK89, Cur18, Dor75b, Fey66, FS24, Giu88, GS08b, Goo59a, Hor94, HHS02, KZH^{+}22, Kle70, Kno14, Leu14, MHB^{+}23, NTW20, Rea04, Rud64, Sch08b, Tha09, Thé16, Wil69, vF06, ABC06, BD99, Hun06, Kuh62, Pas16]. Structures [Bok08a, Bye69, Fri15b, Wol12]. Struggle [Amu99, Gay98]. Studies [ADS84, Chi57, Con03, Dew75, Mum13, PK70a, PK70b, Hoo84, Wit58c, Hog86, Pag00]. Study [AGTZ14, Cla53, Gri96, Lac70, Low04, Mac03, MC12, Miz23, Mor86, Nug85, Pot10, Rus51, Sch00c, Sta17, Tal16, Tuc11, Whu56, Why52b, Gel59, Mac99]. Stuff [ELLH17]. Sturgeon [Lev01]. Styer [Bar01]. style [Mat99]. Subdivisions [Ram13]. Subject [Gal57, Pol58]. Subjective [Bux78, Den53, Ell73, Gil72b, Hac66, Hub05, Moo83, Wil99]. Subjectivism [Gal03]. Subjects [CG09]. Substance [Nee10, Nee17, vF06, Haw99]. Substantivalism [EN87]. Substitution [Pen21]. Substitutional [Byr00, Par82, Lan98]. Substratum [Gri77a]. Subsumption [BM05]. Subsystems [Ket01]. Subtraction [Ros10]. Success [Dou23, ER96, Kel00, Mag03, SV11, Sch09]. Succession [How75c]. Successive [Max68]. Such [Row23]. Sudden [Ses07]. Sufficient [Ner71]. Suggested [Goo70, Sab54]. Sullivan [She01]. Super [Van94b]. Super-task [Van94b]. Supercomputer [Scr63]. Superdeterminism [BL23]. Superdeterministic [COS23]. Superfluous [NTW20]. Superiority [Cur83, Lan01]. Superposition [Ada94]. Supersede [Pro74, Zah73a, Zah73b]. Superselection [Rob94]. supersensibility [But84b]. Superstition [SS65]. Supertask [EN98, Lar97]. Supertasks [Lar05, Lar99]. Supervenience [Cur84, Dar12, Yos12]. Suppes [Dor77]. Support [Coh66a, Coh66b, Coh67b, How84, HF85, How85, LW15, Lev79, Nic87, Riv94, Sho03, Wal07]. Supposedly [Lar04]. Surfaces [Pap05]. Surplus [NTW20]. Surprise [Lev79, Ols02b]. Survey [Mac99]. Survival [Amu99, Elt03, Tro93, Gay98]. Susan [Kus01]. Sustaining [Goo13]. Sutherland [Fai57]. Sutton [Hoe00]. SUTVA [Boe22]. Swann [Set90]. Swanson [Far67].

Swinburne [Grü05]. Switching [GSK13]. Symbolic [Gro09, JLZ05]. Symmetries [CB12b, Dew19, GW14, Hea09, Ram23, Rob08b, Kos00]. Symmetry [Bel05b, Ber55b, BB04, Bud97, Das16, Daw85, Fri15a, Hun23, Hut95a, Hug99]. Symplectic [Thé12]. Synchronisation [Dor68, Lep76, Pod79, Sto86]. Synchronisms [Fra80, Qui74, Rox75]. Syndrome [Cla53]. Syntactic [Bon98]. Synthesis [LB18, Plu05, Van88]. Synthetic [Bab22, Goo09]. System [Col00, Cur83, Dor19, PS86, Von50, Ada01]. Systematic [Sie10b]. Systematicity [But93]. Systematized [Bor16]. Systems [Cha02b, CB69, Cle70a, Fri97a, Fri04, HH93, Kro89, Mat17, Naj24, Pop55b, SGSC18, Woo14, Yos12, BV05]. Szasz [Wal56c].

T

[Car12, DiS99, Gar99, LM85, Mar00, Rus00]. T. [Kat84]. tacit [Dav89]. Tacking [Cha07b, SS20]. Tainted [Ale16]. take [Sch00b]. Takes [But92]. Taking [Lar03, Zyc89]. Talk [Sti81]. Taming [Pot96]. Task [Col08, VC24, Van94b]. Taxonomy [Hul65a, Hul65b]. Taylor [C.56a, C.56c]. Teaches [Cla07]. Teaching [Mac93, Wig61]. Technical [Dur02, VH03]. Technically [Lai84, Mor86]. Technological [Dur02]. Technologically [MS07]. Technology [Koe04, Lel93, Ver05, Wis56a]. Teleological [Can64, Can65, Leh65, She73, Wri68]. Teleology [Ker87a, Mac69d, Sch59a]. Teleportation [Tim06]. Telescopic [Cha85a]. Television [Tuc58]. Tell [DTW17, FS94]. Tells [Fra88a]. Temperature [Gil09b, Cha04]. Temporal [BY06, Hoo93, Hut95a, Hut95b, Phi09, Wil50, Le 07]. Temporality [HL11]. Tenet [Fri09b]. Tennant [HR95]. Tense [Tei99]. Terms [Ach63, Bla79, DEO19, HW18, Jab92, SS93, Sim85, Win67]. Terrell [Tom99]. Territory [Cla01]. Test [Bro93, FAB^{+}89, Gie72b, Gil90, Hut83, Sta09]. Testability [Fri72, Pop59a, Pop59b, SG73, Was78]. Testable [Max88]. Testimonial [Coh82]. Testimonies [O’N82]. Testimony [Hol98, Ols02a, Ols02b]. Testing [Ale58b, Aro89, Bub94, Gho06, Gra78, Gre11, MS06a, MR94, Spi73, Tro93, Dic01b]. Tests [De 82, Joh87, Joh88, Lai88, Pop60, Red74a]. Textbooks [Bla09]. Their [Cha03, Mil74b, Pop52, Ram23, BH94]. Theism [Pit08]. Theistic [Grü04]. them [Han05, Gly10a]. Themes [Put01, Hoo00, Hea99]. Themselves [Grü69a]. theological [Grü00]. Theorem [Fin64, Gre10, Mil18, Oka08, Sch09, Sle82, SS13, But92, Cav10, Cuf17, Goo69, Pla00]. Theorems [Cuf18, DE76, Goo63, Han71, KP99]. Theoretic [Ber23, Eva17, Hun81, Par11, Skl67]. Theoretical [Ach63, Ach65, Bra93, Cof14, Dem03, Dem11, Gar72a, LC15, MS06b, Miz23, New17, Pro86, Ryn83, SS65, Sch09, SS93, Sim85, Wil51, Win67]. Theories [Ack66, BD86c, BD86b, BV63, Bor54, BH87, BF11, BH94, Cal59, Car85, Car64, Col00, Cur86, Din67, Dor68, Elb93, Ell57, Ern75, Fei22, Fle51, FS94, Fra21, FV11, Gre58a, Gre58b, Gre63b, Gre64, Hac71a, Hal51, Hes52, Hes58a, Hes58c, How95, Hut51c, Kar12, Kör67, Lav23, Lep69, Leu14, Lew06, Mil74b, Mor88, Mus74, Oka02, Pap76, Pik54, Pik55, Pit11, Pos60b, Raw51c, Red78, Res13, RS96, Set72, SS93, Shi84, Sie10b, SG73, Sim83, Sim85, Skl00, Smi98, Spe65, Sta09, Van63, Wac87, Was78, Whi51a, Whi51h, Why51c, Why51d, Wil73a, CM99, Rue11, DiS99, Mar00, Wal14]. Theorists [Bel95, Lan65]. Theorizing [Irv16]. Theory [Ahm05, Ahm14, Ale58b, AGTZ08, AGTZ11, And10, Azz04, Bac95, Bak09b, Bar72b, Bar13, BF23, Bel81, Bel05a, Bet50, Bet51c, Bet51d, Bha86, Blo70, Boh62, Bon92, Bra17, Bro93, Buc59, Buc65, Bun68a, Cav10, Cha79c, Cha81, Cha83, Cha64, Chi87, Chr10, CH01, Coh87, CEH21, Coo03a, CL19, Cro08, CK89, CT23, Dau88, Dav05, DEO19, DiF23, Din50b, Din64a, Dow08, ER96, Edw79, Far67, Fei54, Fle16, Fra91, FAB^{+}89, Fra20, Fri04, FA14, Gal14, Gar72b, Gar08, GS08a, Gia06b, Gil72b, Gil91b, Gil05, Gil92b, Giu88, Gly04, Goo59b, Goo10, Gra78, Gre11, Gri08a, Gri07, Grü76b, Grü86, Haa76a, Hal79c, Hal79d, Han05, HR95, Har13, Hea12, Hea15, Hor94, How85, Hue09, Hug00, Hug04, Hut83, HZ13, Jef55, Job80, Kre53b, Kre60b, Kri84]. Theory [LP18, Lan65, Leb89, Lee94, Lin93, Loe98, Lup18, LS94, Mar60, Mar57, Mar89a, Mat05, Max88, May77b, May77a, McA86, Mcl90, Mic94, Mił23, Mil74c, Mil18, Mor94, Mor00, Mül05b, Mun53, Mun54a, Mun89, Nol87, Nug85, Nun93, Nyh88, Ott94, PG13, Pet56, Pir52b, Pir53, Pit11, Pod79, Pos62a, Pri94, Pro68, Puc89, Red74a, RT92, Ree88, Ric16, Riv94, Roh88, Rue05, SV11, Sab54, Sac64, Sac76, Sam75, San98, San80, Sär68, Sca10, Sch74a, Sch08b, Sch53, Sch54b, Sha03, Siz00, Smy56, Spe66a, Spe66b, Spi73, Sta09, SW18, Tan89, Tel69, Ten94, Ten97, Tuc58, Vai15, Van08b, Ves20, Von50, Wer19a, Whi72, Why51a, WMR^{+}22, Wil78b, WS23, Wis63a, Wit61a, WY72, Yos12, dR96a, BD99, Hau99, Hun06, Joy99, Lan98, Mor99]. theory [Pri04, Sob11, Wal12, Woo03, Zah83a, BD86a, Byr00, Dob61, Dup13, Eel00, Kre60a, New73, Bel01, HH92]. Theory-Assessment [McA86]. Theory-Construction [Ale58b]. Theory-Dependent [Gal14]. Theory-Laden [Bro93, FAB^{+}89]. Theory-Testing [Ale58b]. Therapy [Erw82, Gre81, Gre84]. There [Dic95, FV11, Hes62, Nor09, Pin16, Sam18, Sch52a, Sch52b, Sko14, Sko15, Sud73, Tri94, Whi53b, Wor07, Zem87, Leh11, Max68].

Thermodynamic [Alb94, Nee13]. Thermodynamics [Den53, Dun11, Nor02, HS22, Rob23]. Theses [Gol58, Gol59, Wat59c]. Thesis [Ari84, But09, Cot03, Gar73a, McK14, NR13, Pic11, Row93, SW19a, Sol75, RS07]. Thick [Psi04]. Thin [Psi04]. Things [Bla20, FV11, Gly04, Mel80, PR95, Gel59, Woo03, Qui61]. Think [Bun56a, Bun56b, Fri02a, Mac05]. Thinking [Car04a, Gia94, Gia06a, Puc67, Pum53, Pum54, Car04b]. Third [Coh80b, Nii98, Pop58a, Wat59d]. Thomas [Bro02, HH92, O’M14]. Thompson [Cla66]. Thought [Gen98, Hea99, Hut53a, Mac05, Res09, Row94, Sha59, Ste03b, Car04b, PP78, Gia06a, She06, BG16b, Woo00b]. Thoughts [Buc51, Mus71, Sch59a, Why63, Zah81, ZG21]. Thousand [Hul65a, Hul65b]. Thread [Che91]. Three [Coh16, Dem11, Far16a, Fis24, HR22, Leu12, Max93b, Moo78, She85, Whi55e, Zha13]. Threshold [Van08b]. Tienson [Gar99]. Tiles [Roc86]. Tim [Bir01, Hey16b, Ney11]. Time [Abr58, AG09, Bar82, BE21, Bel05b, Ber68, Bla52, Bla53, Bun58a, Cal95, Cha57, Cha58, CBC^{+}22, Chr81, Den53, Dob51a, Dob51b, Dob53, Dob56, Dob58a, Far20, Grü54, Grü69b, GT16, Hea02, Hin88, Hoo71, Hoo03, Hor75, Hut93, Jam23, Mun53, Mun54a, Nar65, Nee10, Nor02, Phi09, Pla60, Pri02, Rox77, Sav94, Sav96, Sha59, Smi97, Smy58a, Ste03a, Swi80a, Tal16, Tei99, Thé12, Vie20, Wei04, Whi54b, Why52b, Why53a, Why53b, Why54b, You86, Goo59a, KP99, Le 03, Le 07, Pla00, Sac72, Tor84, Bri06, HS22, Sac93, Smi00b, Haw99, Grü53a]. Time-Asymmetric [CBC^{+}22]. Time-Asymmetry [Nor02]. Time-reversible [Hut93]. times [Bel05a, Gal00b, Mül05b]. Tipler [Cra88]. Titelbaum [Wil14a]. Tolerance [GC19]. Tollens [Wag04]. Too [BFHS02, Pri89, VC24]. Topologies [Hea88]. Topology [Fle16, Gro85a]. Topos [Eva17, Mcl90]. Tortoise [Bar62]. Total [Goo67b]. Toy [Ngu21, RHH18]. Traces [Hoe00]. Tracking [Ech16, Psi04, vFH17]. Tractability [Col15]. Tractatus [Pro59]. Trading [O’C17]. Tradition [Har59a]. Traditional [Mic94]. Traditions [Van94a]. Training [Bun86]. Traits [Ram13, Mat99]. Transcendent [Sie10b]. Transcendental [Loc03]. Transfer [Suz85]. Transfiguration [Roz69]. Transfinite [AB10, Gru66]. Transformation [Bor61, Kes60]. Transformational [D’A76]. Transformations [BB04, Ros78]. Transforming [Mul91]. Transiency [Fer70]. Transiently [Grü69a]. Transit [Why53b]. Transitional [Buc16]. Transitions [GSK13]. Transitivity [Fra80, Lep76, Qui74, Sho03]. Translation [Nis70, Pea83, She73]. Translations [Pan63]. Transplantation [Mun94]. Transport [Pod79]. Travel [Ber68, Smi97]. Travels [Le 03, Bri06]. Treat [Bla09]. Treatment [Lan06, Pla63, Str01, Str05]. Tree [Nea95b, Vel10]. Trials [Ful19]. Tristram [Sma86]. Trouble [Pos73]. True [AB16, Bla78, BFHS02, Jon82a, NS20]. Truesdell [Bun86]. Trust [IK19, Wil13a]. Trustworthiness [Mor93b]. Truth [BJ04, Bel01, Bet16, But89, DF89, Fra88a, Gel75, Haa76a, Hoo00, Jen84, LR07, Mac94b, Mag03, Raa05, Ram07a, RS96, Smi98, Sor07, Pri01, Put01]. Truthlikeness [GL23, Nii79, Pea83]. Truths [SF13, Mit09]. Trying [Ewe11, Lee82]. Tswett [LE22]. Tugby [McK16]. Tuning [Kop05, Mon06, Wei05]. Turetzky [Smi00b]. Turing [Bea89, But09, Cla72, Cot03, Gil90, Pic11]. Turn [Gri96]. Turning [Bai83]. Twelve [Coh86b]. twentieth [Gin84, Rai86]. twentieth-century [Gin84, Rai86]. Twice [Das16]. Twining [Tei13]. Twins [Kes85]. Two [Bar72b, Bet24, BV63, But09, Car64, Car02b, Fey68, Gar72b, Gol58, Gol59, Gre58a, Gre58b, Gre63b, Gre64, Gro85a, Hul65a, Hul65b, Jon82a, KGB^{+}22, Kui83, LB13, Max93c, Mea10, MG65, Mil85a, Mun53, Mun54a, O’N89, Pop55b, Puc89, Ram13, Sam75, San08, Spr18, ST22, Wat59c, Car99a, Mel99b, Zah83a]. Two-Dimensional [Mun53, Mun54a]. Two-level [Sam75]. Two-Slit [Mil85a]. Two-Stage [Jon82a]. Tycho [Fie86]. Type [Kri18, LP18, Pur23, GPW05]. Type-Level [Kri18]. Types [Pap76, Skl67, Wat52a]. Typical [WMR^{+}22]. Typicality [WMR^{+}22].

Ubiquity [Fri02a]. Uebel [Obe99]. Ugly [Sch22b]. Ullin [Bir01]. Ultimate [Vic21]. Un-debunking [Gła23]. Uncertain [BM10, CFM08, Hub08, TP19, TP21]. Uncertainties [FRZ12]. Uncertainty [Bra93, Bro20, Lev85, SW08a, Sch79a, SC18, Tap11, Wat55c, Wer19b, BM10]. Unconceived [Egg16, Enf08, Sta06a, Sta06b]. Unconscious [Gre51, MK56]. Undecidability [Kum69]. Undefeated [Sam80]. Underdetermination [Kie01, Kov19, Mag05, Pit11]. Underdetermined [Tuc11]. Undermines [Kle02]. Undermining [Wei15]. Understand [SW19b, Tal15, Wei13]. Understanding [Bel23, Bel98, Bir22, Cha21, Cur21, Del21, Fei18, GG86, Goo09, Gou23, Gre10, Gri06, Hal01, Hun23, Kha13b, Mac91, Mor12, New17, RHH18, Ste95, Ste00, Sul22, Wer10, Str03, Ver99, Fri16, Pri01, Til00]. Unexpected [Hol88]. Unfashionable [Kus01]. Unflattering [Buc23]. Unification [CH17, Gro85a, Kap58, Kar12, Kör64, Leu14, Mor90, Mor06, Nug85, Plu05, Red84, Rue05]. Unificatory [Nat17]. Unified [FS94]. Uniform [Rox77]. Uniformity [Gra74, Row93]. unique [Ste12b, Lew14]. Uniqueness [Giu01]. Unitary [Le 52].

Units [Dew24]. Unity [Gla24, Lan59, Mor06, Wei11]. Universal [Ell63, God69, Sni67, Ste97]. Universalia [Van94c]. Universality [Kne61, Bat00]. Universals [San80]. Universe [Abr54, CBC^{+}22, Cra92, Cra93, Dav54b, Dav56, EL72, Gra55, Hut55a, LE70, Mac71b, Ney14, Öpi54, Poo55, Rob08a, Sch54a, Sch11a, Scr54, Set89, Sma00, Whi54a, Whi55a, Dor05, Car12]. Universities [May60]. University [And86, Bai86, BD86a, Bri06, Coh86a, Fie86, Gia06b, Gib86, Hog86, Kör86, Rai86, Roc86, Rub86, Saf86, Sco61, Tau06, PP78]. Unmixing [ZG21]. Unobservability [Cha03]. Unobservable [Cha03, Mas07, Cha07a, Cho11]. Unpicturability [Pau71]. Unpredictability [Eag05, Sch74c, Sch76, Wer09]. Unpredictable [Mar89b]. Unravelling [Puc78]. Unreal [Wor84]. Unreasonable [Mol23]. Unsavory [Ber23]. Unscathed [Set72]. Unsimple [Mit09, SF13]. Unsought [Van94a]. Unsuccessful [Hau94]. Untidy [KT93]. untitled [Act57a, Ale59b, Apt84a, Bar60a, Bra65a, Cal82a, Car67a, Cou53a, Cou62a, Cro52a, Cro52b, Dav86, Din54, Dun62a, Ecc54, Ecc55a, Ell66, Fra88b, Fre84, Gil96b, Hal79a, Hal86, Hod80a, Hut51a, Hut53c, Jar79a, Jef54, Kne65a, Kör59a, Lau66, Llo66a, Mas54, Mil86b, Mor97, Nee66, Nol80a, Pap80a, Rav60a, Ric52a, Rus55a, Spi60a, Spi60b, Sti66, Wat54, Whi51c, Whi51d, Whi58a, Why52c, Wis54b, Wis57a, Woo58a, You66, YHF68]. Untruth [SV11]. Update [Sch15]. Updating [CMS22, Wei15, vFH17]. upon [Hal55b, Hut99]. Urbach [How97b, Dea76, Tiz76]. Ursu [Fie86]. Usage [Rus53, Ste16b]. Use [Bla79, Col66, Mac62, SW18]. Use-Novelty [SW18]. Users [Fri15b]. Uses [Eas14, Kre56, Mcl90, Swa22]. Using [Ell88, Gly10a, Hog04, LaC22, Oka16, Tal15, Wei13]. Utilising [Tre04]. Utility [Mic66, Pik54, Pik55, Wei86]. UV [CL19]. UV-Completion [CL19].

v [Gar72a]. Vacuity [PR95]. Vacuous [Sch87]. Vagueness [FC18, Sai95, Swi69]. Vaio [Str09]. Valence [Car23]. Valid [Urb78]. Validation [Kra09, Mas05]. Validity [Cle70a]. Value [Car23, Cla16, Cur23, Das23, Fel23, JL22, Jon82a, Kla64, Lea68, Mee64, Odd97, Ram90, Ros55, Sta20, Til00, Wor88, Coh86a]. Value-Judgments [Mee64, Ros55]. Values [Ell88, Pet22, Til00, Tum07, Vol91]. Values-Based [Ell88]. Variable [Cle79]. Variables [Bha86, Bub68, Edw79, Fei15, Fis79, FH16, Men54, Shi84, Sob98]. Variance [Gie70, Lep69, Mac71a, Mar71, Mar72, For99]. Variant [Kre53b]. Variation [Gro09, Kru97, JLZ05]. Variations [Ste12a]. Varieties [Bir20, Gil00b, GS08b, OL23, Hea86]. Vasilaki [Tei13]. Vectors [BP89]. Vegetative [Kle17, SB10]. Velocity [NS84, NS87, Nor86, Øhr86]. Venetian [Sob01]. Venus [Cha85a]. Verbal [Col07]. Verisimilar [BH87]. Verisimilitude [BH91, Har74, Jon73, McC23, Mee04, Mil74c, Mil76b, Nii82, Nii98, Odd81, Odd90, Sch22b, Tic74, Tic76, Urb83, Vas21a, Vol95]. Verismilitude [Pop76]. Verlag [Bun86]. Vermaas [Hal01, Pre03]. Versions [But09]. Verstehen [Mar69]. Versus [Aga57, Aga59a, Bla55, BM93, HS04, Jon73, Max93a, Max93c, Max93b, Sch74a, Set77a, Sie10b, Tri88, Zah87, AB16, Coh80a, DW17, Hoe98, Hol86, Llo67, Loe98, Mus74, Saa11, Set89, Set90]. Very [Wat89]. via [DTW17, LaC22]. Viability [Bin17]. Vicissitudes [Wei92]. Vickers [Fri16]. Victoria [Sco61]. Vienna [Obe98, Ueb96, Gat01]. View [Aga58a, Cal98b, Gre17, Gro85b, Lev79, Mac60a, May96, Poo55, Ric19, Sha84, Smi07b, Ste01]. Views [Bon55b, Das58, Dem11, Hal55b, Puc77, San08]. viii [Jen86, Kör86]. Vincent [Pri09]. Vinci [McC86, Pir52b]. Vindication [Hun70]. vindicatus [Luc69a]. Violate [Elb93]. Virtual [Pet96]. Virtues [Gly15, Miz23, Pap94b, Sti94]. virtuous [Nol99]. Vision [AH14, Dav18, Fra91, Pen21, Pir53, vEDS06]. Visual [Gia94, Gia06a, Car04b, Ech16]. Vol [Rai86]. Volterra [KL17]. Volumes [Mel70]. vs [Cra88, Elt03, Fer70, Til92, Wet94].

W [Sch59b, Ste00]. W. [May61]. Waiting [San19]. Wajsberg [Zah91]. Wallace [Vai15, BM10, Kle85, SW08b, Tap08]. Ward [Tom99]. Wars [Koe04]. Was [Cur76, Sle88, Sor86, Lev14]. Wason [VC24]. Watch [Dob57, Kap57b]. Water [Mik04, Nyr21]. Waterman [Str05]. Watkins [Wor99, Dup00, Gol59, O’H86, Ree78]. Watson [Sch69a]. Wave [Bar11, Cal15, CT16, Gao22, Liu94, Max88, Red77, She85, Squ89, vEDS06, Bic98, NA13, Bon00]. Wave/Particle [Max88, Squ89]. wavefunction [CM99]. WavePacket [Kri88]. Waves [Ach90, ET97, Lew07]. Way [HN14, NS84, NS87, Noo97, Nor86, Øhr86, Ric19, Why55b, HM99]. Weak [HN14]. Weber [AGTZ08]. Weight [Coh86b, Gil90, Ram90, Mee99]. Weisberg [Tal15]. Welfare [Sin00]. Well [Ale18, BN94, Her23]. Well-Being [Ale18, Her23]. Wellington [Sco61]. went [LM99]. Wesley [Cam00]. Western [Jen89, Tri88]. Weyl [H56, Sie10b]. Wheel [Haa78]. Wheeler [Lee94, Lee95, Pri94, Wil12]. Whence [Byr07, Fra21]. Where [Tah21, LM99]. Whewell [Hea53, Hea54, Wet94]. Which [LU09, Pra24, Sud73, Wal06a]. White [Fre88, Goo67c, Goo68b, Hem67, Sim91, Dod13, Set72]. Whitehead [Grü53b]. Whither [But96]. Whitrow [Bel79a, Cra79, Pop78]. Whittaker [W56, Bor54]. Who [McC84, Wei07]. Whom [Hud01]. Why-Questions [San98]. Whyte [Whi73]. Wide [Lee21]. Wideawake [Haa91]. Wigan [Puc89]. Wightman [May61, Wal56c]. Wiley [Hog86]. Wilks [Urb77]. Will [Dew73, EL72, Goo71a, LE70, Pla60, Tha70, FS94]. William [Bon01, Tei13, Hal15, Hea53, Hea54, Spr11, Wet94]. Williams [Aga78, Ber78, Van88]. Williamson [Gly09]. Wilson [Bra15, Dup96]. Winding [Dob57, Kap57b]. Wine [Mik04]. Wine/Water [Mik04]. Wisconsin [Mat99]. Wisdom [Das58]. Within [Ezr56, Ney11, Spi57, Wis63a, Bel01, Mau07, Was78]. Without [Bau64, Bau68, Car04a, Wat60a, Woo51c, Bet16, Bro20, Dew18, Har13, MP08, Par11, Pas15, Rob13, Sho00]. Wittgenstein [Gil82b, Kre58, Kre60a, Kre60b]. Wolfgang [Pro86]. Won’t [Oka01, Set70]. Woodger [Pan60, Pop81]. Woodward [Gly04, Car02b, Ste06a]. Word [Wis65]. Words [Car04a, Gel59, Qui61]. Work [Bel95, Cur21, LM85, McC57, Mor00, Ste95]. Works [Pan63]. World [Ale14, BL23, Bar01, BEL92, Bra11, Coh80b, Ech16, Fri02a, Hei99, Hof01, Kuk08, MB21, MK56, Puc78, Roe04, RS96, Sta90, Sud73, Tho88, Tho52, Tos16, Tro93, Tuo89, Cha12, Ryc99, Ste03b, Sty00, Wei13, She06, Fel01, Pol14, Rad00, Tal15, Sch07a]. Worlds [AGTZ11, But84a, CBR90, Mel00, Sti85, Chi98]. Worse [Hoe98, Pap96b]. Worthiness [Nyr21]. Wray [RW13]. Wright [Mor06]. Write [Dup13, Wil75b, Sob11]. Writing [Bra76]. Wrong [CBR90, Coo04, Gly10a, Lad00, Lei13, Sch08a, Hüt04, LM99]. Wrote [McC84].

X [Bro86, Hea86, Hog86, Rai86, D’A86, Gas75, MG17]. X-ray [MG17]. X-Rays [Gas75]. xi [Bai86, Bro86, Gib86]. Xian [Tha09]. xii [Hoo86, O’H86]. xiv [Mos86]. xv [Pro86]. xvi [Cap86]. xvii [And86]. xviii [Bun86, Coh86a]. xxii [Roc86].

Yale [Bai86]. Yang [Pit11]. Years [Gat01, Hul65a, Hul65b, MG17, Ein50a, Ein50b, Cou52a]. York [Den86]. Yourgrau [Pro86].

Z [Cal15]. Zahar [Mil78, Fey74, Fey80, Pod76, Tho92]. Zajonc [Hal99]. Zande [Jen89, Kei93]. Zeeman [Hea88]. Zeev [Wri86]. Zeno [Ang01, Ang02]. Zero [How73, Pop66c, Pri76]. Zhores [Jos10]. Zombie [Cla07].

References

A:1956:OMM

[A.56a] H. B. A. Obituary: Margaret Macdonald. British Journal for the Philosophy of Science, VII(26):184a–184, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                                   tronic). URL http://bjps.
                                                   oxfordjournals.org/content/
                                                   VII/26/184-a.full.pdf+html;
                                                    http://www.jstor.org/
                                                   stable/685710.

A:1956:PSC

[A.56b] H. B. A. Philosophy of science conference. British Journal for the Philosophy of Science, VII(26):184b–184, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                                   tronic). URL http://bjps.
                                                   oxfordjournals.org/content/
                                                   VII/26/184-b.full.pdf+html.

Adams:1987:PSC

[AA87]  Ernest  W.

Adams and Williams Y. Adams. Purpose and scientific concept formation. British Journal for the Philosophy of Science, 38(4): 419–440, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/419.full.pdf+html;
http://www.jstor.org/
stable/687352.

Aitken:2010:NPP

[AB10] Wayne Aitken and Jeffrey A.

Barrett. A note on the physical possibility of transfinite computation. British Journal for the Philosophy of Science, 61(4): 867–874, December 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/4/867.full.pdf+html.

Azzouni:2016:TNR

[AB16] Jody Azzouni and Otávio Bueno.

True nominalism: Referring versus coding. British Journal for the Philosophy of Science, 67(3):781–816, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/781.

Andersen:2024:WPA

[ABB^{+}24] Line Edslev Andersen, Rosemary Bailey, Udit Bhatia, Joanna Burch-Brown, Liam Kofi Bright, Jim Brooke, Peter Cameron, Catrin Campbell-Moore, Joe Cornelli, Stephen Crowley, Haixin Dang, Rachel Fraser, Isobel Falconer, Adam Ferner, Monica Gamez, Mikkel Gerken, Katherine Hawley, Matthias Heyman, Klemens Kappel, Ethan Landes, Hélène Landemore, Benedikt Loewe, Dan Lunt, Ursula Martin, Chris Meyns, Louise Millard, Dann Mitchell, Samir Okasha, Andrew Peet, Richard Pettigrew, Jeroen de Ridder, Robert Ross, Pierre Saint-Germier, Mona Simion, Anya Skatova, Fenner Tanswell, Daniel Whiting, Alan Wilson, K. Brad Wray, Dilara Yesilova, and Joshua Habgood-Coote. What’s the point of authors? British Journal for the Philosophy of Science, 75(2): 487–517, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Andersen:2006:CSS

[ABC06] Hanne Andersen, Peter Barker, and Xiang Chen. The cognitive structure of scientific revolutions. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85575-6. xvii + 199 pp. LCCN Q174.8 .A53

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0633/2005021713-d.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2005021713-b.html;   http:/
/www.loc.gov/catdir/toc/

ecip0516/2005021713.html.

Aboites:2002:SRA

[Abo02] Vicente Aboites. Some remarks about Newton’s demonstrations in optics: Newton’s missing experiment. British Journal for the Philosophy of Science, 53(3): 455–458, September 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/455.full.pdf+html;
http://www.jstor.org/
stable/3541890.

Abramenko:1954:AU

[Abr54] B. Abramenko. The age of the universe. British Journal for the Philosophy of Science, V(19):237–252, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/237.full.pdf+html;
http://www.jstor.org/
stable/685719.

Abramenko:1958:DCP

[Abr58] B. Abramenko. On dimensionality and continuity of physical space and time. British Journal for the Philosophy of Science, IX (34):89–109, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/34/89.full.pdf+html;
http://www.jstor.org/
stable/685146.

Achinstein:1962:IH

[Ach62] Peter Achinstein. The identity hypothesis. British Journal for the Philosophy of Science, XIII(50):167–171, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/50/167.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685970.

Achinstein:1963:TTP

[Ach63] Peter Achinstein. Theoretical terms and partial interpretation. British Journal for the Philosophy of Science, XIV(54):89–105, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIV/54/89.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685428.

Achinstein:1965:TM

[Ach65] Peter Achinstein. Theoretical models. British Journal for the Philosophy of Science, XVI(62): 102–120, August 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XVI/62/102.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686152.

Achinstein:1968:R

[Ach68]  Peter Achinstein.  Reviews.

British Journal for the Philosophy of Science, 19(2):159–168, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/159.full.pdf+
html;   http://www.jstor.
org/stable/686796.

Achinstein:1983:NE

[Ach83] Peter Achinstein. The nature of explanation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-503215-2. ix + 385 pp. LCCN BD237 .A25 1983.

Achinstein:1990:HPW

[Ach90]  Peter Achinstein.  Hypothe-

ses, probability, and waves. British Journal for the Philosophy of Science, 41(1):73–102, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/1/73.full.pdf+
html;   http://www.jstor.
org/stable/688004.

Achinstein:1995:EEC

[Ach95] Peter Achinstein. Are empirical evidence claims a priori? British Journal for the Philosophy of Science, 46(4):447–473, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/4/447.full.pdf+
html;   http://www.jstor.
org/stable/687893.

Ackermann:1965:R

[Ack65] Robert Ackermann. Reviews. British Journal for the Philosophy of Science, XVI(62): 161–163, August 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XVI/62/161.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686156.

Ackermann:1966:CMT

[Ack66] Robert Ackermann. Confirmatory models of theories. British Journal for the Philosophy of Science, XVI(64):312–326, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/64/312.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686832.

Ackermann:1989:NE

[Ack89] Robert Ackermann. The new experimentalism. British Journal for the Philosophy of Science, 40(2):185–190, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/2/185.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687509.

Acton:1956:R

[Act56] H. B. Acton. Reviews. British Journal for the Philosophy of Science, VI(24):338–340, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/338.full.pdf+
html;   http://www.jstor.
org/stable/685416.

Acton:1957:RU

[Act57a] H. B. Acton. Review: [untitled].

British Journal for the Philosophy of Science, VIII(29):84, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/685392.

Acton:1957:R

[Act57b] H. B. Acton. Reviews. British Journal for the Philosophy of Science, VIII(29):84a–84, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
VIII/29/84-a.full.pdf+html.

Adams:1994:MS

[Ada94]  Ernest  W.  Adams.

On the method of superposition. British Journal for the Philosophy of Science, 45(2):693–708, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/693.full.pdf+
html;   http://www.jstor.
org/stable/687691.

Adams:2001:BRB

[Ada01] Frederick Adams. Book review: Dynamics in Action: Intentional Behavior as a Complex System, Alicia Juarrero. British Journal for the Philosophy of Science, 52(3):635–640, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/635.full.pdf+html;
http://www.jstor.org/
stable/3541935.

Adkins:1951:H

[Adk51] B. M. Adkins. The homeostat. British Journal for the Philosophy of Science, II(7): 248, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/248.full.pdf+html;
http://www.jstor.org/
stable/685295.

Adkins:1952:DD

[Adk52] B. M. Adkins. The dictum of Descartes. British Journal for the Philosophy of Science, III(11):259–260, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/259.full.pdf+html;
http://www.jstor.org/
stable/685269.

Al-Daffa:1984:SES

[ADS84] Ali Abd Al-Daffa and John J.

Stroyls. Studies in the exact sciences in Medieval Islam. University of Petroleum and Minerals, Dhahran, Saudi Arabia, 1984. ISBN 0-471-90320-5. x + 243 pp. LCCN Q127.A5 D34 1984. US$48.00 (U.S.).

Arntzenius:2009:TRC

[AG09] Frank Arntzenius and Hilary Greaves. Time reversal in classical electromagnetism. British Journal for the Philosophy of Science, 60(3):557–584, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/60/3/557.full.pdf+
html;   http://www.jstor.
org/stable/25592019.

Agassi:1957:DVG

[Aga57] J. Agassi. Duhem versus Galileo.

British Journal for the Philosophy of Science, VIII(31):237– 248, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/31/237.full.pdf+html;
http://www.jstor.org/
stable/685760.

Agassi:1958:HVC

[Aga58a] J. Agassi. A Hegelian view of complementarity. British Journal for the Philosophy of Science, IX(33):57–63, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
IX/33/57.full.pdf+html;
http://www.jstor.org/
stable/686103.

Agassi:1958:KHC

[Aga58b] J. Agassi. Koyré on the history of cosmology. British Journal for the Philosophy of Science, IX(35):234–245, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/35/234.full.pdf+
html;   http://www.jstor.
org/stable/685655.

Agassi:1959:CVI

[Aga59a] J. Agassi. Corroboration versus induction. British Journal for the Philosophy of Science, IX(36):311–317, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/311.full.pdf+html;
http://www.jstor.org/
stable/685191.

Agassi:1959:EAS

[Aga59b] J. Agassi. Epistemology as an aid to science: Comments on Dr Buchdahl’s paper. British Journal for the Philosophy of Science, X(38):135–146, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/38/135.full.pdf+html;
http://www.jstor.org/
stable/686004.

Agassi:1960:R

[Aga60] J. Agassi. Reviews. British Journal for the Philosophy of Science, XI(41):83–84, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/41/83.full.pdf+html;
http://www.jstor.org/
stable/685831.

Agassi:1963:BMM

[Aga63] Joseph Agassi. Between micro and macro. British Journal for the Philosophy of Science, XIV(53):26–31, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/53/26.full.pdf+html;
http://www.jstor.org/
stable/685348.

Agassi:1978:BRW

[Aga78] Joseph Agassi. Book review:

Williams Dodges, Agassi’s Criticism. British Journal for the Philosophy of Science, 29(3): 248–252, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/248.full.pdf+html;
http://www.jstor.org/
stable/687012.

Agassi:1987:R

[Aga87] J. Agassi. Reviews. British Journal for the Philosophy of Science, 38(3):414–418, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/3/414.full.pdf+
html;   http://www.jstor.
org/stable/687501.

Agassi:1990:ISI

[Aga90] Joseph Agassi. Induction and stochastic independence. British Journal for the Philosophy of Science, 41(1):141–142, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/1/141.full.pdf+
html;   http://www.jstor.
org/stable/688009.

Allori:2008:CSB

[AGTZ08] Valia Allori, Sheldon Goldstein, Roderich Tumulka, and Nino Zanghì. On the common structure of Bohmian mechanics and the Ghirardi–Rimini–Weber theory: Dedicated to GianCarlo Ghirardi on the occasion of his 70th birthday. British Journal for the Philosophy of Science, 59(3):353–389, September 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/353.full.pdf+html;
http://www.jstor.org/
stable/40072290.

Allori:2011:MWS

[AGTZ11] Valia Allori, Sheldon Goldstein, Roderich Tumulka, and Nino Zanghì. Many worlds and Schrödinger’s first quantum theory. British Journal for the Philosophy of Science, 62(1):1–27, March 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/1.full.pdf+html.

Allori:2014:PPO

[AGTZ14] Valia Allori, Sheldon Goldstein, Roderich Tumulka, and Nino Zanghì. Predictions and primitive ontology in quantum foundations: a study of examples. British Journal for the Philosophy of Science, 65(2):323–352, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/65/2/323.

Arntzenius:2003:WWK

[AH03] Frank Arntzenius and Ned Hall.

On what we know about chance. British Journal for the Philosophy of Science, 54(2):171–179, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/171.full.pdf+
html;   http://www.jstor.
org/stable/3541963.

Akins:2014:ECM

[AH14] Kathleen A. Akins and Martin Hahn. Editor’s choice: More than mere colouring: The role of spectral information in human vision. British Journal for the Philosophy of Science, 65(1): 125–171, January 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/125.full.pdf+html.

Ahmed:2005:EDT

[Ahm05] Arif Ahmed. Evidential decision theory and medical Newcomb problems. British Journal for the Philosophy of Science, 56(2):191–198, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/2/191.full.pdf+html;
http://www.jstor.org/
stable/3541699.

Ahmed:2014:CDT

[Ahm14] Arif Ahmed. Causal decision theory and the fixity of the past. British Journal for the Philosophy of Science, 65(4): 665–685, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/665.

Aronson:1995:RRH

[AHW95] Jerrold L. Aronson, Rom Harré, and Eileen Cornell Way. Realism rescued: how scientific progress is possible. Open Court Publishing Co., La Salle, IL, USA, 1995. ISBN 0-8126-9288-8, 0-8126-9289-6 (paperback). viii + 213 pp. LCCN Q175.32.R42 A76 1995.

Ainsworth:2009:NO

[Ain09] Peter M. Ainsworth. Newman’s objection. British Journal for the Philosophy of Science, 60 (1):135–171, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/135.full.pdf+html;
http://www.jstor.org/
stable/25591991.

Allis:1991:SPI

[AK91] Victor Allis and Teunis Koetsier. On some paradoxes of the infinite. British Journal for the Philosophy of Science, 42 (2):187–194, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/187.full.pdf+html;
http://www.jstor.org/
stable/688052.

Allis:1995:SPI

[AK95] Victor Allis and Teun Koetsier.

On some paradoxes of the infinite II. British Journal for the Philosophy of Science, 46(2):235–247, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/235.full.pdf+
html;   http://www.jstor.
org/stable/687814.

Akeroyd:1986:CFL

[Ake86] F. Michael Akeroyd. A challenge to the followers of Lakatos. British Journal for the Philosophy of Science, 37(3):359– 362, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687294.

Akeroyd:1988:RPE

[Ake88] F. M. Akeroyd. Research programmes and empirical results. British Journal for the Philosophy of Science, 39(1):51–58, March 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/39/1/51.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687113.

Akeroyd:1990:CLR

[Ake90a] F. Michael Akeroyd. The challenge to Lakatos restated. British Journal for the Philosophy of Science, 41(3):437–439, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/41/3/437.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687831.

Akeroyd:1990:OMG

[Ake90b] F. Michael Akeroyd. An oscillatory model of the growth of scientific knowledge. British Journal for the Philosophy of Science, 41(3):407–414, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/3/407.full.pdf+
html;   http://www.jstor.
org/stable/687828.

Akeroyd:1991:PEG

[Ake91] F. M. Akeroyd. A practical example of Grue. British Journal for the Philosophy of Science, 42(4):535–539, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/4/535.full.pdf+html;
http://www.jstor.org/
stable/687707.

Akeroyd:1993:LPS

[Ake93] F. Michael Akeroyd. Laudan’s problem solving model. British Journal for the Philosophy of Science, 44(4):785–788, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/4/785.full.pdf+
html;   http://www.jstor.
org/stable/688045.

Akkerman:1994:SAC

[Akk94] Abraham Akkerman. Sameness of age cohorts in the mathematics of population growth. British Journal for the Philosophy of Science, 45(2):679–691, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/2/679.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687690.

Ariew:2004:CF

[AL04] André Ariew and R. C. Lewontin. The confusions of fitness. British Journal for the Philosophy of Science, 55(2):347–363, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/2/347.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541695.

Albertson:1962:SNQ

[Alb62] James Albertson. The statistical nature of quantum mechanics. British Journal for the Philosophy of Science, XIII(51): 229–233, November 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/51/229.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685840.

Albert:1994:FQM

[Alb94] David Z. Albert. The foundations of quantum mechanics and the approach to thermodynamic equilibrium. British Journal for the Philosophy of Science, 45(2):669–677, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
45/2/669.full.pdf+html.

Albert:2002:RNP

[Alb02] Max Albert. Resolving Neyman’s Paradox. British Journal for the Philosophy of Science, 53(1):69–76, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/69.full.pdf+html;
http://www.jstor.org/
stable/3541641.

Alexander:1952:Ra

[Ale52a]  Peter Alexander.  Reviews.

British Journal for the Philosophy of Science, III(11):284– 286, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/284.full.pdf+html;
http://www.jstor.org/
stable/685281.

Alexander:1952:Rb

[Ale52b]  Peter Alexander.  Reviews.

British Journal for the Philosophy of Science, III(11):287– 289, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/287.full.pdf+html;
http://www.jstor.org/
stable/685283.

Alexander:1954:R

[Ale54]  Peter Alexander.  Reviews.

British Journal for the Philosophy of Science, V(17):83–85, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/17/83.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685958.

Alexander:1957:R

[Ale57]  Peter Alexander.  Reviews.

British Journal for the Philosophy of Science, VIII(29):76–78, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/29/76.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685386.

Alexander:1958:PC

[Ale58a] H. G. Alexander. The paradoxes of confirmation. British Journal for the Philosophy of Science, IX(35):227–233, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/35/227.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685654.

Alexander:1958:TCT

[Ale58b]  Peter Alexander.  Theory-

construction and theory-testing. British Journal for the Philosophy of Science, IX(33):29–38, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/29.full.pdf+
html;   http://www.jstor.
org/stable/686098.

Alexander:1959:PCR

[Ale59a] H. G. Alexander. The paradoxes of confirmation — a reply to Dr Agassi. British Journal for the Philosophy of Science, X(39):229–234, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/229.full.pdf+html;
http://www.jstor.org/
stable/686057.

Alexander:1959:RU

[Ale59b] H. G. Alexander. Review: [untitled]. British Journal for the Philosophy of Science, X(39): 257–258, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686070.

Alexander:1959:R

[Ale59c] H. G. Alexander. Reviews.

British Journal for the Philosophy of Science, X(39):257b– 258, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/257-b.full.pdf+html.

Alexander:1960:R

[Ale60] P. Alexander. Reviews. British Journal for the Philosophy of Science, XI(43):255–257, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/43/255.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685316.

Alexander:1962:Ra

[Ale62a] H. G. Alexander. Reviews. British Journal for the Philosophy of Science, XII(48):348– 349, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XII/48/348.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685407.

Alexander:1962:Rb

[Ale62b] H. G. Alexander. Reviews. British Journal for the Philosophy of Science, XIII(49):82–85, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/49/82.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685234.

Alexander:1967:R

[Ale67]  Peter Alexander.  Reviews.

British Journal for the Philosophy of Science, 18(3):263– 264, November 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/3/263.full.pdf+html;
http://www.jstor.org/
stable/686603.

Alexander:2014:LSD

[Ale14] J. McKenzie Alexander. Learning to signal in a dynamic world. British Journal for the Philosophy of Science, 65(4): 797–820, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/797.

Alexandrova:2016:BRK

[Ale16] Anna Alexandrova. Book review: Kristin Shrader-Frechette, Tainted: How Philosophy of Science Can Expose Bad Science. British Journal for the Philosophy of Science, 67(3):901– 905, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/901.

Alexandrova:2018:CSW

[Ale18] Anna Alexandrova. Can the science of well-being be objective? British Journal for the Philosophy of Science, 69(2):421–445, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/421/2669793.

Allen:1992:MC

[All92] Colin Allen. Mental content. British Journal for the Philosophy of Science, 43(4):537–553, October 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/4/537.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687840.

Alliksaar:2001:MLR

[All01] Mark Alliksaar. Metabolism in a life: Reply to Boden. British Journal for the Philosophy of Science, 52(1):131–135, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/52/1/131.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541947.

Altham:1968:NGP

[Alt68] J. E. J. Altham. A note on Goodman’s Paradox. British Journal for the Philosophy of Science, 19(3):257, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        19/3/257.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686246.

Amundson:1998:R

[Amu98] Ron Amundson. Review. British Journal for the Philosophy of Science, 49(3):515–521, September 1998. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/515.full.pdf+
html;   http://www.jstor.
org/stable/688089.

Amundsom:1999:BRB

[Amu99] R. Amundsom. Book review:

Darwinism’s Struggle for Survival: Heredity and the Hypothesis of Natural Selection. J. Gayon. British Journal for the Philosophy of Science, 50(4): 761–767, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/761.full.pdf+html;
http://www.jstor.org/
stable/40072266.

Andersson:1978:R

[And78] Gunnar Andersson. Reviews.

British Journal for the Philosophy of Science, 29(4):389– 392, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/389.full.pdf+html;
http://www.jstor.org/
stable/687103.

Anderson:1986:GRM

[And86] James Anderson. Book review: Gregory, Richard [1984]: Mind in Science. Cambridge University Press. xvii + 641 pp. British Journal for the Philosophy of Science, 37(4):

525–529, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687381.

Andreasen:1998:NPR

[And98] Robin O. Andreasen. A new perspective on the race debate. British Journal for the Philosophy of Science, 49(2):199–225, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/2/199.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688112.

Andersen:2010:BRE

[And10] Hanne Andersen. Book review: Edwin H.-C. Hung, Beyond Kuhn. Scientific Explanation, Theory Structure, Incommensurability and Physical Necessity. British Journal for the Philosophy of Science, 61 (1):237–239, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/1/237.full.pdf+html.

Andersen:2018:CCC

[And18] Holly Andersen. Complements, not competitors: Causal and mathematical explanations. British Journal for the Philosophy of Science, 69(2):485–508, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/485/2669791.

Angel:2001:PMZ

[Ang01]  Leonard Angel.  A physi-

cal model of Zeno’s Dichotomy. British Journal for the Philosophy of Science, 52(2):347–358, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/2/347.full.pdf+
html;   http://www.jstor.
org/stable/3541900.

Angel:2002:ZAN

[Ang02] Leonard Angel. Zeno’s Arrow, Newton’s Mechanics, and Bell’s Inequalities. British Journal for the Philosophy of Science, 53(2):161–182, June 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/2/161.full.pdf+html;
http://www.jstor.org/
stable/3541762.

Angel:2005:EON

[Ang05] Leonard Angel. Evens and odds in Newtonian collision mechanics. British Journal for the Philosophy of Science, 56 (1):179–188, March 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/1/179.full.pdf+html;
http://www.jstor.org/
stable/3541856.

Anonymous:1986:IA

[Ano86a] Anonymous. Index of authors. British Journal for the Philosophy of Science, 37(2):vii–viii, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:1986:RP

[Ano86b] Anonymous. Recent publications. British Journal for the Philosophy of Science, 37(3): 385–393, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Anon:1999:E

[Ano99] Anon. Erratum. British Journal for the Philosophy of Science, 50(2):324, June 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/2/324.full.pdf+html.

Anonymous:2013:A

[Ano13] Anonymous. Acknowledgements. British Journal for the Philosophy of Science, 64(4):675– 679, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/4/675.full.pdf+html.

Anonymous:2021:E

[Ano21a] Anonymous. Editorial. British Journal for the Philosophy of Science, 71(4):1155–1156, ????

2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2021:FM

[Ano21b] Anonymous. Front matter.

British Journal for the Philosophy of Science, 72(1):??, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2022:FMa

[Ano22a] Anonymous. Front matter.

British Journal for the Philosophy of Science, 72(4):??, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2022:FMb

[Ano22b] Anonymous. Front matter.

British Journal for the Philosophy of Science, 73(1):??, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2022:FMc

[Ano22c] Anonymous. Front matter.

British Journal for the Philosophy of Science, 73(2):??, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2022:FMd

[Ano22d] Anonymous. Front matter.

British Journal for the Philosophy of Science, 73(3):??, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2023:FMa

[Ano23a] Anonymous. Front matter. British Journal for the Philosophy of Science, 73(4):??, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2023:FMb

[Ano23b] Anonymous. Front matter. British Journal for the Philosophy of Science, 74(1):??, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2023:FMc

[Ano23c] Anonymous. Front matter. British Journal for the Philosophy of Science, 74(2):??, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2023:FMd

[Ano23d] Anonymous. Front matter. British Journal for the Philosophy of Science, 74(3):??, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2023:FMe

[Ano23e] Anonymous. Front matter. British Journal for the Philosophy of Science, 74(4):??, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2024:FMa

[Ano24a] Anonymous. Front matter.

British Journal for the Philosophy of Science, 75(1):??, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Anonymous:2024:FMb

[Ano24b] Anonymous. Front matter.

British Journal for the Philosophy of Science, 75(2):??, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Apt:1984:RU

[Apt84a] A. J. Apt. Review: [untitled]. British Journal for the Philosophy of Science, 35(3):303– 307, September 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687482.

Apt:1984:R

[Apt84b] A. J. Apt. Reviews. British Journal for the Philosophy of Science, 35(3):303, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/303.full.pdf+html.

Aranyosi:2011:NAM

[Ara11] István Aranyosi. A new argument for mind-brain identity. British Journal for the Philosophy of Science, 62(3):489– 517, September 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/489.full.pdf+html.

Archibald:1958:R

[Arc58]  G. C. Archibald.  Reviews.

British Journal for the Philosophy of Science, IX(34):173–175, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/173.full.pdf+
html;   http://www.jstor.
org/stable/685161.

Archibald:1959:RSE

[Arc59] G. C. Archibald. Reviews: The state of economic science. British Journal for the Philosophy of Science, X(37):58–69, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/58.full.pdf+html;
http://www.jstor.org/
stable/685779.

Archibald:1967:RC

[Arc67] G. C. Archibald. Refutation or comparison? British Journal for the Philosophy of Science, 17(4):279–296, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/279.full.pdf+html;
http://www.jstor.org/
stable/686770.

Ariew:1984:DT

[Ari84] Roger Ariew. The Duhem thesis. British Journal for the Philosophy of Science, 35(4): 313–325, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/313.full.pdf+html;
http://www.jstor.org/
stable/687336.

Armitage:1958:R

[Arm58] A. Armitage. Reviews. British Journal for the Philosophy of Science, IX(33):72–73, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/72.full.pdf+
html;   http://www.jstor.
org/stable/686109.

Armitage:1959:Ra

[Arm59a] A. Armitage. Reviews. British Journal for the Philosophy of Science, X(38):162–163, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/162.full.pdf+
html;   http://www.jstor.
org/stable/686010.

Armitage:1959:Rb

[Arm59b] A. Armitage. Reviews. British Journal for the Philosophy of Science, X(39):254–255, November 1959. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/39/254.full.pdf+
html;   http://www.jstor.
org/stable/686067.

Armstrong:1970:R

[Arm70] D. M. Armstrong. Reviews. British Journal for the Philosophy of Science, 21(1):117– 119, February 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/1/117.full.pdf+html;
http://www.jstor.org/
stable/686476.

Armendt:1993:MES

[Arm93] Brad Armendt. Marinoff on evolutionarily stable strategies. British Journal for the Philosophy of Science, 44(4):789– 793, December 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/4/789.full.pdf+html;
http://www.jstor.org/
stable/688046.

Arntzenius:1990:CPS

[Arn90] Frank Arntzenius. Causal paradoxes in Special Relativity. British Journal for the Philosophy of Science, 41(2):223–243, June 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/41/2/223.full.pdf+
html;   http://www.jstor.
org/stable/687773.

Arntzenius:1994:SC

[Arn94] Frank Arntzenius. Spacelike connections. British Journal for the Philosophy of Science, 45 (1):201–217, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/201.full.pdf+html;
http://www.jstor.org/
stable/687968.

Aronson:1989:TCR

[Aro89] Jerrold L. Aronson. Testing for convergent realism. British Journal for the Philosophy of Science, 40(2):255–259, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/255.full.pdf+html;
http://www.jstor.org/
stable/687517.

Ariew:2015:ASE

[ARR15] André Ariew, Collin Rice, and Yasha Rohwer. Autonomousstatistical explanations and natural selection. British Journal for the Philosophy of Science, 66(3):635–658, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/635.

Arthur:1994:SRN

[Art94] Richard Arthur. Space and Relativity in Newton and Leibniz. British Journal for the Philosophy of Science, 45(1):219–240, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/1/219.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687969.

Ahmed:2019:DLN

[AS19] Arif Ahmed and Bernhard Salow. Don’t look now. British Journal for the Philosophy of Science, 70(2):327–350, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/2/327/4098120.

Ashby:1952:CMC

[Ash52] W. Ross Ashby. Can a mechanical chess-player outplay its designer? British Journal for the Philosophy of Science, III(9):44– 57, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/9/44.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685473.

Ashby:1955:R

[Ash55]  W. Ross Ashby.  Reviews.

British Journal for the Philosophy of Science, VI(23):255– 256, November 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/23/255.full.pdf+html;
http://www.jstor.org/
stable/685688.

Ashby:1957:PR

[Ash57a] W. Ross Ashby. Pavlov reconditioned. British Journal for the Philosophy of Science, VIII (31):249–252, November 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/31/249.full.pdf+html;
 http://www.jstor.org/
stable/685761.

Ashby:1957:Ra

[Ash57b] W. Ross Ashby. Reviews.

British Journal for the Philosophy of Science, VIII(29):80–81, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/80.full.
pdf+html;   http://www.
jstor.org/stable/685388.

Ashby:1957:Rb

[Ash57c]  W. Ross Ashby.  Reviews.

British Journal for the Philosophy of Science, VIII(29):83–84, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/83.full.
pdf+html;   http://www.
jstor.org/stable/685391.

Ashby:1959:R

[Ash59] W. Ross Ashby. Reviews. British Journal for the Philosophy of Science, X(39):258, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/39/258.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686071.

Ashby:1963:R

[Ash63]  W. Ross Ashby.  Reviews.

British Journal for the Philosophy of Science, XIII(52):341– 342, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/52/341.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685341.

Autzen:2020:MMF

[Aut20] Bengt Autzen. Musing on means: Fitness, expectation, and the principles of natural selection. British Journal for the Philosophy of Science, 71(1):373–389, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/1/373/4091080.

Avin:2019:CFE

[Avi19] Shahar Avin. Centralized funding and epistemic exploration. British Journal for the Philosophy of Science, 70(3):629– 656, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
629/4675327.

Avramides:2002:KOO

[Avr02] Anita Avramides. Knowing our own minds. British Journal for the Philosophy of Science, 53(3):465–471, September 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/465.full.pdf+html;
http://www.jstor.org/
stable/3541892.

Ayers:1970:R

[Aye70] M. R. Ayers. Reviews. British Journal for the Philosophy of Science, 21(2):215–216, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/215.full.pdf+
html;   http://www.jstor.
org/stable/686709.

Azzouni:2004:TOS

[Azz04] Jody Azzouni. Theory, observation and scientific realism. British Journal for the Philosophy of Science, 55(3):371– 392, September 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/3/371.full.pdf+html;
http://www.jstor.org/
stable/3541667.

Babcock:2022:SGP

[Bab22] Gunnar Babcock. Are synthetic genomes parts of a genetic lineage? British Journal for the Philosophy of Science, 72(4): 995–1011, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Backhouse:1995:EPE

[Bac95] Roger E. Backhouse. An empirical philosophy of economic theory. British Journal for the Philosophy of Science, 46 (1):111–121, March 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/1/111.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687635.

Baetu:2012:GPM

[Bae12] Tudor M. Baetu. Genomic programs as mechanism schemas: a non-reductionist interpretation. British Journal for the Philosophy of Science, 63(3):649– 671, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/3/649.full.pdf+html.

Baetu:2016:BPP

[Bae16] Tudor M. Baetu. The ‘big picture’: The problem of extrapolation in basic research. British Journal for the Philosophy of Science, 67(4):941–964, December 2016. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/4/941.

Baetu:2019:PCE

[Bae19] Tudor Baetu. On the possibility of crucial experiments in biology. British Journal for the Philosophy of Science, 70(2):407–429, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/70/2/407/4102128.

Baillie:1969:CMY

[Bai69] Patricia Baillie. That confirmation may yet be a probability. British Journal for the Philosophy of Science, 20(1):41–51, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/41.full.pdf+
html;   http://www.jstor.
org/stable/686717.

Baillie:1971:CPR

[Bai71] Patricia Baillie. Confirmation and probability: a reply to Settle. British Journal for the Philosophy of Science, 22(3): 285–286, August 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/3/285.full.pdf+html;
http://www.jstor.org/
stable/686748.

Baillie:1973:CDB

[Bai73] Patricia Baillie. Confirmation and the Dutch Book argument. British Journal for the Philosophy of Science, 24(4):393– 397, December 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        24/4/393.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686619.

Baird:1983:FPC

[Bai83] Davis Baird. The Fisher/Pearson chi-squared controversy: a turning point for inductive inference. British Journal for the Philosophy of Science, 34(2):105–118, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/2/105.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687444.

Baird:1986:RU

[Bai86]  Davis Baird.  Book review:

Skyrms, Brian [1984]: Pragmatics and Empiricism. Yale University Press. Pp. xi + 143. £16.95. (ISBN 0-300-03174-2). British Journal for the Philosophy of Science, 37(4):514– 516, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687377.

Baird:1987:EFA

[Bai87] Davis Baird. Exploratory factor analysis, instruments and the logic of discovery. British Journal for the Philosophy of Science, 38(3):319–337, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/319.full.pdf+html;
http://www.jstor.org/
stable/687488.

Baker:2003:QPE

[Bak03] Alan Baker. Quantitative parsimony and explanatory power. British Journal for the Philosophy of Science, 54(2):245–259, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/245.full.pdf+
html;   http://www.jstor.
org/stable/3541966.

Baker:2009:MES

[Bak09a] Alan Baker. Mathematical explanation in science. British Journal for the Philosophy of Science, 60(3):611–633, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/60/3/611.full.pdf+
html;   http://www.jstor.
org/stable/25592021.

Baker:2009:AFI

[Bak09b] David John Baker. Against field interpretations of quantum field theory. British Journal for the Philosophy of Science, 60(3):585–609, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/3/585.full.pdf+html;
http://www.jstor.org/
stable/25592020.

Baker:2015:BRC

[Bak15]  Alan Baker.  Book review:

Christopher Pincock, Mathematics and Scientific Representation. British Journal for the Philosophy of Science, 66(3):695– 699, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/695.

Balint:1957:CFG

[Bal57] Michael Balint. Criticism of Fairbairn’s generalisation about object-relations. British Journal for the Philosophy of Science, VII(28):323–324, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/323.full.pdf+html;
http://www.jstor.org/
stable/685112.

Baldwin:1971:R

[Bal71]  Thomas  Bald-

win. Reviews. British Journal for the Philosophy of Science, 22 (1):68–70, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/68.full.pdf+html;
http://www.jstor.org/
stable/686653.

Balashov:1992:ENL

[Bal92] Yury V. Balashov. On the evolution of natural laws. British Journal for the Philosophy of Science, 43(3):343–370, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/3/343.full.pdf+
html;   http://www.jstor.
org/stable/687763.

Ballarin:2014:DEL

[Bal14] Roberta Ballarin. Disjunctive effects and the logic of causation. British Journal for the Philosophy of Science, 65 (1):21–38, January 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/21.full.pdf+html.

Bamford:1993:PEA

[Bam93] Greg Bamford. Popper’s explications of ad hoc-ness: Circularity, empirical content, and scientific practice. British Journal for the Philosophy of Science, 44(2):335–355, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        44/2/335.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687653.

Bangu:2013:IE

[Ban13] Sorin Bangu. Indispensability and explanation. British Journal for the Philosophy of Science, 64(2):255–277, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/2/255.full.pdf+html.

Bartlett:1960:RU

[Bar60a] M. S. Bartlett. Review: [untitled]. British Journal for the Philosophy of Science, XI(41):77, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

                                        jstor.org/stable/685825.

Bartlett:1960:R

[Bar60b] M. S. Bartlett. Reviews. British Journal for the Philosophy of Science, XI(41):77a–77, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XI/41/77-a.full.pdf+html.

Bartley:1962:ATE

[Bar62] W. W. Bartley III. Achilles, the tortoise, and explanation in science and history. British Journal for the Philosophy of Science, XIII(49):15–33, May 1962.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/49/15.full.pdf+html;
http://www.jstor.org/
stable/685221.

Barnard:1963:R

[Bar63]  G. A. Barnard.  Reviews.

British Journal for the Philosophy of Science, XIII(52):343– 344, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/343.full.pdf+html;
 http://www.jstor.org/
stable/685343.

Barnard:1964:EA

[Bar64] G. A. Barnard. Each and all. British Journal for the Philosophy of Science, XIV(56): 351, February 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/56/351.full.pdf+html;
http://www.jstor.org/
stable/685641.

Barbour:1969:R

[Bar69] Ian G. Barbour. Reviews. British Journal for the Philosophy of Science, 20(1):89–92, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/1/89.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/686725.

Barnes:1970:R

[Bar70] J. A. Barnes. Reviews. British Journal for the Philosophy of Science, 21(4):399–401, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/4/399.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686568.

Barrett:1971:R

[Bar71] Cyril Barrett. Reviews. British Journal for the Philosophy of Science, 22(2):187–198, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/2/187.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686488.

Barnard:1972:LSI

[Bar72a] G. A. Barnard. The logic of statistical inference. British Journal for the Philosophy of Science, 23(2):123–132, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/2/123.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686437.

Barnard:1972:TPT

[Bar72b] G. A. Barnard. Two points in the theory of statistical inference. British Journal for the Philosophy of Science, 23(4): 329–331, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/329.full.pdf+html;
http://www.jstor.org/
stable/686732.

Barlow:1979:R

[Bar79] Horace Barlow. Reviews. British Journal for the Philosophy of Science, 30(2):204, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/2/204.full.pdf+html;
http://www.jstor.org/
stable/686957.

Barbour:1982:RCS

[Bar82] Julian B. Barbour. Relational concepts of space and time. British Journal for the Philosophy of Science, 33(3):251– 274, September 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/251.full.pdf+html;
http://www.jstor.org/
stable/687224.

BarbosaDeOliveira:1985:PIN

[Bar85a] Marcos Barbosa De Oliveira.

The problem of induction: a new approach. British Journal for the Philosophy of Science, 36(2):129–145, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/2/129.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687030.

Barker:1985:R

[Bar85b] Andrew Barker. Reviews. British Journal for the Philosophy of Science, 36(4):465–469, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/4/465.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687406.

Barnes:1998:PEP

[Bar98] Eric Christian Barnes. Probabilities and epistemic pluralism. British Journal for the Philosophy of Science, 49(1):31–47, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/1/31.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688142.

Barnes:1999:QPO

[Bar99] E. C. Barnes. The quantitative problem of old evidence. British Journal for the Philosophy of Science, 50(2):249–264, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/50/2/249.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072222.

Barker:2000:SMP

[Bar00]  Andrew Barker.  Scientific

method in Ptolemy’s Harmonics. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-55372-5 (hardcover). viii + 281 pp. LCCN MT5.5.P823 B37 2000.

Barrett:2001:BRB

[Bar01] Jeffrey A. Barrett. Book review:

The Strange World of Quantum Mechanics, Daniel F. Styer. British Journal for the Philosophy of Science, 52(2):393–396, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/2/393.full.pdf+
html;   http://www.jstor.
org/stable/3541904.

Bartha:2004:CAF

[Bar04] Paul Bartha. Countable additivity and the de Finetti lottery. British Journal for the Philosophy of Science, 55(2):301–321, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/2/301.full.pdf+
html;   http://www.jstor.
org/stable/3541693.

Barnes:2005:PP

[Bar05a] Eric Christian Barnes. Predictivism for pluralists. British Journal for the Philosophy of Science, 56(3):421–450, September 2005. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/3/421.full.pdf+
html;   http://www.jstor.
org/stable/3541713.

Barrett:2005:PBP

[Bar05b]  Jeffrey  A.  Bar-

rett. The preferred-basis problem and the quantum mechanics of everything. British Journal for the Philosophy of Science, 56(2):199–220, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/2/199.full.pdf+html;
http://www.jstor.org/
stable/3541700.

Barnes:2008:PP

[Bar08a] Eric C. (Eric Christian) Barnes. The paradox of predictivism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87962-0 (hardcover). x + 265 pp. LCCN Q175 .B2186 2008.

Barnes:2008:ELC

[Bar08b] Eric Christian Barnes. Evidence and leverage: Comment on Roush. British Journal for the Philosophy of Science, 59(3): 549–557, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/549.full.pdf+html;
http://www.jstor.org/
stable/40072298.

Barrett:2011:FIP

[Bar11] Jeffrey A. Barrett. On the faithful interpretation of pure wave mechanics. British Journal for the Philosophy of Science, 62(4):693–709, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/693.full.pdf+html.

Barrios:2012:KGC

[Bar12] Edison Barrios. Knowledge of grammar and concept possession. British Journal for the Philosophy of Science, 63(3):577– 606, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/3/577.full.pdf+html.

Barrett:2013:ECM

[Bar13]  David  A.

Barrett. Editor’s choice: Multiple realizability, identity theory, and the gradual reorganization principle. British Journal for the Philosophy of Science, 64(2):325–346, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/325.full.pdf+html.

Barrett:2015:SCM

[Bar15] Thomas William Barrett. On the structure of classical mechanics. British Journal for the Philosophy of Science, 66(4): 801–828, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Bartol:2016:BK

[Bar16]  Jordan Bartol.  Biochemical

kinds. British Journal for the Philosophy of Science, 67 (2):531–551, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/531.

Barack:2019:CR

[Bar19a] David L. Barack. Cognitive recycling. British Journal for the Philosophy of Science, 70 (1):239–268, March 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
239/4096301.

Barker:2019:EPI

[Bar19b] Matthew J. Barker. Eliminative pluralism and integrative alternatives: The case of species. British Journal for the Philosophy of Science, 70(3): 657–681, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
657/4671054.

Baron:2019:MEL

[Bar19c] Sam Baron. Mathematical explanation by law. British Journal for the Philosophy of Science, 70(3):683–717, September 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
683/4671057.

Barrett:2019:EIF

[Bar19d] Thomas William Barrett. Equivalent and inequivalent formulations of classical mechanics. British Journal for the Philosophy of Science, 70(4):1167– 1199, December 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1167/4867963.

Barrett:2023:SAG

[Bar23]  Jeffrey A. Barrett.  Self-

assembling games and the evolution of salience. British Journal for the Philosophy of Science, 74 (1):75–89, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bastin:1964:R

[Bas64] Ted Bastin. Reviews. British Journal for the Philosophy of Science, XV(59):259–264, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/259.full.pdf+
html;   http://www.jstor.
org/stable/686528.

Bass:1992:SPP

[Bas92] Ludvik Bass. Schrödinger: a philosopher in Planck’s chair. British Journal for the Philosophy of Science, 43(1):111–127, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/1/111.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687887.

Batens:1971:SOK

[Bat71] Diderik Batens. Some objections to Keith Lehrer’s rule IR(1). British Journal for the Philosophy of Science, 22(4): 357–362, November 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        22/4/357.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686317.

Batterman:2000:MRU

[Bat00]  R. W. Batterman.  Multi-

ple realizability and universality. British Journal for the Philosophy of Science, 51(1):115–145, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/1/115.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541751.

Batterman:2002:ARM

[Bat02] Robert W. Batterman. Asymptotics and the role of minimal models. British Journal for the Philosophy of Science, 53 (1):21–38, March 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/21.full.pdf+html;
http://www.jstor.org/
stable/3541639.

Batterman:2007:SSF

[Bat07] Robert W. Batterman. On the specialness of special functions (the nonrandom effusions of the divine mathematician). British Journal for the Philosophy of Science, 58(2):263–286, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/2/263.full.pdf+
html;   http://www.jstor.
org/stable/30115226.

Batterman:2010:ERM

[Bat10] Robert W. Batterman. On the explanatory role of mathematics in empirical science. British Journal for the Philosophy of Science, 61(1):1–25, March 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/1.full.pdf+html.

Batterman:2015:BRL

[Bat15] Robert W. Batterman. Book review: Lawrence Sklar, Philosophy and the Foundations of Dynamics. British Journal for the Philosophy of Science, 66(3): 701–705, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/3/701.

Baumer:1964:CP

[Bau64] William H. Baumer. Confirmation without paradoxes. British Journal for the Philosophy of Science, XV(59):177–195, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/59/177.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686522.

Baumer:1968:CSP

[Bau68] William H. Baumer. Confirmation still without paradoxes. British Journal for the Philosophy of Science, 19(1):57–63, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/19/1/57.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686848.

Brading:2004:GST

[BB04] Katherine Brading and Harvey R. Brown. Are gauge symmetry transformations observable? British Journal for the Philosophy of Science, 55(4): 645–665, December 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/645.full.pdf+html;
http://www.jstor.org/
stable/3541620.

Bigelow:1993:BBB

[BCP93] John Bigelow, John Collins, and Robert Pargetter. The big bad bug: What are the Humean’s chances? British Journal for the Philosophy of Science, 44(3): 443–462, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/443.full.pdf+html;
http://www.jstor.org/
stable/688015.

Balzer:1986:KJH

[BD86a] W. Balzer and C. M. Dawe.

Book review: Kyburg Jr., H. E. [1984]: Theory and Measurement. Cambridge University Press. British Journal for the Philosophy of Science, 37(4): 506–510, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687375.

Balzer:1986:SCGb

[BD86b] W. Balzer and C. M. Dawe.

Structure and comparison of genetic theories: (2) the reduction of character-factor genetics to molecular genetics. British Journal for the Philosophy of Science, 37(2):177–191, June 1986.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686976.

Balzer:1986:SCGa

[BD86c] W. Balzer and C. M. Dawe. Structure and comparison of genetic theories: (I) classical genetics. British Journal for the Philosophy of Science, 37 (1):55–69, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686997.

Balzer:1999:SST

[BD99] W. Balzer and V. Dreier. The structure of the spatial theory of elections. British Journal for the Philosophy of Science, 50(4):613–638, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/4/613.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072259.

Baron:2021:WSA

[BE21] Sam Baron and Peter W. Evans. What’s so spatial about time anyway? British Journal for the Philosophy of Science, 72(1): 159–183, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Beattie:1966:R

[Bea66] J. H. M. Beattie. Reviews. British Journal for the Philosophy of Science, XVI(64):340– 341, February 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/64/340.full.pdf+html;
http://www.jstor.org/
stable/686836.

Beausoleil:1989:MPE

[Bea89]  Jean-Roch  Beau-

soleil. The metamathematics– Popperian epistemology connection and its relation to the logic of Turing’s programme. British Journal for the Philosophy of Science, 40(3):307–322, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/3/307.full.pdf+
html;   http://www.jstor.
org/stable/687779.

Becker:2004:NDB

[Bec04] Lon Becker. That von Neumann did not believe in a physical collapse. British Journal for the Philosophy of Science, 55 (1):121–135, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/121.full.pdf+html;
http://www.jstor.org/
stable/3541837.

Beck:2015:AMR

[Bec15] Jacob Beck. Analogue magnitude representations: a philosophical introduction. British Journal for the Philosophy of Science, 66(4):829–855, December 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Bechtel:2019:ANM

[Bec19] William Bechtel. Analysing network models to make discoveries about biological mechanisms. British Journal for the Philosophy of Science, 70(2):459–484, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/2/459/4100193.

Beebee:1998:R

[Bee98] Helen Beebee. Review. British Journal for the Philosophy of Science, 49(1):181–184, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/1/181.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688151.

Beebee:2009:BRJ

[Bee09] Helen Beebee. Book review: John Foster: The Divine Lawmaker. British Journal for the Philosophy of Science, 60 (2):453–457, June 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/2/453.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592010.

Bell:1951:R

[Bel51] Peter Bell. Reviews. British Journal for the Philosophy of Science, I(4):336–339, February 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/4/336.full.pdf+
html;   http://www.jstor.
org/stable/685103.

Belehradek:1961:R

[Bel61] Jan Belehradek. Reviews. British Journal for the Philosophy of Science, XI(44):346–348, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/44/346.full.pdf+
html;   http://www.jstor.
org/stable/685141.

Belsey:1971:R

[Bel71] Andrew Belsey. Reviews. British Journal for the Philosophy of Science, 22(1):78–80, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/78.full.pdf+html;
http://www.jstor.org/
stable/686660.

Bell:1972:R

[Bel72] J. L. Bell. Reviews. British Journal for the Philosophy of Science, 23(4):363–366, November 1972. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/4/363.full.pdf+
html;   http://www.jstor.
org/stable/686739.

Bell:1975:R

[Bel75] J. L. Bell. Reviews. British Journal for the Philosophy of Science, 26(2):165–170, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/2/165.full.pdf+html;
http://www.jstor.org/
stable/686516.

Bell:1976:R

[Bel76]  John Bell.  Reviews.

British Journal for the Philosophy of Science, 27(2):187–191, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/2/187.full.pdf+
html;   http://www.jstor.
org/stable/686170.

Bell:1977:R

[Bel77] John Bell. Reviews. British Journal for the Philosophy of Science, 28(1):95, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/1/95.full.pdf+html;
http://www.jstor.org/
stable/686454.

Beloff:1978:MA

[Bel78]  John Beloff.  Is mind au-

tonomous? British Journal for the Philosophy of Science, 29(3):265–273, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/265.full.pdf+html;
http://www.jstor.org/
stable/687017.

Bell:1979:IPR

[Bel79a] John Bell. The infinite past regained: a reply to Whitrow. British Journal for the Philosophy of Science, 30(2):161–165, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/161.full.pdf+
html;   http://www.jstor.
org/stable/686946.

Bell:1979:R

[Bel79b] John Bell. Reviews. British Journal for the Philosophy of Science, 30(3):306–309, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/306.full.pdf+
html;   http://www.jstor.
org/stable/687148.

Bell:1981:CTF

[Bel81] J. L. Bell. Category theory and the foundations of mathematics. British Journal for the Philosophy of Science, 32(4):

349–358, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/349.full.pdf+html;
http://www.jstor.org/
stable/687305.

Bell:1982:R

[Bel82] J. L. Bell. Reviews. British Journal for the Philosophy of Science, 33(1):95–97, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/1/95.full.pdf+html;
http://www.jstor.org/
stable/687244.

Bell:1986:NAQ

[Bel86] J. L. Bell. A new approach to quantum logic. British Journal for the Philosophy of Science, 37(1):83–99, March 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686999.

Bigelow:1992:WOK

[BEL92] John Bigelow, Brian Ellis, and Caroline Lierse. The world as one of a kind: Natural necessity and laws of nature. British Journal for the Philosophy of Science, 43(3):371–388, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/3/371.full.pdf+

html;   http://www.jstor.
org/stable/687764.

Belot:1995:NWC

[Bel95] Gordon Belot. New work for counterpart theorists: Determinism. British Journal for the Philosophy of Science, 46 (2):185–195, June 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/2/185.full.pdf+html;
http://www.jstor.org/
stable/687811.

Belsey:1996:R

[Bel96] Andrew Belsey. Review. British Journal for the Philosophy of Science, 47(2):335–338, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/335.full.pdf+
html;   http://www.jstor.
org/stable/687956.

Belsey:1997:R

[Bel97] Andrew Belsey. Reviews. British Journal for the Philosophy of Science, 48(2):281–283, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/281.full.pdf+
html;   http://www.jstor.
org/stable/687749.

Belot:1998:UE

[Bel98] Gordon Belot. Understanding electromagnetism. British Journal for the Philosophy of Science, 49(4):531–555, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/4/531.full.pdf+html;
http://www.jstor.org/
stable/688130.

Belot:2000:GM

[Bel00] Gordon Belot. Geometry and motion. British Journal for the Philosophy of Science, 51(4): 561–595, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541608.

Belot:2001:BRB

[Bel01] Gordon Belot. Book review: Theory and Truth: Philosophical Critique within Foundational Science, Lawrence Sklar. British Journal for the Philosophy of Science, 52(3):647–650, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/3/647.full.pdf+
html;   http://www.jstor.
org/stable/3541937.

Belnap:2005:TCC

[Bel05a] Nuel Belnap. A theory of causation: Causae causantes (originating causes) as Inus conditions in branching space–times. British Journal for the Philosophy of Science, 56(2):221–253, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/2/221.full.pdf+
html;   http://www.jstor.
org/stable/3541701.

Belot:2005:DTS

[Bel05b] Gordon Belot. Dust, time and symmetry. British Journal for the Philosophy of Science, 56 (2):255–291, June 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/2/255.full.pdf+html;
http://www.jstor.org/
stable/3541702.

Belot:2017:CCF

[Bel17a] Gordon Belot. Correction: Curve-Fitting for Bayesians? British Journal for the Philosophy of Science, 68 (3):927, September 1, 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL https://academic.oup.
com/bjps/article/68/3/927/

4087231. See [Bel17b].

Belot:2017:CFB

[Bel17b] Gordon Belot. Curve-fitting for Bayesians? British Journal for the Philosophy of Science, 68(3):689–702, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/

                                        article/68/3/689/2669739.

See correction [Bel17a].

Belkoniene:2023:GU

[Bel23] Miloud Belkoniene. Grasping in understanding. British Journal for the Philosophy of Science, 74 (3):603–617, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bennett:1956:SAP

[Ben56] Jonathan Bennett. Some aspects of probability and induction (I). British Journal for the Philosophy of Science, VII (27):220–230, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/27/220.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685874.

Bennett:1957:SAP

[Ben57] Jonathan Bennett. Some aspects of probability and induction (II). British Journal for the Philosophy of Science, VII(28): 316–322, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/28/316.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685111.

Bennett:1963:E

[Ben63a] Jonathan Bennett. Erratum. British Journal for the Philosophy of Science, XIV(55):264s– 264, November 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/55/264-s.full.pdf+html.

Bennett:1963:ESD

[Ben63b] Jonathan Bennett. Erratum: The status of determinism. British Journal for the Philosophy of Science, XIV (55):264, November 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685246.

Bennett:1963:SD

[Ben63c] Jonathan Bennett. The status of determinism. British Journal for the Philosophy of Science, XIV(54):106–119, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/54/106.full.pdf+html;
http://www.jstor.org/
stable/685429.

Bennett:1990:MMR

[Ben90] Laura J. Bennett. Modularity of mind revisited. British Journal for the Philosophy of Science, 41(3):429–436, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/3/429.full.pdf+
html;   http://www.jstor.
org/stable/687830.

Bennett:2012:SSS

[Ben12] David J. Bennett. Seeing shape: Shape appearances and shape constancy. British Journal for the Philosophy of Science, 63(3): 487–518, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/3/487.full.pdf+html.

Bernal:1955:RHM

[Ber55a] J. D. Bernal. Reviews: Has history a meaning? British Journal for the Philosophy of Science, VI (22):164–169, August 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/22/164.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685525.

Bernal:1955:S

[Ber55b] J. D. Bernal. Symmetry. British Journal for the Philosophy of Science, V(20):335–341, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/20/335.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685738.

Bernays:1959:R

[Ber59] Paul Bernays. Reviews. British Journal for the Philosophy of Science, IX(36):321–327, February 1959. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/36/321.full.pdf+
html;   http://www.jstor.
org/stable/685194.

Berger:1968:CPT

[Ber68] George Berger. The conceptual possibility of time travel. British Journal for the Philosophy of Science, 19(2):152–155, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/152.full.pdf+
html;   http://www.jstor.
org/stable/686793.

Bernstein:1969:R

[Ber69] David L. Bernstein. Reviews.

British Journal for the Philosophy of Science, 20(1):86–87, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/86.full.pdf+
html;   http://www.jstor.
org/stable/686723.

Berkson:1978:RPW

[Ber78] William Berkson. Reply to L.

Pearce Williams. British Journal for the Philosophy of Science, 29(3):243–248, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/243.full.pdf+html;

http://www.jstor.org/
stable/687011.

Belot:1999:HIL

[BER99]  G.  Belot,

J. Earman, and L. Ruetsche. The Hawking information loss paradox: the anatomy of controversy. British Journal for the Philosophy of Science, 50(2):189–229, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/2/189.full.pdf+
html;   http://www.jstor.
org/stable/40072220.

Berry:2010:BRA

[Ber10] Michael Berry. Book review: Alisa Bokulich, Reexamining the Quantum-Classical Relation: Beyond Reductionism and Pluralism. British Journal for the Philosophy of Science, 61(4):889–895, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/4/889.full.pdf+html.

Bergeron:2016:FIC

[Ber16] Vincent Bergeron. Functional independence and cognitive architecture. British Journal for the Philosophy of Science, 67(3): 817–836, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/817.

Berry:2023:SIE

[Ber23] Sharon Berry. Σ_{1}^{0} soundness isn’t enough: Number-theoretic indeterminacy’s unsavory physical commitments. British Journal for the Philosophy of Science, 74(2):469–484, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Beth:1950:CED

[Bet50] Evert W. Beth. Critical epochs in the development of the theory of science. British Journal for the Philosophy of Science, I(1):27– 42, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/1/27.full.pdf+
html;   http://www.jstor.
org/stable/685668.

Beth:1951:Ea

[Bet51a] E. W. Beth. Errata. British Journal for the Philosophy of Science, I(4):302, February 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/4/302.full.pdf+html.

Beth:1951:Eb

[Bet51b] E. W. Beth. Errata. British Journal for the Philosophy of Science, II(5):1s–1, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/5/1-s.full.pdf+html.

Beth:1951:EFF

[Bet51c] E. W. Beth. Errata: Fundamental features of contemporary theory of science. British Journal for the Philosophy of Science, II(5):i–iv, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685604.

Beth:1951:FFC

[Bet51d] Evert W. Beth. Fundamental features of contemporary theory of science. British Journal for the Philosophy of Science, I(4): 291–302, February 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/4/291.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685097.

Beth:1951:R

[Bet51e] Evert W. Beth. Reviews. British Journal for the Philosophy of Science, I(4):333–335, February 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/I/4/333.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685101.

Beth:1952:R

[Bet52a] E. W. Beth. Reviews. British Journal for the Philosophy of Science, II(8):329–330, February 1952. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/329.full.pdf+
html;   http://www.jstor.
org/stable/686027.

Beth:1952:PRF

[Bet52b] Evert W. Beth. The prehistory of research into foundations. British Journal for the Philosophy of Science, III(9):58– 81, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/58.full.pdf+
html;   http://www.jstor.
org/stable/685474.

Beth:1959:Ra

[Bet59a] E. W. Beth. Reviews. British Journal for the Philosophy of Science, X(38):159–160, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/159.full.pdf+
html;   http://www.jstor.
org/stable/686008.

Beth:1959:Rb

[Bet59b] E. W. Beth. Reviews. British Journal for the Philosophy of Science, X(39):249–250, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/39/249.full.pdf+
html;   http://www.jstor.
org/stable/686063.

Betz:2016:TET

[Bet16] Gregor Betz. Truth in evidence and truth in arguments without logical omniscience. British Journal for the Philosophy of Science, 67(4):1117–1137, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/4/1117.

Betzler:2024:TSN

[Bet24] Riana J. Betzler. Two sources of normativity in enthusiastic accounts of kinds. British Journal for the Philosophy of Science, 75 (1):127–152, March 2024. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Beveridge:1952:R

[Bev52] W. I. B. Beveridge. Reviews. British Journal for the Philosophy of Science, III(11):291– 293, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/11/291.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685285.

Baird:1990:SIS

[BF90] Davis Baird and Thomas Faust. Scientific instruments, scientific progress and the cyclotron. British Journal for the Philosophy of Science, 41(2):147–175, June 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/2/147.full.pdf+
html;   http://www.jstor.
org/stable/687770.

Bueno:2000:BRB

[BF00] O. Bueno and S. French. Book review: An Essay on Contraction. A. Fuhrmann. British Journal for the Philosophy of Science, 51(3):513–517, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/513.full.pdf+html;
http://www.jstor.org/
stable/3541780.

Bueno:2011:HTR

[BF11] Otávio Bueno and Steven French. How theories represent. British Journal for the Philosophy of Science, 62(4):857–894, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/857.full.pdf+html.

Bueno:2012:CMEa

[BF12a]  Otávio  Bueno

and Steven French. Can mathematics explain physical phenomena? British Journal for the Philosophy of Science, 63 (1):85–113, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/1/85.full.pdf+html.

Bueno:2012:CMEb

[BF12b]  Otávio  Bueno

and Steven French. Can mathematics explain physical phenomena? British Journal for the Philosophy of Science, 63(2):455, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/2/455.full.pdf+html.

Baumgartner:2023:BDM

[BF23]  Michael  Baumgart-

ner and Christoph Falk. Boolean difference-making: a modern regularity theory of causation. British Journal for the Philosophy of Science, 74(1):171–197, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bovens:2002:TOT

[BFHS02] Luc Bovens, Branden Fitelson, Stephan Hartmann, and Josh Snyder. Too odd (not) to be true? a reply to Olsson. British Journal for the Philosophy of Science, 53(4):539–563, December 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/53/4/539.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541823.

Bassi:1999:DMA

[BG99] A. Bassi and G-C Ghirardi. Discussion. more about dynamical reduction and the enumeration principle. British Journal for the Philosophy of Science, 50(4): 719–734, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/719.full.pdf+html;
http://www.jstor.org/
stable/40072263.

Bassi:2001:CMR

[BG01] Angelo Bassi and GianCarlo Ghirardi. Counting marbles: Reply to Clifton and Monton. British Journal for the Philosophy of Science, 52(1):125–130, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/125.full.pdf+
html;   http://www.jstor.
org/stable/3541946.

Baumgartner:2016:CRM

[BG16a]  Michael  Baum-

gartner and Alexander Gebharter. Constitutive relevance, mutual manipulability, and fathandedness. British Journal for the Philosophy of Science, 67(3):731–756, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/731.

Bortolotti:2016:BRP

[BG16b]  Lisa  Bortolotti

and Rachel Gunn. Book review: Philip Gerrans, The Measure of Madness: Philosophy of Mind, Cognitive Neuroscience, and Delusional Thought. British Journal for the Philosophy of Science, 67(3):919–923, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/3/919.

Bradley:2023:MDL

[BGH^{+}23] Adam Bradley, Quinn Hiroshi Gibson, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Thomas Müller, Tomasz Placek, Matteo Colombo, Peggy Seriès, Jeffrey A. Barrett, Brian Skyrms, Porter Williams, and Isaac Wilhelm. Monothematic delusions and the limits of rationality. British Journal for the Philosophy of Science, 74(3): 811–835, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Babic:2024:NER

[BGTW24] Boris Babic, Anil Gaba, Ilia Tsetlin, and Robert L. Winkler. Normativity, epistemic rationality, and noisy statistical evidence. British Journal for the Philosophy of Science, 75(1): 153–176, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bar-Hillel:1953:NCI

[BH53] Yehoshua Bar-Hillel. A note on comparative inductive logic.

British Journal for the Philosophy of Science, III(12):308– 310, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/308.full.pdf+html;
http://www.jstor.org/
stable/685444.

Bar-Hillel:1955:CDC

[BH55] Yehoshua Bar-Hillel. Comments on ‘degree of confirmation’ by Professor K. R. Popper. British Journal for the Philosophy of Science, VI(22):155–157, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/22/155.full.pdf+
html;   http://www.jstor.
org/stable/685523.

Bar-Hillel:1956:CDC

[BH56] Y. Bar-Hillel. Content and degree of confirmation: Further comments on probability and confirmation a rejoinder to Professor Popper. British Journal for the Philosophy of Science, VII(27):245–248, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/245.full.
pdf+html;   http://www.
jstor.org/stable/685882.

Brink:1987:VOT

[BH87] Chris Brink and Johannes Heidema. A verisimilar ordering of theories phrased in a propositional language. British Journal for the Philosophy of Science, 38(4):533–549, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/533.full.pdf+html;
http://www.jstor.org/
stable/687358.

Brink:1991:VPR

[BH91] Chris Brink and Johannes Heidema. Verisimilitude by power relations: a response to Oddie. British Journal for the Philosophy of Science, 42(1):101–104, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/1/101.full.pdf+
html;   http://www.jstor.
org/stable/687609.

Baigrie:1992:CSS

[BH92] Brian S. Baigrie and J. N. Hattiangadi. On consensus and stability in science. British Journal for the Philosophy of Science, 43 (4):435–458, October 1992. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/4/435.full.pdf+html;
http://www.jstor.org/
stable/687835.

Burger:1994:CTT

[BH94] Isabella C. Burger and Johannes Heidema. Comparing theories by their positive and negative contents. British Journal for the Philosophy of Science, 45 (2):605–630, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/605.full.pdf+html;
http://www.jstor.org/
stable/687686.

Baker:2010:A

[BH10] David John Baker and Hans Halvorson. Antimatter. British Journal for the Philosophy of Science, 61(1):93–121, March 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/93.full.pdf+html.

Brossel:2015:BCM

[BH15] Peter Brössel and Franz Huber.

Bayesian confirmation: a means with no end. British Journal for the Philosophy of Science, 66(4): 737–749, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Bhave:1986:SHV

[Bha86] S. V. Bhave. Separable hidden variables theory to explain Einstein–Podolsky–Rosen paradox. British Journal for the Philosophy of Science, 37(4): 467–475, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687372.

Bhave:1991:BIQ

[Bha91] S. V. Bhave. Bell’s Inequalities and quantum non-separability. British Journal for the Philosophy of Science, 42(4):541– 545, December 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/4/541.full.pdf+html;
http://www.jstor.org/
stable/687708.

Bar-Hillel:1953:SI

[BHC53] Yehoshua Bar-Hillel and Rudolf Carnap. Semantic information. British Journal for the Philosophy of Science, IV(14):147–157, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/147.full.pdf+
html;   http://www.jstor.
org/stable/685989.

Bar-Hillel:1972:NPR

[BHM72] Maya Bar-Hillel and Avishai Margalit. Newcomb’s Paradox revisited. British Journal for the Philosophy of Science, 23(4): 295–304, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/295.full.pdf+html;
http://www.jstor.org/
stable/686730.

Baker:2015:CP

[BHS15] David John Baker, Hans Halvorson, and Noel Swanson. The conventionality of parastatistics. British Journal for the Philosophy of Science, 66(4):929– 976, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Benci:2018:IP

[BHW18] Vieri Benci, Leon Horsten, and Sylvia Wenmackers. Infinitesimal probabilities. British Journal for the Philosophy of Science, 69(2):509–552, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/509/2669779.

Bicchieri:1987:RPE

[Bic87] Cristina Bicchieri. Rationality and predictability in economics. British Journal for the Philosophy of Science, 38(4): 501–513, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/501.full.pdf+html;
http://www.jstor.org/
stable/687355.

Bickle:1998:PRN

[Bic98] John Bickle. Psychoneural reduction: the new wave. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-02432-2. xi + 240 pp. LCCN BF38 .B48 1998; BF38 .B583p 1998.

Bicchieri:2006:GSN

[Bic06] Cristina Bicchieri. The grammar of society: the nature and dynamics of social norms. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-57372-6 (hardcover), 0-521-57490-0 (paperback). xvi + 260 pp. LCCN HM676 .B53

                                        2006. URL http://www.loc.
                                        gov/catdir/toc/ecip059/
                                        2005006200.html.

Billinge:2000:DAC

[Bil00] H. Billinge. Discussion. applied constructive mathematics: on Hellman’s ’mathematical constructivism in spacetime’. British Journal for the Philosophy of Science, 51(2):299–318, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/2/299.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541807.

Binet:1954:NRL

[Bin54] F. E. Binet. Notes on a remark by Lord Russell. British Journal for the Philosophy of Science, V(17):67–70, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/17/67.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685953.

Binns:1971:R

[Bin71] Peter Binns. Reviews. British Journal for the Philosophy of Science, 22(2):198–200, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/198.full.pdf+
html;   http://www.jstor.
org/stable/686489.

Binns:1990:EEP

[Bin90] P. Binns. Experimental evidence and psychotherapy. British Journal for the Philosophy of Science, 41(4):531–552, December 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/4/531.full.pdf+
html;   http://www.jstor.
org/stable/687864.

Binmore:2009:RD

[Bin09] K. G. Binmore. Rational decisions. Gorman lectures in economics. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13074-4 (hardcover). x + 200 pp. LCCN QA279.5 .B56 2009.

Binkoski:2017:VGR

[Bin17] James P. Binkoski. On the viability of Galilean relationalism. British Journal for the Philosophy of Science, 68(4):1183–1204, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/1183/2669749.

Binkoski:2019:GFS

[Bin19]  James Binkoski.  Geometry,

fields, and spacetime. British Journal for the Philosophy of Science, 70(4):1097–1117, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/4/1097/4819209.

Birks:1956:TSSa

[Bir56a] G. A. Birks. Towards a science of social relations (I). British Journal for the Philosophy of Science, VII(26):117–128, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/26/117.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685695.

Birks:1956:TSSb

[Bir56b] G. A. Birks. Towards a science of social relations (II). British Journal for the Philosophy of Science, VII(27):206– 211, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/27/206.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685872.

Birnbaum:1956:R

[Bir56c] Norman Birnbaum. Reviews. British Journal for the Philosophy of Science, VII(25):114–116, May 1956. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/25/114.full.
pdf+html;   http://www.
jstor.org/stable/685944.

Birner:1985:R

[Bir85] Jack Birner. Reviews. British Journal for the Philosophy of Science, 36(2):215–221, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/2/215.full.pdf+
html;   http://www.jstor.
org/stable/687038.

Bird:1996:R

[Bir96] Alexander Bird. Review. British Journal for the Philosophy of Science, 47(1):149–151, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/149.full.pdf+
html;   http://www.jstor.
org/stable/1215836.

Bird:1997:R

[Bir97] Alexander Bird. Review. British Journal for the Philosophy of Science, 48(2):284–286, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/284.full.pdf+
html;   http://www.jstor.
org/stable/687750.

Bird:2001:DAC

[Bir01] Alexander Bird. David Armstrong, Charlie Martin, and Ullin Place, edited by Tim Crane. Dispositions: a debate; Stephen Mumford dispositions. British Journal for the Philosophy of Science, 52(1):137–149, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/137.full.pdf+
html;   http://www.jstor.
org/stable/3541948.

Bird:2012:BSS

[Bir12] Alexander Bird. The Structure of Scientific Revolutions and its significance: an essay review of the fiftieth anniversary edition. British Journal for the Philosophy of Science, 63(4): 859–883, December 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/4/859.full.pdf+html.

Birch:2013:BRS

[Bir13] Jonathan Birch. Book review: Samir Okasha and Ken Binmore (eds), Evolution and Rationality: Decisions, Cooperation, and Strategic Behaviour. British Journal for the Philosophy of Science, 64(3):669– 673, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
64/3/669.full.pdf+html.

Birch:2014:ECH

[Bir14] Jonathan Birch. Editor’s choice:

Hamilton’s rule and its discontents. British Journal for the Philosophy of Science, 65 (2):381–411, June 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/381.

Birch:2020:KSG

[Bir20] Jonathan Birch. Kin selection, group selection, and the varieties of population structure. British Journal for the Philosophy of Science, 71(1):259–286, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/71/1/259/4100192.

Bird:2022:URC

[Bir22] Alexander Bird. Understanding the replication crisis as a base rate fallacy. British Journal for the Philosophy of Science, 72(4): 965–993, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Balashov:2003:PR

[BJ03]  Yuri  Balashov

and Michel Janssen. Presentism and Relativity. British Journal for the Philosophy of Science, 54(2):327–346, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/2/327.full.pdf+html;
http://www.jstor.org/
stable/3541972.

Bailer-Jones:2004:RMT

[BJ04] Daniela Bailer-Jones. Review: Making truth: Metaphor in science. British Journal for the Philosophy of Science, 55(4): 811–815, December 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/811.full.pdf+html;
http://www.jstor.org/
stable/3541634.

Bondi:1959:RIL

[BK59] H. Bondi and C. W. Kilmister. Reviews: the impact of Logik Der Forschung (German) [Logic of research]. British Journal for the Philosophy of Science, X(37):55–57, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/55.full.pdf+html;
http://www.jstor.org/
stable/685778.

Bird:2005:E

[BL05] Alexander Bird and James Ladyman. Editorial. British Journal for the Philosophy of Science, 56(1):1–2, March 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
56/1/1.full.pdf+html;
http://www.jstor.org/
stable/3541847.

Bird:2006:E

[BL06] Alexander Bird and James Ladyman. Editorial. British Journal for the Philosophy of Science, 57(2):273, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/273.full.pdf+html;
http://www.jstor.org/
stable/3873467.

Brading:2014:BRD

[BL14] Katherine Brading and Xavi Lanao. Book review: Don Ross, James Ladyman, and Harold Kincaid (eds), Scientific Metaphysics. British Journal for the Philosophy of Science, 65(4): 899–903, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/899.

Baas:2023:WDW

[BL23] Augustin Baas and Baptiste Le Bihan. What does the world look like according to superdeterminism? British Journal for the Philosophy of Science, 74(3): 555–572, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Blair:1950:SAS

[Bla50] G. W. Scott Blair. Some aspects of the search for invariants. British Journal for the Philosophy of Science, I(3):230– 244, November 1950. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/3/230.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685369.

Blair:1952:TPP

[Bla52] G. W. Scott Blair. Time of psychology and of physics. British Journal for the Philosophy of Science, III(9):82–85, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/9/82.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685475.

Blair:1953:TPP

[Bla53]  G. W. Scott Blair.  Time

of psychology and of physics. British Journal for the Philosophy of Science, III(12):359a– 359, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/359-a.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685455.

Black:1955:LVF

[Bla55] Virginia Black. Laboratory versus field research in psychology and the social sciences. British Journal for the Philosophy of Science, V(20):319–330, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/319.full.pdf+
html;   http://www.jstor.
org/stable/685733.

Black:1958:R

[Bla58a] Max Black. Reviews. British Journal for the Philosophy of Science, IX(33):80–81, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/80.full.pdf+
html;   http://www.jstor.
org/stable/686112.

Blair:1958:R

[Bla58b] G. W. Scott Blair. Reviews.

British Journal for the Philosophy of Science, IX(33):66–67, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/66.full.pdf+
html;   http://www.jstor.
org/stable/686105.

Blair:1959:R

[Bla59] G. W. Scott Blair. Reviews.

British Journal for the Philosophy of Science, X(37):81–83, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/37/81.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685783.

Black:1960:R

[Bla60a] Max Black. Reviews. British Journal for the Philosophy of Science, XI(41):79–80, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/79.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685828.

Blair:1960:R

[Bla60b] G. W. Scott Blair. Reviews. British Journal for the Philosophy of Science, XI(41):78–79, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/78.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685827.

Blair:1962:R

[Bla62] G. W. Scott Blair. Reviews. British Journal for the Philosophy of Science, XIII(49):85–86, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/49/85.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685235.

Blair:1963:DPF

[Bla63a] G. W. Scott Blair. Discussion of Professor F. A. Paneth’s second article. British Journal for the Philosophy of Science, XIV(53): 40, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/40.full.pdf+
html;   http://www.jstor.
org/stable/685351.

Blair:1963:R

[Bla63b] G. W. Scott Blair. Reviews.

British Journal for the Philosophy of Science, XIII(52):339– 340, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/339.full.pdf+html;
http://www.jstor.org/
stable/685340.

Black:1966:RDI

[Bla66] Max Black. The raison d’Être of inductive argument. British Journal for the Philosophy of Science, 17(3):177–204, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/3/177.full.pdf+
html;   http://www.jstor.
org/stable/686756.

Blackburn:1969:R

[Bla69] S. Blackburn. Reviews. British Journal for the Philosophy of Science, 19(4):371–373, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/4/371.full.pdf+
html;   http://www.jstor.
org/stable/686297.

Blackburn:1970:R

[Bla70] S. Blackburn. Reviews. British Journal for the Philosophy of Science, 21(4):385–386, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/385.full.pdf+
html;   http://www.jstor.
org/stable/686560.

Blackburn:1971:R

[Bla71]  Simon Blackburn.  Reviews.

British Journal for the Philosophy of Science, 22(2):202–205, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/202.full.pdf+
html;   http://www.jstor.
org/stable/686491.

Blackmore:1978:DNP

[Bla78] John T. Blackmore. Discussion notes: Is Planck’s ‘Principle’ true? British Journal for the Philosophy of Science, 29(4):347–349, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

29/4/347.full.pdf+html;
http://www.jstor.org/
stable/687097.

Blackmore:1979:IUT

[Bla79] John Blackmore. On the inverted use of the terms ‘realism’ and ‘idealism’ among scientists and historians of science. British Journal for the Philosophy of Science, 30(2):125–134, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/125.full.pdf+
html;   http://www.jstor.
org/stable/686944.

Blaug:1982:R

[Bla82] Mark Blaug. Reviews. British Journal for the Philosophy of Science, 33(3):323–324, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/3/323.full.pdf+
html;   http://www.jstor.
org/stable/687231.

Blair:1984:R

[Bla84] Gordon Blair. Reviews. British Journal for the Philosophy of Science, 35(2):179–184, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/179.full.pdf+
html;   http://www.jstor.
org/stable/686969.

Blackmore:1985:HNE

[Bla85] John Blackmore. An historical note on Ernst Mach. British Journal for the Philosophy of Science, 36(3):299–305, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/3/299.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687573.

Blackmore:1989:EML

[Bla89] John Blackmore. Ernst Mach leaves ‘The Church of Physics’. British Journal for the Philosophy of Science, 40(4):519– 540, December 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/4/519.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687739.

Black:1998:CCP

[Bla98] Robert Black. Chance, credence, and the principal principle. British Journal for the Philosophy of Science, 49(3):371– 385, September 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        49/3/371.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688080.

Blachowicz:2009:HST

[Bla09] James Blachowicz. How science textbooks treat scientific method: a philosopher’s perspective. British Journal for the Philosophy of Science, 60 (2):303–344, June 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/2/303.full.pdf+html;
http://www.jstor.org/
stable/25592003.

Blanchard:2020:EAG

[Bla20] Thomas Blanchard. Explanatory abstraction and the Goldilocks problem: Interventionism gets things just right. British Journal for the Philosophy of Science, 71(2):633–663, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
633/5053917.

Bloor:1970:OTM

[Blo70] David Bloor. Is the official theory of mind absurd? British Journal for the Philosophy of Science, 21(2):167–183, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/2/167.full.pdf+html;
http://www.jstor.org/
stable/686697.

Blum:1970:LIS

[Blu70] Alex Blum. Laws and instantial statements. British Journal for the Philosophy of Science, 21(4):371–378, November 1970.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/4/371.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686558.

Blum:1989:ADM

[Blu89] Alex Blum. An anomaly in the D–N model of explanation. British Journal for the Philosophy of Science, 40(3):365– 367, September 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/3/365.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687783.

Ben-Menahem:1990:ED

[BM90] Yemima Ben-Menahem. Equivalent descriptions. British Journal for the Philosophy of Science, 41(2):261–279, June 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        41/2/261.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687775.

Brown:1993:LSC

[BM93] Harvey R. Brown and Adolfo Maia, Jr. Light-speed constancy versus light-speed invariance in the derivation of relativistic kinematics. British Journal for the Philosophy of Science, 44(3):381–407, September 1993.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/381.full.pdf+html;
http://www.jstor.org/
stable/688012.

Ben-Menahem:2001:CPS

[BM01] Yemima Ben-Menahem. Convention: Poincaré and some of his critics. British Journal for the Philosophy of Science, 52(3): 471–513, September 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/471.full.pdf+html;
http://www.jstor.org/
stable/3541926.

Braddon-Mitchell:2005:SR

[BM05] David Braddon-Mitchell. The subsumption of reference. British Journal for the Philosophy of Science, 56(1):157–178, March 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/1/157.full.pdf+
html;   http://www.jstor.
org/stable/3541855.

Belnap:2010:BUS

[BM10] Nuel Belnap and Thomas Müller.

Branching with uncertain semantics: Discussion note on Saunders and Wallace, ‘Branching and Uncertainty’. British Journal for the Philosophy of Science, 61(3):681–696, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/3/681.full.pdf+html.

Bloch-Mullins:2018:BGB

[BM18] Corinne L. Bloch-Mullins. Bridging the gap between similarity and causality: an integrated approach to concepts. British Journal for the Philosophy of Science, 69(3):605–632, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/3/605/3605017.

Baird:1994:FWP

[BN94]  Davis  Baird  and

Alfred Nordmann. Facts-well-put. British Journal for the Philosophy of Science, 45(1):37–77, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/1/37.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687961.

Boden:1964:R

[Bod64] Margaret A. Boden. Reviews. British Journal for the Philosophy of Science, XIV(56):366– 368, February 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/56/366.full.pdf+html;

http://www.jstor.org/
stable/685645.

Boden:1966:R

[Bod66] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 17(2):162–165, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/2/162.full.pdf+
html;   http://www.jstor.
org/stable/686353.

Boden:1968:R

[Bod68] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 19(3):271– 274, November 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/3/271.full.pdf+html;
http://www.jstor.org/
stable/686255.

Boden:1969:R

[Bod69] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 20(2):176–177, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/176.full.pdf+
html;   http://www.jstor.
org/stable/686187.

Boden:1970:Ra

[Bod70a] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 21(1):119– 121, February 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/1/119.full.pdf+html;
http://www.jstor.org/
stable/686477.

Boden:1970:Rb

[Bod70b] Margaret A. Boden. Reviews. British Journal for the Philosophy of Science, 21(2):212–214, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/212.full.pdf+
html;   http://www.jstor.
org/stable/686707.

Boden:1971:R

[Bod71] Margaret A. Boden. Reviews. British Journal for the Philosophy of Science, 22(3):309–312, August 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/3/309.full.pdf+
html;   http://www.jstor.
org/stable/686752.

Boden:1973:HAA

[Bod73] Margaret A. Boden. How artificial is artificial intelligence? British Journal for the Philosophy of Science, 24(1):61–72, March 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/24/1/61.full.pdf+
html;   http://www.jstor.
org/stable/686670.

Boden:1974:R

[Bod74] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 25(1):85–88, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/85.full.pdf+
html;   http://www.jstor.
org/stable/686505.

Boden:1975:R

[Bod75] Margaret A. Boden. Reviews.

British Journal for the Philosophy of Science, 26(3):269– 273, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/269.full.pdf+html;
http://www.jstor.org/
stable/686286.

Boden:1999:MN

[Bod99] M. A. Boden. Is metabolism necessary? British Journal for the Philosophy of Science, 50 (2):231–248, June 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/231.full.pdf+html;
http://www.jstor.org/
stable/40072221.

Boesche:2022:RQE

[Boe22] Tom Boesche. Reassessing quasiexperiments: Policy evaluation, induction, and SUTVA. British Journal for the Philosophy of Science, 73(1):1–22, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bogen:1977:FDS

[Bog77] Joseph E. Bogen. Further discussion on split-brains and hemispheric capabilities. British Journal for the Philosophy of Science, 28(3):281–286, September 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/28/3/281.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686810.

Bohm:1961:RBM

[Boh61] David Bohm. On the relationship between methodology in scientific research and the content of scientific knowledge. British Journal for the Philosophy of Science, XII(46):103–116, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/46/103.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685490.

Bohm:1962:CNC

[Boh62] David Bohm. Classical and nonclassical concepts in the quantum theory: an answer to Heisenberg’s Physics and Philosophy. British Journal for the Philosophy of Science, XII(48): 265–280, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/48/265.full.pdf+html;
http://www.jstor.org/
stable/685396.

Bohm:1966:R

[Boh66] David Bohm. Reviews. British Journal for the Philosophy of Science, 17(3):253–256, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/3/253.full.pdf+
html;   http://www.jstor.
org/stable/686763.

Bohm:1968:R

[Boh68] David Bohm. Reviews. British Journal for the Philosophy of Science, 19(2):173–175, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/173.full.pdf+
html;   http://www.jstor.
org/stable/686799.

Bokulich:2008:CCS

[Bok08a] Alisa Bokulich. Can classical structures explain quantum phenomena? British Journal for the Philosophy of Science, 59(2):217–235, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/2/217.full.pdf+html;
http://www.jstor.org/
stable/40072279.

Bokulich:2008:RQC

[Bok08b] Alisa Bokulich. Reexamining the quantum-classical relation: beyond reductionism and pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp. LCCN QC174.12 .B65

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0834/2008018902-b.html;
http://www.loc.gov/catdir/
enhancements/fy0834/
2008018902-d.html;
http://www.loc.gov/catdir/
enhancements/fy0834/
2008018902-t.html.

Boland:1983:BSD

[Bol83] Lawrence A. Boland. On the best strategy for doing philosophy of economics. British Journal for the Philosophy of Science, 34(4):387–392, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/387.full.pdf+html;
http://www.jstor.org/
stable/687531.

Bolton:1996:R

[Bol96] Derek Bolton. Review. British Journal for the Philosophy of Science, 47(3):474–475, September 1996. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/3/474.full.pdf+
html;   http://www.jstor.
org/stable/687991.

Bondi:1955:CLH

[Bon55a] H. Bondi. Comment on Lord Halsbury’s remarks. British Journal for the Philosophy of Science, VI(23):244, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/244.full.pdf+
html;   http://www.jstor.
org/stable/685682.

Bondi:1955:DPK

[Bon55b] H. Bondi. A discussion of Professor Kapp’s views. British Journal for the Philosophy of Science, VI(23):239–241, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/239.full.pdf+
html;   http://www.jstor.
org/stable/685679.

Bonnor:1958:IR

[Bon58] W. B. Bonnor. Instrumentalism and Relativity. British Journal for the Philosophy of Science, VIII(32):291–294, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/291.full.

                                        pdf+html;   http://www.
                                        jstor.org/stable/685854.

Bonnor:1965:R

[Bon65] W. B. Bonnor. Reviews. British Journal for the Philosophy of Science, XV(60):349–352, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/60/349.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686547.

Boniolo:1992:TEC

[Bon92] Giovanni Boniolo. Theory and experiment. the case of Eötvös’ experiments. British Journal for the Philosophy of Science, 43(4): 459–486, October 1992. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        43/4/459.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687836.

Bontly:1998:INS

[Bon98] Thomas Bontly. Individualism and the nature of syntactic states. British Journal for the Philosophy of Science, 49(4): 557–574, December 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        49/4/557.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688131.

Bontly:2000:BRB

[Bon00] Thomas Bontly. Book review:

John Bickle — Psychoneural Reduction: the New Wave. British Journal for the Philosophy of Science, 51(4):901–905, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/4/901.full.pdf+
html;   http://www.jstor.
org/stable/3541734.

BonJour:2001:BRM

[Bon01] Laurence BonJour. Book review:

Michael Depaul and William Ramsey (eds), Rethinking Intuition: the Psychology of Intuition and its Role in Philosophical Inquiry. British Journal for the Philosophy of Science, 52 (1):151–158, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/151.full.pdf+html;
http://www.jstor.org/
stable/3541949.

Boon:1979:R

[Boo79] Louis Boon. Reviews. British Journal for the Philosophy of Science, 30(2):195–199, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/195.full.pdf+
html;   http://www.jstor.
org/stable/686954.

Boone:2024:RM

[Boo24] Trey Boone. Robustness and modularity. British Journal for the Philosophy of Science, 75(2): 417–442, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Born:1953:IQM

[Bor53]  Max Born.  The interpre-

tation of quantum mechanics. British Journal for the Philosophy of Science, IV(14):95–106, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IV/14/95.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685986;
                                        http://www.jstor.org/
                                        stable/pdfplus/685986.pdf.

Born:1954:BRB

[Bor54] Max Born. Book review: A History of the Theories of Aether and Electricity. The Modern Theories, 1900–1926, by Edmund Whittaker. British Journal for the Philosophy of Science, 5(19):261–263, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/19/261.full.pdf+html;
                                        http://www.jstor.org/

stable/685722. Special issue on The Age of the Universe.

Born:1961:MAL

[Bor61] Max Born. More about Lorentz transformation equations.

British Journal for the Philosophy of Science, XII(46):150–151, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/150.full.
pdf+html;   http://www.
jstor.org/stable/685494.

Bortolotti:2016:EBE

[Bor16] Lisa Bortolotti. Epistemic benefits of elaborated and systematized delusions in schizophrenia. British Journal for the Philosophy of Science, 67(3):879– 900, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/879.

Bostock:2006:BRS

[Bos06] Simon Bostock. Book review:

Stephen Mumford, Laws in Nature, London, Routledge, 2004 Hardback £60.00 ISBN 0-415-31128-4. British Journal for the Philosophy of Science, 57 (2):449–452, June 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/449.full.pdf+html;
http://www.jstor.org/
stable/3873475.

Botterill:1994:BFD

[Bot94] George Botterill. Beliefs, functionally discrete states, and connectionist networks: a comment on Ramsey, Stich, and Garon. British Journal for the Philosophy of Science, 45(3): 899–906, September 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/3/899.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687799.

Botterill:1996:R

[Bot96] George Botterill. Review. British Journal for the Philosophy of Science, 47(2):328–330, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/2/328.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687954.

Bourne:2004:BI

[Bou04] Craig Bourne. Becoming inflated. British Journal for the Philosophy of Science, 55(1): 107–119, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/1/107.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541836.

Bourrat:2015:HRH

[Bou15]  Pierrick Bourrat.  How to

read ‘heritability’ in the recipe approach to natural selection. British Journal for the Philosophy of Science, 66(4):883– 903, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Bigelow:1989:VC

[BP89] John Bigelow and Robert Pargetter. Vectors and change. British Journal for the Philosophy of Science, 40(3):289– 306, September 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/289.full.pdf+html;
http://www.jstor.org/
stable/687778.

Bub:1968:MPI

[BR68] Jeffrey Bub and Michael Radner. Miller’s Paradox of information. British Journal for the Philosophy of Science, 19(1):63– 67, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/1/63.full.pdf+
html;   http://www.jstor.
org/stable/686849.

Benditt:1976:NP

[BR76] T. M. Benditt and David J.

Ross. Newcomb’s ‘Paradox’. British Journal for the Philosophy of Science, 27(2):161–164, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/2/161.full.pdf+
html;   http://www.jstor.
org/stable/686167.

Bouchard:2004:FPP

[BR04] Frédéric Bouchard and Alex Rosenberg. Fitness, probability and the principles of natural selection. British Journal for the Philosophy of Science, 55(4):693–712, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/4/693.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541623.

Bruner:2020:IFP

[BR20] Justin P. Bruner and Hannah Rubin. Inclusive fitness and the problem of honest communication. British Journal for the Philosophy of Science, 71 (1):115–137, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/71/1/
                                        115/5046214.

Bradley:1955:OIC

[Bra55a] John Bradley. On the operational interpretation of classical chemistry. British Journal for the Philosophy of Science, VI(21):32–42, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/21/32.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685572.

Bradley:1955:R

[Bra55b] John Bradley. Reviews. British Journal for the Philosophy of Science, VI(23):253–254, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/253.full.pdf+
html;   http://www.jstor.
org/stable/685686.

Brain:1960:SSD

[Bra60] W. Russell Brain. Space and sense-data. British Journal for the Philosophy of Science, XI (43):177–191, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/177.full.pdf+html;
http://www.jstor.org/
stable/685308.

Bradley:1962:DIM

[Bra62a] R. D. Bradley. Determinism or indeterminism in microphysics. British Journal for the Philosophy of Science, XIII(51):193– 215, November 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/51/193.full.pdf+html;
http://www.jstor.org/
stable/685838.

Brain:1962:RPS

[Bra62b]  Brain.  Reply to Professor

Smythies. British Journal for the Philosophy of Science, XIII (50):165–167, August 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/50/165.full.pdf+html;
 http://www.jstor.org/
stable/685969.

Bradley:1963:DPF

[Bra63a] John Bradley. Discussion of Professor F. A. Paneth’s second article. British Journal for the Philosophy of Science, XIV(53):39– 40, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/39.full.pdf+
html;   http://www.jstor.
org/stable/685350.

Bradley:1963:DPP

[Bra63b] John Bradley. Discussion of Professor Paneth’s article. British Journal for the Philosophy of Science, XIII(52):316–317, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/316.full.
pdf+html;   http://www.
jstor.org/stable/685331.

Bradley:1965:RU

[Bra65a] John Bradley. Review: [untitled]. British Journal for the Philosophy of Science, XVI (61):65–66, May 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686146.

Bradley:1965:R

[Bra65b] John Bradley. Reviews. British Journal for the Philosophy of Science, XVI(61):65b–66, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
XVI/61/65-b.full.pdf+html.

Braithwaite:1970:BRP

[Bra70] R. B. Braithwaite. Bertrand Russell as philosopher of science. British Journal for the Philosophy of Science, 21(2):129–132, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/129.full.pdf+
html;   http://www.jstor.
org/stable/686694.

Bradie:1976:WPH

[Bra76] Michael Bradie. On writing Popperian history. British Journal for the Philosophy of Science, 27(4):398, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/398.full.pdf+html;
http://www.jstor.org/
stable/686867.

Bracken:1979:R

[Bra79] Harry M. Bracken. Reviews.

British Journal for the Philosophy of Science, 30(3):297– 301, September 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/3/297.full.pdf+html;
http://www.jstor.org/
stable/687144.

Bracken:1980:R

[Bra80] Harry M. Bracken. Reviews. British Journal for the Philosophy of Science, 31(3):307– 310, September 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/3/307.full.pdf+html;
http://www.jstor.org/
stable/686926.

Bracken:1984:MLE

[Bra84] Harry M. Bracken. Mind and language: essays on Descartes and Chomsky. Foris Publications, Dordrecht, Holland, 1984. ISBN 90-6765-020-X (paper), 90-6765-021-8 (bound). xiv + 154 pp. LCCN P106 .B721 1984.

Brady:1993:JMK

[Bra93] Michael Emmett Brady. J. M. Keynes’s theoretical approach to decision-making under conditions of risk and uncertainty. British Journal for the Philosophy of Science, 44(2):357–376, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/44/2/357.full.pdf+
html;   http://www.jstor.
org/stable/687654.

Bradley:1999:BRB

[Bra99] R. Bradley. Book review: Explorations in Economic Methodology: from Lakatos to Empirical Philosophy of Science. R. Backhouse. British Journal for the Philosophy of Science, 50 (2):316–318, June 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/316.full.pdf+html;
http://www.jstor.org/
stable/40072230.

Bradley:2011:CBW

[Bra11] D. J. Bradley. Confirmation in a branching world: the Everett interpretation and Sleeping Beauty. British Journal for the Philosophy of Science, 62 (2):323–342, June 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/323.full.pdf+html.

Bradley:2015:ECS

[Bra15] Darren Bradley. Everettian confirmation and Sleeping Beauty: Reply to Wilson. British Journal for the Philosophy of Science, 66(3):683–693, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        66/3/683.

Bradley:2017:CRT

[Bra17] Seamus Bradley. Constraints on rational theory choice. British Journal for the Philosophy of Science, 68(3):639–661, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/639/2669744.

Bradley:2021:SEO

[Bra21] Darren Bradley. Should explanations omit the details? British Journal for the Philosophy of Science, 71(3):827–853, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bradley:2022:NEE

[Bra22]  Clara  Bradley.

The non-equivalence of Einstein and Lorentz. British Journal for the Philosophy of Science, 72(4): 1039–1059, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Bridgman:1951:NSOb

[Bri51a] P. W. Bridgman. The nature of some of our physical concepts — II. British Journal for the Philosophy of Science, II(5):25– 44, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/5/25.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685606.

Bridgman:1951:NSOa

[Bri51b] P. W. Bridgman. The nature of some of our physical concepts: I. British Journal for the Philosophy of Science, I(4): 257–272, February 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/4/257.full.pdf+html;
http://www.jstor.org/
stable/685095.

Bridgman:1951:NSOc

[Bri51c] P. W. Bridgman. The nature of some of our physical concepts III. British Journal for the Philosophy of Science, II(6):142–160, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/142.full.pdf+
html;   http://www.jstor.
org/stable/685507.

BritoDaCunha:1971:R

[Bri71]  A. Brito Da Cunha.  Re-

views. British Journal for the Philosophy of Science, 22(1): 62–63, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/62.full.pdf+html;
http://www.jstor.org/
stable/686649.

Briskman:1982:LLB

[Bri82] Larry Briskman. From logic to logics (and back again).

British Journal for the Philosophy of Science, 33(1):77–94, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/1/77.full.pdf+
html;   http://www.jstor.
org/stable/687243.

Brighouse:1997:DM

[Bri97] Carolyn Brighouse. Determinism and modality. British Journal for the Philosophy of Science, 48(4):465–481, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/4/465.full.pdf+html;
http://www.jstor.org/
stable/688095.

Bricker:2006:BRR

[Bri06] Phillip Bricker. Book review: Robin Le Poidevin, Travels in Four Dimensions: the Enigmas of Space and Time Oxford, Oxford University Press, 2003 Hardback £18.00 ISBN 0-19-875254-7. British Journal for the Philosophy of Science, 57 (2):453–458, June 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/453.full.pdf+html;
http://www.jstor.org/
stable/3873476.

Brown:1951:R

[Bro51] G. Burniston Brown. Reviews. British Journal for the Philosophy of Science, II(7):261– 262, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/261.full.pdf+html;
http://www.jstor.org/
stable/685301.

Brown:1952:Ra

[Bro52a] G. Burniston Brown. Reviews.

British Journal for the Philosophy of Science, III(10):200, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/200.full.
pdf+html;   http://www.
jstor.org/stable/685561.

Brown:1952:Rb

[Bro52b] G. Burniston Brown. Reviews.

British Journal for the Philosophy of Science, III(11):275– 276, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/275.full.pdf+html;
http://www.jstor.org/
stable/685274.

Brown:1952:Rc

[Bro52c] G. Burniston Brown. Reviews.

British Journal for the Philosophy of Science, III(11):282– 283, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        III/11/282.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685280.

Brown:1954:R

[Bro54] G. Burniston Brown. Reviews. British Journal for the Philosophy of Science, V(19):268– 269, November 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/19/268.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685725.

Brown:1955:AD

[Bro55] G. Burniston Brown. Achilles and dimensions. British Journal for the Philosophy of Science, VI(23):251–252, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/23/251.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685685.

Brown:1956:R

[Bro56] John Brown. Reviews. British Journal for the Philosophy of Science, VI(24):346–347, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/24/346.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685420.

Bronowski:1958:SP

[Bro58] J. Bronowski. The scandal of philosophy. British Journal for the Philosophy of Science, VIII(32):329–334, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/329.full.
pdf+html;   http://www.
jstor.org/stable/685861.

Brookes:1962:R

[Bro62a] B. C. Brookes. Reviews. British Journal for the Philosophy of Science, XIII(50):180–181, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/180.full.
pdf+html;   http://www.
jstor.org/stable/685973.

Brophy:1962:R

[Bro62b] Brigid Brophy. Reviews. British Journal for the Philosophy of Science, XIII(51):253–254, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/253.full.
pdf+html;   http://www.
jstor.org/stable/685845.

Brown:1964:R

[Bro64a] R. Brown. Reviews. British Journal for the Philosophy of Science, XV(58):161–164, August 1964. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/161.full.pdf+
html;   http://www.jstor.
org/stable/685259.

Brown:1964:RA

[Bro64b] Robert Brown. The revolution in anthropology. British Journal for the Philosophy of Science, XV(58):143–150, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/58/143.full.pdf+html;
http://www.jstor.org/
stable/685256.

Brookes:1965:R

[Bro65] B. C. Brookes. Reviews. British Journal for the Philosophy of Science, XVI(62):159–160, August 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/62/159.full.
pdf+html;   http://www.
jstor.org/stable/686155.

Brooke:1969:R

[Bro69] John Hedley Brooke. Reviews. British Journal for the Philosophy of Science, 20(2):183–184, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/183.full.pdf+
html;   http://www.jstor.
org/stable/686191.

Brown:1978:R

[Bro78] James M. Brown. Reviews.

British Journal for the Philosophy of Science, 29(2):201–204, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/201.full.pdf+
html;   http://www.jstor.
org/stable/687131.

Brown:1979:R

[Bro79] Harvey R. Brown. Reviews.

British Journal for the Philosophy of Science, 30(2):187–191, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/187.full.pdf+
html;   http://www.jstor.
org/stable/686952.

Brown:1984:R

[Bro84] James M. Brown. Reviews.

British Journal for the Philosophy of Science, 35(1):86–90, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/86.full.pdf+
html;   http://www.jstor.
org/stable/687564.

Brown:1986:HJC

[Bro86] Harvey R. Brown. Book review: Hendry, John [1984]: The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. D. Reidel Publishing Company. xi + 177 pp. (ISBN 9O-277-1648-X). British Journal for the Philosophy of Science, 37(4):497–506, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687374.

Brooke:1987:R

[Bro87] John Hedley Brooke. Reviews. British Journal for the Philosophy of Science, 38(3):397– 399, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/397.full.pdf+html;
http://www.jstor.org/
stable/687495.

Brown:1990:CSR

[Bro90] Harold I. Brown. Cherniak on scientific realism. British Journal for the Philosophy of Science, 41(3):415–427, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/3/415.full.pdf+html;
http://www.jstor.org/
stable/687829.

Brown:1993:TLO

[Bro93] Harold I. Brown. A theoryladen observation can test the theory. British Journal for the Philosophy of Science, 44(3): 555–559, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/555.full.pdf+html;
http://www.jstor.org/
stable/688024.

Brown:1996:MQP

[Bro96]  Harvey R. Brown.  Mind-

ful of quantum possibilities. British Journal for the Philosophy of Science, 47(2):189–199, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/189.full.pdf+
html;   http://www.jstor.
org/stable/687941.

Brown:1997:PP

[Bro97] James Robert Brown. Proofs and pictures. British Journal for the Philosophy of Science, 48(2):161–180, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/2/161.full.pdf+html;
http://www.jstor.org/
stable/687743.

Brown:2002:TK

[Bro02] James Robert Brown. Thomas Kuhn. British Journal for the Philosophy of Science, 53(1): 143–149, March 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/143.full.pdf+html;
http://www.jstor.org/
stable/3541646.

Brown:2009:RQM

[Bro09a] Matthew J. Brown. Relational quantum mechanics and the determinacy problem. British Journal for the Philosophy of Science, 60(4):679–695, December 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/60/4/679.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/25592030.

Brown:2009:BRD

[Bro09b] Robin Brown. Book review: Dominic Murphy, Psychiatry in the Scientific Image. British Journal for the Philosophy of Science, 60(3):673–678, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/3/673.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592026.

Broadbent:2013:PE

[Bro13] Alex Broadbent. Philosophy of Epidemiology. New directions in the philosophy of science. Palgrave Macmillan, Basingstoke, Hampshire, 2013. ISBN 0-230-35512-9 (hardcover), 1-137-31560-1 (e-book). xxii + 203 pp. LCCN RA652 .B76 2013.

Brown:2014:WER

[Bro14] Rachael L. Brown. What evolvability really is. British Journal for the Philosophy of Science, 65(3):549–572, September 2014.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/3/549.

Broadbent:2019:HSP

[Bro19] Alex Broadbent. Health as a secondary property. British Journal for the Philosophy of Science, 70(2):609–627, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/2/
609/4102132.

Brochhagen:2020:SUU

[Bro20] Thomas Brochhagen. Signalling under uncertainty: Interpretative alignment without a common prior. British Journal for the Philosophy of Science, 71 (2):471–496, June 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
471/4671055.

Barrett:1992:ERA

[BS92] Martin Barrett and Elliott Sober. Is entropy relevant to the asymmetry between retrodiction and prediction? British Journal for the Philosophy of Science, 43(2):141–160, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/2/141.full.pdf+html;
http://www.

                                        jstor.org/stable/687870. See

[Den94].

Barrett:1994:SLP

[BS94]  Martin  Barrett  and

Elliott Sober. The Second Law of Probability Dynamics. British Journal for the Philosophy of Science, 45(4):941–953, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/4/941.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687615.

Brocklehurst:2000:BRB

[BS00] A. Brocklehurst and M. Suárez. Book review: Quantum State Diffusion. I. Percival. British Journal for the Philosophy of Science, 51(3):527–530, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/3/527.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541782.

Barrett:2017:SAG

[BS17a] Jeffrey A. Barrett and Brian Skyrms. Self-assembling games. British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/2/329/3044196/
                                        Self-assembling-Games.

Bradley:2017:CD

[BS17b] Richard Bradley and H. Orri Stefánsson. Counterfactual desirability. British Journal for the Philosophy of Science, 68(2):68– ??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/2/485/3063100/
Counterfactual-
Desirability.

Barrett:2019:SAN

[BSM19]  Jeffrey  A.

Barrett, Brian Skyrms, and Aydin Mohseni. Self-assembling networks. British Journal for the Philosophy of Science, 70 (1):301–325, March 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
301/4091083.

Bub:1968:HVC

[Bub68] Jeffrey Bub. Hidden variables and the Copenhagen interpretation — a reconciliation. British Journal for the Philosophy of Science, 19(3):185–210, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/185.full.pdf+
html;   http://www.jstor.
org/stable/686240.

Bub:1973:USB

[Bub73] Jeffrey Bub. Under the spell of Bohr. British Journal for the Philosophy of Science, 24 (1):78–90, March 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/1/78.full.pdf+html;
http://www.jstor.org/
stable/686672.

Bub:1979:SRQ

[Bub79] Jeffrey Bub. Some reflections on quantum logic and Schrödinger’s Cat. British Journal for the Philosophy of Science, 30(1):27–39, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/1/27.full.pdf+
html;   http://www.jstor.
org/stable/687415.

Bub:1989:PQM

[Bub89]  Jeffrey Bub.  The philos-

ophy of quantum mechanics. British Journal for the Philosophy of Science, 40(2):191–211, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/191.full.pdf+
html;   http://www.jstor.
org/stable/687510.

Bub:1994:TMC

[Bub94] Jeffrey Bub. Testing models of cognition through the analysis of brain-damaged performance. British Journal for the Philosophy of Science, 45(3): 837–855, September 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/3/837.full.pdf+html;
http://www.jstor.org/
stable/687796.

Bub:2000:IEC

[Bub00]  Jeffrey Bub.  Indeterminacy

and entanglement: The challenge of quantum mechanics. British Journal for the Philosophy of Science, 51(4):597– 615, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541609.

Buchdahl:1951:SLS

[Buc51] G. Buchdahl. Science and logic:

Some thoughts on Newton’s Second Law of Motion in classical mechanics. British Journal for the Philosophy of Science, II(7):217–235, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/217.full.pdf+html;
http://www.jstor.org/
stable/685293.

Buchdahl:1959:SSA

[Buc59] Gerd Buchdahl. Sources of scepticism in atomic theory. British Journal for the Philosophy of Science, X(38):120–134, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/38/120.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686003.

Buchdahl:1961:R

[Buc61] Gerd Buchdahl. Reviews. British Journal for the Philosophy of Science, XII(45):79– 82, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/45/79.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685547.

Buchdahl:1963:R

[Buc63] G. Buchdahl. Reviews. British Journal for the Philosophy of Science, XIV(55):262–263, November 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIV/55/262.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685244.

Buchdahl:1965:CCL

[Buc65]  Gerd Buchdahl.  Causality,

causal laws and scientific theory in the philosophy of Kant. British Journal for the Philosophy of Science, XVI(63):187– 208, November 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XVI/63/187.full.pdf+html.

Buchdahl:1972:HPN

[Buc72] Gerd Buchdahl. Hegel’s philosophy of nature. British Journal for the Philosophy of Science, 23 (3):257–266, August 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/3/257.full.pdf+html;
http://www.jstor.org/
stable/686689.

Buck:1983:R

[Buc83] Peter Buck. Reviews. British Journal for the Philosophy of Science, 34(1):88–97, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/1/88.full.pdf+
html;   http://www.jstor.
org/stable/686939.

Buckner:2016:TGM

[Buc16] Cameron Buckner. Transitional gradation in the mind: Rethinking psychological kindhood. British Journal for the Philosophy of Science, 67(4):1091– 1115, December 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1091.

Buckner:2023:BBU

[Buc23] Cameron Buckner. Black boxes or unflattering mirrors? Comparative bias in the science of machine behaviour. British Journal for the Philosophy of Science, 74 (3):681–712, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Budden:1997:GSS

[Bud97] Tim Budden. Galileo’s ship and spacetime symmetry. British Journal for the Philosophy of Science, 48(4):483–516, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/483.full.pdf+
html;   http://www.jstor.
org/stable/688096.

Bunge:1955:SACa

[Bun55a] Mario Bunge. Strife about complementarity (I). British Journal for the Philosophy of Science, VI(21):1–12, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/21/1.full.pdf+html;
http://www.jstor.org/
stable/685570.

Bunge:1955:SACb

[Bun55b] Mario Bunge. Strife about complementarity (II). British Journal for the Philosophy of Science, VI(22):141–154, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/22/141.full.pdf+html;

http://www.jstor.org/
stable/685522.

Bunge:1956:DCTa

[Bun56a] Mario Bunge. Do computers think? (I). British Journal for the Philosophy of Science, VII(26):139–148, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/26/139.full.pdf+html;
http://www.jstor.org/
stable/685697.

Bunge:1956:DCTb

[Bun56b] Mario Bunge. Do computers think? (II). British Journal for the Philosophy of Science, VII(27):212–219, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/212.full.
pdf+html;   http://www.
jstor.org/stable/685873.

Bunge:1958:MDT

[Bun58a] Mario Bunge. On multidimensional time. British Journal for the Philosophy of Science, IX(33):39, May 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/33/39.full.pdf+html;
http://www.jstor.org/
stable/686099.

Bunge:1958:Ra

[Bun58b] Mario Bunge. Reviews. British Journal for the Philosophy of Science, IX(33):74–77, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/33/74.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686110.

Bunge:1958:Rb

[Bun58c] Mario Bunge. Reviews. British Journal for the Philosophy of Science, IX(33):77–79, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/33/77.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686111.

Bunge:1960:R

[Bun60] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XI(42):153–156, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/42/153.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685591.

Bunge:1963:R

[Bun63] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XIV(53):65– 69, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/65.full.pdf+
html;   http://www.jstor.
org/stable/685355.

Bunge:1964:R

[Bun64] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XV(58):168–170, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/168.full.pdf+
html;   http://www.jstor.
org/stable/685262.

Bunge:1968:AQT

[Bun68a] Mario Bunge. Analogy in quantum theory: from insight to nonsense. British Journal for the Philosophy of Science, 18(4): 265–286, February 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/4/265.full.pdf+html;
http://www.jstor.org/
stable/686357.

Bunge:1968:R

[Bun68b] Mario Bunge. Reviews. British Journal for the Philosophy of Science, 18(4):350–352, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/4/350.full.pdf+
html;   http://www.jstor.
org/stable/686370.

Bunge:1973:R

[Bun73] Mario Bunge. Reviews. British Journal for the Philosophy of Science, 24(4):409–410, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/24/4/409.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686623.

Bunge:1986:BRT

[Bun86] Mario Bunge. Book review: Truesdell, Clifford [1984]: An Idiot’s Fugitive Essays on Science. Methods, Criticism, Training, Circumstances. Springer-Verlag. xviii + 654 pp. $58.00. British Journal for the Philosophy of Science, 37(4):520– 523, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687379.

Bunzl:2003:R

[Bun03] Martin Bunzl. Review. British Journal for the Philosophy of Science, 54(4):623–625, December 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/54/4/623.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541684.

Burgers:1954:ED

[Bur54] J. M. Burgers. Entropy and disorder. British Journal for the Philosophy of Science, V(17):70– 71, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/17/70.full.pdf+
html;   http://www.jstor.
org/stable/685954.

Bursten:2018:STB

[Bur18] Julia R. Bursten. Smaller than a breadbox: Scale and natural kinds. British Journal for the Philosophy of Science, 69(1):1– 23, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/1/2669788.

Bush:1952:R

[Bus52a] Douglas Bush. Reviews. British Journal for the Philosophy of Science, III(9):89–91, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/9/89.full.pdf+html;
http://www.jstor.org/
stable/685478.

Bush:1952:SLC

[Bus52b] Douglas Bush. Science and literary criticism. British Journal for the Philosophy of Science, III (10):195–196, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/10/195.full.pdf+html.

Butterfield:1984:RPW

[But84a] J. Butterfield. Relationism and possible worlds. British Journal for the Philosophy of Science, 35(2):101–113, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        35/2/101.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686961.

Butts:1984:KDG

[But84b] Robert E. Butts. Kant and the double government methodology: supersensibility and method in Kant’s philosophy of science, volume 24 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, 1984. ISBN 90-277-1760-5. xvi + 339 pp. LCCN B2798 .B83 1984.

Butterfield:1989:HT

[But89] Jeremy Butterfield. The hole truth. British Journal for the Philosophy of Science, 40 (1):1–28, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687461.

Butterfield:1992:BTW

[But92] Jeremy Butterfield. Bell’s Theorem: What it takes. British Journal for the Philosophy of Science, 43(1):41–83, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/1/41.full.pdf+
html;   http://www.jstor.
org/stable/687884.

Butler:1993:CSC

[But93] Keith Butler. On Clark on systematicity and connectionism. British Journal for the Philosophy of Science, 44(1):37–44, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/37.full.pdf+
html;   http://www.jstor.
org/stable/687848.

Butterfield:1996:WM

[But96] Jeremy Butterfield. Whither the minds? British Journal for the Philosophy of Science, 47(2):200–221, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/2/200.full.pdf+html;
http://www.jstor.org/
stable/687942.

Butterfield:2002:CN

[But02] Jeremy Butterfield. Critical notice. British Journal for the Philosophy of Science, 53 (2):289–330, June 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

53/2/289.full.pdf+html;
http://www.jstor.org/
stable/3541768.

Butterfield:2006:RDA

[But06a] J. Butterfield. The rotating discs argument defeated. British Journal for the Philosophy of Science, 57(1):1–45, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/1.full.pdf+html;
http://www.jstor.org/
stable/3541651.

Butterfield:2006:APA

[But06b] Jeremy Butterfield. Against pointillisme about mechanics. British Journal for the Philosophy of Science, 57(4):709– 753, December 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/4/709.full.pdf+html;
http://www.jstor.org/
stable/4489092.

Butterfield:2007:SEL

[But07] Jeremy Butterfield. Stochastic Einstein locality revisited. British Journal for the Philosophy of Science, 58(4):805– 867, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/4/805.full.pdf+html;
http://www.jstor.org/
stable/30115175.

Button:2009:SCT

[But09] Tim Button. Sad computers and two versions of the Church– Turing thesis. British Journal for the Philosophy of Science, 60(4):765–792, December 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/765.full.pdf+html;
http://www.jstor.org/
stable/25592034.

Buxton:1978:IJS

[Bux78] Richard Buxton. The interpretation and justification of the subjective Bayesian approach to statistical inference. British Journal for the Philosophy of Science, 29(1):25–38, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/1/25.full.pdf+html;
http://www.jstor.org/
stable/686389.

Bock:1963:TET

[BV63] Walter J. Bock and Gerd Von Wahlert. Two evolutionary theories — a discussion. British Journal for the Philosophy of Science, XIV(54):140–146, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/54/140.full.
pdf+html;   http://www.
jstor.org/stable/685431.

Bertuglia:2005:NCC

[BV05] Cristoforo Sergio Bertuglia and Franco Vaio. Nonlinearity, chaos, and complexity: the dynamics of natural and social systems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856790-1 (hardcover), 0-19-856791-X (paperback). xv + 387 pp. LCCN Q172.5.C45 2005.

Ben-Yami:2006:CTO

[BY06] Hanoch Ben-Yami. Causality and temporal order in Special Relativity. British Journal for the Philosophy of Science, 57(3): 459–479, September 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        57/3/459.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3873479.

Byerly:1969:MSM

[Bye69] Henry Byerly. Model-structures and model-objects. British Journal for the Philosophy of Science, 20(2):135–144, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/2/135.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686175.

Byrd:2000:BRR

[Byr00] M. Byrd. Book review: Russell’s Hidden Substitutional Theory. G. Landini. British Journal for the Philosophy of Science, 51(2):357–362, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/2/357.full.pdf+html;
http://www.jstor.org/
stable/3541813.

Byron:2007:WPB

[Byr07] Jason M. Byron. Whence philosophy of biology? British Journal for the Philosophy of Science, 58(3):409–422, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/3/409.full.pdf+
html;   http://www.jstor.
org/stable/30115182.

Ben-Zeev:1984:PAS

[BZ84] Aaron Ben-Zeev. The passivity assumption of the sensation– perception distinction. British Journal for the Philosophy of Science, 35(4):327–343, December 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/4/327.full.pdf+
html;   http://www.jstor.
org/stable/687337.

C:1956:FST

[C.56a] A. C. C. Frank Sherwood Taylor. British Journal for the Philosophy of Science, VII(26): 183b–183, August 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/26/183-b.full.pdf+html.

C:1956:GAL

[C.56b] A. C. C. George Alfred Leon Sarton. British Journal for the Philosophy of Science, VII(26): 183a–183, August 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/26/183-a.full.pdf+html.

C:1956:OFS

[C.56c] A. C. C. Obituary: Frank Sherwood Taylor. British Journal for the Philosophy of Science, VII(26):183–184, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685709.

C:1956:OGA

[C.56d] A. C. C. Obituary: George Alfred Leon Sarton. British Journal for the Philosophy of Science, VII(26):183, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685708.

Cabrera:2020:DIR

[Cab20] Frank Cabrera. Does IBE require a ‘model’ of explanation? British Journal for the Philosophy of Science, 71(2):727–750, June 2020. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/71/2/727/4815702.

Caldin:1950:SPI

[Cal50] E. F. Caldin. Science and philosophy: Implications or presuppositions? British Journal for the Philosophy of Science, I(3): 196–210, November 1950. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/3/196.full.pdf+html;
http://www.jstor.org/
stable/685367.

Caldin:1952:R

[Cal52] E. F. Caldin. Reviews. British Journal for the Philosophy of Science, II(8):330–331, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/330.full.pdf+
html;   http://www.jstor.
org/stable/686028.

Caldin:1958:R

[Cal58] E. F. Caldin. Reviews. British Journal for the Philosophy of Science, IX(35):245–246, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/35/245.full.pdf+
html;   http://www.jstor.
org/stable/685656.

Caldin:1959:TDC

[Cal59] E. F. Caldin. Theories and the development of chemistry. British Journal for the Philosophy of Science, X(39):209– 222, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        X/39/209.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686054.

Caldin:1967:R

[Cal67] E. F. Caldin. Reviews. British Journal for the Philosophy of Science, 18(1):83–84, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/1/83.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686901.

Callinicos:1982:RU

[Cal82a] A. T. Callinicos. Review: [untitled]. British Journal for the Philosophy of Science, 33 (1):97–101, March 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687245.

Callinicos:1982:R

[Cal82b]  A.  T.  Call-

inicos. Reviews. British Journal for the Philosophy of Science, 33(1):97–100, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/1/97.full.pdf+html.

Callender:1995:MTR

[Cal95] Craig Callender. The metaphysics of time reversal: Hutchison on classical mechanics. British Journal for the Philosophy of Science, 46(3):331–340, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/3/331.full.pdf+
html;   http://www.jstor.
org/stable/687660.

Callender:1997:R

[Cal97]  Craig Callender.  Reviews.

British Journal for the Philosophy of Science, 48(1):117–120, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/117.full.pdf+
html;   http://www.jstor.
org/stable/687723.

Callender:1998:R

[Cal98a] Craig Callender. Review. British Journal for the Philosophy of Science, 49(2):332–337, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/332.full.pdf+
html;   http://www.jstor.
org/stable/688120.

Callender:1998:VNW

[Cal98b] Craig Callender. The view from no-when. British Journal for the Philosophy of Science, 49 (1):135–159, March 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        49/1/135.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688147.

Callender:2004:MEP

[Cal04] Craig Callender. Measures, explanations and the past: Should ‘special’ initial conditions be explained? British Journal for the Philosophy of Science, 55 (2):195–217, June 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/2/195.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541688.

Calcott:2009:LEE

[Cal09] Brett Calcott. Lineage explanations: Explaining how biological mechanisms change. British Journal for the Philosophy of Science, 60(1):51–78, March 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/60/1/51.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/25591988.

Callender:2015:AND

[Cal15] Craig Callender. Book review:

Alyssa Ney and David Z. Albert, The Wave Function: Essays on the Metaphysics of Quantum Mechanics. British Journal for the Philosophy of Science, 66(4): 1025–1028, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Callender:2023:QMK

[Cal23]  Craig Callender.  Quantum

mechanics: Keeping it real? British Journal for the Philosophy of Science, 74(4):837–851, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Campaner:2000:BRM

[Cam00] Raffaella Campaner. Book review: Maria Carla Galavotti and Alessandro Pagnini (eds)

— Experience, Reality, and Scientific Explanation: Essays in Honour of Merrilee and Wesley Salmon. British Journal for the Philosophy of Science, 51(4):941–945, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/941.full.pdf+html;
http://www.jstor.org/
stable/3541741.

Canfield:1964:TEB

[Can64] John Canfield. Teleological explanation in biology. British Journal for the Philosophy of Science, XIV(56):285–295, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/56/285.full.
pdf+html;   http://www.
jstor.org/stable/685635.

Canfield:1965:TEB

[Can65] John Canfield. Teleological explanation in biology: a reply. British Journal for the Philosophy of Science, XV(60): 327b–331, February 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/60/327-b.full.pdf+html;
 http://www.jstor.org/
stable/686540.

Capozzi:1986:BRK

[Cap86] Mirella Capozzi. Book review: Butts, Robert E. [1984]: Kant and the Double Government Methodology. D. Reidel. Pp. xvi + 339. (Hardback £20.25) (ISBN 90-277-1760-5). British Journal for the Philosophy of Science, 37(3):371–375, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687298.

Carter:1952:CC

[Car52] G. S. Carter. Concepts out of context. British Journal for the Philosophy of Science, III(9):86– 87, May 1952. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/86.full.pdf+
html;   http://www.jstor.
org/stable/685476.

Carnap:1953:CCC

[Car53] Rudolf Carnap. On the comparative concept of confirmation. British Journal for the Philosophy of Science, III(12): 311–318, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/311.full.pdf+html;
http://www.jstor.org/
stable/685445.

Carnap:1956:CDC

[Car56] Rudolf Carnap. Content and degree of confirmation: Remarks on Popper’s note on content and degree of confirmation. British Journal for the Philosophy of Science, VII(27): 243–244, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/243.full.pdf+html;
http://www.jstor.org/
stable/685880.

Carter:1964:TET

[Car64] G. S. Carter. Two evolutionary theories, by M. Grene; a further discussion. British Journal for the Philosophy of Science, XIV(56):345–349, February 1964. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIV/56/345.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685639.

Cargile:1966:R

[Car66] James Cargile. Reviews. British Journal for the Philosophy of Science, XVI(64):344–346, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/64/344.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686838.

Cargile:1967:RU

[Car67a] James Cargile. Review: [untitled]. British Journal for the Philosophy of Science, 17 (4):347, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686785.

Cargile:1967:R

[Car67b] James Cargile. Reviews. British Journal for the Philosophy of Science, 17(4):347a–347, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        17/4/347-a.full.pdf+html.

Cargile:1969:SP

[Car69] James Cargile. The Sorites Paradox. British Journal for the Philosophy of Science, 20(3): 193–202, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/3/193.full.pdf+html;
http://www.jstor.org/
stable/686414.

Cargile:1975:NP

[Car75] James Cargile. Newcomb’s Paradox. British Journal for the Philosophy of Science, 26(3):234– 239, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/234.full.pdf+html;
http://www.jstor.org/
stable/686281.

Carr:1976:R

[Car76] Brian Carr. Reviews. British Journal for the Philosophy of Science, 27(4):403–405, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/4/403.full.pdf+
html;   http://www.jstor.
org/stable/686869.

Carr:1977:R

[Car77] Brian Carr. Reviews. British Journal for the Philosophy of Science, 28(3):287–289, September 1977. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/3/287.full.pdf+
html;   http://www.jstor.
org/stable/686811.

Cartwright:1983:R

[Car83] Nancy Cartwright. Reviews. British Journal for the Philosophy of Science, 34(3):310– 312, September 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/3/310.full.pdf+html;
http://www.jstor.org/
stable/687330.

Carloye:1985:NSE

[Car85] Jack C. Carloye. Normal science and the extension of theories. British Journal for the Philosophy of Science, 36(3):241– 256, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/241.full.pdf+html;
http://www.jstor.org/
stable/687569.

Carroll:1997:R

[Car97] John W. Carroll. Review. British Journal for the Philosophy of Science, 48(4):625–627, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/625.full.pdf+

html;   http://www.jstor.
org/stable/688109.

Carruthers:1998:R

[Car98]  Peter Carruthers.  Review.

British Journal for the Philosophy of Science, 49(1):189–193, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/1/189.full.pdf+
html;   http://www.jstor.
org/stable/688153.

Carroll:1999:TDD

[Car99a] John W. Carroll. The two dams and that damned paresis. British Journal for the Philosophy of Science, 50(1):65–81, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/1/65.full.pdf+
html;   http://www.jstor.
org/stable/40072208.

Carruthers:1999:BRJ

[Car99b] P. Carruthers. Book review:

José Luis Bermúdez. British Journal for the Philosophy of Science, 50(3):483–486, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/3/483.full.pdf+
html;   http://www.jstor.
org/stable/40072246.

Carruthers:2002:HCC

[Car02a] Peter Carruthers. Human creativity: its cognitive basis, its evolution, and its connections with childhood pretence. British Journal for the Philosophy of Science, 53(2):225–249, June 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/2/225.full.pdf+
html;   http://www.jstor.
org/stable/3541765.

Cartwright:2002:AMC

[Car02b] Nancy Cartwright. Against modularity, the causal Markov condition, and any link between the two: Comments on Hausman and Woodward. British Journal for the Philosophy of Science, 53(3):411–453, September 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/3/411.full.pdf+
html;   http://www.jstor.
org/stable/3541889.

Carter:2003:RCC

[Car03] K. Codell (Kay Codell) Carter. The rise of causal concepts of disease: case histories. History of medicine in context. Ashgate, Aldershot, Hampshire, England, 2003. ISBN 0-7546-0678-3. ix + 237 pp. LCCN RB151 CAR 2003.

Carruthers:2004:RTW

[Car04a]  Peter Carruthers.  Review:

Thinking without words. British Journal for the Philosophy of Science, 55(4):807–810, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/4/807.full.pdf+
html;   http://www.jstor.
org/stable/3541633.

Carsetti:2004:STK

[Car04b] Arturo Carsetti, editor. Seeing, thinking and knowing: meaning and self-organization in visual cognition and thought, volume 38 of Theory and decision library. Series A, Philosophy and methodology of the social sciences. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-2080-5, 1-4020-2081-3 (e-book). 356 pp. LCCN BD161 .S385

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0821/2004045949-d.html;
http://www.loc.gov/catdir/
enhancements/fy0821/
2004045949-t.html.

Cartwright:2006:MMC

[Car06] Nancy Cartwright. From metaphysics to method: Comments on manipulability and the causal Markov condition. British Journal for the Philosophy of Science, 57(1):197–218, March 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        57/1/197.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541658.

Carroll:2012:BRJ

[Car12] John W. Carroll. Book review: John T. Roberts, The Law-Governed Universe. British Journal for the Philosophy of Science, 63(4):895–901, December 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        63/4/895.full.pdf+html.

Carruthers:2023:VIR

[Car23] Peter Carruthers. On valence: Imperative or representation of value? British Journal for the Philosophy of Science, 74(3): 533–553, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Castillejo:1950:R

[Cas50] D. Castillejo. Reviews. British Journal for the Philosophy of Science, I(2):171, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/2/171.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685810.

Cassirer:1957:R

[Cas57] Eva Cassirer. Reviews. British

Journal  for  the  Phi-

losophy of Science, VIII(29):70– 73, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/70.full.
pdf+html;   http://www.
jstor.org/stable/685384.

Cassirer:1958:MQP

[Cas58] Eva Cassirer. Methodology and quantum physics. British Journal for the Philosophy of Science, VIII(32):334–341, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/334.full.
pdf+html;   http://www.
jstor.org/stable/685862.

Cassirer:1964:Ra

[Cas64a] Eva Cassirer. Reviews. British Journal for the Philosophy of Science, XIV(56):359–366, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/56/359.full.
pdf+html;   http://www.
jstor.org/stable/685644.

Cassirer:1964:Rb

[Cas64b] Eva Cassirer. Reviews. British Journal for the Philosophy of Science, XV(58):166–168, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/XV/58/166.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685261.

Cassirer:1969:R

[Cas69] Eva Cassirer. Reviews. British Journal for the Philosophy of Science, 19(4):359–371, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/19/4/359.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686296.

Castell:1995:R

[Cas95] Paul Castell. Review. British Journal for the Philosophy of Science, 46(4):628–630, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/4/628.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687906.

Castell:1998:CRP

[Cas98] Paul Castell. A consistent restriction of the principle of indifference. British Journal for the Philosophy of Science, 49(3): 387–395, September 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        49/3/387.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688081.

Casini:2016:CIR

[Cas16] Lorenzo Casini. Can interventions rescue Glennan’s mechanistic account of causality? British Journal for the Philosophy of Science, 67(4):1155–1183, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/4/1155.

Cavalcanti:2010:CDT

[Cav10] Eric G. Cavalcanti. Causation, decision theory, and Bell’s Theorem: a quantum analogue of the Newcomb problem. British Journal for the Philosophy of Science, 61(3):569–597, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/569.full.pdf+html.

Churchman:1969:DIS

[CB69] C. West Churchman and Bruce G. Buchanan. On the design of inductive systems: Some philosophical problems. British Journal for the Philosophy of Science, 20(4):311–323, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/311.full.pdf+
html;   http://www.jstor.
org/stable/686259.

Caulton:2012:KIL

[CB12a] Adam Caulton and Jeremy Butterfield. On kinds of indiscernibility in logic and metaphysics. British Journal for the Philosophy of Science, 63 (1):27–84, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/1/27.full.pdf+html.

Caulton:2012:SPM

[CB12b] Adam Caulton and Jeremy Butterfield. Symmetries and paraparticles as a motivation for structuralism. British Journal for the Philosophy of Science, 63(2):233–285, June 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/233.full.pdf+html.

Casini:2023:PAI

[CB23] Lorenzo Casini and Michael Baumgartner. The PC algorithm and the inference to constitution. British Journal for the Philosophy of Science, 74(2): 405–429, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Chen:2022:QMT

[CBC^{+}22] Eddy Keming Chen, Clara Bradley, Eddy Keming Chen, Alexander Bird, and Daniel Williams. Quantum mechanics in a time-asymmetric universe: On the nature of the initial quantum state. British Journal for the Philosophy of Science, 72(4):1155–1183, ????

2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Clifton:1990:NIP

[CBR90] Robert K. Clifton, Jeremy N.

Butterfield, and Michael L. G. Redhead. Nonlocal influences and possible worlds — a Stapp in the wrong direction. British Journal for the Philosophy of Science, 41(1):5–58, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/1/5.full.pdf+html;
http://www.jstor.org/
stable/688002.

Chihara:1993:BOC

[CC93] Charles Chihara and Carol Chihara. A biological objection to constructive empiricism. British Journal for the Philosophy of Science, 44(4):653–658, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/4/653.full.pdf+
html;   http://www.jstor.
org/stable/688036.

Carel:2013:HID

[CC13] Havi Carel and Rachel Valerie Cooper, editors. Health, illness and disease: philosophical essays. Acumen, Durham, UK, 2013. ISBN 1-84465-543-1, 1-84465-544-X (e-book). xi + 255 pp. LCCN R723 .H42 2013.

Crupi:2013:NAP

[CCT13] Vincenzo Crupi, Nick Chater, and Katya Tentori. New axioms for probability and likelihood ratio measures. British Journal for the Philosophy of Science, 64 (1):189–204, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/189.full.pdf+html.

Craver:2024:EA

[CDCH^{+}24] Carl F. Craver, Talia Dan-Cohen, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Thomas Müller, Tomasz Placek, Ludwig von Bertalanffy, Joseph Y. Halpern, Judea Pearl, Matteo Mossio, Cristian Saborido, Alvaro Moreno, and Torsten Wilholt. Experimental artefacts. British Journal for the Philosophy of Science, 75(1): 253–274, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Cook:2004:BRK

[CE04] Roy T. Cook and Philip A. Ebert. Book review: Kit Fine, The Limits of Abstraction Oxford, Clarendon Press, 2002, cloth £18.99/US $25.00 ISBN 0-19-924618-1. British Journal for the Philosophy of Science, 55(4):791–800, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

55/4/791.full.pdf+html;
http://www.jstor.org/
stable/3541631.

Colombo:2021:BRA

[CEH21] Matteo Colombo, Lee Elkin, and Stephan Hartmann. Being realist about Bayes, and the predictive processing theory of mind. British Journal for the Philosophy of Science, 72(1):185–220, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Crupi:2008:BCU

[CFM08] Vincenzo Crupi, Roberto Festa, and Tommaso Mastropasqua. Bayesian confirmation by uncertain evidence: a reply to Huber [2005]. British Journal for the Philosophy of Science, 59 (2):201–211, June 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/2/201.full.pdf+html;
http://www.jstor.
org/stable/40072277. See re-

ply [Hub08].

Culbertson:2009:LBS

[CG09] Jennifer Culbertson and Steven Gross. Are linguists better subjects? British Journal for the Philosophy of Science, 60(4): 721–736, December 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/721.full.pdf+html;

http://www.jstor.org/
stable/25592032.

Clifton:2001:RQR

[CH01] Rob Clifton and Hans Halvorson. Are Rindler quanta real? Inequivalent particle concepts in quantum field theory. British Journal for the Philosophy of Science, 52(3):417–470, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/3/417.full.pdf+
html;   http://www.jstor.
org/stable/3541925.

Colombo:2017:BCS

[CH17] Matteo Colombo and Stephan Hartmann. Bayesian cognitive science, unification, and explanation. British Journal for the Philosophy of Science, 68(2):68– ??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/2/451/3044190/
Bayesian-Cognitive-Science-
Unification-and.

Chari:1957:NMD

[Cha57] C. T. K. Chari. A note on multi-dimensional time. British Journal for the Philosophy of Science, VIII(30):155–158, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/30/155.full.
pdf+html;   http://www.
jstor.org/stable/685794.

Chari:1958:STH

[Cha58]  C. T. K. Chari.  On the

‘space’ and ‘time’ of hallucinations. British Journal for the Philosophy of Science, VIII(32): 302–306, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/302.full.pdf+html;
http://www.jstor.org/
stable/685856.

Chari:1964:ETR

[Cha64] C. T. K. Chari. ESP and the ‘theory of resonance’. British Journal for the Philosophy of Science, XV(58):137–140, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/137.full.pdf+
html;   http://www.jstor.
org/stable/685252.

Chalmers:1970:CP

[Cha70a] A. F. Chalmers. Curie’s principle. British Journal for the Philosophy of Science, 21(2): 133–148, May 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/2/133.full.pdf+html;
http://www.jstor.org/
stable/686695.

Chalmers:1970:R

[Cha70b] A. F. Chalmers. Reviews.

British Journal for the Philosophy of Science, 21(3):307–308, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/307.full.pdf+
html;   http://www.jstor.
org/stable/686210.

Chalmers:1973:LOM

[Cha73] A. F. Chalmers. On learning from our mistakes. British Journal for the Philosophy of Science, 24(2):164–173, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/2/164.full.pdf+html;
http://www.jstor.org/
stable/686607.

Chalmers:1979:Ra

[Cha79a] A. F. Chalmers. Reviews. British Journal for the Philosophy of Science, 30(1):96–97, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/96.full.pdf+html;
http://www.jstor.org/
stable/687422.

Chalmers:1979:Rb

[Cha79b] A. F. Chalmers. Reviews. British Journal for the Philosophy of Science, 30(1):97–99, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/97.full.pdf+html;

http://www.jstor.org/
stable/687423.

Chalmers:1979:TOA

[Cha79c] Alan Chalmers. Towards an objectivist account of theory change. British Journal for the Philosophy of Science, 30(3): 227–233, September 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/3/227.full.pdf+html;
http://www.jstor.org/
stable/687137.

Chalmers:1981:PDC

[Cha81] Alan Chalmers. Planetary distances in Copernican theory. British Journal for the Philosophy of Science, 32(4):374– 375, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/374.full.pdf+html;
http://www.jstor.org/
stable/687308.

Chalmers:1982:R

[Cha82] Alan Chalmers. Reviews. British Journal for the Philosophy of Science, 33(1):101–104, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/1/101.full.pdf+
html;   http://www.jstor.
org/stable/687246.

Chalmers:1983:PDC

[Cha83] Alan Chalmers. Planetary distances and Copernican theory: a reply. British Journal for the Philosophy of Science, 34(4): 372–374, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        34/4/372.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687528.

Chalmers:1985:GTO

[Cha85a] Alan Chalmers. Galileo’s telescopic observations of Venus and Mars. British Journal for the Philosophy of Science, 36 (2):175–184, June 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/2/175.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687033.

Chalmers:1985:R

[Cha85b] Alan Chalmers. Reviews. British Journal for the Philosophy of Science, 36(2):228–230, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/228.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687041.

Champernowne:1985:R

[Cha85c] D. G. Champernowne. Reviews. British Journal for the Philosophy of Science, 36(3): 348–352, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/348.full.pdf+html;
http://www.jstor.org/
stable/687582.

Chalmers:1989:HDS

[Cha89] Alan Chalmers. How to defend science against scepticism: a reply to Barry Gower. British Journal for the Philosophy of Science, 40(2):249–253, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/249.full.pdf+
html;   http://www.jstor.
org/stable/687516.

Chalmers:1995:R

[Cha95] Alan Chalmers. Review. British Journal for the Philosophy of Science, 46(1):142–144, March 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/1/142.full.pdf+
html;   http://www.jstor.
org/stable/687638.

Chart:1996:R

[Cha96] David Chart. Review. British Journal for the Philosophy of Science, 47(3):479–482, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/3/479.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687993.

Chalmers:1997:R

[Cha97a] Alan Chalmers. Review. British Journal for the Philosophy of Science, 48(3):444–446, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/3/444.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688074.

Chang:1997:R

[Cha97b] Ha·ok Chang. Review. British Journal for the Philosophy of Science, 48(3):455–459, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/3/455.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688076.

Chart:2000:DSG

[Cha00] D. Chart. Discussion. Schulte and Goodman’s riddle. British Journal for the Philosophy of Science, 51(1):147–149, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/1/147.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541752.

Charland:2002:NKS

[Cha02a] Louis C. Charland. The natural kind status of emotion. British Journal for the Philosophy of Science, 53(4):511–537, December 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/4/511.full.pdf+
html;   http://www.jstor.
org/stable/3541822.

Chart:2002:WFE

[Cha02b] David Chart. What functions explain: Functional explanation and self-reproducing systems. British Journal for the Philosophy of Science, 53(4): 593–596, December 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/4/593.full.pdf+html;
http://www.jstor.org/
stable/3541828.

Chakravartty:2003:RUO

[Cha03] Anjan Chakravartty. The reality of the unobservable: Observability, unobservability and their impact on the issue of scientific realism. British Journal for the Philosophy of Science, 54(2):359–363, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/2/359.full.pdf+html;
http://www.jstor.org/
stable/3541975.

Chang:2004:ITM

[Cha04] Hasok Chang. Inventing temperature: measurement and scientific progress. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-517127-6. xviii + 286 pp. LCCN QC271.6 .C46

                                        2004. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0614/2003058489-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0726/
                                        2003058489-b.html.

Chakravartty:2007:MSR

[Cha07a] Anjan Chakravartty. A metaphysics for scientific realism: knowing the unobservable. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-511-36768-6 (e-book), 0-521-87649-4, 0-511-36646-9. xvii + 251 pp. LCCN Q175.32.R42 ELL 2009.

                                        URL http://www.netLibrary.
                                        com/urlapi.asp?action=
                                        summary\%26v=1\%26bookid=
                                        21441.

Chandler:2007:STP

[Cha07b] Jake Chandler. Solving the tacking problem with contrast classes. British Journal for the Philosophy of Science, 58(3): 489–502, September 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/3/489.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115185.

Chandler:2010:LPG

[Cha10] Jake Chandler. The lottery paradox generalized? British Journal for the Philosophy of Science, 61(3):667–679, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/667.full.pdf+html.

Chalmers:2011:DPL

[Cha11] Alan Chalmers. Drawing philosophical lessons from Perrin’s experiments on Brownian motion: a response to van Fraassen. British Journal for the Philosophy of Science, 62(4):711– 732, December 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/711.full.pdf+html.

Chalmers:2012:CW

[Cha12] David John Chalmers. Constructing the world. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-960857-1. xxvi + 494 pp. LCCN BD161 .C385

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1403/2012532280-b.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012532280-d.html;
http://www.loc.gov/catdir/
enhancements/fy1403/
2012532280-t.html.

Chakravartty:2013:BRS

[Cha13] Anjan Chakravartty. Book review: Stephen Mumford and Rani Lill Anjum, Getting Causes from Powers. British Journal for the Philosophy of Science, 64(4):895–899, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/4/895.full.pdf+html.

Charbonneau:2021:UCF

[Cha21] Mathieu Charbonneau. Understanding cultural fidelity. British Journal for the Philosophy of Science, 71(4):1209–1233, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Cheng:1991:CTI

[Che91] Peter C.-H. Cheng. The common thread of induction. British Journal for the Philosophy of Science, 42(2):269–272, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/42/2/269.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688059.

Cheyne:1997:R

[Che97] Colin Cheyne. Review. British Journal for the Philosophy of Science, 48(1):140–142, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/140.full.pdf+
html;   http://www.jstor.
org/stable/687727.

Chitty:1957:PSS

[Chi57] Dennis Chitty. Population studies and scientific methodology. British Journal for the Philosophy of Science, VIII(29):64–66, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/64.full.
pdf+html;   http://www.
jstor.org/stable/685382.

Chidgey:1969:R

[Chi69] J. R. Chidgey. Reviews. British Journal for the Philosophy of Science, 20(2):181–182, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/181.full.pdf+
html;   http://www.jstor.
org/stable/686190.

Chidgey:1970:R

[Chi70] J. R. Chidgey. Reviews. British Journal for the Philosophy of Science, 21(4):404–405, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/404.full.pdf+
html;   http://www.jstor.
org/stable/686572.

Chihara:1987:SPB

[Chi87] Charles S. Chihara. Some problems for Bayesian confirmation theory. British Journal for the Philosophy of Science, 38(4): 551–560, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/4/551.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687359.

Chihara:1998:WPM

[Chi98] Charles S. Chihara. The worlds of possibility: modal realism and the semantics of modal logic. Clarendon Press, New York, NY, USA, 1998. ISBN 0-19-823767-7. xii + 342 pp. LCCN BC199.M6 C48

                                        1998. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0606/98006633-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0606/
                                        98006633-t.html.

Chirimuuta:2018:ECN

[Chi18] M. Chirimuuta. Explanation in computational neuroscience: Causal and non-causal. British Journal for the Philosophy of Science, 69(3):849–880, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/3/849/3069966.

Choi:2011:ACM

[Cho11] Sungho Choi. Book review: Anjan Chakravartty, A Metaphysics for Scientific Realism: Knowing the Unobservable. British Journal for the Philosophy of Science, 62(2):443–451, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/443.full.pdf+html.

Chow:2015:MMH

[Cho15] Sheldon J. Chow. Many meanings of ‘heuristic’. British Journal for the Philosophy of Science, 66(4):977–1016, December 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Christensen:1977:HEN

[Chr77] Ferrel Christensen. How to establish non-conventional isochrony. British Journal for the Philosophy of Science, 28(1):49–54, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/1/49.full.pdf+
html;   http://www.jstor.
org/stable/686446.

Christensen:1981:SRS

[Chr81]  Ferrel Christensen.  Special

Relativity and space-like time. British Journal for the Philosophy of Science, 32(1):37–53, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/32/1/37.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687385.

Christensen:2010:DRT

[Chr10] Wayne Christensen. The decoupled representation theory of the evolution of cognition — a critical assessment. British Journal for the Philosophy of Science, 61(2):361–405, June 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/2/361.full.pdf+html.

Colombo:2015:MMC

[CHvI15] Matteo Colombo, Stephan Hartmann, and Robert van Iersel. Models, mechanisms, and coherence. British Journal for the Philosophy of Science, 66 (1):181–212, March 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/1/181.

Culp:1989:TST

[CK89] Sylvia Culp and Philip Kitcher. Theory structure and theory change in contemporary molecular biology. British Journal for the Philosophy of Science, 40(4):459–483, December 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/4/459.full.pdf+html;

http://www.jstor.org/
stable/687736.

Craver:2020:MDB

[CK20] Carl F. Craver and David M.

Kaplan. Are more details better? On the norms of completeness for mechanistic explanations. British Journal for the Philosophy of Science, 71 (1):287–319, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
287/4816342.

Cutland:1988:CNI

[CKKR88] Nigel Cutland, Christoph Kessler, Ekkehard Kopp, and David Ross. On Cauchy’s notion of infinitesimal. British Journal for the Philosophy of Science, 39(3):375–378, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/375.full.pdf+
html;   http://www.jstor.
org/stable/687214.

Crowther:2019:RFF

[CL19] Karen Crowther and Niels Linnemann. Renormalizability, fundamentality, and a final theory: The role of UV-completion in the search for quantum gravity. British Journal for the Philosophy of Science, 70(2):377–406, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/2/377/4101972.

Clarke:1953:GAS

[Cla53] John R. Clarke. The general adaptation syndrome in the study of animal populations. British Journal for the Philosophy of Science, III(12): 350–352, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/350.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685451.

Clack:1966:CMC

[Cla66] Robert J. Clack. Can a machine be conscious?: Discussion of Dennis Thompson. British Journal for the Philosophy of Science, 17(3):232–234, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/3/232.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686759.

Clarke:1972:TMM

[Cla72]  J. J. Clarke.  Turing ma-

chines and the mind-body problem. British Journal for the Philosophy of Science, 23(1): 1–12, February 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/1/1.full.pdf+html;

http://www.jstor.org/
stable/686217.

Clark:1976:EHC

[Cla76]  Peter  Clark.

Elkana on Helmholtz and conservation of energy. British Journal for the Philosophy of Science, 27(2):165–176, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/2/165.full.pdf+html;
http://www.jstor.org/
stable/686168.

Clark:1982:MMI

[Cla82a] Peter Clark. Matter, motion and irreversibility. British Journal for the Philosophy of Science, 33(2):165–185, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/2/165.full.pdf+html.

Clark:1982:RMM

[Cla82b] Peter Clark. Review: Matter, motion and irreversibility. British Journal for the Philosophy of Science, 33(2):165–186, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/687277.

Clark:1987:R

[Cla87] Peter Clark. Reviews. British Journal for the Philosophy of Science, 38(1):122–126, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/1/122.full.pdf+
html;   http://www.jstor.
org/stable/687088.

Clark:1990:CCE

[Cla90] Andy Clark. Connectionism, competence, and explanation. British Journal for the Philosophy of Science, 41(2):195–222, June 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/2/195.full.pdf+
html;   http://www.jstor.
org/stable/687772.

Clark:2000:E

[Cla00a] Peter Clark. Editorial. British Journal for the Philosophy of Science, 51(4):559, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/3541607.

Clark:2000:RBJ

[Cla00b] Peter Clark. Referees for the British Journal for the Philosophy of Science (1 June 1998– 1 June 2000). British Journal for the Philosophy of Science, 51(4):963–966, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/963.full.pdf+html.

Clarke:2001:DTE

[Cla01] Steve Clarke. Defensible territory for entity realism. British Journal for the Philosophy of Science, 52(4):701–722, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/4/701.full.pdf+
html;   http://www.jstor.
org/stable/3541915.

Clark:2007:WRT

[Cla07] Andy Clark. What reaching teaches: Consciousness, control, and the inner zombie. British Journal for the Philosophy of Science, 58(3):563–594, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/3/563.full.pdf+
html;   http://www.jstor.
org/stable/30115188.

Clarke:2014:BRJ

[Cla14] Ellen Clarke. Book review: John Dupré, Processes of Life: Essays in the Philosophy of Biology. British Journal for the Philosophy of Science, 65(1): 173–177, January 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/173.full.pdf+html.

Clarke:2016:MLS

[Cla16] Christopher Clarke. Multi-level selection and the explanatory value of mathematical decompositions. British Journal for the Philosophy of Science, 67 (4):1025–1055, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1025.

Clement:1955:SH

[Cle55] W. C. Clement. Seeing and hearing. British Journal for the Philosophy of Science, VI(21):61–63, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/21/61.full.pdf+
html;   http://www.jstor.
org/stable/685576.

Cleave:1970:NVL

[Cle70a] J. P. Cleave. The notion of validity in logical systems with inexact predicates. British Journal for the Philosophy of Science, 21 (3):269–274, August 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/3/269.full.pdf+html;
http://www.jstor.org/
stable/686197.

Cleave:1970:R

[Cle70b] John P. Cleave. Reviews. British Journal for the Philosophy of Science, 21(4):403–404, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/403.full.pdf+
html;   http://www.jstor.
org/stable/686571.

Cleave:1971:CCC

[Cle71] John P. Cleave. Cauchy, convergence and continuity. British Journal for the Philosophy of Science, 22(1):27–37, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/27.full.pdf+html;
http://www.jstor.org/
stable/686641.

Cleave:1979:CVN

[Cle79] John P. Cleave. The concept of ‘variable’ in Nineteenth Century analysis. British Journal for the Philosophy of Science, 30(3):266–278, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/3/266.full.pdf+html;
http://www.jstor.org/
stable/687141.

Cleland:2011:PEH

[Cle11] Carol E. Cleland. Prediction and explanation in historical natural science. British Journal for the Philosophy of Science, 62(3):551–582, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/551.full.pdf+html.

Clifton:1995:IMK

[Cli95]  Rob Clifton.  Independently

motivating the Kochen–Dieks modal interpretation of quantum mechanics. British Journal for the Philosophy of Science, 46(1):33–57, March 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/1/33.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687632.

Clifton:1996:PMI

[Cli96] Rob Clifton. The properties of modal interpretations of quantum mechanics. British Journal for the Philosophy of Science, 47(3):371–398, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        47/3/371.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687984.

Clowes:1966:R

[Clo66] Max Clowes. Reviews. British Journal for the Philosophy of Science, 17(2):165–167, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/2/165.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686354.

Cohen:1973:HNN

[CM73] Barry Cohen and Edward H.

Madden. Harré and nonlogical necessity. British Journal for the Philosophy of Science, 24(2):176–182, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/2/176.full.pdf+html;
http://www.jstor.org/
stable/686609.

Clifton:1999:DLY

[CM99] R. Clifton and B. Monton. Discussion. losing your marbles in wavefunction collapse theories. British Journal for the Philosophy of Science, 50(4):697– 717, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/697.full.pdf+html;
http://www.jstor.org/
stable/40072262.

Clifton:2000:DCM

[CM00] R. Clifton and B. Monton. Discussion. counting marbles with ’accessible’ mass density: a reply to Bassi and Ghirardi. British Journal for the Philosophy of Science, 51(1):155–164, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/1/155.full.pdf+
html;   http://www.jstor.
org/stable/3541754.

Campbell-Moore:2022:AUR

[CMS22] Catrin Campbell-Moore and Bernhard Salow. Accurate updating for the risk-sensitive. British Journal for the Philosophy of Science, 73(3):751–776, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Coats:1974:SDE

[Coa74] A. W. Coats. Situational determinism in economics: the implications of Lastis’s argument for the historian of economics. British Journal for the Philosophy of Science, 25(3): 285–288, September 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/3/285.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686910.

Coddington:1975:CSB

[Cod75] Alan Coddington. Creaking semaphore and beyond: a consideration of Shackle’s ‘epistemics and economics’. British Journal for the Philosophy of Science, 26(2):151–163, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/2/151.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686515.

Coffey:2014:ECT

[Cof14] Kevin Coffey. Editor’s choice: Theoretical equivalence as interpretative equivalence. British Journal for the Philosophy of Science, 65(4):821–844, December 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/65/4/821.

Cohen:1963:AIS

[Coh63a] Percy S. Cohen. The aims and interests of sociology. British Journal for the Philosophy of Science, XIV(55):246–261, November 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/55/246.full.
pdf+html;   http://www.
jstor.org/stable/685243.

Cohen:1963:R

[Coh63b] Percy S. Cohen. Reviews.

British Journal for the Philosophy of Science, XIV(53):72–73, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/72.full.pdf+
html;   http://www.jstor.
org/stable/685358.

Cohen:1966:LESa

[Coh66a] L. Jonathan Cohen. A logic for evidential support. British Journal for the Philosophy of Science, 17(1):21–43, May 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686402.

Cohen:1966:LESb

[Coh66b] L. Jonathan Cohen. A logic for evidential support (II). British Journal for the Philosophy of Science, 17(2):105–126, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/2/105.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686343.

Cohen:1967:C

[Coh67a] L. Jonathan Cohen. Corrigendum. British Journal for the Philosophy of Science, 17(4): 351s–352, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/4/351-s.full.pdf+html.

Cohen:1967:CLE

[Coh67b] L. Jonathan Cohen. Corrigendum: a logic for evidential support. British Journal for the Philosophy of Science, 17 (4):352, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686787.

Cohen:1967:R

[Coh67c] L. Jonathan Cohen. Reviews. British Journal for the Philosophy of Science, 18(1):70–72, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/70.full.pdf+
html;   http://www.jstor.
org/stable/686893.

Cohen:1968:NC

[Coh68] L. Jonathan Cohen. A note on consilience. British Journal for the Philosophy of Science, 19(1):70–71, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/1/70.full.pdf+html;
http://www.jstor.org/
stable/686851.

Cohen:1970:R

[Coh70] L. Jonathan Cohen. Reviews.

British Journal for the Philosophy of Science, 21(2):203–206, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/203.full.pdf+
html;   http://www.jstor.
org/stable/686703.

Cohen:1972:R

[Coh72] L. Jonathan Cohen. Reviews.

British Journal for the Philosophy of Science, 23(3):292–294, August 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/3/292.full.pdf+
html;   http://www.jstor.
org/stable/686693.

Cohen:1973:PSE

[Coh73] L. Jonathan Cohen. Is the progress of science evolutionary? British Journal for the Philosophy of Science, 24(1):41–61, March 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/24/1/41.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686669.

Cohen:1974:PHE

[Coh74] L. Jonathan Cohen. Professor Hull and the evolution of science. British Journal for the Philosophy of Science, 25(4): 334–336, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/4/334.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686461.

Cohen:1979:R

[Coh79] L. Jonathan Cohen. Reviews. British Journal for the Philosophy of Science, 30(3):293– 297, September 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/3/293.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687143.

Cohen:1980:BVB

[Coh80a] L. Jonathan Cohen. Bayesianism versus Baconianism in the evaluation of medical diagnoses. British Journal for the Philosophy of Science, 31(1):45–62, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/1/45.full.pdf+
html;   http://www.jstor.
org/stable/687251.

Cohen:1980:SCT

[Coh80b] L. Jonathan Cohen. Some comments on third world epistemology. British Journal for the Philosophy of Science, 31 (2):175–180, June 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/2/175.full.pdf+html;
http://www.jstor.org/
stable/687185.

Cohen:1981:R

[Coh81] L. Jonathan Cohen. Reviews.

British Journal for the Philosophy of Science, 32(1):95–98, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/1/95.full.pdf+
html;   http://www.jstor.
org/stable/687390.

Cohen:1982:WNT

[Coh82] L. Jonathan Cohen. What is necessary for testimonial corroboration? British Journal for the Philosophy of Science, 33 (2):161–164, June 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/2/161.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687276.

Cohen:1986:LID

[Coh86a] L. Jonathan Cohen. Book review: Levi, Isaac [1984]: Decisions and Revisions: Philosophical Essays on Knowledge and Value. Cambridge University Press. £27.50. Pp. xviii + 294. British Journal for the Philosophy of Science, 37 (2):252–255, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686990.

Cohen:1986:TQA

[Coh86b] L. Jonathan Cohen. Twelve questions about Keynes’s concept of weight. British Journal for the Philosophy of Science, 37(3):263–278, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687287.

Cohen:1987:NET

[Coh87] L. Jonathan Cohen. A note on the evolutionary theory of software development. British Journal for the Philosophy of Science, 38(3):381–384, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/3/381.full.pdf+
html;   http://www.jstor.
org/stable/687491.

Cohen:2002:ANE

[Coh02] Jonathan Cohen. On an alleged non-equivalence between dispositions and disjunctive properties. British Journal for the Philosophy of Science, 53(1):77–81, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/1/77.full.pdf+
html;   http://www.jstor.
org/stable/3541642.

Cohen:2016:TME

[Coh16] Michael P. Cohen. On three measures of explanatory power with axiomatic representations. British Journal for the Philosophy of Science, 67(4):1077– 1089, December 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1077.

Cole:1956:R

[Col56] D. F. Cole. Reviews. British Journal for the Philosophy of Science, VII(26):179, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/26/179.full.
pdf+html;   http://www.
jstor.org/stable/685705.

Cole:1957:R

[Col57] D. F. Cole. Reviews. British Journal for the Philosophy of Science, VIII(31):253–254, November 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/31/253.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685763.

Collins:1966:USE

[Col66] Arthur W. Collins. The use of statistics in explanation. British Journal for the Philosophy of Science, 17(2):127–140, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/2/127.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686344.

Collier:1997:R

[Col97] John Collier. Review. British Journal for the Philosophy of Science, 48(2):287–289, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/2/287.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687751.

Colyvan:1998:R

[Col98] Mark Colyvan. Review. British Journal for the Philosophy of Science, 49(4):652–656, December 1998. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/4/652.full.pdf+
html;   http://www.jstor.
org/stable/688138.

Cole:2000:BRB

[Col00] J. Cole. Book review: Foundations of Complex-System Theories in Economics, Evolutionary Biology, and Statistical Physics. S. Y. Auyang. British Journal for the Philosophy of Science, 51 (1):187–190, March 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/1/187.full.pdf+html;
http://www.jstor.org/
stable/3541758.

Collins:2007:MSE

[Col07] John Collins. Meta-scientific eliminativism: a reconsideration of Chomsky’s review of Skinner’s verbal behavior. British Journal for the Philosophy of Science, 58(4):625–658, December 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/4/625.full.pdf+
html;   http://www.jstor.
org/stable/30115168.

Colombetti:2008:SMH

[Col08] Giovanna Colombetti. The somatic marker hypotheses, and what the Iowa gambling task does and does not show. British

Journal  for  the  Phi-

losophy of Science, 59(1):51–71, March 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/1/51.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/30116993.

Colombo:2015:ECF

[Col15]  Matteo Colombo.  Editor’s

choice: For a few neurons more: Tractability and neurally informed economic modelling. British Journal for the Philosophy of Science, 66(4): 713–736, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Connell:2003:APB

[Con03] Sophia Connell. Aristotle’s philosophy of biology: Studies in the origins of life sciences. British Journal for the Philosophy of Science, 54(3):509–513, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/54/3/509.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541798.

Cooper:1965:CP

[Coo65] Neil Cooper. The concept of probability. British Journal for the Philosophy of Science, XVI (63):226–238, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
XVI/63/226.full.pdf+html;
http://www.jstor.org/
stable/686304.

Cooper:1979:R

[Coo79] Neil Cooper. Reviews. British Journal for the Philosophy of Science, 30(3):301–302, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/301.full.pdf+
html;   http://www.jstor.
org/stable/687145.

Cooper:2002:CMP

[Coo02] Rachel Valerie Cooper. Classifying madness: a philosophical examination of the diagnostic and statistical manual of mental disorders. Ph.D. thesis, Cambridge University, Cambridge, UK, 2002. vii + 172 pp.

Cook:2003:FMT

[Coo03a] Roy T. Cook. The foundations of mathematics in the theory of sets. British Journal for the Philosophy of Science, 54 (2):347–352, June 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/2/347.full.pdf+html;
http://www.jstor.org/
stable/3541973.

Cooke:2003:ISF

[Coo03b] Martin C. Cooke. Infinite sequences: Finitist consequence.

British Journal for the Philosophy of Science, 54(4):591– 599, December 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/4/591.full.pdf+html;
http://www.jstor.org/
stable/3541680.

Cooper:2004:WHW

[Coo04] Rachel Cooper. Why Hacking is wrong about human kinds. British Journal for the Philosophy of Science, 55(1):73–85, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/1/73.full.pdf+
html;   http://www.jstor.
org/stable/3541834.

Cooper:2005:CMP

[Coo05] Rachel Cooper. Classifying Madness: a Philosophical Examination of the Diagnostic and Statistical Manual of Mental Disorders, volume 86 of Philosophy and Medicine. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-4020-3345-1. 300 pp. LCCN RC455.2.C4 C68 2005.

Cook:2012:CSA

[Coo12] Roy T. Cook. Conservativeness, stability, and abstraction. British Journal for the Philosophy of Science, 63(3):673– 696, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/3/673.full.pdf+html.

Copleston:1951:R

[Cop51] Frederick C. Copleston. Reviews. British Journal for the Philosophy of Science, I(4): 328–332, February 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/4/328.full.pdf+html;
http://www.jstor.org/
stable/685099.

Copi:1980:R

[Cop80]  Irving M. Copi.  Reviews.

British Journal for the Philosophy of Science, 31(1):81–85, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/1/81.full.pdf+
html;   http://www.jstor.
org/stable/687254.

Corsi:1997:R

[Cor97] Giovanna Corsi. Review. British Journal for the Philosophy of Science, 48(2):290–291, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/290.full.pdf+
html;   http://www.jstor.
org/stable/687752.

Campbell:2011:CNJ

[COS11] Joseph Keim Campbell, Michael O’Rourke, and Matthew H. Slater, editors. Carving nature at its joints: natural kinds in metaphysics and science. Topics in contemporary philosophy. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-01593-5 (hardcover), 0-262-51626-8 (paperback), 0-262-29878-3 (e-book). viii + 355 pp. LCCN Q175.32.N38 C37

2011. URL http://mitpress-
ebooks.mit.edu/product/
carving-nature-at-its-
joints;   http://muse.jhu.
edu/books/9780262298780/.

Ciepielewski:2023:SRS

[COS23] Gerardo S. Ciepielewski, Elias Okon, and Daniel Sudarsky. On superdeterministic rejections of settings independence. British Journal for the Philosophy of Science, 74(2): 435–467, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Cotogno:2003:HPC

[Cot03] Paolo Cotogno. Hypercomputation and the physical Church– Turing thesis. British Journal for the Philosophy of Science, 54(2):181–223, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/2/181.full.pdf+html;
http://www.jstor.org/
stable/3541964.

Coulson:1952:RML

[Cou52a] C. A. Coulson. Book review:

Out of My Later Years, by Albert Einstein. British Journal for the Philosophy of Science, III(9):92– 93, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/92.full.pdf+
html;   http://links.jstor.
org/sici?sici=0007-0882%
28195205%293%3A9%3C92%
3AOOMLY%3E2.0.CO%3B2-8;
http://www.jstor.org/
stable/685480.

Coulson:1952:Ra

[Cou52b] C. A. Coulson. Reviews. British Journal for the Philosophy of Science, II(8):333–334, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/333.full.pdf+
html;   http://www.jstor.
org/stable/686031.

Coulson:1952:Rb

[Cou52c] C. A. Coulson. Reviews. British Journal for the Philosophy of Science, III(9):91–92, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/9/91.full.pdf+html;
http://www.jstor.org/
stable/685479.

Coulson:1953:RU

[Cou53a] C. A. Coulson. Review: [untitled]. British Journal for the Philosophy of Science, III(12): 382–386, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685464.

Coulson:1953:R

[Cou53b] C. A. Coulson. Reviews. British Journal for the Philosophy of Science, III(12):382b– 386, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/382-b.full.pdf+html.

Coulson:1962:RU

[Cou62a] W. F. Coulson. Review: [untitled]. British Journal for the Philosophy of Science, XIII(50): 184–185, August 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685979.

Coulson:1962:R

[Cou62b] W. F. Coulson. Reviews. British Journal for the Philosophy of Science, XIII(50):184c–185, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/184-c.full.
                                        pdf+html.

Cowling:1952:R

[Cow52]  T. G. Cowling.  Reviews.

British Journal for the Philosophy of Science, III(11):281– 282, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/281.full.pdf+html;
http://www.jstor.org/
stable/685279.

Cowie:1998:MDN

[Cow98] Fiona Cowie. Mad dog nativism.

British Journal for the Philosophy of Science, 49(2):227–252, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/227.full.pdf+
html;   http://www.jstor.
org/stable/688113.

Cox:1975:MDP

[Cox75] J. W. Roxbee Cox. Mackie on dispositional properties. British Journal for the Philosophy of Science, 26(3):232–234, September 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/3/232.full.pdf+
html;   http://www.jstor.
org/stable/686280.

Calcott:2020:SMD

[CPG20]  Brett  Calcott,  Arnaud

Pocheville, and Paul Griffiths. Signals that make a difference. British Journal for the Philosophy of Science, 71(1):233–258, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/1/233/4636564.

Cranfield:1968:R

[Cra68] R. Cranfield. Reviews. British Journal for the Philosophy of Science, 18(4):346–348, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/4/346.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686368.

Craig:1969:PG

[Cra69a] E. J. Craig. Phenomenal geometry. British Journal for the Philosophy of Science, 20 (2):121–134, August 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/2/121.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686174.

Craig:1969:Ra

[Cra69b] E. J. Craig. Reviews. British Journal for the Philosophy of Science, 20(3):272–274, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/272.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686424.

Craig:1969:Rb

[Cra69c] E. J. Craig. Reviews. British Journal for the Philosophy of Science, 20(3):274–275, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/3/274.full.pdf+
html;   http://www.jstor.
org/stable/686425.

Cranfield:1969:R

[Cra69d] R. Cranfield. Reviews. British Journal for the Philosophy of Science, 20(2):169–170, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/169.full.pdf+
html;   http://www.jstor.
org/stable/686181.

Craig:1970:Ra

[Cra70a] E. J. Craig. Reviews. British Journal for the Philosophy of Science, 21(2):211–212, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/211.full.pdf+
html;   http://www.jstor.
org/stable/686706.

Craig:1970:Rb

[Cra70b] E. J. Craig. Reviews. British Journal for the Philosophy of Science, 21(4):397–399, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/397.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/686567.

Craig:1971:R

[Cra71] E. J. Craig. Reviews. British Journal for the Philosophy of Science, 22(1):77–78, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/77.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686659.

Craig:1979:WPI

[Cra79] William Lane Craig. Whitrow and Popper on the impossibility of an infinite past. British Journal for the Philosophy of Science, 30(2):165–170, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/2/165.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686947.

Craig:1986:GCM

[Cra86] William Lane Craig. God, creation and Mr Davies. British Journal for the Philosophy of Science, 37(2):163–175, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686975.

Craig:1988:BTA

[Cra88] William Lane Craig. Barrow and Tipler on the anthropic principle vs. divine design. British Journal for the Philosophy of Science, 39(3):389–395, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/389.full.pdf+
html;   http://www.jstor.
org/stable/687216.

Craig:1990:WPT

[Cra90] William Lane Craig. ‘What place, then, for a creator?’: Hawking on God and creation. British Journal for the Philosophy of Science, 41(4):473– 491, December 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/473.full.pdf+html;
http://www.jstor.org/
stable/687861.

Craig:1992:OCU

[Cra92] William Lane Craig. The origin and creation of the universe: a reply to Adolf Grünbaum. British Journal for the Philosophy of Science, 43(2):233–240, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/233.full.pdf+
html;   http://www.jstor.
org/stable/687874.

Craig:1993:CBU

[Cra93] Wm. Lane Craig. The caused beginning of the universe: a response to Quentin Smith. British Journal for the Philosophy of Science, 44(4):623–639, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/4/623.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688034.

Crane:1996:R

[Cra96]  Tim Crane.  Review.

British Journal for the Philosophy of Science, 47(1):146–148, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/1/146.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/1215835.

Craig:2001:PGN

[Cra01]  William Lane Craig.  Prof.

Grünbaum on the ‘normalcy of nothingness’ in the Leibnizian and Kalam cosmological arguments. British Journal for the Philosophy of Science, 52 (2):371–386, June 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/2/371.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541902.

Crombie:1952:RUb

[Cro52a] A. C. Crombie. Review: [untitled]. British Journal for the Philosophy of Science, II(8): 326–329, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686026.

Crombie:1952:RUg

[Cro52b] A. C. Crombie. Review: [untitled]. British Journal for the Philosophy of Science, III(11): 279, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685277.

Crombie:1952:Ra

[Cro52c] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, II(8):322–325, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/322.full.pdf+
html;   http://www.jstor.
org/stable/686023.

Crombie:1952:Rb

[Cro52d]  A. C. Crombie.  Reviews.

British Journal for the Philosophy of Science, II(8):326b– 329, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/8/326-b.full.pdf+html.

Crombie:1952:Rc

[Cro52e] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, III(9):93–97, May 1952.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/9/93.full.pdf+html;
http://www.jstor.org/
stable/685481.

Crombie:1952:Rd

[Cro52f]  A. C. Crombie.  Reviews.

British Journal for the Philosophy of Science, III(9):99–101, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/99.full.pdf+
html;   http://www.jstor.
org/stable/685483.

Crombie:1952:Re

[Cro52g] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, III(10):201–202, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/201.full.
pdf+html;   http://www.
jstor.org/stable/685562.

Crombie:1952:Rf

[Cro52h] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, III(10):210–211, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/210.full.
pdf+html;   http://www.
jstor.org/stable/685566.

Crombie:1952:Rg

[Cro52i]  A. C. Crombie.  Reviews.

British Journal for the Philosophy of Science, III(11):279a– 279, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/279-a.full.pdf+html.

Crombie:1953:NDC

[Cro53a] A. C. Crombie. A note on the descriptive conception of motion in the Fourteenth Century. British Journal for the Philosophy of Science, IV(13):46–51, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/13/46.full.pdf+
html;   http://www.jstor.
org/stable/686045.

Crombie:1953:Ra

[Cro53b] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, III(12):374–375, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/12/374.full.
pdf+html;   http://www.
jstor.org/stable/685458.

Crombie:1953:Rb

[Cro53c] A. C. Crombie. Reviews. British Journal for the Philosophy of Science, IV(14):168–169, August 1953. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/168.full.pdf+
html;   http://www.jstor.
org/stable/685994.

Crombie:1954:R

[Cro54]  A. C. Crombie.  Reviews.

British Journal for the Philosophy of Science, V(18):173–174, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/173.full.pdf+
html;   http://www.jstor.
org/stable/685183.

Crombie:1959:R

[Cro59]  A. C. Crombie.  Reviews.

British Journal for the Philosophy of Science, X(38):164–165, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/164.full.pdf+
html;   http://www.jstor.
org/stable/686011.

Crossley:1962:R

[Cro62] J. N. Crossley. Reviews. British Journal for the Philosophy of Science, XIII(51):251–253, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/251.full.
pdf+html;   http://www.
jstor.org/stable/685844.

Crosland:1963:R

[Cro63]  M. P. Crosland.  Reviews.

British Journal for the Philosophy of Science, XIV(54):172, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/54/172.full.
pdf+html;   http://www.
jstor.org/stable/685439.

Crozier:2008:RCS

[Cro08] G. K. D. Crozier. Reconsidering cultural selection theory. British Journal for the Philosophy of Science, 59(3):455–479, September 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/3/455.full.pdf+
html;   http://www.jstor.
org/stable/40072295.

Cook:1998:HRP

[CS98] Roy Cook and Stewart Shpiro.

Hintikka’s revolution: the principles of mathematics revisited. British Journal for the Philosophy of Science, 49(2):309–316, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/309.full.pdf+
html;   http://www.jstor.
org/stable/688117.

Colombo:2012:BBB

[CS12] Matteo Colombo and Peggy Seriès. Bayes in the brainon Bayesian modelling in neuroscience. British Journal for the Philosophy of Science, 63(3): 697–723, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/3/697.full.pdf+html.

Colombo:2016:BRD

[CS16] Camilla Colombo and Katie Steele. Book review: Daniel Steel, Philosophy and the Precautionary Principle: Science, Evidence, and Environmental Policy. British Journal for the Philosophy of Science, 67 (4):1195–1200, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1195.

Cowan:2016:EWF

[CT16] Charles Wesley Cowan and Roderich Tumulka. Epistemology of wave function collapse in quantum physics. British Journal for the Philosophy of Science, 67(2):405–434, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/405.

Curtis-Trudel:2023:WDW

[CT23] André Curtis-Trudel. Why do we need a theory of implementation? British Journal for the Philosophy of Science, 73(4):

1067–1091, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Cuffaro:2017:SGK

[Cuf17] Michael E. Cuffaro. On the significance of the Gottesman–Knill Theorem. British Journal for the Philosophy of Science, 68(1):68– ??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/1/91/2995239/On-
the-Significance-of-the-
Gottesman-Knill-Theorem.

Cuffaro:2018:RNG

[Cuf18] Michael E. Cuffaro. Reconsidering no-go theorems from a practical perspective. British Journal for the Philosophy of Science, 69(3):633–655, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/633/3002298.

Cullwick:1981:ESR

[Cul81] E. G. Cullwick. Einstein and Special Relativity. British Journal for the Philosophy of Science, 32(2):167–176, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/2/167.full.pdf+html;
http://www.jstor.org/
stable/687198.

Cummiskey:1992:RFS

[Cum92] D. Cummiskey. Reference failure and scientific realism: a response to the meta-induction. British Journal for the Philosophy of Science, 43(1):21–40, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/1/21.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687883.

Currie:1975:R

[Cur75] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 26(2):180–182, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/2/180.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686519.

Currie:1976:WFL

[Cur76] Gregory Currie. Was Frege a linguistic philosopher? British Journal for the Philosophy of Science, 27(1):79–92, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/27/1/79.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686382.

Currie:1979:R

[Cur79] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 30(4):418–422, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/4/418.full.pdf+
html;   http://www.jstor.
org/stable/687550.

Currie:1981:Ra

[Cur81a] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 32(2):197–200, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/2/197.full.pdf+
html;   http://www.jstor.
org/stable/687201.

Currie:1981:Rb

[Cur81b] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 32(2):200–206, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/2/200.full.pdf+
html;   http://www.jstor.
org/stable/687202.

Currie:1982:FL

[Cur82a] Gregory Currie. Frege’s letters. British Journal for the Philosophy of Science, 33(1): 65–76, March 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/1/65.full.pdf+html.

Currie:1982:RFL

[Cur82b] Gregory Currie. Review: Frege’s letters. British Journal for the Philosophy of Science, 33 (1):65–77, March 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687242.

Currie:1982:R

[Cur82c] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 33(4):435–437, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/4/435.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687174.

Curd:1983:SCS

[Cur83] Martin Curd. The superiority of the Copernican system: a reply to Chalmers. British Journal for the Philosophy of Science, 34(4):367–369, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/4/367.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687526.

Currie:1984:IGS

[Cur84] Gregory Currie. Individualism and global supervenience. British Journal for the Philosophy of Science, 35(4):345–358, December 1984. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/4/345.full.pdf+
html;   http://www.jstor.
org/stable/687338.

Currie:1985:R

[Cur85] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 36(4):475–479, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/4/475.full.pdf+
html;   http://www.jstor.
org/stable/687409.

Curtis:1986:MTS

[Cur86] Ronald C. Curtis. Are methodologies theories of scientific rationality? British Journal for the Philosophy of Science, 37 (2):135–161, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687007.

Currie:1987:R

[Cur87] Gregory Currie. Reviews. British Journal for the Philosophy of Science, 38(1):127–132, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/1/127.full.pdf+
html;   http://www.jstor.
org/stable/687089.

Currie:2013:CE

[Cur13] Adrian Currie. Convergence as evidence. British Journal for the Philosophy of Science, 64(4):763–786, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/4/763.full.pdf+html.

Curiel:2014:CML

[Cur14] Erik Curiel. Classical mechanics is Lagrangian; it is not Hamiltonian. British Journal for the Philosophy of Science, 65 (2):269–321, June 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/2/269.

Currie:2016:BRM

[Cur16] Adrian Currie. Book review: Melinda Fagan, Philosophy of Stem Cell Biology: Knowledge in Flesh and Blood. British Journal for the Philosophy of Science, 67(2):651–655, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/2/651.

Currie:2017:HBG

[Cur17] Adrian Currie. Hot-blooded gluttons: Dependency, coherence, and method in the historical sciences. British Journal for the Philosophy of Science, 68(4):929–952, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/929/2669759.

Curiel:2018:ESS

[Cur18]  Erik Curiel.  On the exis-

tence of spacetime structure. British Journal for the Philosophy of Science, 69(2):447–483, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/447/2669781.

Currie:2021:BUR

[Cur21] Adrian Currie. Bottled understanding: The role of lab work in ecology. British Journal for the Philosophy of Science, 71(3): 905–932, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Currie:2023:EEA

[Cur23] Adrian Currie. Epistemic engagement, aesthetic value, and scientific practice. British Journal for the Philosophy of Science, 74(2):313–334, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Cushing:1990:SMI

[Cus90] James T. Cushing. Is scientific methodology interestingly atemporal? British Journal for the Philosophy of Science, 41 (2):177–194, June 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/2/177.full.pdf+html;
http://www.jstor.org/
stable/687771.

Cusbert:2022:HSO

[Cus22] John Cusbert. How stable is objective chance? British Journal for the Philosophy of Science, 73(3):613–629, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Campbell:1983:NC

[CV83]  Richmond  Camp-

bell and Thomas Vinci. Novel confirmation. British Journal for the Philosophy of Science, 34(4):315–341, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/315.full.pdf+html;
http://www.jstor.org/
stable/687523.

DAgostino:1976:RTL

[D’A76] F. B. D’Agostino. Rethinking transformational linguistics. British Journal for the Philosophy of Science, 27(3):275– 287, September 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/3/275.full.pdf+html;
http://www.jstor.org/
stable/686126.

DAgostino:1977:KL

[D’A77] F. B. D’Agostino. Knowledge of language. British Journal for the Philosophy of Science, 28(1):74–80, March 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/1/74.full.pdf+html;
http://www.jstor.org/
stable/686450.

DAgostino:1978:SD

[D’A78] F. B. D’Agostino. Sampson’s ‘Dilemma’. British Journal for the Philosophy of Science, 29 (2):183–184, June 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/2/183.full.pdf+html;
http://www.jstor.org/
stable/687126.

DAgostino:1984:R

[D’A84]  Fred D’Agostino.  Reviews.

British Journal for the Philosophy of Science, 35(4):408– 411, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/408.full.pdf+html;
http://www.jstor.org/
stable/687346.

DAgostino:1986:BHM

[D’A86] Fred D’Agostino. Book review: Bracken, Harry M. [1984]: Mind and Language: Essays on Descartes and Chomsky. Foris Publications. ISBN 90-6765-020-X. British Journal for the Philosophy of Science, 37(2): 249–251, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686989.

DAlessandro:2020:MEB

[D’A20] William D’Alessandro. Mathematical explanation beyond explanatory proof. British Journal for the Philosophy of Science, 71(2):581–603, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
581/4815701.

Dacey:2023:EDR

[Dac23] Mike Dacey. Evidence in default: Rejecting default models of animal minds. British Journal for the Philosophy of Science, 74(2): 291–312, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Dale:1979:R

[Dal79] A. J. Dale. Reviews. British Journal for the Philosophy of Science, 30(3):304–306, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/304.full.pdf+
html;   http://www.jstor.
org/stable/687147.

Dale:1989:ARL

[Dal89] A. J. Dale. Anti-realism and logic. British Journal for the Philosophy of Science, 40(2): 213–217, June 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/213.full.pdf+html;
http://www.jstor.org/
stable/687511.

Dale:1990:R

[Dal90] A. J. Dale. Review. British Journal for the Philosophy of Science, 41(4):575–578, December 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/4/575.full.pdf+
html;   http://www.jstor.
org/stable/687866.

Danks:2005:SCF

[Dan05] David Danks. Scientific coherence and the fusion of experimental results. British Journal for the Philosophy of Science, 56(4):791–807, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/791.full.pdf+html;
http://www.jstor.org/
stable/3541867.

Darlington:1954:R

[Dar54] C. D. Darlington. Reviews.

British Journal for the Philosophy of Science, V(19):271, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/19/271.full.pdf+
html;   http://www.jstor.
org/stable/685728.

Darden:1983:R

[Dar83] Lindley Darden. Reviews. British Journal for the Philosophy of Science, 34(4):399–403, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/399.full.pdf+html;
http://www.jstor.org/
stable/687534.

Darby:2012:RHH

[Dar12] George Darby. Relational holism and Humean supervenience. British Journal for the Philosophy of Science, 63(4):773–788, December 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/4/773.full.pdf+html.

Das:1958:INI

[Das58] R. Das. Induction and noninstantial hypothesis: (a review of Dr Wisdom’s views). British Journal for the Philosophy of Science, VIII(32):317– 325, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/317.full.pdf+html;
http://www.jstor.org/
stable/685859.

Dasgupta:2016:SEN

[Das16] Shamik Dasgupta. Symmetry as an epistemic notion (twice over). British Journal for the Philosophy of Science, 67(3): 837–878, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/837.

Das:2023:VBI

[Das23] Nilanjan Das. The value of biased information. British Journal for the Philosophy of Science, 74(1):25–55, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Dauben:1988:CST

[Dau88] Joseph W. Dauben. Cantorian set theory and limitations of size. British Journal for the Philosophy of Science, 39(4): 541–550, November 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/541.full.pdf+html;
http://www.jstor.org/
stable/687440.

Davies:1954:R

[Dav54a] E. Rowan Davies. Reviews.

British Journal for the Philosophy of Science, V(19):270– 271, November 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/270.full.pdf+html;
http://www.jstor.org/
stable/685727.

Davies:1954:AU

[Dav54b] J. T. Davies. The age of the universe. British Journal for the Philosophy of Science, V(19):191–202, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/191.full.pdf+html;
http://www.jstor.org/
stable/685715.

Davies:1955:R

[Dav55] E. Rowan Davies. Reviews. British Journal for the Philosophy of Science, V(20):358– 359, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/358.full.pdf+html;
http://www.jstor.org/
stable/685752.

Davies:1956:ESR

[Dav56] J. T. Davies. On extrapolation, with special reference to the ‘age of the universe’. British Journal for the Philosophy of Science, VII(26):129–138, August 1956. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/26/129.full.
pdf+html;   http://www.
jstor.org/stable/685696.

Davies:1957:R

[Dav57] J. T. Davies. Reviews. British Journal for the Philosophy of Science, VIII(29):78– 79, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/78.full.
pdf+html;   http://www.
jstor.org/stable/685387.

Davies:1959:R

[Dav59] E. Rowan Davies. Reviews.

British Journal for the Philosophy of Science, X(39):255– 256, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/255.full.pdf+html;
http://www.jstor.org/
stable/686068.

Davidson:1962:PAC

[Dav62] W. Davidson. Philosophical aspects of cosmology. British Journal for the Philosophy of Science, XIII(50):120–129, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/50/120.full.pdf+html;

                                         http://www.jstor.org/
                                        stable/685965.

Davies:1986:RU

[Dav86] Martin Davies. Review: [untitled]. British Journal for the Philosophy of Science, 37 (1):130–134, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687006.

Davies:1989:CMT

[Dav89] Martin Davies. Connectionism, modularity, and tacit knowledge. British Journal for the Philosophy of Science, 40(4): 541–555, December 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/4/541.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687740.

Davies:2001:BIM

[Dav01] E. B. Davies. Building infinite machines. British Journal for the Philosophy of Science, 52(4):671–682, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/4/671.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541913.

Davey:2003:MRN

[Dav03] Kevin Davey. Is mathematical rigor necessary in physics?

British Journal for the Philosophy of Science, 54(3):439– 463, September 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/3/439.full.pdf+html;
http://www.jstor.org/
stable/3541794.

Davies:2005:SRF

[Dav05] E. B. Davies. Some remarks on the foundations of quantum theory. British Journal for the Philosophy of Science, 56(3):521–539, September 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/521.full.pdf+html;
http://www.jstor.org/
stable/3541717.

Davies:2018:CVS

[Dav18] Will Davies. Colour vision and seeing colours. British Journal for the Philosophy of Science, 69(3):657–690, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/657/2959522.

Dawson:1965:R

[Daw65] E. E. Dawson. Reviews. British Journal for the Philosophy of Science, XV(60):341–345, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/XV/60/341.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686544.

Dawson:1970:R

[Daw70] E. E. Dawson. Reviews. British Journal for the Philosophy of Science, 21(4):407–408, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/4/407.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686574.

Dawid:1985:PSF

[Daw85] A. P. Dawid. Probability, symmetry and frequency. British Journal for the Philosophy of Science, 36(2):107–128, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/107.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687029.

Dingle:1952:SLC

[DB52] Herbert Dingle and Douglas Bush. Science and literary criticism. British Journal for the Philosophy of Science, III(10): 194–195, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/10/194.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685559.

Dizadji-Bahmani:2015:PPE

[DB15]  Foad  Dizadji-

Bahmani. The probability problem in Everettian quantum mechanics persists. British Journal for the Philosophy of Science, 66(2):257–283, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/2/257.

deCanson:2022:OMA

[dC22] Chloé de Canson. Objectivity and the method of arbitrary functions. British Journal for the Philosophy of Science, 73(3): 663–684, September 2022. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

DeMonchaux:1959:R

[De 59]  Cecily De Monchaux.  Re-

views. British Journal for the Philosophy of Science, X(39): 252–254, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/252.full.pdf+html;
http://www.jstor.org/
stable/686066.

DeMare:1972:CGF

[De 72] Jacques De Maré. A comment on Gillies’s Falsifying Rule for Probability Statements. British Journal for the Philosophy of Science, 23(4):335, November 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/335-a.full.pdf+html;
http://www.
jstor.org/stable/686734. See

[Gil71].

DeAMartins:1982:ETI

[De 82] Roberto De A.Martins. Experimental tests of isometry hypotheses. British Journal for the Philosophy of Science, 33(3): 296–304, September 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/296.full.pdf+html;
http://www.jstor.org/
stable/687228.

Dorling:1976:ABC

[DE76] Jon Dorling and Dorothy Edgington. The applicability of Bayesian convergence-of-opinion theorems to the case of actual scientific inference. British Journal for the Philosophy of Science, 27(2):160–161, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/2/160-b.full.pdf+html;
http://www.jstor.org/
stable/686166.

Dear:1961:DCP

[Dea61] G. F. Dear. Determinism in classical physics. British Journal for the Philosophy of Science, XI(44):289–304, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/289.full.pdf+html;
http://www.jstor.org/
stable/685130.

Deakin:1976:UAI

[Dea76] Michael A. B. Deakin. On Urbach’s analysis of the ‘IQ Debate’. British Journal for the Philosophy of Science, 27 (1):60–65, March 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/1/60.full.pdf+html;
http://www.jstor.org/
stable/686379.

Dellsen:2021:BEU

[Del21] Finnur Dellsén. Beyond explanation: Understanding as dependency modelling. British Journal for the Philosophy of Science, 71 (4):1261–1286, ???? 2021. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Demopoulos:2001:BRB

[Dem01] William Demopoulos. Book review: Reason’s Nearest Kin: Philosophies of Arithmetic from Kant to Carnap, Michael Potter. British Journal for the Philosophy of Science, 52(3):599– 612, September 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/599.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/3541931.

Demopoulos:2003:RRO

[Dem03] William Demopoulos. On the rational reconstruction of our theoretical knowledge. British Journal for the Philosophy of Science, 54(3):371–403, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/3/371.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541791.

Demopoulos:2011:TVT

[Dem11] William Demopoulos. Three views of theoretical knowledge. British Journal for the Philosophy of Science, 62(1):177–205, March 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        62/1/177.full.pdf+html.

Demopoulos:2013:LPL

[Dem13] William Demopoulos. Logicism and its philosophical legacy. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-02980-5. xii + 272 pp. LCCN QA9 .D38

                                        2013. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1213/2012023164-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1213/
                                        2012023164-d.html;
                                        http://www.loc.gov/catdir/

enhancements/fy1213/
2012023164-t.html.

Denbigh:1953:TSS

[Den53] K. G. Denbigh. Thermodynamics and the subjective sense of time. British Journal for the Philosophy of Science, IV(15): 183–191, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/183.full.pdf+html;
http://www.jstor.org/
stable/685895.

Denyer:1958:R

[Den58] G. E. Denyer. Reveiws. British Journal for the Philosophy of Science, VIII(32):341–342, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/341.full.
pdf+html;   http://www.
jstor.org/stable/685863.

Denyer:1959:ACP

[Den59a] G. E. Denyer. Annual conference of the philosophy of science group. British Journal for the Philosophy of Science, IX(36):341–342, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/341.full.pdf+html;
http://www.jstor.org/
stable/685201.

Denyer:1959:R

[Den59b] G. E. Denyer. Reviews. British Journal for the Philosophy of Science, IX(36):333–335, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/36/333.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685197.

Denyer:1960:ACB

[Den60a] G. E. Denyer. Annual conference of the British Society for the Philosophy of Science. British Journal for the Philosophy of Science, XI(41):86b–88, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/86-b.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685835.

Denyer:1960:R

[Den60b] G. E. Denyer. Reviews. British Journal for the Philosophy of Science, XI(41):80–81, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/80.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685829.

Dent:1969:R

[Den69] N. J. H. Dent. Reviews. British Journal for the Philosophy of Science, 20(3):283–284, October 1969. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/3/283.full.pdf+
html;   http://www.jstor.
org/stable/686430.

Denbigh:1982:R

[Den82] K. G. Denbigh. Reviews. British Journal for the Philosophy of Science, 33(3):325–329, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/3/325.full.pdf+
html;   http://www.jstor.
org/stable/687232.

Denbigh:1985:R

[Den85] Kenneth Denbigh. Reviews.

British Journal for the Philosophy of Science, 36(3):352– 354, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/352.full.pdf+html;
http://www.jstor.org/
stable/687583.

Denbigh:1986:KJB

[Den86] Kenneth Denbigh. Book review:

Kane, Jeffrey [1984]: Beyond Empiricism: Michael Polanyi Reconsidered. Peter Lang, New York. Pp. 263. (ISBN 0-8204-0118-8). British Journal for the Philosophy of Science, 37(3): 375–377, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687299.

Denbigh:1987:Ra

[Den87a] Kenneth Denbigh. Reviews. British Journal for the Philosophy of Science, 38(2):259–261, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/2/259.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687051.

Denbigh:1987:Rb

[Den87b] Kenneth Denbigh. Reviews. British Journal for the Philosophy of Science, 38(3):404– 406, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/3/404.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687497.

Denbigh:1989:MFI

[Den89a] K. G. Denbigh. The many faces of irreversibility. British Journal for the Philosophy of Science, 40(4):501–518, December 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/4/501.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687738.

Denbigh:1989:NED

[Den89b] K. G. Denbigh. Note on entropy, disorder and disorganization. British Journal for the Philosophy of Science, 40(3):323– 332, September 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/323.full.pdf+html;
http://www.jstor.org/
stable/687780.

Denbigh:1994:CBS

[Den94]  K. G. Denbigh.  Comment

on Barrett and Sober’s paper on the relevance of entropy to retrodiction and prediction. British Journal for the Philosophy of Science, 45(2):709–711, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/709.full.pdf+
html;   http://www.
jstor.org/stable/687692. See

[BS92].

Deckert:2019:PPO

[DEO19] Dirk-André Deckert, Michael Esfeld, and Andrea Oldofredi. A persistent particle ontology for quantum field theory in terms of the Dirac sea. British Journal for the Philosophy of Science, 70(3):747–770, September 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
747/4669819.

DeRose:2002:KL

[DeR02] Keith DeRose. Knowledge and its limits. British Journal for the Philosophy of Science, 53(4): 573–577, December 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        53/4/573.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541825.

Deutsch:1996:CL

[Deu96] David Deutsch. Comment on Lockwood. British Journal for the Philosophy of Science, 47 (2):222–228, June 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        47/2/222.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687943.

Devitt:1991:NR

[Dev91] Michael Devitt. Naturalistic representation. British Journal for the Philosophy of Science, 42(3):425–443, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/3/425.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687918.

DeVito:1997:GPC

[DeV97] Scott DeVito. A gruesome problem for the curve-fitting solution. British Journal for the Philosophy of Science, 48(3):391– 396, September 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/3/391.full.pdf+html;
http://www.jstor.org/
stable/688069.

Devitt:2006:IL

[Dev06] Michael Devitt. Intuitions in linguistics. British Journal for the Philosophy of Science, 57(3): 481–513, September 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/3/481.full.pdf+html;
http://www.jstor.org/
stable/3873480.

Devitt:2010:LIR

[Dev10] Michael Devitt. Linguistic intuitions revisited. British Journal for the Philosophy of Science, 61(4):833–865, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/4/833.full.pdf+html.

Dewitt:1973:HAM

[Dew73] Larry W. Dewitt. The hidden assumption in Mackay’s logical paradox concerning free will. British Journal for the Philosophy of Science, 24(4):402– 405, December 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/4/402.full.pdf+html;
http://www.jstor.org/
stable/686621.

Dewitt:1975:CMS

[Dew75] Larry W. Dewitt. Consciousness, mind, and self: the implications of the split-brain studies. British Journal for the Philosophy of Science, 26(1):41–47, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/1/41.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686331.

Dewhurst:2018:IR

[Dew18] Joe Dewhurst. Individuation without representation. British Journal for the Philosophy of Science, 69(1):103–116, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/1/103/2669787.

Dewar:2019:SAS

[Dew19] Neil Dewar. Sophistication about symmetries. British Journal for the Philosophy of Science, 70 (2):485–521, June 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/2/
                                        485/4111183.

Dewar:2024:AU

[Dew24] Neil Dewar. On absolute units. British Journal for the Philosophy of Science, 75(1):1–30, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Finetti:1981:RDB

[dF81] Bruno de Finetti. The role of ‘Dutch Books’ and of ‘proper scoring rules’. British Journal for the Philosophy of Science, 32(1):55–56, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/1/55.full.pdf+html;
http://www.jstor.org/
stable/687386.

DaCosta:1989:PTL

[DF89] Newton C. A. Da Costa and Steven French. Pragmatic truth and the logic of induction. British Journal for the Philosophy of Science, 40(3):333– 356, September 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/333.full.pdf+html;
http://www.jstor.org/
stable/687781.

Dawid:2022:ESE

[DF22] Richard Dawid and Simon Friederich. Epistemic separability and Everettian branches: a critique of Sebens and Carroll. British Journal for the Philosophy of Science, 73(3):711–721, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic). See [SC18].

Daston:2010:OHT

[DG10] Lorraine J. Daston and Peter Louis Galison. Objectivity.

Zone Books, New York, NY, USA, 2010. ISBN 1-890951-79-X (paperback). 501 pp. LCCN BD220.

Dowe:2007:BBW

[DGO07] David L. Dowe, Steve Gardner, and Graham Oppy. Bayes not bust! why simplicity is no problem for Bayesians. British Journal for the Philosophy of Science, 58(4):709–754, December 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/4/709.full.pdf+html;
http://www.jstor.org/
stable/30115171.

Dunn:1986:DCA

[DH86] J. Michael Dunn and Geoffrey Hellman. Dualling: a critique of an argument of Popper and Miller. British Journal for the Philosophy of Science, 37(2):220–223, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686981.

Dawid:2015:NAA

[DHS15] Richard Dawid, Stephan Hartmann, and Jan Sprenger. The no alternatives argument. British Journal for the Philosophy of Science, 66(1):213–234, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/1/213.

Diamond:1963:EA

[Dia63] R. J. Diamond. Each and all. British Journal for the Philosophy of Science, XIII(52): 278–286, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/278.full.pdf+html;
http://www.jstor.org/
stable/685325.

Diamond:1964:RGB

[Dia64] R. J. Diamond. Reply to G.

A. Barnard. British Journal for the Philosophy of Science, XV(58):141b–142, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/58/141-b.full.pdf+html;
http://www.jstor.org/
stable/685255.

Dickson:1995:TRN

[Dic95] W. Michael Dickson. Is there really no projection postulate in the modal interpretation? British Journal for the Philosophy of Science, 46(2):197–218, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/197.full.pdf+
html;   http://www.jstor.
org/stable/687812.

Dickson:1998:R

[Dic98]  Michael Dickson.  Review.

British Journal for the Philosophy of Science, 49(2):329–331, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/329.full.pdf+
html;   http://www.jstor.
org/stable/688119.

Dickson:2001:BRB

[Dic01a] Michael Dickson. Book review: From Physics to Philosophy, Jeremy Butterfield, Constantine Pagonis. British Journal for the Philosophy of Science, 52(2):397–399, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/397.full.pdf+html;
http://www.jstor.org/
stable/3541905.

Dickson:2001:PGT

[Dic01b] Michael Dickson. Book review: Partha Ghose, Testing Quantum Mechanics on New Ground. British Journal for the Philosophy of Science, 52(1):207–209, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/207.full.pdf+
html;   http://www.jstor.
org/stable/3541959.

Dicken:2007:CEM

[Dic07] Paul Dicken. Constructive empiricism and the metaphysics of modality. British Journal for the Philosophy of Science, 58(3):

605–612, September 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/3/605.full.pdf+html;
http://www.jstor.org/
stable/30115190.

Dietl:1967:PAA

[Die67] Paul Dietl. Paresis and the alleged asymmetry between explanation and prediction. British Journal for the Philosophy of Science, 17(4):313–318, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/313.full.pdf+
html;   http://www.jstor.
org/stable/686772.

Dieterle:1999:BRB

[Die99] J. M. Dieterle. Book review:

Platonism and Anti-Platonism in Mathematics. M. Balaguer. British Journal for the Philosophy of Science, 50(4):775– 780, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/775.full.pdf+html;
http://www.jstor.org/
stable/40072268.

DiFrisco:2019:KBI

[DiF19] James DiFrisco. Kinds of biological individuals: Sortals, projectibility, and selection. British Journal for the Philosophy of Science, 70(3):845–875, September 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/3/845/4815699.

DiFrisco:2023:TTH

[DiF23] James DiFrisco. Toward a theory of homology: Development and the De-coupling of morphological and molecular evolution. British Journal for the Philosophy of Science, 74(3):771–810, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Dilman:1976:R

[Dil76] İlham Dilman. Reviews. British Journal for the Philosophy of Science, 27(2):191–197, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/27/2/191.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686171.

Dillinger:1984:R

[Dil84] Mike Dillinger. Reviews. British Journal for the Philosophy of Science, 35(3):301–303, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/35/3/301.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687481.

Dingle:1950:R

[Din50a] Herbert Dingle. Reviews. British Journal for the Philosophy of Science, I(3):245–248, November 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/3/245.full.pdf+
html;   http://www.jstor.
org/stable/685370.

Dingle:1950:TM

[Din50b] Herbert Dingle. A theory of measurement. British Journal for the Philosophy of Science, I(1):5– 26, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/1/5.full.pdf+
html;   http://www.jstor.
org/stable/685667.

Dingle:1951:SO

[Din51] Herbert Dingle. The scientific outlook in 1851 and in 1951. British Journal for the Philosophy of Science, II(6):85–104, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/85.full.pdf+
html;   http://www.jstor.
org/stable/685504.

Dingle:1953:R

[Din53a] Herbert Dingle. Reviews. British Journal for the Philosophy of Science, IV(16):350–352, February 1953. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
IV/16/350.full.pdf+html.

Dingle:1953:RSR

[Din53b] Herbert Dingle. Reviews: Science and religion. British Journal for the Philosophy of Science, IV(15):235–246, November 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/15/235.full.pdf+
html;   http://www.jstor.
org/stable/685901.

Dingle:1954:RU

[Din54] Herbert Dingle. Review: [untitled]. British Journal for the Philosophy of Science, IV(16): 350–352, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686089.

Dingle:1956:DBB

[Din56] Herbert Dingle. On Dr Burniston Brown’s note: a correction. British Journal for the Philosophy of Science, VI (24):335, February 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/24/335.full.pdf+html;
http://www.jstor.org/
stable/685414.

Dingle:1959:FLF

[Din59] Herbert Dingle. The falsifiability of the Lorentz–Fitzgerald contraction hypothesis. British Journal for the Philosophy of Science, X(39):228–229, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/39/228.full.pdf+
html;   http://www.jstor.
org/stable/686056.

Dingle:1960:DEFa

[Din60a] Herbert Dingle. The Doppler Effect and the foundations of physics (I). British Journal for the Philosophy of Science, XI (41):11–31, May 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/41/11.full.pdf+html;
http://www.jstor.org/
stable/685816.

Dingle:1960:DEFb

[Din60b] Herbert Dingle. The Doppler Effect and the foundations of physics (II). British Journal for the Philosophy of Science, XI (42):113–129, August 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/42/113.full.pdf+html;
http://www.jstor.org/
stable/685584.

Dingle:1960:RPG

[Din60c] Herbert Dingle. Reply to Professor Grünbaum. British Journal for the Philosophy of Science, XI(42):145a–145, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XI/42/145-a.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685587.

Dingle:1961:RDV

[Din61a] Herbert Dingle. A reply to Dr Van heerden. British Journal for the Philosophy of Science, XII(45):71, May 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XII/45/71.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685543.

Dingle:1961:RPG

[Din61b] Herbert Dingle. A reply to Professor Grünbaum’s rejoinder. British Journal for the Philosophy of Science, XII(46):156–157, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/46/156.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685497.

Dingle:1964:RER

[Din64a] Herbert Dingle. Reason and experiment in relation to the Spe-

cial  Relativity  the-

ory. British Journal for the Philosophy of Science, XV(57):41– 61, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/57/41.full.pdf+
html;   http://www.jstor.
org/stable/685626.

Dingle:1964:R

[Din64b] Herbert Dingle. Reviews. British Journal for the Philosophy of Science, XV(59):270–271, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/270.full.pdf+
html;   http://www.jstor.
org/stable/686532.

Dingle:1965:NMK

[Din65a] Herbert Dingle. Note on Mr Keswani’s articles, Origin and Concept of Relativity. British Journal for the Philosophy of Science, XVI(63):242–246, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/63/242.full.
pdf+html;   http://links.
jstor.org/sici?sici=0007-
0882%28196511%2916%3A63%
3C242%3ANOMKAO%3E2.0.CO%
3B2-0;   http://www.
jstor.org/stable/686307. See

[Kes65a, Kes65b].

Dingle:1965:RDS

[Din65b] Herbert Dingle. Reply to Dr Schleichert. British Journal for the Philosophy of Science, XV (60):331b–332, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XV/60/331-b.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/686542.

Dingle:1967:PFT

[Din67] Herbert Dingle. Particle and field theories of gravitation. British Journal for the Philosophy of Science, 18(1):57–64, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/1/57.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686890.

Dingle:1970:CSM

[Din70] Herbert Dingle. Causality and statistics in modern physics. British Journal for the Philosophy of Science, 21(3):223–246, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/223.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686194.

Disalle:1994:DIS

[Dis94] Robert Disalle. On dynamics, indiscernibility, and spacetime ontology. British Journal for the Philosophy of Science, 45 (1):265–287, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/265.full.pdf+html;
http://www.jstor.org/
stable/687972.

DiSalle:1999:BRB

[DiS99] R. DiSalle. Book review: Philosophical Concepts in Physics: the Historical Relation Between Philosophy and Scientific Theories. J. T. Cushing. British Journal for the Philosophy of Science, 50(4):747–759, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/4/747.full.pdf+
html;   http://www.jstor.
org/stable/40072265.

Dobbs:1951:RBTa

[Dob51a] H. A. C. Dobbs. The relation between the time of psychology and the time of physics Part I. British Journal for the Philosophy of Science, II(6):122–141, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/122.full.pdf+
html;   http://www.jstor.
org/stable/685506.

Dobbs:1951:RBTb

[Dob51b] H. A. C. Dobbs. The relation between the time of psychology and the time of physics Part II. British Journal for the Philosophy of Science, II(7): 177–192, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/7/177.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685291.

Dobbs:1953:TPP

[Dob53] H. A. C. Dobbs. The time of psychology and of physics. British Journal for the Philosophy of Science, IV(14):161–164, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IV/14/161.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685992.

Dobbs:1956:TPP

[Dob56] H. A. C. Dobbs. The time of physics and psychology. British Journal for the Philosophy of Science, VII(26):156–160, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/26/156.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685699.

Dobbs:1957:DSW

[Dob57] H. A. C. Dobbs. Diathesis, the self-winding watch, and photosynthesis. British Journal for the Philosophy of Science, VIII (30):140–150, August 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/30/140.full.pdf+html;
 http://www.jstor.org/
stable/685792.

Dobbs:1958:MT

[Dob58a] H. A. C. Dobbs. Multidimensional time. British Journal for the Philosophy of Science, IX(35):225–227, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/225.full.pdf+html;
http://www.jstor.org/
stable/685653.

Dobbs:1958:RPR

[Dob58b] H. A. C. Dobbs. Reply to Professor R. O. Kapp. British Journal for the Philosophy of Science, VIII(32):306–309, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/306.full.
pdf+html;   http://www.
jstor.org/stable/685857.

Dobbs:1961:CTR

[Dob61]  H. A. C. Dobbs.  Com-

ments on ‘Theory of Resonance’: Comments on Dr Ninian Marshall’s ‘Theory of Resonance’. British Journal for the Philosophy of Science, XII(45):65–68, May 1961. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/45/65.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685540.

Dobbs:1969:PP

[Dob69] H. A. C. Dobbs. The ‘present’ in physics. British Journal for the Philosophy of Science, 19(4): 317–324, February 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        19/4/317.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686291.

Dobbs:1970:RPG

[Dob70] H. A. C. Dobbs. Reply to Professor Grünbaum. British Journal for the Philosophy of Science, 21 (3):275–278, August 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/3/275.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686198.

Doboszewski:2021:EHD

[Dob21] Juliusz Doboszewski. Epistemic holes and determinism in classical general relativity. British Journal for the Philosophy of Science, 71(3):1093–1111, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Dodd:2013:RWA

[Dod13] Dylan Dodd. Roger White’s argument against imprecise credences. British Journal for the Philosophy of Science, 64 (1):69–77, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/69.full.pdf+html.

Dorling:1968:LCC

[Dor68] Jon Dorling. Length contraction and clock synchronisation: the empirical equivalence of the Einsteinian and Lorentzian theories. British Journal for the Philosophy of Science, 19(1):67– 69, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/1/67.full.pdf+
html;   http://www.jstor.
org/stable/686850.

Dorling:1971:EIP

[Dor71] Jon Dorling. Einstein’s introduction of photons: Argument by analogy or deduction from the phenomena? British Journal for the Philosophy of Science, 22(1):1–8, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/1.full.pdf+html;
http://www.jstor.org/
stable/686638.

Dorling:1972:BRS

[Dor72] Jon Dorling. Bayesianism and the rationality of scientific inference. British Journal for the Philosophy of Science, 23 (2):181–190, May 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/2/181.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686442.

Dorling:1975:R

[Dor75a] Jon Dorling. Reviews. British Journal for the Philosophy of Science, 26(4):355–358, December 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/4/355.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686682.

Dorling:1975:SSI

[Dor75b] Jon Dorling. The structure of scientific inference. British Journal for the Philosophy of Science, 26(1):61–71, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/1/61.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686335.

Dorling:1976:CL

[Dor76a] Jon Dorling. Comments on Landé. British Journal for the Philosophy of Science, 27(2):160, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/2/160-a.full.
pdf+html;   http://www.
jstor.org/stable/686165.

Dorling:1976:Ra

[Dor76b] Jon Dorling. Reviews. British Journal for the Philosophy of Science, 27(3):295–297, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/295.full.pdf+
html;   http://www.jstor.
org/stable/686129.

Dorling:1976:Rb

[Dor76c] Jon Dorling. Reviews. British Journal for the Philosophy of Science, 27(3):297–299, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/297.full.pdf+
html;   http://www.jstor.
org/stable/686130.

Dorling:1976:Rc

[Dor76d] Jon Dorling. Reviews. British Journal for the Philosophy of Science, 27(3):299–302, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/299.full.pdf+
html;   http://www.jstor.
org/stable/686131.

Dorling:1977:EMF

[Dor77] Jon Dorling. The eliminability of masses and forces in Newtonian particle mechanics: Suppes reconsidered. British Journal for the Philosophy of Science, 28(1):55–57, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/1/55.full.pdf+html;
http://www.jstor.org/
stable/686447.

Dorling:1978:DEN

[Dor78]  Jon Dorling.  Did Einstein

need General Relativity to solve the problem of absolute space? Or had the problem already been solved by Special Relativity? British Journal for the Philosophy of Science, 29(4): 311–323, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/311.full.pdf+html;
http://www.jstor.org/
stable/687095.

Dorato:1998:R

[Dor98] Mauro Dorato. Review. British Journal for the Philosophy of Science, 49(2):338–347, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/338.full.pdf+
html;   http://www.jstor.
org/stable/688121.

Dorato:2005:SUI

[Dor05] Mauro Dorato. The software of the universe: an introduction to the history and philosophy of laws of nature. Ashgate new critical thinking in philosophy. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-7546-3994-0 (hardcover). xv + 158 pp. LCCN BD581 .D65

2005. URL http://www.loc.
gov/catdir/toc/ecip058/
2005004113.html.

Dorst:2019:TBP

[Dor19] Chris Dorst. Towards a best predictive system account of laws of nature. British Journal for the Philosophy of Science, 70(3): 877–900, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
877/4816343.

Douven:2002:NSP

[Dou02] Igor Douven. A new solution to the paradoxes of rational acceptability. British Journal for the Philosophy of Science, 53(3):391–410, September 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/391.full.pdf+html;
http://www.jstor.org/
stable/3541888.

Douglas:2013:BRP

[Dou13] Heather Douglas. Book review: Philip Kitcher, Science in a Democratic Society. British Journal for the Philosophy of Science, 64(4):901–905, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
64/4/901.full.pdf+html.

Douven:2023:ESI

[Dou23] Igor Douven. Explaining the success of induction. British Journal for the Philosophy of Science, 74 (2):381–404, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Dowe:1993:RPC

[Dow93] Phil Dowe. On the reduction of process causality to statistical relations. British Journal for the Philosophy of Science, 44(2):325–327, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/2/325.full.pdf+html;
http://www.jstor.org/
stable/687651.

Dowe:2000:RAC

[Dow00] P. Dowe. Review article. causality and explanation. British Journal for the Philosophy of Science, 51(1):165–174, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/1/165.full.pdf+

html;   http://www.jstor.
org/stable/3541755.

Dowe:2005:GDH

[Dow05]  Phil Dowe.  Galileo, Dar-

win, and Hawking: the interplay of science, reason, and religion. William B. Eerdmans Pub. Co., Grand Rapids, MI, USA, 2005. ISBN 0-8028-2696-2 (paperback). viii + 205 pp. LCCN BL240.3 .D69 2005.

Dowding:2008:BRD

[Dow08] Keith Dowding. Book review:

Don Ross, Economic Theory and Cognitive Science: Microexplanation. British Journal for the Philosophy of Science, 59(3): 569–573, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/569.full.pdf+html;
http://www.jstor.org/
stable/40072300.

deRegt:1996:PKT

[dR96a] Henk W. de Regt. Philosophy and the kinetic theory of gases. British Journal for the Philosophy of Science, 47(1):31–62, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/31.full.pdf+
html;   http://www.jstor.
org/stable/1215825.

deRegt:1996:R

[dR96b] Henk W. de Regt. Review.

British Journal for the Philosophy of Science, 47(4):654– 657, December 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/4/654.full.pdf+html;
http://www.jstor.org/
stable/687935.

Dretske:1985:R

[Dre85]  Fred I. Dretske.  Reviews.

British Journal for the Philosophy of Science, 36(1):79–81, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/79.full.pdf+
html;   http://www.jstor.
org/stable/687159.

Drewery:2000:BRB

[Dre00] A. Drewery. Book review: On the Contrary: Critical Essays, 1987–1997. P. M. Churchland, P. S. Churchland. British Journal for the Philosophy of Science, 51(3):507–511, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/507.full.pdf+html;
http://www.jstor.org/
stable/3541779.

Drewery:2001:DCP

[Dre01] Alice Drewery. Dispositions and ceteris paribus laws. British Journal for the Philosophy of Science, 52(4):723–733, December 2001. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/4/723.full.pdf+
html;   http://www.jstor.
org/stable/3541916.

Driscoll:2011:FAW

[Dri11] Catherine Driscoll. Fatal attraction? why Sperber’s attractors do not prevent cumulative cultural evolution. British Journal for the Philosophy of Science, 62(2):301–322, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/301.full.pdf+html.

Dawid:2011:EIE

[DTV11] Philip. Dawid, William L. Twining, and Mimi Vasilaki, editors. Evidence, inference and enquiry, volume 171 of Proceedings of the British Academy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-726484-0. LCCN BC173 .E95

2011. URL http://www.loc.
gov/catdir/enhancements/
fy1214/2011294480-b.html;
http://www.loc.gov/catdir/
enhancements/fy1214/
2011294480-d.html;
http://www.loc.gov/catdir/
enhancements/fy1214/
2011294480-t.html.

Dardashti:2017:CAS

[DTW17] Radin Dardashti, Karim P. Y.

Thébault, and Eric Winsberg. Confirmation via analogue simulation: What dumb holes could tell us about gravity. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/1/55/2995149/
                                        Confirmation-via-Analogue-
                                        Simulation-What-Dumb.

Dunham:1962:RU

[Dun62a] B. Dunham. Review: [untitled]. British Journal for the Philosophy of Science, XIII (50):184, August 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685978.

Dunham:1962:Ra

[Dun62b] B. Dunham. Reviews. British Journal for the Philosophy of Science, XIII(50):182–183, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/182.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685975.

Dunham:1962:Rb

[Dun62c] B. Dunham. Reviews. British Journal for the Philosophy of Science, XIII(50):184b–184, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/184-b.full.
                                        pdf+html.

Dunn:1967:R

[Dun67] John Dunn. Reviews. British Journal for the Philosophy of Science, 17(4):337–339, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/337.full.pdf+
html;   http://www.jstor.
org/stable/686780.

Dunmore:1987:R

[Dun87] Caroline Dunmore. Reviews.

British Journal for the Philosophy of Science, 38(4):573– 575, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/573.full.pdf+html;
http://www.jstor.org/
stable/687363.

Dunn:2011:FET

[Dun11] Jeffrey Dunn. Fried eggs, thermodynamics, and the special sciences. British Journal for the Philosophy of Science, 62 (1):71–98, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/71.full.pdf+html.

Dupre:1994:R

[Dup94] John Dupré. Review. British Journal for the Philosophy of Science, 45(4):1084–1087, December 1994. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/4/1084.full.pdf+
html;   http://www.jstor.
org/stable/687627.

Dupre:1996:PRR

[Dup96] John Dupré. Promiscuous realism: Reply to Wilson. British Journal for the Philosophy of Science, 47(3):441–444, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/3/441.full.pdf+
html;   http://www.jstor.
org/stable/687987.

Dupre:1997:R

[Dup97] John Dupré. Review. British Journal for the Philosophy of Science, 48(2):292–296, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/292.full.pdf+
html;   http://www.jstor.
org/stable/687753.

Dupre:2000:DDB

[Dup00] J. Dupré. Discussion. in defence of the Baldwin effect: a reply to Watkins. British Journal for the Philosophy of Science, 51(3):477–479, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/477.full.pdf+html;

http://www.jstor.org/
stable/3541774.

Dupre:2013:BRE

[Dup13] John Dupré. Book review: Elliott Sober, Did Darwin Write The Origin Backwards? Philosophical Essays on Darwin’s Theory. British Journal for the Philosophy of Science, 64(4): 889–893, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/889.full.pdf+html.

Dupre:2014:PLE

[Dup14]  John Dupré.  Processes of

life: essays in the philosophy of biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-870122-5 (paperback), 0-19-969198-3 (hardcover). viii + 350 pp. LCCN QH331 .D85 2014.

Duran:1989:GDE

[Dur89] Jane Duran. Glymour on deoccamization and the epistemology of geometry. British Journal for the Philosophy of Science, 40 (1):127–134, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/1/127.full.pdf+html;
http://www.jstor.org/
stable/687469.

Durbin:2002:TFD

[Dur02] Paul T. Durbin. Technical fouls:

Democratic dilemmas and technological change. British Journal for the Philosophy of Science, 53 (1):157–159, March 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/157.full.pdf+html;
http://www.jstor.org/
stable/3541648.

Douven:2006:GLP

[DW06]  Igor  Douven  and

Timothy Williamson. Generalizing the lottery paradox. British Journal for the Philosophy of Science, 57(4):755–779, December 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/4/755.full.pdf+
html;   http://www.jstor.
org/stable/4489093.

Douven:2017:IBE

[DW17] Igor Douven and Sylvia Wenmackers. Inference to the best explanation versus Bayes’s rule in a social setting. British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/2/535/3044188/
Inference-to-the-Best-
Explanation-versus-Bayes-s.

Eagle:2005:RU

[Eag05] Antony Eagle. Randomness is unpredictability. British Journal for the Philosophy of Science, 56(4):749–790, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/749.full.pdf+html;
http://www.jstor.org/
stable/3541866.

Easwaran:2014:WPU

[Eas14] Kenny Easwaran. Why physics uses second derivatives. British Journal for the Philosophy of Science, 65(4):845–862, December 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/65/4/845.

Eccles:1952:OSC

[Ecc52]  J. C. Eccles.  Obituary:

Sir Charles Sherrington, O.M., F.R.S., 1857–1952. British Journal for the Philosophy of Science, III(11):298–301, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/298.full.
pdf+html;   http://www.
jstor.org/stable/685288.

Eccles:1953:R

[Ecc53] J. C. Eccles. Reviews. British Journal for the Philosophy of Science, IV(16):345–347, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        IV/16/345.full.pdf+html.

Eccles:1954:RU

[Ecc54] J. C. Eccles. Review: [untitled]. British Journal for the Philosophy of Science, IV(16): 345–347, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686086.

Eccles:1955:RU

[Ecc55a] J. C. Eccles. Review: [untitled]. British Journal for the Philosophy of Science, V(20): 351–352, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685746.

Eccles:1955:R

[Ecc55b] J. C. Eccles. Reviews. British Journal for the Philosophy of Science, V(20):351b–352, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        V/20/351-b.full.pdf+html.

Echeverri:2016:IWV

[Ech16] Santiago Echeverri. Indexing the world? Visual tracking, modularity, and the perception cognition interface. unknown, 67(1): 215–245, March 2016. CODEN ???? ISSN ???? URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/1/215.

Eder:2023:EPC

[Ede23] Anna-Maria Asunta Eder. Evidential probabilities and credences. British Journal for the Philosophy of Science, 74(1):1– 23, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Edgington:1997:MCC

[Edg97] Dorothy Edgington. Mellor on chance and causation. British Journal for the Philosophy of Science, 48(3):411–433, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/3/411.full.pdf+
html;   http://www.jstor.
org/stable/688071.

Board:1950:F

[Edi50] The Editorial Board. Foreword.

British Journal for the Philosophy of Science, I(1):1–4, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
I/1/1.full.pdf+html.

Edwards:1979:HVP

[Edw79] Jim Edwards. Hidden variables and the propensity theory of probability. British Journal for the Philosophy of Science, 30(4):315–328, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/4/315.full.pdf+html;

http://www.jstor.org/
stable/687542.

Eells:1988:AII

[Eel88] Ellery Eells. On the alleged impossibility of inductive probability. British Journal for the Philosophy of Science, 39 (1):111–116, March 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/1/111.full.pdf+html;
http://www.jstor.org/
stable/687117.

Eells:2000:BRB

[Eel00] Ellery Eells. Book review: James M. Joyce — The Foundations of Causal Decision Theory. British Journal for the Philosophy of Science, 51(4):893–900, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/4/893.full.pdf+
html;   http://www.jstor.
org/stable/3541733.

Eggenberger:1980:R

[Egg80] Peter Eggenberger. Reviews. British Journal for the Philosophy of Science, 31(3):299– 301, September 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/3/299.full.pdf+html;
http://www.jstor.org/
stable/686924.

Egg:2016:EOG

[Egg16] Matthias Egg. Expanding our grasp: Causal knowledge and the problem of unconceived alternatives. unknown, 67(1):115– 141, March 2016. CODEN ????

ISSN ???? URL http://bjps.
oxfordjournals.org/content/
67/1/115.

Esfeld:2014:OBM

[EHLD14] Michael Esfeld, Mario Hubert, Dustin Lazarovici, and Detlef Dürr. The ontology of Bohmian mechanics. British Journal for the Philosophy of Science, 65(4): 773–796, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/773.

Ehring:1984:PCP

[Ehr84] Douglas Ehring. Probabilistic causality and preemption. British Journal for the Philosophy of Science, 35(1):55–57, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/55.full.pdf+
html;   http://www.jstor.
org/stable/687557.

Ehring:1987:PCA

[Ehr87] Douglas Ehring. Papineau on causal asymmetry. British Journal for the Philosophy of Science, 38(1):81–87, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/1/81.full.pdf+html;
                                        http://www.
                                        jstor.org/stable/687082. See

response [Pap88].

Einstein:1950:MLYa

[Ein50a] Albert Einstein. Out of my later years. Philosophical Library, New York, NY, USA, 1950. vii + 282 pp. LCCN QC16.E5 A3.

Einstein:1950:MLYb

[Ein50b] Albert Einstein. Out of my later years. Thames and Hudson, London, UK, 1950. viii + 282 pp. LCCN QC16.E5 A3.

Evans:1972:FWM

[EL72] D. A. Evans and P. T. Landsberg. Free will in a mechanistic universe? an extension. British Journal for the Philosophy of Science, 23(4): 336–343, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/4/336.full.pdf+html.

Elby:1993:WLR

[Elb93] Andrew Elby. Why local realistic theories violate, nontrivially, the quantum mechanical EPR perfect correlations. British Journal for the Philosophy of Science, 44(2):213–230, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/

44/2/213.full.pdf+html;
http://www.jstor.org/
stable/687645.

Elby:1994:CCI

[Elb94] Andrew Elby. Contentious contents: for inductive probability. British Journal for the Philosophy of Science, 45(1):193–200, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/193.full.pdf+
html;   http://www.jstor.
org/stable/687967.

Elkana:1971:NEC

[Elk71] Yehuda Elkana. Newtonianism in the Eighteenth Century. British Journal for the Philosophy of Science, 22(3):297–306, August 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/3/297.full.pdf+
html;   http://www.jstor.
org/stable/686750.

Ellis:1957:CPN

[Ell57] Brian Ellis. A comparison of process and non-process theories in the physical sciences. British Journal for the Philosophy of Science, VIII(29):45–56, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/29/45.full.pdf+html.

Ellis:1963:UDF

[Ell63] Brian Ellis. Universal and differential forces. British Journal for the Philosophy of Science, XIV (55):177–194, November 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/55/177.full.pdf+html.

Ellis:1966:RU

[Ell66] G. F. R. Ellis. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):83–85, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686410.

Ellis:1967:R

[Ell67] G. F. R. Ellis. Reviews. British Journal for the Philosophy of Science, 18(3):257–259, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/3/257.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686600.

Ellis:1973:LSP

[Ell73] Brian Ellis. The logic of subjective probability. British Journal for the Philosophy of Science, 24(2):125–152, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        24/2/125.full.pdf+html;

http://www.jstor.org/
stable/686605.

Ellis:1988:SPI

[Ell88] Brian Ellis. Solving the problem of induction using a values-based epistemology. British Journal for the Philosophy of Science, 39 (2):141–160, June 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/2/141.full.pdf+html;
http://www.jstor.org/
stable/687263.

Ellis:1996:R

[Ell96] Brian Ellis. Review. British Journal for the Philosophy of Science, 47(4):636–639, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/636.full.pdf+
html;   http://www.jstor.
org/stable/687932.

Elliott:2022:RP

[Ell22] Steve Elliott. Research problems. British Journal for the Philosophy of Science, 72(4):1013– 1037, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Esfeld:2017:PMP

[ELLH17] Michael Esfeld, Dustin Lazarovici, Vincent Lam, and Mario Hubert. The physics and metaphysics of primitive stuff. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL https://academic.oup.
com/bjps/article/68/1/133/
2995312/The-Physics-and-
Metaphysics-of-Primitive-
Stuff.

Elton:2003:DVG

[Elt03] Matthew Elton. Dawkins vs. Gould: Survival of the fittest. British Journal for the Philosophy of Science, 54(2):365–369, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/365.full.pdf+
html;   http://www.jstor.
org/stable/3541976.

Elzinga:1983:R

[Elz83] Aant Elzinga. Reviews. British Journal for the Philosophy of Science, 34(3):295–303, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/3/295.full.pdf+
html;   http://www.jstor.
org/stable/687327.

Emery:2015:CPE

[Eme15] Nina Emery. Chance, possibility, and explanation. British Journal for the Philosophy of Science, 66(1):95–120, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/1/95.

Emmet:1968:R

[Emm68] Dorothy Emmet. Reviews.

British Journal for the Philosophy of Science, 19(1):83, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/1/83.full.pdf+
html;   http://www.jstor.
org/stable/686854.

Earman:1987:WPS

[EN87] John Earman and John Norton.

What price spacetime substantivalism? the hole story. British Journal for the Philosophy of Science, 38(4):515–525, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/4/515.full.pdf+
html;   http://www.jstor.
org/stable/687356.

Earman:1998:CLC

[EN98] John Earman and John D. Norton. Comments on Laraudogoitia’s ‘classical particle dynamics, indeterminism and a supertask’. British Journal for the Philosophy of Science, 49 (1):123–133, March 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/1/123.full.pdf+html;
http://www.
jstor.org/stable/688146. See

[Lar97].

Enc:1994:R

[Enç94] Berent Enç. Review. British Journal for the Philosophy of Science, 45(4):1070–1077, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/4/1070.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687625.

Enfield:2008:BRP

[Enf08] Patrick Enfield. Book review: P. Kyle Stanford, Exceeding Our Grasp: Science, History, and the Problem of Unconceived Alternatives. British Journal for the Philosophy of Science, 59(4): 881–895, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/4/881.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072316.

Evans:2013:NSE

[EPW13] P. W. Evans, Huw Price, and K. B. Wharton. New slant on the EPR– Bell experiment. British Journal for the Philosophy of Science, 64(2):297–324, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/2/297.full.pdf+html.

Earman:1996:WET

[ER96] John Earman and Miklós Rédei.

Why ergodic theory does not explain the success of equilibrium statistical mechanics. British Journal for the Philosophy of Science, 47(1):63–78, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/63.full.pdf+
html;   http://www.jstor.
org/stable/1215826.

Ernest:1975:CSF

[Ern75] Paul Ernest. A critique of some formal theories of meaning. British Journal for the Philosophy of Science, 26(4): 319–330, December 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/4/319.full.pdf+html;
http://www.jstor.org/
stable/686678.

Ernest:1990:MME

[Ern90] Paul Ernest. The meaning of mathematical expressions: Does philosophy shed any light on psychology? British Journal for the Philosophy of Science, 41(4):443–460, December 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/443.full.pdf+html;
http://www.jstor.org/
stable/687859.

Ernst:2001:ESC

[Ern01] Zachary Ernst. Explaining the social contract. British Journal for the Philosophy of Science, 52(1):1–24, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541940.

Eronen:2012:BRJ

[Ero12] Markus I. Eronen. Book review: Jakob Hohwy and Jesper Kallestrup (eds), Being Reduced: New Essays on Reduction, Explanation, and Causation. British Journal for the Philosophy of Science, 63(1): 227–231, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/1/227.full.pdf+html.

Erwin:1982:BTR

[Erw82] Edward Erwin. Behaviour therapy: a reply to an incautious discussion. British Journal for the Philosophy of Science, 33 (1):53–60, March 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/1/53.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687240.

Erwin:1984:SP

[Erw84] Edward Erwin. The standing of psychoanalysis. British Journal for the Philosophy of Science, 35(2):115–128, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/2/115.full.pdf+html;
http://www.jstor.org/
stable/686962.

Ebersole:1959:OEO

[ES59] Frank B. Ebersole and Marvin M. Shrewsbury. Origin explanations and the origin of life. British Journal for the Philosophy of Science, X(38):103–119, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/103.full.pdf+
html;   http://www.jstor.
org/stable/686002.

Erwin:1989:CD

[ES89]  Edward  Erwin

and Harvey Siegel. Is confirmation differential? British Journal for the Philosophy of Science, 40 (1):105–119, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/1/105.full.pdf+html;
http://www.jstor.org/
stable/687467.

Eva:2019:CEP

[ES19]  Benjamin  Eva

and Reuben Stern. Causal explanatory power. British Journal for the Philosophy of Science, 70 (4):1029–1050, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1029/4831125.

Eliasmith:1997:WPE

[ET97] Chris Eliasmith and Paul Thagard. Waves, particles, and explanatory coherence. British Journal for the Philosophy of Science, 48(1):1–19, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/1/1.full.pdf+html;
http://www.jstor.org/
stable/687714.

Evans:1968:R

[Eva68] J. D. G. Evans. Reviews. British Journal for the Philosophy of Science, 19(3):265–267, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/265.full.pdf+
html;   http://www.jstor.
org/stable/686250.

Evans:1971:R

[Eva71] Gareth Evans. Reviews. British Journal for the Philosophy of Science, 22(1):74–75, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/74.full.pdf+html;
http://www.jstor.org/
stable/686656.

Evans:1998:R

[Eva98] Dylan Evans. Review. British Journal for the Philosophy of Science, 49(2):348–350, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/348.full.pdf+
html;   http://www.jstor.
org/stable/688122.

Evans:2002:SHE

[Eva02] Dylan Evans. The search hypothesis of emotion. British Journal for the Philosophy of Science, 53(4):497–509, December 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/4/497.full.pdf+
html;   http://www.jstor.
org/stable/3541821.

Eva:2017:TTQ

[Eva17] Benjamin Eva. Topos theoretic quantum realism. British Journal for the Philosophy of Science, 68(4):1149–1181, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/1149/2669655.

Evans:2018:QCM

[Eva18] Peter W. Evans. Quantum causal models, faithfulness, and retrocausality. British Journal for the Philosophy of Science, 69(3):745–774, September 1, 2018. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/3/745/3066275.

Everitt:1987:R

[Eve87]  Nicholas Everitt.  Reviews.

British Journal for the Philosophy of Science, 38(2):277–279, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/2/277.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687056.

Ewens:2011:WGT

[Ewe11] Warren J. Ewens. What is the gene trying to do? British Journal for the Philosophy of Science, 62(1):155–176, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/1/155.full.pdf+html.

Ezriel:1956:EWP

[Ezr56] Henry Ezriel. Experimentation within the psycho-analytic session. British Journal for the Philosophy of Science, VII(25):29– 48, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/25/29.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685934.

Ezriel:1957:RMS

[Ezr57] Henry Ezriel. Reply to Mr Spilsbury. British Journal for the Philosophy of Science, VII(28): 342–347, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/342.full.pdf+html;
http://www.
jstor.org/stable/685117. See

[Spi57].

Frost-Arnold:2014:CPI

[FA14] Greg Frost-Arnold. Can the pessimistic induction be saved from semantic anti-realism about scientific theory? British Journal for the Philosophy of Science, 65(3):521–548, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/3/521.

Franklin:1989:CTL

[FAB^{+}89] A. Franklin, M. Anderson, D. Brock, S. Coleman, J. Downing, A. Gruvander, J. Lilly, J. Neal, D. Peterson, M. Price, R. Rice, L. Smith, S. Speirer, and D. Toering. Can a theory-laden observation test the theory? British Journal for the Philosophy of Science, 40(2):229–231, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/229.full.pdf+html;
http://www.jstor.org/
stable/687514.

Fagan:2013:PSC

[Fag13] Melinda Bonnie Fagan. Philosophy of stem cell biology: knowledge in flesh and blood. New directions in the philosophy of science. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2013. ISBN 0-230-36827-1, 0-230-20210-1 (series). xx + 274 pp. LCCN QH331 .F25 2013; QH588.S83 F35

                                        2013. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1310/2012540513-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1310/
                                        2012540513-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1310/
                                        2012540513-t.html.

Fairbairn:1956:CEC

[Fai56] W. Ronald D. Fairbairn. A critical evaluation of certain basic psycho-analytical conceptions. British Journal for the Philosophy of Science, VII (25):49–60, May 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/25/49.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685935.

Fairbairn:1957:FRC

[Fai57] W. Ronald D. Fairbairn. Fairbairn’s reply to the comments of Balint, Foulkes, and Sutherland. British Journal for the Philosophy of Science, VII(28): 333–338, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/333.full.pdf+html;
http://www.jstor.org/
stable/685115.

Fales:1979:RE

[Fal79] Evan Fales. Relative essentialism. British Journal for the Philosophy of Science, 30(4): 349–370, December 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/4/349.full.pdf+html;
http://www.jstor.org/
stable/687544.

Fallis:2000:RRA

[Fal00] D. Fallis. The reliability of randomized algorithms. British Journal for the Philosophy of Science, 51(2):255–271, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/2/255.full.pdf+
html;   http://www.jstor.
org/stable/3541804.

Farre:1967:RST

[Far67]  G. L. Farre.  Remarks on

Swanson’s theory of models. British Journal for the Philosophy of Science, 18(2):140–144, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/18/2/140.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686581.

Fara:2000:BRN

[Far00] Patricia Fara. Book review: Niccolò Guicciardini — Reading the Principia: the Debate on Newton’s Mathematical Methods for Natural Philosophy from 1687 to 1736. British Journal for the Philosophy of Science, 51(4):935–939, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/4/935.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541740.

Farina:2016:TAH

[Far16a] Mirko Farina. Three approaches to human cognitive development: Neo-nativism, neuroconstructivism, and dynamic enskillment. British Journal for the Philosophy of Science, 67 (2):617–641, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/2/617.

Farr:2016:BRM

[Far16b] Matt Farr. Book review: Mathias Frisch, Causal Reasoning in Physics. British Journal for the Philosophy of Science, 67 (4):1207–1213, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1207.

Farr:2020:CTR

[Far20] Matt Farr. Causation and time reversal. British Journal for the Philosophy of Science, 71 (1):177–204, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
177/4092770.

Faye:1995:R

[Fay95] Jan Faye. Review. British Journal for the Philosophy of Science, 46(2):275–279, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/2/275.full.pdf+html;
http://www.jstor.org/
stable/687819.

Frigg:2001:BRP

[FB01]  Roman  Frigg  and

Joseph Berkovitz. Book review: Peter Smith, Explaining Chaos. British Journal for the Philosophy of Science, 52(1):201–205, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/201.full.pdf+
html;   http://www.jstor.
org/stable/3541958.

Franke:2018:VII

[FC18] Michael Franke and José Pedro Correia. Vagueness and impre-

cise  imitation  in

signalling games. British Journal for the Philosophy of Science, 69(4):1037–1067, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/1037/3952633.

Feibleman:1954:TIL

[Fei54] James K. Feibleman. Theory of integrative levels. British Journal for the Philosophy of Science, V(17):59–66, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/17/59.full.pdf+html;
http://www.jstor.org/
stable/685952.

Feintzeig:2015:HVI

[Fei15] Benjamin Feintzeig. Hidden variables and incompatible observables in quantum mechanics. British Journal for the Philosophy of Science, 66(4): 905–927, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Feintzeig:2018:TUP

[Fei18] Benjamin Feintzeig. Toward an understanding of parochial observables. British Journal for the Philosophy of Science, 69(1): 161–191, March 1, 2018. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/161/2669774.

Feintzeig:2022:REC

[Fei22] Benjamin H. Feintzeig. Reductive explanation and the construction of quantum theories. British Journal for the Philosophy of Science, 73(2):457–486, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Feldman:2001:BRA

[Fel01] Richard Feldman. Book review:

Alvin Goldman, Knowledge in a Social World. British Journal for the Philosophy of Science, 52 (1):163–168, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/163.full.pdf+html;
http://www.jstor.org/
stable/3541951.

Fellowes:2023:VCP

[Fel23] Sam Fellowes. The value of categorical polythetic diagnoses in psychiatry. British Journal for the Philosophy of Science, 73(4): 941–963, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Ferre:1970:GVD

[Fer70] Frederick Ferré. Grünbaum vs.

Dobbs: the need for physical transiency. British Journal for the Philosophy of Science, 21 (3):278–280, August 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
21/3/278.full.pdf+html;
http://www.jstor.org/
stable/686199.

Festa:2006:BRG

[Fes06] Roberto Festa. Book review: George Kampis, Ladislav Kvasz, and Michael Stöltzner: Appraising Lakatos. Mathematics, Methodology and the Man, Dordrecht/Boston/London: Kluwer Academic Publishers, 2002 cloth £90/US$143/EUR130, hardback ISBN 1-4020-0226-2 (hardback). British Journal for the Philosophy of Science, 57(1):247–253, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/1/247.full.pdf+
html;   http://www.jstor.
org/stable/3541661.

Feyerabend:1956:R

[Fey56] P. K. Feyerabend. Reviews. British Journal for the Philosophy of Science, VII(27):260– 263, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/260.full.pdf+html;
http://www.jstor.org/
stable/685885.

Feyerabend:1958:R

[Fey58] P. K. Feyerabend. Reviews. British Journal for the Philosophy of Science, VIII(32):343– 347, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/343.full.pdf+html;
 http://www.jstor.org/
stable/685864.

Feyerabend:1959:R

[Fey59] P. K. Feyerabend. Reviews.

British Journal for the Philosophy of Science, IX(36):336– 337, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/336.full.pdf+html;
http://www.jstor.org/
stable/685198.

Feyerabend:1960:PBP

[Fey60]  P. K. Feyerabend.  Profes-

sor Bohm’s philosophy of nature. British Journal for the Philosophy of Science, X(40): 321–338, February 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/321.full.pdf+html;
http://www.jstor.org/
stable/685921.

Feyerabend:1961:R

[Fey61] P. K. Feyerabend. Reviews.

British Journal for the Philosophy of Science, XII(47):245– 250, November 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        XII/47/245.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685211.

Feyerabend:1963:R

[Fey63] P. K. Feyerabend. Reviews. British Journal for the Philosophy of Science, XIII(52):319– 323, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/52/319.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685333.

Feyerabend:1964:R

[Fey64] Paul K. Feyerabend. Reviews. British Journal for the Philosophy of Science, XV(59):244– 254, November 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XV/59/244.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686526.

Feyerabend:1966:SS

[Fey66] Paul K. Feyerabend. The structure of science. British Journal for the Philosophy of Science, 17(3):237–249, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/3/237.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686761.

Feyerabend:1967:R

[Fey67] Paul K. Feyerabend. Reviews.

British Journal for the Philosophy of Science, 17(4):326– 328, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/326.full.pdf+html;
http://www.jstor.org/
stable/686776.

Feyerabend:1968:NTP

[Fey68] Paul K. Feyerabend. A note on two ‘problems’ of induction. British Journal for the Philosophy of Science, 19(3): 251–253, November 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/3/251.full.pdf+html;
http://www.jstor.org/
stable/686244.

Feyerabend:1974:ZE

[Fey74] P. K. Feyerabend. Zahar on Einstein. British Journal for the Philosophy of Science, 25 (1):25–28, March 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/1/25.full.pdf+html;
http://www.jstor.org/
stable/686500.

Feyerabend:1975:IL

[Fey75] Paul Feyerabend. Imre Lakatos.

British Journal for the Philosophy of Science, 26(1):1–18, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/1/1.full.pdf+
html;   http://www.jstor.
org/stable/686328.

Feyerabend:1976:LLP

[Fey76] Paul Feyerabend. Logic, literacy, and Professor Gellner. British Journal for the Philosophy of Science, 27(4):381–391, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/4/381.full.pdf+
html;   http://www.jstor.
org/stable/686864.

Feyerabend:1977:CPR

[Fey77] Paul Feyerabend. Changing patterns of reconstruction. British Journal for the Philosophy of Science, 28(4):351–369, December 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/4/351.full.pdf+
html;   http://www.jstor.
org/stable/686879.

Feyerabend:1980:ZME

[Fey80] Paul Feyerabend. Zahar on Mach, Einstein and modern science. British Journal for the Philosophy of Science, 31(3):273– 282, September 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/3/273.full.pdf+html;
http://www.jstor.org/
stable/686921.

Feyerabend:1981:MCE

[Fey81] Paul K. Feyerabend. More clothes from the emperor’s bargain basement. British Journal for the Philosophy of Science, 32(1):57–71, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/1/57.full.pdf+html;
http://www.jstor.org/
stable/687387.

Feyerabend:1987:PI

[Fey87] Paul Feyerabend. Putnam on incommensurability. British Journal for the Philosophy of Science, 38(1):75b–81, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/75-b.full.pdf+html;
http://www.jstor.org/
stable/687081.

Fenton-Glynn:2017:PPE

[FG17] Luke Fenton-Glynn. A proposed probabilistic extension of the Halpern and Pearl definition of ‘actual cause’. British Journal for the Philosophy of Science, 68(4):1061–1124, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/1061/2669651.

Fenton-Glynn:2015:RQE

[FGK15] Luke Fenton-Glynn and Thomas Kroedel. Relativity, quantum entanglement, counterfactuals, and causation. British Journal for the Philosophy of Science, 66 (1):45–67, March 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/1/45.

Franklin-Hall:2016:HLE

[FH16]  L. R. Franklin-Hall.  High-

level explanation and the interventionist’s ‘variables problem’. British Journal for the Philosophy of Science, 67(2):553–577, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/2/553.

Field:1986:JBH

[Fie86]  J. V. Field.  Book review:

Jardine, N. [1984]: The birth of history and philosophy of science: Kepler’s ‘A Defence of Tycho against Ursus’ with essays on its provenance and significance. Cambridge University Press. £32.50. Pp. ix + 301. British Journal for the Philosophy of Science, 37(2): 255–257, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686991.

Fine:1964:NGS

[Fin64] Arthur I. Fine. Note on Goodstein’s ‘the significance of incompleteness theorems’. British Journal for the Philosophy of Science, XV(58):140–141, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/140.full.pdf+
html;   http://www.
jstor.org/stable/685253. See

[Goo63, Goo64].

Fine:1968:Ra

[Fin68a] Arthur I. Fine. Reviews. British Journal for the Philosophy of Science, 18(4):335–339, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
18/4/335.full.pdf+html.

Fine:1968:Rb

[Fin68b] Arthur I. Fine. Reviews. British Journal for the Philosophy of Science, 18(4):342–344, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/4/342.full.pdf+
html;   http://www.jstor.
org/stable/686365.

Fine:1969:R

[Fin69] Arthur Fine. Reviews. British Journal for the Philosophy of Science, 20(4):357–359, December 1969. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/357.full.pdf+
html;   http://www.jstor.
org/stable/686264.

Fine:1971:Ra

[Fin71a] K. Fine. Reviews. British Journal for the Philosophy of Science, 22(2):205–206, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/2/205.full.pdf+html;
http://www.jstor.org/
stable/686492.

Fine:1971:Rb

[Fin71b] K. Fine. Reviews. British Journal for the Philosophy of Science, 22(2):206–207, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/2/206.full.pdf+html;
http://www.jstor.org/
stable/686493.

Fine:1973:PIQ

[Fin73] Arthur Fine. Probability and the interpretation of quantum mechanics. British Journal for the Philosophy of Science, 24 (1):1–37, March 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/1/1.full.pdf+html;
http://www.jstor.org/
stable/686667.

Finocchiaro:1975:PAG

[Fin75]  Maurice  A.  Finoc-

chiaro. Philosophising about Galileo. British Journal for the Philosophy of Science, 26(3): 255–264, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/255.full.pdf+html;
http://www.jstor.org/
stable/686283.

Finch:1981:RDS

[Fin81] Peter D. Finch. On the role of description in statistical enquiry. British Journal for the Philosophy of Science, 32(2):127–144, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/2/127.full.pdf+
html;   http://www.jstor.
org/stable/687194.

Finocchiaro:1996:R

[Fin96] Maurice A. Finocchiaro. Review.

British Journal for the Philosophy of Science, 47(2):325–327, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/325.full.pdf+
html;   http://www.jstor.
org/stable/687953.

Fine:2002:LA

[Fin02] Kit Fine. The limits of abstraction. Clarendon Press, New York, NY, USA, 2002. ISBN 0-19-924618-1 (hardcover). x + 203 pp. LCCN QA8.4 .F56 2002.

Fisk:1967:DPE

[Fis67] Milton Fisk. A defence of the principle of event causality. British Journal for the Philosophy of Science, 18(2):89–108, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/89.full.pdf+
html;   http://www.jstor.
org/stable/686578.

Fisher:1979:CVO

[Fis79] Gordon Fisher. Cauchy’s variables and orders of the infinitely small. British Journal for the Philosophy of Science, 30(3):261–265, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/3/261.full.pdf+html;
http://www.jstor.org/
stable/687140.

Fisher:2006:HOF

[Fis06] Justin C. Fisher. On higherorder and free-floating chances. British Journal for the Philosophy of Science, 57(4):691– 707, December 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/4/691.full.pdf+html;
http://www.jstor.org/
stable/4489091.

Fischer:2024:TCA

[Fis24] Enno Fischer. Three concepts of actual causation. British Journal for the Philosophy of Science, 75 (1):77–98, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Fitzgerald:2010:LI

[Fit10] Gareth Fitzgerald. Linguistic intuitions. British Journal for the Philosophy of Science, 61 (1):123–160, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/123.full.pdf+html.

Flew:1951:TR

[Fle51] Antony Flew. Theories of Relativity. British Journal for the Philosophy of Science, II(5):78– 79, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/78.full.pdf+
html;   http://www.jstor.
org/stable/685616.

Flew:1952:R

[Fle52] Antony Flew. Reviews. British Journal for the Philosophy of Science, III(11):290–291, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/290.full.
pdf+html;   http://www.
jstor.org/stable/685284.

Flew:1979:RFS

[Fle79] Antony Flew. A religious form of scientific life. British Journal for the Philosophy of Science, 30(2):183–186, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/2/183.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686951.

Fletcher:2016:STP

[Fle16] Samuel C. Fletcher. Similarity, topology, and physical significance in relativity theory. British Journal for the Philosophy of Science, 67(2):365–389, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/2/365.

French:2011:E

[FM11] Steven French and Michela Massimi. Editorial. British Journal for the Philosophy of Science, 62(4):691–692, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/4/691.full.pdf+html.

Fazekas:2021:AAS

[FN21] Peter Fazekas and Bence Nanay. Attention is amplification, not selection. British Journal for the Philosophy of Science, 72(1): 299–324, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Fodor:2001:BRB

[Fod01] Jerry Fodor. Book review: Evolution and the Human Mind: Modularity, Language and Meta-Cognition, Peter Carruthers and Andrew Chamberlin. British Journal for the Philosophy of Science, 52(3):623–628, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/3/623.full.pdf+
html;   http://www.jstor.
org/stable/3541933.

Folina:1999:DPP

[Fol99] J. Folina. Discussion. pictures, proofs, and ’mathematical practice’: reply to James Robert Brown. British Journal for the Philosophy of Science, 50(3): 425–429, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/425.full.pdf+html;
http://www.jstor.org/
stable/40072240.

Folina:2000:RAO

[Fol00] J. Folina. Review article. ontology, logic, and mathematics. British Journal for the Philosophy of Science, 51(2):319–332, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/51/2/319.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541808.

Forrest:1962:R

[For62] D. W. Forrest. Reviews. British Journal for the Philosophy of Science, XIII(50):185–186, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/185.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685980.

Forster:1995:BBS

[For95a] Malcolm R. Forster. Bayes and bust: Simplicity as a problem for a probabilist’s approach to confirmation. British Journal for the Philosophy of Science, 46(3):399–424, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/3/399.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687666.

Forster:1995:GDR

[For95b] Malcolm R. Forster. The Golfer’s Dilemma: a reply to Kukla on curve-fitting. British Journal for the Philosophy of Science, 46(3):348–360, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/3/348.full.pdf+html;

http://www.jstor.org/
stable/687662.

Forrest:1997:R

[For97] Peter Forrest. Review. British Journal for the Philosophy of Science, 48(1):149–150, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/149.full.pdf+
html;   http://www.jstor.
org/stable/687729.

Forster:1999:MSS

[For99] Malcolm R. Forster. Model selection in science: the problem of language variance. British Journal for the Philosophy of Science, 50(1):83–102, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/83.full.pdf+html;
http://www.jstor.org/
stable/40072209.

Foster:1969:FSG

[Fos69] Lawrence Foster. Feyerabend’s solution to the Goodman paradox. British Journal for the Philosophy of Science, 20(3): 259–260, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/3/259.full.pdf+html;
http://www.jstor.org/
stable/686420.

Foster:2004:DLL

[Fos04] John Foster. The divine lawmaker: lectures on induction, laws of nature, and the existence of God. Clarendon Press, New York, NY, USA, 2004. ISBN 0-19-925059-6. viii + 170 pp. LCCN BT103 .F67 2004; BD555; BD555 .F67 2004. URL http://

                                        www.oxfordscholarship.com/
                                        oso/public/content/
                                        philosophy/0199250596/toc.

Foulkes:1957:CFP

[Fou57] S. H. Foulkes. Comments on Fairbairn’s paper. British Journal for the Philosophy of Science, VII(28):324–329, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/28/324.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685113.

Fowler:1983:IEE

[Fow83] D. H. Fowler. Investigating Euclid’s elements. British Journal for the Philosophy of Science, 34(1):57–70, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        34/1/57.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686934.

Fox:1975:R

[Fox75] John Fox. Reviews. British Journal for the Philosophy of Science, 26(1):85–87, March 1975.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/1/85.full.pdf+html;
http://www.jstor.org/
stable/686339.

Fozzy:1963:PMM

[Foz63] P. J. Fozzy. Professor Mackay on machines. British Journal for the Philosophy of Science, XIV(54):154–157, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/54/154.full.pdf+html;
http://www.jstor.org/
stable/685434.

Frankfurt:1966:FAB

[FP66] Harry G. Frankfurt and Brian Poole. Functional analyses in biology. British Journal for the Philosophy of Science, 17 (1):69–72, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686406.

French:1988:QPI

[FR88] Steven French and Michael Redhead. Quantum physics and the identity of indiscernibles. British Journal for the Philosophy of Science, 39(2):233–246, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/2/233.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687269.

Frank:1950:MISa

[Fra50a] Philipp Frank. Metaphysical interpretations of science: Part I. British Journal for the Philosophy of Science, I(1):60–74, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/I/1/60.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685670.

Frank:1950:MISb

[Fra50b] Philipp Frank. Metaphysical interpretations of science: Part II. British Journal for the Philosophy of Science, I(2):77–91, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/I/2/77.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685804.

Francis:1980:ITN

[Fra80] R. Francis. On the interpretation and transitivity of non-standard synchronisms. British Journal for the Philosophy of Science, 31(2):165–173, June 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/2/165.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687184.

Franklin:1981:WMG

[Fra81] A. D. Franklin. What makes a ‘good’ experiment? British Journal for the Philosophy of Science, 32(4):367–374, December 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/4/367.full.pdf+
html;   http://www.jstor.
org/stable/687307.

Franklin:1983:R

[Fra83] Allan Franklin. Reviews. British Journal for the Philosophy of Science, 34(2):190–192, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/190.full.pdf+
html;   http://www.jstor.
org/stable/687453.

Franklin:1984:PI

[Fra84a] Allan Franklin. Are paradigms incommensurable? British Journal for the Philosophy of Science, 35(1):57–60, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/1/57.full.pdf+html;
http://www.jstor.org/
stable/687558.

Franklin:1984:EE

[Fra84b] Allan Franklin. The epistemology of experiment. British Journal for the Philosophy of Science, 35(4):381–390, December 1984.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/381.full.pdf+html;
http://www.jstor.org/
stable/687342.

Franklin:1987:NDL

[Fra87] James Franklin. Non-deductive logic in mathematics. British Journal for the Philosophy of Science, 38(1):1–18, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/1.full.pdf+html;
http://www.jstor.org/
stable/687075.

Franklin:1988:HNC

[Fra88a] Allan Franklin. How Nancy Cartwright tells the truth. British Journal for the Philosophy of Science, 39(4):527–529, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/527.full.pdf+html;
http://www.jstor.org/
stable/687437.

Franklin:1988:RU

[Fra88b] Allan Franklin. Review: [untitled]. British Journal for the Philosophy of Science, 39(3):411– 414, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687219.

Francescotti:1991:EMT

[Fra91] Robert M. Francescotti. Externalism and Marr’s theory of vision. British Journal for the Philosophy of Science, 42 (2):227–238, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/227.full.pdf+html;
http://www.jstor.org/
stable/688054.

Franklin:1992:Ra

[Fra92a] Allan Franklin. Review. British Journal for the Philosophy of Science, 43(2):283–285, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/283.full.pdf+
html;   http://www.jstor.
org/stable/687879.

Franklin:1992:Rb

[Fra92b] Allan Franklin. Review. British Journal for the Philosophy of Science, 43(3):431–433, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/3/431.full.pdf+
html;   http://www.jstor.
org/stable/687768.

Franks:1995:ECS

[Fra95] Bradley Franks. On explanation in the cognitive sciences: Competence, idealization, and the failure of the classical cascade. British Journal for the Philosophy of Science, 46(4): 475–502, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/4/475.full.pdf+html;
http://www.jstor.org/
stable/687894.

Franks:1999:DIC

[Fra99] B. Franks. Discussion. Idealizations, competence and explanation: a response to Patterson. British Journal for the Philosophy of Science, 50(4): 735–746, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/735.full.pdf+html;
http://www.jstor.org/
stable/40072264.

Fraser:2020:RPP

[Fra20] James D. Fraser. The real problem with perturbative quantum field theory. British Journal for the Philosophy of Science, 71(2):391–413, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
391/4091085.

Franklin:2021:WEE

[Fra21] Alexander Franklin. Whence the effectiveness of effective field theories? British Journal for the Philosophy of Science, 71(4):

1235–1259, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Freed:1953:DL

[Fre53] Lan Freed. Dualism and language. British Journal for the Philosophy of Science, III(12): 327–338, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/327.full.pdf+html;
http://www.jstor.org/
stable/685447.

Freundlich:1978:CP

[Fre78]  Yehudah  Freundlich.

Copenhagenism and Popperism. British Journal for the Philosophy of Science, 29(2):145–177, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/145.full.pdf+
html;   http://www.jstor.
org/stable/687124.

French:1984:RU

[Fre84] Steven French. Review: [untitled]. British Journal for the Philosophy of Science, 35(3):297– 300, September 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/297.full.pdf+html;
http://www.jstor.org/
stable/687480.

French:1988:GPJ

[Fre88] Steven French. A green parrot is just as much a red herring as a white shoe: a note on confirmation, background knowledge and the logico-probabilistic approach. British Journal for the Philosophy of Science, 39(4): 531–535, November 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        39/4/531.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687438.

French:1995:R

[Fre95] Steven French. Review. British Journal for the Philosophy of Science, 46(3):436–439, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/3/436.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687668.

French:1996:R

[Fre96] Steven French. Review. British Journal for the Philosophy of Science, 47(4):658–662, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/4/658.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687936.

French:1998:R

[Fre98] Steven French. Review. British Journal for the Philosophy of Science, 49(2):355–358, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/355.full.pdf+
html;   http://www.jstor.
org/stable/688124.

Freeman:2010:BRR

[Fre10] Norman Henry Freeman. Book review: Robert C. Richardson, Evolutionary Psychology as Maladapted Psychology. British Journal for the Philosophy of Science, 61(4): 883–888, December 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/4/883.full.pdf+html.

French:2011:BRE

[Fre11] Steven French. Book review:

Emily R. Grosholz, Representation and Productive Ambiguity in Mathematics and the Sciences. British Journal for the Philosophy of Science, 62(4): 895–898, December 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/895.full.pdf+html.

Friedman:1972:EST

[Fri72] Kenneth S. Friedman. Empirical simplicity as testability. British Journal for the Philosophy of Science, 23(1):

25–33, February 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/1/25.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686219.

Friedman:1982:IRF

[Fri82]  Kenneth Friedman.  Is in-

tertheoretic reduction feasible? British Journal for the Philosophy of Science, 33(1):17–40, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/1/17.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687238.

Fricke:1997:WAH

[Fri97a] Martin Frické. What are the advantages of hyperproof-like reasoning systems? British Journal for the Philosophy of Science, 48(3):397–410, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        48/3/397.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688070.

Fricker:1997:R

[Fri97b] Miranda Fricker. Review. British Journal for the Philosophy of Science, 48(4):618–620, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/618.full.pdf+
html;   http://www.jstor.
org/stable/688107.

Frisch:2000:DSP

[Fri00] M. Frisch. (Dis-)solving the puzzle of the arrow of radiation. British Journal for the Philosophy of Science, 51(3):381– 410, September 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/381.full.pdf+html;
http://www.jstor.org/
stable/3541771.

Frigg:2002:USH

[Fri02a] Roman Frigg. Ubiquity. the science of history … or why the world is simpler than we think. British Journal for the Philosophy of Science, 53(4): 585–591, December 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/4/585.full.pdf+html;
http://www.jstor.org/
stable/3541827.

Frisch:2002:NLC

[Fri02b] Mathias Frisch. Non-locality in classical electrodynamics. British Journal for the Philosophy of Science, 53(1):1–19, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/53/1/1.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541638.

Frigg:2004:WSK

[Fri04] Roman Frigg. In what sense is the Kolmogorov–Sinai entropy a measure for chaotic behaviour?

— bridging the gap between dynamical systems theory and communication theory. British Journal for the Philosophy of Science, 55(3):411–434, September 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/3/411.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541669.

Frisch:2009:CDR

[Fri09a] Mathias Frisch. Causality and dispersion: a reply to John Norton. British Journal for the Philosophy of Science, 60(3):487– 495, September 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/3/487.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592016.

Frisch:2009:MST

[Fri09b] Mathias Frisch. ‘The most sacred tenet’? causal reasoning in physics. British Journal for the Philosophy of Science, 60(3):459–474, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/3/459.full.pdf+html;
http://www.jstor.org/
stable/25592014.

Friederich:2015:SEE

[Fri15a] Simon Friederich. Symmetry, empirical equivalence, and identity. British Journal for the Philosophy of Science, 66(3):537– 559, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/537.

Frisch:2015:USR

[Fri15b] Mathias Frisch. Users, structures, and representation. British Journal for the Philosophy of Science, 66(2):285–306, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/2/285.

Frisch:2016:BRP

[Fri16] Mathias Frisch. Book review:

Peter Vickers, Understanding Inconsistent Science. British Journal for the Philosophy of Science, 67(3):913–918, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/913.

Frohlich:1969:R

[Frö69] H. Fröhlich. Reviews. British Journal for the Philosophy of Science, 20(2):167–168, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/167.full.pdf+
html;   http://www.jstor.
org/stable/686180.

Forster:2018:FIC

[FRSW18] Malcolm Forster, Garvesh Raskutti, Reuben Stern, and Naftali Weinberger. The frugal inference of causal relations. British Journal for the Philosophy of Science, 69(3):821–848, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/821/3739772.

Feduzi:2012:FIR

[FRZ12] Alberto Feduzi, Jochen Runde, and Carlo Zappia. De Finetti on the insurance of risks and uncertainties. British Journal for the Philosophy of Science, 63 (2):329–356, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/329.full.pdf+html.

Forster:1994:HTW

[FS94] Malcolm Forster and Elliott Sober. How to tell when simpler, more unified, or less ad hoc theories will provide more accurate predictions. British Journal for the Philosophy of Science, 45(1):1–35, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/1.full.pdf+html;
http://www.jstor.org/
stable/687960.

Frey:2021:RIA

[FS21] Daniel Frey and Dunja Seselja.

Robustness and idealizations in agent-based models of scientific interaction. British Journal for the Philosophy of Science, 71(4): 1411–1437, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Franklin:2024:PMS

[FS24]  Alexander  Franklin

and Vanessa A. Seifert. The problem of molecular structure just is the measurement problem. British Journal for the Philosophy of Science, 75(1):31– 59, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Fuller:2019:CQC

[Ful19]  Jonathan Fuller.  The con-

founding question of confounding causes in randomized trials. British Journal for the Philosophy of Science, 70(3):901– 926, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
901/4819269.

Futch:2005:LC

[Fut05] Michael J. Futch. Leibnizian causation. British Journal for the Philosophy of Science, 56(3): 451–467, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/3/451.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541714.

French:2011:TNT

[FV11] Steven French and Peter Vickers. Are there no things that are scientific theories? British Journal for the Philosophy of Science, 62(4):771–804, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/4/771.full.pdf+html.

Fitelson:2005:BCA

[FW05] Branden Fitelson and Andrew Waterman. Bayesian confirmation and auxiliary hypotheses revisited: a reply to Strevens. British Journal for the Philosophy of Science, 56(2):293–302, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/56/2/293.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541703.

Fitelson:2007:CBC

[FW07]  Branden  Fitelson  and

Andrew Waterman. Comparative Bayesian confirmation and the Quine–Duhem problem: a rejoinder to Strevens. British Journal for the Philosophy of Science, 58(2):333–338, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/2/333.full.pdf+html.

Frigg:2021:CSP

[FW21] Roman Frigg and Charlotte Werndl. Can somebody please say what Gibbsian statistical mechanics says? British Journal for the Philosophy of Science, 72(1): 105–129, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gabor:1957:R

[Gab57]  D. Gabor.  Reviews.

British Journal for the Philosophy of Science, VIII(29):82–83, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/82.full.
pdf+html;   http://www.
jstor.org/stable/685390.

Gabor:1958:R

[Gab58] D. Gabor. Reviews. British Journal for the Philosophy of Science, IX(34):168–170, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/168.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/685157.

Gabor:1966:R

[Gab66]  D.  Ga-

bor. Reviews. British Journal for the Philosophy of Science, XVI(64):351–352, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/64/351.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686842.

Gallie:1957:WMS

[Gal57] W. B. Gallie. What makes a subject scientific? British Journal for the Philosophy of Science, VIII(30):118–139, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/30/118.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685791.

Gallie:1959:RDH

[Gal59] W. B. Gallie. Reply to David Harrah’s discussion note. British Journal for the Philosophy of Science, X(39):239–240, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/39/239.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686059.

Galison:1981:KQC

[Gal81] Peter Galison. Kuhn and the quantum controversy. British Journal for the Philosophy of Science, 32(1):71–85, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/1/71.full.pdf+
html;   http://www.jstor.
org/stable/687388.

Gale:2000:BRB

[Gal00a] G. Gale. Book review: The Cambridge Companion to Galileo. P Machamer [ed]. British Journal for the Philosophy of Science, 51(3):539–543, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/539.full.pdf+html;
http://www.jstor.org/
stable/3541785.

Gallie:2000:BRB

[Gal00b] R. Gallie. Book review: Other times: Philosophical Perspectives on Past, Present and Future. D. Cockburn. British Journal for the Philosophy of Science, 51(2):371–373, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/2/371.full.pdf+html;
http://www.jstor.org/
stable/3541815.

Galavotti:2003:HJP

[Gal03] Maria Carla Galavotti. Harold Jeffreys’ Probabilistic Epistemology: Between Logicism and Subjectivism. British Journal for the Philosophy of Science, 54 (1):43–57, March 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/1/43.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541879.

Gallow:2014:HLT

[Gal14] J. Dmitri Gallow. How to learn from theory-dependent evidence; or commutativity and holism: a solution for conditionalizers. British Journal for the Philosophy of Science, 65(3): 493–519, September 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/3/493.

Ganz:1971:R

[Gan71] J. S. Ganz. Reviews. British Journal for the Philosophy of Science, 22(1):70–72, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/70.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686654.

Gandevia:1986:HFG

[Gan86] S. C. Gandevia. A human factor in ‘good’ experiments. British Journal for the Philosophy of Science, 37(4):463– 466, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687371.

Gannett:2004:BRR

[Gan04] Lisa Gannett. The biological reification of race. British Journal for the Philosophy of Science, 55(2):323–345, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/2/323.full.pdf+html;
http://www.jstor.org/
stable/3541694.

Gangopadhyay:2011:AGS

[Gan11] Nivedita Gangopadhyay. Book review: Alvin I. Goldman, Simulating Minds: the Philosophy, Psychology and Neuroscience of Mindreading. British Journal for the Philosophy of Science, 62 (2):437–441, June 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/437.full.pdf+html.

Gandenberger:2015:NPL

[Gan15] Greg Gandenberger. A new proof of the likelihood principle. British Journal for the Philosophy of Science, 66(3):475– 503, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/3/475.

Gao:2022:PMR

[Gao22] Shan Gao. Protective measurements and the reality of the wave function. British Journal for the Philosophy of Science, 73(3): 777–794, September 2022. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gardner:1962:NPP

[Gar62] Martin Gardner. A new prediction paradox. British Journal for the Philosophy of Science, XIII(49):51a–51, May 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/49/51-a.full.pdf+html.

Gardiner:1967:R

[Gar67]  Patrick Gardiner.  Reviews.

British Journal for the Philosophy of Science, 17(4):328– 331, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/4/328.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686777.

Gardner:1972:QTR

[Gar72a] Michael R. Gardner. Quantumtheoretical realism: Popper and Einstein v. Kochen and Specker. British Journal for the Philosophy of Science, 23(1):13–23, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/1/13.full.pdf+
html;   http://www.jstor.
org/stable/686218.

Gardner:1972:TDL

[Gar72b] Michael R. Gardner. Two deviant logics for quantum theory: Bohr and Reichenbach. British Journal for the Philosophy of Science, 23(2):89–109, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/2/89.full.pdf+html;
http://www.jstor.org/
stable/686433.

Gardner:1973:ACB

[Gar73a] Michael R. Gardner. Apparent conflicts between Quine’s indeterminacy thesis and his philosophy of science. British Journal for the Philosophy of Science, 24(4):381–393, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/4/381.full.pdf+html;
http://www.jstor.org/
stable/686618.

Gardner:1973:LF

[Gar73b] Michael R. Gardner. A Landé Festschrift. British Journal for the Philosophy of Science, 24 (1):72–78, March 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/1/72.full.pdf+html;
http://www.jstor.org/
stable/686671.

Garner:1974:R

[Gar74]  Jennifer Garner.  Reviews.

British Journal for the Philosophy of Science, 25(3):295– 297, September 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/3/295.full.pdf+html;
http://www.jstor.org/
stable/686913.

Gardner:1977:RR

[Gar77] Michael R. Gardner. Relationism and Relativity. British Journal for the Philosophy of Science, 28(3):215–233, September 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/3/215.full.pdf+html;
http://www.jstor.org/
stable/686807.

Gardner:1982:PNF

[Gar82] Michael R. Gardner. Predicting novel facts. British Journal for the Philosophy of Science, 33(1):1–15, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/1/1.full.pdf+html;

http://www.jstor.org/
stable/687237.

Gardenfors:1984:DBB

[Gär84] Peter Gärdenfors. The dynamics of belief as a basis for logic. British Journal for the Philosophy of Science, 35(1):1–10, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/1.full.pdf+
html;   http://www.jstor.
org/stable/687553.

Garson:1999:BRB

[Gar99] J. W. Garson. Book review:

Connectionism and the Philosophy of Psychology. T Horgan, J Tienson. British Journal for the Philosophy of Science, 50 (2):319–323, June 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/319.full.pdf+html;
http://www.jstor.org/
stable/40072231.

Garvey:2007:PB

[Gar07] Brian Garvey. Philosophy of biology. Philosophy and science (Montréal, Québec). McGill-Queen’s University Press, Montréal, Québec, Canada, 2007. ISBN 0-7735-3342-7 (bound), 0-7735-3343-5 (paperback). xiii + 274 pp. LCCN QH331 .G37 2007. URL

http://www.loc.gov/catdir/
toc/fy0805/2008428220.html.

Garzon:2008:TGT

[Gar08] Francisco Calvo Garzón. Towards a general theory of antirepresentationalism. British Journal for the Philosophy of Science, 59(3):259–292, September 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/3/259.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072285.

Garson:2018:HFP

[Gar18] Justin Garson. How to be a function pluralist. British Journal for the Philosophy of Science, 69(4):1101–1122, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/4/1101/3952641.

Garson:2023:PHB

[Gar23] Justin Garson. Putting history back into mechanisms. British Journal for the Philosophy of Science, 74(4):921–940, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gasman:1975:MXR

[Gas75] L. D. Gasman. Myths and Xrays. British Journal for the Philosophy of Science, 26(1):51–60, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/1/51.full.pdf+

html;   http://www.jstor.
org/stable/686334.

Gattei:2001:BRK

[Gat01] Stefano Gattei. Book review:

Karl Popper — The Formative Years, 1902–1945: Politics and Philosophy in Interwar Vienna. British Journal for the Philosophy of Science, 52(4):815–825, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/4/815.full.pdf+html;
http://www.jstor.org/
stable/3541922.

Gauld:1966:CMP

[Gau66] Alan Gauld. Could a machine perceive? British Journal for the Philosophy of Science, 17(1):44–58, May 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686403.

Gayon:1998:DSS

[Gay98] Jean Gayon. Darwinism’s struggle for survival: heredity and the hypothesis of natural selection. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56250-3 (hardcover). xvi + 516 pp. LCCN QH375 .G288d 1998; QH375 .G3813 1998.

Ghirardi:1999:DDR

[GB99] G. C. Ghirardi and A. Bassi.

Do dynamical reduction models imply that arithmetic does not apply to ordinary macroscopic objects? British Journal for the Philosophy of Science, 50(1):49–64, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/49.full.pdf+html;
http://www.jstor.org/
stable/40072207.

Graham:1961:MPR

[GBW61a] A. Graham, G. W. Scott Blair, and R. F. J. Withers. A methodological problem in rheology. British Journal for the Philosophy of Science, XI(44): 265–288, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685129.

Graseam:1961:MPR

[GBW61b] A. Graseam, G. W. Scoot Blair, and R. F. J. Withers. A methodological problem in rheology: I: Experimental evidence on the rheology of complex alloys and its philosophical significance. British Journal for the Philosophy of Science, XI(44): 265–280, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/265.full.pdf+html.

Gross:2011:RLI

[GC11] Steven Gross and Jennifer Culbertson. Revisited linguistic intuitions. British Journal for the Philosophy of Science, 62(3): 639–656, September 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/639.full.pdf+html.

Gonzalez-Cabrera:2019:STE

[GC19] Ivan Gonzalez-Cabrera. On social tolerance and the evolution of human normative guidance. British Journal for the Philosophy of Science, 70(2):523–549, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/70/2/523/4345800.

Gebharter:2023:QPL

[GE23]  Alexander  Gebharter

and Markus I. Eronen. Quantifying proportionality and the limits of higher-level causation and explanation. British Journal for the Philosophy of Science, 74(3): 573–601, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Geach:1965:MBP

[Gea65] Peter Geach. Man, beast, and philosophical psychology. British Journal for the Philosophy of Science, XVI(63):248, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/63/248.full.

                                        pdf+html;   http://www.
                                        jstor.org/stable/686309.

Geach:1969:R

[Gea69] P. T. Geach. Reviews. British Journal for the Philosophy of Science, 20(3):277–279, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/277.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686427.

Gebharter:2022:CBN

[Geb22] Alexander Gebharter. A causal Bayes net analysis of Glennan’s mechanistic account of higherlevel causation (and some consequences). British Journal for the Philosophy of Science, 73(1): 185–210, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gellner:1959:WTC

[Gel59] Ernest Gellner. Words and things: a critical account of linguistic philosophy and a study in ideology. Victor Gollancz, London, UK, 1959. 270 pp. With an introduction by Bertram Russell.

Gellner:1961:R

[Gel61] Ernest Gellner. Reviews. British Journal for the Philosophy of Science, XI(44):345–346, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/44/345.full.pdf+

html;   http://www.jstor.
org/stable/685140.

Gellner:1975:BTF

[Gel75] Ernest Gellner. Beyond truth and falsehood. British Journal for the Philosophy of Science, 26(4):331–342, December 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/4/331.full.pdf+html;
http://www.jstor.org/
stable/686679.

Gellner:1983:R

[Gel83] Ernest Gellner. Reviews. British Journal for the Philosophy of Science, 34(2):173–175, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/173.full.pdf+
html;   http://www.jstor.
org/stable/687448.

Gellner:1985:RSS

[Gel85] Ernest Gellner. Relativism and the social sciences. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26530-4. viii + 200 pp. LCCN HM26 .G45 1985. URL http://www.

loc.gov/catdir/description/
cam022/84017042.html;
http://www.loc.gov/catdir/
toc/cam023/84017042.html.

Gemes:1989:RPI

[Gem89] Ken Gemes. A refutation of Popperian inductive scepticism.

British Journal for the Philosophy of Science, 40(2):183–184, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/183.full.pdf+
html;   http://www.jstor.
org/stable/687508.

Gendler:1998:GIS

[Gen98] Tamar Szabó Gendler. Galileo and the indispensability of scientific thought experiment. British Journal for the Philosophy of Science, 49(3):397–424, September 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/397.full.pdf+
html;   http://www.jstor.
org/stable/688082.

George:1952:R

[Geo52] Philip George. Reviews. British Journal for the Philosophy of Science, III(11):273–275, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/273.full.
pdf+html;   http://www.
jstor.org/stable/685273.

George:1953:R

[Geo53] Philip George. Reviews. British Journal for the Philosophy of Science, IV(14):165–168, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/165.full.pdf+
html;   http://www.jstor.
org/stable/685993.

Gardner:1986:RDU

[GG86] R. Allen Gardner and Beatrix T.

Gardner. Review: Discovering and understanding the meaning of primate signals. British Journal for the Philosophy of Science, 37(4):477–495, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687373.

Glymour:2016:RIN

[GH16] Clark Glymour and Catherine Hanson. Reverse inference in neuropsychology. British Journal for the Philosophy of Science, 67(4):1139–1153, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/4/1139.

Ghose:2006:TQM

[Gho06] Partha Ghose. Testing quantum mechanics on new ground. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-55463-2 (hardback) 0-521-02659-8 (paperback). xiv + 208 pp. LCCN QC174.12 2006. US$22.99.

URL http://www.gbv.de/dms/
bowker/toc/9780521026598.

Giaquinto:1983:HPM

[Gia83] Marcus Giaquinto. Hilbert’s philosophy of mathematics. British Journal for the Philosophy of Science, 34(2):119–132, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/119.full.pdf+
html;   http://www.jstor.
org/stable/687445.

Giaquinto:1994:EVT

[Gia94] Marcus Giaquinto. Epistemology of visual thinking in elementary real analysis. British Journal for the Philosophy of Science, 45(3):789–813, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/789.full.pdf+
html;   http://www.jstor.
org/stable/687794.

Giardino:2006:BRA

[Gia06a] Valeria Giardino. Book review: Arturo Carsetti, Seeing, Thinking and Knowing: Meaning and Self-Organisation in Visual Cognition and Thought, Dordrecht, the Netherlands: Kluwer Academic Publishers, 2004 Hardback £97.00 ISBN 1-4020-2080-5. British Journal for the Philosophy of Science, 57(3):623– 625, September 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

57/3/623.full.pdf+html;
http://www.jstor.org/
stable/3873485.

Giardino:2006:BRM

[Gia06b] Valeria Giardino. Book review:

Mohan Matthen, Seeing, Doing and Knowing: a Philosophical Theory of Sense Perception, Oxford: Oxford University Press, 2005 Hardback £40.00 ISBN 0-19-926850-9. British Journal for the Philosophy of Science, 57(4):781–783, December 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/4/781.full.pdf+html;
http://www.jstor.org/
stable/4489094.

Gibson:1969:RRL

[Gib69] L. Gibson. On ‘ravens and relevance’ and a likelihood solution of the paradox of confirmation. British Journal for the Philosophy of Science, 20(1):75–80, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/75.full.pdf+
html;   http://www.jstor.
org/stable/686720.

Gibbon:1981:R

[Gib81] Guy Gibbon. Reviews. British Journal for the Philosophy of Science, 32(1):106–109, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/1/106.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687393.

Gibbins:1982:R

[Gib82] Peter Gibbins. Reviews. British Journal for the Philosophy of Science, 33(2):209–217, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/209.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687280.

Gibbins:1984:RNC

[Gib84] Peter Gibbins. Review: Nancy Cartwright’s new philosophy of physics. British Journal for the Philosophy of Science, 35(4): 390–402, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        35/4/390.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687343.

Gibbins:1986:HCE

[Gib86] Peter Gibbins. Book review: Hookway, Christopher ( ed. ) [1984]: Minds, Machines and Evolution. Cambridge University Press. Pp. xi + 177 (ISBN 0-521-26547-9). British Journal for the Philosophy of Science, 37(3):369–371, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687297.

Gibson:1989:R

[Gib89] Roger F. Gibson, Jr. Review. British Journal for the Philosophy of Science, 40(4): 557–567, December 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/4/557.full.pdf+html;
http://www.jstor.org/
stable/687741.

Giedymin:1963:RI

[Gie63]  Jerzy Giedymin.  Reliability

of informants. British Journal for the Philosophy of Science, XIII(52):287–302, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/287.full.
pdf+html;   http://www.
jstor.org/stable/685326.

Giedymin:1970:PMV

[Gie70] Jerzy Giedymin. The paradox of meaning variance. British Journal for the Philosophy of Science, 21(3):257–268, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/257.full.pdf+
html;   http://www.jstor.
org/stable/686196.

Giedymin:1971:CI

[Gie71] Jerzy Giedymin. Consolations for the irrationalist? British Journal for the Philosophy of Science, 22(1):39–48, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/39.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686642.

Giedymin:1972:QPN

[Gie72a] Jerzy Giedymin. Quine’s philosophical naturalism. British Journal for the Philosophy of Science, 23(1):45–55, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        23/1/45.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686221.

Giere:1972:STC

[Gie72b] Ronald N. Giere. The significance test controversy. British Journal for the Philosophy of Science, 23(2):170–181, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/23/2/170.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686441.

Giedymin:1973:LCC

[Gie73a] Jerzy Giedymin. Logical comparability and conceptual disparity between Newtonian and relativistic mechanics. British Journal for the Philosophy of Science, 24(3):270–276, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/3/270.full.pdf+
html;   http://www.jstor.
org/stable/686232.

Giedymin:1973:R

[Gie73b]  Jerzy Giedymin.  Reviews.

British Journal for the Philosophy of Science, 24(3):307– 313, September 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/3/307.full.pdf+html;
http://www.jstor.org/
stable/686236.

Giere:1973:HPS

[Gie73c] Ronald N. Giere. History and philosophy of science: Intimate relationship or marriage of convenience? British Journal for the Philosophy of Science, 24(3):282–297, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/3/282.full.pdf+html;
http://www.jstor.org/
stable/686234.

Giedymin:1974:R

[Gie74]  Jerzy Giedymin.  Reviews.

British Journal for the Philosophy of Science, 25(2):189–195, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/25/2/189.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686823.

Giedymin:1975:ACP

[Gie75] Jerzy Giedymin. Antipositivism in contemporary philosophy of social science and humanities. British Journal for the Philosophy of Science, 26(4):275– 301, December 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/4/275.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686676.

Gillies:1968:R

[Gil68] D. A. Gillies. Reviews. British Journal for the Philosophy of Science, 18(4):348–349, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/4/348.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686369.

Gillies:1971:FRP

[Gil71] Donald A. Gillies. A falsifying rule for probability statements. British Journal for the Philosophy of Science, 22(3):231–261, August 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/3/231.full.pdf+
html;   http://www.
jstor.org/stable/686745. See

comment [De 72].

Gillies:1972:RM

[Gil72a] D. A. Gillies. Reply to De Maré. British Journal for the Philosophy of Science, 23(4): 335–336, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/335-b.full.pdf+html;
http://www.jstor.org/
stable/686735.

Gillies:1972:STP

[Gil72b] D. A. Gillies. The subjective theory of probability. British Journal for the Philosophy of Science, 23(2):138–157, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/2/138.full.pdf+html;
http://www.jstor.org/
stable/686439.

Gillies:1978:DM

[Gil78] Donald Gillies. The development of mathematics. British Journal for the Philosophy of Science, 29(1):68–87, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/1/68.full.pdf+html;
http://www.jstor.org/
stable/686394.

Gillies:1980:PIM

[Gil80] D. A. Gillies. Phenomenology and the infinite in mathematics. British Journal for the Philosophy of Science, 31(3): 289–298, September 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/3/289.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686923.

Gillies:1981:KMP

[Gil81] Donald Gillies. Karl Menger as a philosopher. British Journal for the Philosophy of Science, 32(2):183–196, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/2/183.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687200.

Gillies:1982:R

[Gil82a] D. A. Gillies. Reviews. British Journal for the Philosophy of Science, 33(2):217–220, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/217.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687281.

Gillies:1982:WR

[Gil82b] D. A. Gillies. Wittgenstein and revisionism. British Journal for the Philosophy of Science, 33(4):422–433, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/4/422.full.pdf+html;
http://www.jstor.org/
stable/687173.

Gillies:1988:KM

[Gil88] Donald Gillies. Keynes as a methodologist. British Journal for the Philosophy of Science, 39 (1):117–129, March 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/1/117.full.pdf+html;
http://www.jstor.org/
stable/687118.

Gillett:1989:PN

[Gil89] Grant Gillett. Perception and neuroscience. British Journal for the Philosophy of Science, 40 (1):83–103, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/1/83.full.pdf+html;
http://www.jstor.org/
stable/687466.

Gillies:1990:TGW

[Gil90] Donald Gillies. The Turing– Good weight of evidence function and Popper’s measure of the severity of a test. British Journal for the Philosophy of Science, 41(1):143–146, March 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        41/1/143.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688010.

Gillies:1991:BPH

[Gil91a] Donald Gillies. A Bayesian proof of a Humean principle. British Journal for the Philosophy of Science, 42(2):255–256, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/42/2/255.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688056.

Gillies:1991:IPC

[Gil91b]  Donald Gillies.  Intersubjec-

tive probability and confirmation theory. British Journal for the Philosophy of Science, 42(4):513–533, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/4/513.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687706.

Gillies:1992:R

[Gil92a] Donald A. Gillies. Review. British Journal for the Philosophy of Science, 43(2):263–278, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/2/263.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687877.

Gilman:1992:WTD

[Gil92b] Daniel Gilman. What’s a theory to do … with seeing? or some empirical considerations for observation and theory. British Journal for the Philosophy of Science, 43(3):287–309, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/3/287.full.pdf+
html;   http://www.jstor.
org/stable/687761.

Gillett:1993:EIC

[Gil93] Grant Gillett. Explaining intentions: Critical review of explaining behaviour. British Journal for the Philosophy of Science, 44 (1):157–165, March 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/1/157.full.pdf+html;
http://www.jstor.org/
stable/687855.

Gillies:1996:R

[Gil96a] Donald Gillies. Review. British Journal for the Philosophy of Science, 47(3):476, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
47/3/476.full.pdf+html.

Gillies:1996:RU

[Gil96b] Donald Gillies. Review: [untitled]. British Journal for the Philosophy of Science, 47(3):476– 478, September 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687992.

Gill:2000:BRB

[Gil00a] M. L. Gill. Book review: The Order of Nature in Aristotle’s Physics: Place and the Elements. H. S. Lang. British Journal for the Philosophy of Science, 51(3):549–557, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/549.full.pdf+html;
http://www.jstor.org/
stable/3541787.

Gillies:2000:VP

[Gil00b] Donald Gillies. Varieties of propensity. British Journal for the Philosophy of Science, 51(4): 807–835, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/807.full.pdf+html;
http://www.jstor.org/
stable/3541728.

Gillies:2001:BRB

[Gil01]  Donald Gillies.  Book re-

view: Causality: Models, Reasoning, and Inference, Judea Pearl. British Journal for the Philosophy of Science, 52(3): 613–622, September 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/613.full.pdf+html;
http://www.jstor.org/
stable/3541932.

Gillies:2003:RMD

[Gil03] Donald Gillies. Review: Michael Dummett: La natura e il futuro della filosofia, ISBN 88-7018-386-6. (Italian) [Nature and the future of philosophy]. British Journal for the Philosophy of Science, 54(3):501–507, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/3/501.full.pdf+
html;   http://www.jstor.
org/stable/3541797.

Gillies:2005:ART

[Gil05] Donald Gillies. An action-related theory of causality. British Journal for the Philosophy of Science, 56(4):823–842, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/823.full.pdf+html;
http://www.jstor.org/
stable/3541869.

Gillies:2007:BRK

[Gil07] Donald Gillies. Book review:

K. Codell Carter, The Rise of Causal Concepts of Disease. Case Histories, Ashgate, 2003. ISBN 0-7546-0678-3. British Journal for the Philosophy of Science, 58(2):365–377, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/58/2/365.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/30115234.

Gildenhuys:2009:ECD

[Gil09a] Peter Gildenhuys. An explication of the causal dimension of drift. British Journal for the Philosophy of Science, 60(3): 521–555, September 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/3/521.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592018.

Gillies:2009:BRH

[Gil09b] Donald Gillies. Book review: Hasok Chang, Inventing Temperature: Measurement and Scientific Progress. British Journal for the Philosophy of Science, 60 (1):221–228, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/1/221.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25591994.

Gillies:2013:BRD

[Gil13] Donald Gillies. Book review: Darrell P. Rowbottom, Popper’s Critical Rationalism: a Philosophical Investigation. British Journal for the Philosophy of Science, 64(1):229–232, March 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
64/1/229.full.pdf+html.

Gildenhuys:2014:ACC

[Gil14] Peter Gildenhuys. Arbitrariness and causation in classical population genetics. British Journal for the Philosophy of Science, 65(3):429–444, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/3/429.

Gimbel:2004:RAA

[Gim04] Steven Gimbel. Restoring ambiguity to Achinstein’s account of evidence. British Journal for the Philosophy of Science, 55 (2):269–285, June 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/2/269.full.pdf+html;
http://www.jstor.org/
stable/3541691.

Gingerich:1984:ATC

[Gin84] Owen Gingerich, editor. Astrophysics and twentieth-century astronomy to 1950, volume 4 of The General history of astronomy. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-24256-8. ???? pp. LCCN QB15 .G38 1984 vol. 4

QB461. URL http://www.loc.

gov/catdir/enhancements/
fy1002/83010164-d.html;
http://www.loc.gov/catdir/
enhancements/fy1002/
83010164-t.html.

Ginev:1993:DWN

[Gin93]  D. Ginev.  Do we need

fixed methodological principles? British Journal for the Philosophy of Science, 44(2):329–334, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/2/329.full.pdf+
html;   http://www.jstor.
org/stable/687652.

Giray:1976:IBA

[Gir76] Erol F. Giray. An integrated biological approach to the species problem. British Journal for the Philosophy of Science, 27(4): 317–328, December 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/317.full.pdf+html;
http://www.jstor.org/
stable/686861.

Giunti:1988:HTS

[Giu88] Marco Giunti. Hattiangadi’s theory of scientific problems and the structure of standard epistemologies. British Journal for the Philosophy of Science, 39(4): 421–439, November 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

39/4/421.full.pdf+html;
http://www.jstor.org/
stable/687430.

Giulini:2001:US

[Giu01] Domenico Giulini. Uniqueness of simultaneity. British Journal for the Philosophy of Science, 52(4):651–670, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/4/651.full.pdf+html;
http://www.jstor.org/
stable/3541912.

Glass:1988:MME

[GJ88] J. C. Glass and W. Johnson. Metaphysics, msrp and economics. British Journal for the Philosophy of Science, 39(3): 313–329, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/313.full.pdf+html;
http://www.jstor.org/
stable/687209.

Garcia-Lapena:2023:TQD

[GL23] Alfonso García-Lapeña. Truthlikeness for quantitative deterministic laws. British Journal for the Philosophy of Science, 74(3): 649–679, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Glassen:1984:OAP

[Gla84] Peter Glassen. O’Hear on an argument of Popper’s. British Journal for the Philosophy of Science, 35(4):375–377, December 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/4/375.full.pdf+
html;   http://www.jstor.
org/stable/687340.

Glackin:2014:BRH

[Gla14] Shane N. Glackin. Book review: Havi Carel and Rachel Cooper, Health, Illness, and Disease: Philosophical Essays. British Journal for the Philosophy of Science, 65(2):413–417, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/65/2/413.

Glass:2021:CEH

[Gla21] David H. Glass. Coherence, explanation, and hypothesis selection. British Journal for the Philosophy of Science, 72(1):1– 26, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gladziejewski:2023:DOO

[Gła23]  Paweł Gładziejewski.  Un-

debunking ordinary objects with the help of predictive processing. British Journal for the Philosophy of Science, 74(4):1047– 1068, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Glazier:2024:USM

[Gla24]  Martin Glazier.  The unity

of science and the mentaculus. British Journal for the Philosophy of Science, 75(2):275– 297, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Glick:2020:GMO

[Gli20] David Glick. Generalism and the metaphysics of ontic structural realism. British Journal for the Philosophy of Science, 71(2):751–772, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
751/4815700.

Glymour:1994:MCN

[Gly94] Clark Glymour. On the methods of cognitive neuropsychology. British Journal for the Philosophy of Science, 45(3):815– 835, September 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/3/815.full.pdf+html;
http://www.jstor.org/
stable/687795.

Glymour:2004:BRJ

[Gly04]  Clark Glymour.  Book re-

view: James Woodward, Making Things Happen: A Theory of Causal Explanation, Oxford, 2003, 418pp, $65.00 (hbk) ISBN 0-19-515527-0. British Journal for the Philosophy of Science, 55(4):779–790, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/779.full.pdf+html;
http://www.jstor.org/
stable/3541630.

Glymour:2008:CIG

[Gly08] Bruce Glymour. Correlated interaction and group selection. British Journal for the Philosophy of Science, 59(4):835– 855, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/835.full.pdf+html;
http://www.jstor.org/
stable/40072314.

Glymour:2009:JWB

[Gly09] Clark Glymour. Jon Williamson: Bayesian nets and causality. British Journal for the Philosophy of Science, 60(4):849– 855, December 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/849.full.pdf+html;
http://www.jstor.org/
stable/25592038. See response

[Wil09a].

Glymour:2010:WRB

[Gly10a] Clark Glymour. What is right with ‘Bayes Net Methods’ and what is wrong with ‘Hunting Causes and Using Them’? British Journal for the Philosophy of Science, 61 (1):161–211, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/161.full.pdf+html.

Glynn:2010:DC

[Gly10b]  Luke Glynn.  Deterministic

chance. British Journal for the Philosophy of Science, 61 (1):51–80, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/51.full.pdf+html.

Glynn:2011:BRD

[Gly11a] Luke Glynn. Book review: D.

H. Mellor,The Matter of Chance. British Journal for the Philosophy of Science, 62(4):899– 906, December 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/899.full.pdf+html.

Glynn:2011:PAC

[Gly11b] Luke Glynn. A probabilistic analysis of causation. British Journal for the Philosophy of Science, 62(2):343–392, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/343.full.pdf+html.

Glymour:2015:PEV

[Gly15] Clark Glymour. Probability and the explanatory virtues. British Journal for the Philosophy of Science, 66(3):591–604, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/66/3/591.

Griffiths:2018:EDD

[GM18] Paul E. Griffiths and John Matthewson. Evolution, dysfunction, and disease: a reappraisal. British Journal for the Philosophy of Science, 69(2):301–327, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/2/301/2670197.

Godlovitch:1969:UBI

[God69] S. Godlovitch. Universal, basic and instantial statements in the logic of scientific discovery. British Journal for the Philosophy of Science, 20(4): 355–356, December 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/4/355.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686263.

Goldstein:1958:TTM

[Gol58] Leon J. Goldstein. The two theses of methodological individualism. British Journal for the Philosophy of Science, IX(33):1– 11, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/1.full.pdf+
html;   http://www.jstor.
org/stable/686096.

Goldstein:1959:MWT

[Gol59] Leon J. Goldstein. Mr Watkins on the two theses. British Journal for the Philosophy of Science, X(39):240–241, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/240.full.pdf+html;
http://www.jstor.org/
stable/686060.

Goldie:1998:R

[Gol98] Peter Goldie. Review. British Journal for the Philosophy of Science, 49(4):642–648, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/4/642.full.pdf+
html;   http://www.jstor.
org/stable/688136.

Goldman:2006:SMP

[Gol06] Alvin I. Goldman. Simulating Minds: The Philosophy, Psychology, and Neuroscience of Mindreading. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-513892-9. 336 pp. LCCN BF323.S63

                                        G65 2006. URL http://www.
                                        oxfordscholarship.com/oso/
                                        public/content/philosophy/
                                        9780195138924/toc.htm.

Gombrich:1954:R

[Gom54]  E. H. Gombrich.  Reviews.

British Journal for the Philosophy of Science, V(18):168–169, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/18/168.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685179.

Goodstein:1952:ML

[Goo52] R. L. Goodstein. Mathematical logic. British Journal for the Philosophy of Science, III (10):197–200, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/10/197.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685560.

Good:1958:R

[Goo58]  I.  J.

Good. Reviews. British Journal for the Philosophy of Science, IX(35):251–255, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/35/251.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685660.

Good:1959:LSS

[Goo59a] I. J. Good. Lattice structure of space–time. British Journal for the Philosophy of Science, IX(36):317–319, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/317.full.pdf+html;
http://www.jstor.org/
stable/685192.

Good:1959:TC

[Goo59b] I. J. Good. A theory of causality. British Journal for the Philosophy of Science, IX(36): 307–310, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/307.full.pdf+html;
http://www.jstor.org/
stable/685190.

Good:1960:PC

[Goo60a] I. J. Good. The paradox of confirmation. British Journal for the Philosophy of Science, XI(42):145b–145, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/42/145-b.full.pdf+html;
http://www.jstor.org/
stable/685588.

Gooddy:1960:R

[Goo60b] William Gooddy. Reviews.

British Journal for the Philosophy of Science, XI(41):84–85, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/84.full.pdf+
html;   http://www.jstor.
org/stable/685832.

Goodstein:1960:R

[Goo60c] R. L. Goodstein. Reviews. British Journal for the Philosophy of Science, XI(42):163–164, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/163.full.pdf+
html;   http://www.jstor.
org/stable/685596.

Good:1961:CC

[Goo61a] I. J. Good. A causal calculus (I). British Journal for the Philosophy of Science, XI(44): 305–318, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/305.full.pdf+html;
http://www.jstor.org/
stable/685131.

Good:1961:CCI

[Goo61b] I. J. Good. A causal calculus (II). British Journal for the Philosophy of Science, XII(45):43– 51, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/43.full.pdf+

html;   http://www.jstor.
org/stable/685537.

Good:1961:PCI

[Goo61c] I. J. Good. The paradox of confirmation (II). British Journal for the Philosophy of Science, XII(45):63–64, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/45/63.full.pdf+html;
http://www.jstor.org/
stable/685539.

Good:1962:EC

[Goo62] I. J. Good. Errata and corrigenda. British Journal for the Philosophy of Science, XIII(49): 88, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/88.full.
pdf+html;   http://www.
jstor.org/stable/685237.

Goodstein:1963:SIT

[Goo63] R. L. Goodstein. The significance of incompleteness theorems. British Journal for the Philosophy of Science, XIV(55): 208–220, November 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/55/208.full.pdf+html;
http://www.
jstor.org/stable/685241. See

note [Fin64, Goo64].

Goodstein:1964:RMF

[Goo64] R. L. Goodstein. Reply to Mr Fine’s note. British Journal for the Philosophy of Science, XV(58):141, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XV/58/141-a.full.pdf+html;
                                         http://www.
                                        jstor.org/stable/685254. See

[Goo63, Fin64].

Good:1967:HML

[Goo67a] I. J. Good. Human and machine logic. British Journal for the Philosophy of Science, 18 (2):144–147, August 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/2/144.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686582.

Good:1967:PTE

[Goo67b] I. J. Good. On the principle of total evidence. British Journal for the Philosophy of Science, 17(4):319–321, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/4/319.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686773.

Good:1967:WSR

[Goo67c] I. J. Good. The white shoe is a red herring. British Journal for the Philosophy of Science, 17(4):322, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/322.full.pdf+html;
http://www.jstor.org/
stable/686774.

Good:1968:CEE

[Goo68a] I. J. Good. Corroboration, explanation, evolving probability, simplicity and a sharpened razor. British Journal for the Philosophy of Science, 19(2):123–143, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/123.full.pdf+
html;   http://www.jstor.
org/stable/686791.

Good:1968:WSQ

[Goo68b] I. J. Good. The white shoe qua herring is pink. British Journal for the Philosophy of Science, 19 (2):156–157, August 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/2/156.full.pdf+html;
http://www.jstor.org/
stable/686795.

Good:1969:GTR

[Goo69] I. J. Good. Gödel’s Theorem is a red herring. British Journal for the Philosophy of Science, 19(4):357–358, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        19/4/357.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686295.

Good:1970:SRM

[Goo70]  I. J. Good.  A suggested

resolution of Miller’s Paradox. British Journal for the Philosophy of Science, 21(3):288–289, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/288.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686203.

Good:1971:FWS

[Goo71a] I. J. Good. Free will and speed of computation. British Journal for the Philosophy of Science, 22 (1):48–50, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        22/1/48.full.pdf+html.

Good:1971:R

[Goo71b] I. J. Good. Reviews. British Journal for the Philosophy of Science, 22(4):382–387, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/4/382.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686322.

Good:1974:CCC

[Goo74a] I. J. Good. A correction concerning complexity. British Journal for the Philosophy of Science, 25(3):289, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/3/289.full.pdf+html;
http://www.jstor.org/
stable/686911.

Good:1974:LLC

[Goo74b] I. J. Good. A little learning can be dangerous. British Journal for the Philosophy of Science, 25(4):340–342, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/4/340.full.pdf+
html;   http://www.jstor.
org/stable/686463.

Good:1976:R

[Goo76] I. J. Good. Reviews. British Journal for the Philosophy of Science, 27(3):289–292, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/289.full.pdf+
html;   http://www.jstor.
org/stable/686127.

Good:1984:BAP

[Goo84] I. J. Good. A Bayesian approach in the philosophy of inference. British Journal for the Philosophy of Science, 35(2):161–166, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/161.full.pdf+
html;   http://www.jstor.
org/stable/686966.

Good:1985:HCC

[Goo85] I. J. Good. A historical comment concerning novel confirmation. British Journal for the Philosophy of Science, 36(2):184–185, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/2/184.full.pdf+
html;   http://www.jstor.
org/stable/687034.

Gooding:1994:IS

[Goo94] David Gooding. Imaginary science. British Journal for the Philosophy of Science, 45(4):1029– 1045, December 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/4/1029.full.pdf+html;
http://www.jstor.org/
stable/687620.

Goodwin:2009:SUS

[Goo09] William Goodwin. Scientific understanding and synthetic design. British Journal for the Philosophy of Science, 60(2): 271–301, June 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
60/2/271.full.pdf+html;
http://www.jstor.org/
stable/25592002.

Goodwin:2010:HDS

[Goo10] William Goodwin. How do structural formulas embody the theory of organic chemistry? British Journal for the Philosophy of Science, 61(3):621– 633, September 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/621.full.pdf+html.

Goodwin:2013:SCN

[Goo13] William Goodwin. Sustaining a controversy: The non-classical ion debate. British Journal for the Philosophy of Science, 64(4):787–816, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/787.full.pdf+html.

Gorovitz:1970:IOE

[Gor70] Samuel Gorovitz. Inscriptionalism and the objects of explanation. British Journal for the Philosophy of Science, 21 (3):247–256, August 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/3/247.full.pdf+html;
http://www.jstor.org/
stable/686195.

Gould:1957:PHS

[Gou57] R. P. Gould. The place of historical statements in biology. British Journal for the Philosophy of Science, VIII(31):192– 210, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/31/192.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685757.

Goudge:1958:CEN

[Gou58] T. A. Goudge. Causal explanations in natural history. British Journal for the Philosophy of Science, IX(35):194– 202, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/35/194.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685649.

Gould:1960:R

[Gou60] R. P. Gould. Reviews. British Journal for the Philosophy of Science, XI(43):259–260, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/43/259.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685318.

Gough:2023:PEM

[Gou23] Joe Gough. On the proper epistemology of the mental for psychiatry: What’s the point of understanding and explaining? British Journal for the Philosophy of Science, 74(4):975–998, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Gower:1970:R

[Gow70] B. S. Gower. Reviews. British Journal for the Philosophy of Science, 21(4):408–410, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/408.full.pdf+
html;   http://www.jstor.
org/stable/686575.

Gower:1988:CM

[Gow88] Barry Gower. Chalmers on method. British Journal for the Philosophy of Science, 39 (1):59–65, March 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/1/59.full.pdf+html;
http://www.jstor.org/
stable/687114.

Gower:1991:HP

[Gow91] Barry Gower. Hume on probability. British Journal for the Philosophy of Science, 42(1):1–19, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/1/1.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687604.

Gower:1992:R

[Gow92] Barry Gower. Review. British Journal for the Philosophy of Science, 43(4):555–559, October 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/4/555.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687841.

Gower:1993:OI

[Gow93] Barry Gower. O’Hear’s introduction. British Journal for the Philosophy of Science, 44(4): 743–758, December 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        44/4/743.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688042.

Gower:1997:R

[Gow97] Barry Gower. Review. British Journal for the Philosophy of Science, 48(1):126–131, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/1/126.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687725.

Garson:2014:FMP

[GP14] Justin Garson and Gualtiero Piccinini. Functions must be performed at appropriate rates in appropriate situations. British Journal for the Philosophy of Science, 65(1):1–20, January 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/1.full.pdf+html.

Grasshoff:2005:MAD

[GPW05] Gerd Graßhoff, Samuel Portmann, and Adrian Wüthrich. Minimal assumption derivation of a Bell-type inequality. British Journal for the Philosophy of Science, 56(4):663–680, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/4/663.full.pdf+
html;   http://www.jstor.
org/stable/3541862.

Gyenis:2015:DBP

[GR15] Zalán Gyenis and Miklós Rédei.

Defusing Bertrand’s Paradox. British Journal for the Philosophy of Science, 66(2):349–373, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/2/349.

Grant:1955:SCA

[Gra55]  C. K. Grant.  Some com-

ments on ‘the age of the universe’. British Journal for the Philosophy of Science, VI(23): 248–251, November 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/23/248.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685684.

Graves:1973:R

[Gra73]  John C. Graves.  Reviews.

British Journal for the Philosophy of Science, 24(2):183–190, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/24/2/183.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686610.

Graves:1974:UI

[Gra74] John C. Graves. Uniformity and induction. British Journal for the Philosophy of Science, 25(4):301–318, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/4/301.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686457.

Graves:1978:NPT

[Gra78]  Spencer Graves.  On the

Neyman–Pearson theory of testing. British Journal for the Philosophy of Science, 29(1):1–24, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/29/1/1.full.pdf+

html;   http://www.jstor.
org/stable/686388.

Graham:1983:R

[Gra83] Keith Graham. Reviews. British Journal for the Philosophy of Science, 34(1):84–88, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/1/84.full.pdf+
html;   http://www.jstor.
org/stable/686938.

Grantham:2001:BRK

[Gra01] Todd A. Grantham. Book review: K. Sterelny and P. E. Griffiths, Sex and Death: an Introduction to Philosophy of Biology. British Journal for the Philosophy of Science, 52(1):175–179, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/175.full.pdf+
html;   http://www.jstor.
org/stable/3541953.

Gregory:1951:CMU

[Gre51] Joshua C. Gregory. The concept of mind and the unconscious. British Journal for the Philosophy of Science, II(5):52– 57, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/52.full.pdf+
html;   http://www.jstor.
org/stable/685608.

Gregory:1953:PME

[Gre53] R. L. Gregory. On physical model explanations in psychology. British Journal for the Philosophy of Science, IV(15): 192–197, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/15/192.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685896.

Grene:1958:TETa

[Gre58a] Marjorie Grene. Two evolutionary theories (1). British Journal for the Philosophy of Science, IX (34):110–127, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/34/110.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685147.

Grene:1958:TETb

[Gre58b] Marjorie Grene. Two evolutionary theories (II). British Journal for the Philosophy of Science, IX(35):185–193, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/35/185.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685648.

Grene:1961:SS

[Gre61] Marjorie Grene. Statistics and selection. British Journal for the Philosophy of Science, XII (45):25–42, May 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/45/25.full.pdf+html;
http://www.jstor.org/
stable/685536.

Grene:1962:Ra

[Gre62a] Marjorie Grene. Reviews. British Journal for the Philosophy of Science, XIII(50):172–176, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/172.full.
pdf+html;   http://www.
jstor.org/stable/685971.

Grene:1962:Rb

[Gre62b] Marjorie Grene. Reviews. British Journal for the Philosophy of Science, XIII(51):255–256, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/255.full.
pdf+html;   http://www.
jstor.org/stable/685846.

Grene:1963:R

[Gre63a] Marjorie Grene. Reviews. British Journal for the Philosophy of Science, XIII(52):329–332, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/52/329.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685336.

Grene:1963:TET

[Gre63b] Marjorie Grene. Two evolutionary theories: a reply. British Journal for the Philosophy of Science, XIV(54):152–154, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIV/54/152.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685433.

Grene:1964:TET

[Gre64] Marjorie Grene. Two evolutionary theories: Reply to Dr. Carter. British Journal for the Philosophy of Science, XIV(56): 349–351, February 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/56/349.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685640.

Gretton:1970:R

[Gre70] M. Gretton. Reviews. British Journal for the Philosophy of Science, 21(4):394–396, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/4/394.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686565.

Gregory:1974:R

[Gre74] R. L. Gregory. Reviews. British Journal for the Philosophy of Science, 25(3):299–300, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/3/299.full.pdf+
html;   http://www.jstor.
org/stable/686915.

Grenander:1979:PDR

[Gre79]  M.  E.  Grenander.

Psychosis: Disease or rhetoric? British Journal for the Philosophy of Science, 30(2):177–182, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/177.full.pdf+
html;   http://www.jstor.
org/stable/686950.

Grenander:1981:COB

[Gre81] M. E. Grenander. A cautious overview of behaviour therapy. British Journal for the Philosophy of Science, 32(1):85–94, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/1/85.full.pdf+
html;   http://www.jstor.
org/stable/687389.

Grenander:1984:BTR

[Gre84] M. E. Grenander. Behaviour therapy: a response to Edward Erwin. British Journal for the Philosophy of Science, 35 (2):149–156, June 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/2/149.full.pdf+html;
http://www.jstor.org/
stable/686964.

Grene:1997:R

[Gre97] Marjorie Grene. Review. British Journal for the Philosophy of Science, 48(1):113–116, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/113.full.pdf+
html;   http://www.jstor.
org/stable/687722.

Greaves:2010:TGU

[Gre10] Hilary Greaves. Towards a geometrical understanding of the CPT theorem. British Journal for the Philosophy of Science, 61(1):27–50, March 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/27.full.pdf+html.

Greco:2011:STT

[Gre11] Daniel Greco. Significance testing in theory and practice. British Journal for the Philosophy of Science, 62(3):607– 637, September 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
62/3/607.full.pdf+html.

Green:2017:LVS

[Gre17] E. J. Green. A layered view of shape perception. British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL https://academic.oup.
com/bjps/article/68/2/355/
3044195/A-Layered-View-of-
Shape-Perception.

Grice:1968:R

[Gri68] G. R. Grice. Reviews. British Journal for the Philosophy of Science, 19(2):179–180, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/179.full.pdf+
html;   http://www.jstor.
org/stable/686802.

Griffin:1970:R

[Gri70] Nicholas Griffin. Reviews. British Journal for the Philosophy of Science, 21(4):402–403, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/402.full.pdf+html;
http://www.jstor.org/
stable/686570.

Grieder:1977:RCS

[Gri77a]  Alfons Grieder.  Relativity,

causality and the ‘substratum’.

British Journal for the Philosophy of Science, 28(1):35–48, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/1/35.full.pdf+
html;   http://www.jstor.
org/stable/686445.

Grieder:1977:R

[Gri77b] Alfons Grieder. Reviews. British Journal for the Philosophy of Science, 28(3):289–291, September 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/3/289.full.pdf+
html;   http://www.jstor.
org/stable/686812.

Grimes:1988:SEP

[Gri88] Thomas R. Grimes. Statistical explanation, probability, and counteracting conditions. British Journal for the Philosophy of Science, 39(4):495–503, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/4/495.full.pdf+
html;   http://www.jstor.
org/stable/687434.

Griffiths:1993:FAP

[Gri93] Paul E. Griffiths. Functional analysis and proper functions. British Journal for the Philosophy of Science, 44(3):409– 422, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/409.full.pdf+html;
http://www.jstor.org/
stable/688013.

Griffiths:1995:CC

[Gri95] Paul E. Griffiths. The Cronin controversy. British Journal for the Philosophy of Science, 46 (1):122–138, March 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/1/122.full.pdf+html;
http://www.jstor.org/
stable/687636.

Griffiths:1996:HTS

[Gri96] Paul E. Griffiths. The historical turn in the study of adaptation. British Journal for the Philosophy of Science, 47(4): 511–532, December 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/4/511.full.pdf+html;
http://www.jstor.org/
stable/687922.

Grimm:2006:USK

[Gri06] Stephen R. Grimm. Is understanding a species of knowledge? British Journal for the Philosophy of Science, 57(3): 515–535, September 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        57/3/515.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3873481.

Grinbaum:2007:RQT

[Gri07] Alexei Grinbaum. Reconstruction of quantum theory. British Journal for the Philosophy of Science, 58(3):387–408, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/58/3/387.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/30115181.

Griffiths:2008:BRJ

[Gri08a] Paul E. Griffiths. Book review: Jesse Prinz, Gut Reactions: A Perceptual Theory of Emotion. British Journal for the Philosophy of Science, 59(3): 559–567, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/3/559.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072299.

Grimm:2008:EIN

[Gri08b] Stephen R. Grimm. Explanatory inquiry and the need for explanation. British Journal for the Philosophy of Science, 59(3):481–497, September 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/3/481.full.pdf+html;

http://www.jstor.org/
stable/40072296.

Grllner:1957:R

[Grl57] Ernest Grllner. Reviews. British Journal for the Philosophy of Science, VII(28):360–362, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/28/360.full.
pdf+html;   http://www.
jstor.org/stable/685123.

Grosholz:1985:TEU

[Gro85a] E. R. Grosholz. Two episodes in the unification of logic and topology. British Journal for the Philosophy of Science, 36 (2):147–157, June 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/2/147.full.pdf+html;
http://www.jstor.org/
stable/687031.

Grosholz:1985:NVM

[Gro85b] Emily Grosholz. A new view of mathematical knowledge. British Journal for the Philosophy of Science, 36(1):71–78, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/71.full.pdf+
html;   http://www.jstor.
org/stable/687158.

Grobler:1990:BRR

[Gro90] Adam Grobler. Between rationalism and relativism. on Larry Laudan’s model of scientific rationality. British Journal for the Philosophy of Science, 41(4): 493–507, December 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/493.full.pdf+html;
http://www.jstor.org/
stable/687862.

Groisman:2008:ESB

[Gro08]  Berry Groisman.  The end

of Sleeping Beauty’s nightmare. British Journal for the Philosophy of Science, 59(3):409– 416, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/409.full.pdf+html;
http://www.jstor.org/
stable/40072292.

Grose:2009:BRE

[Gro09] Jonathan Grose. Book review: Eva Jablonbka and Marion J. Lamb, Evolution in Four Dimensions: Genetic, Epigenetic, Behavioral and Symbolic Variation in the History of Life. British Journal for the Philosophy of Science, 60(3):667–672, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/60/3/667.full.pdf+

html;   http://www.jstor.
org/stable/25592025.

Grose:2011:MSN

[Gro11] Jonathan Grose. Book review:

Margaret Schabas, The Natural Origins of Economics. British Journal for the Philosophy of Science, 62(2):433–436, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/433.full.pdf+html.

Grunbaum:1953:MST

[Grü53a] Adolf Grünbaum. E. A. Milne’s Scales of Time. British Journal for the Philosophy of Science, IV(16):329–331, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/329.full.pdf+html.

Grunbaum:1953:WME

[Grü53b] Adolf Grünbaum. Whitehead’s method of extensive abstraction. British Journal for the Philosophy of Science, IV(15): 215–226, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/215.full.pdf+html;
http://www.jstor.org/
stable/685898.

Grunbaum:1954:MST

[Grü54] Adolf Grünbaum. E. A. Milne’s scales of time. British Journal for the Philosophy of Science, IV(16):329–331, February 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686082.

Grunbaum:1956:HDS

[Grü56] Adolf Grünbaum. Historical determinism, social activism, and predictions in the social sciences. British Journal for the Philosophy of Science, VII(27): 236–240, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/27/236.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685878.

Grunbaum:1959:DTF

[Grü59] Adolf Grünbaum. Discussions: The falsifiability of the Lorentz– Fitzgerald contraction hypothesis. British Journal for the Philosophy of Science, X(37):48–50, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/37/48.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685775.

Grunbaum:1960:FLF

[Grü60] Adolf Grünbaum. The falsifiability of the Lorentz–Fitzgerald contraction hypothesis: a rejoinder to Professor Dingle. British Journal for the Philosophy of Science, XI(42):143–145, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/143.full.pdf+
html;   http://www.jstor.
org/stable/685586.

Grunbaum:1961:PDF

[Grü61] Adolf Grünbaum. Professor Dingle on falsifiability: a second rejoinder. British Journal for the Philosophy of Science, XII (46):153–156, August 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/46/153.full.pdf+html;
http://www.jstor.org/
stable/685496.

Gruender:1966:APT

[Gru66] David C. Gruender. The Achilles Paradox and transfinite numbers. British Journal for the Philosophy of Science, 17(3): 219–231, November 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/3/219.full.pdf+html;
http://www.jstor.org/
stable/686758.

Grunbaum:1969:PET

[Grü69a] A. Grünbaum. Are physical events themselves transiently past, present and future? a reply to H. A. C. Dobbs. British Journal for the Philosophy of Science, 20(2):145–153, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/2/145.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686176.

Grunbaum:1969:CIO

[Grü69b] Adolf Grünbaum. Can an infinitude of operations be performed in a finite time? British Journal for the Philosophy of Science, 20(3):203–218, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/203.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686415.

Grunbaum:1976:AHA

[Grü76a] Adolf Grünbaum. Ad hoc auxiliary hypotheses and falsificationism. British Journal for the Philosophy of Science, 27(4): 329–362, December 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/4/329.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686862.

Grunbaum:1976:CTA

[Grü76b] Adolf Grünbaum. Can a theory answer more questions than one of its rivals? British Journal for the Philosophy of Science, 27(1):1–23, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/1/1.full.pdf+html;
http://www.jstor.org/
stable/686375.

Grunbaum:1976:MBC

[Grü76c] Adolf Grünbaum. Is the method of bold conjectures and attempted refutations justifiably the method of science? British Journal for the Philosophy of Science, 27(2):105–136, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/2/105.full.pdf+
html;   http://www.jstor.
org/stable/686162.

Grunbaum:1986:CTA

[Grü86] Adolf Grünbaum. Can a theory answer more questions than one of its rivals? British Journal for the Philosophy of Science, 37(1):1–23, March 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
37/1/1.full.pdf+html;
http://www.jstor.org/
stable/686995.

Grunbaum:2000:NCT

[Grü00] A. Grünbaum. A new critique of theological interpretations of physical cosmology. British Journal for the Philosophy of Science, 51(1):1–43, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541747.

Grunbaum:2004:PTC

[Grü04] Adolf Grünbaum. The poverty of theistic cosmology. British Journal for the Philosophy of Science, 55(4):561–614, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/4/561.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541618.

Grunbaum:2005:RRS

[Grü05] Adolf Grünbaum. Rejoinder to Richard Swinburne’s ‘Second Reply to Grünbaum’. British Journal for the Philosophy of Science, 56(4):927–938, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/4/927.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541874. See [Swi05].

Gerrans:2008:GPC

[GS08a] Philip Gerrans and Valerie E. Stone. Generous or parsimonious cognitive architecture? cognitive neuroscience and theory of mind.

British Journal for the Philosophy of Science, 59(2):121–141, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/2/121.full.pdf+
html;   http://www.jstor.
org/stable/40072274.

Godfrey-Smith:2008:VPS

[GS08b] Peter Godfrey-Smith. Varieties of population structure and the levels of selection. British Journal for the Philosophy of Science, 59(1):25–50, March 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/1/25.full.pdf+html;
http://www.jstor.org/
stable/30116992.

Godfrey-Smith:2013:ECG

[GSK13] Peter Godfrey-Smith and Benjamin Kerr. Editor’s choice: Gestalt-switching and the evolutionary transitions. British Journal for the Philosophy of Science, 64(1):205–222, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/205.full.pdf+html.

Gryb:2016:ECT

[GT16] Sean Gryb and Karim P. Y.

Thébault. Editor’s choice: Time remains. British Journal for the Philosophy of Science, 67(3): 663–705, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/663.

Guala:2005:MEE

[Gua05] Francesco Guala. The methodology of experimental economics. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-85340-0 (hardcover), 0-521-61861-4 (paperback). xiv + 286 pp. LCCN HB131 .G83

2005. URL http://www.loc.
gov/catdir/toc/ecip055/
2004031079.html.

Guala:2007:BRC

[Gua07] Francesco Guala. Book review: Cristina Bicchieri, The Grammar of Society: the Nature and Dynamics of Social Norms. British Journal for the Philosophy of Science, 58(3):613–618, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/3/613.full.pdf+
html;   http://www.jstor.
org/stable/30115191.

Gunderson:1968:RCP

[Gun68] Keith Gunderson. Robots, consciousness, and programmed behaviour. British Journal for the Philosophy of Science, 19 (2):109–122, August 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/2/109.full.pdf+html;

http://www.jstor.org/
stable/686790.

Guthrie:1953:R

[Gut53] W. K. C. Guthrie. Reviews.

British Journal for the Philosophy of Science, IV(15):252– 254, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/252.full.pdf+html;
http://www.jstor.org/
stable/685905.

Gutting:2000:BRB

[Gut00] G. Gutting. Book review: A House Built on Sand: Exposing Postmodernist Myths About Science. N. Koertge [ed]. British Journal for the Philosophy of Science, 51(1):191–195, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/1/191.full.pdf+
html;   http://www.jstor.
org/stable/3541759.

Greaves:2014:ECS

[GW14] Hilary Greaves and David Wallace. Empirical consequences of symmetries. British Journal for the Philosophy of Science, 65(1):59–89, January 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/59.full.pdf+html.

H:1956:HW

[H56] E. H. H. Hermann Weyl. British Journal for the Philosophy of Science, VII(26):182–183, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/26/182.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685707.

Haack:1971:R

[Haa71] R. J. Haack. Reviews. British Journal for the Philosophy of Science, 22(1):67–68, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/67.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686652.

Haack:1976:PTT

[Haa76a] Susan Haack. The pragmatist theory of truth. British Journal for the Philosophy of Science, 27(3):231–249, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/3/231.full.pdf+html.

Haack:1976:R

[Haa76b] Susan Haack. Reviews. British Journal for the Philosophy of Science, 27(1):93–96, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/

content/27/1/93.full.pdf+
html;   http://www.jstor.
org/stable/686383.

Haack:1977:R

[Haa77] Susan Haack. Reviews. British Journal for the Philosophy of Science, 28(1):82–86, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/1/82.full.pdf+
html;   http://www.jstor.
org/stable/686452.

Haack:1978:WB

[Haa78] Susan Haack. The wheel and beyond. British Journal for the Philosophy of Science, 29 (2):185–188, June 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/2/185.full.pdf+html;
http://www.jstor.org/
stable/687127.

Haack:1980:R

[Haa80] Susan Haack. Reviews. British Journal for the Philosophy of Science, 31(4):411–413, December 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/4/411.full.pdf+
html;   http://www.jstor.
org/stable/687071.

Haack:1991:WPE

[Haa91] S. Haack. What is ‘The Problem of the Empirical Basis’, and does Johnny Wideawake solve it? British Journal for the Philosophy of Science, 42(3):369– 389, September 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/3/369.full.pdf+html;
http://www.jstor.org/
stable/687914.

Hacking:1964:FS

[Hac64] Ian Hacking. On the foundations of statistics. British Journal for the Philosophy of Science, XV(57):1–26, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/57/1.full.pdf+html;
http://www.jstor.org/
stable/685624.

Hacking:1965:Ra

[Hac65a] Ian Hacking. Reviews. British Journal for the Philosophy of Science, XV(60):345–346, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/345.full.pdf+
html;   http://www.jstor.
org/stable/686545.

Hacking:1965:Rb

[Hac65b] Ian Hacking. Reviews. British Journal for the Philosophy of Science, XV(60):347–349, February 1965. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/347.full.pdf+
html;   http://www.jstor.
org/stable/686546.

Hacking:1966:SP

[Hac66] Ian Hacking. Subjective probability. British Journal for the Philosophy of Science, XVI(64): 334–339, February 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/64/334.full.pdf+html;
http://www.jstor.org/
stable/686835.

Hacker:1967:R

[Hac67] P. M. S. Hacker. Reviews. British Journal for the Philosophy of Science, 18(1):74–76, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/74.full.pdf+html;
http://www.jstor.org/
stable/686895.

Hacking:1971:ETP

[Hac71a] Ian Hacking. Equipossibility theories of probability. British Journal for the Philosophy of Science, 22(4):339–355, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/4/339.full.pdf+html;
http://www.jstor.org/
stable/686316.

Hacking:1971:JBA

[Hac71b] Ian Hacking. Jacques Bernoulli’s art of conjecturing. British Journal for the Philosophy of Science, 22(3):209–229, August 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/3/209.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686744.

Hacking:1972:L

[Hac72] Ian Hacking. Likelihood. British Journal for the Philosophy of Science, 23(2):132–137, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/23/2/132.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686438.

Hacking:1979:ILP

[Hac79] Ian Hacking. Imre Lakatos’s philosophy of science. British Journal for the Philosophy of Science, 30(4):381–402, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/4/381.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687547.

Hacking:1981:KPH

[Hac81] Ian Hacking. Karl Pearson’s history of statistics. British Journal for the Philosophy of Science, 32(2):177–183, June 1981.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/2/177.full.pdf+html;
http://www.jstor.org/
stable/687199.

Hacking:1988:PIL

[Hac88] Ian Hacking. The participant irrealist at large in the laboratory. British Journal for the Philosophy of Science, 39(3): 277–294, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/277.full.pdf+html;
http://www.jstor.org/
stable/687207.

Hajek:2008:AAP

[Háj08] Alan Hájek. Arguments for–or against–probabilism? British Journal for the Philosophy of Science, 59(4):793–819, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/4/793.full.pdf+
html;   http://www.jstor.
org/stable/40072312.

Haldane:1951:TR

[Hal51] J. B. S. Haldane. Theories of Relativity. British Journal for the Philosophy of Science, II(5):73– 74, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/II/5/73.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685614.

Haldane:1952:MCP

[Hal52a] J. B. S. Haldane. The mechanical chess–player. British Journal for the Philosophy of Science, III (10):189–191, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/10/189.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685557.

Haldane:1952:R

[Hal52b] J. B. S. Haldane. Reviews. British Journal for the Philosophy of Science, III(9):103–105, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/9/103.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685485.

Haldane:1955:LBG

[Hal55a] J. B. S. Haldane. A logical basis for genetics? British Journal for the Philosophy of Science, VI(23):245–248, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/23/245.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685683.

Halsbury:1955:RUP

[Hal55b] Halsbury. Remarks upon Professor Bondi’s views. British Journal for the Philosophy of Science, VI(23):243–244, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/243.full.pdf+
html;   http://www.jstor.
org/stable/685681.

Halsey:1961:R

[Hal61] A. H. Halsey. Reviews. British Journal for the Philosophy of Science, XII(45):82– 83, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/82.full.pdf+
html;   http://www.jstor.
org/stable/685548.

Hall:1962:R

[Hal62] P. G. Hall. Reviews. British Journal for the Philosophy of Science, XII(48):342–344, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/48/342.full.
pdf+html;   http://www.
jstor.org/stable/685403.

Hall:1967:R

[Hal67] Roland Hall. Reviews. British Journal for the Philosophy of Science, 18(3):255–257, November 1967. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/255.full.pdf+
html;   http://www.jstor.
org/stable/686599.

Hall:1970:KCR

[Hal70] Richard J. Hall. Kuhn and the Copernican revolution. British Journal for the Philosophy of Science, 21(2):196–197, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/196.full.pdf+
html;   http://www.jstor.
org/stable/686700.

Hall:1972:R

[Hal72] A. Rupert Hall. Reviews. British Journal for the Philosophy of Science, 23(1):80–82, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
23/1/80.full.pdf+html;
http://www.jstor.org/
stable/686225.

Hall:1975:R

[Hal75a] A. Rupert Hall. Reviews. British Journal for the Philosophy of Science, 26(3):265–267, September 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/3/265.full.pdf+
html;   http://www.jstor.
org/stable/686284.

Hallett:1975:R

[Hal75b] M. F. Hallett. Reviews. British Journal for the Philosophy of Science, 26(2):170–174, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/2/170.full.pdf+
html;   http://www.jstor.
org/stable/686517.

Hallam:1978:R

[Hal78] A. Hallam. Reviews. British Journal for the Philosophy of Science, 29(4):392–393, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/4/392.full.pdf+
html;   http://www.jstor.
org/stable/687104.

Hall:1979:RU

[Hal79a] Marie Boas Hall. Review: [untitled]. British Journal for the Philosophy of Science, 30 (1):100, March 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687424.

Hall:1979:R

[Hal79b]  Marie  Boas

Hall. Reviews. British Journal for the Philosophy of Science, 30(1):100a–100, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
30/1/100-a.full.pdf+html.

Hallett:1979:TTMa

[Hal79c] Michael Hallett. Towards a theory of mathematical research programmes (I). British Journal for the Philosophy of Science, 30(1):1–25, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/1.full.pdf+html;
http://www.jstor.org/
stable/687414.

Hallett:1979:TTMb

[Hal79d] Michael Hallett. Towards a theory of mathematical research programmes (II). British Journal for the Philosophy of Science, 30(2):135–159, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/2/135.full.pdf+html;
http://www.jstor.org/
stable/686945.

Hallett:1981:RJL

[Hal81] Michael Hallett. Russell, Jourdain and ‘limitation of size’. British Journal for the Philosophy of Science, 32(4):381– 399, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/381.full.pdf+html;
http://www.jstor.org/
stable/687310.

Hall:1982:NR

[Hal82] A. Rupert Hall. Newton’s revolution. British Journal for the Philosophy of Science, 33(3): 305–315, September 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/305.full.pdf+html;
http://www.jstor.org/
stable/687229.

Hall:1986:RU

[Hal86] John A. Hall. Review: [untitled]. British Journal for the Philosophy of Science, 37 (1):126–130, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687005.

Hall:1987:R

[Hal87] J. A. Hall. Reviews. British Journal for the Philosophy of Science, 38(2):279–281, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/2/279.full.pdf+
html;   http://www.jstor.
org/stable/687057.

Hale:1998:R

[Hal98] Bob Hale. Review. British Journal for the Philosophy of Science, 49(1):161–167, March 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

49/1/161.full.pdf+html;
http://www.jstor.org/
stable/688148.

Hall:1999:BRB

[Hal99] N. Hall. Book review: The Quantum Challenge. G. Greenstein, A. G. Zajonc. British Journal for the Philosophy of Science, 50(2):313–315, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/313.full.pdf+html;
http://www.jstor.org/
stable/40072229.

Halvorson:2001:BRB

[Hal01] Hans Halvorson. Book review: A Philosopher’s Understanding of Quantum Mechanics: Possibilities and Impossibilities of a Modal Interpretation, Pieter Vermaas. British Journal for the Philosophy of Science, 52 (2):387–391, June 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/387.full.pdf+html;
http://www.jstor.org/
stable/3541903.

Hale:2015:BRW

[Hal15] Bob Hale. Book review: William Demopoulos, Logicism and Its Philosophical Legacy. British Journal for the Philosophy of Science, 66(2):459–463, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/2/459.

Hamblin:1965:R

[Ham65] C. L. Hamblin. Reviews. British Journal for the Philosophy of Science, XV(60):360–362, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/360.full.pdf+
html;   http://www.jstor.
org/stable/686552.

Hamlyn:1967:R

[Ham67] D. W. Hamlyn. Reviews. British Journal for the Philosophy of Science, 17(4):335–336, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/335.full.pdf+
html;   http://www.jstor.
org/stable/686779.

Hamlyn:1968:R

[Ham68] D. W. Hamlyn. Reviews. British Journal for the Philosophy of Science, 19(1):81–83, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/1/81.full.pdf+html;
http://www.jstor.org/
stable/686853.

Hamami:2023:PRP

[Ham23] Yacin Hamami. Proofs, reliable processes, and justification in mathematics. British Journal for the Philosophy of Science, 74 (4):1027–1045, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hanson:1961:DP

[Han61] Norwood Russell Hanson. Discovering the positron (I). British Journal for the Philosophy of Science, XII(47):194–214, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/47/194.full.
pdf+html;   http://www.
jstor.org/stable/685207.

Hanson:1962:DPI

[Han62a] Norwood Russell Hanson. Discovering the positron (II). British Journal for the Philosophy of Science, XII(48):299–313, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/48/299.full.
pdf+html;   http://www.
jstor.org/stable/685398.

Hanson:1962:RDH

[Han62b] Norwood Russell Hanson. A reply to Dr Hutten. British Journal for the Philosophy of Science, XIII(49):55, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/49/55.full.pdf+html;

http://www.jstor.org/
stable/685228.

Hanson:1971:MGT

[Han71] William H. Hanson. Mechanism and Gödel’s theorems. British Journal for the Philosophy of Science, 22(1):9–16, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/1/9.full.pdf+
html;   http://www.jstor.
org/stable/686639.

Hands:1992:MEP

[Han92] D. Wade Hands. Metaphysics, economics and progress: a comment on Glass and Johnson. British Journal for the Philosophy of Science, 43(2):241–244, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/241.full.pdf+
html;   http://www.jstor.
org/stable/687875.

Hanna:2004:CLW

[Han04] Joseph F. Hanna. Contra Ladyman: What really is right with constructive empiricism. British Journal for the Philosophy of Science, 55(4):767– 777, December 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/767.full.pdf+html;

http://www.jstor.org/
stable/3541629. See [Lad04].

Hanoch:2005:OTF

[Han05] Yaniv Hanoch. One theory to fit them all: the search hypothesis of emotion revisited. British Journal for the Philosophy of Science, 56(1):135–145, March 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/1/135.full.pdf+
html;   http://www.jstor.
org/stable/3541853.

Harrah:1959:GST

[Har59a] David Harrah. Gallie and the scientific tradition. British Journal for the Philosophy of Science, X(39):234–239, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/234.full.pdf+html;
http://www.jstor.org/
stable/686058.

Harre:1959:DNP

[Har59b] R. Harré. Discussions: Notes on P. K. Feyerabend’s criticism of positivism. British Journal for the Philosophy of Science, X(37):43–48, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/43.full.pdf+html;
http://www.jstor.org/
stable/685774.

Harrod:1960:DNA

[Har60] Roy Harrod. Discussions: New argument for induction: Reply to Professor Popper. British Journal for the Philosophy of Science, X(40):309–312, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/40/309.full.pdf+
html;   http://www.jstor.
org/stable/685917.

Harre:1961:R

[Har61a] R. Harré. Reviews. British Journal for the Philosophy of Science, XI(44):348–349, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/44/348.full.pdf+
html;   http://www.jstor.
org/stable/685142.

Harris:1961:R

[Har61b] F. T. C. Harris. Reviews. British Journal for the Philosophy of Science, XII(46):172–173, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/172.full.
pdf+html;   http://www.
jstor.org/stable/685501.

Harre:1962:PAC

[Har62] R. Harré. Philosophical aspects of cosmology. British Journal for the Philosophy of Science, XIII(50):104–119, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/50/104.full.pdf+html;
 http://www.jstor.org/
stable/685964.

Harre:1963:R

[Har63] R. Harré. Reviews. British Journal for the Philosophy of Science, XIII(52):337–338, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/337.full.
pdf+html;   http://www.
jstor.org/stable/685339.

Harre:1967:R

[Har67] R. Harré. Reviews. British Journal for the Philosophy of Science, 17(4):346, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/346.full.pdf+html;
http://www.jstor.org/
stable/686784.

Harre:1968:R

[Har68a] R. Harré. Reviews. British Journal for the Philosophy of Science, 18(4):352–353, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/4/352.full.pdf+
html;   http://www.jstor.
org/stable/686371.

Harris:1968:SF

[Har68b] Errol E. Harris. Simultaneity and the future. British Journal for the Philosophy of Science, 19(3):254–256, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/3/254.full.pdf+html;
http://www.jstor.org/
stable/686245.

Harre:1970:P

[Har70] R. Harré. Powers. British Journal for the Philosophy of Science, 21(1):81–101, February 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/1/81.full.pdf+
html;   http://www.jstor.
org/stable/686473.

Harrison:1971:R

[Har71]  Andrew Harrison.  Reviews.

British Journal for the Philosophy of Science, 22(1):59–60, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/1/59.full.pdf+
html;   http://www.jstor.
org/stable/686647.

Harris:1972:EP

[Har72] Errol E. Harris. Epicyclic Popperism. British Journal for the Philosophy of Science, 23 (1):55–67, February 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/1/55.full.pdf+html;
http://www.jstor.org/
stable/686222.

Harris:1974:PDV

[Har74] John H. Harris. Popper’s definitions of ‘verisimilitude’. British Journal for the Philosophy of Science, 25(2):160–166, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/2/160.full.pdf+
html;   http://www.jstor.
org/stable/686820.

Hart:1984:R

[Har84] W. D. Hart. Reviews. British Journal for the Philosophy of Science, 35(4):411–414, December 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/4/411.full.pdf+
html;   http://www.jstor.
org/stable/687347.

Harre:1991:O

[Har91] Rom Harré. Obituary. British Journal for the Philosophy of Science, 42(1):145–146, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/1/145.full.pdf+

html;   http://www.jstor.
org/stable/687612.

Hardin:1993:VBN

[Har93] Clyde L. Hardin. Van Brakel and the not-so-naked emperor. British Journal for the Philosophy of Science, 44(1):137–150, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/137.full.pdf+
html;   http://www.jstor.
org/stable/687853.

Hardcastle:1994:IO

[Har94] Valerie Gray Hardcastle. The image of observables. British Journal for the Philosophy of Science, 45(2):585–597, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/585.full.pdf+html;
http://www.jstor.org/
stable/687684.

Hardin:1996:R

[Har96a] C. L. Hardin. Review. British Journal for the Philosophy of Science, 47(2):339–343, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/339.full.pdf+
html;   http://www.jstor.
org/stable/687957.

Harre:1996:R

[Har96b]  Rom Harré.  Review.

British Journal for the Philosophy of Science, 47(1):141–145, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/141.full.pdf+
html;   http://www.jstor.
org/stable/1215834.

Harker:2006:APC

[Har06] David Harker. Accommodation and prediction: the case of the persistent head. British Journal for the Philosophy of Science, 57(2):309–321, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/309.full.pdf+html;
http://www.jstor.org/
stable/3873469.

Harker:2008:PP

[Har08] David Harker. On the predilections for predictions. British Journal for the Philosophy of Science, 59(3):429–453, September 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/3/429.full.pdf+
html;   http://www.jstor.
org/stable/40072294.

Harker:2011:BRE

[Har11] David Harker. Book review: Eric Christian Barnes: The Paradox of Predictivism. British Journal for the Philosophy of Science, 62 (1):219–223, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/219.full.pdf+html.

Harker:2013:HST

[Har13] David Harker. How to split a theory: Defending selective realism and convergence without proximity. British Journal for the Philosophy of Science, 64(1):79–106, March 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/79.full.pdf+html.

Harding:2022:EQM

[Har22] Jacqueline Harding. Everettian quantum mechanics and the metaphysics of modality. British Journal for the Philosophy of Science, 72(4): 939–964, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hauptli:1994:RUE

[Hau94] Bruce W. Hauptli. Rescher’s unsuccessful evolutionary argument. British Journal for the Philosophy of Science, 45(1): 295–301, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/295.full.pdf+html;
http://www.jstor.org/
stable/687974.

Hausman:1998:CA

[Hau98] Daniel M. Hausman. Causal asymmetries. Cambridge studies in probability, induction, and decision theory. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-62289-1 (hardcover). xv + 300 pp. LCCN BD541 .H38 1998. URL http://www.

                                        loc.gov/catdir/description/
                                        cam028/97042437.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam025/97042437.html.

Hausman:1999:RAM

[Hau99] D. M. Hausman. Review article. the mathematical theory of causation. British Journal for the Philosophy of Science, 50 (1):151–162, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/1/151.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072213.

Hausman:2011:OPB

[Hau11] Daniel M. Hausman. Is an overdose of paracetamol bad for one’s health? British Journal for the Philosophy of Science, 62(3):657–668, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/3/657.full.pdf+html.

Haufe:2013:NCB

[Hau13] Chris Haufe. From necessary chances to biological laws. British Journal for the Philosophy of Science, 64(2):279–295, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/279.full.pdf+html.

Havstad:2018:MCK

[Hav18]  Joyce C. Havstad.  Messy

chemical kinds. British Journal for the Philosophy of Science, 69(3):719–743, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/719/3603393.

Hawley:1997:R

[Haw97]  Katherine Hawley.  Review.

British Journal for the Philosophy of Science, 48(2):297–299, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/297.full.pdf+
html;   http://www.jstor.
org/stable/687754.

Hawley:1999:BRB

[Haw99] K. Hawley. Book review: The possibility of metaphysics; substance, identity and time. E. J. Lowe. British Journal for the Philosophy of Science, 50(3): 478–482, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        50/3/478.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072245.

Hayek:1955:DE

[Hay55] F. A. Hayek. Degrees of explanation. British Journal for the Philosophy of Science, VI (23):209–225, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/23/209.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685677.

Heathcote:1953:WWP

[Hea53]  A. W. Heathcote.  William

Whewell’s Philosophy of Science. British Journal for the Philosophy of Science, IV(16): 302–314, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/16/302.full.pdf+html.

Heathcote:1954:WWP

[Hea54]  A. W. Heathcote.  William

Whewell’s philosophy of science. British Journal for the Philosophy of Science, IV(16): 302–314, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686079.

Heal:1986:RU

[Hea86] Jane Heal. Book review: Strawson, P. F. [1985]: Scepticism and Naturalism: Some Varieties. Methuen. x + 98 pp. £10.95. British Journal for the Philosophy of Science, 37(4): 523–525, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687380.

Heathcote:1988:ZGT

[Hea88] Adrian Heathcote. Zeeman– Göbel topologies. British Journal for the Philosophy of Science, 39(2):247–261, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/2/247.full.pdf+html;
http://www.jstor.org/
stable/687270.

Heal:1999:BRB

[Hea99] J. Heal. Book review: Language and Thought: Interdisciplinary Themes. P. Carruthers, J. Boucher [eds]. British Journal for the Philosophy of Science, 50(2):305–308, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/305.full.pdf+html;
http://www.jstor.org/
stable/40072227.

Healey:2002:AT

[Hea02] Richard Healey. The arguments of time. British Journal for the Philosophy of Science, 53(3):459–463, September 2002.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        53/3/459.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541891.

Healey:2009:PS

[Hea09] Richard Healey. Perfect symmetries. British Journal for the Philosophy of Science, 60(4): 697–720, December 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/4/697.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592031.

Healey:2012:QTP

[Hea12] Richard Healey. Quantum theory: a pragmatist approach. British Journal for the Philosophy of Science, 63(4):729– 771, December 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/4/729.full.pdf+html.

Healey:2015:HQT

[Hea15] Richard A. Healey. How quantum theory helps us explain. British Journal for the Philosophy of Science, 66(1):1–43, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/66/1/1.

Heisenberg:1958:PPR

[Hei58] Werner Heisenberg. Physics and philosophy: the revolution in modern science, volume 19 of World perspectives. Harper, New York, NY, USA, 1958. 206 pp. LCCN QC6.H473; QC6 .H473.

Heisenberg:1962:PPR

[Hei62] Werner Heisenberg. Physics and philosophy; the revolution in modern science, volume TB 549 of Harper torchbooks: The science library. Harper & Row, New York, NY, USA, 1962. ISBN 0-06-130549-9. 213 pp. LCCN QC6.H473; QC6 .H473 1962. Introduction by F. S. C. Northrop.

Heilig:1978:CFB

[Hei78] Klaus Heilig. Carnap and de Finetti on Bets and the Probability of Singular Events: the Dutch Book Argument Reconsidered. British Journal for the Philosophy of Science, 29(4): 325–346, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/325.full.pdf+html;
http://www.jstor.org/
stable/687096.

Heidelberger:1983:R

[Hei83] Michael Heidelberger. Reviews.

British Journal for the Philosophy of Science, 34(4):406– 410, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
34/4/406.full.pdf+html;
http://www.jstor.org/
stable/687536.

Heil:1999:BRB

[Hei99] J. Heil. Book review: Mind in a Physical World: an Assay on the Mind-Body Problem and Mental Causation. J. Kim. British Journal for the Philosophy of Science, 50(4):769–773, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/4/769.full.pdf+
html;   http://www.jstor.
org/stable/40072267.

Helman:1987:RAA

[Hel87] David H. Helman. Realism and antirealism in artificial intelligence. British Journal for the Philosophy of Science, 38 (1):19–26, March 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/19.full.pdf+html;
http://www.jstor.org/
stable/687076.

Hellman:1998:MCS

[Hel98] Geoffrey Hellman. Mathematical constructivism in spacetime. British Journal for the Philosophy of Science, 49(3):425– 450, September 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
49/3/425.full.pdf+html;
http://www.jstor.org/
stable/688083.

Helgeson:2018:MDR

[Hel18] Casey Helgeson. Modus Darwin reconsidered. British Journal for the Philosophy of Science, 69(1): 193–213, March 1, 2018. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/193/2669783.

Hempel:1967:WSN

[Hem67] Carl G. Hempel. The white shoe: No red herring. British Journal for the Philosophy of Science, 18(3):239–240, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/239.full.pdf+
html;   http://www.jstor.
org/stable/686596.

Henry:1966:R

[Hen66]  Desmond Paul Henry.  Re-

views. British Journal for the Philosophy of Science, XVI(64): 349–351, February 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/64/349.full.pdf+html;
http://www.jstor.org/
stable/686841.

Hendry:1981:PP

[Hen81] John Hendry. Pauli as philosopher. British Journal for the Philosophy of Science, 32(3):277– 282, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/3/277.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687591.

Hendry:1982:R

[Hen82] John Hendry. Reviews. British Journal for the Philosophy of Science, 33(3):329–331, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/3/329.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687233.

Hendry:1984:CQM

[Hen84] John Hendry. The creation of quantum mechanics and the Bohr–Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Hendry:1997:R

[Hen97] Robin Findlay Hendry. Review. British Journal for the Philosophy of Science, 48(1):109–112, March 1997. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/109.full.pdf+
html;   http://www.jstor.
org/stable/687721.

Henderson:2003:NER

[Hen03]  Leah Henderson.  The von

Neumann entropy: a reply to Shenker. British Journal for the Philosophy of Science, 54 (2):291–296, June 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/2/291.full.pdf+html;
http://www.jstor.org/
stable/3541968.

Henderson:2014:BIB

[Hen14] Leah Henderson. Bayesianism and inference to the best explanation. British Journal for the Philosophy of Science, 65(4): 687–715, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/687.

Herzberg:2007:ILP

[Her07] Frederik Herzberg. Internal laws of probability, generalized likelihoods and Lewis’ infinitesimal chances — a response to Adam Elga. British Journal for the Philosophy of Science, 58(1):25–43, March 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/58/1/25.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/4489099.

Hersch:2023:WBC

[Her23] Gil Hersch. Well-being coherentism. British Journal for the Philosophy of Science, 73(4):1045– 1065, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hesse:1952:ODA

[Hes52] Mary Hesse. Operational definition and analogy in physical theories. British Journal for the Philosophy of Science, II(8): 281–294, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/8/281.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686017.

Hesse:1953:MP

[Hes53a] Mary B. Hesse. Models in physics. British Journal for the Philosophy of Science, IV (15):198–214, November 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/15/198.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685897.

Hesse:1953:Ra

[Hes53b]  Mary B. Hesse.  Reviews.

British Journal for the Philosophy of Science, IV(15):254– 256, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/254.full.pdf+html;
http://www.jstor.org/
stable/685906.

Hesse:1953:Rb

[Hes53c]  Mary B. Hesse.  Reviews.

British Journal for the Philosophy of Science, IV(15):262– 264, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/262.full.pdf+html;
http://www.jstor.org/
stable/685910.

Hesse:1956:R

[Hes56] Mary B. Hesse. Reviews. British Journal for the Philosophy of Science, VII(27):269, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/269.full.
pdf+html;   http://www.
jstor.org/stable/685889.

Hessb:1957:R

[Hes57] Mary B. Hessb. Reviews. British Journal for the Philosophy of Science, VII(28):359–360, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/VII/28/359.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685122.

Hesse:1958:NTD

[Hes58a] Mary B. Hesse. A note on ‘theories, dictionaries, and observation’. British Journal for the Philosophy of Science, IX(34): 128–129, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/34/128.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685148.

Hesse:1958:R

[Hes58b]  Mary B. Hesse.  Reviews.

British Journal for the Philosophy of Science, IX(35):248– 249, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/35/248.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685658.

Hesse:1958:TDO

[Hes58c]  Mary B. Hesse.  Theories,

dictionaries, and observation. British Journal for the Philosophy of Science, IX(33):12–28, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/33/12.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686097.

Hesse:1959:Ra

[Hes59a] Mary B. Hesse. Reviews. British Journal for the Philosophy of Science, X(37):69–73, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/69.full.pdf+html;
http://www.jstor.org/
stable/685780.

Hesse:1959:Rb

[Hes59b] Mary B. Hesse. Reviews. British Journal for the Philosophy of Science, X(37):73–78, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/73.full.pdf+html;
http://www.jstor.org/
stable/685781.

Hesse:1960:Ra

[Hes60a] M. B. Hesse. Reviews. British Journal for the Philosophy of Science, XI(42):164–166, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/164.full.pdf+
html;   http://www.jstor.
org/stable/685597.

Hesse:1960:GH

[Hes60b] Mary B. Hesse. Gilbert and the historians (I). British Journal for the Philosophy of Science, XI(41):1–10, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/41/1.full.pdf+html;
http://www.jstor.org/
stable/685815.

Hesse:1960:GHI

[Hes60c] Mary B. Hesse. Gilbert and the historians (II). British Journal for the Philosophy of Science, XI (42):130–142, August 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/42/130.full.pdf+html;
http://www.jstor.org/
stable/685585.

Hesse:1960:Rb

[Hes60d]  Mary B. Hesse.  Reviews.

British Journal for the Philosophy of Science, XI(43):252– 255, November 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/252.full.pdf+html;
http://www.jstor.org/
stable/685315.

Hesse:1961:R

[Hes61] Mary B. Hesse. Reviews. British Journal for the Philosophy of Science, XII(46):173–176, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/173.full.
pdf+html;   http://www.
jstor.org/stable/685502.

Hesse:1962:WTP

[Hes62] Mary Hesse. On what there is in physics. British Journal for the Philosophy of Science, XIII(51):234–244, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/234.full.
pdf+html;   http://www.
jstor.org/stable/685841.

Hesse:1967:Ra

[Hes67a] Mary Hesse. Reviews. British Journal for the Philosophy of Science, 18(1):67–70, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/67.full.pdf+html;
http://www.jstor.org/
stable/686892.

Hesse:1967:Rb

[Hes67b] Mary Hesse. Reviews. British Journal for the Philosophy of Science, 18(2):160–161, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/160.full.pdf+
html;   http://www.jstor.
org/stable/686586.

Hesse:1968:Ra

[Hes68a] Mary Hesse. Reviews. British Journal for the Philosophy of Science, 18(4):330–335, February 1968. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
18/4/330.full.pdf+html.

Hesse:1968:Rb

[Hes68b] Mary Hesse. Reviews. British Journal for the Philosophy of Science, 19(2):176–177, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/176.full.pdf+
html;   http://www.jstor.
org/stable/686800.

Hesse:1969:EP

[Hes69a] Mary Hesse. The encyclopedia of philosophy. British Journal for the Philosophy of Science, 20 (3):263–269, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/3/263.full.pdf+html;
http://www.jstor.org/
stable/686422.

Hesse:1969:RG

[Hes69b] Mary Hesse. Ramifications of ‘Grue’. British Journal for the Philosophy of Science, 20(1):13– 25, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/13.full.pdf+
html;   http://www.jstor.
org/stable/686715.

Hesse:1970:R

[Hes70] Mary Hesse. Reviews. British Journal for the Philosophy of Science, 21(3):303–304, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/303.full.pdf+
html;   http://www.jstor.
org/stable/686207.

Heyd:1996:R

[Hey96] David Heyd. Review. British Journal for the Philosophy of Science, 47(1):129–132, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/129.full.pdf+
html;   http://www.jstor.
org/stable/1215831.

Hey:2016:HMH

[Hey16a] Spencer Phillips Hey. Heuristics and meta-heuristics in scientific judgement. British Journal for the Philosophy of Science, 67(2):471–495, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/471.

Heyes:2016:BRT

[Hey16b] Cecilia Heyes. Book review:

Tim Lewens, Cultural Evolution. British Journal for the Philosophy of Science, 67(4):1189– 1193, December 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1189.

Howson:1985:BAE

[HF85]  Colin  Howson

and Allan Franklin. A Bayesian analysis of excess content and the localisation of support. British Journal for the Philosophy of Science, 36(4):425–431, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/4/425.full.pdf+
html;   http://www.jstor.
org/stable/687401.

Howson:1994:BCP

[HF94]  Colin  How-

son and Allan Franklin. Bayesian conditionalization and probability kinematics. British Journal for the Philosophy of Science, 45(2):451–466, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/451.full.pdf+html;
http://www.jstor.org/
stable/687675.

Humphreys:1996:GL

[HF96] Paul Humphreys and David Freedman. The grand leap. British Journal for the Philosophy of Science, 47(1):113–123, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/113.full.pdf+
html;   http://www.jstor.
org/stable/1215829. See re-

marks [KW97, SGS97].

Harris:1963:ACB

[HH63] D. G. Harris and F. T. C. Harris.

Annual conference of the British Society for the Philosophy of Science. British Journal for the Philosophy of Science, XIV(53):76– 77, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/76.full.pdf+
html;   http://www.jstor.
org/stable/685361.

Harris:1964:EAC

[HH64] F. T. C. Harris and D. G. Harris. Eighth annual conference of the British Society for the Philosophy of Science. British Journal for the Philosophy of Science, XV(57):83–85, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/57/83.full.pdf+html.

Hoyningen-Huene:1992:IBP

[HH92] Paul Hoyningen-Huene. The interrelations between the philosophy, history and sociology of science in Thomas Kuhn‘s Theory of Scientific Development. British Journal for the Philosophy of Science, 43(4):487–501, October 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/4/487.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687837.

Holt:1993:RND

[HH93] D. Lynn Holt and R. Glynn Holt. Regularity in nonlinear dynamical systems. British Journal for the Philosophy of Science, 44(4):711–727, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        44/4/711.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688040.

Halpern:2015:GCD

[HH15] Joseph Y. Halpern and Christopher Hitchcock. Graded causation and defaults. British Journal for the Philosophy of Science, 66(2):413–457, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/2/413.

Hartenstein:2021:WFD

[HH21] Vera Hartenstein and Mario Hubert. When fields are not degrees of freedom. British Journal for the Philosophy of Science, 72(1): 245–275, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hoyningen-Huene:2002:RSS

[HHS02]  Paul  Hoyningen-

Huene and Howard Sankey. The road since ‘structure’: Philosophical essays, 1970–1993, with an autobiographical interview. British Journal for the Philosophy of Science, 53(1):137–142, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/1/137.full.pdf+
html;   http://www.jstor.
org/stable/3541645.

Hick:1953:MCP

[Hic53] W. E. Hick. The mechanical chess-player. British Journal for the Philosophy of Science, III(12):353–355, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/353.full.pdf+html;
http://www.jstor.org/
stable/685453.

Hicks:2018:DH

[Hic18] Michael Townsen Hicks. Dynamic humeanism. British Journal for the Philosophy of Science, 69(4):983–1007, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/983/3859152.

Hiddleston:2005:CP

[Hid05] Eric Hiddleston. Causal powers. British Journal for the Philosophy of Science, 56(1):27–59, March 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/56/1/27.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541849.

Hinde:1956:EMC

[Hin56] R. A. Hinde. Ethological models and the concept of ‘drive’. British Journal for the Philosophy of Science, VI(24):321– 331, February 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/24/321.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685412.

Hinckfuss:1988:ARS

[Hin88] Ian Hinckfuss. Absolutism and relationism in space and time: a false dichotomy. British Journal for the Philosophy of Science, 39(2):183–192, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        39/2/183.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687265.

Hinckfuss:1995:R

[Hin95] Ian Hinckfuss. Review. British Journal for the Philosophy of Science, 46(3):425–435, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/3/425.full.pdf+

html;   http://www.jstor.
org/stable/687667.

Hirst:1962:R

[Hir62] R. J. Hirst. Reviews. British Journal for the Philosophy of Science, XIII(50):186, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/186.full.
pdf+html;   http://www.
jstor.org/stable/685981.

Hirst:1963:R

[Hir63] R. J. Hirst. Reviews. British Journal for the Philosophy of Science, XIV(53):69– 71, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/69.full.pdf+
html;   http://www.jstor.
org/stable/685356.

Hirst:1965:R

[Hir65] R. J. Hirst. Reviews. British Journal for the Philosophy of Science, XV(60):358–360, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/358.full.pdf+
html;   http://www.jstor.
org/stable/686551.

Hirschmann:1966:R

[Hir66] David Hirschmann. Reviews.

British Journal for the Philosophy of Science, 17(2):155–159, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/2/155.full.pdf+
html;   http://www.jstor.
org/stable/686350.

Hirschmann:1967:R

[Hir67] David Hirschmann. Reviews. British Journal for the Philosophy of Science, 18(1):80–81, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/80.full.pdf+
html;   http://www.jstor.
org/stable/686899.

Hirschmann:1978:R

[Hir78] David Hirschmann. Reviews. British Journal for the Philosophy of Science, 29(2):195–197, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/195.full.pdf+
html;   http://www.jstor.
org/stable/687129.

Hitchcock:1999:CED

[Hit99] C. Hitchcock. Contrastive explanation and the demons of determinism. British Journal for the Philosophy of Science, 50(4):585–612, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/585.full.pdf+html;

http://www.jstor.org/
stable/40072258.

Hitchock:2000:BRB

[Hit00]  C. Hitchock.  Book review:

Causal Asymmetries. D. M. Hausman. British Journal for the Philosophy of Science, 51 (1):175–179, March 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/1/175.full.pdf+html;
http://www.jstor.org/
stable/3541756.

Hitchcock:2003:HB

[Hit03]  Christopher Hitchcock.  Of

Humean bondage. British Journal for the Philosophy of Science, 54(1):1–25, March 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/1/1.full.pdf+html;
http://www.jstor.org/
stable/3541877.

Harper:1968:DJC

[HK68] William L. Harper and Henry E.

Kyburg. Discussions: the Jones Case. British Journal for the Philosophy of Science, 19(3): 247–251, November 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/3/247.full.pdf+html;
http://www.jstor.org/
stable/686243.

Hohwy:2008:BRN

[HK08]  Jakob  Ho-

hwy and Jesper Kallestrup, editors. Being reduced: new essays on reduction, explanation, and causation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-921153-1 (hardcover). viii + 312 pp. LCCN BD418.3.

Huttemann:2011:ARE

[HL11] Andreas Hüttemann and Alan C. Love. Aspects of reductive explanation in biological science: Intrinsicality, fundamentality, and temporality. British Journal for the Philosophy of Science, 62(3):519–549, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/3/519.full.pdf+html.

Halperin:2024:WIA

[HL24] Topaz Halperin and Arnon Levy. What, if anything, is biological altruism? British Journal for the Philosophy of Science, 75(2): 465–486, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hawthorne:2017:PPI

[HLWW17] James Hawthorne, Jürgen Landes, Christian Wallmann, and Jon Williamson. The principal principle implies the principle of indifference. British Journal for the Philosophy of Science, 68(1): 68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/1/123/2995353/
The-Principal-Principle-
Implies-the-Principle-of.

Hendry:1999:BRB

[HM99] R. F. Hendry and D. J. Mossley. Book review: Realism rescued: how scientific progress is possible. Jerrold L Aronson, R Harré, Eileen Cornell Way. British Journal for the Philosophy of Science, 50(1):175–179, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/1/175.full.pdf+
html;   http://www.jstor.
org/stable/40072216.

Hendry:2007:PRC

[HN07] Robin Findlay Hendry and Paul Needham. Le Poidevin on the reduction of chemistry. British Journal for the Philosophy of Science, 58(2):339–353, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/2/339.full.pdf+
html;   http://www.jstor.
org/stable/30115231.

Huggett:2014:WDQ

[HN14] Nick Huggett and Josh Norton.

Weak discernibility for quanta, the right way. British Journal for the Philosophy of Science, 65(1):39–58, January 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/1/39.full.pdf+html.

Howson:1979:MCP

[HO79]  Colin  Howson  and

Graham Oddie. Miller’s so-called paradox of information. British Journal for the Philosophy of Science, 30(3):253–261, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/3/253.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687139.

Hobbs:1994:LDP

[Hob94] Jesse Hobbs. A limited defense of the pessimistic induction. British Journal for the Philosophy of Science, 45(1): 171–191, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/1/171.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687966.

Hochstein:2017:WDF

[Hoc17] Eric Hochstein. When does ‘folk psychology’ count as folk psychological? British Journal for the Philosophy of Science, 68(4):1125–1147, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/1125/2669637.

Hodges:1978:R

[Hod78] Wilfrid Hodges. Reviews. British Journal for the Philosophy of Science, 29(2):207–208, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/207.full.pdf+
html;   http://www.jstor.
org/stable/687133.

Hodge:1980:RU

[Hod80a] M. J. S. Hodge. Review: [untitled]. British Journal for the Philosophy of Science, 31 (1):85–89, March 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687255.

Hodge:1980:R

[Hod80b] M. J. S. Hodge. Reviews. British Journal for the Philosophy of Science, 31(1):85–88, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/1/85.full.pdf+html.

Hodge:1997:R

[Hod97] Jon Hodge. Review. British Journal for the Philosophy of Science, 48(3):435–438, September 1997. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/3/435.full.pdf+
html;   http://www.jstor.
org/stable/688072.

Hoering:1969:ICP

[Hoe69] Walter Hoering. Indeterminism in classical physics. British Journal for the Philosophy of Science, 20(3):247–255, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/3/247.full.pdf+
html;   http://www.jstor.
org/stable/686418.

Hoefer:1998:AVR

[Hoe98] Carl Hoefer. Absolute versus relational spacetime: for better or worse, the debate goes on. British Journal for the Philosophy of Science, 49(3):451– 467, September 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/3/451.full.pdf+html;
http://www.jstor.org/
stable/688084.

Hoerl:2000:BRJ

[Hoe00] Christoph Hoerl. Book review: John Sutton — Philosophy and Memory Traces: Descartes to Connectionism. British Journal for the Philosophy of Science, 51(4):923–926, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/923.full.pdf+html;
http://www.jstor.org/
stable/3541738.

Hofweber:2001:BRB

[Hof01] Thomas Hofweber. Book review: Philosophy of Mathematics: an Introduction to the World of Proofs and Pictures, James Robert Brown. British Journal for the Philosophy of Science, 52(2):413–416, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/413.full.pdf+html;
http://www.jstor.org/
stable/3541908.

Hogendijk:1986:ADA

[Hog86] Jan P. Hogendijk. Book review: Al-Daffa, Ali A. and Stroyls, John J. [1984]: Studies in the Exact Sciences in Medieval Islam. University of Petroleum and Minerals (Dhahran, Saudi Arabia) and John Wiley and Sons. x + 243 pp. (ISBN 0-471-90320-5). British Journal for the Philosophy of Science, 37(4):516–520, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687378.

Hogarth:2004:DAU

[Hog04] Mark Hogarth. Deciding arithmetic using sad computers. British Journal for the Philosophy of Science, 55(4):681–691, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/4/681.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541622.

Holdcroft:1979:R

[Hol79]  David Holdcroft.  Reviews.

British Journal for the Philosophy of Science, 30(4):411– 418, December 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/4/411.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687549.

Holcomb:1986:LVC

[Hol86] Harmon R. Holcomb, III. Latency versus complementarity: Margenau and Bohr on quantum mechanics. British Journal for the Philosophy of Science, 37(2):193–206, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686977.

Holtzman:1988:NSC

[Hol88] Jack M. Holtzman. A note on Schrödinger’s Cat and the unexpected hanging paradox. British Journal for the Philosophy of Science, 39(3):397–401, September 1988. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/397.full.pdf+
html;   http://www.jstor.
org/stable/687217.

Holgate:1994:MNR

[Hol94] Philip Holgate. Mathematical notes on Ross’s Paradox. British Journal for the Philosophy of Science, 45(1):302–304, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/302.full.pdf+
html;   http://www.jstor.
org/stable/687975.

Holder:1998:HMB

[Hol98] Rodney D. Holder. Hume on miracles: Bayesian interpretation, multiple testimony, and the existence of God. British Journal for the Philosophy of Science, 49(1):49–65, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/1/49.full.pdf+html;
http://www.jstor.org/
stable/688143.

Hon:1991:CNJ

[Hon91] Giora Hon. A critical note on J.

S. Mill’s classification of fallacies. British Journal for the Philosophy of Science, 42(2):263–268, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/42/2/263.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688058.

Hooker:1971:RDS

[Hoo71] Clifford A. Hooker. The relational doctrines of space and time. British Journal for the Philosophy of Science, 22(2):97–130, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/2/97.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686482.

Hookway:1984:MME

[Hoo84] Christopher Hookway, editor. Minds, machines, and evolution: philosophical studies. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26547-9 (hardcover). xi + 177 pp. LCCN B29 .M535

                                        1984. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/84011321.html.

Hookway:1985:P

[Hoo85] Christopher Hookway. Peirce. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1985. ISBN 0-7100-9715-8 (US). xii + 301 pp. LCCN B945.P44 H66 1985. US$50.00.

Hooker:1986:BRS

[Hoo86] C. A. Hooker. Book review: Stich, Stephen P. [1983]: From Folk Psychology to Cognitive Science. MIT Press (a Bradford Book). xii + 266 pp. ISBN 0-262-19215-2. British Journal for the Philosophy of Science, 37 (2):238–242, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686986.

Hooker:1991:PPI

[Hoo91] C. A. Hooker. Projection, physical intelligibility, objectivity and completeness: the divergent ideals of Bohr and Einstein. British Journal for the Philosophy of Science, 42(4):491–511, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/4/491.full.pdf+
html;   http://www.jstor.
org/stable/687705.

Hoover:1993:CTO

[Hoo93] Evin D. Hoover. Causality and temporal order in macroeconomics or why even economists don’t know how to get causes from probabilities. British Journal for the Philosophy of Science, 44(4):693–710, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/4/693.full.pdf+html;
http://www.jstor.org/
stable/688039.

Hooker:1996:R

[Hoo96] C. A. Hooker. Review. British Journal for the Philosophy of Science, 47(4):647–653, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/4/647.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687934.

Hookway:1999:P

[Hoo99] Christopher Hookway. Peirce. The arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1999. ISBN 0-415-20382-1, 0-415-20392-9 (set). xii + 301 pp. LCCN B945.P44 H66

                                        1999. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0649/00269382-d.html.

Hookway:2000:TRP

[Hoo00] Christopher Hookway. Truth, rationality, and pragmatism: themes from Peirce. Clarendon Press, New York, NY, USA, 2000. ISBN 0-19-823836-3. vi + 313 pp. LCCN B945.P44 H67

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/99089742-d.html.

Hoover:2003:NTS

[Hoo03] Kevin D. Hoover. Nonstationary time series, cointegration, and the principle of the common cause. British Journal for the Philosophy of Science, 54(4):527–551, December 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/

54/4/527.full.pdf+html;
http://www.jstor.org/
stable/3541677.

Hooker:2004:ARE

[Hoo04] C. A. Hooker. Asymptotics, reduction and emergence. British Journal for the Philosophy of Science, 55(3):435–479, September 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/3/435.full.pdf+
html;   http://www.jstor.
org/stable/3541670.

Horowitz:1965:DSK

[Hor65] Irving Louis Horowitz. In defence of the sociology of knowledge. British Journal for the Philosophy of Science, XVI(63): 239–241, November 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/63/239.full.pdf+html;
http://www.jstor.org/
stable/686305.

Horwich:1975:GMS

[Hor75] Paul Horwich. Grünbaum on the metric of space and time. British Journal for the Philosophy of Science, 26(3):199–211, September 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/3/199.full.pdf+
html;   http://www.jstor.
org/stable/686277.

Horton:1978:APN

[Hor78a] Deryck Horton. Accuracy of prediction: a note on David Miller’s problem. British Journal for the Philosophy of Science, 29(2):179–183, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/2/179.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687125.

Horwich:1978:PCN

[Hor78b] Paul Horwich. A peculiar consequence of Nicod’s criterion. British Journal for the Philosophy of Science, 29(3):262– 263, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/3/262.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687016.

Horwood:1979:R

[Hor79] Rob Horwood. Reviews. British Journal for the Philosophy of Science, 30(1):86–90, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/1/86.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687419.

Horwich:1993:R

[Hor93] Paul Horwich. Review. British Journal for the Philosophy of Science, 44(3):579–583, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/3/579.full.pdf+
html;   http://www.jstor.
org/stable/688028.

Horgan:1994:IBN

[Hor94] James Horgan. Icon and Bild:

a note on the analogical structure of models — the role of models in experiment and theory. British Journal for the Philosophy of Science, 45(2):599–604, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/599.full.pdf+
html;   http://www.jstor.
org/stable/687685.

Horgan:1996:KMB

[Hor96] Terence Horgan. Kim on the mind–body problem. British Journal for the Philosophy of Science, 47(4):579–607, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/579.full.pdf+
html;   http://www.jstor.
org/stable/687926.

Hossack:2000:PC

[Hos00] K. Hossack. Plurals and complexes. British Journal for the Philosophy of Science, 51(3): 411–443, September 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/3/411.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541772.

Hotopf:1963:R

[Hot63] W. H. N. Hotopf. Reviews. British Journal for the Philosophy of Science, XIII(52):342– 343, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/52/342.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685342.

Howard:1958:R

[How58] I. P. Howard. Reviews. British Journal for the Philosophy of Science, IX(34):175–176, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/34/175.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685162.

Howson:1969:R

[How69] Colin Howson. Reviews. British Journal for the Philosophy of Science, 20(4):359–361, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/4/359.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686265.

Howson:1971:R

[How71] C. Howson. Reviews. British Journal for the Philosophy of Science, 22(1):72–73, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/72.full.pdf+html;
http://www.jstor.org/
stable/686655.

Howson:1972:PMG

[How72] Colin Howson. The plain man’s guide to probability. British Journal for the Philosophy of Science, 23(2):157–170, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/2/157.full.pdf+
html;   http://www.jstor.
org/stable/686440.

Howson:1973:MLP

[How73] C. Howson. Must the logical probability of laws be zero? British Journal for the Philosophy of Science, 24(2):153–163, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/2/153.full.pdf+
html;   http://www.jstor.
org/stable/686606.

Howard:1975:PM

[How75a] V. A. Howard. The pragmatic maximum. British Journal for the Philosophy of Science, 26(4):

343–351, December 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/4/343.full.pdf+html;
http://www.jstor.org/
stable/686680.

Howson:1975:ERI

[How75b] Colin Howson. The end of the road for inductive logic? British Journal for the Philosophy of Science, 26(2):143–149, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/2/143.full.pdf+
html;   http://www.jstor.
org/stable/686514.

Howson:1975:RSI

[How75c] Colin Howson. The rule of succession, inductive logic, and probability logic. British Journal for the Philosophy of Science, 26(3):187–198, September 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/187.full.pdf+html;
http://www.jstor.org/
stable/686276.

Howson:1978:PC

[How78] Colin Howson. The prehistory of chance. British Journal for the Philosophy of Science, 29(3):274–280, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/274.full.pdf+html;
http://www.jstor.org/
stable/687018.

Howard:1980:R

[How80] Nigel Howard. Reviews. British Journal for the Philosophy of Science, 31(2):199–203, June 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/2/199.full.pdf+
html;   http://www.jstor.
org/stable/687189.

Howson:1984:BSN

[How84] Colin Howson. Bayesianism and support by novel facts. British Journal for the Philosophy of Science, 35(3):245–251, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/3/245.full.pdf+
html;   http://www.jstor.
org/stable/687475.

Howson:1985:SRO

[How85] Colin Howson. Some recent objections to the Bayesian theory of support. British Journal for the Philosophy of Science, 36(3):305–309, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/305.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687574.

Howson:1987:PPP

[How87a] Colin Howson. Popper, prior probabilities, and inductive inference. British Journal for the Philosophy of Science, 38 (2):207–224, June 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/2/207.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687048.

Howson:1987:R

[How87b] Colin Howson. Reviews. British Journal for the Philosophy of Science, 38(2):268–272, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/2/268.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687054.

Howson:1991:OEP

[How91] Colin Howson. The ‘old evidence’ problem. British Journal for the Philosophy of Science, 42(4):547–555, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/4/547.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687709.

Howson:1995:TP

[How95] Colin Howson. Theories of probability. British Journal for the Philosophy of Science, 46 (1):1–32, March 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/1/1.full.pdf+html;
http://www.jstor.org/
stable/687631.

Howson:1997:LP

[How97a] Colin Howson. Logic and probability. British Journal for the Philosophy of Science, 48(4): 517–531, December 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/4/517.full.pdf+html;
http://www.jstor.org/
stable/688097.

Howson:1997:CHU

[How97b] Colin Howson. On Chihara’s ‘The Howson–Urbach Proofs of Bayesian Principles’. British Journal for the Philosophy of Science, 48(1):83–90, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/83.full.pdf+
html;   http://www.jstor.
org/stable/687719.

Howson:2008:FCA

[How08]  Colin Howson.  De Finetti,

countable additivity, consistency and coherence. British Journal for the Philosophy of Science, 59(1):1–23, March 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/1/1.full.pdf+html;
http://www.jstor.org/
stable/30116991.

Hemmo:2003:PNM

[HP03] Meir Hemmo and Itamar Pitowsky. Probability and nonlocality in many minds interpretations of quantum mechanics. British Journal for the Philosophy of Science, 54(2):225–243, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/225.full.pdf+
html;   http://www.jstor.
org/stable/3541965.

Halpern:2005:CESa

[HP05a] Joseph Y. Halpern and Judea Pearl. Causes and explanations: a structural-model approach. Part I: Causes. British Journal for the Philosophy of Science, 56(4):843–887, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/4/843.full.pdf+
html;   http://www.jstor.
org/stable/3541870.

Halpern:2005:CESb

[HP05b] Joseph Y. Halpern and Judea Pearl. Causes and explanations: a structural-model approach. Part II: Explanations. British Journal for the Philosophy of Science, 56(4):889– 911, December 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/889.full.pdf+html;
http://www.jstor.org/
stable/3541871.

Hooker:1992:CCC

[HPE92] C. A. Hooker, H. B. Penfold, and R. J. Evans. Control, connectionism and cognition: Towards a new regulatory paradigm. British Journal for the Philosophy of Science, 43(4): 517–536, October 1992. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/4/517.full.pdf+html;
http://www.jstor.org/
stable/687839.

Hoefer:1992:R

[HR92] Carl Hoefer and Christopher Ray. Review. British Journal for the Philosophy of Science, 43 (4):573–580, October 1992. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/4/573.full.pdf+html;
http://www.jstor.org/
stable/687843.

Hansson:1995:HCT

[HR95] Sven Ove Hansson and Hans Rott. How not to change the theory of theory change: a reply to Tennant. British Journal for the Philosophy of Science, 46(3):361–380, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/3/361.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687663.

HellemAaby:2022:TKN

[HR22] Bendik Hellem Aaby and Grant Ramsey. Three kinds of niche construction. British Journal for the Philosophy of Science, 73(2): 351–372, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hooker:1968:RR

[HS68] C. A. Hooker and D. Stove. Relevance and the ravens. British Journal for the Philosophy of Science, 18(4):305–315, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/4/305.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686359.

Hitchcock:2004:PVA

[HS04] Christopher Hitchcock and Elliott Sober. Prediction versus accommodation and the risk of overfitting. British Journal for the Philosophy of Science, 55(1):1–34, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/1.full.pdf+html;
http://www.jstor.org/
stable/3541832.

Hicks:2017:DPF

[HS17] Michael Townsen Hicks and Jonathan Schaffer. Derivative properties in fundamental laws. British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/2/411/3044192/
Derivative-Properties-in-
Fundamental-Laws.

Hemmo:2022:SLT

[HS22] Meir Hemmo and Orly Shenker.

The Second Law of Thermodynamics and the psychological Arrow of Time. British Journal for the Philosophy of Science, 73(1): 85–107, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hofer-Szabo:1999:RCC

[HSRS99] Gábor Hofer-Szabó, Miklós Rédei, and László E. Szabó. On Reichenbach’s common cause principle and Reichenbach’s notion of common cause. British Journal for the Philosophy of Science, 50(3):377–399, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/50/3/377.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072237.

Huber:2005:SPB

[Hub05] Franz Huber. Subjective probabilities as basis for scientific reasoning? British Journal for the Philosophy of Science, 56 (1):101–116, March 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/1/101.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541851.

Huber:2008:RCA

[Hub08] Franz Huber. Reply to Crupi et al.’s ‘Bayesian Confirmation by Uncertain Evidence’ ([2008]). British Journal for the Philosophy of Science, 59(2):213–215, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/2/213.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072278.  See

[CFM08].

Huber:2011:LCS

[Hub11] Franz Huber. Lewis causation is a special case of Spohn causation. British Journal for the Philosophy of Science, 62 (1):207–210, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
62/1/207.full.pdf+html.

Hudson:2001:DWW

[Hud01] Robert G. Hudson. Discoveries, when and by whom? British Journal for the Philosophy of Science, 52(1):75–93, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/75.full.pdf+
html;   http://www.jstor.
org/stable/3541944.

Hudson:2020:RJP

[Hud20]  Robert Hudson.  The real-

ity of Jean Perrin’s atoms and molecules. British Journal for the Philosophy of Science, 71 (1):33–58, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
33/4683640.

Huemer:2009:EAT

[Hue09] Michael Huemer. Explanationist aid for the theory of inductive logic. British Journal for the Philosophy of Science, 60(2):345–375, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/2/345.full.pdf+html;
http://www.jstor.org/
stable/25592004.

Hughes:1981:AR

[Hug81] Martin Hughes. Absolute rotation. British Journal for the Philosophy of Science, 32(4): 359–366, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/4/359.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687306.

Hughes:1992:NHB

[Hug92] M. Hughes. Newton, Hermes and Berkeley. British Journal for the Philosophy of Science, 43(1):1–19, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        43/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687882.

Huggett:1999:SPS

[Hug99] N. Huggett. On the significance of permutation symmetry. British Journal for the Philosophy of Science, 50(3):325– 347, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/3/325.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072235.

Huggett:2000:PFQ

[Hug00] Nick Huggett. Philosophical foundations of quantum field theory. British Journal for the Philosophy of Science, 51(4): 617–637, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541610.

Huggett:2004:CSD

[Hug04] Nick Huggett. Cartesian spacetime: Descartes’ physics and the relational theory of space and motion. British Journal for the Philosophy of Science, 55 (1):189–193, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/189.full.pdf+html;
http://www.jstor.org/
stable/3541844.

Huggett:2008:WPA

[Hug08] Nick Huggett. Why the parts of absolute space are immobile. British Journal for the Philosophy of Science, 59(3):391– 407, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/391.full.pdf+html;
http://www.jstor.org/
stable/40072291.

Hull:1965:EETa

[Hul65a]  David L. Hull.  The effect

of essentialism on taxonomy — two thousand years of stasis (I). British Journal for the Philosophy of Science, XV(60):

314–326, February 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XV/60/314.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686538.

Hull:1965:EETb

[Hul65b] David L. Hull. The effect of essentialism on taxonomy — two thousand years of stasis (II). British Journal for the Philosophy of Science, XVI(61):1–18, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/61/1.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686135.

Hull:1970:MBR

[Hul70] David L. Hull. Morphospecies and biospecies: a reply to Ruse. British Journal for the Philosophy of Science, 21(3):280–282, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/280.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686200.

Hull:1974:MBS

[Hul74] David L. Hull. Are the ‘members’ of biological species ‘similar’ to each other? British Journal for the Philosophy of Science, 25(4):332–334, December 1974. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/4/332.full.pdf+
html;   http://www.jstor.
org/stable/686460.

Hull:1977:LEL

[Hul77] David L. Hull. A logical empiricist looks at biology. British Journal for the Philosophy of Science, 28(2):181–189, June 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/2/181.full.pdf+
html;   http://www.jstor.
org/stable/686634.

Hull:1998:R

[Hul98] David L. Hull. Review. British Journal for the Philosophy of Science, 49(3):511–514, September 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/511.full.pdf+
html;   http://www.jstor.
org/stable/688088.

Humphreys:1983:R

[Hum83]  Paul Humphreys.  Reviews.

British Journal for the Philosophy of Science, 34(4):410– 412, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/410.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687537.

Humphreys:2004:SCC

[Hum04] Paul Humphreys. Some considerations on conditional chances. British Journal for the Philosophy of Science, 55(4):667– 680, December 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/4/667.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541621.

Hunt:1969:FRG

[Hun69a] G. M. K. Hunt. Further ramifications of ‘Grue’. British Journal for the Philosophy of Science, 20 (3):257–259, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/3/257.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686419.

Hunt:1969:R

[Hun69b] G. M. K. Hunt. Reviews. British Journal for the Philosophy of Science, 20(3):276–277, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/276.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686426.

Hunt:1970:CVS

[Hun70] G. M. K. Hunt. A conditional vindication of the straight rule. British Journal for the Philosophy of Science, 21(2):198–199, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/198.full.pdf+
html;   http://www.jstor.
org/stable/686701.

Hung:1981:NNC

[Hun81] H.-C. Hung. Nomic necessity is cross-theoretic. British Journal for the Philosophy of Science, 32(3):219–236, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/219.full.pdf+html;
http://www.jstor.org/
stable/687588.

Hung:2006:BKS

[Hun06] Edwin H.-C. Hung. Beyond Kuhn: scientific explanation, theory structure, incommensurability and physical necessity. Ashgate new critical thinking in philosophy. Ashgate, Aldershot, Hampshire, England, 2006. ISBN 0-7546-5347-1. xv + 146 pp. LCCN Q175 .H925

2006. URL http://www.loc.
gov/catdir/toc/ecip059/
2005007947.html.

Hunt:2023:EDU

[Hun23] Josh Hunt. Epistemic dependence and understanding: Reformulating through symmetry. British Journal for the Philosophy of Science, 74(4):941–974, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Hutten:1951:RU

[Hut51a] E. H. Hutten. Review: [untitled]. British Journal for the Philosophy of Science, II(5):81, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

                                        jstor.org/stable/685618.

Hutten:1951:R

[Hut51b] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, II(6):172–174, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/6/172.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685515.

Hutten:1951:TR

[Hut51c] E. H. Hutten. Theories of Relativity. British Journal for the Philosophy of Science, II (5):81a–81, May 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/5/81-a.full.pdf+html.

Hutten:1951:PAC

[Hut51d] Ernest H. Hutten. On the principle of action by contact. British Journal for the Philosophy of Science, II(5):45–51, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/5/45.full.pdf+html;
http://www.jstor.org/
stable/685607.

Hutten:1952:Ra

[Hut52a] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, II(8):325–326, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/325.full.pdf+
html;   http://www.jstor.
org/stable/686024.

Hutten:1952:Rb

[Hut52b] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, II(8):331–332, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/331.full.pdf+
html;   http://www.jstor.
org/stable/686029.

Hutten:1952:Rc

[Hut52c] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, III(11):286, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/286.full.

                                        pdf+html;   http://www.
                                        jstor.org/stable/685282.

Hutchison:1953:BQP

[Hut53a] T. W. Hutchison. Berkeley’s querist and its place in the economic thought of the Eighteenth Century. British Journal for the Philosophy of Science, IV(13):52–77, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/13/52.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686046.

Hutten:1953:OHR

[Hut53b] E. H. Hutten. Obituary: Hans Reichenbach. British Journal for the Philosophy of Science, IV (14):93–94, August 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/14/93.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685984.

Hutten:1953:RU

[Hut53c] E. H. Hutten. Review: [untitled]. British Journal for the Philosophy of Science, III (12):382, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685463.

Hutten:1953:Ra

[Hut53d] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, III(12):382a–382, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/382-a.full.pdf+html.

Hutten:1953:RS

[Hut53e] E. H. Hutten. Reviews: Semantics. British Journal for the Philosophy of Science, IV (15):229–235, November 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/229.full.pdf+html;
http://www.jstor.org/
stable/685900.

Hutten:1953:RMP

[Hut53f] E. H. Hutten. The Rôle of models in physics. British Journal for the Philosophy of Science, IV(16):284–301, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/284.full.pdf+html.

Hutten:1953:Rb

[Hut53g] Ernest H. Hutten. Reviews.

British Journal for the Philosophy of Science, IV(15):258– 260, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/258.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685908.

Hutten:1954:I

[Hut54a] E. H. Hutten. Indeterminism. British Journal for the Philosophy of Science, V(18):159–164, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/18/159.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685176.

Hutten:1954:MAP

[Hut54b] E. H. Hutten. Methodological aspects of physics and geometry. British Journal for the Philosophy of Science, V(19): 256–261, November 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/19/256.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685721.

Hutten:1954:Ra

[Hut54c] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, V(17):76–80, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/17/76.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685956.

Hutten:1954:Rb

[Hut54d] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, V(18):169–171, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/169.full.pdf+
html;   http://www.jstor.
org/stable/685180.

Hutten:1954:RMP

[Hut54e] E. H. Hutten. The Rôle of models in physics. British Journal for the Philosophy of Science, IV(16):284–301, February 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686078.

Hutten:1955:NAU

[Hut55a] E. H. Hutten. A note on ‘the age of the universe’. British Journal for the Philosophy of Science, VI(21):58–61, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/21/58.full.pdf+html;
http://www.jstor.org/
stable/685575.

Hutten:1955:Ra

[Hut55b] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, V(20):352–354, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/352.full.pdf+
html;   http://www.jstor.
org/stable/685747.

Hutten:1955:Rc

[Hut55c] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, VI(23):256–257, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/23/256.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685689.

Hutten:1955:MPA

[Hut55d] Ernest H. Hutten. The methodology of psycho-analysis. British Journal for the Philosophy of Science, VI(21):78–88, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/21/78.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685579.

Hutten:1955:Rb

[Hut55e] Ernest H. Hutten. Reviews. British Journal for the Philosophy of Science, V(20):354– 355, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/20/354.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685748.

Hutten:1956:EPP

[Hut56a] Ernest H. Hutten. On explanation in psychology and in physics. British Journal for the Philosophy of Science, VII (25):73–85, May 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/25/73.full.pdf+html;
http://www.jstor.org/
stable/685937.

Hutten:1956:R

[Hut56b] Ernest H. Hutten. Reviews.

British Journal for the Philosophy of Science, VII(25):97–100, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/25/97.full.pdf+
html;   http://www.jstor.
org/stable/685939.

Hutten:1958:R

[Hut58a] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, VIII(32):347–348, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/347.full.
pdf+html;   http://www.
jstor.org/stable/685865.

Hutten:1958:PI

[Hut58b] Ernest H. Hutten. Probability and induction. British Journal for the Philosophy of Science, IX(33):43–51, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        IX/33/43.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686101.

Hutten:1959:Ra

[Hut59a] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, IX(36):328–330, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/36/328.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685195.

Hutten:1959:Rb

[Hut59b] Ernest H. Hutten. Reviews. British Journal for the Philosophy of Science, X(37):78–81, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/37/78.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685782.

Hutten:1960:Ra

[Hut60a] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, X(40):344–346, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/40/344.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685924.

Hutten:1960:Rd

[Hut60b] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, XI(43):261–262, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/43/261.full.pdf+
html;   http://www.jstor.
org/stable/685320.

Hutten:1960:Re

[Hut60c] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, XI(43):262–263, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/43/262.full.pdf+
html;   http://www.jstor.
org/stable/685321.

Hutten:1960:PB

[Hut60d] Ernest H. Hutten. Physics and biology. British Journal for the Philosophy of Science, XI(42): 101–108, August 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/42/101.full.pdf+html;
http://www.jstor.org/
stable/685582.

Hutten:1960:Rc

[Hut60e] Ernest H. Hutten. Reviews.

British Journal for the Philosophy of Science, XI(43):257– 259, November 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/257.full.pdf+html;
http://www.jstor.org/
stable/685317.

Hutten:1960:Rb

[Hut60f]  Joan  M.

Hutten. Reviews. British Journal for the Philosophy of Science, X(40):351, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/351.full.pdf+html;
http://www.jstor.org/
stable/685927.

Hutten:1962:DP

[Hut62] Ernest H. Hutten. Discovering the positron. British Journal for the Philosophy of Science, XIII(49):54–55, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/49/54.full.pdf+html;
http://www.jstor.org/
stable/685227.

Hutten:1963:Rc

[Hut63a] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, XIV(54):166–169, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/54/166.full.
pdf+html;   http://www.
jstor.org/stable/685437.

Hutten:1963:Ra

[Hut63b] Ernest H. Hutten. Reviews.

British Journal for the Philosophy of Science, XIV(53):63–65, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/63.full.pdf+
html;   http://www.jstor.
org/stable/685354.

Hutten:1963:Rb

[Hut63c] Ernest H. Hutten. Reviews.

British Journal for the Philosophy of Science, XIV(53):71–72, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/71.full.pdf+
html;   http://www.jstor.
org/stable/685357.

Hutten:1965:LPS

[Hut65a] E. H. Hutten. The logic and psychology of science. British Journal for the Philosophy of Science, XVI(61):51–52, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/61/51.full.pdf+html;
http://www.jstor.org/
stable/686140.

Hutten:1965:R

[Hut65b] E. H. Hutten. Reviews. British Journal for the Philosophy of Science, XVI(62):165–166, August 1965. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/62/165.full.
pdf+html;   http://www.
jstor.org/stable/686158.

Hutchison:1983:PDT

[Hut83] Keith Hutchison. Planetary distances as a test for the Copernican theory. British Journal for the Philosophy of Science, 34(4):369–371, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/369.full.pdf+html;
http://www.jstor.org/
stable/687527.

Hutchison:1993:CMR

[Hut93] Keith Hutchison. Is classical mechanics really time-reversible and deterministic? British Journal for the Philosophy of Science, 44(2):307–323, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/2/307.full.pdf+html;
http://www.jstor.org/
stable/687650.

Hutchison:1995:DCT

[Hut95a] Keith Hutchison. Differing criteria for temporal symmetry. British Journal for the Philosophy of Science, 46(3):341– 347, September 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/3/341.full.pdf+html;
http://www.jstor.org/
stable/687661.

Hutchison:1995:TAC

[Hut95b] Keith Hutchison. Temporal asymmetry in classical mechanics. British Journal for the Philosophy of Science, 46(2): 219–234, June 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/2/219.full.pdf+html;
http://www.jstor.org/
stable/687813.

Hutchison:1999:WCP

[Hut99] K. Hutchison. What are conditional probabilities conditional upon? British Journal for the Philosophy of Science, 50(4): 665–695, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/4/665.full.pdf+html;
http://www.jstor.org/
stable/40072261.

Huttemann:2004:WWM

[Hüt04] Andreas Hüttemann. What’s wrong with microphysicalism? International library of philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2004. ISBN 0-415-32794-6 (cloth). vii + 141 pp. LCCN QC173 .H87 2004. URL http:

//www.loc.gov/catdir/

                                        enhancements/fy0651/
                                        2003014129-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip045/2003014129.html.

Hausman:1999:IIC

[HW99] D. M. Hausman and J. Woodward. Independence, invariance and the causal Markov condition. British Journal for the Philosophy of Science, 50(4): 521–583, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/4/521.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072257.

Hausman:2004:MCM

[HW04] Daniel M. Hausman and James Woodward. Modularity and the causal Markov condition: a restatement. British Journal for the Philosophy of Science, 55 (1):147–161, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/1/147.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541839. See comment

[Ste06a].

Haggqvist:2018:NKN

[HW18] Sören Häggqvist and Åsa Wikforss. Natural kinds and natural kind terms: Myth and reality. British Journal for the Philosophy of Science, 69(4):911–933, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/911/3748467.

Hilborn:2002:IPQ

[HY02] Robert C. Hilborn and Candice L. Yuca. Identical particles in quantum mechanics revisited. British Journal for the Philosophy of Science, 53(3): 355–389, September 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/355.full.pdf+html;
http://www.jstor.org/
stable/3541887.

Huttegger:2013:MBG

[HZ13] Simon M. Huttegger and Kevin J. S. Zollman. Methodology in biological game theory. British Journal for the Philosophy of Science, 64(3):637– 658, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/637.full.pdf+html.

Irzik:1995:CKA

[IG95] Gürol Irzik and Teo Grünberg.

Carnap and Kuhn: Arch enemies or close allies? British Journal for the Philosophy of Science, 46(3):285–307, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/3/285.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687658.

Irzik:2019:WEP

[IK19] Gürol Irzik and Faik Kurtulmus. What is epistemic public trust in science? British Journal for the Philosophy of Science, 70 (4):1145–1166, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/4/
                                        1145/4819210.

Irzik:2024:DEJ

[IK24] Gürol Irzik and Faik Kurtulmus. Distributive epistemic justice in science. British Journal for the Philosophy of Science, 75 (2):325–345, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Irvine:2012:OPN

[Irv12] Elizabeth Irvine. Old problems with new measures in the science of consciousness. British Journal for the Philosophy of Science, 63(3):627–648, September 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/3/627.full.pdf+html.

Irvine:2013:CSC

[Irv13] Elizabeth Irvine. Consciousness as a scientific concept: a philosophy of science perspective, volume 5 of Studies in brain and mind. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5172-9, 94-007-5173-7 (e-book). xi + 182 pp. LCCN QP411 .I78 2013.

Irvine:2016:MBT

[Irv16] Elizabeth Irvine. Model-based theorizing in cognitive neuroscience. unknown, 67(1):143– 168, March 2016. CODEN ????

ISSN ???? URL http://bjps.
oxfordjournals.org/content/
67/1/143.

Irzik:1985:PPE

[Irz85] Gürol Irzik. Popper’s piecemeal engineering: What is good for science is not always good for society. British Journal for the Philosophy of Science, 36 (1):1–10, March 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/1/1.full.pdf+html;
http://www.jstor.org/
stable/687152.

Irzik:1996:CCR

[Irz96] Gürol Irzik. Can causes be reduced to correlations? British Journal for the Philosophy of Science, 47(2):249–270, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/249.full.pdf+
html;   http://www.jstor.
org/stable/687947.

Incurvati:2024:ED

[IS24] Luca Incurvati and Giorgio Sbardolini. The evolution of denial. British Journal for the Philosophy of Science, 75(1):177– 208, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Isaac:2019:SLS

[Isa19] Alistair M. C. Isaac. The semantics latent in Shannon information. British Journal for the Philosophy of Science, 70 (1):103–125, March 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/1/
                                        103/4092772.

Ismael:1996:WCC

[Ism96] Jenann Ismael. What chances could not be. British Journal for the Philosophy of Science, 47(1):79–91, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        47/1/79.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/1215827.

Jabs:1992:IFQ

[Jab92] Arthur Jabs. An interpretation of the formalism of quantum mechanics in terms of epistemological realism. British Journal for the Philosophy of Science, 43(3):405–421, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/3/405.full.pdf+html;
http://www.jstor.org/
stable/687766.

Jacob:1953:R

[Jac53] E. F. Jacob. Reviews. British Journal for the Philosophy of Science, IV(14):173–175, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/173.full.pdf+
html;   http://www.jstor.
org/stable/685996.

Jammer:1963:FE

[Jam63] Max Jammer. The factorisation of energy. British Journal for the Philosophy of Science, XIV(54):160–166, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/54/160.full.pdf+html;
http://www.jstor.org/
stable/685436.

James:2023:NPT

[Jam23] Lucy James. A new perspective on time and physical laws. British Journal for the Philosophy of Science, 73(4):849–877, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Jarvie:1961:NAM

[Jar61a] I. C. Jarvie. Nadel on the aims and methods of social anthropology. British Journal for the Philosophy of Science, XII(45):1–24, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/1.full.pdf+
html;   http://www.jstor.
org/stable/685535.

Jarvie:1961:Ra

[Jar61b] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XII(45):73– 77, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/73.full.pdf+
html;   http://www.jstor.
org/stable/685545.

Jarvie:1961:Rb

[Jar61c] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XII(45):77– 79, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/77.full.pdf+
html;   http://www.jstor.
org/stable/685546.

Jarvie:1963:R

[Jar63] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XIII(52):327–329, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/XIII/52/327.full.
pdf+html;   http://www.
jstor.org/stable/685335.

Jarvie:1964:ESS

[Jar64a] I. C. Jarvie. Explanation in social science. British Journal for the Philosophy of Science, XV(57):62–72, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/57/62.full.pdf+html;
http://www.jstor.org/
stable/685627.

Jarvie:1964:KPS

[Jar64b] I. C. Jarvie. Key problems of sociology. British Journal for the Philosophy of Science, XIV(56): 352–355, February 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/56/352.full.pdf+html;
http://www.jstor.org/
stable/685642.

Jarvie:1964:Ra

[Jar64c] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XV(58):151–158, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/151.full.pdf+
html;   http://www.jstor.
org/stable/685257.

Jarvie:1964:Rb

[Jar64d] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XV(59):264–266, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/59/264.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686529.

Jarvie:1964:Rc

[Jar64e] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, XV(59):266–268, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/59/266.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686530.

Jarvie:1965:RPH

[Jar65] L. C. Jarvie. Reply to Professor Horowitz. British Journal for the Philosophy of Science, XVI(63):241–242, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/63/241.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686306.

Jardine:1967:CHB

[Jar67a] N. Jardine. The concept of homology in biology. British Journal for the Philosophy of Science, 18(2):125–139, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/125.full.pdf+
html;   http://www.jstor.
org/stable/686580.

Jardine:1967:R

[Jar67b] N. Jardine. Reviews. British Journal for the Philosophy of Science, 18(1):78–79, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/78.full.pdf+html;
http://www.jstor.org/
stable/686897.

Jardine:1969:R

[Jar69] N. Jardine. Reviews. British Journal for the Philosophy of Science, 20(2):172, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/2/172.full.pdf+html;
http://www.jstor.org/
stable/686183.

Jarvie:1973:R

[Jar73] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, 24(2):193–199, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/2/193.full.pdf+
html;   http://www.jstor.
org/stable/686612.

Jarvie:1974:R

[Jar74] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, 25(2):195–198, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/25/2/195.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686824.

Jarvie:1979:RU

[Jar79a] I. C. Jarvie. Review: [untitled]. British Journal for the Philosophy of Science, 30(1): 100–104, March 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687425.

Jarvie:1979:R

[Jar79b] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, 30(1):100b–104, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/1/100-b.full.pdf+html.

Jarvie:1980:R

[Jar80] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, 31(4):396–401, December 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/31/4/396.full.pdf+

html;   http://www.jstor.
org/stable/687067.

Jarvie:1983:R

[Jar83] I. C. Jarvie. Reviews. British Journal for the Philosophy of Science, 34(3):289–295, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/3/289.full.pdf+
html;   http://www.jstor.
org/stable/687326.

Jardine:2012:BRL

[Jar12] Nick Jardine. Book review: Lorraine Daston and Peter Galison, Objectivity. British Journal for the Philosophy of Science, 63(4):885–893, December 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/4/885.full.pdf+html.

Jeffreys:1952:R

[Jef52] Harold Jeffreys. Reviews. British Journal for the Philosophy of Science, III(11):276–277, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/276.full.
pdf+html;   http://www.
jstor.org/stable/685275.

Jeffreys:1953:R

[Jef53] Harold Jeffreys. Reviews. British Journal for the Philosophy of Science, IV(16):348–349, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
IV/16/348.full.pdf+html.

Jeffreys:1954:RU

[Jef54] Harold Jeffreys. Review: [untitled]. British Journal for the Philosophy of Science, IV(16): 348–349, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686087.

Jeffreys:1955:PPP

[Jef55] Harold Jeffreys. The present position in probability theory. British Journal for the Philosophy of Science, V(20):275– 289, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/275.full.pdf+html;
http://www.jstor.org/
stable/685730.

Jeffreys:1956:R

[Jef56] Harold Jeffreys. Reviews. British Journal for the Philosophy of Science, VII(26):174–175, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/26/174.full.
pdf+html;   http://www.
jstor.org/stable/685701.

Jennings:1984:TRS

[Jen84] Richard C. Jennings. Truth, rationality and the sociology of science. British Journal for the Philosophy of Science, 35(3): 201–211, September 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/201.full.pdf+html;
http://www.jstor.org/
stable/687472.

Jennings:1986:BRR

[Jen86] Richard C. Jennings. Book review: Richards, Stewart [1983]: Philosophy & Sociology of Science. An Introduction. Basil Blackwell. viii+210 pp. (Hardback £17.50) (Paperback £6.50). British Journal for the Philosophy of Science, 37(2):246–249, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686988.

Jennings:1988:LAP

[Jen88] Richard C. Jennings. Leviathan and the air-pump: Hobbes, Boyle, and the experimental life. British Journal for the Philosophy of Science, 39(3):403– 410, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/403.full.pdf+html;
http://www.jstor.org/
stable/687218.

Jennings:1989:ZLW

[Jen89] Richard C. Jennings. Zande logic and Western logic. British Journal for the Philosophy of Science, 40(2):275–285, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/2/275.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687519.

Jennings:1992:R

[Jen92] Richard C. Jennings. Review. British Journal for the Philosophy of Science, 43(4):561–571, October 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/4/561.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687842.

Jenkins:2005:KA

[Jen05] C. S. Jenkins. Knowledge of arithmetic. British Journal for the Philosophy of Science, 56(4): 727–747, December 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/4/727.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541865.

Jenson:2016:BI

[Jen16] J. Christopher Jenson. The belief illusion. British Journal for the Philosophy of Science, 67(4):965–995, December 2016.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/965.

Jacobs:2023:CAI

[JHB^{+}23] Caspar Jacobs, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Ludwig von Bertalanffy, Joseph Y. Halpern, Judea Pearl, Matteo Mossio, Cristian Saborido, and Alvaro Moreno. The coalescence approach to inequivalent representation: Pre-QM_{∞} parallels. British Journal for the Philosophy of Science, 74(4):1069–1090, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Jaramillo:2022:CIV

[JL22] José Luis Jaramillo and Vincent Lam. Counterfactuals in the initial value formulation of general relativity. British Journal for the Philosophy of Science, 72(4): 1111–1128, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Jablonka:2005:EFD

[JLZ05] Eva Jablonka, Marion J. Lamb, and Anna Zeligowski. Evolution in four dimensions: genetic, epigenetic, behavioral, and symbolic variation in the history of life. Life and mind: philosophical issues in biology and psychology. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-10107-6 (hardcover), 0-262-60069-2 (paperback). x + 462 pp. LCCN QH366.2 .J322 2005.

Jobe:1980:RTR

[Job80] Evan K. Jobe. Reichenbach’s theory of reasonable assertion. British Journal for the Philosophy of Science, 31(4):375– 384, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/4/375.full.pdf+html;
http://www.jstor.org/
stable/687064.

Johnstone:1987:TSF

[Joh87] D. J. Johnstone. Tests of significance following R. A. Fisher. British Journal for the Philosophy of Science, 38(4):481– 499, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/481.full.pdf+html;
http://www.jstor.org/
stable/687354.

Johnstone:1988:HTC

[Joh88] D. J. Johnstone. Hypothesis tests and confidence intervals in the single case. British Journal for the Philosophy of Science, 39(3):353–360, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/353.full.pdf+html;
http://www.jstor.org/
stable/687211.

Johnstone:1989:NRS

[Joh89] D. J. Johnstone. On the necessity for random sampling. British Journal for the Philosophy of Science, 40(4):443–457, December 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/4/443.full.pdf+
html;   http://www.jstor.
org/stable/687735.

Johansen:2003:R

[Joh03] Thomas Kjeller Johansen. Review. British Journal for the Philosophy of Science, 54(4): 617–621, December 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/4/617.full.pdf+html;
http://www.jstor.org/
stable/3541683.

John:2015:BRA

[Joh15] Stephen John. Book review:

Alex Broadbent, Philosophy of Epidemiology. British Journal for the Philosophy of Science, 66(3):707–711, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/707.

Jones:1956:PN

[Jon56] Ernest Jones. Prefatory note.

British Journal for the Philosophy of Science, VII(25):1, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

                                        jstor.org/stable/685930.

Jones:1966:R

[Jon66] Peter Jones. Reviews. British Journal for the Philosophy of Science, 17(2):153–155, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/2/153.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686349.

Jones:1969:NAG

[Jon69a] Bill Jones. A note on Alan Gauld’s ‘could a machine perceive?’. British Journal for the Philosophy of Science, 20(3): 261–262, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/3/261.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686421.

Jones:1969:R

[Jon69b] Peter Jones. Reviews. British Journal for the Philosophy of Science, 20(3):279–280, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/279.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686428.

Jones:1971:R

[Jon71]  P. Lloyd Jones.  Reviews.

British Journal for the Philosophy of Science, 22(2):183–187, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/183.full.pdf+
html;   http://www.jstor.
org/stable/686487.

Jones:1973:VVP

[Jon73] Keith E. Jones. Verisimilitude versus probable verisimilitude. British Journal for the Philosophy of Science, 24(2):174–176, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/2/174.full.pdf+
html;   http://www.jstor.
org/stable/686608.

Jones:1980:R

[Jon80] Roger Jones. Reviews. British Journal for the Philosophy of Science, 31(3):311–315, September 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/3/311.full.pdf+
html;   http://www.jstor.
org/stable/686927.

Jones:1982:SCT

[Jon82a] Gareth Jones. Scientific consistency, two-stage priors and the true value of a parameter. British Journal for the Philosophy of Science, 33(2):133–160, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/133.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687275.

Jones:1982:R

[Jon82b] K. E. Jones. Reviews. British Journal for the Philosophy of Science, 33(3):331–334, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/3/331.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687234.

Jones:1986:KS

[Jon86] Keith Jones. Is Kuhn a sociologist? British Journal for the Philosophy of Science, 37(4):443–452, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687369.

Jones:2018:IMR

[Jon18] Nicholaos Jones. Inference to the more robust explanation. British Journal for the Philosophy of Science, 69(1):75–102, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/1/75/2669770.

Joshi:1982:CEH

[Jos82] V. M. Joshi. A counter example of Hacking against the Long Run Rule. British Journal for the Philosophy of Science, 33(3):287–289, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/287.full.pdf+html;
http://www.jstor.org/
stable/687226.

Josephson:2010:LLZ

[Jos10] Paul R. Josephson. Lenin’s laureate: Zhores Alferov’s life in communist science. Transformations. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01458-0 (hardcover). 307 pp. LCCN QC16.A3417 J67 2010.

Joyce:1999:FCD

[Joy99] James M. Joyce. The foundations of causal decision theory. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-64164-0. xii + 268 pp. LCCN BD184 .J68

1999. URL http://www.loc.
gov/catdir/description/
cam029/98036028.html;
http://www.loc.gov/catdir/
toc/cam025/98036028.html.

Jozsa:1986:AMP

[Joz86] Richard Jozsa. An approach to the modelling of the physical continuum. British Journal for the Philosophy of Science, 37(4):395–404, December 1986.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687366.

Jansson:2019:EA

[JS19]  Lina  Jans-

son and Juha Saatsi. Explanatory abstractions. British Journal for the Philosophy of Science, 70(3):817–844, September 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
817/4682638.

Jansson:2017:QPP

[JT17] Lina Jansson and Jonathan Tallant. Quantitative parsimony: Probably for the better. British Journal for the Philosophy of Science, 68(3):781–803, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/3/781/2669747.

Juhl:1994:SOR

[Juh94] Cory F. Juhl. The speedoptimality of Reichenbach’s straight rule of induction. British Journal for the Philosophy of Science, 45(3):857–863, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/857.full.pdf+
html;   http://www.jstor.
org/stable/687797.

Kahane:1967:BCP

[Kah67] Howard Kahane. Baumer on the Confirmation Paradoxes. British Journal for the Philosophy of Science, 18(1):52–56, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/52.full.pdf+html;
http://www.jstor.org/
stable/686889.

Kahane:1969:ECP

[Kah69]  Howard Kahane.  Elimina-

tive confirmation and paradoxes. British Journal for the Philosophy of Science, 20(2):160–162, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/160.full.pdf+
html;   http://www.jstor.
org/stable/686178.

Kamminga:1983:R

[Kam83] Harmke Kamminga. Reviews.

British Journal for the Philosophy of Science, 34(4):393– 395, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/393.full.pdf+html;
http://www.jstor.org/
stable/687532.

Kamminga:1985:R

[Kam85] Harmke Kamminga. Reviews.

British Journal for the Philosophy of Science, 36(1):85–89, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/85.full.pdf+
html;   http://www.jstor.
org/stable/687161.

Kane-:1984:BEM

[Kan84] Jeffrey Kane. Beyond empiricism: Michael Polanyi reconsidered, volume 6 of American university studies. Series XIV, Education. P. Lang, New York, NY, USA, 1984. ISBN 0-8204-0118-8 (paperback). ISSN 0740-4565. 263 pp. LCCN B945.P584 K36 1984. US$26.00.

Kantorovich:1988:PSJ

[Kan88] Aharon Kantorovich. Philosophy of science: from justification to explanation. British Journal for the Philosophy of Science, 39(4):469–494, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/4/469.full.pdf+
html;   http://www.jstor.
org/stable/687433.

Kapp:1953:R

[Kap53] Reginald O. Kapp. Reviews. British Journal for the Philosophy of Science, IV(14):169–173, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/169.full.pdf+
html;   http://www.jstor.
org/stable/685995.

Kapp:1954:LLM

[Kap54] Reginald O. Kapp. Living and lifeless machines. British Journal for the Philosophy of Science, V (18):91–103, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/91.full.pdf+html;
http://www.jstor.org/
stable/685167.

Kapp:1955:NIG

[Kap55a] R. O. Kapp. A new interpretation of gravitation. British Journal for the Philosophy of Science, V(20):331–332, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/331.full.pdf+
html;   http://www.jstor.
org/stable/685734.

Kapp:1955:RPB

[Kap55b] R. O. Kapp. A reply to Professor Bondi. British Journal for the Philosophy of Science, VI(23):241–243, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/23/241.full.pdf+html;
http://www.jstor.org/
stable/685680.

Kapp:1955:HAO

[Kap55c] Reginald O. Kapp. Hypotheses about the origin and disappearance of matter. British Journal for the Philosophy of Science, VI(23):177–185, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/177.full.pdf+
html;   http://www.jstor.
org/stable/685675.

Kapp:1955:R

[Kap55d] Reginald O. Kapp. Reviews. British Journal for the Philosophy of Science, VI(22):172–173, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/22/172.full.pdf+
html;   http://www.jstor.
org/stable/685528.

Kapp:1956:Ra

[Kap56a] R. O. Kapp. Reviews. British Journal for the Philosophy of Science, VI(24):350–351, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/350.full.pdf+
html;   http://www.jstor.
org/stable/685422.

Kapp:1956:Rb

[Kap56b] Reginald O. Kapp. Reviews. British Journal for the Philosophy of Science, VII(25):100–105, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/25/100.full.

pdf+html;   http://www.
jstor.org/stable/685940.

Kapp:1957:Ra

[Kap57a] R. O. Kapp. Reviews. British Journal for the Philosophy of Science, VIII(29):67– 70, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/67.full.
pdf+html;   http://www.
jstor.org/stable/685383.

Kapp:1957:MDD

[Kap57b] Reginald O. Kapp. Mr Dobbs on ‘diathesis, the self-winding watch, and photosynthesis‘. British Journal for the Philosophy of Science, VIII(30):159–160, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/30/159.full.pdf+html;
http://www.jstor.org/
stable/685795.

Kapp:1957:Rb

[Kap57c] Reginald O. Kapp. Reviews.

British Journal for the Philosophy of Science, VIII(29):81–82, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/29/81.full.
pdf+html;   http://www.
jstor.org/stable/685389.

Kapp:1958:ORU

[Kap58] Reginald O. Kapp. Ockam’s Razor and the unification of physical science. British Journal for the Philosophy of Science, VIII(32):265–280, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/265.full.
pdf+html;   http://www.
jstor.org/stable/685852.

Kapp:1959:S

[Kap59]  Reginald O. Kapp.  Space.

British Journal for the Philosophy of Science, X(37):1–15, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/37/1.full.pdf+
html;   http://www.jstor.
org/stable/685771.

Kapp:1960:RCG

[Kap60]  Reginald O. Kapp.  Reply

to criticisms by G. Schlesinger. British Journal for the Philosophy of Science, XI(41):59–62, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/59.full.pdf+
html;   http://www.
jstor.org/stable/685821. See

[Sch60].

Kapp:1961:RNG

[Kap61] Reginald O. Kapp. Reply to note by G. Schlesinger. British Journal for the Philosophy of Science, XI(44):329–331, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/44/329.full.pdf+
html;   http://www.jstor.
org/stable/685134.

Kapp:1965:LPS

[Kap65] Reginald O. Kapp. The logic and psychology of science. British Journal for the Philosophy of Science, XV(60):333–341, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/333.full.pdf+
html;   http://www.jstor.
org/stable/686543.

Karaca:2012:ECK

[Kar12] Koray Karaca. Editor’s choice:

Kitcher’s explanatory unification, Kaluza–Klein theories, and the normative aspect of higher dimensional unification in physics. British Journal for the Philosophy of Science, 63(2):287–312, June 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/287.full.pdf+html.

Katouzian:1984:KFS

[Kat84] Homa Katouzian. T. S. Kuhn, functionalism, and sociology of knowledge. British Journal for the Philosophy of Science, 35 (2):166–173, June 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        35/2/166.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686967.

Katzav:2013:HMC

[Kat13]  Joel Katzav.  Hybrid mod-

els, climate models, and inference to the best explanation. British Journal for the Philosophy of Science, 64(1): 107–129, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/1/107.full.pdf+html.

Kaveh:2023:EFD

[Kav23] Shahin Kaveh. Escaping the fundamental dichotomy of scientific realism. British Journal for the Philosophy of Science, 74(4): 999–1025, December 2023. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Khalifa:2021:CEP

[KDM21] Kareem Khalifa, Gabriel Doble, and Jared Millson. Counterfactuals and explanatory pluralism. British Journal for the Philosophy of Science, 71(4):1439– 1460, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Keenan:1966:R

[Kee66]  C. W. Keenan.  Reviews.

British Journal for the Philosophy of Science, XVI(64):346– 348, February 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/64/346.full.pdf+html;
http://www.jstor.org/
stable/686839.

Keenan:1970:R

[Kee70] C. W. Keenan. Reviews. British Journal for the Philosophy of Science, 21(1):115–117, February 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/1/115.full.pdf+
html;   http://www.jstor.
org/stable/686475.

Keita:1993:JZL

[Kei93]  Lansana Keita.  Jen-

nings and Zande logic: a note. British Journal for the Philosophy of Science, 44(1):151–156, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/151.full.pdf+
html;   http://www.jstor.
org/stable/687854.

Kekes:1981:R

[Kek81] John Kekes. Reviews. British Journal for the Philosophy of Science, 32(4):411–416, December 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/4/411.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687312.

Kelly:2000:LS

[Kel00] Kevin T. Kelly. The logic of success. British Journal for the Philosophy of Science, 51(4): 639–666, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/3541611.

Keller:2010:MSB

[Kel10] Evelyn Fox Keller. The mirage of a space between nature and nurture. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-4714-8 (hardcover), 0-8223-4731-8 (paperback). 107 pp. LCCN BF341 .K36 2010.

Kennedy:1970:R

[Ken70] Barbara A. Kennedy. Reviews. British Journal for the Philosophy of Science, 21(4):401– 402, November 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/4/401.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686569.

Kennedy:1994:R

[Ken94] J. B. Kennedy. Review. British Journal for the Philosophy of Science, 45(4):1059–1062, December 1994. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/4/1059.full.pdf+
html;   http://www.jstor.
org/stable/687622.

Kernohan:1987:TLF

[Ker87a] Andrew Kernohan. Teleology and logical form. British Journal for the Philosophy of Science, 38(1):27–34, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/27.full.pdf+html;
http://www.jstor.org/
stable/687077.

Kerszberg:1987:AEB

[Ker87b] Pierre Kerszberg. On the alleged equivalence between Newtonian and relativistic cosmology. British Journal for the Philosophy of Science, 38(3):347– 380, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/347.full.pdf+html;
http://www.jstor.org/
stable/687490.

Keswani:1960:MAL

[Kes60]  G.  H.  Keswani.

More about Lorentz transformation equations. British Journal for the Philosophy of Science, XI(41):50–55, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XI/41/50.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685819.

Keswani:1961:RPB

[Kes61] G. H. Keswani. Reply to Professor Born. British Journal for the Philosophy of Science, XII(46):151–153, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XII/46/151.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685495.

Keswani:1965:OCRa

[Kes65a] G. H. Keswani. Origin and concept of Relativity (I). British Journal for the Philosophy of Science, XV(60):286–306, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/60/286.full.pdf+
                                        html;   http://links.jstor.
                                        org/sici?sici=0007-0882%
                                        28196502%2915%3A60%3C286%
                                        3AOACOR%28%3E2.0.CO%3B2-F;
                                         http://www.
                                        jstor.org/stable/686536. See

note [Din65a].

Keswani:1965:OCRb

[Kes65b]  G.  H.  Keswani.

Origin and concept of Relativity (II). British Journal for the Philosophy of Science, XVI(61):19– 32, May 1965. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/61/19.full.pdf+
html;   http://www.jstor.
org/stable/686136.

Keswani:1966:OCRa

[Kes66a] G. H. Keswani. Origin and concept of Relativity (III). British Journal for the Philosophy of Science, XVI(64):273–294, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/64/273.full.
pdf+html;   http://www.
jstor.org/stable/686829.

Keswani:1966:OCRb

[Kes66b] G. H. Keswani. Origin and concept of Relativity (parts I and II): Reply to Professor Dingle and Mr Levinson. British Journal for the Philosophy of Science, 17(2):149–152, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/2/149.full.pdf+
html;   http://www.jstor.
org/stable/686348.

Keswani:1966:OCRc

[Kes66c] G. H. Keswani. Origin and concept of Relativity: Reply to Professor Popper. British Journal for the Philosophy of Science, 17(3):234–236, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/3/234.full.pdf+html;
                                        http://www.jstor.
                                        org/stable/686760. See reply

[Pod75].

Keswani:1985:AT

[Kes85] G. H. Keswani. Accelerated twins. British Journal for the Philosophy of Science, 36 (1):53–61, March 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/1/53.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687155.

Ketland:2000:BRB

[Ket00] J. Ketland. Book review: Popper: Philosophy, Politics and Scientific Method. G. Stokes. British Journal for the Philosophy of Science, 51(2):363–369, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/2/363.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541814.

Ketland:2001:BRS

[Ket01]  Jeffrey Ketland.  Book re-

view: Stephen G. Simpson, Subsystems of Second-Order Arithmetic. British Journal for the Philosophy of Science, 52(1): 191–195, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/191.full.pdf+html;
http://www.jstor.org/
stable/3541956.

Ketland:2004:EAR

[Ket04] Jeffrey Ketland. Empirical adequacy and ramsification. British Journal for the Philosophy of Science, 55(2):287–300, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/2/287.full.pdf+
html;   http://www.jstor.
org/stable/3541692.

King-Farlow:1965:MBP

[KFH65] John King-Farlow and Elton A.

Hall. Man, beast, and philosophical psychology. British Journal for the Philosophy of Science, XVI(62):81–101, August 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/62/81.full.pdf+html;
http://www.jstor.org/
stable/686151.

Karaca:2022:TSE

[KGB^{+}22] Koray Karaca, John C. Granbery, W. J. Wouter Botzen, Olivier Deschenes, Mark Sanders, Mark S. Granovetter, James S. Coleman, John W. Meyer, and Brian Rowan. Two senses of experimental robustness: Result robustness and procedure robustness. British Journal for the Philosophy of Science, 73(1):279–298, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Kuehni:2010:CM

[KH10] Rolf G. Kuehni and C. L. Hardin. Churchland’s metamers. British Journal for the Philosophy of Science, 61(1):81–92, March 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/1/81.full.pdf+html.

Khalidi:2002:NNC

[Kha02] Muhammad Ali Khalidi. Nature and nurture in cognition. British Journal for the Philosophy of Science, 53(2):251–272, June 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/53/2/251.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541766.

Khalidi:2010:IK

[Kha10] Muhammad Ali Khalidi. Interactive kinds. British Journal for the Philosophy of Science, 61(2):335–360, June 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/2/335.full.pdf+html.

Khalidi:2013:NCH

[Kha13a] Muhammad Ali Khalidi. Natural categories and human kinds: classification in the natural and social sciences. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-01274-0 (hardcover). xvi + 250 pp. LCCN BD331 .K43 2013.

Khalifa:2013:REU

[Kha13b] Kareem Khalifa. The role of explanation in understanding. British Journal for the Philosophy of Science, 64(1): 161–187, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/161.full.pdf+html.

Khosrowi:2020:GSA

[Kho20] Donal Khosrowi. Getting serious about shared features. British Journal for the Philosophy of Science, 71(2):523–546, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/71/2/523/5046218.

Kieseppa:1997:AIC

[Kie97] I. A. Kieseppä. Akaike information criterion, curve-fitting, and the philosophical problem of simplicity. British Journal for the Philosophy of Science, 48(1):21–48, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
48/1/21.full.pdf+html;
http://www.jstor.org/
stable/687715.

Kieseppa:2001:SMS

[Kie01] I. A. Kieseppä. Statistical model selection criteria and the philosophical problem of underdetermination. British Journal for the Philosophy of Science, 52(4): 761–794, December 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/4/761.full.pdf+html;
http://www.jstor.org/
stable/3541918.

Kilmister:1962:R

[Kil62]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, XII(48):347– 348, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/48/347.full.pdf+html;
http://www.jstor.org/
stable/685406.

Kilmister:1981:Ra

[Kil81a]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 32(2):206–210, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/2/206.full.pdf+
html;   http://www.jstor.
org/stable/687203.

Kilmister:1981:Rb

[Kil81b]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 32(3):303– 305, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/303.full.pdf+html;
http://www.jstor.org/
stable/687594.

Kilmister:1983:Ra

[Kil83a]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 34(2):197–200, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/197.full.pdf+
html;   http://www.jstor.
org/stable/687456.

Kilmister:1983:Rb

[Kil83b]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 34(2):200–201, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/200.full.pdf+
html;   http://www.jstor.
org/stable/687457.

Kilmister:1983:Rc

[Kil83c]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 34(4):396– 398, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/396.full.pdf+html;
http://www.jstor.org/
stable/687533.

Kilmister:1984:R

[Kil84]  C. W. Kilmister.  Reviews.

British Journal for the Philosophy of Science, 35(4):403– 404, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/403.full.pdf+html;
http://www.jstor.org/
stable/687344.

Kilmister:1991:WRI

[Kil91] C. W. Kilmister. Why is Relativity interesting? British Journal for the Philosophy of Science, 42(3):413–423, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/3/413.full.pdf+
html;   http://www.jstor.
org/stable/687917.

Kilmister:1993:R

[Kil93] C. W. Kilmister. Review. British Journal for the Philosophy of Science, 44(3):577–578, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/3/577.full.pdf+
html;   http://www.jstor.
org/stable/688027.

Kingma:2010:PPP

[Kin10]  Elselijn Kingma.  Paraceta-

mol, poison, and polio: Why Boorse’s account of function fails to distinguish health and disease. British Journal for the Philosophy of Science, 61 (2):241–264, June 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/2/241.full.pdf+html.

Kingma:2016:SSD

[Kin16]  Elselijn Kingma.  Situation-

specific disease and dispositional function. British Journal for the Philosophy of Science, 67 (2):391–404, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/391.

Kirk:2001:BRG

[Kir01] Robert Kirk. Book review:

George Botterill and Peter Carruthers, The Philosophy of Psychology. British Journal for the Philosophy of Science, 52 (1):159–162, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/159.full.pdf+html;
http://www.jstor.org/
stable/3541950.

Kistler:2006:CLN

[Kis06] Max Kistler. Causation and laws of nature, volume 2 of Routledge studies in contemporary philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2006. ISBN 0-415-39859-2. 264 pp. LCCN BD541 .K5713

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0654/2006002909-d.html;
http://www.loc.gov/catdir/
toc/ecip067/2006002909.
html.

Kitcher:1982:G

[Kit82] Philip Kitcher. Genes. British Journal for the Philosophy of Science, 33(4):337–359, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/4/337.full.pdf+
html;   http://www.jstor.
org/stable/687168.

Kitchener:1987:GEP

[Kit87] Richard F. Kitchener. Is genetic epistemology possible? British Journal for the Philosophy of Science, 38(3):283–299, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/3/283.full.pdf+
html;   http://www.jstor.
org/stable/687486.

Kitchener:1991:R

[Kit91] Richard F. Kitchener. Review. British Journal for the Philosophy of Science, 42(2):285–290, June 1991. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/2/285.full.pdf+
html;   http://www.jstor.
org/stable/688061.

Kitcher:2011:SDS

[Kit11] Philip Kitcher. Science in a democratic society. Prometheus Books, Amherst, NY, USA, 2011. ISBN 1-61614-407-6 (hardcover). 270 pp. LCCN Q175.5 .K524 2011.

Kneale:1965:CNC

[KK65] W. C. Kneale and S. Körner.

Committee notice: Change of editorship. British Journal for the Philosophy of Science, XV (60):368a–368, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/60/368-a.full.pdf+html.

Keswani:1983:IRR

[KK83] G. H. Keswani and C. W.

Kilmister. Intimations of Relativity: Relativity before Einstein. British Journal for the Philosophy of Science, 34(4): 343–354, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/343.full.pdf+html;
http://www.jstor.org/
stable/687524.

Kaiser:2017:MCM

[KK17]  Marie  I.  Kaiser

and Beate Krickel. The metaphysics of constitutive mechanistic phenomena. British Journal for the Philosophy of Science, 68(3):745–779, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/745/2669661.

Kampis:2002:ALM

[KKS02] György Kampis, Ladislav Kvasz, and Michael Stöltzner, editors. Appraising Lakatos: mathematics, methodology, and the man, volume 1 of Vienna Circle Institute library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0226-2. LCCN Q175 .A685 2002.

Knuuttila:2017:MIR

[KL17] Tarja Knuuttila and Andrea Loettgers. Modelling as indirect representation? The Lotka–Volterra model revisited. British Journal for the Philosophy of Science, 68(4):1007–1036, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/4/1007/2669647.

Klappholz:1964:VJE

[Kla64] Kurt Klappholz. Value judgments and economics. British Journal for the Philosophy of Science, XV(58):97–114, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/97.full.pdf+
html;   http://www.jstor.
org/stable/685249.

Klappholz:1969:R

[Kla69]  Kurt Klappholz.  Reviews.

British Journal for the Philosophy of Science, 20(2):177–179, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/177.full.pdf+
html;   http://www.jstor.
org/stable/686188.

Klein:1953:OOE

[Kle53] Martin J. Klein. Order, organisation and entropy. British Journal for the Philosophy of Science, IV (14):158–160, August 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/14/158.full.pdf+html;
http://www.jstor.org/
stable/685990.

Kleiner:1970:ELS

[Kle70] Scott A. Kleiner. Erotetic logic and the structure of scientific revolution. British Journal for the Philosophy of Science, 21 (2):149–165, May 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
21/2/149.full.pdf+html;
http://www.jstor.org/
stable/686696.

Kleiner:1985:DWR

[Kle85] Scott A. Kleiner. Darwin’s and Wallace’s revolutionary research programme. British Journal for the Philosophy of Science, 36(4):367–392, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/4/367.full.pdf+html;
http://www.jstor.org/
stable/687398.

Kleiner:1987:R

[Kle87] Scott A. Kleiner. Reviews. British Journal for the Philosophy of Science, 38(2):261–265, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/2/261.full.pdf+
html;   http://www.jstor.
org/stable/687052.

Klee:1992:AMC

[Kle92]  Robert Klee.  Anomalous

monism, ceteris paribus and psychological explanation. British Journal for the Philosophy of Science, 43(3):389–403, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/3/389.full.pdf+
html;   http://www.jstor.
org/stable/687765.

Klee:2002:RPH

[Kle02] Robert Klee. The revenge of Pythagoras: How a mathematical sharp practice undermines the contemporary design argument in astrophysical cosmology. British Journal for the Philosophy of Science, 53(3):331– 354, September 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/331.full.pdf+html;
http://www.jstor.org/
stable/3541886.

Klein:2010:IEN

[Kle10] Colin Klein. Images are not the evidence in neuroimaging. British Journal for the Philosophy of Science, 61(2):265–278, June 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
61/2/265.full.pdf+html.

Klein:2017:CIC

[Kle17] Colin Klein. Consciousness, intention, and command-following in the vegetative state. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/1/27/2995217/
Consciousness-Intention-
and-Command-Following-in.

Kline:1987:PPP

[Kli87] Paul Kline. Philosophy, psychology and psychoanalysis. British Journal for the Philosophy of Science, 38(1):106–116, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/1/106.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687085.

Kuorikoski:2010:EMR

[KLM10] Jaakko Kuorikoski, Aki Lehtinen, and Caterina Marchionni. Economic modelling as robustness analysis. British Journal for the Philosophy of Science, 61(3):541–567, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/3/541.full.pdf+html.

Kahn:1992:NC

[KLS92] James A. Kahn, Steven E. Landsburg, and Alan C. Stockman. On novel confirmation. British Journal for the Philosophy of Science, 43(4):503–516, October 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/4/503.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687838.

Kneale:1954:R

[Kne54]  William  Kneale.

Reviews. British Journal for the Philosophy of Science, V(17):85– 87, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/17/85.full.pdf+
html;   http://www.jstor.
org/stable/685959.

Kneale:1957:R

[Kne57a] William Kneale. Reviews. British Journal for the Philosophy of Science, VIII(31):254–256, November 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/31/254.full.
pdf+html;   http://www.
jstor.org/stable/685764.

Kneale:1957:SAP

[Kne57b] William Kneale. Some aspects of probability and induction: a reply to Mr. Bennett. British Journal for the Philosophy of Science, VIII(29):57–63, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/29/57.full.pdf+html;
http://www.jstor.org/
stable/685381.

Kneale:1960:R

[Kne60] William Kneale. Reviews. British Journal for the Philosophy of Science, XI(43):260–261, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/43/260.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685319.

Kneale:1961:UN

[Kne61] William Kneale. Universality and necessity. British Journal for the Philosophy of Science, XII (46):89–102, August 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XII/46/89.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685489.

Kneale:1962:R

[Kne62] William Kneale. Reviews. British Journal for the Philosophy of Science, XIII(50):184a–184, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/50/184-a.full.pdf+
                                        html.

Kneale:1963:Ra

[Kne63a] William Kneale. Reviews. British Journal for the Philosophy of Science, XIII(52):334–337, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/52/334.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685338.

Kneale:1963:Rb

[Kne63b] William Kneale. Reviews.

British Journal for the Philosophy of Science, XIV(53):73–74, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/73.full.pdf+
html;   http://www.jstor.
org/stable/685359.

Kneale:1965:RU

[Kne65a] William Kneale. Review: [untitled]. British Journal for the Philosophy of Science, XVI(61): 65, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/686145.

Kneale:1965:R

[Kne65b] William Kneale. Reviews. British Journal for the Philosophy of Science, XVI(61):65a–65, May 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/61/65-a.full.pdf+html.

Kneebone:1965:R

[Kne65c] G. T. Kneebone. Reviews.

British Journal for the Philosophy of Science, XVI(61):63–64, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/61/63.full.pdf+

html;   http://www.jstor.
org/stable/686144.

Kneale:1966:R

[Kne66]  William Kneale.  Reviews.

British Journal for the Philosophy of Science, 17(2):160–162, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/2/160.full.pdf+
html;   http://www.jstor.
org/stable/686352.

Kneale:1967:R

[Kne67] William Kneale. Reviews. British Journal for the Philosophy of Science, 18(3):261–262, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/3/261.full.pdf+html;
http://www.jstor.org/
stable/686602.

Kneale:1968:R

[Kne68a] William Kneale. Reviews. British Journal for the Philosophy of Science, 18(4):340–342, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/4/340.full.pdf+html;
http://www.jstor.org/
stable/686364.

Kneale:1968:SRE

[Kne68b] William Kneale. Scientific revolution for ever? British Journal for the Philosophy of Science, 19(1):27–42, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/1/27.full.pdf+html;
http://www.jstor.org/
stable/686846.

Kneale:1971:RPS

[Kne71]  William C. Kneale.  Rus-

sell’s Paradox and some others. British Journal for the Philosophy of Science, 22(4): 321–338, November 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/4/321.full.pdf+html;
http://www.jstor.org/
stable/686315.

Kneale:1972:NN

[Kne72] William C. Kneale. Numbers and numerals. British Journal for the Philosophy of Science, 23 (3):191–206, August 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/3/191.full.pdf+html;
http://www.jstor.org/
stable/686685.

Knox:2010:FOC

[Kno10] Eleanor Knox. Flavour-oscillation clocks and the geometricity of General Relativity. British Journal for the Philosophy of Science, 61(2):433–452, June 2010.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/2/433.full.pdf+html.

Knox:2014:NSS

[Kno14] Eleanor Knox. Newtonian spacetime structure in light of the equivalence principle. British Journal for the Philosophy of Science, 65(4):863–880, December 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/65/4/863.

Koertge:1972:AM

[Koe72] Noretta Koertge. For and against method. British Journal for the Philosophy of Science, 23 (3):274–290, August 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/3/274.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686691.

Koertge:1973:R

[Koe73]  Noretta Koertge.  Reviews.

British Journal for the Philosophy of Science, 24(3):313– 316, September 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        24/3/313.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686237.

Koertge:1974:R

[Koe74] Noretta Koertge. Reviws. British Journal for the Philosophy of Science, 25(3):297–299, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/3/297.full.pdf+
html;   http://www.jstor.
org/stable/686914.

Koertge:1976:R

[Koe76]  Noretta Koertge.  Reviews.

British Journal for the Philosophy of Science, 27(4):405– 408, December 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/405.full.pdf+html;
http://www.jstor.org/
stable/686870.

Koertge:1978:R

[Koe78]  Noretta Koertge.  Reviews.

British Journal for the Philosophy of Science, 29(2):204–207, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/204.full.pdf+
html;   http://www.jstor.
org/stable/687132.

Koertge:1980:Ra

[Koe80a] Noretta Koertge. Reviews.

British Journal for the Philosophy of Science, 31(2):193–195, June 1980. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/2/193.full.pdf+
html;   http://www.jstor.
org/stable/687187.

Koertge:1980:Rb

[Koe80b] Noretta Koertge. Reviews. British Journal for the Philosophy of Science, 31(4):385– 390, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/4/385.full.pdf+html;
http://www.jstor.org/
stable/687065.

Koertge:1983:R

[Koe83]  Noretta Koertge.  Reviews.

British Journal for the Philosophy of Science, 34(2):185–188, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/185.full.pdf+
html;   http://www.jstor.
org/stable/687451.

Koertge:1992:EP

[Koe92] Noretta Koertge. Explanation and its problems. British Journal for the Philosophy of Science, 43(1):85–98, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/1/85.full.pdf+html;

http://www.jstor.org/
stable/687885.

Koertge:1999:BRB

[Koe99] N. Koertge. Book review: Belief and resistance: dynamics of contemporary intellectual controversy. Barbara Herrnstein Smith. British Journal for the Philosophy of Science, 50(3):508– 513, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/508.full.pdf+html;
http://www.jstor.org/
stable/40072252.

Koertge:2000:NAP

[Koe00] Noretta Koertge. ’New age’ philosophies of science: Constructivism, feminism and postmodernism. British Journal for the Philosophy of Science, 51(4): 667–683, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541612.

Koertge:2004:SWD

[Koe04]  Noretta Koertge.  The sci-

ence wars: Debating scientific knowledge and technology. British Journal for the Philosophy of Science, 55(1):185–188, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/1/185.full.pdf+
html;   http://www.jstor.
org/stable/3541843.

Koperski:2001:CBE

[Kop01] Jeffrey Koperski. Has chaos been explained? British Journal for the Philosophy of Science, 52(4):683–700, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/4/683.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541914.

Koperski:2005:SWC

[Kop05] Jeffrey Koperski. Should we care about fine-tuning? British Journal for the Philosophy of Science, 56(2):303–319, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/2/303.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541704.

Korner:1959:RU

[Kör59a] S. Körner. Review: [untitled]. British Journal for the Philosophy of Science, X(39): 257, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686069.

Korner:1959:Ra

[Kör59b] S. Körner. Reviews. British Journal for the Philosophy of Science, IX(36):337–339, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/36/337.full.pdf+
html;   http://www.jstor.
org/stable/685199.

Korner:1959:Rb

[Kör59c]  S. Körner.  Reviews.

British Journal for the Philosophy of Science, X(38):153–156, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/153.full.pdf+
html;   http://www.jstor.
org/stable/686006.

Korner:1959:Rc

[Kör59d] S. Körner. Reviews. British Journal for the Philosophy of Science, X(39):257a–257, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/257-a.full.pdf+html.

Korner:1964:DUI

[Kör64] S. Körner. Deductive unification and idealisation. British Journal for the Philosophy of Science, XIV(56):274–284, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/56/274.full.
pdf+html;   http://www.
jstor.org/stable/685634.

Korner:1967:SRB

[Kör67] S. Körner. Some relations between philosophical and scientific theories. British Journal for the Philosophy of Science, 17(4):265–278, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/265.full.pdf+html;
http://www.jstor.org/
stable/686769.

Korner:1968:RMK

[Kör68] S. Körner. Reply to Mr Kumar. British Journal for the Philosophy of Science, 18(4): 323–324, February 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/4/323.full.pdf+html;
http://www.jstor.org/
stable/686362.

Korner:1986:GER

[Kör86] S. Körner. Book review: Gellner, Ernest [1985]: Relativism and the Social Sciences. Cambridge University Press. £22.50. Pp. viii + 200 (ISBN 0-521-26530-4). British Journal for the Philosophy of Science, 37(3): 367–369, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687296.

Korb:1991:ES

[Kor91] Kevin B. Korb. Explaining science. British Journal for the Philosophy of Science, 42 (2):239–253, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/239.full.pdf+html;
http://www.jstor.org/
stable/688055.

Kosso:1988:DO

[Kos88] Peter Kosso. Dimensions of observability. British Journal for the Philosophy of Science, 39(4):449–467, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/449.full.pdf+html;
http://www.jstor.org/
stable/687432.

Kosso:2000:ESS

[Kos00] P. Kosso. The empirical status of symmetries in physics. British Journal for the Philosophy of Science, 51(1):81–98, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/1/81.full.pdf+
html;   http://www.jstor.
org/stable/3541749.

Koskinen:2021:DRA

[Kos21] Inkeri Koskinen. Defending a risk account of scientific objectivity. British Journal for the Philosophy of Science, 71(4):1187– 1207, ???? 2021. CODEN ????

ISSN 0007-0882 (print), 1464-3537 (electronic).

Kotzen:2012:SBL

[Kot12] Matthew Kotzen. Selection biases in likelihood arguments. British Journal for the Philosophy of Science, 63(4):825– 839, December 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/4/825.full.pdf+html.

Knobe:2006:PII

[KOV06] Joshua Knobe, Ken D. Olum, and Alexander Vilenkin. Philosophical implications of inflationary cosmology. British Journal for the Philosophy of Science, 57(1):47–67, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        57/1/47.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541652.

Kovaka:2019:UED

[Kov19] Karen Kovaka. Underdetermination and evidence in the developmental plasticity debate. British Journal for the Philosophy of Science, 70(1):127–152, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/1/127/4098115.

Kozuch:2014:BRE

[Koz14] Benjamin Kozuch. Book review: Elizabeth Irvine, Consciousness as a Scientific Concept: a Philosophy of Science Perspective. British Journal for the Philosophy of Science, 65(3):651–655, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/3/651.

Kowalski:1999:OBS

[KP99] T. Kowalski and T. Placek. Outcomes in branching space–time and GHZ–Bell theorems. British Journal for the Philosophy of Science, 50(3):349–375, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/3/349.full.pdf+
html;   http://www.jstor.
org/stable/40072236.

Krajewski:1978:RP

[Kra78] Wladyslaw Krajewski. Reduction in physics. British Journal for the Philosophy of Science, 29(3):280–283, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/280.full.pdf+html;
http://www.jstor.org/
stable/687019.

Kragh:2009:BRM

[Kra09] Helge Kragh. Book review:

Michela Massimi, Pauli’s Exclusion Principle: The Origin and Validation of a Scientific Principle. British Journal for the Philosophy of Science, 60 (1):235–238, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/235.full.pdf+html;
http://www.jstor.org/
stable/25591996.

Kreisel:1953:DMF

[Kre53a] G. Kreisel. The diagonal method in formalized arithmetic. British Journal for the Philosophy of Science, III(12):364–374, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/12/364.full.
pdf+html;   http://www.
jstor.org/stable/685457.

Kreisel:1953:VHT

[Kre53b] G. Kreisel. A variant to Hilbert’s theory of the foundations of arithmetic. British Journal for the Philosophy of Science, IV (14):107–129, August 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/14/107.full.pdf+html;
http://www.jstor.org/
stable/685987.

Kreisel:1956:SUM

[Kre56] G. Kreisel. Some uses of mathematics. British Journal for the Philosophy of Science, VII(26): 161–173, August 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/26/161.full.pdf+html;
http://www.jstor.org/
stable/685700.

Kreisel:1958:WRF

[Kre58] G. Kreisel. Wittgenstein’s remarks on the foundations of mathematics. British Journal for the Philosophy of Science, IX (34):135–158, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/34/135.full.pdf+html;
http://www.jstor.org/
stable/685151.

Kreisel:1960:BRW

[Kre60a]  G. Kreisel.  Book review:

Wittgenstein’s Theory and Practice of Philosophy. British Journal for the Philosophy of Science, XI(43):238–252, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/238.full.pdf+html.

Kreisel:1960:RWT

[Kre60b] G. Kreisel. Review: Wittgenstein’s theory and practice of philosophy. British Journal for the Philosophy of Science, XI (43):238–251, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/

685313. The Blue and Brown Books by Ludwig Wittgenstein. Remarks on the Foundations of Mathematics.

Krishna:1961:LCR

[Kri61] Daya Krishna. ‘Lying’ and the compleat robot. British Journal for the Philosophy of Science, XII(46):146–149, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/46/146.full.pdf+html;
http://www.jstor.org/
stable/685493.

Krips:1984:PPQ

[Kri84] Henry Krips. Popper, propensities, and quantum theory. British Journal for the Philosophy of Science, 35(3):253–274, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/3/253.full.pdf+
html;   http://www.jstor.
org/stable/687476.

Krips:1988:RCW

[Kri88]  Henry Krips.  Realism and

the collapse of the WavePacket. British Journal for the Philosophy of Science, 39(2):225–232, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/2/225.full.pdf+

html;   http://www.jstor.
org/stable/687268.

Krips:1992:R

[Kri92] Henry Krips. Review. British Journal for the Philosophy of Science, 43(1):137–139, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/1/137.full.pdf+
html;   http://www.jstor.
org/stable/687889.

Krickel:2018:RAM

[Kri18] Beate Krickel. A regularist approach to mechanistic typelevel explanation. British Journal for the Philosophy of Science, 69(4):1123–1153, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/1123/3980253.

Kroes:1989:SAB

[Kro89] Peter Kroes. Structural analogies between physical systems. British Journal for the Philosophy of Science, 40(2):145–154, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/145.full.pdf+
html;   http://www.jstor.
org/stable/687505.

Kronz:2002:PP

[Kro02] Frederick M. Kronz. The philosophy of physics. British Journal for the Philosophy of Science, 53(3):473–475, September 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/3/473.full.pdf+html;
http://www.jstor.org/
stable/3541893.

Kruse:1997:VAP

[Kru97] Michael Kruse. Variation and the accuracy of predictions. British Journal for the Philosophy of Science, 48(2):181–193, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/181.full.pdf+
html;   http://www.jstor.
org/stable/687744.

Koscholke:2019:CCC

[KS19] Jakob Koscholke and Michael Schippers. Coherence and common causes: Against relevancesensitive measures of coherence. British Journal for the Philosophy of Science, 70(3):771– 785, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
771/4682635.

Kamminga:1993:HUG

[KT93] Harmke Kamminga and Reza Tavakol. How untidy is God’s mind? a note on the dynamical implications of Nancy Cartwright’s metaphysics. British Journal for the Philosophy of Science, 44(3):549–553, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/549.full.pdf+html;
http://www.jstor.org/
stable/688023.

Kuhn:1962:SSR

[Kuh62] Thomas S. Kuhn. The structure of scientific revolutions. University of Chicago Press, Chicago, IL, USA, 1962. xv + 172 pp. LCCN Q175 .K95.

Kuhn:1966:R

[Kuh66] Thomas S. Kuhn. Reviews.

British Journal for the Philosophy of Science, 17(3):256– 258, November 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/3/256.full.pdf+html;
http://www.jstor.org/
stable/686764.

Kuhn:1967:R

[Kuh67] Thomas S. Kuhn. Reviews.

British Journal for the Philosophy of Science, 18(2):148–154, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/148.full.pdf+
html;   http://www.jstor.
org/stable/686583.

Kuhn:1980:HBP

[Kuh80] Thomas S. Kuhn. The halt and the blind: Philosophy and history of science. British Journal for the Philosophy of Science, 31(2):181–192, June 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/2/181.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687186.

Kuipers:1983:NIE

[Kui83] Theo A. F. Kuipers. Noninductive explication of two inductive intuitions. British Journal for the Philosophy of Science, 34(3):209–223, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        34/3/209.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687320.

Kuklick:1987:R

[Kuk87] Bruce Kuklick. Reviews. British Journal for the Philosophy of Science, 38(1):117–119, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/1/117.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687086.

Kukla:1992:CMR

[Kuk92] Rebecca Kukla. Cognitive models and representation. British Journal for the Philosophy of Science, 43(2):219–232, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/219.full.pdf+
html;   http://www.jstor.
org/stable/687873.

Kukla:1994:SRS

[Kuk94] André Kukla. Scientific realism, scientific practice, and the natural ontological attitude. British Journal for the Philosophy of Science, 45(4):955–975, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/4/955.full.pdf+
html;   http://www.jstor.
org/stable/687616.

Kukla:1995:FSC

[Kuk95] André Kukla. Forster and Sober on the curve-fitting problem. British Journal for the Philosophy of Science, 46(2):248–252, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/248.full.pdf+
html;   http://www.jstor.
org/stable/687815.

Kukla:1997:R

[Kuk97] André Kukla. Review. British Journal for the Philosophy of Science, 48(3):447–454, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/3/447.full.pdf+
html;   http://www.jstor.
org/stable/688075.

Kukla:2008:OWO

[Kuk08] André Kukla. The one world, one science argument. British Journal for the Philosophy of Science, 59(1):73–88, March 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/1/73.full.pdf+html;
http://www.jstor.org/
stable/30116994.

Kulka:1977:SPC

[Kul77] Tomas Kulka. Some problems concerning rational reconstruction: Comments on Elkana and Lakatos. British Journal for the Philosophy of Science, 28(4): 325–344, December 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/4/325.full.pdf+html;
http://www.jstor.org/
stable/686877.

Kumar:1967:LIP

[Kum67] Dharmendra Kumar. Logic and inexact predicates. British Journal for the Philosophy of Science, 18(3):211–222, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/3/211.full.pdf+html;

http://www.jstor.org/
stable/686591.

Kumar:1969:NCU

[Kum69] Dharmendra Kumar. Neutrality, contingency and undecidability. British Journal for the Philosophy of Science, 19(4): 353–356, February 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/4/353.full.pdf+html;
http://www.jstor.org/
stable/686294.

Kuorikoski:2024:CEI

[Kuo24] Jaakko Kuorikoski. Contrastive evidence and inductive risk. British Journal for the Philosophy of Science, 75(1):61–76, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Kusch:2001:BRS

[Kus01] Martin Kusch. Book review:

Susan Haack, Manifesto of a Passionate Moderate: Unfashionable Essays. British Journal for the Philosophy of Science, 52 (1):169–173, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/169.full.pdf+html;
http://www.jstor.org/
stable/3541952.

Kutach:1998:R

[Kut98] Douglas Kutach. Review. British Journal for the Philosophy of Science, 49(4):649–651, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/4/649.full.pdf+
html;   http://www.jstor.
org/stable/688137.

Kutach:2000:BRR

[Kut00]  Douglas Kutach.  Book re-

view: Richard A. Healey and Geoffrey Hellman ( eds ) — Quantum Measurement: Beyond Paradox. British Journal for the Philosophy of Science, 51(4): 947–951, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/947.full.pdf+html;
http://www.jstor.org/
stable/3541742.

Kuykendall:2024:DAP

[Kuy24] Davis Kuykendall. In defence of the agent and patient distinction: The case from molecular biology and chemistry. British Journal for the Philosophy of Science, 75 (2):443–463, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Korb:1997:SPS

[KW97] Kevin B. Korb and Chris S. Wallace. In search of the Philosopher’s Stone: Remarks on Humphreys and Freedman’s critique of causal discovery. British Journal for the Philosophy of Science, 48(4):543–553, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/543.full.pdf+
html;   http://www.
jstor.org/stable/688099. See

[HF96, SGS97].

Kyburg:1958:RBB

[Kyb58] Henry E. Kyburg, Jr. R. B.

Braithwaite on probability and induction. British Journal for the Philosophy of Science, IX (35):203–220, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/203.full.pdf+html;
http://www.jstor.org/
stable/685650.

Kyberg:1974:PP

[Kyb74] Henry E. Kyberg. Propensities and probabilities. British Journal for the Philosophy of Science, 25(4):358–375, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/4/358.full.pdf+
html;   http://www.jstor.
org/stable/686466.

Kummerfeld:2016:ECC

[KZ16] Erich Kummerfeld and Kevin J. S. Zollman. Editor’s choice: Conservatism and the scientific state of nature. British Journal for the Philosophy of Science, 67(4):1057–1076, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/4/1057.

Karakostas:2022:SFS

[KZH^{+}22] Vassilios Karakostas, Elias Zafiris, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Marc Lange, Matteo Colombo, Peggy Seriès, Jeffrey A. Barrett, and Brian Skyrms. On the structure and function of scientific perspectivism in categorical quantum mechanics. British Journal for the Philosophy of Science, 73(3):811–848, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lacey:1969:R

[Lac69] Hugh M. Lacey. Reviews. British Journal for the Philosophy of Science, 20(1):88–89, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/1/88.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686724.

Lacey:1970:SIA

[Lac70] Hugh M. Lacey. The scientific intelligibility of absolute space: a study of Newtonian argument. British Journal for the Philosophy of Science, 21(4): 317–342, November 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/317.full.pdf+html;
http://www.jstor.org/
stable/686555.

LaCroix:2022:ULE

[LaC22] Travis LaCroix. Using logic to evolve more logic: Composing logical operators via selfassembly. British Journal for the Philosophy of Science, 73(2): 407–437, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lad:1984:CQ

[Lad84] Frank Lad. The calibration question. British Journal for the Philosophy of Science, 35(3): 213–221, September 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/213.full.pdf+html;
http://www.jstor.org/
stable/687473.

Ladyman:1996:R

[Lad96] James Ladyman. Review. British Journal for the Philosophy of Science, 47(3):487–490, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/3/487.full.pdf+
html;   http://www.jstor.
org/stable/687995.

Ladyman:1999:BRB

[Lad99] J. Ladyman. Book review: A novel defense of scientific realism. Jarrett Leplin. British Journal for the Philosophy of Science, 50(1):181–188, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/181.full.pdf+html;
http://www.jstor.org/
stable/40072217.

Ladyman:2000:WRW

[Lad00] James Ladyman. What’s really wrong with constructive empiricism? Van Fraassen and the metaphysics of modality. British Journal for the Philosophy of Science, 51(4):837–856, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/4/837.full.pdf+
html;   http://www.jstor.
org/stable/3541729.

Ladyman:2004:CEM

[Lad04]  James Ladyman.  Construc-

tive empiricism and modal metaphysics: a reply to Monton and van Fraassen. British Journal for the Philosophy of Science, 55(4):755–765, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/755.full.pdf+html;
http://www.jstor.org/
stable/3541628. See [Han04].

Lai:1984:PRT

[Lai84] Tyrone Lai. The philosophical relevance of ‘technically good’ experiments. British Journal for the Philosophy of Science, 35 (2):156–159, June 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/2/156.full.pdf+html;
http://www.jstor.org/
stable/686965.

Lai:1988:ETO

[Lai88] Tyrone Lai. Empirical tests are only auxiliary devices. British Journal for the Philosophy of Science, 39(2):211–223, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/2/211.full.pdf+
html;   http://www.jstor.
org/stable/687267.

Lai:1991:DPI

[Lai91] Tyrone Lai. Discovery as a problem for the inventor. British Journal for the Philosophy of Science, 42(3):327–337, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/3/327.full.pdf+
html;   http://www.jstor.
org/stable/687911.

Laimann:2021:CK

[Lai21] Jessica Laimann. Capricious kinds. British Journal for the Philosophy of Science, 71(3): 1043–1068, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lakatos:1963:PR

[Lak63a] I. Lakatos. Proofs and refutations (I). British Journal for the Philosophy of Science, XIV(53):1–25, May 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/53/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685347.

Lakatos:1963:PRIa

[Lak63b] I. Lakatos. Proofs and refutations (II). British Journal for the Philosophy of Science, XIV(54):120–139, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/54/120.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685430.

Lakatos:1963:PRIb

[Lak63c] I. Lakatos. Proofs and refutations (III). British Journal for the Philosophy of Science, XIV (55):221–245, November 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/55/221.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685242.

Lakatos:1964:PRI

[Lak64] I. Lakatos. Proofs and refutations (IV). British Journal for the Philosophy of Science, XIV(56):296–342, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/56/296.full.
pdf+html;   http://www.
jstor.org/stable/685636.

Lakatos:1976:RER

[Lak76] Imre Lakatos. A renaissance of empiricism in the recent philosophy of mathematics. British Journal for the Philosophy of Science, 27(3):201–223, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/201.full.pdf+
html;   http://www.jstor.
org/stable/686119.

Lamont:1962:R

[Lam62] Corliss Lamont. Reviews. British Journal for the Philosophy of Science, XIII(51):244–248, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/244.full.
pdf+html;   http://www.
jstor.org/stable/685842.

Landsberg:1951:R

[Lan51] P. T. Landsberg. Reviews.

British Journal for the Philosophy of Science, II(7):255– 256, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/255.full.pdf+html;
http://www.jstor.org/
stable/685298.

Lande:1956:LQ

[Lan56] Alfred Lande. The logic of quanta. British Journal for the Philosophy of Science, VI(24): 300–320, February 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/24/300.full.pdf+html;
http://www.jstor.org/
stable/685411.

Lande:1957:R

[Lan57] Alfred Landé. Reviews. British Journal for the Philosophy of Science, VII(28):357–359, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/28/357.full.
pdf+html;   http://www.
jstor.org/stable/685121.

Lande:1959:DUQ

[Lan59] Alfred Landé. From dualism to unity in quantum mechanics. British Journal for the Philosophy of Science, X(37):16–24, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/37/16.full.pdf+
html;   http://www.jstor.
org/stable/685772.

Lande:1961:R

[Lan61] Alfred Landé. Reviews. British Journal for the Philosophy of Science, XII(47):259–260, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/47/259.full.
pdf+html;   http://www.
jstor.org/stable/685215.

Lande:1965:WDQ

[Lan65] Alfred Landé. Why do quantum theorists ignore the quantum theory? British Journal for the Philosophy of Science, XV(60):307–313, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/60/307.full.pdf+html;
http://www.jstor.org/
stable/686537.

Lande:1976:LBQ

[Lan76] Alfred Landé. The laws behind the quantum laws. British Journal for the Philosophy of Science, 27(1):43–50, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/1/43.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/686377.

Langford:1983:R

[Lan83]  Glenn Langford.  Reviews.

British Journal for the Philosophy of Science, 34(2):188–190, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/2/188.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687452.

Lange:1995:SP

[Lan95] Marc Lange. Spearman’s Principle. British Journal for the Philosophy of Science, 46(4): 503–521, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/4/503.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687895.

Landini:1998:RHS

[Lan98] Gregory Landini. Russell’s hidden substitutional theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-511683-6. xi + 337 pp. LCCN B1649.R94 L36 1998; B1649.R94 L36 1998eb; B1649.R94.

Lange:2001:ASP

[Lan01] Marc Lange. The apparent superiority of prediction to accommodation as a side effect: a reply to Maher. British Journal for the Philosophy of Science, 52(3):575–588, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/575.full.pdf+html;
http://www.jstor.org/
stable/3541929.

Lange:2006:DCR

[Lan06] Marc Lange. Do chances receive equal treatment under the laws? or: Must chances be probabilities? British Journal for the Philosophy of Science, 57(2):383–403, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/383.full.pdf+html;
http://www.jstor.org/
stable/3873472.

Lange:2013:WMS

[Lan13] Marc Lange. What makes a scientific explanation distinctively mathematical? British Journal for the Philosophy of Science, 64(3):485–511, September 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/485.full.pdf+html.

LaPorte:2003:R

[LaP03] Joseph LaPorte. Review. British Journal for the Philosophy of Science, 54(4):627–630, December 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/54/4/627.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541685.

Larmore:1981:R

[Lar81]  Charles Larmore.  Reviews.

British Journal for the Philosophy of Science, 32(3):318– 325, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/3/318.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687599.

Laraudogoitia:1997:CPD

[Lar97] Jon Pérez Laraudogoitia. Classical particle dynamics, indeterminism and a supertask. British Journal for the Philosophy of Science, 48(1):49–54, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/1/49.full.pdf+
                                        html;   http://www.
                                        jstor.org/stable/687716. See

comments [EN98].

Laraudogoitia:1999:DEN

[Lar99] J. P. Laraudogoitia. Discussion. Earman and Norton on supertasks that generate indeterminism. British Journal for the Philosophy of Science, 50 (1):137–141, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/137.full.pdf+html;
http://www.jstor.org/
stable/40072211.

Larvor:2000:BRI

[Lar00]  Brendan Larvor.  Book re-

view: Imre Lakatos and Paul Feyerabend — For and Against Method: Including Lakatos’s Lectures on Scientific Method and the Lakatos–Feyerabend Correspondence. British Journal for the Philosophy of Science, 51(4):919–922, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/919.full.pdf+html;
http://www.jstor.org/
stable/3541737.

Laraudogoitia:2003:TSE

[Lar03] Jon Pérez Laraudogoitia. Taking self-excitations seriously: on Angel’s initial condition. British Journal for the Philosophy of Science, 54(2):319–326, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/319.full.pdf+
html;   http://www.jstor.
org/stable/3541971.

Laraudogoitia:2004:SPE

[Lar04] Jon Pérez Laraudogoitia. On a (supposedly) plausible extension of Newtonian collision dynamics. British Journal for the Philosophy of Science, 55(2):365–370, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/2/365.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541696.

Laraudogoitia:2005:IFC

[Lar05] Jon Pérez Laraudogoitia. An interesting fallacy concerning dynamical supertasks. British Journal for the Philosophy of Science, 56(2):321–334, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/56/2/321.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541705.

Latsis:1972:SDE

[Lat72] Spiro J. Latsis. Situational determinism in economics. British Journal for the Philosophy of Science, 23(3):207–245, August 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/23/3/207.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686686.

Latsis:1976:LSE

[Lat76] Spiro J. Latsis. The limitations of single-exit models: Reply to Machlup. British Journal for the Philosophy of Science, 27(1):51–60, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/1/51.full.pdf+html;
http://www.jstor.org/
stable/686378.

Laudan:1966:RU

[Lau66] Laurens Laudan. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):76–80, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686408.

Laudan:1967:R

[Lau67] L. Laudan. Reviews. British Journal for the Philosophy of Science, 18(2):154–157, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/154.full.pdf+
html;   http://www.jstor.
org/stable/686584.

Laudan:1968:R

[Lau68] L. Laudan. Reviews. British Journal for the Philosophy of Science, 19(3):264–265, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/264.full.pdf+
html;   http://www.jstor.
org/stable/686249.

Laudan:1969:R

[Lau69] L. Laudan. Reviews. British Journal for the Philosophy of Science, 20(2):180–181, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/180.full.pdf+
html;   http://www.jstor.
org/stable/686189.

Laudan:1970:R

[Lau70]  Laurens Laudan.  Reviews.

British Journal for the Philosophy of Science, 21(3):311–312, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/311.full.pdf+
html;   http://www.jstor.
org/stable/686212.

Laudan:1989:IIA

[Lau89] Larry Laudan. If it ain’t broke, don’t fix it. British Journal for the Philosophy of Science, 40(3):369–375, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/369.full.pdf+html;
http://www.jstor.org/
stable/687784.

Laudisa:1995:EBN

[Lau95]  Federico Laudisa.  Einstein,

Bell, and nonseparable realism. British Journal for the Philosophy of Science, 46(3):309– 329, September 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
46/3/309.full.pdf+html;
http://www.jstor.org/
stable/687659.

Lavis:1977:RSM

[Lav77] David Lavis. The role of statistical mechanics in classical physics. British Journal for the Philosophy of Science, 28(3): 255–279, September 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/3/255.full.pdf+html;
http://www.jstor.org/
stable/686809.

Lavine:1995:R

[Lav95]  Shaughan Lavine.  Review.

British Journal for the Philosophy of Science, 46(2):267–274, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/267.full.pdf+
html;   http://www.jstor.
org/stable/687818.

Lavelle:2023:WCB

[Lav23] Jane Suilin Lavelle. When a crisis becomes an opportunity: The role of replications in making better theories. British Journal for the Philosophy of Science, 73 (4):965–986, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lawson:1985:CPP

[Law85] Tony Lawson. The context of prediction (and the paradox of confirmation). British Journal for the Philosophy of Science, 36(4):393–407, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/4/393.full.pdf+html;
http://www.jstor.org/
stable/687399.

Lawrence:1994:R

[Law94] Sklar Lawrence. Review. British Journal for the Philosophy of Science, 45(3):933–934, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/933.full.pdf+
html;   http://www.jstor.
org/stable/687806.

Lazerowitz:1958:R

[Laz58] Morris Lazerowitz. Reviews. British Journal for the Philosophy of Science, IX(33):68–70, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/68.full.pdf+
html;   http://www.jstor.
org/stable/686107.

Linquist:2013:TMA

[LB13] Stefan Linquist and Jordan Bartol. Two myths about somatic markers. British Journal for the Philosophy of Science, 64(3):455–484, September 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/455.full.pdf+html.

Lu:2018:EGE

[LB18] Qiaoying Lu and Pierrick Bourrat. The evolutionary gene and the extended evolutionary synthesis. British Journal for the Philosophy of Science, 69(3):775–800, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/775/3744978.

Levy:2015:ECM

[LC15] Arnon Levy and Adrian Currie. Editor’s choice: Model organisms are not (theoretical) models. British Journal for the Philosophy of Science, 66 (2):327–348, June 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/2/327.

LeVay:1952:HSU

[Le 52] David Le Vay. Hans Selye and a unitary conception of disease. British Journal for the Philosophy of Science, III(10):157–168, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/157.full.
pdf+html;   http://www.
jstor.org/stable/685554.

LePoidevin:1998:R

[Le 98] Robin Le Poidevin. Review. British Journal for the Philosophy of Science, 49(2):365–369, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/2/365.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688126.

LePoidevin:2003:TFD

[Le 03] Robin Le Poidevin. Travels in four dimensions: the enigmas of space and time. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-875254-7 (hardcover), 0-19-875255-5 (paperback). xviii + 275 pp. LCCN BD632 .L46 2003.

LePoidevin:2005:MEM

[Le 05] Robin Le Poidevin. Missing elements and missing premises: a combinatorial argument for the ontological reduction of chemistry. British Journal for the Philosophy of Science, 56(1): 117–134, March 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/1/117.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541852.

LePoidevin:2007:ITE

[Le 07] Robin Le Poidevin. The images of time: an essay on temporal representation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-957551-7, 0-19-926589-5. xiv + 193 pp. LCCN BD638 .L33 2007.

Landsberg:1970:FWM

[LE70] P. T. Landsberg and D. A.

Evans. Free will in a mechanistic universe? British Journal for the Philosophy of Science, 21(4):343–358, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/343.full.pdf+html.

Livengood:2022:CCH

[LE22] Jonathan Livengood and Adam Edwards. Calibrating chromatography: How Tswett broke the experimenters’ regress. British Journal for the Philosophy of Science, 73(3):685–710, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Leach:1968:EVN

[Lea68] James Leach. Explanation and value neutrality. British Journal for the Philosophy of Science, 19(2):93–108, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/2/93.full.pdf+html;
http://www.jstor.org/
stable/686789.

Leach:1969:R

[Lea69] Edmund Leach. Reviews. British Journal for the Philosophy of Science, 20(2):175–176, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/175.full.pdf+
html;   http://www.jstor.
org/stable/686186.

Leblanc:1960:CDC

[Leb60] Hugues Leblanc. On so-called degrees of confirmation. British Journal for the Philosophy of Science, X(40):312–315, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/40/312.full.pdf+
html;   http://www.jstor.
org/stable/685918.

Leblanc:1989:APT

[Leb89] Hugues Leblanc. The autonomy of probability theory (notes on Kolmogorov, Rényi, and Popper). British Journal for the Philosophy of Science, 40 (2):167–181, June 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/167.full.pdf+html;
http://www.jstor.org/
stable/687507.

Lederman:1987:GA

[Led87] Muriel Lederman. ‘Genes’ amplified. British Journal for the Philosophy of Science, 38(4): 561–566, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/561.full.pdf+html;
http://www.jstor.org/
stable/687360.

Lees:1982:TMS

[Lee82] R. B. Lees. Trying to make sense.

British Journal for the Philosophy of Science, 33(2):194–208, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/2/194.full.pdf+
html;   http://www.jstor.
org/stable/687279.

Leeds:1994:PWF

[Lee94]  Stephen Leeds.  Price on

the Wheeler–Feynman theory. British Journal for the Philosophy of Science, 45(1):288–294, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/288.full.pdf+
html;   http://www.jstor.
org/stable/687973.

Leeds:1995:WFA

[Lee95] Stephen Leeds. Wheeler–Feynman again: a reply to Price. British Journal for the Philosophy of Science, 46(3):381–383, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/3/381.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687664.

Lee:2021:MWF

[Lee21] Jonny Lee. Mechanisms, wide functions, and content: Towards a computational pluralism. British Journal for the Philosophy of Science, 72(1):221– 244, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lehman:1965:TEB

[Leh65] Hugh S. Lehman. Teleological explanation in biology. British Journal for the Philosophy of Science, XV(60):327, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        XV/60/327-a.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/686539.

Lehrer:1970:IRC

[Leh70] Keith Lehrer. Induction, reason and consistency. British Journal for the Philosophy of Science, 21(1):103–114, February 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/1/103.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686474.

Lehmkuhl:2011:MEM

[Leh11] Dennis Lehmkuhl. Mass–energy– momentum: Only there because of spacetime? British Journal for the Philosophy of Science, 62(3):453–488, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/453.full.pdf+html.

Leitgeb:2013:CIS

[Lei13] Hannes Leitgeb. Criteria of identity: Strong and wrong. British Journal for the Philosophy of Science, 64(1):61–68, March 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/61.full.pdf+html.

Lejewski:1954:LE

[Lej54] Czesław Lejewski. Logic and existence. British Journal for the Philosophy of Science, V(18): 104–119, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/104.full.pdf+html;
http://www.jstor.org/
stable/685168.

Lejewski:1955:R

[Lej55]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, V(20):343– 344, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/343.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685740.

Lejewski:1956:R

[Lej56]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, VI(24):345– 346, February 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/24/345.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685419.

Lejewski:1957:R

[Lej57]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, VIII(30): 172–173, August 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/30/172.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685800.

Lejewski:1958:Rb

[Lej58a] C. Lejewski. Reviews. British Journal for the Philosophy of Science, IX(35):249–251, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/35/249.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685659.

Lejewski:1958:Ra

[Lej58b]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, VIII(32):349– 350, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/349.full.pdf+html;
http://www.jstor.org/
stable/685867.

Lejewski:1959:R

[Lej59]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, IX(36):331– 333, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/331.full.pdf+html;
http://www.jstor.org/
stable/685196.

Lejewski:1963:R

[Lej63]  Czesław Lejewski.  Reviews.

British Journal for the Philosophy of Science, XIII(52):323– 326, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/323.full.pdf+html;
http://www.jstor.org/
stable/685334.

Lelas:1993:ST

[Lel93] Srdjan Lelas. Science as technology. British Journal for the Philosophy of Science, 44(3):

423–442, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/423.full.pdf+html;
http://www.jstor.org/
stable/688014.

Lenhard:2006:MSI

[Len06]  Johannes Lenhard.  Models

and statistical inference: the controversy between Fisher and Neyman–Pearson. British Journal for the Philosophy of Science, 57(1):69–91, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/69.full.pdf+html;
http://www.jstor.org/
stable/3541653.

Leng:2010:MR

[Len10] Mary Leng. Mathematics and reality. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-928079-7. x + 278 pp. LCCN QA8.4 .L446 2010.

Leplin:1969:MVC

[Lep69] Jarrett Leplin. Meaning variance and the comparability of theories. British Journal for the Philosophy of Science, 20(1):69–75, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/69.full.pdf+
html;   http://www.jstor.
org/stable/686719.

Leplin:1976:SRT

[Lep76] Jarrett Leplin. Synchronisation rules and transitivity. British Journal for the Philosophy of Science, 27(4):399–402, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/4/399.full.pdf+
html;   http://www.jstor.
org/stable/686868.

Leplin:1982:AAH

[Lep82] Jarrett Leplin. The assessment of auxiliary hypotheses. British Journal for the Philosophy of Science, 33(3):235–250, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/3/235.full.pdf+
html;   http://www.jstor.
org/stable/687223.

Leplin:1987:R

[Lep87] Jarrett Leplin. Reviews. British Journal for the Philosophy of Science, 38(4):575–577, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/4/575.full.pdf+
html;   http://www.jstor.
org/stable/687364.

Leuridan:2012:TPM

[Leu12] Bert Leuridan. Three problems for the mutual manipulability account of constitutive relevance in mechanisms. British Journal for the Philosophy of Science, 63(2):399–427, June 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/399.full.pdf+html.

Leuridan:2014:SST

[Leu14] Bert Leuridan. The structure of scientific theories, explanation, and unification. a causalstructural account. British Journal for the Philosophy of Science, 65(4):717–771, December 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/717.

Levi:1963:CRA

[Lev63] I. Levi. Corroboration and rules of acceptance. British Journal for the Philosophy of Science, XIII(52):307–313, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/307.full.
pdf+html;   http://www.
jstor.org/stable/685328.

Levinson:1965:OCR

[Lev65]  H. B. Levinson.  The ori-

gin and concept of Relativity. British Journal for the Philosophy of Science, XVI(63):246– 248, November 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/63/246.full.pdf+html;
http://www.jstor.org/
stable/686308.

Levi:1967:PK

[Lev67] Isaac Levi. Probability kinematics. British Journal for the Philosophy of Science, 18(3): 197–209, November 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/3/197.full.pdf+html;
http://www.jstor.org/
stable/686590.

Levi:1968:Ra

[Lev68a] Isaac Levi. Reviews. British Journal for the Philosophy of Science, 19(1):73–81, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/1/73.full.pdf+html;
http://www.jstor.org/
stable/686852.

Levi:1968:Rb

[Lev68b] Isaac Levi. Reviews. British Journal for the Philosophy of Science, 19(3):259–261, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/259.full.pdf+
html;   http://www.jstor.
org/stable/686247.

Levi:1969:IJO

[Lev69] Isaac Levi. If Jones only knew more! British Journal for the Philosophy of Science, 20 (2):153–159, August 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/2/153.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686177.

Levi:1979:SSJ

[Lev79] Isaac Levi. Support and surprise: L. J. Cohen’s view of inductive probability. British Journal for the Philosophy of Science, 30(3):279–292, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/3/279.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687142.

Levi:1985:IAU

[Lev85] Isaac Levi. Illusions about uncertainty. British Journal for the Philosophy of Science, 36(3): 331–340, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/3/331.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687578.

Levin:1997:RAC

[Lev97a]  Janet Levin.  Review arti-

cle: Consciousness disputed.

British Journal for the Philosophy of Science, 48(1):91–107, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/91.full.pdf+
html;   http://www.jstor.
org/stable/687720.

Levin:1997:PRC

[Lev97b] Michael Levin. Putnam on reference and constructive sets. British Journal for the Philosophy of Science, 48(1):55–67, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/55.full.pdf+
html;   http://www.jstor.
org/stable/687717.

Levine:2001:BRB

[Lev01] Joseph Levine. Book review:

Matters of Mind: Consciousness, Reason, and Nature, Scott Sturgeon. British Journal for the Philosophy of Science, 52(3): 629–634, September 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/629.full.pdf+html;
http://www.jstor.org/
stable/3541934.

Levy:2014:WWH

[Lev14] Arnon Levy. What was Hodgkin and Huxley’s achievement? British Journal for the Philosophy of Science, 65(3):469–492, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/65/3/469.

Lewis:1997:QMO

[Lew97]  Peter J. Lewis.  Quantum

mechanics, orthogonality, and counting. British Journal for the Philosophy of Science, 48(3): 313–328, September 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        48/3/313.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688065.

Lewens:2001:SSR

[Lew01] Tim Lewens. Sex and selection: a reply to Matthen. British Journal for the Philosophy of Science, 52(3):589–598, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/52/3/589.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541930.

Lewis:2003:CMR

[Lew03]  Peter J. Lewis.  Counting

marbles: a reply to critics. British Journal for the Philosophy of Science, 54(1):165–170, March 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/

content/54/1/165.full.pdf+
html;   http://www.jstor.
org/stable/3541883.

Lewis:2004:LCS

[Lew04] Peter J. Lewis. Life in configuration space. British Journal for the Philosophy of Science, 55(4):713–729, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/713.full.pdf+html;
http://www.jstor.org/
stable/3541624.

Lewens:2005:RSP

[Lew05] Tim Lewens. Realism and the strong program. British Journal for the Philosophy of Science, 56(3):559–577, September 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/559.full.pdf+html;
http://www.jstor.org/
stable/3541719.

Lewis:2006:CTQ

[Lew06] Peter J. Lewis. Conspiracy theories of quantum mechanics. British Journal for the Philosophy of Science, 57(2):359–381, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/2/359.full.pdf+
html;   http://www.jstor.
org/stable/3873471.

Lewis:2007:EWB

[Lew07] Peter J. Lewis. Empty waves in Bohmian quantum mechanics. British Journal for the Philosophy of Science, 58(4): 787–803, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/4/787.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115174.

Lewens:2010:NS

[Lew10] Tim Lewens. The natures of selection. British Journal for the Philosophy of Science, 61 (2):313–333, June 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/2/313.full.pdf+html.

Lewens:2014:KSE

[Lew14] Tim Lewens. Book review: Kim Sterelny, The Evolved Apprentice: How Evolution Made Humans Unique. British Journal for the Philosophy of Science, 65 (1):185–189, January 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/1/185.full.pdf+html.

Lieb:1967:R

[Lie67] Irwin C. Lieb. Reviews. British Journal for the Philosophy of Science, 18(3):241–250, November 1967. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/241.full.pdf+
html;   http://www.jstor.
org/stable/686597.

Liggins:2014:AEC

[Lig14] David Liggins. Abstract expressionism and the communication problem. British Journal for the Philosophy of Science, 65(3):599–620, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/3/599.

Lillehammer:1999:BRM

[Lil99] H. Lillehammer. Book review:

from metaphysics to ethics: a defence of conceptual analysis. British Journal for the Philosophy of Science, 50(1):169–173, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/1/169.full.pdf+
html;   http://www.jstor.
org/stable/40072215.

Lindley:1984:R

[Lin84] D. V. Lindley. Reviews. British Journal for the Philosophy of Science, 35(3):293–296, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/3/293.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687479.

Lind:1993:NFT

[Lin93] Hans Lind. A note on fundamental theory and idealizations in economics and physics. British Journal for the Philosophy of Science, 44(3):493–503, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/3/493.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688018.

Linford:2023:KCA

[Lin23] Dan Linford. The Kalām cosmological argument meets the Mentaculus. British Journal for the Philosophy of Science, 74(1):91– 115, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lipton:1982:NR

[Lip82] Peter Lipton. Nagel revisited. British Journal for the Philosophy of Science, 33(2):186–194, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/186.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687278.

Lipton:1993:R

[Lip93] Peter Lipton. Review. British Journal for the Philosophy of Science, 44(3):585–588, September 1993. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/3/585.full.pdf+
html;   http://www.jstor.
org/stable/688029.

Lipton:1994:R

[Lip94] Peter Lipton. Review. British Journal for the Philosophy of Science, 45(3):929–932, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/929.full.pdf+
html;   http://www.jstor.
org/stable/687805.

Lipton:1999:BRB

[Lip99] P. Lipton. Book review: Philosophy of science. Alexander Bird. British Journal for the Philosophy of Science, 50(1):163–167, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/1/163.full.pdf+
html;   http://www.jstor.
org/stable/40072214.

Lipton:2002:HPI

[Lip02] Peter Lipton. Hume’s problem: Induction and the justification of belief. British Journal for the Philosophy of Science, 53(4):579–583, December 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/4/579.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/3541826.

Lipton:2004:IBE

[Lip04] Peter Lipton. Inference to the best explanation. International library of philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, second edition, 2004. ISBN 0-203-47085-0 (e-book), 0-415-24202-9 (hardcover), 0-415-24203-7 (paperback). xii + 219 pp. LCCN Q175.LIP 2004. URL http://

                                        oxford.etailer.dpsl.net/
                                        home/html/moreinfo.asp?
                                        isbn=0203470850.

List:2004:SAM

[Lis04] Christian List. On the significance of the absolute margin. British Journal for the Philosophy of Science, 55(3):521– 544, September 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/3/521.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541673.

Liu:1994:ABE

[Liu94] Chuang Liu. The Aharonow– Bohm effect and the reality of wave packets. British Journal for the Philosophy of Science, 45(4):977–1000, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/4/977.full.pdf+html;

http://www.jstor.org/
stable/687617.

Liu:2022:IPI

[Liu22] Yang Liu. Incomplete preference and indeterminate comparative probabilities. British Journal for the Philosophy of Science, 73(3):795–810, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lizza:1993:MPP

[Liz93] John P. Lizza. Multiple personality and personal identity revisited. British Journal for the Philosophy of Science, 44 (2):263–274, June 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/2/263.full.pdf+html;
http://www.jstor.org/
stable/687648.

Lloyd:1966:RU

[Llo66a] G. E. R. Lloyd. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):87–88, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686412.

Lloyd:1966:R

[Llo66b] G. E. R. Lloyd. Reviews. British Journal for the Philosophy of Science, 17(3):263–264, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/3/263.full.pdf+
html;   http://www.jstor.
org/stable/686767.

Lloyd:1967:PVK

[Llo67] G. E. R. Lloyd. Popper versus Kirk: a controversy in the interpretation of Greek science. British Journal for the Philosophy of Science, 18(1):21–38, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/21.full.pdf+
html;   http://www.jstor.
org/stable/686887.

Lloyd:1976:R

[Llo76] G. E. R. Lloyd. Reviews. British Journal for the Philosophy of Science, 27(4):408–411, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/4/408.full.pdf+
html;   http://www.jstor.
org/stable/686871.

Lavis:1985:WJP

[LM85] D. A. Lavis and P. J. Milligan. The work of E. T. Jaynes on probability, statistics and statistical physics. British Journal for the Philosophy of Science, 36(2):193–210, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/2/193.full.pdf+html;

http://www.jstor.org/
stable/687036.

Laurence:1999:BRB

[LM99] S. Laurence and E. Margolis.

Book review: Concepts: where cognitive science went wrong. Jerry A. Fodor. British Journal for the Philosophy of Science, 50(3):487–491, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/487.full.pdf+html;
http://www.jstor.org/
stable/40072247.

Laurence:2001:PSA

[LM01] Stephen Laurence and Eric Margolis. The poverty of the stimulus argument. British Journal for the Philosophy of Science, 52(2):217–276, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/217.full.pdf+html;
http://www.jstor.org/
stable/3541896.

Loaiza:2024:FE

[Loa24] Juan R. Loaiza. Functionalism and the emotions. British Journal for the Philosophy of Science, 75(1):233–251, March 2024. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lockwood:1996:MMIa

[Loc96a]  Michael Lockwood.  ‘Many

minds’ interpretations of quantum mechanics. British Journal for the Philosophy of Science, 47(2):159–188, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/2/159.full.pdf+html;
http://www.jstor.org/
stable/687940.

Lockwood:1996:MMIb

[Loc96b]  Michael Lockwood.  ‘Many

minds’ interpretations of quantum mechanics: Replies to replies. British Journal for the Philosophy of Science, 47(3): 445–461, September 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/3/445.full.pdf+html;
http://www.jstor.org/
stable/687988.

Lockie:2003:TAA

[Loc03] Robert Lockie. Transcendental arguments against eliminativism. British Journal for the Philosophy of Science, 54(4): 569–589, December 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/4/569.full.pdf+html;
http://www.jstor.org/
stable/3541679.

Loewer:1996:CL

[Loe96] Barry Loewer. Comment on Lockwood. British Journal for the Philosophy of Science, 47 (2):229–232, June 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/2/229.full.pdf+html;
http://www.jstor.org/
stable/687944.

Loewer:1998:CVB

[Loe98] Barry Loewer. Copenhagen versus Bohmian interpretations of quantum theory. British Journal for the Philosophy of Science, 49(2):317–328, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/2/317.full.pdf+html;
http://www.jstor.org/
stable/688118.

Lomas:2002:WPD

[Lom02] Dennis Lomas. What perception is doing, and what it is not doing, in mathematical reasoning. British Journal for the Philosophy of Science, 53(2):205–223, June 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/2/205.full.pdf+
html;   http://www.jstor.
org/stable/3541764.

Lonsdale:1951:R

[Lon51] Kathleen Lonsdale. Reviews.

British Journal for the Philosophy of Science, II(6):165–166, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/6/165.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685510.

Lonsdale:1952:R

[Lon52] Kathleen Lonsdale. Reviews. British Journal for the Philosophy of Science, III(11):278– 279, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/11/278.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685276.

Long:1953:CAP

[Lon53] Wilbur Long. Comments on the alleged proof of epiphenomenalism. British Journal for the Philosophy of Science, III(12): 355–358, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/355.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685454.

Losee:1986:SPN

[Los86] John Losee. Shapere’s project for a nonpresuppositionist philosophy of science. British Journal for the Philosophy of Science, 37(2):223–229, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686982.

Losee:1998:R

[Los98] John Losee. Review. British Journal for the Philosophy of Science, 49(3):508–510, September 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/508.full.pdf+
html;   http://www.jstor.
org/stable/688087.

Louch:1962:SP

[Lou62] A. R. Louch. Science and psychology. British Journal for the Philosophy of Science, XII(48): 314–327, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/48/314.full.pdf+html;
http://www.jstor.org/
stable/685399.

Lowe:1973:R

[Low73] Malcolm Lowe. Reviews. British Journal for the Philosophy of Science, 24(4):413–414, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/4/413.full.pdf+
html;   http://www.jstor.
org/stable/686625.

Lowe:1997:R

[Low97]  E. J. Lowe.  Review.

British Journal for the Philosophy of Science, 48(2):300–301, June 1997. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/300.full.pdf+
html;   http://www.jstor.
org/stable/687755.

Lowe:2000:BRB

[Low00]  E. J. Lowe.  Book review:

Interpreting Bodies: Classical and Quantum Objects in Modern Physics. E. Castellani [ed]. British Journal for the Philosophy of Science, 51(2):353–355, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/2/353.full.pdf+
html;   http://www.jstor.
org/stable/3541812.

Lowe:2004:RPS

[Low04] E. J. Lowe. Review: Powers: a study in metaphysics. British Journal for the Philosophy of Science, 55(4):817–822, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/4/817.full.pdf+
html;   http://www.jstor.
org/stable/3541635.

Ladyman:2018:DHT

[LP18]  James  Lady-

man and Stuart Presnell. Does homotopy type theory provide a foundation for mathematics? British Journal for the Philosophy of Science, 69(2):377–420, June 1, 2018. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/377/2669763.

Lucas:2007:TP

[LR07] John Lucas and Michael Redhead. Truth and provability. British Journal for the Philosophy of Science, 58(2):331–332, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/2/331.full.pdf+
html;   http://www.jstor.
org/stable/30115229.

Lean:2023:DLA

[LRP23] Oliver M. Lean, Luca Rivelli, and Charles H. Pence. Digital literature analysis for empirical philosophy of science. British Journal for the Philosophy of Science, 74(4):875–898, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lyngzeidetson:1994:ACT

[LS94] Albert E. Lyngzeidetson and Martin K. Solomon. Abstract complexity theory and the mindmachine problem. British Journal for the Philosophy of Science, 45(2):549–554, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/549.full.pdf+html;
http://www.jstor.org/
stable/687681.

Lal:2017:CGP

[LT17]  Raymond  Lal  and

Nicholas Teh. Categorical generalization and physical structuralism. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/1/213/2995114/
                                        Categorical-Generalization-
                                        and-Physical.

Linnebo:2009:WAP

[LU09] Øystein Linnebo and Gabriel Uzquiano. Which abstraction principles are acceptable? Some limitative results. British Journal for the Philosophy of Science, 60(2):239–252, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/2/239.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592000.

Lucas:1968:HML

[Luc68]  J. R. Lucas.  Human and

machine logic: a rejoinder. British Journal for the Philosophy of Science, 19(2):155–156, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/19/2/155.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686794.

Lucas:1969:EON

[Luc69a] J. R. Lucas. Euclides ab omni naevo vindicatus. British Journal for the Philosophy of Science, 20(1):1–11, May 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/1/1.full.pdf+html;
http://www.jstor.org/
stable/686714.

Lucas:1969:R

[Luc69b] J. R. Lucas. Reviews. British Journal for the Philosophy of Science, 20(2):170–171, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/170.full.pdf+
html;   http://www.jstor.
org/stable/686182.

Lucas:1978:R

[Luc78] J. R. Lucas. Reviews. British Journal for the Philosophy of Science, 29(3):299–301, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/3/299.full.pdf+
html;   http://www.jstor.
org/stable/687024.

Lucas:1980:R

[Luc80] J. R. Lucas. Reviews. British Journal for the Philosophy of Science, 31(4):414–416, December 1980. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/4/414.full.pdf+
html;   http://www.jstor.
org/stable/687072.

Lucas:1982:R

[Luc82] J. R. Lucas. Reviews. British Journal for the Philosophy of Science, 33(4):441–444, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/4/441.full.pdf+
html;   http://www.jstor.
org/stable/687176.

Lucas:1985:R

[Luc85] J. R. Lucas. Reviews. British Journal for the Philosophy of Science, 36(4):473–475, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/4/473.full.pdf+
html;   http://www.jstor.
org/stable/687408.

Luce:1991:LAA

[Luc91] L. Luce. Literalism and the applicability of arithmetic. British Journal for the Philosophy of Science, 42(4):469–489, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/4/469.full.pdf+
html;   http://www.jstor.
org/stable/687704.

Ludwig:2017:ISK

[Lud17] David Ludwig. Indigenous and scientific kinds. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL https://academic.oup.
com/bjps/article/68/1/187/
2995355/Indigenous-and-
Scientific-Kinds.

Luna:2021:SBE

[Lun21]  Laureano  Luna.

Sleeping Beauty: Exploring a neglected solution. British Journal for the Philosophy of Science, 71 (3):1069–1092, ???? 2021. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lupher:2018:LPE

[Lup18]  Tracy Lupher.  The limits

of physical equivalence in algebraic quantum field theory. British Journal for the Philosophy of Science, 69(2):553–576, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/553/2669785.

Lusk:2021:SD

[Lus21] Greg Lusk. Saving the data.

British Journal for the Philosophy of Science, 72(1):277–298, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Lam:2015:NCS

[LW15]  Vincent  Lam

and Christian Wüthrich. No categorial support for radical ontic structural realism. British Journal for the Philosophy of Science, 66(3):605–634, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/3/605.

Lyngzeidetson:1990:MPD

[Lyn90] Albert E. Lyngzeidetson. Massively parallel distributed processing and a computationalist foundation for cognitive science. British Journal for the Philosophy of Science, 41(1):121–127, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/1/121.full.pdf+
html;   http://www.jstor.
org/stable/688006.

Lyon:1967:C

[Lyo67] Ardon Lyon. Causality. British Journal for the Philosophy of Science, 18(1):1–20, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/1.full.pdf+html;
http://www.jstor.org/
stable/686886.

Lyons:2001:CMN

[Lyo01] Jack C. Lyons. Carving the mind at its (not necessarily modular) joints. British Journal for the Philosophy of Science, 52 (2):277–302, June 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/277.full.pdf+html;
http://www.jstor.org/
stable/3541897.

Lyons:2006:BRP

[Lyo06a] Timothy D. Lyons. Book review:

Peter Lipton: Inference to the Best Explanation, London, Routledge, 2004, 2nd Edition Paperback $33.95 ISBN 0-415-24203-7. British Journal for the Philosophy of Science, 57(1):255–258, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/1/255.full.pdf+
html;   http://www.jstor.
org/stable/3541662.

Lyons:2006:SRS

[Lyo06b] Timothy D. Lyons. Scientific realism and the stratagema de divide et impera. British Journal for the Philosophy of Science, 57(3):537–560, September 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/3/537.full.pdf+html;
http://www.jstor.org/
stable/3873482.

Lyon:2014:WND

[Lyo14] Aidan Lyon. Why are normal distributions normal? British Journal for the Philosophy of Science, 65(3):621–649, September 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/65/3/621.

Mumford:2011:GCP

[MA11] Stephen Mumford and Rani Lill Anjum. Getting causes from powers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-969561-X. xvi + 254 pp. LCCN BD591 .M864

                                        2011. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1210/2011276707-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1210/
                                        2011276707-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1210/
                                        2011276707-t.html.

Martinez:2023:NOR

[MA23] Manolo Martínez and Marc Artiga. Neural oscillations as representations. British Journal for the Philosophy of Science, 74(3): 619–648, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mackay:1951:MBA

[Mac51] D. M. Mackay. Mindlike behaviour in artefacts. British Journal for the Philosophy of Science, II(6):105–121, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/

content/II/6/105.full.pdf+
html;   http://www.jstor.
org/stable/685505.

Mackay:1953:MLB

[Mac53a] D. M. Mackay. Mind-like behaviour in artefacts. British Journal for the Philosophy of Science, III(12):352–353, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/12/352.full.
pdf+html;   http://www.
jstor.org/stable/685452.

Mackay:1953:R

[Mac53b] D. M. Mackay. Reviews.

British Journal for the Philosophy of Science, IV(15):249– 251, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/249.full.pdf+html;
http://www.jstor.org/
stable/685903.

Macdonald:1960:DPC

[Mac60a] D. K. C. Macdonald. Discussion: Physics and chemistry: Comments on Caldin’s view of chemistry. British Journal for the Philosophy of Science, XI (43):222–223, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/222.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685311.

Mackay:1960:Ra

[Mac60b] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, XI(41):82–83, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/82.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685830.

Mackay:1960:Rb

[Mac60c] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, XI(42):166–167, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/42/166.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685598.

Mackay:1962:UBL

[Mac62] D. M. Mackay. The use of behavioural language to refer to mechanical processes. British Journal for the Philosophy of Science, XIII(50):89–103, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/89.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685963.

Mackay:1963:CMR

[Mac63a] D. M. Mackay. Consciousness and mechanism: a reply to Miss Fozzy. British Journal for the Philosophy of Science, XIV(54):157–159, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/54/157.full.pdf+html;
http://www.jstor.org/
stable/685435.

Mackie:1963:PC

[Mac63b] J. L. Mackie. The paradox of confirmation. British Journal for the Philosophy of Science, XIII(52):265–277, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/265.full.
pdf+html;   http://www.
jstor.org/stable/685324.

Mackay:1965:Ra

[Mac65a] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, XVI(62):163–164, August 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/62/163.full.
pdf+html;   http://www.
jstor.org/stable/686157.

Mackay:1965:Rb

[Mac65b] D. M. Mackay. Reviews.

British Journal for the Philosophy of Science, XVI(63):253– 255, November 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XVI/63/253.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686311.

Mackie:1965:R

[Mac65c] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, XVI(61):57– 61, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/61/57.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686142.

Mackie:1966:MCP

[Mac66]  J. L. Mackie.  Miller’s so-

called paradox of information. British Journal for the Philosophy of Science, 17(2):144–147, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/17/2/144.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686346.

Mackay:1967:R

[Mac67] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, 18(3):259–261, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/3/259.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686601.

Macintyre:1968:R

[Mac68a] Alasdair Macintyre. Reviews.

British Journal for the Philosophy of Science, 18(4):344– 345, February 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/4/344.full.pdf+html;
http://www.jstor.org/
stable/686366.

Mackie:1968:Ra

[Mac68b] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, 19(2):168–171, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/168.full.pdf+
html;   http://www.jstor.
org/stable/686797.

Mackie:1968:Rb

[Mac68c] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, 19(3):261–263, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/261.full.pdf+
html;   http://www.jstor.
org/stable/686248.

Macintyre:1969:Ra

[Mac69a] Alasdair Macintyre. Reviews.

British Journal for the Philosophy of Science, 20(2):173–174, August 1969. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/173.full.pdf+
html;   http://www.jstor.
org/stable/686184.

Macintyre:1969:Rb

[Mac69b] Alasdair Macintyre. Reviews. British Journal for the Philosophy of Science, 20(2):174–175, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/174.full.pdf+
html;   http://www.jstor.
org/stable/686185.

Mackie:1969:RCC

[Mac69c] J. L. Mackie. The relevance criterion of confirmation. British Journal for the Philosophy of Science, 20(1):27–40, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/1/27.full.pdf+html;
http://www.jstor.org/
stable/686716.

Macklin:1969:ACT

[Mac69d] Ruth Macklin. Action, causality, and teleology. British Journal for the Philosophy of Science, 19(4):301–316, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/4/301.full.pdf+html;

http://www.jstor.org/
stable/686290.

MacCormac:1971:MVM

[Mac71a] Earl R. MacCormac. Meaning variance and metaphor. British Journal for the Philosophy of Science, 22(2):145–159, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/145.full.pdf+
html;   http://www.jstor.
org/stable/686484.

Mackay:1971:CMU

[Mac71b] D. M. Mackay. Choice in a mechanistic universe: a reply to some critics. British Journal for the Philosophy of Science, 22 (3):275–285, August 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/3/275.full.pdf+html;
http://www.jstor.org/
stable/686747.

Mackay:1971:N

[Mac71c] D. M. Mackay. Note. British Journal for the Philosophy of Science, 22(1):53, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/53.full.pdf+html.

Mackay:1971:R

[Mac71d] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, 22(1):65–67, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/65.full.pdf+html;
http://www.jstor.org/
stable/686651.

Mackie:1971:R

[Mac71e] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, 22(4):389–396, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/4/389.full.pdf+
html;   http://www.jstor.
org/stable/686324.

Mackie:1972:MCS

[Mac72] J. L. Mackie. Metaphysical common sense. British Journal for the Philosophy of Science, 23 (3):247–252, August 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/3/247.full.pdf+html;
http://www.jstor.org/
stable/686687.

Mackay:1973:LIH

[Mac73a] D. M. Mackay. The logical indeterminateness of human choices. British Journal for the Philosophy of Science, 24(4): 405–408, December 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
24/4/405.full.pdf+html;
http://www.jstor.org/
stable/686622.

Mackie:1973:R

[Mac73b] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, 24(2):190–193, June 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/2/190.full.pdf+
html;   http://www.jstor.
org/stable/686611.

Machlup:1974:SDE

[Mac74a] Fritz Machlup. Situational determinism in economics. British Journal for the Philosophy of Science, 25(3):271–284, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/3/271.full.pdf+
html;   http://www.jstor.
org/stable/686909.

Mackie:1974:R

[Mac74b] J. L. Mackie. Reviews. British Journal for the Philosophy of Science, 25(1):83–84, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/83.full.pdf+
html;   http://www.jstor.
org/stable/686504.

Macintyre:1978:R

[Mac78a] Alasdair Macintyre. Reviews. British Journal for the Philosophy of Science, 29(1):105–110, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/29/1/105.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686396.

Mackay:1978:R

[Mac78b] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, 29(4):394–395, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/29/4/394.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687105.

Mackie:1978:FCP

[Mac78c] J. L. Mackie. Failures in criticism: Popper and his commentators. British Journal for the Philosophy of Science, 29(4): 363–375, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/4/363.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687101.

Mackay:1979:R

[Mac79] D. M. Mackay. Reviews. British Journal for the Philosophy of Science, 30(3):302–304, September 1979. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/302.full.pdf+
html;   http://www.jstor.
org/stable/687146.

Machover:1983:TNP

[Mac83] Moshe Machover. Towards a new philosophy of mathematics. British Journal for the Philosophy of Science, 34(1):1–11, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/1/1.full.pdf+
html;   http://www.jstor.
org/stable/686930.

Macrossan:1986:NRB

[Mac86] M. N. Macrossan. A note on Relativity before Einstein. British Journal for the Philosophy of Science, 37(2):232–234, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/686984.

Mackay:1987:R

[Mac87] Donald M. Mackay. Reviews.

British Journal for the Philosophy of Science, 38(3):408– 412, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/408.full.pdf+html;
http://www.jstor.org/
stable/687499.

Macnamara:1991:UI

[Mac91] John Macnamara. Understanding induction. British Journal for the Philosophy of Science, 42(1):21–48, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/1/21.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687605.

Machover:1993:PNA

[Mac93] Moshé Machover. The place of nonstandard analysis in mathematics and in mathematics teaching. British Journal for the Philosophy of Science, 44 (2):205–212, June 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        44/2/205.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687644.

Machover:1994:R

[Mac94a] Moshé Machover. Review. British Journal for the Philosophy of Science, 45(4):1078– 1083, December 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/4/1078.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687626.

Macintosh:1994:PCA

[Mac94b] Duncan Macintosh. Partial convergence and approximate truth. British Journal for the Philosophy of Science, 45(1):153–170, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/153.full.pdf+
html;   http://www.jstor.
org/stable/687965.

MacBride:1996:R

[Mac96a] Fraser MacBride. Review.

British Journal for the Philosophy of Science, 47(3):463– 466, September 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/3/463.full.pdf+html;
http://www.jstor.org/
stable/687989.

Machover:1996:Ra

[Mac96b] Moshé Machover. Review.

British Journal for the Philosophy of Science, 47(1):137–140, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/137.full.pdf+
html;   http://www.jstor.
org/stable/1215833.

Machover:1996:Rb

[Mac96c] Moshé Machover. Review.

British Journal for the Philosophy of Science, 47(2):320–324, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/2/320.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687952.

Macdonald:1998:R

[Mac98] Graham Macdonald. Review. British Journal for the Philosophy of Science, 49(1):175–180, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/1/175.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688150.

MacBride:1999:SAL

[Mac99] F. MacBride. Survey article. listening to fictions: a study of Fieldian nominalism. British Journal for the Philosophy of Science, 50(3):431–455, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/50/3/431.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072241.

MacBride:2003:SSS

[Mac03]  Fraser MacBride.  Speaking

with shadows: a study of neologicism. British Journal for the Philosophy of Science, 54 (1):103–163, March 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/1/103.full.pdf+html;

http://www.jstor.org/
stable/3541882.

Machery:2005:YDK

[Mac05] Edouard Machery. You don’t know how you think: Introspection and language of thought. British Journal for the Philosophy of Science, 56(3):469– 485, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/469.full.pdf+html;
http://www.jstor.org/
stable/3541715.

Machery:2014:DRI

[Mac14] Edouard Machery. In defense of reverse inference. British Journal for the Philosophy of Science, 65(2):251–267, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/251.

Madden:1957:LAP

[Mad57] Edward H. Madden. A logical analysis of ‘psychological isomorphism’. British Journal for the Philosophy of Science, VIII (31):177–191, November 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/31/177.full.pdf+html;
 http://www.jstor.org/
stable/685756.

Maddy:1986:MA

[Mad86] Penelope Maddy. Mathematical alchemy. British Journal for the Philosophy of Science, 37(3):279–314, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687288.

Maddy:2007:SPN

[Mad07] Penelope Maddy. Second philosophy: a naturalistic method. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-927366-9 (paperback). xii + 448 pp. LCCN B828.2 .M33 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0714/2007012570.html.

Magyar:1965:DNL

[Mag65] George Magyar. On the dual nature of light. British Journal for the Philosophy of Science, XVI(61):44–49, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XVI/61/44.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686138.

Magnus:2003:STG

[Mag03] P. D. Magnus. Success, truth and the Galilean strategy. British Journal for the Philosophy of Science, 54(3):465–474, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/

content/54/3/465.full.pdf+
html;   http://www.jstor.
org/stable/3541795.

Magnus:2004:PIQ

[Mag04] P. D. Magnus. The price of insisting that quantum mechanics is complete. British Journal for the Philosophy of Science, 55(2):257–267, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/2/257.full.pdf+html;
http://www.jstor.org/
stable/3541690.

Magnus:2005:RSE

[Mag05] P. D. Magnus. Reckoning the shape of everything: Underdetermination and cosmotopology. British Journal for the Philosophy of Science, 56(3):541– 557, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/541.full.pdf+html;
http://www.jstor.org/
stable/3541718.

Magnus:2010:IRH

[Mag10] P. D. Magnus. Inductions, red herrings, and the best explanation for the mixed record of science. British Journal for the Philosophy of Science, 61(4): 803–819, December 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        61/4/803.full.pdf+html.

Mahoney:1970:Rb

[Mah70a] M. S. Mahoney. Reviews. British Journal for the Philosophy of Science, 21(3):305–307, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/305.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686209.

Mahoney:1970:Ra

[Mah70b] Michael S. Mahoney. Reviews. British Journal for the Philosophy of Science, 21(3):301–302, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/301.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686206.

Maher:1990:WSG

[Mah90] Patrick Maher. Why scientists gather evidence. British Journal for the Philosophy of Science, 41 (1):103–119, March 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        41/1/103.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688005.

Maidens:1992:R

[Mai92] Anna Maidens. Review. British Journal for the Philosophy of Science, 43(1):129–136, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/1/129.full.pdf+
html;   http://www.jstor.
org/stable/687888.

Maidens:1995:R

[Mai95] Anna Maidens. Review. British Journal for the Philosophy of Science, 46(2):253–260, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/253.full.pdf+
html;   http://www.jstor.
org/stable/687816.

Maley:2023:ACR

[Mal23]  Corey J. Maley.  Analogue

computation and representation. British Journal for the Philosophy of Science, 74(3):739–769, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mameli:2004:NSN

[Mam04] Matteo Mameli. Nongenetic selection and nongenetic inheritance. British Journal for the Philosophy of Science, 55 (1):35–71, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/35.full.pdf+html;
http://www.jstor.org/
stable/3541833.

Mandelbaum:1957:SL

[Man57] Maurice Mandelbaum. Societal laws. British Journal for the Philosophy of Science, VIII(31): 211–224, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/31/211.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685758.

Mannu:1974:R

[Man74] Andrea Mannu. Reviews. British Journal for the Philosophy of Science, 25(2):198–206, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/25/2/198.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686825.

Manders:1976:FCS

[Man76] Kenneth L. Manders. Friedman’s criterion for simplicity. British Journal for the Philosophy of Science, 27(4):395–397, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/27/4/395.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686866.

Manier:1978:R

[Man78]  Edward Manier.  Reviews.

British Journal for the Philosophy of Science, 29(1):110–115, March 1978. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/1/110.full.pdf+
html;   http://www.jstor.
org/stable/686397.

Manson:1996:R

[Man96] Neil Manson. Review. British Journal for the Philosophy of Science, 47(4):633–635, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/633.full.pdf+
html;   http://www.jstor.
org/stable/687931.

Manning:1997:BFS

[Man97] Richard N. Manning. Biological function, selection, and reduction. British Journal for the Philosophy of Science, 48(1):69–82, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/1/69.full.pdf+
html;   http://www.jstor.
org/stable/687718.

Manchak:2021:MHM

[Man21a] J. B. Manchak. Malament– Hogarth machines. British Journal for the Philosophy of Science, 71(3):1143–1153, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mandrigin:2021:MIS

[Man21b] Alisa Mandrigin. Multisensory integration and sense modalism.

British Journal for the Philosophy of Science, 72(1):27–49, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Martin:1957:HET

[Mar57] Anne Martin. How economic theory may mislead. British Journal for the Philosophy of Science, VIII(31):225–236, November 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/31/225.full.
pdf+html;   http://www.
jstor.org/stable/685759.

Maruyama:1959:CIR

[Mar59a] Magoroh Maruyama. Communicable and incommunicable realities. British Journal for the Philosophy of Science, X (37):50–54, May 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685777.

Maruyama:1959:DCI

[Mar59b] Magoroh Maruyama. Discussions: Communicable and incommunicable realities. British Journal for the Philosophy of Science, X(37):50b–54, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/50-b.full.pdf+html.

Marshall:1960:EMP

[Mar60] Ninian Marshall. ESP and memory: a physical theory. British Journal for the Philosophy of Science, X(40):265–286, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/40/265.full.pdf+
html;   http://www.jstor.
org/stable/685915.

Marshall:1961:RDH

[Mar61a] Ninian Marshall. Reply to Dr H. A. C. Dobbs. British Journal for the Philosophy of Science, XII(45):68–70, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/45/68.full.pdf+html;
http://www.jstor.org/
stable/685541.

Maruyama:1961:CE

[Mar61b] Magoroh Maruyama. Communicational epistemology (I). British Journal for the Philosophy of Science, XI(44):319–327, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/44/319.full.pdf+
html;   http://www.jstor.
org/stable/685132.

Maruyama:1961:CEIa

[Mar61c] Magoroh Maruyama. Communicational epistemology (II).

British Journal for the Philosophy of Science, XII(45):52–62, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/52.full.pdf+
html;   http://www.jstor.
org/stable/685538.

Maruyama:1961:CEIb

[Mar61d] Magoroh Maruyama. Communicational epistemology (III). British Journal for the Philosophy of Science, XII(46):117–131, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/117.full.
pdf+html;   http://www.
jstor.org/stable/685491.

Martin:1962:ELP

[Mar62] Anne Martin. Epistemology as liberal polemic. British Journal for the Philosophy of Science, XIII(49):65–70, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/49/65.full.pdf+html;
http://www.jstor.org/
stable/685230.

Martin:1964:EPE

[Mar64] Anne Martin. Empirical and a priori in economics. British Journal for the Philosophy of Science, XV(58):123–136, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/123.full.pdf+
html;   http://www.jstor.
org/stable/685251.

Martin:1969:ALD

[Mar69] Jane R. Martin. Another look at the doctrine of Verstehen. British Journal for the Philosophy of Science, 20(1):53–67, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/53.full.pdf+
html;   http://www.jstor.
org/stable/686718.

Martin:1971:RVS

[Mar71] Michael Martin. Referential variance and scientific objectivity. British Journal for the Philosophy of Science, 22(1): 17–26, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/17.full.pdf+html;
http://www.jstor.org/
stable/686640.

Martin:1972:OVS

[Mar72] Michael Martin. Ontological variance and scientific objectivity. British Journal for the Philosophy of Science, 23(3): 252–256, August 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

23/3/252.full.pdf+html;
http://www.jstor.org/
stable/686688.

Martin:1987:R

[Mar87] R. N. D. Martin. Reviews. British Journal for the Philosophy of Science, 38(3):406– 408, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/406.full.pdf+html;
http://www.jstor.org/
stable/687498.

Martin:1989:ACT

[Mar89a] James E. Martin. Aesthetic constraints on theory selection: a critique of Laudan. British Journal for the Philosophy of Science, 40(3):357–364, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/3/357.full.pdf+
html;   http://www.jstor.
org/stable/687782.

Marusic:1989:BFP

[Mar89b] M. Marušiç. Biological foundations of prediction in an unpredictable environment. British Journal for the Philosophy of Science, 40(4):485–499, December 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/4/485.full.pdf+
html;   http://www.jstor.
org/stable/687737.

Marinoff:1990:IES

[Mar90] Louis Marinoff. The inapplicability of evolutionarily stable strategy to the Prisoner’s Dilemma. British Journal for the Philosophy of Science, 41(4): 461–472, December 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/461.full.pdf+html;
http://www.jstor.org/
stable/687860.

Martin:2000:BRB

[Mar00] C. A. Martin. Book review: Conceptual Developments of 20th Century Field Theories. T. Y. Cao. British Journal for the Philosophy of Science, 51(3):519– 525, September 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/519.full.pdf+html;
http://www.jstor.org/
stable/3541781.

Marras:2005:CR

[Mar05] Ausonio Marras. Consciousness and reduction. British Journal for the Philosophy of Science, 56(2):335–361, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/2/335.full.pdf+html;
http://www.jstor.org/
stable/3541706.

Martens:2017:IFM

[Mar17] Johannes Martens. Inclusive fitness and the maximizingagent analogy. British Journal for the Philosophy of Science, 68(3):875–905, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/875/2669750.

Martens:2022:MCA

[Mar22] Niels C. M. Martens. Machian comparativism about mass. British Journal for the Philosophy of Science, 73(2): 325–349, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mascall:1953:R

[Mas53] E. L. Mascall. Reviews. British Journal for the Philosophy of Science, IV(16):349–350, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        IV/16/349.full.pdf+html.

Mascall:1954:RU

[Mas54] E. L. Mascall. Review: [untitled]. British Journal for the Philosophy of Science, IV(16): 349–350, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686088.

Massimi:2001:EPI

[Mas01] Michela Massimi. Exclusion Principle and the identity of indiscernibles: a response to Margenau’s argument. British Journal for the Philosophy of Science, 52(2):303–330, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/303.full.pdf+html;
http://www.jstor.org/
stable/3541898.

Massimi:2005:PEP

[Mas05] Michela Massimi. Pauli’s exclusion principle: the origin and validation of a scientific principle. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83911-4 (hardcover). xiv + 211 pp. LCCN QC174.17.P3 M37

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2005296620-d.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2005296620-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2005296620-b.html.

Massimi:2007:SUP

[Mas07] Michela Massimi. Saving unobservable phenomena. British Journal for the Philosophy of Science, 58(2):235–262, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/58/2/235.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/30115225.

Massimi:2016:E

[Mas16] Michela Massimi. Editorial. British Journal for the Philosophy of Science, 67(4):937– 939, December 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/4/937.

Massin:2017:CF

[Mas17] Olivier Massin. The composition of forces. British Journal for the Philosophy of Science, 68(3):805–846, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/805/2669641.

Matthew:1971:PP

[Mat71] Anthony Matthew. Prediction and predication. British Journal for the Philosophy of Science, 22(2):171–182, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        22/2/171.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686486.

Matthews:1981:R

[Mat81]  P. H. Matthews.  Reviews.

British Journal for the Philosophy of Science, 32(4):416– 419, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/416.full.pdf+html;
http://www.jstor.org/
stable/687313.

Matthews:1982:R

[Mat82] Michael Matthews. Reviews.

British Journal for the Philosophy of Science, 33(2):220–223, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/2/220.full.pdf+
html;   http://www.jstor.
org/stable/687282.

Matthen:1999:DEW

[Mat99] M. Matthen. Discussion. evolution, Wisconsin style: selection and the explanation of individual traits. British Journal for the Philosophy of Science, 50 (1):143–150, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/143.full.pdf+html;
http://www.jstor.org/
stable/40072212.

Matthen:2003:SRN

[Mat03] Mohan Matthen. Is sex really necessary? and other questions for Lewens. British Journal for the Philosophy of Science, 54(2):297–308, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/2/297.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541969.

Matthen:2005:SDK

[Mat05] Mohan Matthen. Seeing, Doing, and Knowing: a Philosophical Theory of Sense Perception. Clarendon Press, New York, NY, USA, 2005. ISBN 0-19-926850-9 (hardcover). xxii + 362 pp. LCCN QP431 .M295 2005.

Matthiessen:2017:MES

[Mat17] Dana Matthiessen. Mechanistic explanation in systems biology: Cellular networks. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/1/1/2995179/
                                        Mechanistic-Explanation-in-
                                        Systems-Biology.

Maudlin:1995:R

[Mau95] Tim Maudlin. Review. British Journal for the Philosophy of Science, 46(1):145–149, March 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/1/145.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687639.

Maudlin:2000:DQC

[Mau00] Tim Maudlin. Dickson on quantum chance and non-locality. British Journal for the Philosophy of Science, 51(4):875– 882, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/875.full.pdf+html;
http://www.jstor.org/
stable/3541731.

Maudlin:2007:MWP

[Mau07] Tim Maudlin. The metaphysics within physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-921821-8, 0-19-957537-1 (paperback). 197 pp. LCCN BD111 .M46 2007. URL http://www.

loc.gov/catdir/toc/ecip077/
2006102303.htm.

Maxwell:1966:PCS

[Max66] Nicholas Maxwell. Physics and common sense. British Journal for the Philosophy of Science, XVI(64):295–311, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/64/295.full.
pdf+html;   http://www.
jstor.org/stable/686831.

Maxwell:1968:CTN

[Max68] Nicholas Maxwell. Can there be necessary connections between successive events? British Journal for the Philosophy of Science, 19(1):1–25, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/1/1.full.pdf+html;
http://www.jstor.org/
stable/686845.

Maxwell:1969:Rb

[Max69a] A. N. Maxwell. Reviews. British Journal for the Philosophy of Science, 20(4):361–364, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/361.full.pdf+
html;   http://www.jstor.
org/stable/686266.

Maxwell:1969:Ra

[Max69b] Nicholas Maxwell. Reviews. British Journal for the Philosophy of Science, 20(1):81–83, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/1/81.full.pdf+
html;   http://www.jstor.
org/stable/686721.

Maxwell:1971:R

[Max71] Nicholas Maxwell. Reviews. British Journal for the Philosophy of Science, 22(4):387– 389, November 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/4/387.full.pdf+html;

http://www.jstor.org/
stable/686323.

Maxwell:1988:QPT

[Max88] Nicholas Maxwell. Quantum propensiton theory: a testable resolution of the wave/particle dilemma. British Journal for the Philosophy of Science, 39 (1):1–50, March 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/1/1.full.pdf+html;
http://www.jstor.org/
stable/687112.

Maxwell:1993:ISRa

[Max93a] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. part one: How to solve the problem of induction. British Journal for the Philosophy of Science, 44(1):61–79, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/61.full.pdf+
html;   http://www.jstor.
org/stable/687850.

Maxwell:1993:ISRc

[Max93b] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. part three: Einstein, aim-oriented empiricism and the discovery of Special and General Relativity. British Journal for the Philosophy of Science, 44(2):275–305, June 1993. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/2/275.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687649.

Maxwell:1993:ISRb

[Max93c] Nicholas Maxwell. Induction and scientific realism: Einstein versus van Fraassen. Part two: Aimoriented empiricism and scientific essentialism. British Journal for the Philosophy of Science, 44(1):81–101, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        44/1/81.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687851.

Maxwell:2004:SN

[Max04] Nicholas Maxwell. Is science neurotic? Imperial College Press, Oxford, UK, 2004. ISBN 1-86094-500-7. xx + 240 pp. LCCN Q175 .M4118 2004.

Mays:1951:HC

[May51] W. Mays. The hypothesis of cybernetics. British Journal for the Philosophy of Science, II(7): 249–250, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/7/249.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685296.

Mays:1952:MLB

[May52] W. Mays. Mind-like behaviour in artefacts and the concept of mind. British Journal for the Philosophy of Science, III(10): 191–193, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/10/191.full.pdf+html;
http://www.jstor.org/
stable/685558.

Mays:1956:Ra

[May56a] W. Mays. Reviews. British Journal for the Philosophy of Science, VII(27):263–267, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/263.full.
pdf+html;   http://www.
jstor.org/stable/685886.

Mays:1956:Rb

[May56b] W. Mays. Reviews. British Journal for the Philosophy of Science, VII(27):267–268, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/267.full.
pdf+html;   http://www.
jstor.org/stable/685887.

Mays:1957:R

[May57] W. Mays. Reviews. British Journal for the Philosophy of Science, VIII(30):164–168, August 1957. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/30/164.full.
pdf+html;   http://www.
jstor.org/stable/685797.

Mays:1959:R

[May59] W. Mays. Reviews. British Journal for the Philosophy of Science, X(39):251–252, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/39/251.full.pdf+
html;   http://www.jstor.
org/stable/686065.

Mays:1960:HPS

[May60] W. Mays. History and philosophy of science in British Commonwealth universities. British Journal for the Philosophy of Science, XI(43):192–211, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/43/192.full.pdf+
html;   http://www.jstor.
org/stable/685309.

Mays:1961:RNW

[May61] W. Mays. Reply to note by W. P. D. Wightman. British Journal for the Philosophy of Science, XI(44):336b–336, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/XI/44/336-b.full.
pdf+html;   http://www.
jstor.org/stable/685137.

Mays:1964:Ra

[May64a] W. Mays. Reviews. British Journal for the Philosophy of Science, XV(57):78–82, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/57/78.full.pdf+html;
http://www.jstor.org/
stable/685629.

Mays:1964:Rb

[May64b] W. Mays. Reviews. British Journal for the Philosophy of Science, XV(59):254–259, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/254.full.pdf+
html;   http://www.jstor.
org/stable/686527.

Mayberry:1977:CPSb

[May77a] John Mayberry. The consistency problem for set theory: an essay on the Cantorian foundations of mathematics (II). British Journal for the Philosophy of Science, 28(2):137–170, June 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/2/137.full.pdf+html;
http://www.jstor.org/
stable/686630.

Mayberry:1977:CPSa

[May77b] John Mayberry. On the consistency problem for set theory: an essay on the Cantorian foundations of classical mathematics (I). British Journal for the Philosophy of Science, 28(1):1–34, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/28/1/1.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686444.

Mayberry:1980:NB

[May80a] John P. Mayberry. A new Begriffsschrift (I). British Journal for the Philosophy of Science, 31(3):213–254, September 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/3/213.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686919.

Mayberry:1980:NBI

[May80b] John P. Mayberry. A new Begriffsschrift (II). British Journal for the Philosophy of Science, 31(4):329–358, December 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/4/329.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687062.

Mayo:1996:DRN

[May96] Deborah G. Mayo. Ducks, rabbits, and normal science: Recasting the Kuhn’s-eye view of Popper’s demarcation of science. British Journal for the Philosophy of Science, 47(2):271–290, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/271.full.pdf+
html;   http://www.jstor.
org/stable/687948.

Mayberry:2000:BRB

[May00] J. P. Mayberry. Book review:

A Primer of Infinitesimal Analysis. J. L. Bell. British Journal for the Philosophy of Science, 51(2):339–345, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/2/339.full.pdf+html;
http://www.jstor.org/
stable/3541810.

Mayo:2008:HDD

[May08] Deborah G. Mayo. How to discount double-counting when it counts: Some clarifications. British Journal for the Philosophy of Science, 59(4):857– 879, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/857.full.pdf+html;
http://www.jstor.org/
stable/40072315.

McCullough-Benner:2021:RWI

[MB21] Colin McCullough-Benner. Representing the world with inconsistent mathematics. British Journal for the Philosophy of Science, 71(4):1331–1358, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Machery:2012:EBS

[MC12] Edouard Machery and Kara Cohen. An evidence-based study of the evolutionary behavioral sciences. British Journal for the Philosophy of Science, 63 (1):177–226, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/1/177.full.pdf+html.

McAllister:1986:TAH

[McA86] James W. McAllister. Theoryassessment in the historiography of science. British Journal for the Philosophy of Science, 37(3):315–333, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687289.

McCrea:1950:R

[McC50] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, I(1):75–76, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/1/75.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685672.

McCrea:1951:Ra

[McC51a] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, II(6):170–171, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/170.full.pdf+
html;   http://www.jstor.
org/stable/685514.

McCrea:1951:Rb

[McC51b] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, II(7):257–260, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/7/257.full.pdf+
html;   http://www.jstor.
org/stable/685300.

McCrea:1952:Ra

[McC52a] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, II(8):332–333, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/332.full.pdf+
html;   http://www.jstor.
org/stable/686030.

McCrea:1952:Rb

[McC52b] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, II(8):342–343, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/342.full.pdf+
html;   http://www.jstor.
org/stable/686035.

McCrea:1953:R

[McC53] W. H. McCrea. Reviews. British Journal for the Philosophy of Science, IV(14):175–177, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/14/175.full.pdf+
html;   http://www.jstor.
org/stable/685997.

McCulloch:1954:TM

[McC54] Warren S. McCulloch. Through the den of the metaphysician. British Journal for the Philosophy of Science, V(17):18–31, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/17/18.full.pdf+
html;   http://www.jstor.
org/stable/685949.

McCrea:1957:OEW

[McC57] W. H. McCrea. On the objective of Einstein’s work. British Journal for the Philosophy of Science, VIII(29):18–29, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/29/18.full.pdf+html;
http://www.jstor.org/
stable/685378.

McCausland:1984:ESR

[McC84] Ian McCausland. Einstein and Special Relativity: Who wrote the added footnotes? British Journal for the Philosophy of Science, 35(1):60–61, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/60.full.pdf+
html;   http://www.jstor.
org/stable/687559.

McCarthy:1986:CVN

[McC86] Jeremiah E. McCarthy. Campbell and Vinci on novel confirmation. British Journal for the Philosophy of Science, 37 (2):216–220, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686980.

McConwell:2016:BRM

[McC16] Alison K. McConwell. Book review: Maureen A. O’Malley, Philosophy of Microbiology. British Journal for the Philosophy of Science, 67(3):931–935, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/3/931.

McCoy:2019:EJM

[McC19] C. D. McCoy. Epistemic justification and methodological luck in inflationary cosmology. British Journal for the Philosophy of Science, 70(4):1003–1028, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/70/4/1003/4819211.

McCutcheon:2023:NVA

[McC23] Randall G. McCutcheon. A note on verisimilitude and accuracy. British Journal for the Philosophy of Science, 74(2):431– 434, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mcdermott:1972:IFB

[Mcd72]  John Mcdermott.  I’m free

because I know that I don’t yet know what I’m going to do? British Journal for the Philosophy of Science, 23(4): 343–346, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/343.full.pdf+html;
http://www.jstor.org/
stable/686737.

McDermott:1995:RC

[McD95] Michael McDermott. Redundant causation. British Journal for the Philosophy of Science, 46(4):523–544, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/4/523.full.pdf+html;
http://www.jstor.org/
stable/687896.

Mcfetridge:1975:R

[Mcf75]  I. G. Mcfetridge.  Reviews.

British Journal for the Philosophy of Science, 26(1):83–85, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/1/83.full.pdf+
html;   http://www.jstor.
org/stable/686338.

Mcgonigle:1970:ESL

[Mcg70] T. G. Mcgonigle. Euclidean space: a lasting philosophical obsession. British Journal for the Philosophy of Science, 21 (2):185–191, May 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/2/185.full.pdf+html;
http://www.jstor.org/
stable/686698.

McGrew:2003:CHE

[McG03] Timothy McGrew. Confirmation, heuristics, and explanatory reasoning. British Journal for the Philosophy of Science, 54(4):553–567, December 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/4/553.full.pdf+html;
http://www.jstor.org/
stable/3541678.

Mckinney:1991:EEP

[Mck91] William J. Mckinney. Experimenting on and experimenting with: Polywater and experimental realism. British Journal for the Philosophy of Science, 42(3):295–307, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/3/295.full.pdf+html;
http://www.jstor.org/
stable/687909.

McKenzie:2014:PPP

[McK14] Kerry McKenzie. Priority and particle physics: Ontic structural realism as a fundamentality thesis. British Journal for the Philosophy of Science, 65 (2):353–380, June 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/353.

McKenzie:2016:BRS

[McK16] Kerry McKenzie. Book review: Stephen Mumford and Matthew Tugby, Metaphysics and Science. British Journal for the Philosophy of Science, 67(2):643–650, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/2/643.

McKenzie:2020:SID

[McK20] Kerry McKenzie. Structuralism in the idiom of determination. British Journal for the Philosophy of Science, 71(2):497–522, June 2020. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/71/2/497/4708244.

Mclarty:1990:UAH

[Mcl90] Colin Mclarty. The uses and abuses of the history of topos theory. British Journal for the Philosophy of Science, 41(3): 351–375, September 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/3/351.full.pdf+html;
http://www.jstor.org/
stable/687825.

McLarty:2007:LMH

[McL07]  Colin McLarty.  The last

mathematician from Hilbert’s Göttingen: Saunders Mac Lane as philosopher of mathematics. British Journal for the Philosophy of Science, 58(1):77–112, March 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/1/77.full.pdf+
html;   http://www.jstor.
org/stable/4489101.

Mcmichael:1985:VFI

[Mcm85] Alan Mcmichael. Van Fraassen’s instrumentalism. British Journal for the Philosophy of Science, 36(3):257–272, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/257.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687570.

McMullin:1997:R

[McM97] Ernan McMullin. Review. British Journal for the Philosophy of Science, 48(4):606–609, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/4/606.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688104.

McWhirter:2016:BDF

[McW16] Gregory McWhirter. Behavioural deception and formal models of communication. British Journal for the Philosophy of Science, 67(3):757–780, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/3/757.

McCaffrey:2022:MPI

[MD22] Joseph McCaffrey and David Danks. Mixtures and psychological inference with resting state fMRI. British Journal for the Philosophy of Science, 73(3): 583–611, September 2022. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Meacham:2010:TMR

[Mea10] Christopher J. G. Meacham. Two mistakes regarding the principal principle. British Journal for the Philosophy of Science, 61(2):407–431, June 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/2/407.full.pdf+html.

Medawar:1954:R

[Med54]  P. B. Medawar.  Reviews.

British Journal for the Philosophy of Science, V(18):172–173, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/172.full.pdf+
html;   http://www.jstor.
org/stable/685182.

Meek:1964:VJE

[Mee64]  Ronald  L.  Meek.

Value-judgments in economics. British Journal for the Philosophy of Science, XV(58):89–96, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/58/89.full.pdf+
html;   http://www.jstor.
org/stable/685248.

Meehl:1999:DHW

[Mee99] P. E. Meehl. Discussion. how to weight scientists’ probabilities is not a big problem: comment on Barnes. British Journal for the Philosophy of Science, 50(2):283–295, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        50/2/283.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072224.

Meehl:2004:CMI

[Mee04]  Paul E. Meehl.  Cliometric

metatheory III: Peircean consensus, verisimilitude and asymptotic method. British Journal for the Philosophy of Science, 55(4):615–643, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/4/615.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541619.

Meehan:2022:NPQ

[Mee22] Alexander Meehan. A new problem for quantum mechanics. British Journal for the Philosophy of Science, 73(3):631–661, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Meier:2022:DBD

[Mei22] Lukas J. Meier. The demise of brain death. British Journal for the Philosophy of Science, 73(2): 487–508, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mellor:1965:CCI

[Mel65]  D.  H.

Mellor. Connectivity, chance, and ignorance. British Journal for the Philosophy of Science, XVI(63):209–225, November 1965. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/63/209.full.
pdf+html;   http://www.
jstor.org/stable/686303.

Mellor:1967:CCI

[Mel67a] D. H. Mellor. Connectivity, chance, and ignorance. British Journal for the Philosophy of Science, 18(3):235–238, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/235.full.pdf+
html;   http://www.jstor.
org/stable/686595.

Mellor:1967:R

[Mel67b] D. H. Mellor. Reviews. British Journal for the Philosophy of Science, 17(4):323–326, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/323.full.pdf+
html;   http://www.jstor.
org/stable/686775.

Mellor:1969:R

[Mel69] D. H. Mellor. Reviews. British Journal for the Philosophy of Science, 20(4):366–371, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/366.full.pdf+
html;   http://www.jstor.
org/stable/686268.

Mellor:1970:IV

[Mel70] D. H. Mellor. Index 1950–69 volumes 1–20. British Journal for the Philosophy of Science, 21(1):1–80, February 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686472.

Mellor:1977:NK

[Mel77] D. H. Mellor. Natural kinds. British Journal for the Philosophy of Science, 28(4):299– 312, December 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        28/4/299.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686875.

Mellor:1980:TCS

[Mel80] D. H. Mellor. On things and causes in spacetime. British Journal for the Philosophy of Science, 31(3):282–288, September 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/31/3/282.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686922.

Mellor:1983:R

[Mel83] D. H. Mellor. Reviews. British Journal for the Philosophy of Science, 34(1):97–104, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/1/97.full.pdf+
html;   http://www.jstor.
org/stable/686940.

Mellor:1984:R

[Mel84] D. H. Mellor. Reviews. British Journal for the Philosophy of Science, 35(2):175–179, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/175.full.pdf+
html;   http://www.jstor.
org/stable/686968.

Mellor:1990:RBB

[Mel90] D. H. Mellor. R. B. Braithwaite (1900–1990). British Journal for the Philosophy of Science, 41(4):579–580, December 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/579.full.pdf+html;
http://www.jstor.org/
stable/687867.

Melia:1999:BRB

[Mel99a] J. Melia. Book review: Realistic rationalism. Jerrold J. Katz. British Journal for the Philosophy of Science, 50(3): 475–477, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        50/3/475.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072244.

Melia:1999:HHT

[Mel99b] J. Melia. Holes, Haecceitism and two conceptions of determinism. British Journal for the Philosophy of Science, 50(4): 639–664, December 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/4/639.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072260.

Melia:2000:BRB

[Mel00] J. Melia. Book review: The Worlds of Possibility. C. S. Chihara. British Journal for the Philosophy of Science, 51 (2):333–337, June 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/2/333.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541809.

Menger:1954:VMN

[Men54] Karl Menger. On variables in mathematics and in natural science. British Journal for the Philosophy of Science, V(18): 134–142, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/18/134.full.pdf+html;

http://www.jstor.org/
stable/685170.

Mercier:1988:R

[Mer88] André Mercier. Review. British Journal for the Philosophy of Science, 39(3):415–419, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/415.full.pdf+
html;   http://www.jstor.
org/stable/687220.

Meyer:2021:NMO

[Mey21] Russell Meyer. The non-mechanistic option: Defending dynamical explanations. British Journal for the Philosophy of Science, 71(3): 959–985, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Meltzer:1965:TFP

[MG65] B. Meltzer and I. J. Good. Two forms of the prediction paradox. British Journal for the Philosophy of Science, XVI(61):50–51, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/61/50.full.pdf+
html;   http://www.jstor.
org/stable/686139.

Meek:1994:CI

[MG94] Christopher Meek and Clark Glymour. Conditioning and intervening. British Journal for the Philosophy of Science, 45 (4):1001–1021, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/4/1001.full.pdf+html;
http://www.jstor.org/
stable/687618.

Mitchell:2017:AFY

[MG17] Sandra D. Mitchell and Angela M. Gronenborn. After fifty years, why are protein X-ray crystallographers still in business? British Journal for the Philosophy of Science, 68(3):703–723, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/3/703/3044198.

Miller:2023:MSE

[MHB^{+}23] Michael E. Miller, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Matteo Colombo, Peggy Seriès, Jeffrey A. Barrett, Brian Skyrms, and Porter Williams. Mathematical structure and empirical content. British Journal for the Philosophy of Science, 74 (2):511–532, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Michalos:1966:EUC

[Mic66] Alex C. Michalos. Estimated utility and corroboration. British Journal for the Philosophy of Science, XVI(64):327–331, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/64/327.full.
pdf+html;   http://www.
jstor.org/stable/686833.

Michod:1981:PHE

[Mic81] Richard E. Michod. Positive heuristics in evolutionary biology. British Journal for the Philosophy of Science, 32(1):1–36, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/32/1/1.full.pdf+
html;   http://www.jstor.
org/stable/687384.

Michell:1994:NQR

[Mic94] Joel Michell. Numbers as quantitative relations and the traditional theory of measurement. British Journal for the Philosophy of Science, 45(2):389–406, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/389.full.pdf+
html;   http://www.jstor.
org/stable/687672.

Midgley:1985:R

[Mid85] Mary Midgley. Reviews. British Journal for the Philosophy of Science, 36(4):459–461, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/4/459.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687404.

Mikkelson:2004:DWW

[Mik04] Jeffrey M. Mikkelson. Dissolving the wine/water paradox. British Journal for the Philosophy of Science, 55(1):137–145, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/1/137.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541838.

Miller:1951:R

[Mil51] A. R. Miller. Reviews. British Journal for the Philosophy of Science, I(4):335–336, February 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/I/4/335.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685102.

Miles:1957:DBM

[Mil57] T. R. Miles. On the difference between the men and machines. British Journal for the Philosophy of Science, VII(28): 277–292, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/28/277.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685108.

Miller:1966:CCP

[Mil66a] David Miller. On a so-called socalled paradox: a reply to Professor J. L. Mackie. British Journal for the Philosophy of Science, 17 (2):147–149, August 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/2/147.full.pdf+html;
http://www.jstor.org/
stable/686347.

Miller:1966:PI

[Mil66b] David Miller. A paradox of information. British Journal for the Philosophy of Science, 17(1):59–61, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686404.

Milligan:1967:R

[Mil67]  D. E. Milligan.  Reviews.

British Journal for the Philosophy of Science, 18(2):157–160, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/2/157.full.pdf+
html;   http://www.jstor.
org/stable/686585.

Miller:1968:SNR

[Mil68] David Miller. The straight and narrow rule of induction. British Journal for the Philosophy of Science, 19(2):145–152, August 1968. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/145.full.pdf+
html;   http://www.jstor.
org/stable/686792.

Milligan:1970:R

[Mil70]  D. E. Milligan.  Reviews.

British Journal for the Philosophy of Science, 21(2):214–215, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/214.full.pdf+
html;   http://www.jstor.
org/stable/686708.

Miller:1972:BA

[Mil72] David Miller. Back to aristotle? British Journal for the Philosophy of Science, 23 (1):69–78, February 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/1/69.full.pdf+html;
http://www.jstor.org/
stable/686223.

Miller:1974:LM

[Mil74a] Arthur I. Miller. On Lorentz’s methodology. British Journal for the Philosophy of Science, 25(1):29–45, March 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/1/29.full.pdf+html;

http://www.jstor.org/
stable/686501.

Miller:1974:CFT

[Mil74b] David Miller. On the comparison of false theories by their bases. British Journal for the Philosophy of Science, 25(2):178–188, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/2/178.full.pdf+
html;   http://www.jstor.
org/stable/686822.

Miller:1974:PQT

[Mil74c] David Miller. Popper’s qualitative theory of verisimilitude. British Journal for the Philosophy of Science, 25(2):166–177, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/2/166.full.pdf+
html;   http://www.jstor.
org/stable/686821.

Miller:1974:R

[Mil74d] David Miller. Reviews. British Journal for the Philosophy of Science, 25(3):291–294, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/3/291.full.pdf+
html;   http://www.jstor.
org/stable/686912.

Miller:1975:PPP

[Mil75] Richard W. Miller. Propensity:

Popper or Peirce? British Journal for the Philosophy of Science, 26(2):123–132, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/2/123.full.pdf+html;
http://www.jstor.org/
stable/686511.

Miller:1976:R

[Mil76a] David Miller. Reviews. British Journal for the Philosophy of Science, 27(3):292–295, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/292.full.pdf+
html;   http://www.jstor.
org/stable/686128.

Miller:1976:VR

[Mil76b] David Miller. Verisimilitude redeflated. British Journal for the Philosophy of Science, 27(4): 363–380, December 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/363.full.pdf+html;
http://www.jstor.org/
stable/686863.

Miller:1978:RSN

[Mil78] Arthur I. Miller. A reply to ’Some New Aspects of Relativity: Comments on Zahar’s Paper’. British Journal for the Philosophy of Science, 29(3):252– 256, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/252-b.full.pdf+html;
http://www.jstor.org/
stable/687014.

Milton:1982:R

[Mil82] J. R. Milton. Reviews. British Journal for the Philosophy of Science, 33(2):223–226, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/2/223.full.pdf+
html;   http://www.jstor.
org/stable/687283.

Milne:1983:NSI

[Mil83a] Peter Milne. A note on scale invariance. British Journal for the Philosophy of Science, 34 (1):49–55, March 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/1/49.full.pdf+html;
http://www.jstor.org/
stable/686933.

Milne:1983:R

[Mil83b] Peter Milne. Reviews. British Journal for the Philosophy of Science, 34(3):312–313, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/3/312.full.pdf+
html;   http://www.jstor.
org/stable/687331.

Milton:1983:R

[Mil83c] J. R. Milton. Reviews. British Journal for the Philosophy of Science, 34(2):193–195, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/2/193.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687454.

Milne:1984:R

[Mil84] Peter Milne. Reviews. British Journal for the Philosophy of Science, 35(1):95–100, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/35/1/95.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687566.

Milne:1985:NPP

[Mil85a] Peter Milne. A note on Popper, propensities, and the twoslit experiment. British Journal for the Philosophy of Science, 36(1):66–70, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/1/66.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687157.

Milton:1985:R

[Mil85b] J. R. Milton. Reviews. British Journal for the Philosophy of Science, 36(1):89–93, March 1985. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/89.full.pdf+
html;   http://www.jstor.
org/stable/687162.

Milne:1986:HFM

[Mil86a] Peter Milne. Hartry Field on measurement and intrinsic explanation. British Journal for the Philosophy of Science, 37(3): 340–346, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687291.

Milne:1986:RU

[Mil86b] Peter Milne. Review: [untitled]. British Journal for the Philosophy of Science, 37(1): 123–126, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687004.

Milton:1987:IBH

[Mil87a] J. R. Milton. Induction before Hume. British Journal for the Philosophy of Science, 38 (1):49–74, March 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/49.full.pdf+html;
http://www.jstor.org/
stable/687079.

Milton:1987:R

[Mil87b] J. R. Milton. Reviews. British Journal for the Philosophy of Science, 38(1):137–139, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/1/137.full.pdf+
html;   http://www.jstor.
org/stable/687091.

Milne:1989:FII

[Mil89]  Peter Milne.  Frege, infor-

mative identities, and logicism. British Journal for the Philosophy of Science, 40(2):155–166, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/155.full.pdf+
html;   http://www.jstor.
org/stable/687506.

Miller:1990:RPI

[Mil90a] David Miller. A restoration of Popperian inductive scepticism. British Journal for the Philosophy of Science, 41(1):137–139, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/1/137.full.pdf+
html;   http://www.jstor.
org/stable/688008.

Miller:1990:SLM

[Mil90b] David W. Miller. Some logical mensuration. British Journal for the Philosophy of Science, 41(2):281–290, June 1990.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/2/281.full.pdf+html;
http://www.jstor.org/
stable/687776.

Milne:1994:PM

[Mil94a] Peter Milne. The physicalization of mathematics. British Journal for the Philosophy of Science, 45 (1):305–340, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/305.full.pdf+html;
http://www.jstor.org/
stable/687976.

Milton:1994:R

[Mil94b] J. R. Milton. Review. British Journal for the Philosophy of Science, 45(3):927–928, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/927.full.pdf+
html;   http://www.jstor.
org/stable/687804.

Milton:1996:R

[Mil96] J. R. Milton. Review. British Journal for the Philosophy of Science, 47(1):125–128, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/125.full.pdf+
html;   http://www.jstor.
org/stable/1215830.

Milne:1997:BFL

[Mil97a] Peter Milne. Bruno de Finetti and the logic of conditional events. British Journal for the Philosophy of Science, 48 (2):195–232, June 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        48/2/195.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687745.

Milton:1997:R

[Mil97b] J. R. Milton. Review. British Journal for the Philosophy of Science, 48(2):302–305, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/2/302.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687756.

Milne:2001:DIE

[Mil01] Peter Milne. The design inference: eliminating chance through small probabilities. British Journal for the Philosophy of Science, 52(4):801–808, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/4/801.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541920.

Miller:2005:N

[Mil05] David Miller. Note. British Journal for the Philosophy of Science, 56(2):419, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/2/419.full.pdf+html;
http://www.jstor.org/
stable/3541710.

Millstein:2006:NSP

[Mil06] Roberta L. Millstein. Natural selection as a population-level causal process. British Journal for the Philosophy of Science, 57(4):627–653, December 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/4/627.full.pdf+html;
http://www.jstor.org/
stable/4489089.

Miller:2018:HTA

[Mil18] Michael E. Miller. Haag’s theorem, apparent inconsistency, and the empirical adequacy of quantum field theory. British Journal for the Philosophy of Science, 69(3):801–820, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/3/801/2669796.

Millhouse:2019:SCP

[Mil19] Tyler Millhouse. A simplicity criterion for physical computation. British Journal for the Philosophy of Science, 70(1):153–178, March 2019. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/1/153/4098119.

Milkowski:2023:CTS

[Mił23] Marcin Miłkowski. Correspondence theory of semantic information. British Journal for the Philosophy of Science, 74(2): 485–510, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mintz:1966:R

[Min66]  Samuel I. Mintz.  Reviews.

British Journal for the Philosophy of Science, 17(3):259– 261, November 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/3/259.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686765.

Mirowski:2004:EES

[Mir04] Philip Mirowski. The effortless economy of science? Science and cultural theory. Duke University Press, Durham, NC, USA, 2004. ISBN 0-8223-3310-4 (cloth), 0-8223-3322-8 (paperback). 463 pp. LCCN Q180.55.E25 M57

                                        2004. URL http://www.loc.
                                        gov/catdir/toc/ecip0415/
                                        2004003274.html.

Miscevic:1997:R

[Mis97] Nenad Miscevic. Review. British Journal for the Philosophy of Science, 48(4):603–605, December 1997. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/603.full.pdf+
html;   http://www.jstor.
org/stable/688103.

Mitchell:2009:UTS

[Mit09] Sandra D. Mitchell. Unsimple truths: science, complexity, and policy. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-53262-3 (hardcover). x + 149 pp. LCCN Q175.32.C65 M585 2009.

Mizrahi:2023:TVS

[Miz23]  Moti Mizrahi.  Theoretical

virtues in scientific practice: an empirical study. British Journal for the Philosophy of Science, 73 (4):879–902, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Money-Kyrle:1956:WUW

[MK56] R. E. Money-Kyrle. The world of the unconscious and the world of commonsense. British Journal for the Philosophy of Science, VII(25):86–96, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/25/86.full.pdf+html;
http://www.jstor.org/
stable/685938.

Machamer:1975:R

[ML75] Peter Machamer and Stephen Lunsford. Reviews. British Journal for the Philosophy of Science, 26(1):81–82, March 1975.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/1/81.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686337.

Menon:2021:CCR

[MLR21] Tushar Menon, Niels Linnemann, and James Read. Clocks and chronogeometry: Rotating spacetimes and the relativistic null hypothesis. British Journal for the Philosophy of Science, 71 (4):1287–1317, ???? 2021. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Molland:1981:R

[Mol81] A. G. Molland. Reviews. British Journal for the Philosophy of Science, 32(3):306–309, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/3/306.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687595.

Mollo:2022:ACP

[Mol22a] Dimitri Coelho Mollo. Against computational perspectivalism. British Journal for the Philosophy of Science, 72(4):1129–1153, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Molter:2022:BSG

[Mol22b] Daniel J. Molter. Bivalent selection and graded Darwinian individuality. British Journal for the Philosophy of Science, 73(1): 73–84, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Molinini:2023:UEP

[Mol23] Daniele Molinini. The unreasonable effectiveness of physics in mathematics. British Journal for the Philosophy of Science, 74(4):853–874, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Montuschi:1994:R

[Mon94] Eleonora Montuschi. Review.

British Journal for the Philosophy of Science, 45(1):375–386, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/375.full.pdf+
html;   http://www.jstor.
org/stable/687979.

Monton:2006:GFT

[Mon06] Bradley Monton. God, finetuning, and the problem of old evidence. British Journal for the Philosophy of Science, 57 (2):405–424, June 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/405.full.pdf+html;
http://www.jstor.org/
stable/3873473.

Moor:1978:TMC

[Moo78] James H. Moor. Three myths of computer science. British Journal for the Philosophy of Science, 29(3):213–222, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/3/213.full.pdf+
html;   http://www.jstor.
org/stable/687009.

Moore:1983:DBS

[Moo83] P. G. Moore. A Dutch Book and subjective probabilities. British Journal for the Philosophy of Science, 34(3):263–266, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/3/263.full.pdf+
html;   http://www.jstor.
org/stable/687323.

Moore:1996:R

[Moo96] A. W. Moore. Review. British Journal for the Philosophy of Science, 47(4):621–626, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/621.full.pdf+
html;   http://www.jstor.
org/stable/687928.

Morrison:1960:PIC

[Mor60] Paul G. Morrison. On partial identity of cause and effect. British Journal for the Philosophy of Science, XI(41):42–49, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/42.full.pdf+
html;   http://www.jstor.
org/stable/685818.

Morton:1985:R

[Mor85] Adam Morton. Reviews. British Journal for the Philosophy of Science, 36(3):341–344, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/3/341.full.pdf+
html;   http://www.jstor.
org/stable/687579.

Morrison:1986:MRB

[Mor86] Margaret Morrison. More on the relationship between technically good and conceptually important experiments: a case study. British Journal for the Philosophy of Science, 37(1): 101–115, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687000.

Mormann:1988:AFT

[Mor88] Thomas Mormann. Are all false theories equally false?: a remark on David Miller’s problem and geometric conventionalism. British Journal for the Philosophy of Science, 39(4): 505–519, November 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/505.full.pdf+html;
http://www.jstor.org/
stable/687435.

Morris:1989:CP

[Mor89] W. T. Morris. Conventionalism in physics. British Journal for the Philosophy of Science, 40 (1):135–136, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/1/135.full.pdf+html;
http://www.jstor.org/
stable/687470.

Morrison:1990:URI

[Mor90] Margaret Morrison. Unification, realism and inference. British Journal for the Philosophy of Science, 41(3):305–332, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/3/305.full.pdf+
html;   http://www.jstor.
org/stable/687823.

Mormann:1993:R

[Mor93a] Thomas Mormann. Review. British Journal for the Philosophy of Science, 44(3):567– 575, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/567.full.pdf+html;
http://www.jstor.org/
stable/688026.

Morton:1993:MMQ

[Mor93b] Adam Morton. Mathematical models: Questions of trustworthiness. British Journal for the Philosophy of Science, 44(4): 659–674, December 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/4/659.full.pdf+html;
http://www.jstor.org/
stable/688037.

Morrison:1994:CCF

[Mor94] Margaret Morrison. Causes and contexts: the foundations of laser theory. British Journal for the Philosophy of Science, 45 (1):127–151, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/127.full.pdf+html;
http://www.jstor.org/
stable/687964.

Morton:1997:RU

[Mor97] Adam Morton. Review: [untitled]. British Journal for the Philosophy of Science, 48 (2):310–312, June 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687758.

Morton:1999:BRB

[Mor99] A. Morton. Book review: Decision theory as philosophy. Mark Kaplan. British Journal for the Philosophy of Science, 50(3):

505–507, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/505.full.pdf+html;
http://www.jstor.org/
stable/40072251.

Morton:2000:SEE

[Mor00] Adam Morton. Saving epistemology from the epistemologists: Recent work in the theory of knowledge. British Journal for the Philosophy of Science, 51(4):685–704, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541613.

Morrison:2002:MPP

[Mor02] Margaret Morrison. Modelling populations: Pearson and Fisher on Mendelism and biometry. British Journal for the Philosophy of Science, 53(1):39–68, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/1/39.full.pdf+
html;   http://www.jstor.
org/stable/3541640.

Morrison:2006:UEE

[Mor06] Margaret Morrison. Unification, explanation and explaining unity: the Fisher–Wright controversy. British Journal for the Philosophy of Science, 57 (1):233–245, March 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/233.full.pdf+html;
http://www.jstor.org/
stable/3541660.

Morris:2009:DFR

[Mor09] Kevin Morris. Does functional reduction need bridge laws? a response to Marras. British Journal for the Philosophy of Science, 60(3):647–657, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/3/647.full.pdf+html;
http://www.jstor.org/
stable/25592023.

Morris:2012:DLU

[Mor12] Kevin Morris. A defense of lucky understanding. British Journal for the Philosophy of Science, 63(2):357–371, June 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/357.full.pdf+html.

Morganti:2013:CSM

[Mor13] Matteo Morganti. Combining science and metaphysics: contemporary physics, conceptual revision, and common sense. New directions in the philosophy of science. Palgrave Macmillan, New York, NY, USA, 2013. ISBN 1-137-00268-9 (hardcover). xi + 211 pp. LCCN B67 .M67 2013.

Morales:2024:ISD

[Mor24] Jorge Morales. Introspection is signal detection. British Journal for the Philosophy of Science, 75(1):99–126, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Moss:1984:R

[Mos84] J. M. B. Moss. Reviews. British Journal for the Philosophy of Science, 35(3):307–310, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/35/3/307.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687483.

Moss:1985:MPC

[Mos85] J. M. B. Moss. The mathematical philosophy of Charles Parsons. British Journal for the Philosophy of Science, 36(4): 437–457, December 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/4/437.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687403.

Moser:1986:BRS

[Mos86] Paul K. Moser. Book review: Stroud, Barry [1984]: The Significance of Philosophical Scepticism. Clarendon Press. xiv + 277 pp. British Journal for the Philosophy of Science, 37 (2):235–238, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686985.

Moser:1988:OO

[Mos88a] Paul K. Moser. Observation and objectivity. British Journal for the Philosophy of Science, 39(4):551–561, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/551.full.pdf+html;
http://www.jstor.org/
stable/687441.

Moser:1988:R

[Mos88b] Paul K. Moser. Review. British Journal for the Philosophy of Science, 39(1):131–140, March 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/1/131.full.pdf+
html;   http://www.jstor.
org/stable/687119.

Moser:1991:JNS

[Mos91] Paul K. Moser. Justification in the natural sciences. British Journal for the Philosophy of Science, 42(4):557–575, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/4/557.full.pdf+
html;   http://www.jstor.
org/stable/687710.

Mott:1983:R

[Mot83] P. L. Mott. Reviews. British Journal for the Philosophy of Science, 34(3):306–310, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/3/306.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687329.

Mounce:1985:R

[Mou85] H. O. Mounce. Reviews. British Journal for the Philosophy of Science, 36(3):344–346, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/3/344.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687580.

Menzies:1993:CSQ

[MP93] Peter Menzies and Huw Price. Causation as a secondary quality. British Journal for the Philosophy of Science, 44(2):187–203, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/2/187.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687643.

Morgan:2008:ESC

[MP08] Gregory J. Morgan and W. Brad Pitts. Evolution without species: the case of mosaic bacteriophages. British Journal for the Philosophy of Science, 59(4): 745–765, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/745.full.pdf+html;
http://www.jstor.org/
stable/40072310.

Muller:2018:DD

[MP18] Thomas Müller and Tomasz Placek. Defining determinism. British Journal for the Philosophy of Science, 69(1):215–252, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/215/2669643.

Mossio:2022:CFA

[MP22] Matteo Mossio and Gaëlle Pontarotti. Conserving functions across generations: Heredity in light of biological organization. British Journal for the Philosophy of Science, 73(1):249–278, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Munson:1994:TNN

[MR94] Ronald Munson and Paul Roth.

Testing normative naturalism: the problem of scientific medicine. British Journal for the Philosophy of Science, 45(2):571–584, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/2/571.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687683.

Milne:1950:O

[MS50] E. A. Milne and F. R. S. Obituary. British Journal for the Philosophy of Science, I(3):256– 255, November 1950. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/3/256.full.pdf+html.

Mckinsey:1955:NIC

[MS55] J. C. C. Mckinsey and Patrick Suppes. On the notion of invariance in classical mechanics. British Journal for the Philosophy of Science, V(20): 290–302, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/20/290.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685731.

Mayo:2006:STB

[MS06a] Deborah G. Mayo and Aris Spanos. Severe testing as a basic concept in a Neyman– Pearson philosophy of induction. British Journal for the Philosophy of Science, 57(2):323–357, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/57/2/323.full.pdf+

html;   http://www.jstor.
org/stable/3873470.

Melia:2006:RTC

[MS06b] Joseph Melia and Juha Saatsi.

Ramseyfication and theoretical content. British Journal for the Philosophy of Science, 57(3): 561–585, September 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/3/561.full.pdf+html;
http://www.jstor.org/
stable/3873483.

Muller:2007:SQM

[MS07] F. A. Muller and M. P. Seevinck.

Is standard quantum mechanics technologically inadequate? British Journal for the Philosophy of Science, 58(3):595– 604, September 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/3/595.full.pdf+html;
http://www.jstor.org/
stable/30115189.

Muller:2008:DF

[MS08] F. A. Muller and Simon Saunders. Discerning fermions. British Journal for the Philosophy of Science, 59(3):499– 548, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/499.full.pdf+html;

http://www.jstor.org/
stable/40072297.

Mossio:2009:OAB

[MSM09] Matteo Mossio, Cristian Saborido, and Alvaro Moreno. An organizational account of biological functions. British Journal for the Philosophy of Science, 60(4):813–841, December 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/813.full.pdf+html;
http://www.jstor.org/
stable/25592036.

Mumford:2013:MS

[MT13] Stephen Mumford and Matthew Tugby. Metaphysics and science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967452-3 (print), 0-19-176001-3 (e-book). LCCN Q175.32.

Martinez:2024:WEB

[MTH^{+}24] Marco  Martinez,  David

Teira, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Thomas Müller, Tomasz Placek, LUDWIG VON BERTALANFFY, Joseph Y. Halpern, Judea Pearl, Matteo Mossio, Cristian Saborido, Alvaro Moreno, and Torsten Wilholt. Why experimental balance is still a reason to randomize. British Journal for the Philosophy of Science, 75(2):519– 535, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Mueller:1969:EEA

[Mue69] Ian Mueller. Euclid’s Elements and the axiomatic method. British Journal for the Philosophy of Science, 20(4):289–309, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/289.full.pdf+
html;   http://www.jstor.
org/stable/686258.

Mulaik:1991:FAI

[Mul91] Stanley A. Mulaik. Factor analysis, information-transforming instruments, and objectivity: a reply and discussion. British Journal for the Philosophy of Science, 42(1):87–100, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/1/87.full.pdf+html;
http://www.jstor.org/
stable/687608.

Muller:2001:SCC

[Mul01] F. A. Muller. Sets, classes, and categories. British Journal for the Philosophy of Science, 52(3):539–573, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/539.full.pdf+html;
http://www.jstor.org/
stable/3541928.

Muller:2005:DBS

[Mul05a] F. A. Muller. The deep black sea: Observability and modality afloat. British Journal for the Philosophy of Science, 56 (1):61–99, March 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/1/61.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541850.

Muller:2005:PTC

[Mül05b] Thomas Müller. Probability theory and causation: a branching space–times analysis. British Journal for the Philosophy of Science, 56(3):487–520, September 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/56/3/487.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541716.

Mumford:2004:LN

[Mum04] Stephen Mumford. Laws in nature. Routledge studies in twentieth century philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2004. ISBN 0-415-31128-4. xv + 230 pp. LCCN B105.L3 M86

                                        2004. URL http://www.loc.
                                        gov/catdir/toc/ecip0413/
                                        2004000286.html.

Mumford:2013:BRM

[Mum13] Stephen Mumford. Book review: Max Kistler, Causation and Laws of Nature: Routledge Studies in Contemporary Philosophy. British Journal for the Philosophy of Science, 64 (1):223–227, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/223.full.pdf+html.

Mundle:1953:MDT

[Mun53] C. W. K. Mundle. Mr Dobbs’ two-dimensional theory of time. British Journal for the Philosophy of Science, IV(16):331– 337, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/331.full.pdf+html.

Mundle:1954:MDT

[Mun54a] C. W. K. Mundle. Mr Dobbs’ two-dimensional theory of time. British Journal for the Philosophy of Science, IV(16):331– 337, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686083.

Munitz:1954:CNC

[Mun54b] Milton K. Munitz. Creation and the “New” cosmology. British Journal for the Philosophy of Science, V(17):32–46, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        V/17/32.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685950.

Mundle:1956:R

[Mun56] C. W. K. Mundle. Reviews. British Journal for the Philosophy of Science, VI(24):336– 338, February 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/24/336.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685415.

Mundle:1957:R

[Mun57] C. W. K. Mundle. Reviews. British Journal for the Philosophy of Science, VIII(29):73–76, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/29/73.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685385.

Munitz:1962:LC

[Mun62] Milton K. Munitz. The logic of cosmology. British Journal for the Philosophy of Science, XIII(49):34–50, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/49/34.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685222.

Mundy:1986:PCM

[Mun86a] Brent Mundy. The physical content of Minkowski geometry. British Journal for the Philosophy of Science, 37 (1):25–54, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686996.

Mundy:1986:SRQ

[Mun86b] Brent Mundy. Special Relativity and quantum measurement. British Journal for the Philosophy of Science, 37 (2):207–212, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686978.

Mundy:1989:DAC

[Mun89] Brent Mundy. Distant action in classical electromagnetic theory. British Journal for the Philosophy of Science, 40(1):39–68, March 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/1/39.full.pdf+
html;   http://www.jstor.
org/stable/687463.

Munzer:1994:TCI

[Mun94] Stephen R. Munzer. Transplantation, chemical inheritance, and the identity of organs. British Journal for the Philosophy of Science, 45(2):555–570, June 1994. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/555.full.pdf+
html;   http://www.jstor.
org/stable/687682.

Murdoch:1974:R

[Mur74] John E. Murdoch. Reviews. British Journal for the Philosophy of Science, 25(1):89–91, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/89.full.pdf+
html;   http://www.jstor.
org/stable/686506.

Murphy:1990:SRP

[Mur90] Nancey Murphy. Scientific realism and postmodern philosophy. British Journal for the Philosophy of Science, 41(3):291– 303, September 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/3/291.full.pdf+html.

Murphy:2006:PSI

[Mur06] Dominic (Dominic Paul) Murphy. Psychiatry in the scientific image. Philosophical psychopathology. Disorders in mind. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-13455-1 (hardcover). xi + 410 pp. LCCN RC437.5.M87 2006; RC437.5.M87.

Musgrave:1969:R

[Mus69] A. E. Musgrave. Reviews. British Journal for the Philosophy of Science, 20(1):92–94, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/1/92.full.pdf+html;
http://www.jstor.org/
stable/686726.

Musgrave:1971:KST

[Mus71] Alan E. Musgrave. Kuhn’s second thoughts. British Journal for the Philosophy of Science, 22 (3):287–297, August 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/3/287.full.pdf+html;
http://www.jstor.org/
stable/686749.

Musgrave:1974:LVH

[Mus74] Alan Musgrave. Logical versus historical theories of confirmation. British Journal for the Philosophy of Science, 25(1):1–23, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/1.full.pdf+
html;   http://www.jstor.
org/stable/686499.

Musgrave:1977:LR

[Mus77] Alan Musgrave. Logicism revisited. British Journal for the Philosophy of Science, 28(2):99–127, June 1977. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/2/99.full.pdf+
html;   http://www.jstor.
org/stable/686628.

Monton:2003:CEM

[MvF03] Bradley Monton and Bas C. van Fraassen. Constructive empiricism and modal nominalism. British Journal for the Philosophy of Science, 54(3):405– 422, September 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/3/405.full.pdf+html;
http://www.jstor.org/
stable/3541792.

Mayo-Wilson:2014:LPC

[MW14] Conor Mayo-Wilson. The limits of piecemeal causal inference. British Journal for the Philosophy of Science, 65(2):213–249, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/65/2/213.

Marczyk:2015:CCC

[MW15] Michal Marczyk and Leszek Wroński. Completion of the causal completability problem. British Journal for the Philosophy of Science, 66(2):307–326, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/2/307.

Myrvold:2003:RQB

[Myr03] Wayne C. Myrvold. Relativistic quantum becoming. British Journal for the Philosophy of Science, 54(3):475–500, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/3/475.full.pdf+
html;   http://www.jstor.
org/stable/3541796.

Myrvold:2009:CC

[Myr09] Wayne C. Myrvold. Chasing chimeras. British Journal for the Philosophy of Science, 60(3): 635–646, September 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/3/635.full.pdf+html;
http://www.jstor.org/
stable/25592022.

Myrvold:2011:NCQ

[Myr11] Wayne C. Myrvold. Nonseparability, classical, and quantum. British Journal for the Philosophy of Science, 62(2): 417–432, June 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/417.full.pdf+html.

Ney:2013:WFE

[NA13] Alyssa Ney and David Z. Albert, editors. The wave function: essays on the metaphysics of quantum mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-979080-9 (hardcover), 0-19-979054-X (paperback). xii + 229 pp. LCCN QC174.125 .W38 2013.

Nadler:1999:BRB

[Nad99] S. Nadler. Book review: The causation debate in modern philosophy, 1637–1739. Kenneth Clatterbaugh. British Journal for the Philosophy of Science, 50(3): 501–504, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/501.full.pdf+html;
http://www.jstor.org/
stable/40072250.

Nado:2008:EMC

[Nad08] Jennifer Nado. Effects of moral cognition on judgments of intentionality. British Journal for the Philosophy of Science, 59(4):709–731, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/709.full.pdf+html;
http://www.jstor.org/
stable/40072308.

Nagel:1959:R

[Nag59] Ernest Nagel. Reviews. British Journal for the Philosophy of Science, X(39):245–249, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/X/39/245.full.pdf+
html;   http://www.jstor.
org/stable/686062.

Najenson:2024:MSM

[Naj24]  Jonathan Najenson.  Mem-

ory systems and the mnemic character of procedural memory. British Journal for the Philosophy of Science, 75(2):299– 323, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Nanay:2013:SS

[Nan13] Bence Nanay. Singularist semirealism. British Journal for the Philosophy of Science, 64 (2):371–394, June 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/371.full.pdf+html.

Narliker:1965:DT

[Nar65] J. V. Narliker. The direction of time. British Journal for the Philosophy of Science, XV(60):281–285, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/60/281.full.pdf+html;
http://www.jstor.org/
stable/686535.

Narlikar:1968:R

[Nar68] J. V. Narlikar. Reviews. British Journal for the Philosophy of Science, 19(3):267–268, November 1968. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/267.full.pdf+
html;   http://www.jstor.
org/stable/686251.

Narlikar:1970:R

[Nar70] J. V. Narlikar. Reviews. British Journal for the Philosophy of Science, 21(2):216–217, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/216.full.pdf+
html;   http://www.jstor.
org/stable/686710.

Nartonis:1974:QFF

[Nar74] D. K. Nartonis. Quantum fact and fiction. British Journal for the Philosophy of Science, 25(4):329–332, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/4/329.full.pdf+html;
http://www.jstor.org/
stable/686459.

Nathan:2014:CC

[Nat14] Marco J. Nathan. Causation by concentration. British Journal for the Philosophy of Science, 65(2):191–212, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/191.

Nathan:2017:UE

[Nat17] Marco J. Nathan. Unificatory explanation. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/1/163/2995271/
Unificatory-Explanation.

Neander:1995:ECA

[Nea95a]  Karen Neander.  Explaining

complex adaptations: a reply to Sober’s ‘reply to Neander’. British Journal for the Philosophy of Science, 46(4):583– 587, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/4/583.full.pdf+html;
http://www.jstor.org/
stable/687899.

Neander:1995:PTL

[Nea95b] Karen Neander. Pruning the tree of life. British Journal for the Philosophy of Science, 46 (1):59–80, March 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/1/59.full.pdf+html;
http://www.jstor.org/
stable/687633.

Needham:1966:RU

[Nee66] R. M. Needham. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):80–83, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686409.

Needham:2010:ST

[Nee10] Paul Needham. Substance and time. British Journal for the Philosophy of Science, 61(3):485– 512, September 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/3/485.full.pdf+html.

Needham:2013:PCT

[Nee13] Paul Needham. Process and change: From a thermodynamic perspective. British Journal for the Philosophy of Science, 64 (2):395–422, June 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/2/395.full.pdf+html.

Needham:2017:DSS

[Nee17] Paul Needham. Determining sameness of substance. British Journal for the Philosophy of Science, 68(4):953–979, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/4/953/3003366.

Nelkin:1994:PR

[Nel94] Norton Nelkin. Phenomena and representation. British Journal for the Philosophy of Science, 45(2):527–547, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/527.full.pdf+html;
http://www.jstor.org/
stable/687680.

Nelson:1996:CEL

[Nel96] David E. Nelson. Confirmation, explanation, and logical strength. British Journal for the Philosophy of Science, 47(3): 399–413, September 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/3/399.full.pdf+html;
http://www.jstor.org/
stable/687985.

Nemeth:2003:SOS

[Nem03] Elisabeth Nemeth. The social origins of science. British Journal for the Philosophy of Science, 54(3):515–520, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/3/515.full.pdf+html;
http://www.jstor.org/
stable/3541799.

Nerlich:1971:PAS

[Ner71]  Graham Nerlich.  A prob-

lem about sufficient conditions.

British Journal for the Philosophy of Science, 22(2):161–170, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/161.full.pdf+
html;   http://www.jstor.
org/stable/686485.

Nerlich:1978:R

[Ner78]  Graham Nerlich.  Reviews.

British Journal for the Philosophy of Science, 29(3):295– 298, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/295.full.pdf+html;
http://www.jstor.org/
stable/687023.

Nerlich:1979:WCG

[Ner79] Graham Nerlich. What can geometry explain? British Journal for the Philosophy of Science, 30(1):69–83, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/69.full.pdf+html;
http://www.jstor.org/
stable/687417.

Nerlich:1982:SRB

[Ner82] Graham Nerlich. Special Relativity is not based on causality. British Journal for the Philosophy of Science, 33(4): 361–388, December 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/4/361.full.pdf+html;
http://www.jstor.org/
stable/687169.

Neta:2008:WED

[Net08] Ram Neta. What evidence do you have? British Journal for the Philosophy of Science, 59(1):89–119, March 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/1/89.full.pdf+html;
http://www.jstor.org/
stable/30116995.

Newburgh:1973:CIM

[New73] Ronald G. Newburgh. Comments on ‘The Incompatibility of Mach’s Principle and the Principle of Equivalence in Current Gravitation Theory’. British Journal for the Philosophy of Science, 24(3):263–264, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/3/263.full.pdf+
html;   http://www.
jstor.org/stable/686230. See

[WY72, WY73].

Newman:2017:TUS

[New17] Mark P. Newman. Theoretical understanding in science. British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/2/571/3044193/
                                        Theoretical-Understanding-
                                        in-Science.

Ney:2009:PCD

[Ney09] Alyssa Ney. Physical causation and difference-making. British Journal for the Philosophy of Science, 60(4):737–764, December 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/60/4/737.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/25592033.

Ney:2011:BRT

[Ney11]  Alyssa Ney.  Book review:

Tim Maudlin, The Metaphysics Within Physics. British Journal for the Philosophy of Science, 62(3):683–689, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/3/683.full.pdf+html.

Ney:2014:BRJ

[Ney14] Alyssa Ney. Book review: John Heil, The Universe As We Find It. British Journal for the Philosophy of Science, 65(4):881– 886, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/4/881.

Nguyen:2021:GAR

[Ngu21] James Nguyen. It’s not a game:

Accurate representation with toy models. British Journal for the Philosophy of Science, 71(3): 1013–1041, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Niaz:2005:ACN

[Nia05] Mansoor Niaz. An appraisal of the controversial nature of the oil drop experiment: Is closure possible? British Journal for the Philosophy of Science, 56(4):681–702, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/681.full.pdf+html;
http://www.jstor.org/
stable/3541863.

Nickles:1987:LHE

[Nic87] Thomas Nickles. Lakatosian heuristics and epistemic support. British Journal for the Philosophy of Science, 38(2):181–205, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/2/181.full.pdf+
html;   http://www.jstor.
org/stable/687047.

Nickles:1995:R

[Nic95] Thomas Nickles. Review. British Journal for the Philosophy of Science, 46(2):261–266, June 1995. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/261.full.pdf+
html;   http://www.jstor.
org/stable/687817.

Nidditch:1971:R

[Nid71] P. H. Nidditch. Reviews. British Journal for the Philosophy of Science, 22(1):64–65, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/64.full.pdf+html;
http://www.jstor.org/
stable/686650.

Nielsen:2021:DCS

[Nie21] Michael Nielsen. Deterministic convergence and strong regularity. British Journal for the Philosophy of Science, 71(4):1461– 1491, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Niiniluoto:1979:DTS

[Nii79] Ilkka Niiniluoto. Degrees of truthlikeness: from singular sentences to generalisations. British Journal for the Philosophy of Science, 30(4):371–376, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/4/371.full.pdf+
html;   http://www.jstor.
org/stable/687545.

Niiniluoto:1982:EVR

[Nii82] Ilkka Niiniluoto. On explicating verisimilitude: a reply to Oddie. British Journal for the Philosophy of Science, 33(3): 290–296, September 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/290.full.pdf+html;
http://www.jstor.org/
stable/687227.

Niiniluoto:1983:NFB

[Nii83] Ilkka Niiniluoto. Novel facts and Bayesianism. British Journal for the Philosophy of Science, 34(4):375–379, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/375.full.pdf+html;
http://www.jstor.org/
stable/687529.

Niiniluoto:1998:VTP

[Nii98] Ilkka Niiniluoto. Verisimilitude:

the third period. British Journal for the Philosophy of Science, 49(1):1–29, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/1/1.full.pdf+html;
http://www.jstor.org/
stable/688141.

Nissen:1970:CFT

[Nis70] Lowell Nissen. Canfield’s functional translation schema.

British Journal for the Philosophy of Science, 21(2):193–195, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/193.full.pdf+
html;   http://www.jstor.
org/stable/686699.

Niall:1983:R

[NM83] R. Niall and D. Martin. Reviews. British Journal for the Philosophy of Science, 34(2):195–197, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/195.full.pdf+
html;   http://www.jstor.
org/stable/687455.

Nola:1980:RU

[Nol80a] Robert Nola. Review: [untitled]. British Journal for the Philosophy of Science, 31 (1):91–98, March 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687257.

Nola:1980:R

[Nol80b] Robert Nola. Reviews. British Journal for the Philosophy of Science, 31(1):91–97, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/1/91.full.pdf+html.

Nola:1984:R

[Nol84] Robert Nola. Reviews. British Journal for the Philosophy of Science, 35(2):184–188, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/184.full.pdf+
html;   http://www.jstor.
org/stable/686970.

Nola:1987:SPT

[Nol87] Robert Nola. The status of Popper’s theory of scientific method. British Journal for the Philosophy of Science, 38(4):441– 480, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/441.full.pdf+html;
http://www.jstor.org/
stable/687353.

Nola:1989:R

[Nol89] Robert Nola. Review. British Journal for the Philosophy of Science, 40(3):419–427, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/3/419.full.pdf+
html;   http://www.jstor.
org/stable/687790.

Nola:1996:R

[Nol96] Robert Nola. Review. British Journal for the Philosophy of Science, 47(3):467–473, September 1996. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/3/467.full.pdf+
html;   http://www.jstor.
org/stable/687990.

Nolan:1997:QP

[Nol97] Daniel Nolan. Quantitative parsimony. British Journal for the Philosophy of Science, 48(3): 329–343, September 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/3/329.full.pdf+html;
http://www.jstor.org/
stable/688066.

Nolan:1999:FVO

[Nol99] D. Nolan. Is fertility virtuous in its own right? British Journal for the Philosophy of Science, 50(2):265–282, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/265.full.pdf+html;
http://www.jstor.org/
stable/40072223.

Noordhof:1997:MCF

[Noo97] Paul Noordhof. Making the change: the functionalist’s way. British Journal for the Philosophy of Science, 48(2):233–250, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/233.full.pdf+

html;   http://www.jstor.
org/stable/687746.

North:1966:R

[Nor66] J. D. North. Reviews. British Journal for the Philosophy of Science, XVI(64):341–344, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/64/341.full.
pdf+html;   http://www.
jstor.org/stable/686837.

North:1967:R

[Nor67] J. D. North. Reviews. British Journal for the Philosophy of Science, 17(4):341–343, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/341.full.pdf+
html;   http://www.jstor.
org/stable/686782.

North:1971:R

[Nor71] J. North. Reviews. British Journal for the Philosophy of Science, 22(2):207–208, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/2/207.full.pdf+html;
http://www.jstor.org/
stable/686494.

Norton:1986:QOW

[Nor86] John Norton. The quest for the one way velocity of light.

British Journal for the Philosophy of Science, 37(1):118–120, March 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/687002.

North:2002:WPA

[Nor02] Jill North. What is the problem about the time-asymmetry of thermodynamics? — a reply to Price. British Journal for the Philosophy of Science, 53 (1):121–136, March 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/121.full.pdf+html.

Norman:2004:RPS

[Nor04] Jesse Norman. Review: the philosophical status of diagrams. British Journal for the Philosophy of Science, 55(4):801– 805, December 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/801.full.pdf+html;
http://www.jstor.org/
stable/3541632.

Norton:2007:PD

[Nor07] John D. Norton. Probability disassembled. British Journal for the Philosophy of Science, 58(2):141–171, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
58/2/141.full.pdf+html;
http://www.jstor.org/
stable/30115222.

Norton:2008:WCR

[Nor08] John D. Norton. Why constructive Relativity fails. British Journal for the Philosophy of Science, 59(4):821–834, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/4/821.full.pdf+
html;   http://www.jstor.
org/stable/40072313.

Norton:2009:TIP

[Nor09] John D. Norton. Is there an independent principle of causality in physics? British Journal for the Philosophy of Science, 60(3):475–486, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/3/475.full.pdf+html;
http://www.jstor.org/
stable/25592015.

Norton:2019:DIC

[Nor19] John D. Norton. A demonstration of the incompleteness of calculi of inductive inference. British Journal for the Philosophy of Science, 70(4):1119– 1144, December 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1119/4823516.

Nissan-Rozen:2013:JCP

[NR13]  Ittay Nissan-Rozen.  Jeffrey

conditionalization, the principal principle, the desire as belief thesis, and Adams’s thesis. British Journal for the Philosophy of Science, 64(4): 837–850, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/4/837.full.pdf+html.

Nerlich:1967:PLN

[NS67] G. C. Nerlich and W. A. Suchting. Popper on law and natural necessity. British Journal for the Philosophy of Science, 18(3):233–235, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/3/233.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686594.

Newton-Smith:1969:R

[NS69] W. Newton-Smith. Reviews. British Journal for the Philosophy of Science, 20(3):281–283, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/281.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686429.

Newton-Smith:1978:R

[NS78] W. Newton-Smith. Reviews. British Journal for the Philosophy of Science, 29(3):289– 293, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/289.full.pdf+html;
http://www.jstor.org/
stable/687021.

Nissim-Sabat:1984:GMO

[NS84] Charles Nissim-Sabat. A Gedanken experiment to measure the one way velocity of light. British Journal for the Philosophy of Science, 35(1):62–64, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/1/62.full.pdf+html;
http://www.
jstor.org/stable/687560. See

[Øhr86, NS87].

Nissim-Sabat:1987:RGE

[NS87] Charles Nissim-Sabat. A retraction of ‘A Gedanken Experiment to Measure the One-Way Velocity of Light’. British Journal for the Philosophy of Science, 38(1):75a–75, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/1/75-a.full.pdf+html;
http://www.
jstor.org/stable/687080. See

[NS84, Øhr86].

Novick:2020:PIT

[NS20] Aaron Novick and Raphael Scholl. Presume it not: True causes in the search for the basis of heredity. British Journal for the Philosophy of Science, 71(1):59–86, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/1/59/4815698.

Nielsen:2021:PDP

[NS21] Michael Nielsen and Rush T. Stewart. Persistent disagreement and polarization in a Bayesian setting. British Journal for the Philosophy of Science, 72(1):51– 78, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Nguyen:2020:WSS

[NTW20] James Nguyen, Nicholas J. Teh, and Laura Wells. Why surplus structure is not superfluous. British Journal for the Philosophy of Science, 71(2):665–695, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/2/665/4955877.

Nugayev:1985:STU

[Nug85] Renat Nugayev. A study of theory unification. British Journal for the Philosophy of Science, 36(2):159–173, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
36/2/159.full.pdf+html;
http://www.jstor.org/
stable/687032.

Nunan:1993:HNA

[Nun93] Richard Nunan. Heuristic novelty and the asymmetry problem in Bayesian confirmation theory. British Journal for the Philosophy of Science, 44(1):17–36, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/17.full.pdf+
html;   http://www.jstor.
org/stable/687847.

NunodelaRosa:2022:CPE

[NV22] Laura Nuño de la Rosa and Cristina Villegas. Chances and propensities in evo–devo. British Journal for the Philosophy of Science, 73(2):509–533, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Nielsen:2022:SOI

[NW22] Michael Nielsen and Eric Wofsey.

Speed-optimal induction and dynamic coherence. British Journal for the Philosophy of Science, 73 (2):439–455, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Nyhof:1988:POK

[Nyh88]  John Nyhof.  Philosophical

objections to the kinetic theory. British Journal for the Philosophy of Science, 39(1):

81–109, March 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/1/81.full.pdf+html.

Nyrup:2021:WDA

[Nyr21] Rune Nyrup. ‘Of water drops and atomic nuclei: Analogies and pursuit worthiness in science’. British Journal for the Philosophy of Science, 71(3):881–903, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

OBrien:1969:R

[O’B69] D. O’Brien. Reviews. British Journal for the Philosophy of Science, 20(2):163–167, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/2/163.full.pdf+
html;   http://www.jstor.
org/stable/686179.

Okasha:2012:ERD

[OB12] Samir Okasha and K. G. Binmore, editors. Evolution and rationality: decisions, co-operation and strategic behaviour. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-00499-3 (hardcover). x + 281 pp. LCCN HM495 .E96

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1212/2012019850-b.html;
http://www.loc.gov/catdir/
enhancements/fy1212/

2012019850-d.html;
http://www.loc.gov/catdir/
enhancements/fy1212/
2012019850-t.html.

Oberdan:1998:VCA

[Obe98] Thomas Oberdan. The Vienna Circle’s ‘anti-foundationalism’. British Journal for the Philosophy of Science, 49(2):297–308, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/297.full.pdf+
html;   http://www.jstor.
org/stable/688116.

Oberdan:1999:DDP

[Obe99] T. Oberdan. Discussion. deconstructing protocols: reply to Uebel. British Journal for the Philosophy of Science, 50 (2):301–304, June 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/2/301.full.pdf+html;
http://www.jstor.org/
stable/40072226.

OConnor:1954:R

[O’C54]  D. J. O’Connor.  Reviews.

British Journal for the Philosophy of Science, V(18):166–167, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/166.full.pdf+
html;   http://www.jstor.
org/stable/685178.

OConnor:1956:Ra

[O’C56a] D. J. O’Connor. Reviews. British Journal for the Philosophy of Science, VI(24):347–349, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/24/347.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685421.

OConnor:1956:Rb

[O’C56b] D. J. O’Connor. Reviews. British Journal for the Philosophy of Science, VII(26):176–177, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/26/176.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685703.

OConnor:1956:Rc

[O’C56c] D. J. O’Connor. Reviews. British Journal for the Philosophy of Science, VII(27):270, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/27/270.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685890.

OConnor:1957:DP

[O’C57] D. J. O’Connor. Determinism and predictability. British Journal for the Philosophy of Science, VII(28):310–315, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/310.full.pdf+html;
http://www.jstor.org/
stable/685110.

OConnor:1963:R

[O’C63]  D. J. O’Connor.  Reviews.

British Journal for the Philosophy of Science, XIV(55):263– 264, November 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/55/263.full.pdf+html;
http://www.jstor.org/
stable/685245.

OConnor:1972:R

[O’C72] D. J. O’Connor. Reviews. British Journal for the Philosophy of Science, 23(1):79–80, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
23/1/79.full.pdf+html;
http://www.jstor.org/
stable/686224.

OConnor:1975:R

[O’C75] D. J. O’Connor. Reviews. British Journal for the Philosophy of Science, 26(4):353–355, December 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/4/353.full.pdf+
html;   http://www.jstor.
org/stable/686681.

OConnor:2017:EGT

[O’C17] Cailin O’Connor. Evolving to generalize: Trading precision for speed. British Journal for the Philosophy of Science, 68(2):68– ??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/2/389/3044191/
                                        Evolving-to-Generalize-
                                        Trading-Precision-for-
                                        Speed.

OConnor:2019:GK

[O’C19] Cailin O’Connor. Games and kinds. British Journal for the Philosophy of Science, 70(3): 719–745, September 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/3/
                                        719/4636565.

ODonnell:1990:EJM

[O’D90] Rod O’Donnell. The epistemology of J. M. Keynes. British Journal for the Philosophy of Science, 41(3):333–350, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/41/3/333.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687824.

Oddie:1981:VR

[Odd81] Graham Oddie. Verisimilitude reviewed. British Journal for the Philosophy of Science, 32(3):

237–265, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/237.full.pdf+html;
http://www.jstor.org/
stable/687589.

Oddie:1987:R

[Odd87] Graham Oddie. Reviews. British Journal for the Philosophy of Science, 38(2):272–276, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/2/272.full.pdf+
html;   http://www.jstor.
org/stable/687055.

Oddie:1990:VPR

[Odd90] Graham Oddie. Verisimilitude by power relations. British Journal for the Philosophy of Science, 41(1):129–135, March 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/1/129.full.pdf+html;
http://www.jstor.org/
stable/688007.

Oddie:1997:CCC

[Odd97] Graham Oddie. Conditionalization, cogency, and cognitive value. British Journal for the Philosophy of Science, 48(4): 533–541, December 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        48/4/533.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688098.

Oddie:2019:WAC

[Odd19] Graham Oddie. What accuracy could not be. British Journal for the Philosophy of Science, 70(2):551–580, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/2/
                                        551/4101970.

Oftedal:2007:BRN

[Oft07] Gry Oftedal. Book review: Neven Sesardic, Making Sense of Heritability. British Journal for the Philosophy of Science, 58(3):619–623, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/3/619.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115192.

Ogden:1951:R

[Ogd51] C. K. Ogden. Reviews. British Journal for the Philosophy of Science, II(6):164–165, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/6/164.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685509.

OHear:1983:R

[O’H83]  Anthony O’Hear.  Reviews.

British Journal for the Philosophy of Science, 34(3):303– 306, September 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/3/303.full.pdf+html;
http://www.jstor.org/
stable/687328.

OHear:1984:RG

[O’H84] Anthony O’Hear. Reply to Glassen. British Journal for the Philosophy of Science, 35(4): 377–380, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/377.full.pdf+html;
http://www.jstor.org/
stable/687341.

OHear:1985:R

[O’H85]  Anthony O’Hear.  Reviews.

British Journal for the Philosophy of Science, 36(3):354– 357, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/354.full.pdf+html;
http://www.jstor.org/
stable/687584.

OHear:1986:RU

[O’H86] Anthony O’Hear. Book review:

Watkins, John [1984]: Science and Scepticism. Hutchinson. Pp. xii + 387. £25.00. (ISBN 0-09-158010-2). British Journal for the Philosophy of Science, 37(3):363–367, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687295.

OHear:1993:SR

[O’H93] Anthony O’Hear. Science and religion. British Journal for the Philosophy of Science, 44(3): 505–516, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/505.full.pdf+html;
http://www.jstor.org/
stable/688019.

Ohrstrom:1986:NSO

[Øhr86] Peter Øhrstrøm. Nissim-Sabat on the one-way velocity of light. British Journal for the Philosophy of Science, 37(1): 120–122, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687003. See

[NS84, NS87].

Ord:2005:DMH

[OK05] Toby Ord and Tien D. Kieu. The diagonal method and hypercomputation. British Journal for the Philosophy of Science, 56 (1):147–156, March 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
56/1/147.full.pdf+html;
http://www.jstor.org/
stable/3541854.

Okasha:1997:LLE

[Oka97] Samir Okasha. Laudan and Leplin on empirical equivalence. British Journal for the Philosophy of Science, 48(2):251–256, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/251.full.pdf+
html;   http://www.jstor.
org/stable/687747.

Okasha:2000:BRB

[Oka00] S. Okasha. Book review: Genetics and Reductionism. S. Sarkar. British Journal for the Philosophy of Science, 51(1):181–185, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/1/181.full.pdf+
html;   http://www.jstor.
org/stable/3541757.

Okasha:2001:WWG

[Oka01] Samir Okasha. Why won’t the group selection controversy go away? British Journal for the Philosophy of Science, 52 (1):25–50, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/25.full.pdf+html;

http://www.jstor.org/
stable/3541941.

Okasha:2002:PTP

[Oka02]  Samir Okasha.  Philo-

sophical theories of probability. British Journal for the Philosophy of Science, 53(1):151–156, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/53/1/151.full.pdf+
html;   http://www.jstor.
org/stable/3541647.

Okasha:2004:MLS

[Oka04] Samir Okasha. Multi-level selection, covariance and contextual analysis. British Journal for the Philosophy of Science, 55(3):481–504, September 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/3/481.full.pdf+html;
http://www.jstor.org/
stable/3541671.

Okasha:2005:AGS

[Oka05] Samir Okasha. Altruism, group selection and correlated interaction. British Journal for the Philosophy of Science, 56(4): 703–725, December 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/703.full.pdf+html;
http://www.jstor.org/
stable/3541864.

Okasha:2008:FFT

[Oka08] Samir Okasha. Fisher’s fundamental theorem of natural selection — a philosophical analysis. British Journal for the Philosophy of Science, 59(3):319– 351, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/319.full.pdf+html;
http://www.jstor.org/
stable/40072289.

Okasha:2016:RBK

[Oka16] Samir Okasha. The relation between kin and multilevel selection: an approach using causal graphs. British Journal for the Philosophy of Science, 67 (2):435–470, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/435.

Osimani:2023:VEV

[OL23] Barbara Osimani and Jürgen Landes. Varieties of error and varieties of evidence in scientific inference. British Journal for the Philosophy of Science, 74(1): 117–170, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Olding:1978:DEL

[Old78] A. Olding. A defence of evolutionary laws. British Journal for the Philosophy of Science, 29(2):131–143, June 1978.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/2/131.full.pdf+html;
http://www.jstor.org/
stable/687123.

Olding:1982:HD

[Old82] A. Olding. The history of Darwinism. British Journal for the Philosophy of Science, 33(3): 315–321, September 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/3/315.full.pdf+html;
http://www.jstor.org/
stable/687230.

Oliver:2000:LMP

[Oli00] Alex Oliver. Logic, mathematics and philosophy. British Journal for the Philosophy of Science, 51(4):857–873, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/857.full.pdf+html;
http://www.jstor.org/
stable/3541730.

Olsson:2002:CTI

[Ols02a] Erik J. Olsson. Corroborating testimony and ignorance: a reply to Bovens, Fitelson, Hartmann and Snyder. British Journal for the Philosophy of Science, 53(4):565–572, December 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
53/4/565.full.pdf+html;
http://www.jstor.org/
stable/3541824.

Olsson:2002:CTP

[Ols02b] Erik J. Olsson. Corroborating testimony, probability and surprise. British Journal for the Philosophy of Science, 53 (2):273–288, June 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/2/273.full.pdf+html;
http://www.jstor.org/
stable/3541767.

Olsson:2003:FB

[Ols03] Erik J. Olsson. Foundations of Bayesianism. British Journal for the Philosophy of Science, 54(3):521–525, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/3/521.full.pdf+html;
http://www.jstor.org/
stable/3541800.

OMalley:2014:BRT

[O’M14] Maureen A. O’Malley. Book review: Thomas Pradeu, The Limits of the Self: Immunology and Biological Identity. British Journal for the Philosophy of Science, 65(1):179–183, January 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/1/179.full.pdf+html.

ONeill:1982:CT

[O’N82] L. J. O’Neill. Corroborating testimonies. British Journal for the Philosophy of Science, 33 (1):60–63, March 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/1/60.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687241.

ONeill:1989:TPI

[O’N89] John O’Neill. Two problems of induction. British Journal for the Philosophy of Science, 40 (1):121–125, March 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/1/121.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687468.

Opik:1954:AU

[Öpi54] Ernst J. Öpik. The age of the universe. British Journal for the Philosophy of Science, V(19):203–214, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/19/203.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685716.

Orde:1983:R

[Ord83] Simon Orde. Reviews. British Journal for the Philosophy of Science, 34(4):404–406, December 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/4/404.full.pdf+
html;   http://www.jstor.
org/stable/687535.

Orde:1987:R

[Ord87] Simon Orde. Reviews. British Journal for the Philosophy of Science, 38(1):133–136, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/1/133.full.pdf+
html;   http://www.jstor.
org/stable/687090.

Otsuka:2016:CFE

[Ots16] Jun Otsuka. Causal foundations of evolutionary genetics. unknown, 67(1):247–269, March 2016. CODEN ???? ISSN

????  URL http://bjps.
oxfordjournals.org/content/
67/1/247.

Otte:1994:SPB

[Ott94] Richard Otte. A solution to a problem for Bayesian confirmation theory. British Journal for the Philosophy of Science, 45(2):764–769, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/764.full.pdf+html;
http://www.jstor.org/
stable/687698.

Padel:1964:R

[Pad64] John Padel. Reviews. British Journal for the Philosophy of Science, XIV(56):356–359, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIV/56/356.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685643.

Page:2000:BRA

[Pag00] James Page. Book review: André Kukla — Studies in Scientific Realism. British Journal for the Philosophy of Science, 51(4): 957–961, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/4/957.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541744.

Pantin:1960:CCC

[Pan60] C. F. A. Pantin. Chairman’s comments on the contributions by Woodger and Hutten. British Journal for the Philosophy of Science, XI(42):109–112, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/42/109.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685583.

Paneth:1962:ESCa

[Pan62a] F. A. Paneth. The epistemological status of the chemical concept of element (I). British Journal for the Philosophy of Science, XIII(49):1–14, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/49/1.full.pdf+html;
http://www.jstor.org/
stable/685220.

Paneth:1962:ESCb

[Pan62b] F. A. Paneth. The epistemological status of the chemical concept of element (II). British Journal for the Philosophy of Science, XIII(50):144–160, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/144.full.
pdf+html;   http://www.
jstor.org/stable/685967.

Paneth:1963:FPW

[Pan63] Eva Paneth. F. A. Paneth’s works and translations. British Journal for the Philosophy of Science, XIII(52):317–318, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/317.full.
pdf+html;   http://www.
jstor.org/stable/685332.

Pap:1952:CAE

[Pap52] Arthur Pap. The concept of absolute emergence. British Journal for the Philosophy of Science, II(8):302–311, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/8/302.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686019.

Papineau:1976:ITE

[Pap76] David Papineau. Ideal types and empirical theories. British Journal for the Philosophy of Science, 27(2):137–146, June 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/2/137.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686163.

Papineau:1978:R

[Pap78]  David Papineau.  Reviews.

British Journal for the Philosophy of Science, 29(3):304– 310, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/3/304.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687026.

Papineau:1980:RU

[Pap80a] David Papineau. Review: [untitled]. British Journal for the Philosophy of Science, 31 (1):98–104, March 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687258.

Papineau:1980:R

[Pap80b] David Papineau. Reviews. British Journal for the Philosophy of Science, 31(1):98–103, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/1/98.full.pdf+html.

Papineau:1982:R

[Pap82]  David Papineau.  Reviews.

British Journal for the Philosophy of Science, 33(4):444– 448, December 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/4/444.full.pdf+html;
http://www.jstor.org/
stable/687177.

Papineau:1985:CA

[Pap85] David Papineau. Causal asymmetry. British Journal for the Philosophy of Science, 36(3): 273–289, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/273.full.pdf+html;
http://www.jstor.org/
stable/687571.

Papineau:1988:REP

[Pap88]  D. Papineau.  Response to

Ehring’s ‘Papineau on Causal Asymmetry’. British Journal for the Philosophy of Science, 39(4):521–525, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/4/521.full.pdf+html;
http://www.
jstor.org/stable/687436. See

[Ehr87].

Papineau:1991:RAC

[Pap91] D. Papineau. Review article: Correlations and causes. British Journal for the Philosophy of Science, 42(3):397–412, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/3/397.full.pdf+
html;   http://www.jstor.
org/stable/687916.

Papineau:1994:E

[Pap94a] David Papineau. Editorial. British Journal for the Philosophy of Science, 45(3):787– 788, September 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687793.

Papineau:1994:VR

[Pap94b] David Papineau. The virtues of randomization. British Journal for the Philosophy of Science, 45(2):437–450, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
45/2/437.full.pdf+html;
http://www.jstor.org/
stable/687674.

Papineau:1996:E

[Pap96a] David Papineau. Editorial.

British Journal for the Philosophy of Science, 47(3):345– 346, September 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/3/345.full.pdf+html;
http://www.jstor.org/
stable/687982.

Papineau:1996:MMN

[Pap96b] David Papineau. Many minds are no worse than one. British Journal for the Philosophy of Science, 47(2):233–241, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/233.full.pdf+
html;   http://www.jstor.
org/stable/687945.

Papineau:1998:E

[Pap98]  David Papineau.  Editorial.

British Journal for the Philosophy of Science, 49(4):531s– 531, December 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/4/531-s.full.pdf+html.

Paparazzo:2005:EPP

[Pap05] Ernesto Paparazzo. The Elder Pliny, Posidonius and surfaces. British Journal for the Philosophy of Science, 56(2):363–376, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/56/2/363.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541707.

Parsons:1971:CMC

[Par71] Kathryn Pyne Parsons. On criteria of meaning change. British Journal for the Philosophy of Science, 22(2):131–144, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/2/131.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686483.

Parsons:1982:SQM

[Par82]  Charles Parsons.  Substitu-

tional quantification and mathematics. British Journal for the Philosophy of Science, 33(4): 409–421, December 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/4/409.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687172.

Parascandola:1997:R

[Par97] Mark Parascandola. Review. British Journal for the Philosophy of Science, 48(4):621– 624, December 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/4/621.full.pdf+html;
http://www.jstor.org/
stable/688108.

Parker:2011:ITS

[Par11] Daniel Parker. Informationtheoretic statistical mechanics without Landauer’s Principle. British Journal for the Philosophy of Science, 62(4):831– 856, December 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/831.full.pdf+html.

Parrott:2016:BMD

[Par16] Matthew Parrott. Bayesian models, delusional beliefs, and epistemic possibilities. unknown, 67 (1):271–296, March 2016. CO-DEN ???? ISSN ???? URL

http://bjps.oxfordjournals.
org/content/67/1/271.

Parker:2017:CSM

[Par17] Wendy S. Parker. Computer simulation, measurement, and data assimilation. British Journal for the Philosophy of Science, 68(1): 68–??, March 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/1/273/2995387/

                                        Computer-Simulation-
                                        Measurement-and-Data.

Parrott:2021:DPE

[Par21] Matthew Parrott. Delusional predictions and explanations. British Journal for the Philosophy of Science, 72(1):325–353, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Paseau:2005:NMA

[Pas05] Alexander Paseau. Naturalism in mathematics and the authority of philosophy. British Journal for the Philosophy of Science, 56(2):377–396, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/2/377.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541708.

Paseau:2015:KMP

[Pas15] Alexander Paseau. Knowledge of mathematics without proof. British Journal for the Philosophy of Science, 66(4):775– 799, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Paseau:2016:BRJ

[Pas16] A. C. Paseau. Book review: John P. Burgess, Rigor and Structure. British Journal for the Philosophy of Science, 67 (4):1185–1187, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1185.

Patterson:1998:CCC

[Pat98] Sarah Patterson. Competence and the classical cascade: a reply to Franks. British Journal for the Philosophy of Science, 49(4):625–636, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/4/625.full.pdf+html;
http://www.jstor.org/
stable/688134.

Paul:1971:HUM

[Pau71]  Anthony M. Paul.  Hanson

on the unpicturability of microentities. British Journal for the Philosophy of Science, 22 (1):50–53, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/1/50.full.pdf+html;
http://www.jstor.org/
stable/686644.

Paul:2016:BRB

[Pau16]  Emily Paul.  Book review:

Bradford Skow, Objective Becoming. British Journal for the Philosophy of Science, 67 (4):1201–1205, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/4/1201.

Pooley:2002:RRC

[PB02] Oliver Pooley and Harvey R. Brown. Relationalism rehabilitated? I: Classical mechanics. British Journal for the Philosophy of Science, 53(2):183–204, June 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/53/2/183.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541763.

Powell:2012:ECR

[PC12] Russell Powell and Steve Clarke. Editor’s choice: Religion as an evolutionary byproduct: a critique of the standard model. British Journal for the Philosophy of Science, 63(3):457– 486, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/3/457.full.pdf+html.

Pearce:1982:SKI

[Pea82] David Pearce. Stegmüller on Kuhn restated. British Journal for the Philosophy of Science, 33(4):389–396, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/4/389.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687170.

Pearce:1983:TTC

[Pea83] David Pearce. Truthlikeness and translation: a comment on Oddie. British Journal for the Philosophy of Science, 34(4): 380–385, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/380.full.pdf+html;
http://www.jstor.org/
stable/687530.

Pearce:2012:ECC

[Pea12] Trevor Pearce. Editor’s choice:

Convergence and parallelism in evolution: a neo-Gouldian account. British Journal for the Philosophy of Science, 63 (2):429–448, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/429.full.pdf+html.

Pecnjak:1989:EMD

[Peć89] Davor Pećnjak. Epiphenomenalism and machines: a discussion of Van Rooijen’s critique of Popper. British Journal for the Philosophy of Science, 40(3):404–408, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/404.full.pdf+html;
http://www.jstor.org/
stable/687787.

Pence:2021:CVT

[Pen21] David Evan Pence. On the capacity for vision through sensory substitution. British Journal for the Philosophy of Science, 72(1): 79–103, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Perovic:2007:BRN

[Per07] Slobodan Perovic. Book review: Nicholas Maxwell, Is Science Neurotic?, London: Imperial College Press, 2004. Hardback price $48/£29 ISBN 1-86094-500-7. British Journal for the Philosophy of Science, 58 (2):361–363, June 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/2/361.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115233.

Peressini:2018:CPC

[Per18] Anthony F. Peressini. Causation, probability, and the continuity bind. British Journal for the Philosophy of Science, 69(3):881–909, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/3/881/2918938.

Peters:1956:FT

[Pet56] R. S. Peters. Freud’s theory. British Journal for the Philosophy of Science, VII(25):4–12, May 1956. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/25/4.full.pdf+
html;   http://www.jstor.
org/stable/685932.

Peters:1967:R

[Pet67] R. S. Peters. Reviews. British Journal for the Philosophy of Science, 17(4):331–334, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/331.full.pdf+
html;   http://www.jstor.
org/stable/686778.

Petkov:1989:SCE

[Pet89] Vesselin Petkov. Simultaneity, conventionality and existence. British Journal for the Philosophy of Science, 40(1):69–76, March 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/1/69.full.pdf+
html;   http://www.jstor.
org/stable/687464.

Pettit:1996:FEV

[Pet96] Philip Pettit. Functional explanation and virtual selection. British Journal for the Philosophy of Science, 47(2):291–302, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/291.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687949.

Pettigrew:2020:PPD

[Pet20] Richard Pettigrew. The principal principle does not imply the principle of indifference. British Journal for the Philosophy of Science, 71(2):605–619, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/2/605/4671056.

Peters:2022:IVC

[Pet22] Uwe Peters. Illegitimate values, confirmation bias, and Mandevillian cognition in science. British Journal for the Philosophy of Science, 72(4):1061–1081, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Peters:2023:SCP

[Pet23]  Uwe Peters.  Science com-

munication and the problematic impact of descriptive norms. British Journal for the Philosophy of Science, 74(3):713–738, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Peyerabend:1957:R

[Pey57] P. K. Peyerabend. Reviews. British Journal for the Philosophy of Science, VII(28):354– 357, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
VII/28/354.full.pdf+html;
http://www.jstor.org/
stable/685120.

Paternotte:2013:SNG

[PG13] Cédric Paternotte and Jonathan Grose. Social norms and game theory: Harmony or discord? British Journal for the Philosophy of Science, 64(3):551– 587, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/551.full.pdf+html.

Phillips:1979:R

[Phi79] D. C. Phillips. Reviews. British Journal for the Philosophy of Science, 30(1):85–86, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/85.full.pdf+html;
http://www.jstor.org/
stable/687418.

Phillips:1980:R

[Phi80] D. C. Phillips. Reviews. British Journal for the Philosophy of Science, 31(1):89–90, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/1/89.full.pdf+
html;   http://www.jstor.
org/stable/687256.

Phillips:2009:BRR

[Phi09] Ian B. Phillips. Book review: Robin Le Poidevin, The Images of Time: an Essay on Temporal Representation. British Journal for the Philosophy of Science, 60(2):439–446, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/2/439.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592008.

Pickering:1985:R

[Pic85] Andrew Pickering. Reviews. British Journal for the Philosophy of Science, 36(2):226–228, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/226.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687040.

Piccinini:2011:PCT

[Pic11] Gualtiero Piccinini. The physical Church–Turing thesis: Modest or bold? British Journal for the Philosophy of Science, 62(4):733–769, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/4/733.full.pdf+html.

Pikler:1954:UTF

[Pik54] Andrew G. Pikler. Utility theories in field physics and mathematical economics (I). British Journal for the Philosophy of Science, V(17):47–58, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/17/47.full.pdf+html;
http://www.jstor.org/
stable/685951.

Pikler:1955:UTF

[Pik55] Andrew G. Pikler. Utility theories in field physics and mathematical economics (II). British Journal for the Philosophy of Science, V(20):303–318, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/303.full.pdf+
html;   http://www.jstor.
org/stable/685732.

Pincock:2011:BRB

[Pin11a] Christopher Pincock. Book review: Bas C. Van Fraassen, Scientific Representation: Paradoxes of Perspective. British Journal for the Philosophy of Science, 62(3):677–682, September 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/677.full.pdf+html.

Pincock:2011:MSR

[Pin11b] Christopher Pincock. Mathematics and scientific representation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-975710-0 (hardcover). xiv + 330 pp. LCCN Q175.32.M38 P56 2011.

Pincock:2011:BER

[Pin11c] Christopher Pincock. On Batterman’s ‘On the Explanatory Role of Mathematics in Empirical Science’. British Journal for the Philosophy of Science, 62 (1):211–217, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/1/211.full.pdf+html.

Pincock:2015:AES

[Pin15] Christopher Pincock. Abstract explanations in science. British Journal for the Philosophy of Science, 66(4):857–882, December 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Pincock:2016:BRI

[Pin16] Christopher Pincock. Book review: Ian Hacking, Why Is There Philosophy of Mathematics At All? British Journal for the Philosophy of Science, 67(3):907–912, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/3/907.

Pincock:2022:CSM

[Pin22] Christopher Pincock. Concrete scale models, essential idealization, and causal explanation.

British Journal for the Philosophy of Science, 73(2):299–323, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Pirenne:1950:DBM

[Pir50] M. H. Pirenne. Descartes and the body-mind problem in physiology. British Journal for the Philosophy of Science, I(1):43– 59, May 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/1/43.full.pdf+
html;   http://www.jstor.
org/stable/685669.

Pirenne:1952:MLB

[Pir52a] M. H. Pirenne. Mind-like behaviour in artefacts and the concept of mind. British Journal for the Philosophy of Science, II (8):315–317, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/8/315.full.pdf+html;
http://www.jstor.org/
stable/686021.

Pirenne:1952:SBL

[Pir52b] M. H. Pirenne. The scientific basis of Leonardo Da Vinci’s theory of perspective. British Journal for the Philosophy of Science, III (10):169–185, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        III/10/169.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685555.

Pirie:1952:CCP

[Pir52c] N. W. Pirie. Concepts out of context: the pied pipers of science. British Journal for the Philosophy of Science, II(8): 269–280, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/8/269.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686016.

Pirenne:1953:PMP

[Pir53] M. H. Pirenne. Physiological mechanisms in the perception of distance by sight and Berkeley’s theory of vision. British Journal for the Philosophy of Science, IV(13):13–21, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/13/13.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686041.

Pirenne:1954:MBP

[Pir54] M. H. Pirenne. The mind-brain problem. British Journal for the Philosophy of Science, V(18): 153–159, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/18/153.full.pdf+html;

http://www.jstor.org/
stable/685175.

Pirie:1956:R

[Pir56] N. W. Pirie. Reviews. British Journal for the Philosophy of Science, VI(24):341–343, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/341.full.pdf+
html;   http://www.jstor.
org/stable/685417.

Pitowsky:1994:GBC

[Pit94] Itamar Pitowsky. George Boole’s ‘Conditions of Possible Experience’ and the quantum puzzle. British Journal for the Philosophy of Science, 45(1):95–125, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/95.full.pdf+
html;   http://www.jstor.
org/stable/687963.

Pitts:2008:WBB

[Pit08] J. Brian Pitts. Why the Big Bang singularity does not help the Kalām cosmological argument for theism. British Journal for the Philosophy of Science, 59(4):675–708, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/675.full.pdf+html;
http://www.jstor.org/
stable/40072307.

Pitts:2011:PUA

[Pit11] J. Brian Pitts. Permanent underdetermination from approximate empirical equivalence in field theory: Massless and massive scalar gravity, neutrino, electromagnetic, Yang–Mills and gravitational theories. British Journal for the Philosophy of Science, 62(2):259–299, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/2/259.full.pdf+html.

Pearson:1970:SHSa

[PK70a] E. S. (Egon Sharpe) Pearson and Maurice G. (Maurice George) Kendall. Studies in the history of statistics and probability. Griffin, London, UK, 1970. ISBN 0-85264-193-1. 10 + 481 pp. LCCN QA276.15 .P43.

Pearson:1970:SHSb

[PK70b] E. S. (Egon Sharpe) Pearson and Maurice G. (Maurice George) Kendall. Studies in the history of statistics and probability. Hafner Pub. Co., Darien, CT, USA, 1970. ISBN 0-85264-193-1. 481 pp. LCCN QA276.15 .P43 1970b.

Plaut:1960:CCF

[Pla60] H. C. Plaut. Condition, cause, free will, and the direction of time. British Journal for the Philosophy of Science, XI(43): 212–221, November 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/43/212.full.pdf+html;
http://www.jstor.org/
stable/685310.

Plaut:1962:ESR

[Pla62] H. C. Plaut. Empiricism, solipsism, and realism. British Journal for the Philosophy of Science, XIII(51):216–228, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/216.full.
pdf+html;   http://www.
jstor.org/stable/685839.

Plaut:1963:STD

[Pla63] H. C. Plaut. Schlick’s treatment of determinism. British Journal for the Philosophy of Science, XIII(52):315–316, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/315.full.
pdf+html;   http://www.
jstor.org/stable/685330.

Platts:1983:EK

[Pla83] Mark Platts. Explanatory kinds.

British Journal for the Philosophy of Science, 34(2):133–148, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/133.full.pdf+
html;   http://www.jstor.
org/stable/687446.

Placek:2000:SOB

[Pla00] T. Placek. Stochastic outcomes in branching space–time: analysis of Bell’s Theorem. British Journal for the Philosophy of Science, 51(3):445–475, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/3/445.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541773.

Planer:2016:GRR

[Pla16] Ronald J. Planer. Are genetic representations read in development? British Journal for the Philosophy of Science, 67(4):997–1023, December 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/4/997.

Pradeu:2024:PSC

[PLKG24] Thomas Pradeu, Maël Lemoine, Mahdi Khelfaoui, and Yves Gingras. Philosophy in science: Can philosophers of science permeate through science and produce scientific knowledge? British Journal for the Philosophy of Science, 75(2):375–416, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Plutynski:2005:EUE

[Plu05] Anya Plutynski. Explanatory unification and the early synthesis. British Journal for the Philosophy of Science, 56(3):595– 609, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/595.full.pdf+html;
http://www.jstor.org/
stable/3541721.

Paton:1994:ESM

[PNSBC94] R. C. Paton, H. S. Nwana, M. J. R. Shave, and T. J. M. Bench-Capon. An examination of some metaphorical contexts for biologically motivated computing. British Journal for the Philosophy of Science, 45 (2):505–525, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/505.full.pdf+html;
http://www.jstor.org/
stable/687679.

Podlaha:1975:SNA

[Pod75] M. F. Podlaha. Some new aspects of relativity: Remarks on Keswani’s paper. British Journal for the Philosophy of Science, 26(2):133–137, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/2/133.full.pdf+html;
http://www.
jstor.org/stable/686512. See

[Kes66a, Kes66b].

Podlaha:1976:SNA

[Pod76] M. F. Podlaha. Some new aspects of Relativity: Remarks on Zahar’s paper. British Journal for the Philosophy of Science, 27(3):261–267, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/3/261.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686124.

Podlaha:1979:LSS

[Pod79] M. F. Podlaha. Light signal synchronisation and clock transport synchronisation in the theory of Relativity. British Journal for the Philosophy of Science, 30(4):376–380, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/4/376.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687546.

Pogany:1985:MIR

[Pog85] George A. Pogany. The methodology of industrial research. British Journal for the Philosophy of Science, 36(3):309– 313, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        36/3/309.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687575.

Polanyi:1952:HC

[Pol52a] Michael Polanyi. The hypothesis of cybernetics. British Journal for the Philosophy of Science, II (8):312–315, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/8/312.full.pdf+html;
http://www.jstor.org/
stable/686020.

Polanyi:1952:SB

[Pol52b] Michael Polanyi. The stability of beliefs. British Journal for the Philosophy of Science, III(11):217–232, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/217.full.pdf+html;
http://www.jstor.org/
stable/685265.

Polanyi:1957:PS

[Pol57] Michael Polanyi. Problem solving. British Journal for the Philosophy of Science, VIII(30): 89–103, August 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/30/89.full.pdf+html;
http://www.jstor.org/
stable/685789.

Pollock:1958:WMS

[Pol58] Seton Pollock. What makes a subject scientific? British Journal for the Philosophy of Science, IX(34):130–132, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        IX/34/130.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685149.

Polanyi:1967:SR

[Pol67] Michael Polanyi. Science and reality. British Journal for the Philosophy of Science, 18(3): 177–196, November 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/3/177.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686589.

Poltawski:1969:R

[Pol69] Andrzej Poltawski. Reviews. British Journal for the Philosophy of Science, 20(1):84–85, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/1/84.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686722.

Polanyi:1971:R

[Pol71]  Michael Polanyi.  Reviews.

British Journal for the Philosophy of Science, 22(3):307–308, August 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/3/307.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686751.

Polger:2014:BRD

[Pol14] Thomas W. Polger. Book review: David J. Chalmers, Constructing the World. British Journal for the Philosophy of Science, 65(2):419–423, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/419.

Poole:1955:CCV

[Poo55] H. W. Poole. A close-clipped view of the universe, with a note on its ‘origin’ and ‘age ’, and our position in it. British Journal for the Philosophy of Science, VI(21):43–50, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/21/43.full.pdf+html;
http://www.jstor.org/
stable/685573.

Popper:1950:IQPa

[Pop50a] Karl R. Popper. Indeterminism in quantum physics and in classical physics. British Journal for the Philosophy of Science, I(2):117–133, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/2/117.full.pdf+html.

Popper:1950:IQPb

[Pop50b] Karl R. Popper. Indeterminism in quantum physics and in classical physics: Part II. British Journal for the Philosophy of Science, I(3):173–195, November 1950. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/I/3/173.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685366.

Popper:1952:NPP

[Pop52] K. R. Popper. The nature of philosophical problems and their roots in science. British Journal for the Philosophy of Science, III (10):124–156, August 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/10/124.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685553.

Popper:1953:NBP

[Pop53] K. R. Popper. A note on Berkeley as precursor of Mach. British Journal for the Philosophy of Science, IV(13):26–36, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IV/13/26.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686043.

Popper:1954:DC

[Pop54] Karl R. Popper. Degree of confirmation. British Journal for the Philosophy of Science, V(18): 143–149, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/

V/18/143.full.pdf+html;
http://www.jstor.org/
stable/685171.

Popper:1955:CDC

[Pop55a] Karl R. Popper. ‘Content’ and ‘degree of confirmation’: a reply to Dr Bar-Hillel. British Journal for the Philosophy of Science, VI (22):157–163, August 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/22/157.full.pdf+html;
http://www.jstor.org/
stable/685524.

Popper:1955:TAA

[Pop55b] Karl R. Popper. Two autonomous axiom systems for the calculus of probabilities. British Journal for the Philosophy of Science, VI(21):51–57, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/21/51.full.pdf+
html;   http://www.jstor.
org/stable/685574.

Popper:1956:ACA

[Pop56a] K. R. Popper. Adequacy and consistency: Adequacy and consistency: a second reply to Dr Bar-Hillel. British Journal for the Philosophy of Science, VII (27):249–256, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/249.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685883.

Popper:1956:CDC

[Pop56b] K. R. Popper. Content and degree of confirmation: Reply to Professor Carnap. British Journal for the Philosophy of Science, VII(27):244–245, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/27/244.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685881.

Popper:1957:IES

[Pop57a] Karl R. Popper. Irreversibility; or, entropy since 1905. British Journal for the Philosophy of Science, VIII(30):151–155, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/30/151.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685793.

Popper:1957:SND

[Pop57b] Karl R. Popper. A second note on degree of confirmation. British Journal for the Philosophy of Science, VII(28): 350–353, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/28/350.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685119.

Popper:1958:TND

[Pop58a] K. R. Popper. A third note on degree of corroboration or confirmation. British Journal for the Philosophy of Science, VIII(32): 294–302, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/294.full.pdf+html;
http://www.jstor.org/
stable/685855.

Popper:1958:MRH

[Pop58b] Karl R. Popper. On Mr Roy Harrod’s new argument for induction. British Journal for the Philosophy of Science, IX (35):221–224, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/221.full.pdf+html;
http://www.jstor.org/
stable/685651.

Popper:1959:DTA

[Pop59a] K. R.

Popper. Discussions: Testability and ‘ ad-hocness ’ of the contraction hypothesis. British Journal for the Philosophy of Science, X(37):50a–50, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/37/50-a.full.pdf+html.

Popper:1959:TAH

[Pop59b] K. R. Popper. Testability and ’ad-hocness’ of the contraction hypothesis. British Journal for the Philosophy of Science, X(37):50, May 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685776.

Popper:1959:PIP

[Pop59c] Karl R. Popper. The propensity interpretation of probability. British Journal for the Philosophy of Science, X(37):25–42, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/37/25.full.pdf+
html;   http://www.jstor.
org/stable/685773.

Popper:1960:PIC

[Pop60] K. R. Popper. Probabilistic independence and Coksobosation by empirical tests. British Journal for the Philosophy of Science, X(40):315–318, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/315.full.pdf+html;
http://www.jstor.org/
stable/685919.

Popper:1962:CNP

[Pop62] K. R. Popper. A comment on the new prediction paradox. British Journal for the Philosophy of Science, XIII(49):51b–51, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/51-b.full.
pdf+html;   http://www.
jstor.org/stable/685224.

Popper:1966:CMN

[Pop66a] Karl R. Popper. A comment on Miller’s new paradox of information. British Journal for the Philosophy of Science, 17 (1):61–69, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686405.

Popper:1966:NDB

[Pop66b] Karl R. Popper. A note on the difference between the Lorentz– Fitzgerald contraction and the Einstein contraction. British Journal for the Philosophy of Science, XVI(64):332–333, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/64/332.full.
pdf+html;   http://www.
jstor.org/stable/686834.

Popper:1966:PZI

[Pop66c] Karl R. Popper. A paradox of zero information. British Journal for the Philosophy of Science, 17 (2):141–143, August 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/2/141.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/686345.

Popper:1968:RDN

[Pop68] Karl R. Popper. A revised definition of natural necessity. British Journal for the Philosophy of Science, 18(4):316–321, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/4/316.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686360.

Popper:1976:NV

[Pop76] Karl Popper. A note on verismilitude. British Journal for the Philosophy of Science, 27 (2):147–159, June 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/2/147.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686164.

Popper:1978:PIP

[Pop78] Karl Popper. On the possibility of an infinite past: a reply to Whitrow. British Journal for the Philosophy of Science, 29(1):47–48, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/1/47.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686391.

Popkin:1979:R

[Pop79] Richard H. Popkin. Reviews.

British Journal for the Philosophy of Science, 30(2):205–209, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/205.full.pdf+
html;   http://www.jstor.
org/stable/686958.

Popper:1981:JHW

[Pop81] Karl Popper. Joseph Henry Woodger. British Journal for the Philosophy of Science, 32(3): 328–330, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/328.full.pdf+html;
http://www.jstor.org/
stable/687601.

Popkin:1983:R

[Pop83]  Richard Popkin.  Reviews.

British Journal for the Philosophy of Science, 34(2):182–185, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/182.full.pdf+
html;   http://www.jstor.
org/stable/687450.

Post:1960:RMB

[Pos60a]  H. R. Post.  Retrospective

miracles or betting after the race. British Journal for the Philosophy of Science, X(40):

338–341, February 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/338.full.pdf+html;
http://www.jstor.org/
stable/685922.

Post:1960:SST

[Pos60b] H. R. Post. Simplicity in scientific theories. British Journal for the Philosophy of Science, XI(41):32–41, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/41/32.full.pdf+html;
http://www.jstor.org/
stable/685817.

Post:1962:CPT

[Pos62a] H. R. Post. A criticism of Popper’s theory of simplicity. British Journal for the Philosophy of Science, XII(48):328–331, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/48/328.full.
pdf+html;   http://www.
jstor.org/stable/685400.

Post:1962:R

[Pos62b] H. R. Post. Reviews. British Journal for the Philosophy of Science, XIII(49):80– 82, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/XIII/49/80.full.
pdf+html;   http://www.
jstor.org/stable/685233.

Post:1971:R

[Pos71] H. R. Post. Reviews. British Journal for the Philosophy of Science, 22(2):208, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/2/208.full.pdf+html;
http://www.jstor.org/
stable/686495.

Post:1973:TQ

[Pos73] Heinz Post. The trouble with quanta. British Journal for the Philosophy of Science, 24(3): 277–282, September 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/3/277.full.pdf+html;
http://www.jstor.org/
stable/686233.

Post:1974:PPL

[Pos74] John F. Post. Propositions, possible languages and the Liar’s Revenge. British Journal for the Philosophy of Science, 25(3): 223–234, September 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/3/223.full.pdf+html;
http://www.jstor.org/
stable/686905.

Post:1975:PA

[Pos75a] Heinz Post. The problem of atomism. British Journal for the Philosophy of Science, 26 (1):19–26, March 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/1/19.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686329.

Post:1975:R

[Pos75b] John F. Post. Reviews. British Journal for the Philosophy of Science, 26(1):73–81, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/1/73.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686336.

Post:1978:R

[Pos78] John F. Post. Reviews. British Journal for the Philosophy of Science, 29(2):189–195, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/29/2/189.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687128.

Potter:1951:NPF

[Pot51] Owen Potter. Note on Philipp Frank’s interpretation of science. British Journal for the Philosophy of Science, II(5):58c–60, May 1951. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/58-c.full.pdf+
html;   http://www.jstor.
org/stable/685610.

Potter:1953:FCF

[Pot53] Owen Potter. On the function of chemical formulae: a Nineteenth-Century discussion. British Journal for the Philosophy of Science, III(12):359b– 363, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/359-b.full.pdf+html;
 http://www.jstor.org/
stable/685456.

Pottage:1969:R

[Pot69] John Pottage. Reviews. British Journal for the Philosophy of Science, 20(4):374–375, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/374.full.pdf+
html;   http://www.jstor.
org/stable/686271.

Potter:1996:TI

[Pot96] Michael Potter. Taming the infinite. British Journal for the Philosophy of Science, 47(4): 609–619, December 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        47/4/609.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687927.

Potter:2000:RNK

[Pot00] Michael D. Potter. Reason’s nearest kin: philosophies of arithmetic from Kant to Carnap. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-825041-X. x + 305 pp. LCCN QA8.4 .P68 2000; QA248 .P67 2000; QA248 .P68 2000.

Potter:2002:RNK

[Pot02] Michael Potter. Reason’s Nearest Kin. Philosophies of Arithmetic from Kant to Carnap. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-925261-0. 315 pp. LCCN QA248. URL http:/

                                        /www.oxfordscholarship.com/
                                        oso/public/content/
                                        philosophy/9780199252619/
                                        toc.

Potochnik:2010:EIE

[Pot10] Angela Potochnik. Explanatory independence and epistemic interdependence: a case study of the optimality approach. British Journal for the Philosophy of Science, 61 (1):213–233, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/1/213.full.pdf+html.

Povich:2018:MMG

[Pov18] Mark Povich. Minimal models and the generalized ontic conception of scientific explanation. British Journal for the Philosophy of Science, 69(1):117– 137, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/117/3044200.

Powell:1992:R

[Pow92] Andrew Powell. Review. British Journal for the Philosophy of Science, 43(2):245–262, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/245.full.pdf+
html;   http://www.jstor.
org/stable/687876.

Powell:2012:FHE

[Pow12] Russell Powell. The future of human evolution. British Journal for the Philosophy of Science, 63 (1):145–175, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/1/145.full.pdf+html.

Pearson:1978:HSC

[PP78] Karl Pearson and E. S.

(Egon Sharpe) Pearson. The history of statistics in the 17th and 18th centuries against the changing background of intellectual, scientific, and religious thought: lectures by Karl Pearson given at University College, London, during the academic sessions, 1921– 1933. C. Griffin, London, UK, 1978. ISBN 0-85264-250-4. xix + 744 pp. LCCN QA276.15 .P44 1978. US$7.50.

Perez-Ramos:1991:FBD

[PR91] Antonio Pérez-Ramos. Francis Bacon and the disputations of the learned. British Journal for the Philosophy of Science, 42(4):577–588, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/4/577.full.pdf+html;
http://www.jstor.org/
stable/687711.

Pietroski:1995:WOT

[PR95] Paul Pietroski and Georges Rey. When other things aren’t equal: Saving ceteris paribus laws from vacuity. British Journal for the Philosophy of Science, 46 (1):81–110, March 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/1/81.full.pdf+html;
http://www.jstor.org/
stable/687634. See comments

[Sch01b].

Pence:2013:ECN

[PR13] Charles H. Pence and Grant Ramsey. Editor’s choice: a new foundation for the propensity interpretation of fitness.

British Journal for the Philosophy of Science, 64(4):851– 881, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/851.full.pdf+html.

Pratt:1972:BC

[Pra72] Vernon Pratt. Biological classification. British Journal for the Philosophy of Science, 23(4): 305–327, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/305.full.pdf+html;
http://www.jstor.org/
stable/686731.

Pratt:1979:R

[Pra79] Vernon Pratt. Reviews. British Journal for the Philosophy of Science, 30(3):309–312, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/309.full.pdf+
html;   http://www.jstor.
org/stable/687149.

Pratt:1984:R

[Pra84] Vernon Pratt. Reviews. British Journal for the Philosophy of Science, 35(1):81–83, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/81.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687562.

Pradeu:2012:LSI

[Pra12] Thomas Pradeu. The limits of the self: immunology and biological identity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-977528-1. ix + 302 pp. LCCN QR181 .P64 2012.

Prasetya:2024:WMS

[Pra24] Yunus Prasetya. Which models of scientific explanation are (in)compatible with inference to the best explanation? British Journal for the Philosophy of Science, 75(1):209–232, March 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Presley:1961:AAM

[Pre61] C. F. Presley. Arguments about meaninglessness. British Journal for the Philosophy of Science, XII(47):225–234, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/47/225.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685209.

Preston:1994:R

[Pre94] John Preston. Review. British Journal for the Philosophy of Science, 45(4):1063–1065, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/4/1063.full.pdf+
html;   http://www.jstor.
org/stable/687623.

Preston:1997:R

[Pre97] John Preston. Review. British Journal for the Philosophy of Science, 48(4):610–612, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/610.full.pdf+
html;   http://www.jstor.
org/stable/688105.

Preston:2001:BRL

[Pre01] John Preston. Book review:

Luciano Floridi, Philosophy and Computing: an Introduction. British Journal for the Philosophy of Science, 52(1):197–200, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/1/197.full.pdf+
html;   http://www.jstor.
org/stable/3541957.

Preston:2003:MBR

[Pre03] Beth Preston. Of marigold beer: a reply to Vermaas and Houkes. British Journal for the Philosophy of Science, 54(4): 601–612, December 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/4/601.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/3541681.

Pringle:1953:R

[Pri53] J. W. S. Pringle. Reviews. British Journal for the Philosophy of Science, III(12):388– 389, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/388.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685466.

Prior:1955:EO

[Pri55] A. N. Prior. English and ontology. British Journal for the Philosophy of Science, VI(21):64–65, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/21/64.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685577.

Price:1959:R

[Pri59] H. H. Price. Reviews. British Journal for the Philosophy of Science, X(38):147–153, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/38/147.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686005.

Price:1960:RSR

[Pri60] H. H. Price. Reviews: Sir Russell Brain on modes of apprehension. British Journal for the Philosophy of Science, XI(41):71–76, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/71.full.pdf+
html;   http://www.jstor.
org/stable/685824.

Primack:1968:R

[Pri68] M. Primack. Reviews. British Journal for the Philosophy of Science, 19(3):269–270, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/269.full.pdf+
html;   http://www.jstor.
org/stable/686253.

Prior:1969:R

[Pri69] A. N. Prior. Reviews. British Journal for the Philosophy of Science, 20(4):372–374, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/372.full.pdf+
html;   http://www.jstor.
org/stable/686270.

Price:1976:PLB

[Pri76] John F. Price. On the probability of laws being zero. British Journal for the Philosophy of Science, 27(4):392–395, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/27/4/392.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686865.

Price:1984:MCS

[Pri84] Huw Price. Mellor, chance and the single case. British Journal for the Philosophy of Science, 35(1):11–23, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        35/1/11.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687554.

Priest:1989:PQS

[Pri89] Graham Priest. Primary qualities are secondary qualities too. British Journal for the Philosophy of Science, 40(1):29–37, March 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/40/1/29.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687462.

Price:1991:APC

[Pri91a] Huw Price. Agency and probabilistic causality. British Journal for the Philosophy of Science, 42(2):157–176, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/2/157.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688050.

Price:1991:PP

[Pri91b] Huw Price. The philosophy of physics. British Journal for the Philosophy of Science, 42 (1):111–144, March 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/1/111.full.pdf+html;
http://www.jstor.org/
stable/687611.

Price:1994:RWF

[Pri94] Huw Price. Reinterpreting the Wheeler–Feynman absorber theory: Reply to Leeds. British Journal for the Philosophy of Science, 45(4):1023–1028, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/4/1023.full.pdf+
html;   http://www.jstor.
org/stable/687619.

Price:1998:R

[Pri98] Huw Price. Reviews. British Journal for the Philosophy of Science, 49(3):522–525, September 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/522.full.pdf+
html;   http://www.jstor.
org/stable/688090.

Price:1999:BRB

[Pri99] H. Price. Book review: The Fabric of Reality. D. Deutsch.

British Journal for the Philosophy of Science, 50(2):309–312, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/2/309.full.pdf+
html;   http://www.jstor.
org/stable/40072228.

Priest:2001:BRS

[Pri01]  Graham Priest.  Book re-

view: Scott Soames, Understanding Truth. British Journal for the Philosophy of Science, 52 (1):211–215, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/211.full.pdf+html;
http://www.jstor.org/
stable/3541960.

Price:2002:BTB

[Pri02] Huw Price. Boltzmann’s time bomb. British Journal for the Philosophy of Science, 53 (1):83–119, March 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/1/83.full.pdf+html.

Prinz:2004:GRP

[Pri04]  Jesse J. Prinz.  Gut reac-

tions: a perceptual theory of emotion. Philosophy of mind series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-515145-3 (hardcover), 0-19-530936-7 (paperback). ix + 271 pp. LCCN BF531 .P75 2004.

Priest:2009:VFH

[Pri09] Graham Priest. Vincent F.

Hendrick’s mainstream and formal epistemology. British Journal for the Philosophy of Science, 60(2):433–437, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/2/433.full.pdf+html;
http://www.jstor.org/
stable/25592007.

Proctor:1959:SLS

[Pro59] George L. Proctor. Scientific laws and scientific objects in the Tractatus. British Journal for the Philosophy of Science, X(39):177–193, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/177.full.pdf+html;
http://www.jstor.org/
stable/686052.

Prokhovnik:1963:CA

[Pro63] S. J. Prokhovnik. The case for an aether. British Journal for the Philosophy of Science, XIV(55):195–207, November 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/55/195.full.
pdf+html;   http://www.
jstor.org/stable/685240.

Prokhovnik:1968:NRP

[Pro68]  S. J. Prokhovnik.  A note

on relativistic phenomena in an ether theory. British Journal for the Philosophy of Science, 18(4):322–323, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/4/322.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686361.

Prokhovnik:1974:DEP

[Pro74] S. J. Prokhovnik. Did Einstein’s programme supersede Lorentz’s? British Journal for the Philosophy of Science, 25(4):336– 340, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/4/336.full.pdf+html;
                                        http://links.jstor.org/
                                        sici?sici=0007-0882%
                                        28197412%2925%3A4%3C336%
                                        3ADEPSL%3E2.0.CO%3B2-U;
                                        http://www.jstor.org/
                                        stable/686462.

Prokhovnik:1977:MR

[Pro77a] S. J. Prokhovnik. Mathematics and reality. British Journal for the Philosophy of Science, 28(2):189–194, June 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        28/2/189.full.pdf+html;

http://www.jstor.org/
stable/686635.

Prokhovnik:1977:R

[Pro77b] S. J. Prokhovnik. Reviews.

British Journal for the Philosophy of Science, 28(4):389– 390, December 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/4/389.full.pdf+html;
http://www.jstor.org/
stable/686882.

Prokhovnik:1986:RU

[Pro86] Simon Prokhovnik. Book review:

Van der Merwe, Alwyn (ed.) [1983]: Old and New Questions in Physics, Cosmology, Philosophy and Theoretical Biology: Essays in Honor of Wolfgang Yourgrau. Plenum Press. $95. Pp. xv + 920. British Journal for the Philosophy of Science, 37(2):257–261, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686992.

Pryce:1960:R

[Pry60] M. H. L. Pryce. Reviews.

British Journal for the Philosophy of Science, XI(42):157–159, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/157.full.pdf+
html;   http://www.jstor.
org/stable/685593.

Pryor:2001:HRE

[Pry01] James Pryor. Highlights of recent epistemology. British Journal for the Philosophy of Science, 52(1):95–124, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/1/95.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541945.

Przelecki:1962:R

[Prz62] M. Przelecki. Reviews. British Journal for the Philosophy of Science, XIII(49):87, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/49/87.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685236.

Petroni:1986:DPD

[PS86] Angelo M. Petroni and Lucio Scolamiero. On the determination of planetary distances in the Copernican system. British Journal for the Philosophy of Science, 37(3):335– 340, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687290.

Purpura:2000:BRB

[PS00] G. Purpura and R. Samuels. Book review: Species of Mind: the Philosophy and Biology of Cognitive Ethology. C. Allen, M.

Bekoff. British Journal for the Philosophy of Science, 51 (2):375–380, June 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/2/375.full.pdf+html;
http://www.jstor.org/
stable/3541816.

Psillos:1994:R

[Psi94] Stathis Psillos. Review. British Journal for the Philosophy of Science, 45(3):923–926, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/923.full.pdf+
html;   http://www.jstor.
org/stable/687803.

Psillos:1996:R

[Psi96] Stathis Psillos. Review. British Journal for the Philosophy of Science, 47(4):663–668, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/663.full.pdf+
html;   http://www.jstor.
org/stable/687937.

Psillos:2000:PSS

[Psi00] Stathis Psillos. The present state of the scientific realism debate. British Journal for the Philosophy of Science, 51(4): 705–728, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/3541614.

Psillos:2004:TRT

[Psi04] Stathis Psillos. Tracking the real: Through thick and thin. British Journal for the Philosophy of Science, 55(3):393–409, September 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/55/3/393.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541668.

Peters:1957:HHM

[PT57] R. S. Peters and H. Tajfel. Hobbes and Hull — metaphysicians of behaviour. British Journal for the Philosophy of Science, VIII(29):30–44, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/29/30.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685379.

Puce:1956:R

[Puc56] H. H. Puce. Reviews. British Journal for the Philosophy of Science, VII(27):257–260, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/27/257.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685884.

Puccetti:1962:MSC

[Puc62] Roland Puccetti. Are metaphysical statements confirmable? British Journal for the Philosophy of Science, XIII(49):52–53, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/52.full.
pdf+html;   http://www.
jstor.org/stable/685225.

Puccetti:1967:TMF

[Puc67] Roland Puccetti. On thinking machines and feeling machines. British Journal for the Philosophy of Science, 18(1):39– 51, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/39.full.pdf+
html;   http://www.jstor.
org/stable/686888.

Puccetti:1969:SP

[Puc69] Roland Puccetti. The sensation of pleasure. British Journal for the Philosophy of Science, 20(3): 239–245, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/3/239.full.pdf+html;
http://www.jstor.org/
stable/686417.

Puccetti:1973:BBP

[Puc73] Roland Puccetti. Brain bisection and personal identity. British Journal for the Philosophy of Science, 24(4):339–355, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/4/339.full.pdf+
html;   http://www.jstor.
org/stable/686615.

Puccetti:1974:PRC

[Puc74] Roland Puccetti. Pattern recognition in computers and the human brain: With special application to chess playing machines. British Journal for the Philosophy of Science, 25(2):137–154, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/2/137.full.pdf+
html;   http://www.jstor.
org/stable/686818.

Puccetti:1976:MSR

[Puc76] Roland Puccetti. The mute self: a reaction to Dewitt’s alternative account of the split-brain data. British Journal for the Philosophy of Science, 27(1):65–73, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/1/65.full.pdf+
html;   http://www.jstor.
org/stable/686380.

Puccetti:1977:RMV

[Puc77] Roland Puccetti. Robinson on my views: a correction. British Journal for the Philosophy of Science, 28(2):171, June 1977.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/2/171.full.pdf+html;
http://www.jstor.org/
stable/686631.

Puccetti:1978:UWK

[Puc78] Roland Puccetti. Unravelling the world knot: Scientists and philosophers on the mind–brain controversy. British Journal for the Philosophy of Science, 29 (1):61–68, March 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/1/61.full.pdf+html;
http://www.jstor.org/
stable/686393.

Puccetti:1981:CMB

[Puc81] Roland Puccetti. Current mind– brain confusions. British Journal for the Philosophy of Science, 32(3):282–286, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/282.full.pdf+html;
http://www.jstor.org/
stable/687592.

Puccetti:1989:TBT

[Puc89] Roland Puccetti. Two brains, two minds? Wigan’s theory of mental duality. British Journal for the Philosophy of Science, 40(2):137–144, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/2/137.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687504.

Puccetti:1993:MDB

[Puc93] Roland Puccetti. Mind with a double brain. British Journal for the Philosophy of Science, 44(4):675–691, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        44/4/675.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/688038.

Pumphrey:1953:ET

[Pum53] R. J. Pumphrey. The evolution of thinking. British Journal for the Philosophy of Science, IV(16):315–327, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/16/315.full.pdf+html.

Pumphrey:1954:ET

[Pum54] R. J. Pumphrey. The evolution of thinking. British Journal for the Philosophy of Science, IV(16):315–327, February 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686080.

Purton:1971:R

[Pur71] A. C. Purton. Reviews. British Journal for the Philosophy of Science, 22(1):81–82, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/81.full.pdf+html;
http://www.jstor.org/
stable/686661.

Purves:2023:CCL

[Pur23] Gordon Michael Purves. Conventionality and causality in Lewis-type evolutionary prediction games. British Journal for the Philosophy of Science, 74(1): 199–219, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Putnam:2001:BRB

[Put01] Ruth Anna Putnam. Book review: Truth, Rationality and Pragmatism: Themes from Peirce, Christopher Hookway. British Journal for the Philosophy of Science, 52(3):641–645, September 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/3/641.full.pdf+
html;   http://www.jstor.
org/stable/3541936.

Putnam:2005:PLQ

[Put05] Hilary Putnam. A philosopher looks at quantum mechanics (again). British Journal for the Philosophy of Science, 56(4):615–634, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/615.full.pdf+html;
http://www.jstor.org/
stable/3541860.

Quine:1957:SLS

[Qui57] W. V. Quine. The scope and language of science. British Journal for the Philosophy of Science, VIII(29):1–17, May 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/29/1.full.pdf+html;
http://www.jstor.org/
stable/685377.

Quinton:1961:BRB

[Qui61] Anthony Quinton. Book review: Words and Things: by Ernest Gellner. Gollancz. 1959. 25s. British Journal for the Philosophy of Science, XI(44): 337–344, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/337.full.pdf+html;
http://www.jstor.org/
stable/685138.

Quinn:1974:TNS

[Qui74] Philip L. Quinn. The transitivity of non-standard synchronisms. British Journal for the Philosophy of Science, 25(1):78–82, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/78.full.pdf+

html;   http://www.jstor.
org/stable/686503.

Quick:1977:R

[Qui77a] Allison Quick. Reviews. British Journal for the Philosophy of Science, 28(4):383–389, December 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/4/383.full.pdf+
html;   http://www.jstor.
org/stable/686881.

Quine:1977:R

[Qui77b] W. V. Quine. Reviews. British Journal for the Philosophy of Science, 28(1):81–82, March 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/1/81.full.pdf+
html;   http://www.jstor.
org/stable/686451.

Quick:1979:R

[Qui79] Allison Quick. Reviews. British Journal for the Philosophy of Science, 30(2):199–202, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/199.full.pdf+
html;   http://www.jstor.
org/stable/686955.

Raatikainen:2005:TPC

[Raa05] Panu Raatikainen. Truth and provability — a comment on Redhead. British Journal for the Philosophy of Science, 56(3): 611–613, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/611.full.pdf+html;
http://www.jstor.org/
stable/3541722.

Radnitzky:1987:EAP

[Rad87] Gerard Radnitzky. The ‘economic’ approach to the philosophy of science. British Journal for the Philosophy of Science, 38(2):159–179, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/2/159.full.pdf+html;
http://www.jstor.org/
stable/687046.

Radder:1991:HGC

[Rad91a] Hans Radder. Heuristics and the generalized correspondence principle. British Journal for the Philosophy of Science, 42 (2):195–226, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/195.full.pdf+html;
http://www.jstor.org/
stable/688053.

Radnitzky:1991:RFM

[Rad91b] Gerard Radnitzky. Refined falsificationism meets the challenge from the relativist philosophy of science. British Journal for the Philosophy of Science, 42 (2):273–284, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/273.full.pdf+html;
http://www.jstor.org/
stable/688060.

Radick:2000:BRB

[Rad00] G. Radick. Book review: Darwin’s Spectre: Evolutionary Biology in the Modern World. M. R. Rose. British Journal for the Philosophy of Science, 51(3): 499–505, September 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/499.full.pdf+html;
http://www.jstor.org/
stable/3541778.

Raine:1984:R

[Rai84] Derek Raine. Reviews. British Journal for the Philosophy of Science, 35(2):188–191, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/188.full.pdf+
html;   http://www.jstor.
org/stable/686971.

Raine:1986:GOE

[Rai86] Derek J. Raine. Book review:

Gingerich, Owen (Ed.) [1984]: Astrophysics and twentiethcentury astronomy to 1950, The General History of Astronomy, Vol. 4A. Cambridge University Press. Pp. x + 198 (ISBN 0-521-24256-8). British Journal for the Philosophy of Science, 37(4):510–513, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687376.

Raidl:2020:OMO

[Rai20] Eric Raidl. Open-minded orthodox Bayesianism by epsilonconditionalization. British Journal for the Philosophy of Science, 71(1):139–176, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/71/1/
                                        139/5237402.

Rakic:1997:PPF

[Rak97] Nataša Rakić. Past, present, future, and Special Relativity. British Journal for the Philosophy of Science, 48(2):257–280, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/2/257.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687748.

Ramsey:1989:MKP

[Ram89] F. P. Ramsey. Mr Keynes on probability. British Journal for the Philosophy of Science, 40(2):219–222, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/

40/2/219.full.pdf+html;
http://www.jstor.org/
stable/687512.

Ramsey:1990:WVK

[Ram90] F. P. Ramsey. Weight or the value of knowledge. British Journal for the Philosophy of Science, 41(1):1–4, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/1/1.full.pdf+html;
http://www.jstor.org/
stable/688001.

Ramsey:2007:TS

[Ram07a] F. P. Ramsey. Truth and simplicity. British Journal for the Philosophy of Science, 58(3):379– 386, September 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/3/379.full.pdf+html;
http://www.jstor.org/
stable/30115180.

Ramsey:2007:RR

[Ram07b] William (William M.) Ramsey.

Representation reconsidered. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85987-5 (hardback). xx + 248 pp. LCCN BF316.6 .R355

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0713/2007002292-d.htm;
http://www.loc.gov/catdir/
enhancements/fy0729/
2007002292-b.htm;

                                        http://www.loc.gov/catdir/
                                        toc/ecip078/2007002292.htm.

Ramsey:2013:OTP

[Ram13]  Grant Ramsey.  Organisms,

traits, and population subdivisions: Two arguments against the causal conception of fitness? British Journal for the Philosophy of Science, 64(3): 589–608, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/3/589.full.pdf+html.

Ramirez:2023:PCLb

[Ram23] Sebastián Murgueitio Ramírez. A puzzle concerning local symmetries and their empirical significance. British Journal for the Philosophy of Science, 73(4): 1021–1044, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rasmussen:1953:BMP

[Ras53] E. Tranekjær Rasmussen. Berkeley and modern psychology. British Journal for the Philosophy of Science, IV(13):2–12, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IV/13/2.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686040.

Rattansi:1967:R

[Rat67] P. M. Rattansi. Reviews. British Journal for the Philosophy of Science, 18(3):250–255, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/250.full.pdf+
html;   http://www.jstor.
org/stable/686598.

Rattansi:1968:R

[Rat68] P. M. Rattansi. Reviews. British Journal for the Philosophy of Science, 19(3):270, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/270.full.pdf+
html;   http://www.jstor.
org/stable/686254.

Ratcliffe:2002:DPC

[Rat02] Matthew Ratcliffe. Dennett’s philosophy: a comprehensive assessment. British Journal for the Philosophy of Science, 53(4): 597–602, December 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
53/4/597.full.pdf+html;
http://www.jstor.org/
stable/3541829.

Ratcliffe:2004:HSP

[Rat04] Matthew Ratcliffe. How scientific practices matter: Reclaiming philosophical naturalism. British Journal for the Philosophy of Science, 55(1): 179–184, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/1/179.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541842.

Ratti:2020:MMR

[Rat20] Emanuele Ratti. ‘Models of’ and ‘models for’: On the relation between mechanistic models and experimental strategies in molecular biology. British Journal for the Philosophy of Science, 71(2):773–797, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/71/2/
                                        773/4939209.

Ravetz:1960:RU

[Rav60a] J. R. Ravetz. Review: [untitled]. British Journal for the Philosophy of Science, XI (41):77–78, May 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685826.

Ravetz:1960:Ra

[Rav60b] J. R. Ravetz. Reviews. British Journal for the Philosophy of Science, XI(41):77b–78, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XI/41/77-b.full.pdf+html.

Ravetz:1960:Rb

[Rav60c] J. R. Ravetz. Reviews. British Journal for the Philosophy of Science, XI(42):156–157, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/156.full.pdf+
html;   http://www.jstor.
org/stable/685592.

Ravetz:1960:Rc

[Rav60d] J. R. Ravetz. Reviews. British Journal for the Philosophy of Science, XI(42):167–168, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/167.full.pdf+
html;   http://www.jstor.
org/stable/685599.

Ravetz:1961:R

[Rav61] J. Ravetz. Reviews. British Journal for the Philosophy of Science, XII(47):250–255, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/47/250.full.
pdf+html;   http://www.
jstor.org/stable/685212.

Ravetz:1962:R

[Rav62] J. Ravetz. Reviews. British Journal for the Philosophy of Science, XII(48):344–346, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/48/344.full.
pdf+html;   http://www.
jstor.org/stable/685404.

Ravetz:1965:R

[Rav65] J. R. Ravetz. Reviews. British Journal for the Philosophy of Science, XV(60):354–356, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/60/354.full.pdf+
html;   http://www.jstor.
org/stable/686549.

Ravetz:1978:R

[Rav78] J. R. Ravetz. Reviews. British Journal for the Philosophy of Science, 29(4):395–397, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/4/395.full.pdf+
html;   http://www.jstor.
org/stable/687106.

Ravetz:1987:R

[Rav87] Jerome R. Ravetz. Reviews. British Journal for the Philosophy of Science, 38(2):265–268, June 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/2/265.full.pdf+
html;   http://www.jstor.
org/stable/687053.

Rawlins:1950:R

[Raw50] Ian Rawlins. Reviews. British Journal for the Philosophy of Science, I(3):249–254, November 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/3/249.full.pdf+
html;   http://www.jstor.
org/stable/685372.

Rawlins:1951:Ra

[Raw51a] Ian Rawlins. Reviews. British Journal for the Philosophy of Science, II(6):167, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/6/167.full.pdf+html;
http://www.jstor.org/
stable/685511.

Rawlins:1951:Rb

[Raw51b] Ian Rawlins. Reviews. British Journal for the Philosophy of Science, II(7):256–257, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/7/256.full.pdf+
html;   http://www.jstor.
org/stable/685299.

Rawlins:1951:TR

[Raw51c] Ian Rawlins. Theories of Relativity. British Journal for the Philosophy of Science, II(5):80– 81, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/80.full.pdf+
html;   http://www.jstor.
org/stable/685617.

Ramsey:2017:CFH

[RD17] Grant Ramsey and Andreas De Block. Is cultural fitness hopelessly confused? British Journal for the Philosophy of Science, 68(2):68–??, June 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/2/305/3044197/
Is-Cultural-Fitness-
Hopelessly-Confused.

Read:2004:RSS

[Rea04] Rupert Read. The road since ‘structure’. British Journal for the Philosophy of Science, 55 (1):175–178, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/175.full.pdf+html;
http://www.jstor.org/
stable/3541841.

Read:2020:FGE

[Rea20] James Read. Functional gravitational energy. British Journal for the Philosophy of Science, 71 (1):205–232, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
205/4098121.

Redhead:1974:NPT

[Red74a] M. L. G. Redhead. On Neyman’s Paradox and the theory of statistical tests. British Journal for the Philosophy of Science, 25(3):265–271, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/3/265.full.pdf+html;
http://www.jstor.org/
stable/686908.

Redhead:1974:OQM

[Red74b] M. L. G. Redhead. Orthodoxy in quantum mechanics. British Journal for the Philosophy of Science, 25(4):352– 357, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/4/352.full.pdf+html.

Redhead:1974:ROQ

[Red74c] M. L. G. Redhead. Review:

Orthodoxy in quantum mechanics. British Journal for the Philosophy of Science, 25(4): 352–358, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686465.

Redhead:1977:RAW

[Red77] M. L. G. Redhead. Review articles: Wave–particle duality. British Journal for the Philosophy of Science, 28(1):65–74, March 1977. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/1/65.full.pdf+
html;   http://www.jstor.
org/stable/686449.

Redhead:1978:DNA

[Red78] M. Redhead. Discussions notes: Ad hocness and the appraisal of theories. British Journal for the Philosophy of Science, 29(4):355–361, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/355.full.pdf+html;
http://www.jstor.org/
stable/687100.

Redhead:1980:MP

[Red80a] Michael Redhead. Models in physics. British Journal for the Philosophy of Science, 31 (2):145–163, June 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/2/145.full.pdf+html;
http://www.jstor.org/
stable/687183.

Redhead:1980:R

[Red80b] Michael Redhead. Reviews. British Journal for the Philosophy of Science, 31(4):404– 408, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/4/404.full.pdf+html;

http://www.jstor.org/
stable/687069.

Redhead:1981:Ra

[Red81a] Michael Redhead. Reviews.

British Journal for the Philosophy of Science, 32(3):309– 311, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/309.full.pdf+html;
http://www.jstor.org/
stable/687596.

Redhead:1981:Rb

[Red81b] Michael Redhead. Reviews.

British Journal for the Philosophy of Science, 32(3):311– 314, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/311.full.pdf+html;
http://www.jstor.org/
stable/687597.

Redhead:1984:US

[Red84] M. L. G. Redhead. Unification in science. British Journal for the Philosophy of Science, 35(3):274–279, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/274.full.pdf+html;
http://www.jstor.org/
stable/687477.

Redhead:1985:IIP

[Red85a] Michael Redhead. On the impossibility of inductive probability. British Journal for the Philosophy of Science, 36(2):185–191, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/185.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687035.

Redhead:1985:R

[Red85b] Michael Redhead. Reviews. British Journal for the Philosophy of Science, 36(1):100–104, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/1/100.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687164.

Redhead:1986:NC

[Red86] Michael Redhead. Novelty and confirmation. British Journal for the Philosophy of Science, 37 (1):115–118, March 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687001.

Redhead:1989:LAL

[Red89] Michael Redhead. The Lakatos Award Lecture: the nature of reality. British Journal for the Philosophy of Science, 40(4): 429–441, December 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/4/429.full.pdf+html;
http://www.jstor.org/
stable/687734.

Redhead:1996:R

[Red96]  Michael Redhead.  Review.

British Journal for the Philosophy of Science, 47(1):133–136, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/133.full.pdf+
html;   http://www.jstor.
org/stable/1215832.

Redhead:1998:R

[Red98]  Michael Redhead.  Review.

British Journal for the Philosophy of Science, 49(3):499– 507, September 1998. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/3/499.full.pdf+html;
http://www.jstor.org/
stable/688086.

Redhead:2000:BRR

[Red00] Michael Redhead. Book review: Roger Penrose — The Large, the Small and the Human Mind. British Journal for the Philosophy of Science, 51(4):913–917, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/913.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/3541736.

Redhead:2004:MM

[Red04] Michael Redhead. Mathematics and the mind. British Journal for the Philosophy of Science, 55(4):731–737, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/4/731.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541625.

Reed:1978:GSM

[Ree78] Edward Reed. Group selection and methodological individualism: a criticism of Watkins. British Journal for the Philosophy of Science, 29(3):256– 262, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/3/256.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687015.

Reeves:1988:TP

[Ree88] T. V. Reeves. A theory of probability. British Journal for the Philosophy of Science, 39 (2):161–182, June 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        39/2/161.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687264.

Reichert:2023:EEB

[Rei23] Paula Reichert. Essentially ergodic behaviour. British Journal for the Philosophy of Science, 74 (1):57–73, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rescher:1958:PE

[Res58] Nicholas Rescher. On prediction and explanation. British Journal for the Philosophy of Science, VIII(32):281–290, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/281.full.
pdf+html;   http://www.
jstor.org/stable/685853.

Rescher:1966:NLP

[Res66] Nicholas Rescher. A new look at the problem of innate ideas. British Journal for the Philosophy of Science, 17(3):205– 218, November 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/3/205.full.pdf+html;
http://www.jstor.org/
stable/686757.

Resnik:1987:R

[Res87] David B. Resnik. Reviews.

British Journal for the Philosophy of Science, 38(1):119–121, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/38/1/119.full.pdf+
html;   http://www.jstor.
org/stable/687087.

Rescorla:2009:CML

[Res09] Michael Rescorla. Cognitive maps and the language of thought. British Journal for the Philosophy of Science, 60(2):377–407, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/60/2/377.full.pdf+
html;   http://www.jstor.
org/stable/25592005.

Rescorla:2013:AST

[Res13] Michael Rescorla. Against structuralist theories of computational implementation. British Journal for the Philosophy of Science, 64(4):681–707, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/681.full.pdf+html.

Reutlinger:2009:BRM

[Reu09] Alexander Reutlinger. Book review: Markus Schrenk, The Metaphysics of Ceteris Paribus Laws. British Journal for the Philosophy of Science, 60(1): 229–233, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/229.full.pdf+html;

http://www.jstor.org/
stable/25591995.

Reydon:2009:GNP

[Rey09] Thomas A. C. Reydon. Gene names as proper names of individuals: an assessment. British Journal for the Philosophy of Science, 60(2):409–432, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/60/2/409.full.pdf+
html;   http://www.jstor.
org/stable/25592006.

Reydon:2015:MHS

[Rey15] Thomas A. C. Reydon. Book review: Matthew H. Slater, Are Species Real? An Essay on the Metaphysics of Species. British Journal for the Philosophy of Science, 66(4):1029–1033, December 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Roll-Hansen:1972:LPC

[RH72] Nils Roll-Hansen. Louis Pasteur — a case against reductionist historiography. British Journal for the Philosophy of Science, 23(4):347–361, November 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/4/347.full.pdf+
html;   http://www.jstor.
org/stable/686738.

Roll-Hansen:1977:R

[RH77]  Nils Roll-Hansen.  Reviews.

British Journal for the Philosophy of Science, 28(3):291– 292, September 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        28/3/291.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686813.

Roll-Hansen:1979:R

[RH79]  Nils Roll-Hansen.  Reviews.

British Journal for the Philosophy of Science, 30(1):93–96, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/1/93.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687421.

Roll-Hansen:1981:R

[RH81]  Nils Roll-Hansen.  Reviews.

British Journal for the Philosophy of Science, 32(1):104–106, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/1/104.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687392.

Reutlinger:2018:UTM

[RHH18] Alexander Reutlinger, Dominik Hangleiter, and Stephan Hartmann. Understanding (with) toy models. British Journal for the Philosophy of Science, 69(4):1069–1099, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/1069/3952637.

Richens:1951:R

[Ric51] R. H. Richens. Reviews. British Journal for the Philosophy of Science, II(6):169–170, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/169.full.pdf+
html;   http://www.jstor.
org/stable/685513.

Richens:1952:RU

[Ric52a] R. H. Richens. Review: [untitled]. British Journal for the Philosophy of Science, III(11): 279–281, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685278.

Richens:1952:R

[Ric52b]  R. H. Richens.  Reviews.

British Journal for the Philosophy of Science, III(11):279b– 281, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/279-b.full.pdf+html.

Richards:1977:EEH

[Ric77] Joan L. Richards. The evolution of empiricism: Hermann von Helmholtz and the foundations of geometry. British Journal for the Philosophy of Science, 28(3):235–253, September 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/3/235.full.pdf+html;
http://www.jstor.org/
stable/686808.

Richardson:1995:R

[Ric95]  Alan Richardson.  Review.

British Journal for the Philosophy of Science, 46(4):617– 620, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/4/617.full.pdf+html;
http://www.jstor.org/
stable/687903.

Richardson:2007:EPM

[Ric07] Robert C. Richardson. Evolutionary psychology as maladapted psychology. Life and mind. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-18260-2 (hardcover). xi + 213 pp. LCCN BF698.95 .R44

2007. URL http://www.loc.
gov/catdir/toc/ecip071/
2006030807.html.

Rickles:2016:BRR

[Ric16] Dean Rickles. Book review: Richard Dawid, String Theory and the Scientific Method. British Journal for the Philosophy of Science, 67(3):925– 929, September 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/3/925.

Rice:2019:MDD

[Ric19] Collin Rice. Models don’t decompose that way: a holistic view of idealized models. British Journal for the Philosophy of Science, 70 (1):179–208, March 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
179/4098118.

Riggs:1991:CMA

[Rig91] Peter J. Riggs. A critique of Mellor’s argument against ‘backwards’ causation. British Journal for the Philosophy of Science, 42(1):75–86, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/1/75.full.pdf+html;
http://www.jstor.org/
stable/687607.

Ritchie:1962:R

[Rit62] A. M. Ritchie. Reviews. British Journal for the Philosophy of Science, XIII(49):75– 80, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/75.full.
pdf+html;   http://www.
jstor.org/stable/685232.

Rivadulla:1994:PSP

[Riv94] Andres Rivadulla. Probabilistic support, probabilistic induction and Bayesian confirmation theory. British Journal for the Philosophy of Science, 45 (2):477–483, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/477.full.pdf+html;
http://www.jstor.org/
stable/687677.

Resnik:1987:EIR

[RK87] Michael D. Resnik and David Kushner. Explanation, independence and realism in mathematics. British Journal for the Philosophy of Science, 38 (2):141–158, June 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/2/141.full.pdf+html;
http://www.jstor.org/
stable/687045.

Ritchie:2019:DBN

[RKK19] J. Brendan Ritchie, David Michael Kaplan, and Colin Klein. Decoding the brain: Neural representation and the limits of multivariate pattern analysis in cognitive neuroscience. British Journal for the Philosophy of Science, 70(2):581–607, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.

oup.com/bjps/article/70/2/
581/4105010.

Rescher:1955:LAG

[RO55] Nicholas Rescher and Paul Oppenheim. Logical analysis of gestalt concepts. British Journal for the Philosophy of Science, VI (22):89–106, August 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VI/22/89.full.pdf+html;
http://www.jstor.org/
stable/685519.

Robinson:1965:R

[Rob65] Richard Robinson. Reviews.

British Journal for the Philosophy of Science, XVI(63):255– 257, November 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/63/255.full.pdf+html;
http://www.jstor.org/
stable/686312.

Robinson:1968:R

[Rob68] Richard E. Robinson. Reviews.

British Journal for the Philosophy of Science, 19(2):177–179, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/177.full.pdf+
html;   http://www.jstor.
org/stable/686801.

Robinson:1976:WSP

[Rob76] Daniel N. Robinson. What sort of persons are hemispheres? Another look at ‘split-brain’ man. British Journal for the Philosophy of Science, 27(1):73–78, March 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/27/1/73.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686381.

Robinson:1994:CSR

[Rob94] Don Robinson. Can superselection rules solve the measurement problem? British Journal for the Philosophy of Science, 45(1):79–93, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/1/79.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687962.

Roberts:2008:LGU

[Rob08a] John T. Roberts. The Law-Governed Universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-955770-5. xii + 407 pp. LCCN B105.L3 R63 2008. URL

                                        http://www.loc.gov/catdir/
                                        toc/fy0903/2008036352.html.

Roberts:2008:PAL

[Rob08b] John T. Roberts. A puzzle about laws, symmetries and measurability. British Journal for the Philosophy of Science, 59 (2):143–168, June 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/2/143.full.pdf+html;
http://www.jstor.org/
stable/40072275.

Roberts:2011:GSR

[Rob11] Bryan W. Roberts. Group structural realism. British Journal for the Philosophy of Science, 62(1):47–69, March 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/47.full.pdf+html.

Roberts:2013:CC

[Rob13] John T. Roberts. Chance without credence. British Journal for the Philosophy of Science, 64(1):33–59, March 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/33.full.pdf+html.

Robertson:2020:AAA

[Rob20] Katie Robertson. Asymmetry, abstraction, and autonomy: Justifying coarse-graining in statistical mechanics. British Journal for the Philosophy of Science, 71(2):547–579, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
547/5045454.

Robertson:2023:SHG

[Rob23] Katie Robertson. In search of the Holy Grail: How to reduce the Second Law of Thermodynamics. British Journal for the Philosophy of Science, 73(4):987– 1020, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rockmore:1986:RU

[Roc86] Tom Rockmore. Book review: Tiles, Mary [1984]: Bachelard: Science and Objectivity. Cambridge University Press. xxii + 242 pp. (ISBN 0-521-24803-5 hard covers; 0-521-28973-4 paperback). British Journal for the Philosophy of Science, 37(4): 529–531, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687382.

Roessler:2004:CW

[Roe04] Johannes Roessler. Consciousness and the world. British Journal for the Philosophy of Science, 55(1):163–173, March 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/1/163.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541840.

Rohrlich:1988:POT

[Roh88] Fritz Rohrlich. Pluralistic ontology and theory reduction in the physical sciences. British Journal for the Philosophy of Science, 39(3):295–312, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/295.full.pdf+
html;   http://www.jstor.
org/stable/687208.

Roland:2007:MMN

[Rol07] Jeffrey W. Roland. Maddy and mathematics: Naturalism or not. British Journal for the Philosophy of Science, 58(3): 423–450, September 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/3/423.full.pdf+html;
http://www.jstor.org/
stable/30115183.

Roshwald:1955:VJS

[Ros55] M. Roshwald. Value-judgments in the social sciences. British Journal for the Philosophy of Science, VI(23):186–208, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/186.full.pdf+
html;   http://www.jstor.
org/stable/685676.

Roshwald:1956:CI

[Ros56] M. Roshwald. The case for indeterminism. British Journal for the Philosophy of Science, VII(27):240–242, November 1956. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/27/240.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685879.

Rosenfeld:1958:R

[Ros58] L. Rosenfeld. Reviews. British Journal for the Philosophy of Science, IX(34):167–168, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/34/167.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685156.

Ross:1970:R

[Ros70]  George M. Ross.  Reviews.

British Journal for the Philosophy of Science, 21(3):298–301, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/3/298.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686205.

Rosser:1978:DNC

[Ros78] W. G. V. Rosser. Discussions notes: the clock hypothesis and the Lorentz transformations. British Journal for the Philosophy of Science, 29(4): 349–353, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/4/349.full.pdf+html;

http://www.jstor.org/
stable/687098.

Rosenkrantz:1994:BCP

[Ros94] R. D. Rosenkrantz. Bayesian confirmation: Paradise regained. British Journal for the Philosophy of Science, 45(2):467–476, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/467.full.pdf+
html;   http://www.jstor.
org/stable/687676.

Rosenberg:1996:FGR

[Ros96] Alex Rosenberg. A field guide to recent species of naturalism. British Journal for the Philosophy of Science, 47(1):1–29, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/1.full.pdf+
html;   http://www.jstor.
org/stable/1215824.

Rosen:1999:BRB

[Ros99] G. Rosen. Book review: Naturalism in mathematics. Penelope Maddy. British Journal for the Philosophy of Science, 50(3):467–474, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/467.full.pdf+html;
http://www.jstor.org/
stable/40072243.

Rosenberg:2001:HBE

[Ros01] Alex Rosenberg. How is biological explanation possible? British Journal for the Philosophy of Science, 52(4):735–760, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/52/4/735.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541917.

Ross:2008:FGM

[Ros08]  Don Ross.  Book review:

Francesco Guala, The Methodology of Experimental Economics. British Journal for the Philosophy of Science, 59(2):247–252, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/2/247.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072281.

Ross:2009:BRP

[Ros09] Don Ross. Book review: Philip Mirowski, The Effortless Economy of Science? British Journal for the Philosophy of Science, 60(3):659–665, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/3/659.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592024.

Roskies:2010:SSR

[Ros10] Adina L. Roskies. Saving subtraction: a reply to Van Orden and Paap. British Journal for the Philosophy of Science, 61(3):635–665, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/635.full.pdf+html.

Rosenberg:2012:ECW

[Ros12] Alex Rosenberg. Editor’s choice:

Why do spatiotemporally restricted regularities explain in the social sciences? British Journal for the Philosophy of Science, 63(1):1–26, March 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/1/1.full.pdf+html.

Rosaler:2019:RPR

[Ros19]  Joshua Rosaler.  Reduction

as an a posteriori relation. British Journal for the Philosophy of Science, 70(1):269–299, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/70/1/269/4092771.

Ross:2021:CCB

[Ros21] Lauren N. Ross. Causal concepts in biology: How pathways differ from mechanisms and why it matters. British Journal for the Philosophy of Science, 72(1): 131–158, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rowan:1993:SRI

[Row93] Michael Rowan. Stove on the rationality of induction and the uniformity thesis. British Journal for the Philosophy of Science, 44(3):561–566, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/561.full.pdf+html;
http://www.jstor.org/
stable/688025.

Rowlands:1994:CLT

[Row94] Mark Rowlands. Connectionism and the language of thought. British Journal for the Philosophy of Science, 45(2):485–503, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/485.full.pdf+
html;   http://www.jstor.
org/stable/687678.

Rowbottom:2011:PCR

[Row11] Darrell P. Rowbottom. Popper’s critical rationalism: a philosophical investigation. Routledge studies in the philosophy of science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2011. ISBN 0-415-99244-3. xii + 185 pp. LCCN B1649.P64 R69

2011. URL http://www.loc.
gov/catdir/enhancements/
fy1114/2010022916-b.html.

Rowbottom:2013:PMC

[Row13] Darrell P. Rowbottom. Popper’s measure of corroboration and P (h|b). British Journal for the Philosophy of Science, 64(4):739–745, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/739.full.pdf+html.

Rowbottom:2023:PRM

[Row23] Darrell P. Rowbottom. Peer review may not be such a bad idea: Response to Heesen and Bright. British Journal for the Philosophy of Science, 73(4):927–940, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Roxburgh:1975:INS

[Rox75] I. W. Roxburgh. The intransitivity of non standard synchronisms. British Journal for the Philosophy of Science, 26 (1):47–49, March 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/1/47.full.pdf+html;
http://www.jstor.org/
stable/686332.

Roxburgh:1977:CUT

[Rox77] Ian W. Roxburgh. The conventionality of uniform time. British Journal for the Philosophy of Science, 28(2):172–177, June 1977. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/2/172.full.pdf+
html;   http://www.jstor.
org/stable/686632.

Rozeboom:1969:NMO

[Roz69] William W. Rozeboom. New mysteries for old: the transfiguration of Miller’s Paradox. British Journal for the Philosophy of Science, 19(4):345– 353, February 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/4/345.full.pdf+html;
http://www.jstor.org/
stable/686293.

Rueger:1996:STM

[RS96] Alexander Rueger and W. David Sharp. Simple theories of a messy world: Truth and explanatory power in nonlinear dynamics. British Journal for the Philosophy of Science, 47(1):93–112, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/1/93.full.pdf+
html;   http://www.jstor.
org/stable/1215828.

Ross:2007:NCR

[RS07] Don Ross and David Spurrett. Notions of cause: Russell’s thesis revisited. British Journal for the Philosophy of Science, 58(1):45–76, March 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/1/45.full.pdf+html;
http://www.jstor.org/
stable/4489100.

Reydon:2015:SDG

[RS15] Thomas A. C. Reydon and Markus Scholz. Searching for Darwinism in generalized Darwinism. British Journal for the Philosophy of Science, 66(3): 561–589, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/561.

Roche:2018:II

[RS18]  William  Roche

and Tomoji Shogenji. Information and inaccuracy. British Journal for the Philosophy of Science, 69(2):577–604, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/577/2433517.

Ruiz:2023:MFM

[RS23] Nadia Ruiz and Armin W.

Schulz. Micro-foundations and methodology: a complexitybased reconceptualization of the debate. British Journal for the Philosophy of Science, 74(2): 359–379, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Redhead:1992:PLT

[RT92]  Michael  Redhead  and

Paul Teller. Particle labels and the theory of indistinguishable particles in quantum mechanics. British Journal for the Philosophy of Science, 43(2):201–218, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/2/201.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687872.

Rubin:1980:R

[Rub80] Ronald Rubin. Reviews. British Journal for the Philosophy of Science, 31(1):104–108, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/31/1/104.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687259.

Ruben:1981:R

[Rub81] David-Hillel Ruben. Reviews. British Journal for the Philosophy of Science, 32(2):210–217, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/2/210.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687204.

Ruben:1982:R

[Rub82] David-Hillel Ruben. Reviews. British Journal for the Philosophy of Science, 33(4):438– 441, December 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/4/438.full.pdf+html;
http://www.jstor.org/
stable/687175.

Ruben:1986:APN

[Rub86] David-Hillel Ruben. Book review: Achinstein, Peter [1983]: The Nature of Explanation. Oxford University Press. ix + 385 pp. (ISBN 0-19-503215-2). British Journal for the Philosophy of Science, 37(3):377– 384, September 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687300.

Rubenstein:2022:GSG

[Rub22a] Ezra Rubenstein. Grounded shadows, groundless ghosts. British Journal for the Philosophy of Science, 73(3):723–750, September 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rubin:2022:CH

[Rub22b] Mark Rubin. The costs of HARKing. British Journal for the Philosophy of Science, 73(2): 535–560, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Rudwick:1964:IFS

[Rud64] M. J. S. Rudwick. The inference of function from structure in fossils. British Journal for the Philosophy of Science, XV(57):27– 40, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/57/27.full.pdf+
html;   http://www.jstor.
org/stable/685625.

Ruetsche:1998:R

[Rue98] Laura Ruetsche. Review. British Journal for the Philosophy of Science, 49(4):637–641, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/4/637.full.pdf+
html;   http://www.jstor.
org/stable/688135.

Rueger:2005:PMT

[Rue05] Alexander Rueger. Perspectival models and theory unification. British Journal for the Philosophy of Science, 56(3):579– 594, September 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/3/579.full.pdf+html;
http://www.jstor.org/
stable/3541720.

Ruetsche:2011:IQT

[Rue11] Laura Ruetsche. Interpreting quantum theories: the art of the possible. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-953540-X, 0-19-172852-7 (e-book), 0-19-161737-7 (e-book). xvii + 379 pp. LCCN QC174.12 .R848

2011. URL http://www.loc.
gov/catdir/enhancements/
fy1117/2011923708-b.html;
http://www.loc.gov/catdir/
enhancements/fy1117/
2011923708-d.html;
http://www.loc.gov/catdir/
enhancements/fy1117/
2011923708-t.html.

Rupert:2007:RCC

[Rup07] Robert D. Rupert. Realization, completers, and ceteris paribus laws in psychology. British Journal for the Philosophy of Science, 58(1):1–11, March 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/1/1.full.pdf+html;
http://www.jstor.org/
stable/4489097.

Russell:1950:DEL

[Rus50] E. S. Russell. The ‘drive’ element in life. British Journal for the Philosophy of Science, I(2):108–116, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/2/108.full.pdf+html;
http://www.jstor.org/
stable/685806.

Russell:1951:CMS

[Rus51] Roger W. Russell. The comparative method in the study of behaviour. British Journal for the Philosophy of Science, II(7):251–255, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/7/251.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685297.

Russell:1952:R

[Rus52] R. W. Russell. Reviews. British Journal for the Philosophy of Science, III(10):202–204, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/10/202.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685563.

Russell:1953:CCU

[Rus53] Bertrand Russell. The cult of ‘common usage’. British Journal for the Philosophy of Science, III(12):303–307, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/303.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685443.

Russell:1954:NPB

[Rus54] L. J. Russell. Note on a passage: Berkeley’s De Motu. British Journal for the Philosophy of Science, V(18):152, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/152.full.pdf+html;
http://www.jstor.org/
stable/685174.

Russell:1955:RU

[Rus55a] L. J. Russell. Review: [untitled]. British Journal for the Philosophy of Science, V (20):351, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685745.

Russell:1955:Ra

[Rus55b] L. J. Russell. Reviews. British Journal for the Philosophy of Science, V(20):346–348, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/346.full.pdf+
html;   http://www.jstor.
org/stable/685742.

Russell:1955:Rb

[Rus55c] L. J. Russell. Reviews. British Journal for the Philosophy of Science, V(20):351a–351, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
V/20/351-a.full.pdf+html.

Russell:1955:Rc

[Rus55d] L. J. Russell. Reviews. British Journal for the Philosophy of Science, VI(23):254–255, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/254.full.pdf+
html;   http://www.jstor.
org/stable/685687.

Russell:1956:R

[Rus56] L. J. Russell. Reviews. British Journal for the Philosophy of Science, VI(24):343–345, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/343.full.pdf+
html;   http://www.jstor.
org/stable/685418.

Russell:1957:R

[Rus57] L. J. Russell. Reviews. British Journal for the Philosophy of Science, VII (28):363a–363, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/363-a.full.pdf+html.

Russell:1960:R

[Rus60] L. J. Russell. Reviews. British Journal for the Philosophy of Science, X(40):349–350, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/40/349.full.pdf+
html;   http://www.jstor.
org/stable/685926.

Ruse:1969:DSB

[Rus69] Michael Ruse. Definitions of species in biology. British Journal for the Philosophy of Science, 20(2):97–119, August 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
20/2/97.full.pdf+html;
http://www.jstor.org/
stable/686173.

Ruse:1971:SPR

[Rus71] Michael Ruse. The species problem: a reply to Hull. British Journal for the Philosophy of Science, 22(4):369–371, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/4/369.full.pdf+
html;   http://www.jstor.
org/stable/686319.

Ruse:1987:BSN

[Rus87] Michael Ruse. Biological species:

Natural kinds, individuals, or what? British Journal for the Philosophy of Science, 38 (2):225–242, June 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/2/225.full.pdf+html;
http://www.jstor.org/
stable/687049.

Ruse:2000:BRB

[Rus00] M. Ruse. Book review: Gay Science: the Ethics of Sexual Orientation Research. T. F. Murphy. British Journal for the Philosophy of Science, 51(3):487– 493, September 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/487.full.pdf+html;
http://www.jstor.org/
stable/3541776.

Russo:2008:CCM

[Rus08] Federica Russo. Causality and causal modelling in the social sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-8816-7. xiv + 235 pp. LCCN H61.R867 2009.

Ruyant:2019:SRM

[Ruy19] Quentin Ruyant. Structural realism or modal empiricism? British Journal for the Philosophy of Science, 70(4):1051– 1072, December 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1051/4935149.

Read:2013:BRB

[RW13] Rupert Read and Jessica Woolley. Book review: Brad Wray: Kuhn’s Evolutionary Social Epistemology. British Journal for the Philosophy of Science, 64(3): 659–664, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
64/3/659.full.pdf+html.

Ryckman:1999:BRC

[Ryc99] T. A. Ryckman. Book review: Carnap’s construction of the world: the Aufbau and the emergence of logical empiricism. British Journal for the Philosophy of Science, 50(3): 497–500, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/497.full.pdf+html;
http://www.jstor.org/
stable/40072249.

Ryder:1981:CSE

[Ryd81] J. M. Ryder. Consequences of a simple extension of the Dutch Book argument. British Journal for the Philosophy of Science, 32(2):164–167, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/2/164.full.pdf+html;
http://www.jstor.org/
stable/687197.

Rynasiewicz:1983:FSE

[Ryn83] Robert A. Rynasiewicz. Falsifiability and the semantic eliminability of theoretical languages. British Journal for the Philosophy of Science, 34(3):225– 241, September 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        34/3/225.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687321.

Rynasiewicz:1994:LHA

[Ryn94] Robert Rynasiewicz. The lessons of the hole argument. British Journal for the Philosophy of Science, 45(2):407–436, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/2/407.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687673.

Saatsi:2011:EIA

[Saa11] Juha Saatsi. The enhanced indispensability argument: Representational versus explanatory role of mathematics in science. British Journal for the Philosophy of Science, 62(1): 143–154, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/1/143.full.pdf+html.

Saatsi:2018:EGM

[Saa18] Juha Saatsi. On explanations from geometry of motion. British Journal for the Philosophy of Science, 69(1):253–273, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/1/253/2669765.

Sabra:1954:NSM

[Sab54] A. I. Sabra. A note on a suggested modification of Newton’s corpuscular theory of light to reconcile it with Foucault’s experiment of 1850. British Journal for the Philosophy of Science, V (18):149–151, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/149.full.pdf+html;
http://www.jstor.org/
stable/685172.

Sabra:1961:R

[Sab61] Nancy Sutton Sabra. Reviews.

British Journal for the Philosophy of Science, XII(45):72–73, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/72.full.pdf+
html;   http://www.jstor.
org/stable/685544.

Sachs:1964:NAT

[Sac64] Mendel Sachs. A new approach to the theory of fundamental processes. British Journal for the Philosophy of Science, XV(59):213–243, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/213.full.pdf+
html;   http://www.jstor.
org/stable/686525.

Sachs:1970:QRN

[Sac70] M. Sachs. Is quantization really necessary? British Journal for the Philosophy of Science, 21(4):359–370, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/4/359.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686557.

Sachs:1972:MPR

[Sac72] Mendel Sachs. On the Mach Principle and relative space– time. British Journal for the Philosophy of Science, 23(2): 117–119, May 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/2/117.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686435.

Sachs:1975:MPG

[Sac75] Mendel Sachs. On the Mach Principle and General Relativity. British Journal for the Philosophy of Science, 26(1):49–51, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/1/49.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686333.

Sachs:1976:LSE

[Sac76] Mendel Sachs. On the logical status of equivalence principles in General Relativity theory. British Journal for the Philosophy of Science, 27(3):225– 229, September 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/3/225.full.pdf+html;
http://www.jstor.org/
stable/686120.

Sachs:1989:OSR

[Sac89] Mendel Sachs. On the origin of spin in Relativity. British Journal for the Philosophy of Science, 40(3):409–412, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/3/409.full.pdf+
html;   http://www.jstor.
org/stable/687788.

Sachs:1993:HBH

[Sac93] Mendel Sachs. On Hawking’s A Brief History of Time and The Present State of Physics. British Journal for the Philosophy of Science, 44(3):543–547, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/3/543.full.pdf+
html;   http://www.jstor.
org/stable/688022.

Safran:1986:BRG

[Saf86]  J.  N.

Safran. Book review: Goodman, N. [1984]: Of Mind and Other Matters. Harvard University Press. 210 pp. British Journal for the Philosophy of Science, 37(2):242–246, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686987.

Sainsbury:1985:R

[Sai85] R. M. Sainsbury. Reviews. British Journal for the Philosophy of Science, 36(2):211–215, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/211.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687037.

Sainsbury:1995:VIM

[Sai95] R. M. Sainsbury. Vagueness, ignorance, and margin for error. British Journal for the Philosophy of Science, 46(4): 589–601, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        46/4/589.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687900.

Salmon:1981:RP

[Sal81] Wesley C. Salmon. Rational prediction. British Journal for the Philosophy of Science, 32 (2):115–125, June 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
32/2/115.full.pdf+html;
http://www.jstor.org/
stable/687193.

Sambursky:1960:SCI

[Sam60]  S. Sambursky.  Some com-

ments on ‘imaginary experiments’. British Journal for the Philosophy of Science, XI (41):62–64, May 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/41/62.full.pdf+html;
http://www.jstor.org/
stable/685822.

Sampson:1975:TCT

[Sam75]  Geoffrey Sampson.  Theory

choice in a two-level science. British Journal for the Philosophy of Science, 26(4):303– 318, December 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/4/303.full.pdf+html;
http://www.jstor.org/
stable/686677.

Sampson:1978:DND

[Sam78]  Geoffrey Sampson.  Discus-

sions notes: a dilemma defended. British Journal for the Philosophy of Science, 29(4): 353–355, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/353.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687099.

Sampson:1980:PLA

[Sam80] Geoffrey Sampson. Popperian language-acquisition undefeated. British Journal for the Philosophy of Science, 31(1):63–67, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/31/1/63.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687252.

Sampson:1981:R

[Sam81] Geoffrey Sampson. Reviews. British Journal for the Philosophy of Science, 32(1):98–104, March 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/1/98.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687391.

Samuels:1998:EPM

[Sam98] Richard Samuels. Evolutionary psychology and the massive modularity hypothesis. British Journal for the Philosophy of Science, 49(4):575–602, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        49/4/575.full.pdf+html.

Samaroo:2018:TNC

[Sam18] Ryan Samaroo. There is no conspiracy of inertia. British Journal for the Philosophy of Science, 69(4):957–982, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/957/3831694.

Santirocco:1961:NAF

[San61] Raymond A. Santirocco. Note on the axiomatic formulation of electrostatics. British Journal for the Philosophy of Science, XII(46):132–145, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/46/132.full.pdf+html;
http://www.jstor.org/
stable/685492.

Sanford:1980:ATU

[San80]  David H. Sanford.  Arm-

strong’s theory of universals. British Journal for the Philosophy of Science, 31(1):69–79, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/1/69.full.pdf+
html;   http://www.jstor.
org/stable/687253.

Sankey:1993:KCC

[San93] Howard Sankey. Kuhn’s changing concept of incommensurability. British Journal for the Philosophy of Science, 44(4): 759–774, December 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/4/759.full.pdf+html;
http://www.jstor.org/
stable/688043.

Sandborg:1998:MET

[San98] David Sandborg. Mathematical explanation and the theory of why-questions. British Journal for the Philosophy of Science, 49(4):603–624, December 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/4/603.full.pdf+html;
http://www.jstor.org/
stable/688133.

Sansom:2008:CKC

[San08]  Roger Sansom.  Countering

Kauffman with connectionism: Two views of gene regulation and the fundamental nature of ontogeny. British Journal for the Philosophy of Science, 59 (2):169–200, June 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/2/169.full.pdf+html;
http://www.jstor.org/
stable/40072276.

Santana:2019:WA

[San19] Carlos Santana. Waiting for the Anthropocene. British Journal for the Philosophy of Science, 70 (4):1073–1096, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1073/4935146.

Sarndal:1968:SAC

[Sär68] Carl-Erik Särndal. Some aspects of Carnap’s theory of inductive inference. British Journal for the Philosophy of Science, 19(3):225–246, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/3/225.full.pdf+html;
http://www.jstor.org/
stable/686242.

Sarkar:2015:ECG

[Sar15] Sahotra Sarkar. Editor’s choice:

The genomic challenge to adaptationism. British Journal for the Philosophy of Science, 66(3): 505–536, September 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/3/505.

Saunders:1964:RDP

[Sau64] L. R. Saunders. Rational deduction in physics: the parallelogram of forces. British Journal for the Philosophy of Science, XIV(56):265–273, February 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/56/265.full.
pdf+html;   http://www.
jstor.org/stable/685633.

Saunders:1987:R

[Sau87] P. T. Saunders. Reviews. British Journal for the Philosophy of Science, 38(3):412–414, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/3/412.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687500.

Saunders:1996:CL

[Sau96] Simon Saunders. Comment on Lockwood. British Journal for the Philosophy of Science, 47 (2):241–248, June 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        47/2/241.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687946.

Savitt:1994:CMT

[Sav94] Steven F. Savitt. Is classical mechanics time reversal invariant? British Journal for the Philosophy of Science, 45(3): 907–913, September 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/3/907.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687800.

Savitt:1996:DT

[Sav96] Steven F. Savitt. The direction of time. British Journal for the Philosophy of Science, 47(3):347–370, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/3/347.full.pdf+html;
http://www.jstor.org/
stable/687983.

Sagal:1993:RPP

[SB93] Paul Sagal and Gunnar Borg.

The range principle and the problem of other minds. British Journal for the Philosophy of Science, 44(3):477–491, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/3/477.full.pdf+
html;   http://www.jstor.
org/stable/688017.

Shea:2010:VSS

[SB10] Nicholas Shea and Tim Bayne.

The vegetative state and the science of consciousness. British Journal for the Philosophy of Science, 61(3):459–484, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
61/3/459.full.pdf+html.

Stefansson:2019:WRA

[SB19] H. Orri Stefánsson and Richard Bradley. What is risk aversion? British Journal for the Philosophy of Science, 70(1):77–102, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/70/1/77/4091082.

Sebens:2018:SLU

[SC18] Charles T. Sebens and Sean M. Carroll. Self-locating uncertainty and the origin of probability in Everettian quantum mechanics. British Journal for the Philosophy of Science, 69(1):25–74, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/1/25/2669754. See

critique [DF22].

Scarantino:2010:IBC

[Sca10] Andrea Scarantino. Insights and blindspots of the cognitivist theory of emotions. British Journal for the Philosophy of Science, 61(4):729–768, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/4/729.full.pdf+html.

Scerri:1991:ECM

[Sce91] Eric R. Scerri. The electronic configuration model, quantum mechanics and reduction. British Journal for the Philosophy of Science, 42(3):309–325, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/42/3/309.full.pdf+

html;   http://www.jstor.
org/stable/687910.

Schrodinger:1952:TQJa

[Sch52a] E. Schrödinger. Are there quantum jumps?: Part I. British Journal for the Philosophy of Science, III(10):109–123, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/109.full.
pdf+html;   http://www.
jstor.org/stable/685552.

Schrodinger:1952:TQJb

[Sch52b] E. Schrödinger. Are there quantum jumps?: Part II. British Journal for the Philosophy of Science, III(11):233–242, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/233.full.
pdf+html;   http://www.
jstor.org/stable/685266.

Schrodinger:1953:RQT

[Sch53]  E. Schrödinger.  Relativistic

quantum theory. British Journal for the Philosophy of Science, IV(16):328–329, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/328.full.pdf+html.

Schlegel:1954:AU

[Sch54a] Richard Schlegel. The age of the universe. British Journal for the Philosophy of Science, V(19):226–236, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/226.full.pdf+html;
http://www.jstor.org/
stable/685718.

Schrodinger:1954:RQT

[Sch54b]  E. Schrödinger.  Relativistic

quantum theory. British Journal for the Philosophy of Science, IV(16):328–329, February 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686081.

Scheffler:1957:EPA

[Sch57] Israel Scheffler. Explanation, prediction and abstraction. British Journal for the Philosophy of Science, VII(28):293–309, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/28/293.full.
pdf+html;   http://www.
jstor.org/stable/685109.

Scheffler:1959:TT

[Sch59a] Israel Scheffler. Thoughts on teleology. British Journal for the Philosophy of Science, IX(36): 265–284, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/265.full.pdf+html;
http://www.jstor.org/
stable/685187.

Schlesinger:1959:BRP

[Sch59b] G. Schlesinger. Book review:

P. W. Bridgman’s Operational Analysis: the Differential Aspect. British Journal for the Philosophy of Science, IX(36): 299–306, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/299.full.pdf+html;
http://www.jstor.org/
stable/685189.

Schlesinger:1960:PMA

[Sch60] G. Schlesinger. The principle of minimum assumption. British Journal for the Philosophy of Science, XI(41):55–59, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/55.full.pdf+
html;   http://www.jstor.
org/stable/685820. See reply

[Kap60].

Schlesinger:1961:PMR

[Sch61a] G. Schlesinger. The prejudice of micro-reduction. British Journal for the Philosophy of Science, XII(47):215–224, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/47/215.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685208.

Schlesinger:1961:SNP

[Sch61b] G. Schlesinger. A second note on the principle of minimum assumption. British Journal for the Philosophy of Science, XI(44): 328–329, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XI/44/328.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685133.

Schlick:1961:CCP

[Sch61c] M. Schlick. Causality in contemporary physics (I). British Journal for the Philosophy of Science, XII(47):177–193, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/47/177.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685206.

Schlick:1962:CCP

[Sch62] M. Schlick. Causality in contemporary physics (II). British Journal for the Philosophy of Science, XII(48):281–298, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XII/48/281.full.

pdf+html;   http://www.
jstor.org/stable/685397.

Schagrin:1964:AJI

[Sch64a] Morton L. Schagrin. An analytic justification of induction. British Journal for the Philosophy of Science, XIV(56):343– 344, February 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/56/343.full.pdf+html;
http://www.jstor.org/
stable/685638.

Schoenberg:1964:COP

[Sch64b] Judith Schoenberg. Confirmation by observation and the paradox of the ravens. British Journal for the Philosophy of Science, XV(59):200–212, November 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/200.full.pdf+
html;   http://www.jstor.
org/stable/686524.

Schleichert:1965:NPD

[Sch65a] Hubert Schleichert. A note on Professor Dingle’s paper on Relativity. British Journal for the Philosophy of Science, XV (60):331, February 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/60/331-a.full.pdf+html;

                                         http://www.jstor.org/
                                        stable/686541.

Schlesinger:1965:R

[Sch65b] G. Schlesinger. Reviews. British Journal for the Philosophy of Science, XV(60):352–354, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/60/352.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686548.

Schaffner:1969:WCM

[Sch69a] Kenneth F. Schaffner. The Watson–Crick model and reductionism. British Journal for the Philosophy of Science, 20(4): 325–348, December 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/4/325.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686260.

Schick:1969:R

[Sch69b] Frederic Schick. Reviews. British Journal for the Philosophy of Science, 20(3):269–272, October 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/3/269.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686423.

Schlesinger:1970:ICC

[Sch70]  G. Schlesinger.  On irrele-

vant criteria of confirmation.

British Journal for the Philosophy of Science, 21(3):282–287, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/282.full.pdf+
html;   http://www.jstor.
org/stable/686201.

Scheffler:1971:EDI

[Sch71] Israel Scheffler. Explanations, desires, and inscriptions. British Journal for the Philosophy of Science, 22(4):362–369, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/4/362.full.pdf+
html;   http://www.jstor.
org/stable/686318.

Schaffner:1974:EVL

[Sch74a] Kenneth F. Schaffner. Einstein versus Lorentz: Research programmes and the logic of comparative theory evaluation. British Journal for the Philosophy of Science, 25(1):45–78, March 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/1/45.full.pdf+
html;   http://links.jstor.
org/sici?sici=0007-0882%
28197403%2925%3A1%3C45%
3AEVLRPA%3E2.0.CO%3B2-B;
http://www.jstor.org/
stable/686502.

Scheibe:1974:PQL

[Sch74b] Erhard Scheibe. Popper and quantum logic. British Journal for the Philosophy of Science, 25(4):319–328, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/4/319.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686458.

Schlesinger:1974:UFC

[Sch74c] G. Schlesinger. The unpredictability of free choices. British Journal for the Philosophy of Science, 25(3):209–221, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/25/3/209.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686904.

Schmitt:1975:R

[Sch75] Charles B. Schmitt. Reviews. British Journal for the Philosophy of Science, 26(3):268– 269, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        26/3/268.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686285.

Schlesinger:1976:URC

[Sch76]  George Schlesinger.  Unpre-

dictability: a reply to Cargile and to Benditt and Ross. British Journal for the Philosophy of Science, 27(3):267–274, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/27/3/267.full.pdf+
html;   http://www.jstor.
org/stable/686125.

Schick:1979:SKU

[Sch79a] Frederic Schick. Self-knowledge, uncertainty, and choice. British Journal for the Philosophy of Science, 30(3):235–252, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/3/235.full.pdf+
html;   http://www.jstor.
org/stable/687138.

Schramm:1979:R

[Sch79b]  Alfred Schramm.  Reviews.

British Journal for the Philosophy of Science, 30(1):90–93, ???? 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/1/90.full.pdf+
html;   http://www.jstor.
org/stable/687420.

Schmitt:1980:R

[Sch80] C. B. Schmitt. Reviews. British Journal for the Philosophy of Science, 31(2):195–199, June 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/31/2/195.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687188.

Schirn:1981:R

[Sch81]  Matthias Schirn.  Reviews.

British Journal for the Philosophy of Science, 32(4):419– 425, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/4/419.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687314.

Schmitt:1982:R

[Sch82] C. B. Schmitt. Reviews. British Journal for the Philosophy of Science, 33(2):226–227, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/226.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687284.

Schmitt:1983:R

[Sch83] Charles B. Schmitt. Reviews. British Journal for the Philosophy of Science, 34(1):71, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/1/71.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686935.

Schaffer:1984:R

[Sch84a] Simon Schaffer. Reviews. British Journal for the Philosophy of Science, 35(2):191–193, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/191.full.pdf+
html;   http://www.jstor.
org/stable/686972.

Schlagel:1984:RRH

[Sch84b] Richard H. Schlagel. A reasonable reply to Hume’s scepticism. British Journal for the Philosophy of Science, 35(4): 359–374, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/359.full.pdf+html;
http://www.jstor.org/
stable/687339.

Schlesinger:1987:DVD

[Sch87] George N. Schlesinger. Is determinism a vacuous doctrine? British Journal for the Philosophy of Science, 38(3):339– 346, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/339.full.pdf+html;
http://www.jstor.org/
stable/687489.

Schlesinger:1990:LR

[Sch90] George N. Schlesinger. Location and range. British Journal for the Philosophy of Science, 41(2):245–260, June 1990.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        41/2/245.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687774.

Schmidt:1998:NPR

[Sch98]  Jan  Hendrik

Schmidt. Newcomb’s Paradox realized with backward causation. British Journal for the Philosophy of Science, 49(1):67–87, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/1/67.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688144.

Schulte:1999:MEE

[Sch99] O. Schulte. Means-ends epistemology. British Journal for the Philosophy of Science, 50 (1):1–31, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/1/1.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072205.

Schulte:2000:BRB

[Sch00a] O. Schulte. Book review: Elements of Scientific Inquiry. E. Martin, D. Osherson. British Journal for the Philosophy of Science, 51(2):347–352, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/2/347.full.pdf+
html;   http://www.jstor.
org/stable/3541811.

Schulte:2000:DWB

[Sch00b] O. Schulte. Discussion. what to believe and what to take seriously: a reply to David Chart concerning the riddle of induction. British Journal for the Philosophy of Science, 51 (1):151–153, March 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/1/151.full.pdf+html;
http://www.jstor.org/
stable/3541753.

Schulte:2000:ICL

[Sch00c] Oliver Schulte. Inferring conservation laws in particle physics: a case study in the problem of induction. British Journal for the Philosophy of Science, 51(4):771–806, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/771.full.pdf+html;
http://www.jstor.org/
stable/3541727.

Schaffer:2001:PC

[Sch01a] Jonathan Schaffer. Physical causation. British Journal for the Philosophy of Science, 52(4): 809–813, December 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/4/809.full.pdf+html;
http://www.jstor.org/
stable/3541921.

Schurz:2001:PRC

[Sch01b] Gerhard Schurz. Pietroski and Rey on Ceteris Paribus Laws. British Journal for the Philosophy of Science, 52(2):359–370, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
52/2/359.full.pdf+html;
http://www.jstor.org/
stable/3541901. See [PR95].

Schaffer:2003:PC

[Sch03] Jonathan Schaffer. Principled chances. British Journal for the Philosophy of Science, 54 (1):27–41, March 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/1/27.full.pdf+html;
http://www.jstor.org/
stable/3541878.

Schabas:2005:NOE

[Sch05] Margaret Schabas. The natural origins of economics. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-73569-9 (cloth). xi + 231 pp. LCCN HB74.S35 S33

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0622/2005014061-b.htm;

http://www.loc.gov/catdir/
enhancements/fy0622/
2005014061-d.htm;   http://
www.loc.gov/catdir/toc/
ecip0513/2005014061.htm.

Schaffer:2007:BRP

[Sch07a] Jonathan Schaffer. Book review: Phil Dowe and Paul Noordhof, Cause and Chance: Causation in an Indeterministic World. British Journal for the Philosophy of Science, 58(4): 869–874, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/4/869.full.pdf+html;
http://www.jstor.org/
stable/30115176.

Schaffer:2007:DC

[Sch07b] Jonathan Schaffer. Deterministic chance? British Journal for the Philosophy of Science, 58(2):113–140, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/2/113.full.pdf+html;
http://www.jstor.org/
stable/30115221.

Schrenk:2007:MCP

[Sch07c] Markus Andreas Schrenk. The metaphysics of Ceteris Paribus laws, volume 16 of Philosophische Analyse; Philosophical analysis; Bd. 16. Ontos, Frankfurt, Germany, 2007. ISBN 3-938793-42-2 (hardcover). 192 pp. LCCN K460 .S37 2007. EUR79.00.

Schaffer:2008:BRA

[Sch08a] Jonathan Schaffer. Book review: Andreas Hüttemann, What’s Wrong With Microphysicalism? British Journal for the Philosophy of Science, 59(2):253–257, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/2/253.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072282.

Schindler:2008:MTE

[Sch08b] Samuel Schindler. Model, theory, and evidence in the discovery of the DNA structure. British Journal for the Philosophy of Science, 59(4):619–658, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/4/619.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072305.

Schulz:2008:SFM

[Sch08c]  Armin W. Schulz.  Struc-

tural flaws: Massive modularity and the argument from design. British Journal for the Philosophy of Science, 59(4): 733–743, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/4/733.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072309.

Schurz:2009:WES

[Sch09] Gerhard Schurz. When empirical success implies theoretical reference: a structural correspondence theorem. British Journal for the Philosophy of Science, 60 (1):101–133, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/101.full.pdf+html;
http://www.jstor.org/
stable/25591990.

Schrenk:2011:MDS

[Sch11a] Markus Schrenk. Book review:

Mauro Dorato: The Software of the Universe: an Introduction to the History and Philosophy of the Laws of Nature. British Journal for the Philosophy of Science, 62(1): 225–232, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/225.full.pdf+html.

Schupbach:2011:NHS

[Sch11b]  Jonah N. Schupbach.  New

hope for Shogenji’s coherence measure. British Journal for the Philosophy of Science, 62 (1):125–142, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/125.full.pdf+html.

Schroeder:2013:RHH

[Sch13] S. Andrew Schroeder. Rethinking health: Healthy or healthier than? British Journal for the Philosophy of Science, 64 (1):131–159, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/1/131.full.pdf+html.

Schulz:2014:BRK

[Sch14] Armin W. Schulz. Book review: Kim Sterelny, Richard Joyce, Brett Calcott, And Ben Fraser (Eds), Cooperation And Its Evolution. British Journal for the Philosophy of Science, 65(4):893–897, December 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/4/893.

Schwarz:2015:BUA

[Sch15] Wolfgang Schwarz. Belief update across fission. British Journal for the Philosophy of Science, 66(3):659–682, September 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/3/659.

Schupbach:2018:RAE

[Sch18] Jonah N. Schupbach. Robustness analysis as explanatory reasoning. British Journal for the Philosophy of Science, 69(1):275– 300, March 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/275/2669768.

Schneider:2022:BFP

[Sch22a] Mike D. Schneider. Betting on future physics. British Journal for the Philosophy of Science, 73 (1):161–183, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Schoenfield:2022:AVG

[Sch22b]  Miriam Schoenfield.  Accu-

racy and verisimilitude: The good, the bad, and the ugly. British Journal for the Philosophy of Science, 73(2):373–406, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Sciama:1968:R

[Sci68] D. W. Sciama. Reviews. British Journal for the Philosophy of Science, 18(4):345–346, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/4/345.full.pdf+
html;   http://www.jstor.
org/stable/686367.

Sciama:1978:R

[Sci78] D. W. Sciama. Reviews. British Journal for the Philosophy of Science, 29(3):294–295, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/3/294.full.pdf+
html;   http://www.jstor.
org/stable/687022.

Scott:1961:MEI

[Sco61]  K. J. Scott.  Methodologi-

cal and epistemological individualism: Victoria University of Wellington. British Journal for the Philosophy of Science, XI(44):331–336, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/331.full.pdf+html;
http://www.jstor.org/
stable/685135.

Scott:2008:BRP

[Sco08] Michael Scott. Book review: Phil Dowe, Galileo, Darwin, and Hawking: the Interplay of Science, Reason, and Religion. British Journal for the Philosophy of Science, 59(3):575– 577, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/575.full.pdf+html;
http://www.jstor.org/
stable/40072301.

Scriven:1954:AU

[Scr54] Michael Scriven. The age of the universe. British Journal for the Philosophy of Science, V(19):181–190, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/181.full.pdf+html;
http://www.jstor.org/
stable/685714.

Scriven:1962:R

[Scr62]  Michael Scriven.  Reviews.

British Journal for the Philosophy of Science, XII(48):339– 342, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/48/339.full.pdf+html;
http://www.jstor.org/
stable/685402.

Scriven:1963:SL

[Scr63] Michael Scriven. The supercomputer as liar. British Journal for the Philosophy of Science, XIII(52):313–315, February 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/52/313.full.
pdf+html;   http://www.
jstor.org/stable/685329.

Stone:1993:CNP

[SD93] Tony Stone and Martin Davies.

Cognitive neuropsychology and the philosophy of mind. British Journal for the Philosophy of Science, 44(4):589–622, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/4/589.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/688033.

Stern:2023:ARI

[SE23] Reuben Stern and Benjamin Eva. Anti-reductionist interventionism. British Journal for the Philosophy of Science, 74(1): 241–267, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Seager:1987:CCE

[Sea87] William Seager. Credibility, confirmation and explanation. British Journal for the Philosophy of Science, 38(3):301– 317, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/3/301.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687487.

Sebens:2022:MGF

[Seb22] Charles T. Sebens. The mass of the gravitational field. British Journal for the Philosophy of Science, 73(1): 211–248, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Segal:1997:R

[Seg97] Gabriel Segal. Review. British Journal for the Philosophy of Science, 48(1):151–156, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/

content/48/1/151.full.pdf+
html;   http://www.jstor.
org/stable/687730.

Segall:2008:FSR

[Seg08] Robert Segall. Fertility and scientific realism. British Journal for the Philosophy of Science, 59(2):237–246, June 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/2/237.full.pdf+html;
http://www.jstor.org/
stable/40072280.

Self:1952:R

[Sel52] Henry Self. Reviews. British Journal for the Philosophy of Science, III(10):207–210, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/207.full.
pdf+html;   http://www.
jstor.org/stable/685565.

Sesardic:1999:RAA

[Ses99] N. Sesardic. Review article.

altruism. British Journal for the Philosophy of Science, 50(3): 457–466, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/457.full.pdf+html;
http://www.jstor.org/
stable/40072242.

Sesardic:2005:MSH

[Ses05] Neven Sesardić. Making sense of heritability. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-82818-X. xiii + 267 pp. LCCN BF341 .S47 2005. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        ecip057/2005002871.html.

Sesardic:2007:SID

[Ses07] Neven Sesardic. Sudden infant death or murder? A royal confusion about probabilities. British Journal for the Philosophy of Science, 58(2):299–329, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/58/2/299.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/30115228.

Settle:1969:PPE

[Set69] T. W. Settle. The point of positive evidence — reply to Professor Feyerabend. British Journal for the Philosophy of Science, 20(4):352–355, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/4/352.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686262.

Settle:1970:CPD

[Set70] T. W. Settle. Confirmation as a probability: Dead but it won’t lie down! British Journal for the Philosophy of Science, 21(2):200–201, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/2/200.full.pdf+html;
http://www.jstor.org/
stable/686702.

Settle:1972:PTP

[Set72] Tom Settle. Propensity theories of probability unscathed: a reply to White. British Journal for the Philosophy of Science, 23(4):331–335, November 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/331.full.pdf+html;
http://www.jstor.org/
stable/686733.

Settle:1977:PVP

[Set77a] Tom Settle. Popper versus Peirce on the probability of single cases. British Journal for the Philosophy of Science, 28(2):177–180, June 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/28/2/177.full.pdf+
html;   http://www.jstor.
org/stable/686633.

Settle:1977:R

[Set77b] Tom Settle. Reviews. British Journal for the Philosophy of Science, 28(1):86–94, March 1977. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/28/1/86.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686453.

Settle:1989:VRM

[Set89] Tom Settle. Van Rooijen and Mayr versus Popper: is the universe causally closed? British Journal for the Philosophy of Science, 40(3):389–403, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/40/3/389.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687786.

Settle:1990:SVP

[Set90] Tom Settle. Swann versus Popper on induction: an arbitration. British Journal for the Philosophy of Science, 41(3):401– 405, September 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        41/3/401.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687827.

Shrader-Frechette:2013:BRS

[SF13] Kristin Shrader-Frechette. Book review: Sandra D. Mitchell, Unsimple Truths: Science, Complexity, and Policy. British Journal for the Philosophy of Science, 64(2):449–453, June 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/449.full.pdf+html.

Sterelny:2017:EMR

[SF17] Kim Sterelny and Ben Fraser. Evolution and moral realism. British Journal for the Philosophy of Science, 68(4):981–1006, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/981/2669734.

Simon:1973:RET

[SG73] Herbert A. Simon and Guy J.

Groen. Ramsey eliminability and the testability of scientific theories. British Journal for the Philosophy of Science, 24(4): 367–380, December 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/4/367.full.pdf+html;
http://www.jstor.org/
stable/686617.

Spirtes:1997:RHF

[SGS97] Peter Spirtes, Clark Glymour, and Richard Scheines. Reply to Humphreys and Freedman’s review of causation, prediction, and search. British Journal for the Philosophy of Science, 48(4):555–568, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        48/4/555.full.pdf+html;
                                        http://www.
                                        jstor.org/stable/688100. See

[HF96, KW97].

Shea:2018:CSS

[SGSC18] Nicholas Shea, Peter Godfrey-Smith, and Rosa Cao. Content in simple signalling systems. British Journal for the Philosophy of Science, 69(4):1009–1035, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/4/1009/3884315.

Serban:2020:CRI

[SH20] Maria Serban and Sune Holm. Constitutive relevance in interlevel experiments. British Journal for the Philosophy of Science, 71(2):697–725, June 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/71/2/
                                        697/5069218.

Shackle:1959:TT

[Sha59] G. L. S. Shackle. Time and thought. British Journal for the Philosophy of Science, IX(36): 285–298, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/36/285.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685188.

Sharma:1982:RMP

[Sha82] C. S. Sharma. The role of mathematics in physics. British Journal for the Philosophy of Science, 33(3):275–286, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/3/275.full.pdf+
html;   http://www.jstor.
org/stable/687225.

Shanon:1984:MCP

[Sha84]  Benny Shanon.  Meno —

a cognitive psychological view. British Journal for the Philosophy of Science, 35(2):129–147, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/2/129.full.pdf+
html;   http://www.jstor.
org/stable/686963.

Shanon:1988:RMM

[Sha88] Benny Shanon. Remarks on the modularity of mind. British Journal for the Philosophy of Science, 39(3):331–352, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/331.full.pdf+
html;   http://www.jstor.
org/stable/687210.

Shapere:1995:R

[Sha95] Dudley Shapere. Review. British Journal for the Philosophy of Science, 46(1):139–141, March 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/1/139.full.pdf+
html;   http://www.jstor.
org/stable/687637.

Shapiro:1997:JR

[Sha97] Lawrence A. Shapiro. Junk representations. British Journal for the Philosophy of Science, 48(3):345–361, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/3/345.full.pdf+html;
http://www.jstor.org/
stable/688067.

Shapiro:2003:PFN

[Sha03] Stewart Shapiro. Prolegomenon to any future neo-logicist set theory: Abstraction and indefinite extensibility. British Journal for the Philosophy of Science, 54(1):59–91, March 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/1/59.full.pdf+html;
http://www.jstor.org/
stable/3541880.

Shackel:2005:FBD

[Sha05] Nicholas Shackel. The form of the Benardete dichotomy. British Journal for the Philosophy of Science, 56(2):397–417, June 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/2/397.full.pdf+
html;   http://www.jstor.
org/stable/3541709.

Shagrir:2012:SRB

[Sha12] Oron Shagrir. Structural representations and the brain. British Journal for the Philosophy of Science, 63(3):519–545, September 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
63/3/519.full.pdf+html.

Shapiro:2017:MBE

[Sha17] Lawrence A. Shapiro. Mechanism or bust? Explanation in psychology. British Journal for the Philosophy of Science, 68(4):1037–1059, December 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/4/1037/2669741.

Shepherdson:1964:R

[She64] J. C. Shepherdson. Reviews.

British Journal for the Philosophy of Science, XV(59):268– 270, November 1964. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XV/59/268.full.pdf+html;
http://www.jstor.org/
stable/686531.

Shea:1967:R

[She67]  William J. Shea.  Reviews.

British Journal for the Philosophy of Science, 17(4):344– 345, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        17/4/344.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686783.

Shea:1970:R

[She70] William R. Shea. Reviews. British Journal for the Philosophy of Science, 21(1):124– 125, February 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        21/1/124.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686479.

Shea:1971:R

[She71] W. R. Shea. Reviews. British Journal for the Philosophy of Science, 22(1):76–77, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/76.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686658.

Shelanski:1973:NTT

[She73] Vivien B. Shelanski. Nagel’s translation of teleological statements: a critique. British Journal for the Philosophy of Science, 24(4):397–401, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/4/397.full.pdf+html;
http://www.jstor.org/
stable/686620.

Sheldon:1985:OWT

[She85] Neil A. Sheldon. One wave or three? A problem for realism. British Journal for the Philosophy of Science, 36(4): 431–436, December 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/4/431.full.pdf+html;
http://www.jstor.org/
stable/687402.

Shenker:1999:KTR

[She99] O. R. Shenker. Is −kTr(ρ ln ρ)

the entropy in quantum mechanics. British Journal for the Philosophy of Science, 50(1):33–48, March 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/1/33.full.pdf+
html;   http://www.jstor.
org/stable/40072206.

Sher:2001:BRJ

[She01] Gila Sher. Book review: Johannes L. Brandl and Peter Sullivan (eds), New Essays on the Philosophy of Michael Dummett. British Journal for the Philosophy of Science, 52(1):185–189, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/52/1/185.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541955.

Shelley:2002:ACA

[She02] Cameron Shelley. Analogy counterarguments and the acceptability of analogical hypotheses. British Journal for the Philosophy of Science, 53(4): 477–496, December 2002. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        53/4/477.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541820.

Shea:2006:BRK

[She06] Nicholas Shea. Book review: Kim Sterelny: Thought in a Hostile World, Oxford: Blackwell, 2003 Paperback £17.99 ISBN 0-631-18887-8. British Journal for the Philosophy of Science, 57 (1):259–266, March 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        57/1/259.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541663.

Shea:2013:IRR

[She13] Nicholas Shea. Inherited representations are read in development. British Journal for the Philosophy of Science, 64 (1):1–31, March 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/1/1.full.pdf+html.

Shimony:1984:CHV

[Shi84] Abner Shimony. Contextual hidden variables theories and Bell’s Inequalities. British Journal for the Philosophy of Science, 35 (1):25–45, March 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/1/25.full.pdf+html;
http://www.jstor.org/
stable/687555.

Shomar:1998:R

[Sho98] Towfic Shomar. Review. British Journal for the Philosophy of Science, 49(1):185–188, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/1/185.full.pdf+
html;   http://www.jstor.
org/stable/688152.

Shogenji:2000:SDJ

[Sho00] T. Shogenji. Self-dependent justification without circularity. British Journal for the Philosophy of Science, 51(2):287–298, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/51/2/287.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541806.

Shogenji:2003:CTP

[Sho03] Tomoji Shogenji. A condition for transitivity in probabilistic support. British Journal for the Philosophy of Science, 54(4): 613–616, December 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/4/613.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541682.

Shogenji:2017:MC

[Sho17]  Tomoji Shogenji.  Mediated

confirmation. British Journal for the Philosophy of Science, 68(3):847–874, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/847/2669646.

Shrader:1985:R

[Shr85]  Douglas Shrader.  Reviews.

British Journal for the Philosophy of Science, 36(2):221–225, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/2/221.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687039.

Shrapnel:2019:DQC

[Shr19]  Sally Shrapnel.  Discovering

quantum causal models. British Journal for the Philosophy of Science, 70(1):1–25, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
1/4098117.

Shumener:2019:LNE

[Shu19] Erica Shumener. Laws of nature, explanation, and semantic circularity. British Journal for the Philosophy of Science, 70(3):787–815, September 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/3/
787/4682636.

Sider:2001:BRB

[Sid01] Theodore Sider. Book review:

Occasions of Identity, André Gallois. British Journal for the Philosophy of Science, 52 (2):401–405, June 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/401.full.pdf+html;
http://www.jstor.org/
stable/3541906.

Siegel:1980:ORI

[Sie80] Harvey Siegel. Objectivity, rationality, incommensurability, and more. British Journal for the Philosophy of Science, 31(4): 359–375, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        31/4/359.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687063.

Siegel:1997:R

[Sie97] Harvey Siegel. Review. British Journal for the Philosophy of Science, 48(1):121–125, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/1/121.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687724.

Siegel:2010:BRP

[Sie10a] Harvey Siegel. Book review: Penelope Maddy, Second Philosophy: a Naturalistic Method. British Journal for the Philosophy of Science, 61(4):897– 903, December 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/4/897.full.pdf+html.

Sieroka:2010:GVT

[Sie10b] Norman Sieroka. Geometrization versus transcendent matter: a systematic historiography of theories of matter following Weyl. British Journal for the Philosophy of Science, 61(4):769–802, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/4/769.full.pdf+html.

Silvers:1991:NSM

[Sil91] Stuart Silvers. On naturalizing the semantics of mental representation. British Journal for the Philosophy of Science, 42(1):49–73, March 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/1/49.full.pdf+html;
http://www.jstor.org/
stable/687606.

Simon:1965:LRD

[Sim65] Herbert A. Simon. The logic of rational decision. British Journal for the Philosophy of Science, XVI(63):169–186, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/63/169.full.
pdf+html;   http://www.
jstor.org/stable/686301.

Simon:1983:FRS

[Sim83] Herbert A. Simon. Fitness requirements for scientific theories. British Journal for the Philosophy of Science, 34(4): 355–365, December 1983. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/4/355.full.pdf+html;
http://www.jstor.org/
stable/687525.

Simon:1985:QTT

[Sim85] Herbert A. Simon. Quantification of theoretical terms and the falsifiability of theories. British Journal for the Philosophy of Science, 36(3):291–298, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/3/291.full.pdf+
html;   http://www.jstor.
org/stable/687572.

Simon:1991:BRW

[Sim91] Herbert A. Simon. Black ravens and a white shoe. British Journal for the Philosophy of Science, 42(3):339–342, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/3/339.full.pdf+html;
http://www.jstor.org/
stable/687912.

Simons:1997:R

[Sim97] Peter Simons. Review. British Journal for the Philosophy of Science, 48(3):439–443, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/3/439.full.pdf+
html;   http://www.jstor.
org/stable/688073.

Simons:2001:BRM

[Sim01] Peter Simons. Book review: Mark Steiner, The Applicability of Mathematics as a Philosophical Problem. British Journal for the Philosophy of Science, 52 (1):181–184, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/1/181.full.pdf+html;
http://www.jstor.org/
stable/3541954.

Singer:1970:R

[Sin70] B. R. Singer. Reviews. British Journal for the Philosophy of Science, 21(4):396–397, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/396.full.pdf+
html;   http://www.jstor.
org/stable/686566.

Singer:1971:R

[Sin71] B. R. Singer. Reviews. British Journal for the Philosophy of Science, 22(2):200–202, May 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/22/2/200.full.pdf+
html;   http://www.jstor.
org/stable/686490.

Singer:2000:BRB

[Sin00] P. Singer. Book review: Darwinian Dominion: Animal Welfare and Human Interests. L. Petrinovich. British Journal for the Philosophy of Science, 51(3):495–498, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        51/3/495.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541777.

Sizer:2000:TCT

[Siz00] Laura Sizer. Towards a computational theory of mood. British Journal for the Philosophy of Science, 51(4):743–770, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        51/4/743.full.pdf+html.

Sklar:1967:TIT

[Skl67] Lawrence Sklar. Types of intertheoretic reduction. British Journal for the Philosophy of Science, 18(2):109–124, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/2/109.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686579.

Sklar:1996:R

[Skl96] Lawrence Sklar. Review. British Journal for the Philosophy of Science, 47(4):630–632, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/4/630.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687930.

Sklar:2000:ITC

[Skl00] Lawrence Sklar. Interpreting theories: The case of statistical mechanics. British Journal for the Philosophy of Science, 51(4):729–742, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/3541615.

Sklar:2013:PFD

[Skl13]  Lawrence Sklar.  Philosophy

and the foundations of dynamics. Cambridge University Press, Cambridge, UK, 2013. ISBN 0-521-88819-0 (hardcover), 0-521-71630-6 (paperback). ix + 272 pp. LCCN Q175.32.D96 S55 2013.

Skolimowski:1972:R

[Sko72] Henryk Skolimowski. Reviews.

British Journal for the Philosophy of Science, 23(4):366– 369, November 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/4/366.full.pdf+html;
http://www.jstor.org/
stable/686740.

Skorupski:1982:R

[Sko82] John Skorupski. Reviews. British Journal for the Philosophy of Science, 33(1):104–111, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/1/104.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687247.

Skow:2007:SM

[Sko07] Bradford Skow. Sklar’s maneuver. British Journal for the Philosophy of Science, 58(4): 777–786, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/4/777.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115173.

Skow:2014:ECT

[Sko14] Bradford Skow. Editor’s choice: Are there non-causal explanations (of particular events)? British Journal for the Philosophy of Science, 65(3):445– 467, September 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/3/445.

Skow:2015:TGP

[Sko15] Bradford Skow. Are there genuine physical explanations of mathematical phenomena? British Journal for the Philosophy of Science, 66(1):69–93, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/66/1/69.

Skyrms:1984:PE

[Sky84] Brian Skyrms. Pragmatics and empiricism. Yale University Press, New Haven, CT, USA, 1984. ISBN 0-300-03174-2. xi + 143 pp. LCCN B816 .S59 1984.

Skyrms:1996:R

[Sky96] Brian Skyrms. Review. British Journal for the Philosophy of Science, 47(4):627–629, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/627.full.pdf+
html;   http://www.jstor.
org/stable/687929.

Skyrms:2010:SEL

[Sky10] Brian Skyrms. Signals: evolution, learning, and information. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-958294-7, 0-19-157455-4 (e-book). 199 pp. LCCN P99 .S59 2010eb.

Skyrms:2012:BRK

[Sky12] Brian Skyrms. Book review:

Ken Binmore, Rational Decisions. British Journal for the Philosophy of Science, 63 (2):449–453, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/449.full.pdf+html.

Sabbarton-Leary:2013:BRJ

[SL13] Nigel Sabbarton-Leary. Book review: Joseph Keim Campbell, Michael O’Rourke and Matthew H. Slater (eds), Carving Nature at Its Joints: Natural Kinds in Metaphysics and Science. British Journal for the Philosophy of Science, 64(4): 907–911, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/907.full.pdf+html.

Slater:1993:PFO

[Sla93]  B. H. Slater.  Probabilistic

foundations for operator logic. British Journal for the Philosophy of Science, 44(3):517– 530, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/517.full.pdf+html;
http://www.jstor.org/
stable/688020.

Slater:2013:SRE

[Sla13] Matthew H. Slater. Are species real?: an essay on the metaphysics of species. New directions in the philosophy of science. Palgrave Macmillan, New York, NY, USA, 2013. ISBN 0-230-39324-1 (hardcover). xi + 214 pp. LCCN QH83 .S55 2013.

Slater:2015:BRM

[Sla15a] Matthew H. Slater. Book review: Muhammad Ali Khalidi, Natural Categories and Human Kinds: Classification in the Natural and Social Sciences. British Journal for the Philosophy of Science, 66 (4):1017–1023, December 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Slater:2015:NK

[Sla15b] Matthew H. Slater. Natural kindness. British Journal for the Philosophy of Science, 66 (2):375–411, June 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/2/375.

Slezak:1982:GTM

[Sle82] Peter Slezak. Gödel’s theorem and the mind. British Journal for the Philosophy of Science, 33(1):41–52, March 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
33/1/41.full.pdf+html;
http://www.jstor.org/
stable/687239.

Slezak:1983:DDD

[Sle83] Peter Slezak. Descartes’s diagonal deduction. British Journal for the Philosophy of Science, 34(1):13–36, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/1/13.full.pdf+html;
http://www.jstor.org/
stable/686931.

Slezak:1988:WDL

[Sle88] Peter Slezak. Was Descartes a liar? diagonal doubt defended. British Journal for the Philosophy of Science, 39(3): 379–388, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/379.full.pdf+html;
http://www.jstor.org/
stable/687215.

Sloman:1966:R

[Slo66] Aaron Sloman. Reviews. British Journal for the Philosophy of Science, 17(3):249–253, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/3/249.full.pdf+
html;   http://www.jstor.
org/stable/686762.

Sloman:1968:R

[Slo68] A. Sloman. Reviews. British Journal for the Philosophy of Science, 19(2):171–173, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/2/171.full.pdf+
html;   http://www.jstor.
org/stable/686798.

Sloman:1970:R

[Slo70] Aaron Sloman. Reviews. British Journal for the Philosophy of Science, 21(2):208–211, May 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/2/208.full.pdf+

html;   http://www.jstor.
org/stable/686705.

Sloan:1985:R

[Slo85]  Phillip R. Sloan.  Reviews.

British Journal for the Philosophy of Science, 36(1):94–99, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/94.full.pdf+
html;   http://www.jstor.
org/stable/687163.

Sloep:1993:MR

[Slo93] Peter B. Sloep. Methodology revitalized? British Journal for the Philosophy of Science, 44(2):231–249, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/2/231.full.pdf+html;
http://www.jstor.org/
stable/687646.

Sluga:1967:R

[Slu67] H. D. Sluga. Reviews. British Journal for the Philosophy of Science, 17(4):339–341, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/4/339.full.pdf+
html;   http://www.jstor.
org/stable/686781.

Sluga:1980:R

[Slu80] Hans D. Sluga. Reviews. British Journal for the Philosophy of Science, 31(4):390–396, December 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/4/390.full.pdf+
html;   http://www.jstor.
org/stable/687066.

Sluga:1985:R

[Slu85] Hans Sluga. Reviews. British Journal for the Philosophy of Science, 36(4):469–473, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/4/469.full.pdf+
html;   http://www.jstor.
org/stable/687407.

Smart:1953:NC

[Sma53] J. J. C. Smart. A note on categories. British Journal for the Philosophy of Science, IV (15):227–228, November 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/227.full.pdf+html;
http://www.jstor.org/
stable/685899.

Smart:1970:R

[Sma70] Ninian Smart. Reviews. British Journal for the Philosophy of Science, 21(4):392–394, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/392.full.pdf+

html;   http://www.jstor.
org/stable/686564.

Smart:1972:SHM

[Sma72] J. J. C. Smart. Science, history and methodology. British Journal for the Philosophy of Science, 23(3):266–274, August 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/3/266.full.pdf+
html;   http://www.jstor.
org/stable/686690.

Smart:1979:PAP

[Sma79] J. J. C. Smart. A physicalist account of psychology. British Journal for the Philosophy of Science, 30(4):403–410, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/4/403.full.pdf+
html;   http://www.jstor.
org/stable/687548.

Small:1986:TSL

[Sma86] Robin Small. Tristram Shandy’s last page. British Journal for the Philosophy of Science, 37 (2):213–216, June 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686979.

Smart:2000:BRN

[Sma00] J. J. C. Smart. Book review: Nicholas Maxwell — The Comprehensibility of the Universe: A New Conception of Science. British Journal for the Philosophy of Science, 51(4): 907–911, December 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/907.full.pdf+html;
http://www.jstor.org/
stable/3541735.

Smead:2015:ECR

[Sme15] Rory Smead. Editor’s choice: The role of social interaction in the evolution of learning. British Journal for the Philosophy of Science, 66(1):161–180, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/1/161.

Smith:1959:PCL

[Smi59] F. V. Smith. Psychological concepts and linguistic restraints. British Journal for the Philosophy of Science, X(39):223– 227, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/223.full.pdf+html;
http://www.jstor.org/
stable/686055.

Smiley:1970:R

[Smi70] T. J. Smiley. Reviews. British Journal for the Philosophy of Science, 21(4):405–406, November 1970. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/405.full.pdf+
html;   http://www.jstor.
org/stable/686573.

Smith:1985:Rb

[Smi85a] John E. Smith. Reviews. British Journal for the Philosophy of Science, 36(4):461–465, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/4/461.full.pdf+
html;   http://www.jstor.
org/stable/687405.

Smith:1985:Ra

[Smi85b] Roger Smith. Reviews. British Journal for the Philosophy of Science, 36(3):347–348, September 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/3/347.full.pdf+
html;   http://www.jstor.
org/stable/687581.

Smith:1994:DBB

[Smi94]  Quentin  Smith.

Did the Big Bang have a cause? British Journal for the Philosophy of Science, 45(2):649–668, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/649.full.pdf+
html;   http://www.jstor.
org/stable/687688.

Smith:1997:BET

[Smi97] Nicholas J. J. Smith. Bananas enough for time travel? British Journal for the Philosophy of Science, 48(3):363–389, September 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/3/363.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688068.

Smith:1998:ATD

[Smi98]  Peter Smith.  Approximate

truth and dynamical theories. British Journal for the Philosophy of Science, 49(2):253–277, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/2/253.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688114.

Smith:2000:BRB

[Smi00a] C. Smith. Book review: The Natural Philosophy of James Clerk Maxwell. P. M. Harman. British Journal for the Philosophy of Science, 51(3):531–534, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/3/531.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541783.

Smith:2000:BRP

[Smi00b] Quentin Smith. Book review: Philip Turetzky — Time. British Journal for the Philosophy of Science, 51(4):927–933, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/4/927.full.pdf+
html;   http://www.jstor.
org/stable/3541739.

Smith:2007:CRC

[Smi07a]  Sheldon R. Smith.  Causa-

tion and its relation to ‘causal laws’. British Journal for the Philosophy of Science, 58(4): 659–688, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/4/659.full.pdf+html;
http://www.jstor.org/
stable/30115169.

Smith:2007:CBI

[Smi07b] Sheldon R. Smith. Continuous bodies, impenetrability, and contact interactions: the view from the applied mathematics of continuum mechanics. British Journal for the Philosophy of Science, 58(3):503–538, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/3/503.full.pdf+
html;   http://www.jstor.
org/stable/30115186.

Smith:2010:GLP

[Smi10] Martin Smith. A generalised lottery paradox for infinite probability spaces. British Journal for the Philosophy of Science, 61(4):821–831, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/4/821.full.pdf+html.

Smithson:2017:PII

[Smi17] Robert Smithson. The principle of indifference and inductive scepticism. British Journal for the Philosophy of Science, 68(1):68–??, March 2017. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

                                        URL https://academic.oup.
                                        com/bjps/article/68/1/253/
                                        2995349/The-Principle-of-
                                        Indifference-and-Inductive.

Saborido:2011:BOC

[SMM11] Cristian Saborido, Matteo Mossio, and Alvaro Moreno. Biological organization and cross-generation functions. British Journal for the Philosophy of Science, 62(3): 583–606, September 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/3/583.full.pdf+html.

Smokler:1977:SQC

[Smo77] Howard Smokler. Semantical questions in Carnap’s inductive logic. British Journal for the Philosophy of Science, 28 (2):129–135, June 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/2/129.full.pdf+html;
http://www.jstor.org/
stable/686629.

Smortchkova:2021:SGD

[Smo21] Joulia Smortchkova. Seeing goaldirectedness: a case for social perception. British Journal for the Philosophy of Science, 71(3): 855–879, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Smythies:1953:MP

[Smy53] J. R. Smythies. The mescaline phenomena. British Journal for the Philosophy of Science, III(12):339–347, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/339.full.pdf+html;
http://www.jstor.org/
stable/685448.

Smythies:1954:AP

[Smy54] J. R. Smythies. Analysis of projection. British Journal for the Philosophy of Science, V(18): 120–133, August 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/120.full.pdf+html;
http://www.jstor.org/
stable/685169.

Smythies:1956:SPE

[Smy56] J. R. Smythies. The stroboscope as providing empirical confirmation of the representative theory of perception. British Journal for the Philosophy of Science, VI(24):332–335, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/332.full.pdf+
html;   http://www.jstor.
org/stable/685413.

Smythies:1958:STI

[Smy58a] J. R. Smythies. On the space and time of images. British Journal for the Philosophy of Science, IX(33):40–42, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/33/40.full.pdf+html;
http://www.jstor.org/
stable/686100.

Smythies:1958:PSS

[Smy58b] J. R. Smythies. ‘Philosophical’ and ‘scientific’ sense-data. British Journal for the Philosophy of Science, IX(35):224– 225, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/224.full.pdf+html;
http://www.jstor.org/
stable/685652.

Smythies:1960:BC

[Smy60a] J. R. Smythies. Brain and consciousness. British Journal for the Philosophy of Science, X(40):341–344, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/341.full.pdf+html;
http://www.jstor.org/
stable/685923.

Smythies:1960:RPP

[Smy60b] J. R. Smythies. Reviews: the problems of perception. British Journal for the Philosophy of Science, XI(43):224–238, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/43/224.full.pdf+
html;   http://www.jstor.
org/stable/685312.

Smythies:1962:SSD

[Smy62] J. R. Smythies. On ‘space and sense-data’: a reply to Lord Brain. British Journal for the Philosophy of Science, XIII(50): 161–165, August 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/50/161.full.pdf+html;
http://www.jstor.org/
stable/685968.

Snider:1967:CCU

[Sni67] Caroline Whitbeck Snider. The confusion concerning universal forces. British Journal for the Philosophy of Science, 18(1):64– 66, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/64.full.pdf+
html;   http://www.jstor.
org/stable/686891.

Sober:2003:CAN

[SO03] Elliott Sober and Steven Hecht Orzack. Common ancestry and natural selection. British Journal for the Philosophy of Science, 54(3):423–437, September 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
54/3/423.full.pdf+html;
http://www.jstor.org/
stable/3541793.

Strand:2019:RCR

[SO19] Anders Strand and Gry Oftedal. Restricted causal relevance. British Journal for the Philosophy of Science, 70 (2):431–457, June 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/2/
431/4101971.

Soal:1954:R

[Soa54] S. G. Soal. Reviews. British Journal for the Philosophy of Science, V(18):164–166, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/164.full.pdf+
html;   http://www.jstor.
org/stable/685177.

Sober:1981:PP

[Sob81] Elliott Sober. The principle of parsimony. British Journal for the Philosophy of Science, 32(2):145–156, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/2/145.full.pdf+html;
http://www.jstor.org/
stable/687195.

Sober:1985:CEP

[Sob85] Elliott Sober. Constructive empiricism and the problem of aboutness. British Journal for the Philosophy of Science, 36 (1):11–18, March 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/1/11.full.pdf+html;
http://www.jstor.org/
stable/687153.

Sober:1987:EC

[Sob87] Elliott Sober. Explanation and causation. British Journal for the Philosophy of Science, 38 (2):243–257, June 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/2/243.full.pdf+html;

http://www.jstor.org/
stable/687050.

Sober:1995:NSD

[Sob95] Elliott Sober. Natural selection and distributive explanation: a reply to Neander. British Journal for the Philosophy of Science, 46(3):384–397, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/3/384.full.pdf+html;
http://www.jstor.org/
stable/687665.

Sober:1998:BBI

[Sob98] Elliott Sober. Black box inference: When should intervening variables be postulated? British Journal for the Philosophy of Science, 49(3):469–498, September 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/3/469.full.pdf+
html;   http://www.jstor.
org/stable/688085.

Sober:2001:VSL

[Sob01] Elliott Sober. Venetian sea levels, British bread prices, and the principle of the common cause. British Journal for the Philosophy of Science, 52(2):331–346, June 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/52/2/331.full.pdf+
html;   http://www.jstor.
org/stable/3541899.

Sober:2011:DDW

[Sob11] Elliott Sober. Did Darwin write The Origin backwards?: philosophical essays on Darwin’s theory. Prometheus Books, Amherst, NY, USA, 2011. ISBN 1-61614-230-8. 230 pp. LCCN QH365.O8 S63

2011. URL http://www.loc.
gov/catdir/enhancements/
fy1307/2010030409-b.html;
http://www.loc.gov/catdir/
enhancements/fy1307/
2010030409-d.html.

Solomon:1975:DAC

[Sol75] Robert C. Solomon. Doubts about the correlation thesis. British Journal for the Philosophy of Science, 26(1):27–39, March 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/1/27.full.pdf+
html;   http://www.jstor.
org/stable/686330.

Sorensen:1984:PEE

[Sor84] Roy A. Sorensen. Problems with electoral evaluations of expert opinions. British Journal for the Philosophy of Science, 35(1):47–53, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/1/47.full.pdf+html;
http://www.jstor.org/
stable/687556.

Sorensen:1986:WDC

[Sor86]  Roy  A.  Sorensen.

Was Descartes’s Cogito a diagonal deduction? British Journal for the Philosophy of Science, 37(3):346–351, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/687292.

Sorensen:2003:PR

[Sor03] Roy Sorensen. Para-reflections. British Journal for the Philosophy of Science, 54(1):93–101, March 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/54/1/93.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541881.

Sorensen:2007:LEC

[Sor07] Roy A. Sorensen. Logically equivalent — but closer to the truth. British Journal for the Philosophy of Science, 58 (2):287–297, June 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        58/2/287.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/30115227.

Steglich-Petersen:2012:ACA

[SP12] Asbjørn Steglich-Petersen. Against the contrastive account of singular causation. British Journal for the Philosophy of Science, 63 (1):115–143, March 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/1/115.full.pdf+html.

Spector:1965:MT

[Spe65] Marshall Spector. Models and theories. British Journal for the Philosophy of Science, XVI (62):121–142, August 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/62/121.full.pdf+html;
http://www.jstor.org/
stable/686153.

Spector:1966:TO

[Spe66a] Marshall Spector. Theory and observation (I). British Journal for the Philosophy of Science, 17(1):1–20, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686401.

Spector:1966:TOI

[Spe66b] Marshall Spector. Theory and observation (II). British Journal for the Philosophy of Science, 17(2):89–104, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/2/89.full.pdf+html;
http://www.jstor.org/
stable/686342.

Spilsbury:1953:Ra

[Spi53a] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, III(12):379–381, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/12/379.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685461.

Spilsbury:1953:Rb

[Spi53b] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, III(12):381–382, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/12/381.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685462.

Spilsbury:1953:Rc

[Spi53c] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, III(12):389–390, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/12/389.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685467.

Spilsbury:1955:Ra

[Spi55a] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, V(20):356–357, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/356.full.pdf+
html;   http://www.jstor.
org/stable/685750.

Spilsbury:1955:Rb

[Spi55b]  R. J. Spilsbury.  Reviews.

British Journal for the Philosophy of Science, VI(22):173, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/22/173.full.pdf+
html;   http://www.jstor.
org/stable/685529.

Spilsbury:1956:R

[Spi56]  R. J. Spilsbury.  Reviews.

British Journal for the Philosophy of Science, VII(27):268– 269, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/268.full.pdf+html;
http://www.jstor.org/
stable/685888.

Spilsbury:1957:EWP

[Spi57] R. J. Spilsbury. Experimentation within the psycho-analytic session. British Journal for the Philosophy of Science, VII(28): 338–342, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/338.full.pdf+html;
http://www.jstor.

                                        org/stable/685116. See reply

[Ezr57].

Spilsbury:1960:RUb

[Spi60a] R. J. Spilsbury. Review: [untitled]. British Journal for the Philosophy of Science, XI(41):86, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

                                        jstor.org/stable/685834.

Spilsbury:1960:RUc

[Spi60b] R. J. Spilsbury. Review: [untitled]. British Journal for the Philosophy of Science, XI (42):169, August 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685600.

Spilsbury:1960:Ra

[Spi60c]  R. J. Spilsbury.  Reviews.

British Journal for the Philosophy of Science, XI(41):85–86, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/41/85.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685833.

Spilsbury:1960:Rb

[Spi60d] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, XI(41):86a–86, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
XI/41/86-a.full.pdf+html.

Spilsbury:1960:Rc

[Spi60e] R. J. Spilsbury. Reviews. British Journal for the Philosophy of Science, XI(42):169a–169, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
XI/42/169-a.full.pdf+html.

Spielman:1973:RNP

[Spi73] Stephen Spielman. A refutation of the Neyman–Pearson theory of testing. British Journal for the Philosophy of Science, 24(3):201–222, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/3/201.full.pdf+html;
http://www.jstor.org/
stable/686228.

Spielman:1974:IGR

[Spi74] Stephen Spielman. On the infirmities of Gillies’s Rule. British Journal for the Philosophy of Science, 25(3):261–265, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/3/261.full.pdf+
html;   http://www.jstor.
org/stable/686907.

Spohn:2006:CA

[Spo06] Wolfgang Spohn. Causation: an alternative. British Journal for the Philosophy of Science, 57(1):93–119, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/93.full.pdf+html;
http://www.jstor.org/
stable/3541654.

Sprott:1953:R

[Spr53] W. J. H. Sprott. Reviews. British Journal for the Philosophy of Science, IV(15):246– 248, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/246.full.pdf+html;
http://www.jstor.org/
stable/685902.

Sprevak:2007:CRP

[Spr07] Mark D. Sprevak. Chinese rooms and program portability. British Journal for the Philosophy of Science, 58(4):755–776, December 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/4/755.full.pdf+
html;   http://www.jstor.
org/stable/30115172.

Sprevak:2011:BRW

[Spr11] Mark Sprevak. Book review: William M. Ramsey, Representation Reconsidered. British Journal for the Philosophy of Science, 62(3):669–675, September 2011.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/3/669.full.pdf+html.

Sprenger:2018:TIR

[Spr18] Jan Sprenger. Two impossibility results for measures of corroboration. British Journal for the Philosophy of Science, 69(1): 139–159, March 1, 2018. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/1/139/2669784.

Spurway:1952:BMC

[Spu52] H. Spurway. Behold, my child, the Nordic dog. British Journal for the Philosophy of Science, III(11):265–272, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/265.full.
pdf+html;   http://www.
jstor.org/stable/685271.

Squires:1989:CMR

[Squ89] Euan J. Squires. A comment on Maxwell’s resolution of the wave/particle dilemma. British Journal for the Philosophy of Science, 40(3):413–417, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/3/413.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/687789.

Squires:1996:R

[Squ96] Euan Squires. Review. British Journal for the Philosophy of Science, 47(1):152–153, March 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/1/152.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/1215837.

Stopes-Roe:1963:R

[SR63] H. V. Stopes-Roe. Reviews. British Journal for the Philosophy of Science, XIV(54): 169–172, August 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/54/169.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685438.

Stopes-Roe:1964:R

[SR64] H. V. Stopes-Roe. Reviews. British Journal for the Philosophy of Science, XV(58):158–161, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/58/158.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685258.

Scheibe:1965:TTC

[SS65] Karl E. Scheibe and Theodore R. Sarbin. Towards a theoretical conceptualisation of superstition. British Journal for the Philosophy of Science, XVI(62): 143–158, August 1965. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/62/143.full.pdf+html;
http://www.jstor.org/
stable/686154.

Shen:1993:FRS

[SS93] Wei-Min Shen and Herbert A. Simon. Fitness requirements for scientific theories containing recursive theoretical terms. British Journal for the Philosophy of Science, 44(4):641–652, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/4/641.full.pdf+
html;   http://www.jstor.
org/stable/688035.

Sober:2013:SCI

[SS13] Elliott Sober and Mike Steel.

Screening-off and causal incompleteness: a no-go theorem. British Journal for the Philosophy of Science, 64(3):513– 550, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/513.full.pdf+html.

Sober:2017:SEC

[SS17] Elliott Sober and Mike Steel.

Similarities as evidence for common ancestry: a likelihood epistemology. British Journal for the Philosophy of Science, 68(3):617–638, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/617/3044199.

Schippers:2020:GCT

[SS20] Michael Schippers and Gerhard Schurz. Genuine confirmation and tacking by conjunction. British Journal for the Philosophy of Science, 71(1):321–352, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /academic.oup.com/bjps/
                                        article/71/1/321/5045847.

Steeger:2022:TFI

[ST22] Jeremy Steeger and Nicholas Teh. Two forms of inconsistency in quantum foundations. British Journal for the Philosophy of Science, 72(4):1083–1110, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Stalnaker:1967:R

[Sta67] Robert C. Stalnaker. Reviews. British Journal for the Philosophy of Science, 18(1):72–74, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/18/1/72.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686894.

Stairs:1985:BQL

[Sta85] Allen Stairs. Bub on quantum logic and continuous geometry. British Journal for the Philosophy of Science, 36(3): 313–325, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/313.full.pdf+html;
http://www.jstor.org/
stable/687576.

Stapp:1990:CNI

[Sta90] Henry P. Stapp. Comments on ‘nonlocal influences and possible worlds’. British Journal for the Philosophy of Science, 41(1):59–72, March 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/1/59.full.pdf+html;
http://www.jstor.org/
stable/688003.

Stanford:2006:DPP

[Sta06a] P. Kyle Stanford. Darwin’s pangenesis and the problem of unconceived alternatives. British Journal for the Philosophy of Science, 57(1):121–144, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/1/121.full.pdf+
html;   http://www.jstor.
org/stable/3541655.

Stanford:2006:EOG

[Sta06b] P. Kyle Stanford. Exceeding our grasp: science, history, and the problem of unconceived alternatives. Oxford University Press, 2006. ISBN 0-19-517408-9 (hardcover). xiv + 234 pp. LCCN Q175.32.I54 S73

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0637/2005051299-d.html.

Stanford:2009:SRA

[Sta09] P. Kyle Stanford. Scientific realism, the atomic theory, and the catch-all hypothesis: Can we test fundamental theories against all serious alternatives? British Journal for the Philosophy of Science, 60(2):253–269, June 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/60/2/253.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/25592001.

Statham:2017:CCC

[Sta17] Georgie Statham. Contrastive causal claims: a case study. British Journal for the Philosophy of Science, 68(3):663–688, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/663/2669724.

Staley:2020:SEV

[Sta20] Kent W. Staley. Securing the empirical value of measurement results. British Journal for the Philosophy of Science, 71 (1):87–113, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
87/4657126.

Steen:1954:R

[Ste54] S. W. P. Steen. Reviews. British Journal for the Philosophy of Science, V(17):80–83, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/17/80.full.pdf+html;
http://www.jstor.org/
stable/685957.

Steen:1962:R

[Ste62] S. W. P. Steen. Reviews. British Journal for the Philosophy of Science, XII(48):332–338, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/48/332.full.
pdf+html;   http://www.
jstor.org/stable/685401.

Stevens:1970:R

[Ste70] Paul Stevens. Reviews. British Journal for the Philosophy of Science, 21(4):387–388, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/4/387.full.pdf+
html;   http://www.jstor.
org/stable/686561.

Steffens:1971:R

[Ste71]  Henry John Steffens.  Re-

views. British Journal for the Philosophy of Science, 22(1): 55–57, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        22/1/55.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686645.

Stein:1974:MCG

[Ste74] Howard Stein. Maurice Clavelin on Galileo’s natural philosophy. British Journal for the Philosophy of Science, 25(4):375– 397, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        25/4/375.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686467.

Stevenson:1978:R

[Ste78]  Leslie Stevenson.  Reviews.

British Journal for the Philosophy of Science, 29(4):398– 399, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/4/398.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687107.

Stevenson:1984:R

[Ste84]  Leslie Stevenson.  Reviews.

British Journal for the Philosophy of Science, 35(1):83–85, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/83.full.pdf+
html;   http://www.jstor.
org/stable/687563.

Sterelny:1995:ULR

[Ste95] Kim Sterelny. Understanding life: Recent work in philosophy of biology. British Journal for the Philosophy of Science, 46(2):155–183, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/2/155.full.pdf+html;
http://www.jstor.org/
stable/687810.

Stephens:1996:MRA

[Ste96a] Christopher Stephens. Modelling reciprocal altruism. British Journal for the Philosophy of Science, 47(4):533–551, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/4/533.full.pdf+html;
http://www.jstor.org/
stable/687923.

Sterelny:1996:R

[Ste96b] Kim Sterelny. Review. British Journal for the Philosophy of Science, 47(4):640–646, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/4/640.full.pdf+
html;   http://www.jstor.
org/stable/687933.

Sterelny:1997:UB

[Ste97] Kim Sterelny. Universal biology. British Journal for the Philosophy of Science, 48(4): 587–601, December 1997. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/4/587.full.pdf+html;
http://www.jstor.org/
stable/688102.

Stevenson:2000:BRB

[Ste00] L. Stevenson. Book review: Possible Experience: Understanding Kant’s Critique of Pure Reason. A. W. Collins. British Journal for the Philosophy of Science, 51(3):535–538, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/3/535.full.pdf+html;
http://www.jstor.org/
stable/3541784.

Stevenson:2001:HFA

[Ste01] Leslie Stevenson. Human freedom after Darwin: a critical rationalist view. British Journal for the Philosophy of Science, 52(4):795–799, December 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/4/795.full.pdf+html;

http://www.jstor.org/
stable/3541919.

Steel:2003:MTS

[Ste03a] Daniel Steel. Making time stand still: a response to Sober’s counter-example to the principle of the common cause. British Journal for the Philosophy of Science, 54(2):309–317, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/309.full.pdf+
html;   http://www.jstor.
org/stable/3541970.

Sterelny:2003:THW

[Ste03b] Kim Sterelny. Thought in a hostile world: the evolution of human cognition. Blackwell, Oxford, UK, 2003. ISBN 0-631-18886-X, 0-631-18887-8 (paperback). xi + 262 pp. LCCN BF698.95.S74 2003.

Steel:2005:ICM

[Ste05] Daniel Steel. Indeterminism and the causal Markov condition. British Journal for the Philosophy of Science, 56(1):3–26, March 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/1/3.full.pdf+
html;   http://www.jstor.
org/stable/3541848.

Steel:2006:CHW

[Ste06a] Daniel Steel. Comment on Hausman & Woodward on the causal Markov condition. British Journal for the Philosophy of Science, 57(1):219–231, March 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/219.full.pdf+html;
http://www.jstor.org/
stable/3541659. See [HW04].

Sterelny:2006:MR

[Ste06b] Kim Sterelny. Memes revisited. British Journal for the Philosophy of Science, 57(1):145–165, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/1/145.full.pdf+
html;   http://www.jstor.
org/stable/3541656.

Stegmann:2009:DII

[Ste09] Ulrich E. Stegmann. Dna, inference, and information. British Journal for the Philosophy of Science, 60(1):1–17, March 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/1.full.pdf+html;
http://www.jstor.org/
stable/25591986.

Stegmann:2010:BRB

[Ste10] Ulrich Stegmann. Book review: Brian Garvey, Philosophy of Biology. British Journal for the Philosophy of Science, 61 (1):235–236, March 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/1/235.full.pdf+html.

Steel:2012:BRF

[Ste12a] Daniel Steel. Book review:

Federica Russo, Causality and Causal Modelling in the Social Sciences: Measuring Variations. British Journal for the Philosophy of Science, 63(3): 725–728, September 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/3/725.full.pdf+html.

Sterelny:2012:EAH

[Ste12b] Kim Sterelny. The evolved apprentice: how evolution made humans unique. Jean Nicod lectures. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01679-6 (hardcover), 0-262-30281-0 (e-book). xvi + 242 pp. LCCN BF698.95 .S735 2012.

Stegenga:2016:PPN

[Ste16a] Jacob Stegenga. Population pluralism and natural selection. British Journal for the Philosophy of Science, 67(1):1–29, March 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/1/1.

Stegmann:2016:GCR

[Ste16b] Ulrich E. Stegmann. ‘Genetic coding’ reconsidered: an analysis of actual usage. British Journal for the Philosophy of Science, 67(3):707–730, September 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/3/707.

Sterelny:2016:ECC

[Ste16c] Kim Sterelny. Editor’s choice: Cooperation, culture, and conflict. British Journal for the Philosophy of Science, 67(1):31–58, March 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/67/1/31.

Stevens:2020:RRC

[Ste20] Syman Stevens. Regularity relationalism and the constructivist project. British Journal for the Philosophy of Science, 71 (1):353–372, March 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/71/1/
                                        353/4096302.

Still:1966:RU

[Sti66] A. W. Still. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):85–86, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686411.

Stich:1979:BCR

[Sti79]  Stephen P. Stich.  Between

Chomskian rationalism and Popperian empiricism. British Journal for the Philosophy of Science, 30(4):329–347, December 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/4/329.full.pdf+html;
http://www.jstor.org/
stable/687543.

Stich:1981:CPL

[Sti81] Stephen P. Stich. Can Popperians learn to talk? British Journal for the Philosophy of Science, 32(2):157–164, June 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/2/157.full.pdf+html;
http://www.jstor.org/
stable/687196.

Stich:1983:FPC

[Sti83] Stephen P. Stich. From folk psychology to cognitive science: the case against belief. MIT Press, Cambridge, MA, USA, 1983. ISBN 0-262-19215-2. xii + 266 pp. LCCN BF773 .S75 1983.

Stillwell:1985:CPP

[Sti85] Shelley Stillwell. Confirmation, paradoxes, and possible worlds. British Journal for the Philosophy of Science, 36(1):19–52, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/19.full.pdf+
html;   http://www.jstor.
org/stable/687154.

Stich:1994:VCI

[Sti94] Stephen P. Stich. The virtues, challenges and implications of connectionism. British Journal for the Philosophy of Science, 45(4):1047–1058, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/4/1047.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687621.

Stone:1964:PEE

[Sto64] Richard Stone. The a priori and the empirical in economics. British Journal for the Philosophy of Science, XV(58):115–122, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/58/115.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685250.

Stove:1966:LDC

[Sto66] D. Stove. On logical definitions of confirmation. British Journal for the Philosophy of Science, XVI(64):265–272, February 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/64/265.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686828.

Stoothoff:1970:R

[Sto70a] Robert H. Stoothoff. Reviews. British Journal for the Philosophy of Science, 21(4):390– 392, November 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/390.full.pdf+html;
http://www.jstor.org/
stable/686563.

Stove:1970:MGR

[Sto70b] D. Stove. Mr Gibson on ravens and relevance. British Journal for the Philosophy of Science, 21 (3):287–288, August 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/3/287.full.pdf+html;
http://www.jstor.org/
stable/686202.

Stoothoff:1979:R

[Sto79]  R. H. Stoothoff.  Reviews.

British Journal for the Philosophy of Science, 30(2):202–204, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/202.full.pdf+
html;   http://www.jstor.
org/stable/686956.

Stolakis:1986:ACP

[Sto86] George Stolakis. Against conventionalism in physics: Absolute synchronisation in a single frame of reference. British Journal for the Philosophy of Science, 37(2):229–232, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686983.

Stone:1991:KMF

[Sto91]  Mark  A.  Stone.

A Kuhnian model of falsifiability. British Journal for the Philosophy of Science, 42(2):177–185, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/42/2/177.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688051.

Stone:1999:BRB

[Sto99] T. Stone. Book review: Interpreting minds: the evolution of a practice. Radu J. Bogdan. British Journal for the Philosophy of Science, 50(3):492– 496, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        50/3/492.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072248.

Strevens:1995:CLN

[Str95] Michael Strevens. A closer look at the ‘new’ principle. British Journal for the Philosophy of Science, 46(4):545–561, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/4/545.full.pdf+

html;   http://www.jstor.
org/stable/687897.

Strevens:1996:QMF

[Str96] Michael Strevens. Quantum mechanics and frequentism: a reply to Ismael. British Journal for the Philosophy of Science, 47(4):575–577, December 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/4/575.full.pdf+html;
http://www.jstor.org/
stable/687925.

Strevens:2001:BTA

[Str01] Michael Strevens. The Bayesian treatment of auxiliary hypotheses. British Journal for the Philosophy of Science, 52(3):515– 537, September 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/3/515.full.pdf+html;
http://www.jstor.org/
stable/3541927.

Strevens:2003:BTC

[Str03] Michael Strevens. Bigger than chaos: understanding complexity through probability. Harvard University Press, Cambridge, MA, USA, 2003. ISBN 0-674-01042-6. xii + 413 pp. LCCN QC174.85.P76 S77 2003.

Strevens:2005:BTA

[Str05] Michael Strevens. The Bayesian treatment of auxiliary hypotheses: Reply to Fitelson and Waterman. British Journal for the Philosophy of Science, 56(4): 913–918, December 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/913.full.pdf+html;
http://www.jstor.org/
stable/3541872.

Strevens:2009:BRC

[Str09] Michael Strevens. Book review: C. S. Bertuglia and F. Vaio, Nonlinearity, Chaos, and Complexity. British Journal for the Philosophy of Science, 60 (2):447–451, June 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/2/447.full.pdf+html;
http://www.jstor.org/
stable/25592009.

Stuart:1961:SD

[Stu61] Alan L. Stuart. The sciences and deity. British Journal for the Philosophy of Science, XII (47):235–245, November 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/47/235.full.pdf+html;
http://www.jstor.org/
stable/685210.

Stuart:1963:EM

[Stu63] Alan L. Stuart. Evolutionary man. British Journal for the Philosophy of Science, XIV(53):41– 53, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/41.full.pdf+
html;   http://www.jstor.
org/stable/685352.

Studd:2016:AR

[Stu16] J. P. Studd. Abstraction reconceived. British Journal for the Philosophy of Science, 67 (2):579–615, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/579.

Styer:2000:SWQ

[Sty00] Daniel F. Styer. The strange world of quantum mechanics. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66104-8 (hardcover), 0-521-66780-1 (paperback). xiv + 154 pp. LCCN QC174.12 .ST 2000.

Suarez:1996:R

[Suá96] Mauricio Suárez. Review. British Journal for the Philosophy of Science, 47(3):483–486, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/3/483.full.pdf+
html;   http://www.jstor.
org/stable/687994.

Suarez:2004:QSP

[Suá04] Mauricio Suárez. Quantum selections, propensities and the prob-

lem  of  measurement.

British Journal for the Philosophy of Science, 55(2):219–255, June 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/55/2/219.full.pdf+
html;   http://www.jstor.
org/stable/3541689.

Suarez:2018:CP

[Suá18] Mauricio Suárez. The chances of propensities. British Journal for the Philosophy of Science, 69(4):1155–1177, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/1155/4061131.

Suchting:1967:R

[Suc67] W. A. Suchting. Reviews. British Journal for the Philosophy of Science, 18(1):81–83, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/1/81.full.pdf+html;
http://www.jstor.org/
stable/686900.

Suchting:1969:PRD

[Suc69] W. A. Suchting. Popper’s revised definition of natural necessity. British Journal for the Philosophy of Science, 20(4): 349–352, December 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
20/4/349.full.pdf+html;
http://www.jstor.org/
stable/686261.

Suchting:1970:R

[Suc70] W. A. Suchting. Reviews. British Journal for the Philosophy of Science, 21(1):121–124, February 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/1/121.full.pdf+
html;   http://www.jstor.
org/stable/686478.

Sudbury:1973:CTE

[Sud73] A. W. Sudbury. Could there exist a world which obeyed no scientific laws? British Journal for the Philosophy of Science, 24(1):39–40, March 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/1/39.full.pdf+html;
http://www.jstor.org/
stable/686668.

Sudbury:1976:SLN

[Sud76] Aidan Sudbury. Scientific laws that are neither deterministic nor probabilistic. British Journal for the Philosophy of Science, 27(4):307–316, December 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/307.full.pdf+html;
http://www.jstor.org/
stable/686860.

Sullivan:2022:UML

[Sul22] Emily Sullivan. Understanding from machine learning models. British Journal for the Philosophy of Science, 73(1):109– 133, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Summerfield:1953:R

[Sum53]  A.  Sum-

merfield. Reviews. British Journal for the Philosophy of Science, III(12):375–377, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        III/12/375.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685459.

Suppes:1998:R

[Sup98] Patrick Suppes. Review. British Journal for the Philosophy of Science, 49(2):351–354, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/49/2/351.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/688123.

Surez:2000:DQC

[Súr00] Mauricio Súrez. Dickson on quantum chance and non-locality. British Journal for the Philosophy of Science, 51(4):882–892, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/4/882.full.pdf+
html;   http://www.jstor.
org/stable/3541732.

Sustar:2007:CNG

[Šus07] Predrag Šustar. Crick’s notion of genetic information and the ‘central dogma’ of molecular biology. British Journal for the Philosophy of Science, 58(1):13–24, March 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/1/13.full.pdf+
html;   http://www.jstor.
org/stable/4489098.

Sutherland:1957:SCF

[Sut57] J. D. Sutherland. Some comments on Fairbairn’s paper. British Journal for the Philosophy of Science, VII(28):329–333, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/28/329.full.
pdf+html;   http://www.
jstor.org/stable/685114.

Suzuki:1985:TP

[Suz85] Yoshindo Suzuki. The transfer principle. British Journal for the Philosophy of Science, 36(1):61–66, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/1/61.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687156.

Saatsi:2011:MSI

[SV11] Juha Saatsi and Peter Vickers. Miraculous success? Inconsistency and untruth in Kirchhoff’s diffraction theory. British Journal for the Philosophy of Science, 62(1):29–46, March 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/1/29.full.pdf+html.

Saunders:2008:BU

[SW08a] Simon Saunders and David Wallace. Branching and uncertainty. British Journal for the Philosophy of Science, 59(3): 293–305, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/3/293.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072286.

Saunders:2008:SWR

[SW08b] Simon Saunders and David Wallace. Saunders and Wallace reply. British Journal for the Philosophy of Science, 59(3):315– 317, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/3/315.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072288.

Steele:2013:ECC

[SW13] Katie Steele and Charlotte Werndl. Editor’s choice: Climate models, calibration, and confirmation. British Journal for the Philosophy of Science, 64(3):609–635, September 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/3/609.full.pdf+html.

Steele:2018:MST

[SW18]  Katie  Steele

and Charlotte Werndl. Modelselection theory: The need for a more nuanced picture of use-novelty and double-counting. British Journal for the Philosophy of Science, 69(2):351–375, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/351/2669792.

Schiemer:2019:STR

[SW19a] Georg Schiemer and John Wigglesworth. The structuralist thesis reconsidered. British Journal for the Philosophy of Science, 70 (4):1201–1226, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
1201/4830144.

Schonherr:2019:BIH

[SW19b] Julius Schönherr and Evan Westra. Beyond ‘interaction’: How to understand social effects on social cognition. British Journal for the Philosophy of Science, 70(1):27–52, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
27/4092773.

Swanson:1967:M

[Swa67] J. W. Swanson. On models. British Journal for the Philosophy of Science, 17(4): 297–311, February 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
17/4/297.full.pdf+html;
http://www.jstor.org/
stable/686771.

Swann:1988:PI

[Swa88] Andrew J. Swann. Popper on induction. British Journal for the Philosophy of Science, 39(3): 367–373, September 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/3/367.full.pdf+html;
http://www.jstor.org/
stable/687213.

Swanson:2021:HRS

[Swa21] Noel Swanson. How to be a relativistic spacetime state realist. British Journal for the Philosophy of Science, 71(3):933– 957, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Swanson:2022:OIW

[Swa22] Noel Swanson. On the Ostrogradski instability; or, why physics really uses second derivatives. British Journal for the Philosophy of Science, 73(1):23– 46, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Swinburne:1969:VII

[Swi69]  R. G. Swinburne.  Vague-

ness, inexactness, and imprecision. British Journal for the Philosophy of Science, 19(4): 281–299, February 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
19/4/281.full.pdf+html;
http://www.jstor.org/
stable/686289.

Swinburne:1970:R

[Swi70] R. G. Swinburne. Reviews.

British Journal for the Philosophy of Science, 21(3):308–311, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/308.full.pdf+
html;   http://www.jstor.
org/stable/686211.

Swinburne:1976:R

[Swi76] Richard Swinburne. Reviews.

British Journal for the Philosophy of Science, 27(4):412– 414, December 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/4/412.full.pdf+html;
http://www.jstor.org/
stable/686872.

Swinburne:1978:R

[Swi78] R. G. Swinburne. Reviews. British Journal for the Philosophy of Science, 29(3):301– 304, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/3/301.full.pdf+html;
http://www.jstor.org/
stable/687025.

Swinburne:1980:CAS

[Swi80a] Richard Swinburne. Conventionalism about space and time. British Journal for the Philosophy of Science, 31(3):255– 272, September 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/3/255.full.pdf+html;
http://www.jstor.org/
stable/686920.

Swinburne:1980:R

[Swi80b] Richard Swinburne. Reviews. British Journal for the Philosophy of Science, 31(4):401– 404, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/4/401.full.pdf+html;

http://www.jstor.org/
stable/687068.

Swinburne:1981:R

[Swi81] Richard Swinburne. Reviews.

British Journal for the Philosophy of Science, 32(3):314– 318, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/314.full.pdf+html;
http://www.jstor.org/
stable/687598.

Swijtink:1987:Ra

[Swi87a] Zeno G. Swijtink. Reviews.

British Journal for the Philosophy of Science, 38(3):392– 396, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/392.full.pdf+html;
http://www.jstor.org/
stable/687494.

Swijtink:1987:Rb

[Swi87b] Zeno G. Swijtink. Reviews.

British Journal for the Philosophy of Science, 38(3):399– 404, September 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/3/399.full.pdf+html;
http://www.jstor.org/
stable/687496.

Swinburne:2000:DRG

[Swi00] R. Swinburne. Discussion. reply to Grünbaum. British Journal for the Philosophy of Science, 51(3):481–485, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/3/481.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541775.

Swinburne:2005:SRG

[Swi05] Richard Swinburne. Second reply to Grünbaum. British Journal for the Philosophy of Science, 56(4):919–925, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        56/4/919.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541873. See rejoinder

[Swi05].

Swoyer:1995:R

[Swo95] Chris Swoyer. Review. British Journal for the Philosophy of Science, 46(4):603–609, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/4/603.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687901.

Synge:1952:R

[Syn52] J. L. Synge. Reviews. British Journal for the Philosophy of Science, III(10):204–207, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/204.full.
pdf+html;   http://www.
jstor.org/stable/685564.

Synge:1962:Ra

[Syn62a] J. L. Synge. Reviews. British Journal for the Philosophy of Science, XIII(50):177–180, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/177.full.
pdf+html;   http://www.
jstor.org/stable/685972.

Synge:1962:Rb

[Syn62b] J. L. Synge. Reviews. British Journal for the Philosophy of Science, XIII(51):248–250, November 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/51/248.full.
pdf+html;   http://www.
jstor.org/stable/685843.

Szasz:1956:CDP

[Sza56] Thomas S. Szasz. Comments on ‘the definition of psychosomatic disorder’. British Journal for the Philosophy of Science, VII(27):231–234, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/27/231.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685875.

Szasz:1958:MM

[Sza58] Thomas S. Szasz. Men and machines. British Journal for the Philosophy of Science, VIII(32): 310–317, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/32/310.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685858.

Szumilewicz:1977:IRD

[Szu77] Irena Szumilewicz. Incommensurability and the rationality of the development of science. British Journal for the Philosophy of Science, 28(4):345–350, December 1977. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/28/4/345.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686878.

Tahko:2021:WDY

[Tah21] Tuomas E. Tahko. Where do you get your protein? or: Biochemical realization. British Journal for the Philosophy of Science, 71 (3):799–825, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Tallant:2013:PMC

[Tal13] Jonathan Tallant. Pretense, mathematics, and cognitive neuroscience. British Journal for the Philosophy of Science, 64(4): 817–835, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/817.full.pdf+html.

Tal:2015:MWS

[Tal15] Eran Tal. Book review: Michael Weisberg, Simulation and Similarity: Using Models to Understand the World. British Journal for the Philosophy of Science, 66(2):469–473, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
66/2/469.

Tal:2016:MTS

[Tal16] Eran Tal. Making time: a study in the epistemology of measurement. unknown, 67(1): 297–335, March 2016. CODEN ???? ISSN ???? URL http:/

/bjps.oxfordjournals.org/
content/67/1/297.

Tanner:1967:R

[Tan67]  Michael  Tanner.

Reviews. British Journal for the Philosophy of Science, 18(1):76– 77, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/76.full.pdf+
html;   http://www.jstor.
org/stable/686896.

Tanji:1989:QTL

[Tan89] Nobuharu Tanji. Quine on theory and language. British Journal for the Philosophy of Science, 40(2):233–247, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        40/2/233.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687515.

Tappenden:2000:IPE

[Tap00] P. Tappenden. Identity and probability in Everett’s multiverse. British Journal for the Philosophy of Science, 51(1): 99–114, March 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/1/99.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541750.

Tappenden:2008:SWE

[Tap08] Paul Tappenden. Saunders and Wallace on Everett and Lewis. British Journal for the Philosophy of Science, 59(3):307– 314, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        59/3/307.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/40072287.

Tappenden:2011:EUE

[Tap11] Paul Tappenden. Evidence and uncertainty in Everett’s multiverse. British Journal for the Philosophy of Science, 62 (1):99–123, March 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/1/99.full.pdf+html.

Taub:2006:BRA

[Tau06] Liba Taub. Book review: Andrew Barker: Scientific Method in Ptolemy’s ‘Harmonics’, Cambridge: Cambridge University Press, 2000 cloth £45.00 ISBN 0-521-55372-5. British Journal for the Philosophy of Science, 57 (1):267–271, March 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/1/267.full.pdf+html;
http://www.jstor.org/
stable/3541664.

Tavakol:1991:FDM

[Tav91] R. K. Tavakol. Fragility and deterministic modelling in the exact sciences. British Journal for the Philosophy of Science, 42(2):147–156, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/147.full.pdf+html;
http://www.jstor.org/
stable/688049.

Taylor:1952:R

[Tay52] F. Sherwood Taylor. Reviews.

British Journal for the Philosophy of Science, III(9):88–89, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/9/88.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685477.

Taylor:2018:APP

[Tay18] Henry Taylor. Attention, psychology, and pluralism. British Journal for the Philosophy of Science, 69(4):935–956, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/4/935/4091079.

Taylor:2023:ED

[Tay23] Elanor Taylor. Explanatory distance. British Journal for the Philosophy of Science, 74(1): 221–239, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Thebault:2018:MI

[TBR18] Karim Thébault, Seamus Bradley, and Alexander Reutlinger. Modelling inequality. British Journal for the Philosophy of Science, 69(3):691–718, September 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/3/691/3056505.

Teh:2016:GG

[Teh16] Nicholas J. Teh. Gravity and gauge. British Journal for the Philosophy of Science, 67 (2):497–530, June 2016. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/497.

Teichmann:1962:PSM

[Tei62] Max Teichmann. Psychoanalysis and scientific method. British Journal for the Philosophy of Science, XIII(49):56– 65, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/56.full.
pdf+html;   http://www.
jstor.org/stable/685229.

Teichmann:1999:BRB

[Tei99] R. Teichmann. Book review:

Questions of Time and Tense. R. Le Poidevin [ed]. British Journal for the Philosophy of Science, 50(4):781–786, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/4/781.full.pdf+
html;   http://www.jstor.
org/stable/40072269.

Teira:2013:BRP

[Tei13]  David Teira.  Book review:

Philip Dawid, William Twining and Mimi Vasilaki (eds), Evidence, Inference, and Enquiry. British Journal for the Philosophy of Science, 64(3): 665–668, September 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        64/3/665.full.pdf+html.

Teller:1969:GTP

[Tel69] Paul Teller. Goodman’s theory of projection. British Journal for the Philosophy of Science, 20 (3):219–238, October 1969. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        20/3/219.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686416.

Teller:1971:Ra

[Tel71a] Paul Teller. Reviews. British Journal for the Philosophy of Science, 22(1):61–62, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        22/1/61.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686648.

Teller:1971:Rb

[Tel71b] Paul Teller. Reviews. British Journal for the Philosophy of Science, 22(4):378–382, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/22/4/378.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686321.

Teller:1986:RHQ

[Tel86] Paul Teller. Relational holism and quantum mechanics. British Journal for the Philosophy of Science, 37(1):71–81, March 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686998.

Tennant:1981:LP

[Ten81] Neil Tennant. From logic to philosophies. British Journal for the Philosophy of Science, 32(3):287–301, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/287.full.pdf+html;
http://www.jstor.org/
stable/687593.

Tennant:1985:MLA

[Ten85] Neil Tennant. Minimal logic is adequate for Popperian science. British Journal for the Philosophy of Science, 36(3):325– 329, September 1985. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
36/3/325.full.pdf+html;
http://www.jstor.org/
stable/687577.

Tennant:1994:CTT

[Ten94] Neil Tennant. Changing the theory of theory change: Towards a computational approach. British Journal for the Philosophy of Science, 45(3):865–897, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/3/865.full.pdf+
html;   http://www.jstor.
org/stable/687798.

Tennant:1997:CTT

[Ten97] Neil Tennant. Changing the theory of theory change: Reply to my critics. British Journal for the Philosophy of Science, 48(4):569–586, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
48/4/569.full.pdf+html;
http://www.jstor.org/
stable/688101.

Titelbaum:2020:PPD

[TH20] Michael G. Titelbaum and Casey Hart. The principal principle does not imply the principle of indifference, because conditioning on biconditionals is counterintuitive. British Journal for the Philosophy of Science, 71 (2):621–632, June 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
621/4815703.

Thalberg:1970:DNL

[Tha70] Irving Thalberg. Discussion: New light on brain physiology and free will? British Journal for the Philosophy of Science, 21(4):379–383, November 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/379.full.pdf+html;
http://www.jstor.org/
stable/686559.

Thalos:1998:MPI

[Tha98] Mariam Thalos. A modest proposal for interpreting structural explanations. British Journal for the Philosophy of Science, 49(2):279–295, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
49/2/279.full.pdf+html;
http://www.jstor.org/
stable/688115.

Thagard:2009:BRH

[Tha09] Paul Thagard. Book review:

Hanne Andersen, Peter Barker and Xian Chen. The Cognitive Structure of Scientific Revolutions. British Journal for the Philosophy of Science, 60(4): 843–847, December 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/843.full.pdf+html;
http://www.jstor.org/
stable/25592037.

Theobald:1970:R

[The70] D. W. Theobald. Reviews.

British Journal for the Philosophy of Science, 21(2):206–208, May 1970. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/21/2/206.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686704.

Thebault:2012:SRP

[Thé12] Karim P. Y. Thébault. Symplectic reduction and the problem of time in nonrelativistic mechanics. British Journal for the Philosophy of Science, 63(4): 789–824, December 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/4/789.full.pdf+html.

Thebault:2016:QGO

[Thé16] Karim P. Y. Thébault. Quantization as a guide to ontic structure. unknown, 67(1):89– 114, March 2016. CODEN ????

                                        ISSN ???? URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/1/89.

Thorpe:1952:WF

[Tho52] W. H. Thorpe. The world of form. British Journal for the Philosophy of Science, II(8): 318–322, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/8/318.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686022.

Thomson:1954:GMP

[Tho54a] George Thomson. Galileo’s methodological problems. British Journal for the Philosophy of Science, V(19):253–256, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/19/253.full.pdf+html;
http://www.jstor.org/
stable/685720.

Thorpe:1954:RI

[Tho54b] W. H. Thorpe. Reviews: Instinct. British Journal for the Philosophy of Science, V(17):72– 76, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/17/72.full.pdf+
html;   http://www.jstor.
org/stable/685955.

Thompson:1965:CMC

[Tho65] Dennis Thompson. Can a machine be conscious? British Journal for the Philosophy of Science, XVI(61):33–43, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/61/33.full.pdf+
html;   http://www.jstor.
org/stable/686137.

Thomson:1966:R

[Tho66a] G. P. Thomson. Reviews.

British Journal for the Philosophy of Science, XVI(64):348– 349, February 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/64/348.full.pdf+html;
http://www.jstor.org/
stable/686840.

Thorpe:1966:R

[Tho66b] W. H. Thorpe. Reviews. British Journal for the Philosophy of Science, 17(3):261–263, November 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/3/261.full.pdf+
html;   http://www.jstor.
org/stable/686766.

Thomas:1975:PMP

[Tho75] Arthur J. Thomas. Puccetti on machine pattern recognition. British Journal for the Philosophy of Science, 26(3):227– 232, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/227.full.pdf+html;
http://www.jstor.org/
stable/686279.

Thompson:1988:RDP

[Tho88] Ian J. Thompson. Real dispositions in the physical world. British Journal for the Philosophy of Science, 39(1):67–79, March 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/1/67.full.pdf+
html;   http://www.jstor.
org/stable/687115.

Thomason:1992:CLE

[Tho92] Neil Thomason. Could Lakatos, even with Zahar’s criterion for novel fact, evaluate the Copernican research programme? British Journal for the Philosophy of Science, 43(2):161–200, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/2/161.full.pdf+
html;   http://www.jstor.
org/stable/687871.

Thomason:1994:PAH

[Tho94a] Neil Thomason. The power of arched hypotheses: Feyerabend’s Galileo as a closet rationalist. British Journal for the Philosophy of Science, 45(1):255–264, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/255.full.pdf+
html;   http://www.jstor.
org/stable/687971.

Thornton:1994:DBD

[Tho94b] M. Thornton. Double brain.

double person? British Journal for the Philosophy of Science, 45(2):761–763, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        45/2/761.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687697.

Thomas:1996:R

[Tho96] Alan Thomas. Review. British Journal for the Philosophy of Science, 47(2):317–319, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/47/2/317.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687951.

Tianji:1985:SRF

[Tia85] Jiang Tianji. Scientific rationality, formal or informal? British Journal for the Philosophy of Science, 36(4):409–423, December 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/36/4/409.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687400.

Tichy:1974:PDV

[Tic74] Pavel Tichý. On Popper’s definitions of verisimilitude. British Journal for the Philosophy of Science, 25(2):155–160, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/25/2/155.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686819.

Tichy:1976:VR

[Tic76] Pavel Tichý. Verisimilitude redefined. British Journal for the Philosophy of Science, 27 (1):25–42, March 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
27/1/25.full.pdf+html;
http://www.jstor.org/
stable/686376.

Tichy:1980:R

[Tic80] Pavel Tichý. Reviews. British Journal for the Philosophy of Science, 31(3):301–307, September 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/3/301.full.pdf+
html;   http://www.jstor.
org/stable/686925.

Tiles:1984:BSO

[Til84] Mary Tiles. Bachelard, science and objectivity. Modern European philosophy. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-24803-5, 0-521-28973-4 (paperback). xxii + 242 pp. LCCN B2430.B254

T55 1984. URL http://www.
loc.gov/catdir/description/
cam022/84005001.html;
http://www.loc.gov/catdir/
toc/cam029/84005001.html.

Tiles:1992:EEV

[Til92] J. E. Tiles. Experimental evidence vs. experimental practice?

British Journal for the Philosophy of Science, 43(1):99–109, March 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/43/1/99.full.pdf+
html;   http://www.jstor.
org/stable/687886.

Tiles:1993:EI

[Til93a] J. E. Tiles. Experiment as intervention. British Journal for the Philosophy of Science, 44(3):463–475, September 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/463.full.pdf+html;
http://www.jstor.org/
stable/688016.

Tiles:1993:NPC

[Til93b] Mary Tiles. The normal and pathological: the concept of a scientific medicine. British Journal for the Philosophy of Science, 44(4):729–742, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/4/729.full.pdf+html;
http://www.jstor.org/
stable/688041.

Tiles:1994:ODE

[Til94] J. E. Tiles. One dimensional experimental science. British Journal for the Philosophy of Science, 45(1):341–352, March 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/1/341.full.pdf+html;
http://www.jstor.org/
stable/687977.

Tiles:2000:BRH

[Til00]  Mary Tiles.  Book review:

Hugh Lacey — Is Science Value Free? Values and Scientific Understanding. British Journal for the Philosophy of Science, 51(4):953–955, December 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/4/953.full.pdf+html;
http://www.jstor.org/
stable/3541743.

Timpson:2006:GT

[Tim06] Christopher G. Timpson. The grammar of teleportation. British Journal for the Philosophy of Science, 57(3):587–621, September 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/3/587.full.pdf+
html;   http://www.jstor.
org/stable/3873484.

Tizard:1957:CIE

[Tiz57] R. H. Tizard. A cybernetic interpretation of economics. British Journal for the Philosophy of Science, VIII(30):161–164, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/30/161.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685796.

Tizard:1976:PDI

[Tiz76] Jack Tizard. Progress and degeneration in the ‘IQ Debate’: Comments on Urbach. British Journal for the Philosophy of Science, 27(3):251–258, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/27/3/251.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686122.

Todd:2009:DMD

[Tod09] Steven J. Todd. A difference that makes a difference: Passing through Dennett’s Stalinesque/ Orwellian impasse. British Journal for the Philosophy of Science, 60(3):497–520, September 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/3/497.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592017.

Tomassi:1999:BRB

[Tom99] P. Tomassi. Book review: The digital phoenix; how computers are changing philosophy. Terrell Ward Bynum, James H Moor (eds). British Journal for the Philosophy of Science, 50(3): 514–519, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/514.full.pdf+html;
http://www.jstor.org/
stable/40072253.

Tooley:1995:R

[Too95] Michael Tooley. Review. British Journal for the Philosophy of Science, 46(2):280–283, June 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/2/280.full.pdf+
html;   http://www.jstor.
org/stable/687820.

Torretti:1984:STP

[Tor84] Roberto Torretti. Space–time physics and the philosophy of science. British Journal for the Philosophy of Science, 35(3): 280–292, September 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/3/280.full.pdf+html;
http://www.jstor.org/
stable/687478.

Torretti:1987:DCF

[Tor87] Roberto Torretti. Do conjunctive forks always point to a common cause? British Journal for the Philosophy of Science, 38(3):384–387, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        38/3/384.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687492.

Torre:2015:RDC

[Tor15] Stephan Torre. Restricted diachronic composition and Special Relativity. British Journal for the Philosophy of Science, 66(2):235–255, June 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/2/235.

Tosh:2016:FFB

[Tos16] Nick Tosh. Finite frequentism in a big world. unknown, 67(1):169– 213, March 2016. CODEN ????

                                        ISSN ???? URL http://bjps.
                                        oxfordjournals.org/content/
                                        67/1/169.

Toulmin:1953:R

[Tou53]  Stephen Toulmin.  Reviews.

British Journal for the Philosophy of Science, IV(15):256– 258, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/15/256.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685907.

Toulmin:1958:R

[Tou58]  Stephen Toulmin.  Reviews.

British Journal for the Philosophy of Science, IX(35):247– 248, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/247.full.pdf+html;
http://www.jstor.org/
stable/685657.

Toulmin:1960:R

[Tou60]  Stephen Toulmin.  Reviews.

British Journal for the Philosophy of Science, X(40):346– 349, February 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/346.full.pdf+html;
http://www.jstor.org/
stable/685925.

Townshend:1956:R

[Tow56]  Hugh Townshend.  Reviews.

British Journal for the Philosophy of Science, VII(26):177–179, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/26/177.full.
pdf+html;   http://www.
jstor.org/stable/685704.

Trpin:2019:IBE

[TP19] Borut Trpin and Max Pellert.

Inference to the best explanation in uncertain evidential situations. British Journal for the Philosophy of Science, 70(4): 977–1001, December 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.

                                        oup.com/bjps/article/70/4/

977/4935150. See corrigendum [TP21].

Trpin:2021:CIB

[TP21] Borut Trpin and Max Pellert. Corrigendum to: Inference to the Best Explanation in Uncertain Evidential Situations. British Journal for the Philosophy of Science, 72(1):??, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic). See [TP19].

Treisman:1962:PEP

[Tre62] Michel Treisman. Psychological explanation: the ‘private data’ hypothesis. British Journal for the Philosophy of Science, XIII(50):130–143, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/50/130.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685966.

Tregear:2004:UEF

[Tre04] Mark Tregear. Utilising explanatory factors in induction? British Journal for the Philosophy of Science, 55(3):505– 519, September 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/3/505.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541672.

Trigg:1987:R

[Tri87] Roger Trigg. Reviews. British Journal for the Philosophy of Science, 38(3):389–392, September 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/3/389.full.pdf+
html;   http://www.jstor.
org/stable/687493.

Triplett:1988:ALV

[Tri88] Timm Triplett. Azande logic versus Western logic? British Journal for the Philosophy of Science, 39(3):361–366, September 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/3/361.full.pdf+
html;   http://www.jstor.
org/stable/687212.

Triplett:1994:TAE

[Tri94] Timm Triplett. Is there anthropological evidence that logic is culturally relative?: Remarks on Bloor, Jennings, and Evans-Pritchard. British Journal for the Philosophy of Science, 45 (2):749–760, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/749.full.pdf+html;
http://www.jstor.org/
stable/687696.

Trout:1993:RIS

[Tro93] J. D. Trout. Robustness and integrative survival in significance testing: the world’s contribution to rationality. British Journal for the Philosophy of Science, 44(1):1–15, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/1/1.full.pdf+html;
http://www.jstor.org/
stable/687846.

Thomson:1953:CMF

[TS53] R. Thomson and W. Sluckin. Cybernetics and mental functioning. British Journal for the Philosophy of Science, IV(14): 130–146, August 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/14/130.full.pdf+html;
http://www.jstor.org/
stable/685988.

Tsou:2010:BRR

[Tso10] Jonathan Y. Tsou. Book review: Rachel Cooper, Classifying Madness: a Philosophical Examination of the Diagnostic and Statistical Manual of Mental Disorders. British Journal for the Philosophy of Science, 61(2):453–457, June 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/2/453.full.pdf+html.

Tucker:1958:TTP

[Tuc58] John Tucker. The television theory of perception. British Journal for the Philosophy of Science, IX(33):51–57, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/33/51.full.pdf+html;
http://www.jstor.org/
stable/686102.

Tucker:2011:HSU

[Tuc11]  Aviezer Tucker.  Historical

science, over- and underdetermined: a study of Darwin’s inference of origins. British Journal for the Philosophy of Science, 62(4):805–829, December 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/4/805.full.pdf+html.

Tulodziecki:2012:EEE

[Tul12] Dana Tulodziecki. Epistemic equivalence and epistemic incapacitation. British Journal for the Philosophy of Science, 63 (2):313–328, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/313.full.pdf+html.

Tumulka:2007:DVQ

[Tum07] Roderich Tumulka. Determinate values for quantum observables. British Journal for the Philosophy of Science, 58(2):355–360, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/2/355.full.pdf+
html;   http://www.jstor.
org/stable/30115232.

Tuomela:1978:SRP

[Tuo78] Raimo Tuomela. Scientific realism and perception. British Journal for the Philosophy of Science, 29(1):87–104, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/1/87.full.pdf+
html;   http://www.jstor.
org/stable/686395.

Tuomela:1989:RMS

[Tuo89] Raimo Tuomela. Ruben and the metaphysics of the social world. British Journal for the Philosophy of Science, 40(2):261–273, June 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/40/2/261.full.pdf+
html;   http://www.jstor.
org/stable/687518.

Turner:1955:MMP

[Tur55] Joseph Turner. Maxwell on the method of physical analogy. British Journal for the Philosophy of Science, VI(23): 226–238, November 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
VI/23/226.full.pdf+html;
http://www.jstor.org/
stable/685678.

Turner:1960:R

[Tur60] H. D. Turner. Reviews. British Journal for the Philosophy of Science, XI(42):150–153, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/150.full.pdf+
html;   http://www.jstor.
org/stable/685590.

Turney:1990:CFP

[Tur90] Peter Turney. The curve fitting problem: a solution. British Journal for the Philosophy of Science, 41(4):509–530, December 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/4/509.full.pdf+
html;   http://www.jstor.
org/stable/687863.

Turney:1991:NPE

[Tur91] Peter Turney. A note on Popper’s equation of simplicity with falsifiability. British Journal for the Philosophy of Science, 42 (1):105–109, March 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/1/105.full.pdf+html;
http://www.jstor.org/
stable/687610.

Tuziak:1993:R

[Tuz93] Roman Tuziak. Review. British Journal for the Philosophy of Science, 44(1):167–170, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/1/167.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687856.

Uebel:1994:R

[Ueb94] Thomas E. Uebel. Review. British Journal for the Philosophy of Science, 45(4):1066– 1069, December 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/4/1066.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687624.

Uebel:1996:AFV

[Ueb96]  Thomas  E.

Uebel. Anti-foundationalism and the Vienna Circle’s revolution in philosophy. British Journal for the Philosophy of Science, 47(3):415–440, September 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        47/3/415.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687986.

Uebel:1997:R

[Ueb97] Thomas E. Uebel. Review. British Journal for the Philosophy of Science, 48(2):306–309, June 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/2/306.full.pdf+
html;   http://www.jstor.
org/stable/687757.

Uebel:1998:R

[Ueb98] Thomas Uebel. Review. British Journal for the Philosophy of Science, 49(2):359–364, June 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/2/359.full.pdf+
html;   http://www.jstor.
org/stable/688125.

Uebel:1999:DPA

[Ueb99] T. E. Uebel. Discussion. protocols, affirmations, and foundations: reply to Oberdan. British Journal for the Philosophy of Science, 50(2):297–300, June 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/50/2/297.full.pdf+
html;   http://www.jstor.
org/stable/40072225.

Uebel:2001:BRB

[Ueb01] Thomas Uebel. Book review:

Knowledge and Reality: an Essay in Positive Philosophy, Paolo Parrini. British Journal for the Philosophy of Science, 52 (2):407–411, June 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
52/2/407.full.pdf+html;
http://www.jstor.org/
stable/3541907.

Uller:2019:NCC

[UH19] Tobias Uller and Heikki Helanterä. Niche construction and conceptual change in evolutionary biology. British Journal for the Philosophy of Science, 70(2):351–375, June 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/2/
351/4098122.

Urbach:1974:PDIa

[Urb74a] Peter Urbach. Progress and degeneration in the ‘IQ Debate’ (I). British Journal for the Philosophy of Science, 25(2):99–135, June 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/25/2/99.full.pdf+
html;   http://www.jstor.
org/stable/686817.

Urbach:1974:PDIb

[Urb74b] Peter Urbach. Progress and degeneration in the ‘IQ Debate’ (II). British Journal for the Philosophy of Science, 25(3):235– 259, September 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/3/235.full.pdf+html;

http://www.jstor.org/
stable/686906.

Urbach:1977:DPW

[Urb77] Peter Urbach. Discussion: Putnam and Wilks and mind and body. British Journal for the Philosophy of Science, 28 (1):59–63, March 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/1/59.full.pdf+html;
http://www.jstor.org/
stable/686448.

Urbach:1978:PAA

[Urb78] Peter Urbach. Is any of Popper’s arguments against historicism valid? British Journal for the Philosophy of Science, 29(2):117–130, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/2/117.full.pdf+html;
http://www.jstor.org/
stable/687122.

Urbach:1980:SP

[Urb80] Peter Urbach. Social propensities. British Journal for the Philosophy of Science, 31(4): 317–328, December 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
31/4/317.full.pdf+html;
http://www.jstor.org/
stable/687061.

Urbach:1981:R

[Urb81] Peter Urbach. Reviews. British Journal for the Philosophy of Science, 32(4):425–426, December 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/32/4/425.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687315.

Urbach:1982:FBP

[Urb82]  Peter Urbach.  Francis Ba-

con as a precursor to Popper. British Journal for the Philosophy of Science, 33(2):113–132, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/113.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687274.

Urbach:1983:ISS

[Urb83] Peter Urbach. Intimations of similarity: the shaky basis of verisimilitude. British Journal for the Philosophy of Science, 34(3):266–275, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        34/3/266.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687324.

Urbach:1988:WLN

[Urb88] Peter Urbach. What is a law of nature? A Humean answer.

British Journal for the Philosophy of Science, 39(2):193–209, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/2/193.full.pdf+
html;   http://www.jstor.
org/stable/687266.

Urbach:1992:RAC

[Urb92] Peter Urbach. Regression analysis: Classical and Bayesian. British Journal for the Philosophy of Science, 43(3):311– 342, September 1992. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/3/311.full.pdf+html;
http://www.jstor.org/
stable/687762.

Urbach:1994:RDP

[Urb94] Peter Urbach. Reply to David Papineau. British Journal for the Philosophy of Science, 45 (2):712–715, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/712.full.pdf+html;
http://www.jstor.org/
stable/687693.

Yi:1994:GE

[uY94] Byeong uk Yi. Glymour on explanation. British Journal for the Philosophy of Science, 45(3):914–917, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/3/914.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687801.

Vaidman:2015:BRD

[Vai15] Lev Vaidman. Book review: David Wallace, The Emergent Multiverse: Quantum Theory According to the Everett Interpretation. British Journal for the Philosophy of Science, 66 (2):465–468, June 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        66/2/465.

Vallinder:2018:IBG

[Val18]  Aron Vallinder.  Im-

precise Bayesianism and global belief inertia. British Journal for the Philosophy of Science, 69(4):1205–1230, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/69/4/1205/4091081.

VanHeerden:1961:PPD

[Van61a] P. J. Van Heerden. A point of Professor Dingle’s. British Journal for the Philosophy of Science, XII(45):70, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XII/45/70.full.pdf+html;

http://www.jstor.org/
stable/685542.

VanLennep:1961:R

[Van61b] D. J. Van Lennep. Reviews.

British Journal for the Philosophy of Science, XI(44):344– 345, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/344.full.pdf+html;
http://www.jstor.org/
stable/685139.

VanValen:1963:ET

[Van63] Leigh Van Valen. On evolutionary theories. British Journal for the Philosophy of Science, XIV(54):146–152, August 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/54/146.full.pdf+html;
http://www.jstor.org/
stable/685432.

VanFraassen:1981:PRI

[Van81] Bas C. Van Fraassen. A problem for relative information minimizers in probability kinematics. British Journal for the Philosophy of Science, 32(4): 375–379, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/375.full.pdf+html;
http://www.jstor.org/
stable/687309.

VanRooijen:1987:IEC

[Van87] Jeroen Van Rooijen. Interactionism and evolution: a critique of Popper. British Journal for the Philosophy of Science, 38(1):87–92, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        38/1/87.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687083.

VanBalen:1988:DSC

[Van88]  G. Van Balen.  The Dar-

winian Synthesis: a critique of the Rosenberg/Williams argument. British Journal for the Philosophy of Science, 39(4): 441–448, November 1988. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        39/4/441.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687431.

VanBrakel:1993:PCC

[Van93a] J. Van Brakel. The plasticity of categories: the case of colour. British Journal for the Philosophy of Science, 44(1):103–135, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/44/1/103.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687852.

VanBrakel:1993:PER

[Van93b] J. Van Brakel. Polywater and experimental realism. British Journal for the Philosophy of Science, 44(4):775–784, December 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/4/775.full.pdf+
html;   http://www.jstor.
org/stable/688044.

VanAndel:1994:AUF

[Van94a] Pek Van Andel. Anatomy of the unsought finding. serendipity: Orgin, history, domains, traditions, appearances, patterns and programmability. British Journal for the Philosophy of Science, 45(2):631–648, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/631.full.pdf+html;
http://www.jstor.org/
stable/687687.

VanBendegem:1994:RPI

[Van94b] Jean Paul Van Bendegem. Ross’ Paradox is an impossible supertask. British Journal for the Philosophy of Science, 45(2): 743–748, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/2/743.full.pdf+html;
http://www.jstor.org/
stable/687695.

VanBrakel:1994:IFS

[Van94c] J. Van Brakel. The ignis fatuus of semantic universalia: the case of colour. British Journal for the Philosophy of Science, 45 (2):770–783, June 1994. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        45/2/770.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687699.

VanFraassen:2008:SRP

[Van08a] Bas C. Van Fraassen. Scientific representation: paradoxes of perspective. Clarendon Press, New York, NY, USA, 2008. ISBN 0-19-927822-9. xiv + 408 pp. LCCN ????

Vanderschraaf:2008:GTM

[Van08b] Peter Vanderschraaf. Game theory meets threshold analysis: Reappraising the paradoxes of anarchy and revolution. British Journal for the Philosophy of Science, 59(4):579–617, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/59/4/579.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072304.

VanFraassen:2010:SRP

[Van10] Bas C. Van Fraassen. Scientific representation: paradoxes of perspective. Clarendon Press, New York, NY, USA, 2010. ISBN 0-19-927823-7 (paperback). xiv + 408 pp. LCCN ???? UK £18.99.

Vassend:2021:VFI

[Vas21a] Olav B. Vassend. A verisimilitude framework for inductive inference, with an application to phylogenetics. British Journal for the Philosophy of Science, 71 (4):1359–1383, ???? 2021. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Vasudevan:2021:EIR

[Vas21b] Anubav Vasudevan. Entropy and insufficient reason: a note on the Judy Benjamin problem. British Journal for the Philosophy of Science, 71(3):1113–1141, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Vassend:2022:JNI

[Vas22] Olav Benjamin Vassend. Justifying the norms of inductive inference. British Journal for the Philosophy of Science, 73(1): 135–160, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Vindrola:2024:BTS

[VC24] Filippo Vindrola and Vincenzo Crupi. Bayesians too should follow Wason: a comprehensive accuracy-based analysis of the selection task. British Journal for the Philosophy of Science, 75 (2):347–373, June 2024. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

vanEck:2006:ENW

[vEDS06] Dingmar van Eck, Huib Looren De Jong, and Maurice K. D. Schouten. Evaluating new wave reductionism: the case of vision. British Journal for the Philosophy of Science, 57(1):167–196, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/57/1/167.full.pdf+
html;   http://www.jstor.
org/stable/3541657.

Velasco:2010:SGT

[Vel10] Joel D. Velasco. Species, genes, and the tree of life. British Journal for the Philosophy of Science, 61(3):599–619, September 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/599.full.pdf+html.

Vermaas:1999:PUQ

[Ver99] Pieter E. Vermaas. A philosopher’s understanding of quantum mechanics: possibilities and impossibilities of a modal interpretation. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-65108-5. xi + 295 pp. LCCN QC174.12 .V46 1999.

Vermaas:2005:TCI

[Ver05] Pieter E. Vermaas. Technology and the conditions on interpretations of quantum mechanics. British Journal for the Philosophy of Science, 56(4):

635–661, December 2005. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
56/4/635.full.pdf+html;
http://www.jstor.org/
stable/3541861.

Vesey:1963:R

[Ves63]  G. N. A. Vesey.  Reviews.

British Journal for the Philosophy of Science, XIV(53):74–75, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIV/53/74.full.pdf+
html;   http://www.jstor.
org/stable/685360.

Vessonen:2020:CPR

[Ves20] Elina Vessonen. The complementarity of psychometrics and the representational theory of measurement. British Journal for the Philosophy of Science, 71 (2):415–442, June 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
415/5050796.

vanFraassen:2006:SSS

[vF06] Bas C. van Fraassen. Structure: its shadow and substance. British Journal for the Philosophy of Science, 57(2):275–307, June 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/57/2/275.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3873468.

vanFraassen:2017:UPT

[vFH17]  Bas  C.  van

Fraassen and Joseph Y. Halpern. Updating probability: Tracking statistics as criterion. British Journal for the Philosophy of Science, 68(3):725–743, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/3/725/2669633.

Vermaas:2003:AFT

[VH03] Pieter E. Vermaas and Wybo Houkes. Ascribing functions to technical artefacts: a challenge to etiological accounts of functions. British Journal for the Philosophy of Science, 54 (2):261–289, June 2003. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        54/2/261.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541967.

VanFraassen:1986:PRI

[VHH86] Bas C. Van Fraassen, R. I. G. Hughes, and Gilbert Harman. A problem for relative information minimizers, continued. British Journal for the Philosophy of Science, 37(4):453– 463, December 1986. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687370.

Vickers:2008:FMB

[Vic08] Peter Vickers. Frisch, Muller, and Belot on an inconsistency in classical electrodynamics. British Journal for the Philosophy of Science, 59(4):767–792, December 2008. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/59/4/767.full.pdf+
html;   http://www.jstor.
org/stable/40072311.

Vicente:2011:CPP

[Vic11] Agustín Vicente. Current physics and ‘the physical’. British Journal for the Philosophy of Science, 62(2):393–416, June 2011. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
62/2/393.full.pdf+html.

Vickers:2021:DUH

[Vic21] Peter Vickers. Disarming the ultimate historical challenge to scientific realism. British Journal for the Philosophy of Science, 71(3):987–1012, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Viera:2020:ST

[Vie20]  Gerardo Viera.  The sense

of time. British Journal for the Philosophy of Science, 71 (2):443–469, June 2020. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/2/
443/4935145.

VanDerSteen:1991:LNH

[VK91] Wim J. Van Der Steen and Harmke Kamminga. Laws and natural history in biology. British Journal for the Philosophy of Science, 42(4): 445–467, December 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/4/445.full.pdf+html;
http://www.jstor.org/
stable/687703.

Vicente:2016:BCP

[VM16] Agustín Vicente and Fernando Martínez Manrique. The big concepts paper: a defence of hybridism. unknown, 67(1):59–88, March 2016. CODEN ???? ISSN ???? URL

http://bjps.oxfordjournals.
org/content/67/1/59.

Vollrath:1991:EEV

[Vol91] John Vollrath. Ellis, epistemic values, and the problem of induction. British Journal for the Philosophy of Science, 42 (2):257–261, June 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/2/257.full.pdf+html;

http://www.jstor.org/
stable/688057.

Volpe:1995:SAC

[Vol95] Giorgio Volpe. A semantic approach to comparative verisimilitude. British Journal for the Philosophy of Science, 46(4): 563–581, December 1995. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
46/4/563.full.pdf+html;
http://www.jstor.org/
stable/687898.

VonBertalanffy:1950:OGS

[Von50] Ludwig Von Bertalanffy. An outline of general system theory. British Journal for the Philosophy of Science, I(2):134–165, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/2/134.full.pdf+
html;   http://www.jstor.
org/stable/685808.

VonWeizsacker:1973:PQM

[Von73] C. F. Von Weizsäcker. Probability and quantum mechanics. British Journal for the Philosophy of Science, 24(4): 321–337, December 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/4/321.full.pdf+html;
http://www.jstor.org/
stable/686614.

VonLeyden:1975:R

[Von75]  W. Von Leyden.  Reviews.

British Journal for the Philosophy of Science, 26(2):174–180, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/2/174.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686518.

VonLeyden:1983:R

[Von83a] W. Von Leyden. Reviews. British Journal for the Philosophy of Science, 34(1):72–77, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/1/72.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686936.

VonPlato:1983:MAF

[Von83b] Jan Von Plato. The method of arbitrary functions. British Journal for the Philosophy of Science, 34(1):37–47, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/1/37.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686932.

VonPeschke:1987:SCR

[Von87] Joachim Von Peschke. Is spin a consequence of Relativity? a comment on Morrison. British Journal for the Philosophy of Science, 38(4):566–567, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/4/566.full.pdf+
html;   http://www.jstor.
org/stable/687361.

vanRiel:2013:IAR

[vR13] Raphael van Riel. Identity, asymmetry, and the relevance of meanings for models of reduction. British Journal for the Philosophy of Science, 64(4): 747–761, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/747.full.pdf+html.

Vranas:2004:HRP

[Vra04] Peter B. M. Vranas. Hempel’s raven paradox: a lacuna in the standard Bayesian solution. British Journal for the Philosophy of Science, 55(3):545– 560, September 2004. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/3/545.full.pdf+html;
http://www.jstor.org/
stable/3541674.

W:1952:R

[W.52] L. L. W. Reviews. British Journal for the Philosophy of Science, II(8):326a–326, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/8/326-a.full.pdf+html.

W:1956:PSE

[W56] G. J. W. Professor Sir Edmund Whittaker, F. R. S. British Journal for the Philosophy of Science, VII(26):180–181, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/26/180.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685706.

Wachbroit:1987:TRP

[Wac87] Robert Wachbroit. Theories of rationality and principles of charity. British Journal for the Philosophy of Science, 38(1):35–47, March 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/38/1/35.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687078.

Wagner:2004:MTP

[Wag04] Carl G. Wagner. Modus tollens probabilized. British Journal for the Philosophy of Science, 55(4):747–753, December 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        55/4/747.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541627.

Wagner:2012:DCE

[Wag12] Elliott O. Wagner. Deterministic chaos and the evolution of meaning. British Journal for the Philosophy of Science, 63(3):547–575, September 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/3/547.full.pdf+html.

Wagner:2015:CSM

[Wag15] Elliott O. Wagner. Conventional semantic meaning in signalling games with conflicting interests. British Journal for the Philosophy of Science, 66(4): 751–773, December 2015. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Walshe:1951:HC

[Wal51] F. M. R. Walshe. The hypothesis of cybernetics. British Journal for the Philosophy of Science, II (6):161–163, October 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/6/161.full.pdf+html;
http://www.jstor.org/
stable/685508.

Walker:1956:DPD

[Wal56a] Nigel Walker. The definition of psychosomatic disorder. British Journal for the Philosophy of Science, VI(24):265–299, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/24/265.full.pdf+
html;   http://www.jstor.
org/stable/685410.

Walker:1956:FH

[Wal56b] Nigel Walker. Freud and homeostasis. British Journal for the Philosophy of Science, VII (25):61–72, May 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/25/61.full.pdf+html;
http://www.jstor.org/
stable/685936.

Walker:1956:PDR

[Wal56c] Nigel Walker. Psychosomatic disorder: a rejoinder to Wightman and Szasz. British Journal for the Philosophy of Science, VII(27):235–236, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/27/235.full.
pdf+html;   http://www.
jstor.org/stable/685877.

Walker:1969:R

[Wal69] R. C. S. Walker. Reviews. British Journal for the Philosophy of Science, 20(4):364–366, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/20/4/364.full.pdf+
html;   http://www.jstor.
org/stable/686267.

Walsh:1996:FF

[Wal96] D. M. Walsh. Fitness and function. British Journal for the Philosophy of Science, 47(4): 553–574, December 1996. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/4/553.full.pdf+html;
http://www.jstor.org/
stable/687924.

Walsh:1997:R

[Wal97] D. M. Walsh. Review. British Journal for the Philosophy of Science, 48(4):613–617, December 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/48/4/613.full.pdf+
html;   http://www.jstor.
org/stable/688106.

Wallace:2006:EQC

[Wal06a] David Wallace. Epistemology quantized: circumstances in which we should come to believe in the Everett interpretation. British Journal for the Philosophy of Science, 57(4): 655–689, December 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/4/655.full.pdf+html;
http://www.jstor.org/
stable/4489090.

Walsh:2006:EE

[Wal06b] Denis Walsh. Evolutionary essentialism. British Journal for the Philosophy of Science, 57 (2):425–448, June 2006. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
57/2/425.full.pdf+html;
http://www.jstor.org/
stable/3873474.

Wallhagen:2007:CAD

[Wal07] Morgan Wallhagen. Consciousness and action: Does cognitive science support (mild) epiphenomenalism? British Journal for the Philosophy of Science, 58(3):539–561, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/3/539.full.pdf+html;
http://www.jstor.org/
stable/30115187.

Wallace:2010:GEC

[Wal10] David Wallace. Gravity, entropy, and cosmology: in search of clarity. British Journal for the Philosophy of Science, 61(3): 513–540, September 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/3/513.full.pdf+html.

Wallace:2012:EMQ

[Wal12]  David Wallace.  The emer-

gent multiverse: quantum theory according to the Everett Interpretation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-954696-7. xv + 530 pp. LCCN QC174.12 .W348

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1308/2012418035-b.html;
http://www.loc.gov/catdir/
enhancements/fy1308/
2012418035-d.html;
http://www.loc.gov/catdir/
enhancements/fy1308/
2012418035-t.html.

Wallace:2014:BRL

[Wal14] David Wallace. Book review:

Laura Ruetsche, Interpreting Quantum Theories. British Journal for the Philosophy of Science, 65(2):425–428, June 2014. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/2/425.

Wallace:2020:FEG

[Wal20] David Wallace. Fundamental and emergent geometry in Newtonian physics. British Journal for the Philosophy of Science, 71(1):1–32, March 2020. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/71/1/
1/4693797.

Ward:2022:RPC

[War22] Zina B. Ward. Registration pluralism and the cartographic approach to data aggregation across brains. British Journal for the Philosophy of Science, 73 (1):47–72, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wassermann:1955:SCM

[Was55] G. D. Wassermann. Some comments on methods and statements in parapsychology and other sciences. British Journal for the Philosophy of Science, VI (22):122–140, August 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VI/22/122.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685521.

Wassermann:1978:TRN

[Was78] Gerhard D. Wassermann. Testability of the role of natural selection within theories of population genetics and evolution. British Journal for the Philosophy of Science, 29(3):223– 242, September 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        29/3/223.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687010.

Watson:1951:AMM

[Wat51] George Watson. Apparent motion and the mind–body problem. British Journal for the Philosophy of Science, II(7):236– 247, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.

oxfordjournals.org/content/
II/7/236.full.pdf+html;
http://www.jstor.org/
stable/685294.

Watkins:1952:ITH

[Wat52a] J. W. N. Watkins. Ideal types and historical explanation. British Journal for the Philosophy of Science, III(9):22–43, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/22.full.pdf+
html;   http://www.jstor.
org/stable/685472.

Watkins:1952:PMI

[Wat52b] J. W. N. Watkins. The principle of methodological individualism. British Journal for the Philosophy of Science, III(10):186–189, August 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/10/186.full.
pdf+html;   http://www.
jstor.org/stable/685556.

Watkins:1952:SS

[Wat52c] J. W. N. Watkins. Social science. British Journal for the Philosophy of Science, III(11): 261–265, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/261.full.pdf+html;
http://www.jstor.org/
stable/685270.

Watkins:1953:R

[Wat53] J. W. N. Watkins. Reviews. British Journal for the Philosophy of Science, IV(16):352– 353, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/16/352.full.pdf+html.

Watkins:1954:RU

[Wat54] J. W. N. Watkins. Review: [untitled]. British Journal for the Philosophy of Science, IV(16): 352–353, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686090.

Watkins:1955:Ra

[Wat55a] J. W. N. Watkins. Reviews. British Journal for the Philosophy of Science, V(20):348– 350, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/20/348.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685743.

Watkins:1955:Rb

[Wat55b] J. W. N. Watkins. Reviews. British Journal for the Philosophy of Science, VI(22):169–171, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/22/169.full.pdf+

html;   http://www.jstor.
org/stable/685526.

Watkins:1955:RDU

[Wat55c] J. W. N. Watkins. Reviews: Decisions and uncertainty. British Journal for the Philosophy of Science, VI(21):66–78, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/21/66.full.pdf+
html;   http://www.jstor.
org/stable/685578.

Watkins:1957:HES

[Wat57a] J. W. N. Watkins. Historical explanation in the social sciences. British Journal for the Philosophy of Science, VIII(30): 104–117, August 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/30/104.full.pdf+html;
 http://www.jstor.org/
stable/685790.

Watkins:1957:R

[Wat57b] J. W. N. Watkins. Reviews.

British Journal for the Philosophy of Science, VII(28):362– 363, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/362.full.pdf+html;
http://www.jstor.org/
stable/685124.

Watkins:1959:Ra

[Wat59a] J. W. N. Watkins. Reviews. British Journal for the Philosophy of Science, X(38):160–162, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/X/38/160.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686009.

Watkins:1959:Rb

[Wat59b] J. W. N. Watkins. Reviews. British Journal for the Philosophy of Science, X(39):250– 251, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        X/39/250.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686064.

Watkins:1959:TTM

[Wat59c] J. W. N. Watkins. ‘The two theses of methodological individualism’. British Journal for the Philosophy of Science, IX(36): 319–320, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/36/319.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685193.

Watkins:1959:TRM

[Wat59d] J. W. N. Watkins. Third reply to Mr Goldstein. British Journal for the Philosophy of Science, X(39):242–244, November 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/39/242.full.pdf+
html;   http://www.jstor.
org/stable/686061.

Watkins:1960:CBK

[Wat60a] J. W. N. Watkins. Confirmation without background knowledge. British Journal for the Philosophy of Science, X(40): 318–320, February 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/40/318.full.pdf+html;
http://www.jstor.org/
stable/685920.

Watkins:1960:WSE

[Wat60b] J. W. N. Watkins. When are statements empirical? British Journal for the Philosophy of Science, X(40):287–308, February 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/40/287.full.pdf+
html;   http://www.jstor.
org/stable/685916.

Watkins:1962:RPP

[Wat62] J. W. N. Watkins. Reply to Professor Puccetti. British Journal for the Philosophy of Science, XIII(49):53–54, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/49/53.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685226.

Watling:1964:R

[Wat64] John Watling. Reviews. British Journal for the Philosophy of Science, XV(58):164–166, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/58/164.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685260.

Watkins:1965:MCR

[Wat65a] J. W. N. Watkins. Mandelbaum’s critical realism. British Journal for the Philosophy of Science, XVI(63):249–252, November 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XVI/63/249.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/686310.

Watling:1965:Ra

[Wat65b] John Watling. Reviews. British Journal for the Philosophy of Science, XV(60):357–358, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/60/357.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686550.

Watling:1965:Rb

[Wat65c] John Watling. Reviews. British Journal for the Philosophy of Science, XVI(61):61– 63, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XVI/61/61.full.pdf+
html;   http://www.jstor.
org/stable/686143.

Watkins:1971:FPA

[Wat71a] J. W. N. Watkins. Freedom and predictability: an amendment to Mackay. British Journal for the Philosophy of Science, 22 (3):263–275, August 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/3/263.full.pdf+html;
http://www.jstor.org/
stable/686746.

Watkins:1971:R

[Wat71b] J. W. N. Watkins. Reviews.

British Journal for the Philosophy of Science, 22(4):373– 377, November 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
22/4/373.full.pdf+html;
http://www.jstor.org/
stable/686320.

Watkins:1972:R

[Wat72]  J. W. N. Watkins.  Re-

views. British Journal for the Philosophy of Science, 23(1):

83–84, February 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
23/1/83.full.pdf+html;
http://www.jstor.org/
stable/686226.

Watkins:1974:ON

[Wat74] J. W. N. Watkins. Otto Neurath. British Journal for the Philosophy of Science, 25(4): 343–352, December 1974. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
25/4/343.full.pdf+html;
http://www.jstor.org/
stable/686464.

Watkins:1975:MAS

[Wat75] J. W. N. Watkins. Metaphysics and the advancement of science. British Journal for the Philosophy of Science, 26(2):91–121, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/26/2/91.full.pdf+
html;   http://www.jstor.
org/stable/686510.

Watkins:1977:MSM

[Wat77] J. W. N. Watkins. Moritz Schlick and the mind-body problem. British Journal for the Philosophy of Science, 28(4): 369–382, December 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/4/369.full.pdf+html;
http://www.jstor.org/
stable/686880.

Waterlow:1984:R

[Wat84]  Sarah Waterlow.  Reviews.

British Journal for the Philosophy of Science, 35(4):404– 408, December 1984. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
35/4/404.full.pdf+html;
http://www.jstor.org/
stable/687345.

Watkins:1985:SS

[Wat85] John William Nevill Watkins.

Science and scepticism. Hutchinson, London, UK, 1985. ISBN 0-09-158010-2. LCCN ????

Watkins:1987:LRP

[Wat87] John Watkins. Lawson on the Raven Paradox and background knowledge. British Journal for the Philosophy of Science, 38(4): 567–571, December 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
38/4/567.full.pdf+html;
http://www.jstor.org/
stable/687362.

Watt:1989:VLR

[Wat89] D. E. Watt. Not very likely: a reply to Ramsey. British Journal for the Philosophy of Science, 40(2):223–227, June 1989.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/223.full.pdf+html;
http://www.jstor.org/
stable/687513.

Watkins:1991:SRP

[Wat91] John Watkins. Scientific rationality and the problem of induction: Responses to criticisms. British Journal for the Philosophy of Science, 42(3): 343–368, September 1991. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
42/3/343.full.pdf+html;
http://www.jstor.org/
stable/687913.

Watkins:1994:KP

[Wat94] John Watkins. Karl Popper (1902–1994). British Journal for the Philosophy of Science, 45 (4):1089–1090, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
45/4/1089.full.pdf+html;
http://www.jstor.org/
stable/687628.

Watkins:1995:R

[Wat95] John Watkins. Review. British Journal for the Philosophy of Science, 46(4):610–616, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/46/4/610.full.pdf+
html;   http://www.jstor.
org/stable/687902.

Watkins:1999:DNB

[Wat99] J. Watkins. Discussion. A note on Baldwin effect. British Journal for the Philosophy of Science, 50(3):417–423, September 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/417.full.pdf+html;
http://www.jstor.org/
stable/40072239.

Wayne:1995:R

[Way95] Andrew Wayne. Review. British Journal for the Philosophy of Science, 46(4):624–627, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/46/4/624.full.pdf+
html;   http://www.jstor.
org/stable/687905.

Whitrow:1953:PCS

[WB53] G. J. Whitrow and H. Bondi.

Is physical cosmology a science?: a discussion. British Journal for the Philosophy of Science, IV(16):271–283, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/271.full.pdf+html.

Whitrow:1954:PCS

[WB54] G. J. Whitrow and H. Bondi.

Is physical cosmology a science? British Journal for the Philosophy of Science, IV(16): 271–283, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686077.

Wheeler:1999:GRR

[WC99] M. Wheeler and A. Clark. Genic representation: reconciling content and causal complexity. British Journal for the Philosophy of Science, 50 (1):103–135, March 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/1/103.full.pdf+html;
http://www.jstor.org/
stable/40072210.

Wuthrich:2017:WBC

[WC17] Christian Wüthrich and Craig Callender. What becomes of a causal set? British Journal for the Philosophy of Science, 68(3):907–925, September 1, 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/68/3/907/2669639.

Wear:1971:R

[Wea71] Andrew Wear. Reviews. British Journal for the Philosophy of Science, 22(1):57–58, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
22/1/57.full.pdf+html;
http://www.jstor.org/
stable/686646.

Weatherall:2018:RHA

[Wea18] James Owen Weatherall. Regarding the ‘hole argument’. British Journal for the Philosophy of Science, 69(2):329–350, June 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/2/329/2669777.

Webster:1970:R

[Web70] C. Webster. Reviews. British Journal for the Philosophy of Science, 21(3):304–305, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/304.full.pdf+
html;   http://www.jstor.
org/stable/686208.

Weber:2009:CCE

[Web09] Marcel Weber. The crux of crucial experiments: Duhem’s problems and inference to the best explanation. British Journal for the Philosophy of Science, 60(1):19–49, March 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/19.full.pdf+html;
http://www.jstor.org/
stable/25591987.

Weingard:1972:RRP

[Wei72] Robert Weingard. Relativity and the reality of past and future events. British Journal for the Philosophy of Science, 23 (2):119–121, May 1972. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        23/2/119.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686436.

Weingard:1979:GRL

[Wei79] Robert Weingard. General Relativity and the length of the past. British Journal for the Philosophy of Science, 30(2):170–172, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/2/170.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686948.

Weiner:1984:R

[Wei84] Joan Weiner. Reviews. British Journal for the Philosophy of Science, 35(1):90–94, March 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/35/1/90.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687565.

Weirich:1986:EUR

[Wei86] Paul Weirich. Expected utility and risk. British Journal for the Philosophy of Science, 37(4):419–442, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687368.

Weingard:1989:SCC

[Wei89] Robert Weingard. Some comments concerning spin and Relativity. British Journal for the Philosophy of Science, 40 (2):287–288, June 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/2/287.full.pdf+html;
http://www.jstor.org/
stable/687520.

Weinert:1992:VLL

[Wei92] Friedel Weinert. Vicissitudes of laboratory life. British Journal for the Philosophy of Science, 43(3):423–429, September 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
43/3/423.full.pdf+html;
http://www.jstor.org/
stable/687767.

Weintraub:1993:FRB

[Wei93] Ruth Weintraub. Fallibilism and rational belief. British Journal for the Philosophy of Science, 44(2):251–261, June 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/2/251.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687647.

Weiss:1994:R

[Wei94] B. Weiss. Review. British Journal for the Philosophy of Science, 45(3):918–922, September 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/45/3/918.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687802.

Weinstein:2001:AQM

[Wei01a]  Steven Weinstein.  Absolute

quantum mechanics. British Journal for the Philosophy of Science, 52(1):67–73, March 2001. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/52/1/67.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541943.

Weintraub:2001:BP

[Wei01b] Ruth Weintraub. A Bayesian paradox. British Journal for the Philosophy of Science, 52 (1):51–66, March 2001. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        52/1/51.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541942.

Weiskopf:2004:PTC

[Wei04] Daniel A. Weiskopf. The place of time in cognition. British Journal for the Philosophy of Science, 55(1):87–105, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/1/87.full.pdf+html;
http://www.jstor.org/
stable/3541835.

Weisberg:2005:FSF

[Wei05]  Jonathan Weisberg.  Firing

squads and fine-tuning: Sober on the design argument. British Journal for the Philosophy of Science, 56(4):809–821, December 2005. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/56/4/809.full.pdf+
html;   http://www.jstor.
org/stable/3541868.

Weisberg:2007:WM

[Wei07] Michael Weisberg. Who is a modeler? British Journal for the Philosophy of Science, 58 (2):207–233, June 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/2/207.full.pdf+html;
http://www.jstor.org/
stable/30115224.

Weisberg:2009:CHD

[Wei09] Jonathan Weisberg. Commutativity or holism? a dilemma for conditionalizers. British Journal for the Philosophy of Science, 60(4):793–812, December 2009.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/4/793.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25592035.

Weiskopf:2011:FUS

[Wei11] Daniel A. Weiskopf. The functional unity of special science kinds. British Journal for the Philosophy of Science, 62 (2):233–258, June 2011. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        62/2/233.full.pdf+html.

Weisberg:2013:SSU

[Wei13] Michael Weisberg. Simulation and similarity: using models to understand the world. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-993366-9 (hardcover), 0-19-993367-7 (ebook). xviii + 190 pp. LCCN Q175.32.M38 W45 2013.

Weir:2014:BRM

[Wei14] Alan Weir. Book review: Mary Leng, Mathematics and Reality. British Journal for the Philosophy of Science, 65(3):657– 664, September 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/3/657.

Weisberg:2015:UUI

[Wei15] Jonathan Weisberg. Updating, undermining, and independence. British Journal for the Philosophy of Science, 66(1):121–159, March 2015. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/66/1/121.

Weinberger:2019:PSE

[Wei19]  Naftali Weinberger.  Path-

specific effects. British Journal for the Philosophy of Science, 70(1):53–76, March 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/1/
53/4091084.

Wells:1993:PAM

[Wel93] Andrew Wells. Parallel architectures and mental computation. British Journal for the Philosophy of Science, 44(3):531– 542, September 1993. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
44/3/531.full.pdf+html;
http://www.jstor.org/
stable/688021.

Welch:2004:POH

[Wel04] P. D. Welch. On the possibility, or otherwise, of hypercomputation. British Journal for the Philosophy of Science, 55(4):739–746, December 2004.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
55/4/739.full.pdf+html;
http://www.jstor.org/
stable/3541626.

Welch:2008:ECM

[Wel08] P. D. Welch. The extent of computation in Malament–Hogarth spacetimes. British Journal for the Philosophy of Science, 59(4): 659–674, December 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/4/659.full.pdf+html;
http://www.jstor.org/
stable/40072306.

Werndl:2009:WNI

[Wer09] Charlotte Werndl. What are the new implications of chaos for unpredictability? British Journal for the Philosophy of Science, 60 (1):195–220, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/1/195.full.pdf+html;
http://www.jstor.org/
stable/25591993.

Werndl:2010:BRM

[Wer10] Charlotte Werndl. Book review: M. Strevens, Bigger Than Chaos: Understanding Complexity Through Probability. British Journal for the Philosophy of Science, 61(4):875–882, December 2010. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
61/4/875.full.pdf+html.

Werndl:2016:ECD

[Wer16]  Charlotte Werndl.  Editor’s

choice: On defining climate and climate change. British Journal for the Philosophy of Science, 67(2):337–364, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
67/2/337.

Werkhoven:2019:DTH

[Wer19a] Sander Werkhoven. A dispositional theory of health. British Journal for the Philosophy of Science, 70(4):927–952, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/academic.oup.com/bjps/
article/70/4/927/4838240.

Werndl:2019:ICD

[Wer19b] Charlotte Werndl.

Initial-condition dependence and initial-condition uncertainty in climate science. British Journal for the Philosophy of Science, 70(4):953–976, December 2019. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://academic.
oup.com/bjps/article/70/4/
953/5045848.

Westfall:1979:NA

[Wes79a] Richard S. Westfall. Newtonus absolutus. British Journal for the Philosophy of Science, 30 (2):173–182, June 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/2/173.full.pdf+html.

Westfall:1979:RNA

[Wes79b] Richard S. Westfall. Review: Newtonus absolutus. British Journal for the Philosophy of Science, 30(2):173–177, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686949.

Westphal:1997:R

[Wes97] Jonathan Westphal. Review. British Journal for the Philosophy of Science, 48(1):143–148, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/1/143.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687728.

Wettersten:1994:WWP

[Wet94]  John Wettersten.  William

Whewell: Problems of induction vs. problems of rationality. British Journal for the Philosophy of Science, 45(2):716–742, June 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/2/716.full.pdf+
html;   http://www.jstor.
org/stable/687694.

Wagner:2013:BRB

[WF13] Elliott O. Wagner and Michael Franke. Book review: Brian Skyrms, Signals: Evolution, Learning, and Information. British Journal for the Philosophy of Science, 64(4):883–887, December 2013. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/883.full.pdf+html.

Wilhelm:2023:CME

[WHB^{+}23a] Isaac Wilhelm, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Matteo Colombo, Peggy Seriès, Jeffrey A. Barrett, Brian Skyrms, and Porter Williams. Comparing mathematical explanations. British Journal for the Philosophy of Science, 74(1):269–290, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wong:2023:NGA

[WHB^{+}23b] T. Y. William Wong, Remco Heesen, Liam Kofi Bright, Charlotte Werndl, Lauren N. Ross, Matteo Colombo, Peggy Seriès, Jeffrey A. Barrett, Brian Skyrms, Malcolm Forster, and Elliott Sober. The neo-Gouldian argument for evolutionary contingency: Mass extinctions. British Journal for the Philosophy of Science, 73(4): 1093–1124, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wheeler:2009:FCC

[Whe09] Gregory Wheeler. Focused correlation and confirmation. British Journal for the Philosophy of Science, 60(1):79–100, March 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/60/1/79.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/25591989.

Wheeler:2012:OBC

[Whe12] Gregory Wheeler. Objective Bayesian calibration and the problem of non-convex evidence. British Journal for the Philosophy of Science, 63(4):841– 850, December 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/4/841.full.pdf+html.

Whitrow:1950:FD

[Whi50] G. J. Whitrow. On the foundations of dynamics. British Journal for the Philosophy of Science, I(2):92–107, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        I/2/92.full.pdf+html;

http://www.jstor.org/
stable/685805.

Whitaker:1951:TR

[Whi51a] Edmund T. Whitaker. Theories of Relativity. British Journal for the Philosophy of Science, II(5):71–73, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/5/71.full.pdf+html;
http://www.jstor.org/
stable/685613.

Whitrow:1951:LI

[Whi51b] G. J. Whitrow. The Law of Inertia. British Journal for the Philosophy of Science, II (5):58b–58, May 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/5/58-b.full.pdf+html.

Whitrow:1951:RUb

[Whi51c] G. J. Whitrow. Review: [untitled]. British Journal for the Philosophy of Science, II (7):266, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685304.

Whitrow:1951:RUc

[Whi51d] G. J. Whitrow. Review: [untitled]. British Journal for the Philosophy of Science, II(7):266– 268, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685305.

Whitrow:1951:Ra

[Whi51e] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, II(7):262–265, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/7/262.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685302.

Whitrow:1951:Rb

[Whi51f] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, II(7):266a– 266, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/7/266-a.full.pdf+html.

Whitrow:1951:Rc

[Whi51g] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, II(7):266b– 268, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        II/7/266-b.full.pdf+html.

Whitrow:1951:TR

[Whi51h] G. J. Whitrow. Theories of Relativity. British Journal for the Philosophy of Science, II(5):61– 68, May 1951. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/61.full.pdf+
html;   http://www.jstor.
org/stable/685611.

Whitrow:1952:R

[Whi52]  G. J. Whitrow.  Reviews.

British Journal for the Philosophy of Science, III(11):272– 273, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/272.full.pdf+html;
http://www.jstor.org/
stable/685272.

Whitrow:1953:BPM

[Whi53a] G. J. Whitrow. Berkeley’s philosophy of motion. British Journal for the Philosophy of Science, IV(13):37–45, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/13/37.full.pdf+html;
http://www.jstor.org/
stable/686044.

Whittaker:1953:TQJ

[Whi53b] Edmund Whittaker. Are there quantum jumps? British Journal for the Philosophy of Science, III(12):348–349, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/348.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685449.

Whitrow:1954:AU

[Whi54a] G. J. Whitrow. The age of the universe. British Journal for the Philosophy of Science, V(19):215–225, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/19/215.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685717.

Whitrow:1954:MST

[Whi54b] G. J. Whitrow. E. A. Milne’s scales of time. British Journal for the Philosophy of Science, V(18):151, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/18/151.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685173.

Whitrow:1954:R

[Whi54c] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, V(19):266–268, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/19/266.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685724.

Whitrow:1955:AU

[Whi55a] G. J. Whitrow. The age of the universe. British Journal for the Philosophy of Science, V(20):334, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/334.full.pdf+html;
http://www.jstor.org/
stable/685736.

Whitrow:1955:Ra

[Whi55b] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, V(20):344–346, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/344.full.pdf+
html;   http://www.jstor.
org/stable/685741.

Whitrow:1955:Rb

[Whi55c] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, V(20):350–351, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/350.full.pdf+
html;   http://www.jstor.
org/stable/685744.

Whitrow:1955:Rc

[Whi55d] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, V(20):355–356, February 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/20/355.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/685749.

Whitrow:1955:WPS

[Whi55e] G. J. Whitrow. Why physical space has three dimensions. British Journal for the Philosophy of Science, VI(21):13–31, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/21/13.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685571.

Whitrow:1957:Ra

[Whi57a] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, VIII(30):168–170, August 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VIII/30/168.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685798.

Whitrow:1957:Rb

[Whi57b] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, VIII(31):257– 258, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/31/257.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685766.

Whitrow:1958:RU

[Whi58a] G. J. Whitrow. Review: [untitled]. British Journal for the Philosophy of Science, IX(34): 172–173, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685160.

Whitrow:1958:Ra

[Whi58b] G. J. Whitrow. Reviews.

British Journal for the Philosophy of Science, IX(33):64–65, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/64.full.pdf+
html;   http://www.jstor.
org/stable/686104.

Whitrow:1958:Rb

[Whi58c] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, IX(34):166–167, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/166.full.pdf+
html;   http://www.jstor.
org/stable/685155.

Whitrow:1958:Rc

[Whi58d] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, IX(34):172b–173, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        IX/34/172-b.full.pdf+html.

Whitrow:1958:Rd

[Whi58e] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, IX(35):255– 257, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/35/255.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685661.

Whitrow:1959:R

[Whi59] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, X(37):83–85, May 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        X/37/83.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685784.

Whitrow:1962:R

[Whi62]  G. J. Whitrow.  Reviews.

British Journal for the Philosophy of Science, XIII(51):256– 257, November 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/51/256.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685847.

Whitrow:1963:Ra

[Whi63a] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, XIII(52):332– 334, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/332.full.pdf+html;
 http://www.jstor.org/
stable/685337.

Whitrow:1963:Rb

[Whi63b] G. J. Whitrow. Reviews.

British Journal for the Philosophy of Science, XIII(52):344– 345, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIII/52/344.full.pdf+html;
 http://www.jstor.org/
stable/685344.

Whitrow:1968:R

[Whi68] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, 19(3):268–269, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/268.full.pdf+
html;   http://www.jstor.
org/stable/686252.

Whitrow:1969:R

[Whi69] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, 20(4):371–372, December 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

                                        content/20/4/371.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686269.

White:1972:PTP

[Whi72] Alan R. White. The propensity theory of probability. British Journal for the Philosophy of Science, 23(1):35–43, February 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        23/1/35.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686220.

Whitrow:1973:W

[Whi73] G. J. Whitrow. L. L. Whyte 1896–1972. British Journal for the Philosophy of Science, 24 (1):91–92, March 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        24/1/91.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686673.

Whitrow:1975:R

[Whi75] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, 26(4):358–362, December 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/4/358.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686683.

Whitbeck:1978:R

[Whi78a] Caroline Whitbeck. Reviews.

British Journal for the Philosophy of Science, 29(4):399– 406, December 1978. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/4/399.full.pdf+html;
http://www.jstor.org/
stable/687108.

Whitrow:1978:IIP

[Whi78b] G. J. Whitrow. On the impossibility of an infinite past. British Journal for the Philosophy of Science, 29(1):39–46, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/1/39.full.pdf+
html;   http://www.jstor.
org/stable/686390.

Whitrow:1980:R

[Whi80] G. J. Whitrow. Reviews. British Journal for the Philosophy of Science, 31(4):408–411, December 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/31/4/408.full.pdf+
html;   http://www.jstor.
org/stable/687070.

Whitrow:1981:HD

[Whi81] G. J. Whitrow. Herbert Dingle. British Journal for the Philosophy of Science, 32(3):327– 328, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/3/327.full.pdf+html;
http://www.jstor.org/
stable/687600.

Whitrow:1983:R

[Whi83]  G. J. Whitrow.  Reviews.

British Journal for the Philosophy of Science, 34(1):78–84, March 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/1/78.full.pdf+
html;   http://www.jstor.
org/stable/686937.

Whttrow:1961:Ra

[Wht61a] G. J. Whttrow. Reviews. British Journal for the Philosophy of Science, XII(46):170–172, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/170.full.
pdf+html;   http://www.
jstor.org/stable/685500.

Whttrow:1961:Rb

[Wht61b] G. J. Whttrow. Reviews. British Journal for the Philosophy of Science, XII(47):255– 258, November 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/47/255.full.pdf+html;

http://www.jstor.org/
stable/685213.

Whukow:1956:SPS

[Whu56] G. J. Whukow. The study of the philosophy of science. British Journal for the Philosophy of Science, VII(27):189– 205, November 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/189.full.pdf+html;
http://www.jstor.org/
stable/685871.

Whyte:1951:IPP

[Why51a] L. L. Whyte. An interpretation of the present position of the theory of particles. British Journal for the Philosophy of Science, I (4):303–327, February 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/4/303.full.pdf+html;
http://www.jstor.org/
stable/685098.

Whyte:1951:LI

[Why51b] L. L. Whyte. The Law of Inertia. British Journal for the Philosophy of Science, II(5):58a– 58, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/5/58-a.full.pdf+
html;   http://www.jstor.
org/stable/685609.

Whyte:1951:TRa

[Why51c] L. L. Whyte. Theories of Relativity. British Journal for the Philosophy of Science, II(5):68– 71, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/5/68.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685612.

Whyte:1951:TRb

[Why51d] L. L. Whyte. Theories of Relativity. British Journal for the Philosophy of Science, II(5):75– 78, May 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/II/5/75.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685615.

Whyte:1952:AFP

[Why52a] L. L. Whyte. Angles in fundamental physics. British Journal for the Philosophy of Science, III(11):256–258, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/11/256.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685268.

Whyte:1952:ECS

[Why52b] L. L. Whyte. The electric current. A study of the role of time in electron physics. British Journal for the Philosophy of Science, III(11):243–255, November 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/11/243.full.
pdf+html;   http://www.
jstor.org/stable/685267.

Whyte:1952:RU

[Why52c] L. L. Whyte. Review: [untitled]. British Journal for the Philosophy of Science, II (8):326, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686025.

Whyte:1952:R

[Why52d] L. L. Whyte. Reviews. British Journal for the Philosophy of Science, III(9):101–103, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/9/101.full.pdf+
html;   http://www.jstor.
org/stable/685484.

Whyte:1953:GST

[Why53a] L. L. Whyte. Geodesics and the space and time of physical observations. British Journal for the Philosophy of Science, IV(16):337–338, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/337.full.pdf+html.

Whyte:1953:SET

[Why53b] L. L. Whyte. Has a single electron a transit time? British Journal for the Philosophy of Science, III(12):349–350, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/12/349.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685450.

Whyte:1953:LSK

[Why53c] L. L. Whyte. Light signal kinematics. British Journal for the Philosophy of Science, IV(14): 160–161, August 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/14/160.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685991.

Whyte:1954:DP

[Why54a] L. L. Whyte. A dimensionless physics? British Journal for the Philosophy of Science, V(17):1– 17, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/V/17/1.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685948.

Whyte:1954:GST

[Why54b] L. L. Whyte. Geodesics and the space and time of physical observations. British Journal for the Philosophy of Science, IV(16):337–338, February 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686084.

Whyte:1955:NSP

[Why55a] L. L. Whyte. Note on the structural philosophy of organism. British Journal for the Philosophy of Science, V(20): 332–334, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/332.full.pdf+html;
http://www.jstor.org/
stable/685735.

Whyte:1955:OWP

[Why55b] L. L. Whyte. One-way processes in physics and biophysics. British Journal for the Philosophy of Science, VI(22):107–121, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/22/107.full.pdf+
html;   http://www.jstor.
org/stable/685520.

Whyte:1957:RPL

[Why57] L. L. Whyte. On the relation of physical laws to the processes of organisms. British Journal for the Philosophy of Science, VII(28):347–350, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        VII/28/347.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685118.

Whyte:1958:Ra

[Why58a] L. L. Whyte. Reviews. British Journal for the Philosophy of Science, IX(33):85–86, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/33/85.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686115.

Whyte:1958:Rb

[Why58b] L. L. Whyte. Reviews. British Journal for the Philosophy of Science, IX(34):170–171, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IX/34/170.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685158.

Whyte:1958:SQM

[Why58c] L. L. Whyte. The scope of quantum mechanics. British Journal for the Philosophy of Science, IX (34):133–134, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IX/34/133.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685150.

Whyte:1960:R

[Why60] L. L. Whyte. Reviews. British Journal for the Philosophy of Science, XI(43):263–264, November 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/43/263.full.pdf+
html;   http://www.jstor.
org/stable/685322.

Whyte:1963:STC

[Why63] Lancelot Law Whyte. Some thoughts on certainty in physical science. British Journal for the Philosophy of Science, XIV(53):32–38, May 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XIV/53/32.full.pdf+html;
http://www.jstor.org/
stable/685349.

Whyte:1967:R

[Why67] Lancelot Law Whyte. Reviews.

British Journal for the Philosophy of Science, 18(1):79–80, May 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/1/79.full.pdf+
html;   http://www.jstor.
org/stable/686898.

Whyte:1969:R

[Why69]  Lancelot Law Whyte.  Re-

views. British Journal for the Philosophy of Science, 20(1):95, May 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/20/1/95.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686727.

Whyte:1971:R

[Why71]  Lancelot Law Whyte.  Re-

views. British Journal for the Philosophy of Science, 22(1): 75–76, February 1971. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        22/1/75.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686657.

Whyte:1972:R

[Why72] Lancelot Law Whyte. Reviews. British Journal for the Philosophy of Science, 23(3):291–292, August 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/23/3/291.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686692.

Whyte:1992:R

[Why92] J. T. Whyte. Review. British Journal for the Philosophy of Science, 43(2):279–282, June 1992. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/43/2/279.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687878.

Whyte:1993:PC

[Why93] J. T. Whyte. Purpose and content. British Journal for the Philosophy of Science, 44(1):45–60, March 1993. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/44/1/45.full.pdf+
html;   http://www.jstor.
org/stable/687849.

Wigetman:1955:R

[Wig55] William P. D. Wigetman. Reviews. British Journal for the Philosophy of Science, V(20): 342–343, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/20/342.full.pdf+html;
http://www.jstor.org/
stable/685739.

Wightman:1956:NDP

[Wig56] William P. D. Wightman. Note on Descartes and psychosomatic medicine. British Journal for the Philosophy of Science, VII (27):234–235, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/27/234.full.pdf+html;
http://www.jstor.org/
stable/685876.

Wightman:1961:NTH

[Wig61] William P. D. Wightman. Note on the teaching of history and philosophy of science. British Journal for the Philosophy of Science, XI(44):336a–336, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/336-a.full.pdf+html.

Wightman:1970:R

[Wig70]  W. P. D. Wightman.  Re-

views. British Journal for the Philosophy of Science, 21(4): 389–390, November 1970. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
21/4/389.full.pdf+html;
http://www.jstor.org/
stable/686562.

Wilkie:1950:PTR

[Wil50] J. S. Wilkie. The problem of the temporal relation of cause and effect. British Journal for the Philosophy of Science, I(3): 211–229, November 1950. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/3/211.full.pdf+html;
http://www.jstor.org/
stable/685368.

Wilkie:1951:CET

[Wil51] J. S. Wilkie. Causation and explanation in theoretical biology. British Journal for the Philosophy of Science, I(4):273– 290, February 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
I/4/273.full.pdf+html;
http://www.jstor.org/
stable/685096.

Wilkie:1953:R

[Wil53] J. S. Wilkie. Reviews. British Journal for the Philosophy of Science, III(12):377–379, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/III/12/377.full.
pdf+html;   http://www.
jstor.org/stable/685460.

Wilkie:1958:R

[Wil58] J. S. Wilkie. Reviews. British Journal for the Philosophy of Science, IX(34):165–166, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/165.full.pdf+
html;   http://www.jstor.
org/stable/685154.

Williams:1960:Ra

[Wil60a] L. Pearce Williams. Reviews.

British Journal for the Philosophy of Science, XI(42):159–161, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/42/159.full.pdf+
html;   http://www.jstor.
org/stable/685594.

Williams:1960:Rb

[Wil60b] L. Pearce Williams. Reviews. British Journal for the Philosophy of Science, XI(42):162–163, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XI/42/162.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685595.

Wilson:1962:R

[Wil62] N. L. Wilson. Reviews. British Journal for the Philosophy of Science, XIII(50):181–182, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XIII/50/181.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685974.

Wilkins:1964:CR

[Wil64a]  Burleigh  Taylor

Wilkins. Collingwood reconsidered. British Journal for the Philosophy of Science, XV(57):72– 78, May 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/XV/57/72.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685628.

Wilson:1964:CP

[Wil64b] P. R. Wilson. On the confirmation paradox. British Journal for the Philosophy of Science, XV(59):196–199, November 1964. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XV/59/196.full.pdf+
html;   http://www.jstor.
org/stable/686523.

Wilkins:1965:WH

[Wil65] Burleigh Taylor Wilkins. What is history? British Journal for the Philosophy of Science, XVI(61):53–57, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XVI/61/53.full.pdf+html;
http://www.jstor.org/
stable/686141.

Williams:1966:R

[Wil66] L. Pearce Williams. Reviews.

British Journal for the Philosophy of Science, 17(2):159–160, August 1966. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/17/2/159.full.pdf+
html;   http://www.jstor.
org/stable/686351.

Williams:1967:MFB

[Wil67] L. Pearce Williams. Michael Faraday: a biography. British Journal for the Philosophy of Science, 18(3):230–233, November 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/18/3/230.full.pdf+

                                        html;   http://www.jstor.
                                        org/stable/686593.

Wilson:1968:DD

[Wil68a] Fred Wilson. Definition and discovery (I). British Journal for the Philosophy of Science, 18(4):287–303, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        18/4/287.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686358.

Wilson:1968:DDI

[Wil68b] Fred Wilson. Definition and discovery (II). British Journal for the Philosophy of Science, 19(1):43–56, May 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        19/1/43.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686847.

Williams:1969:SAE

[Wil69] P. M. Williams. The structure of acceptance and its evidential basis. British Journal for the Philosophy of Science, 19(4):325–344, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        19/4/325.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686292.

Williams:1970:R

[Wil70]  P. M. Williams.  Reviews.

British Journal for the Philosophy of Science, 21(3):291–298, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/21/3/291.full.pdf+
html;   http://www.jstor.
org/stable/686204.

Williams:1973:LRB

[Wil73a] P. M. Williams. On the logical relations between expressions of different theories. British Journal for the Philosophy of Science, 24(4):357–367, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/4/357.full.pdf+
html;   http://www.jstor.
org/stable/686616.

Williams:1973:Ra

[Wil73b] P. M. Williams. Reviews. British Journal for the Philosophy of Science, 24(3):299–307, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/3/299.full.pdf+
html;   http://www.jstor.
org/stable/686235.

Williams:1973:Rb

[Wil73c] P. M. Williams. Reviews. British Journal for the Philosophy of Science, 24(4):411–413, December 1973. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/24/4/411.full.pdf+
html;   http://www.jstor.
org/stable/686624.

Wilks:1975:PCM

[Wil75a] Yorick Wilks. Putnam and Clarke and mind and body. British Journal for the Philosophy of Science, 26(3):213– 225, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/213.full.pdf+html;
http://www.jstor.org/
stable/686278.

Williams:1975:SPA

[Wil75b] L. Pearce Williams. Should philosophers be allowed to write history? British Journal for the Philosophy of Science, 26(3): 241–253, September 1975. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
26/3/241.full.pdf+html;
http://www.jstor.org/
stable/686282.

Wilks:1976:DD

[Wil76] Yorick Wilks. Dreyfus’s disproofs. British Journal for the Philosophy of Science, 27 (2):177–185, June 1976. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

oxfordjournals.org/content/
27/2/177.full.pdf+html;
http://www.jstor.org/
stable/686169.

Williams:1978:RAB

[Wil78a] L. Pearce Williams. Reply to Agassi and Berkson. British Journal for the Philosophy of Science, 29(3):252, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/3/252-a.full.
pdf+html;   http://www.
jstor.org/stable/687013.

Williams:1978:NTE

[Wil78b] P. M. Williams. On a new theory of epistemic probability. British Journal for the Philosophy of Science, 29(4):375–387, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/4/375.full.pdf+
html;   http://www.jstor.
org/stable/687102.

Wilks:1979:R

[Wil79a] Yorick Wilks. Reviews. British Journal for the Philosophy of Science, 30(2):191–195, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/30/2/191.full.pdf+
html;   http://www.jstor.
org/stable/686953.

Wilson:1979:MCG

[Wil79b] Mark Wilson. Maxwell’s condition — Goodman’s problem. British Journal for the Philosophy of Science, 30(2):107–123, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/30/2/107.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686943.

Wilkes:1980:BS

[Wil80a] Kathleen V. Wilkes. Brain states. British Journal for the Philosophy of Science, 31 (2):111–129, June 1980. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/2/111.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687181.

Williams:1980:BCP

[Wil80b] P. M. Williams. Bayesian conditionalisation and the principle of minimum information. British Journal for the Philosophy of Science, 31(2):131–144, June 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/31/2/131.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687182.

Wilkes:1981:MPP

[Wil81a] Kathleen V. Wilkes. Multiple personality and personal identity. British Journal for the Philosophy of Science, 32(4): 331–348, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/331.full.pdf+html;
http://www.jstor.org/
stable/687304.

Williams:1981:R

[Wil81b] L. Pearce Williams. Reviews.

British Journal for the Philosophy of Science, 32(4):426– 428, December 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/426.full.pdf+html;
http://www.jstor.org/
stable/687316.

Wilkes:1982:Ra

[Wil82a] Kathleen V. Wilkes. Reviews.

British Journal for the Philosophy of Science, 33(2):227–230, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/33/2/227.full.pdf+
html;   http://www.jstor.
org/stable/687285.

Wilkes:1982:Rb

[Wil82b] Kathleen V. Wilkes. Reviews.

British Journal for the Philosophy of Science, 33(2):230–234, June 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/2/230.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687286.

Wilks:1982:R

[Wil82c] Yorick Wilks. Reviews. British Journal for the Philosophy of Science, 33(3):335–336, September 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/33/3/335.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687235.

Wilkes:1983:R

[Wil83] Kathleen V. Wilkes. Reviews. British Journal for the Philosophy of Science, 34(2):175–182, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/34/2/175.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687449.

Wilkes:1984:CI

[Wil84] Kathleen V. Wilkes. Is consciousness important? British Journal for the Philosophy of Science, 35(3):223–243, September 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/35/3/223.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687474.

Wilyte:1991:R

[Wil91] J. T. Wilyte. Review. British Journal for the Philosophy of Science, 42(2):291–294, June 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/42/2/291.full.pdf+
html;   http://www.jstor.
org/stable/688062.

Wilson:1994:CAP

[Wil94] Patrick A. Wilson. Carter on anthropic principle predictions. British Journal for the Philosophy of Science, 45(1):241–253, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/241.full.pdf+
html;   http://www.jstor.
org/stable/687970.

Williamson:1996:R

[Wil96a] Timothy Williamson. Review.

British Journal for the Philosophy of Science, 47(2):331–334, June 1996. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/47/2/331.full.pdf+
html;   http://www.jstor.
org/stable/687955.

Wilson:1996:PR

[Wil96b] Robert A. Wilson. Promiscuous realism. British Journal for the Philosophy of Science, 47(2):303–316, June 1996.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
47/2/303.full.pdf+html;
http://www.jstor.org/
stable/687950.

Williamson:1998:CK

[Wil98] Timothy Williamson. Conditionalizing on knowledge. British Journal for the Philosophy of Science, 49(1):89–121, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/1/89.full.pdf+
html;   http://www.jstor.
org/stable/688145.

Williamson:1999:CAS

[Wil99] J. Williamson. Countable additivity and subjective probability. British Journal for the Philosophy of Science, 50(3):401– 416, September 1999. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
50/3/401.full.pdf+html;
http://www.jstor.org/
stable/40072238.

Williamson:2003:ARS

[Wil03] Jon Williamson. Abduction, reason, and science: Processes of discovery and explanation. British Journal for the Philosophy of Science, 54(2):353–358, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/54/2/353.full.pdf+
html;   http://www.jstor.
org/stable/3541974.

Williamson:2007:II

[Wil07a] Jon Williamson. Inductive influence. British Journal for the Philosophy of Science, 58(4): 689–708, December 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/4/689.full.pdf+html;
http://www.jstor.org/
stable/30115170.

Wilson:2007:NF

[Wil07b]  Jessica Wilson.  Newtonian

forces. British Journal for the Philosophy of Science, 58 (2):173–205, June 2007. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
58/2/173.full.pdf+html;
http://www.jstor.org/
stable/30115223.

Williamson:2009:RG

[Wil09a] Jon Williamson. Response to Glymour. British Journal for the Philosophy of Science, 60(4): 857–860, December 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
60/4/857.full.pdf+html;
http://www.jstor.org/
stable/25592039. See [Gly09].

Wilson:2009:DMM

[Wil09b]  Mark Wilson.  Determinism

and the mystery of the missing physics. British Journal for the Philosophy of Science, 60 (1):173–193, March 2009. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        60/1/173.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/25591992.

Wilson:2010:NRP

[Wil10] Jessica Wilson. Non-reductive physicalism and degrees of freedom. British Journal for the Philosophy of Science, 61(2): 279–311, June 2010. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        61/2/279.full.pdf+html.

Williamson:2012:CCR

[Wil12] Jon Williamson. Calibration and convexity: Response to Gregory Wheeler. British Journal for the Philosophy of Science, 63(4):851–857, December 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        63/4/851.full.pdf+html.

Wilholt:2013:ETS

[Wil13a] Torsten Wilholt. Epistemic trust in science. British Journal for the Philosophy of Science, 64 (2):233–253, June 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/233.full.pdf+html.

Wilson:2013:OPE

[Wil13b] Alastair Wilson. Objective probability in Everettian quantum mechanics. British Journal for the Philosophy of Science, 64(4): 709–737, December 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/4/709.full.pdf+html.

Wilson:2014:BRM

[Wil14a] Alastair Wilson. Book review:

Michael G. Titelbaum, Quitting Certainties: a Bayesian Framework Modeling Degrees of Belief. British Journal for the Philosophy of Science, 65(4): 887–891, December 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
65/4/887.

Wilson:2014:ECS

[Wil14b] Alastair Wilson. Everettian confirmation and Sleeping Beauty. British Journal for the Philosophy of Science, 65(3):573– 598, September 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        65/3/573.

Wilson:2017:QDA

[Wil17] Alastair Wilson. The quantum Doomsday argument. British Journal for the Philosophy of Science, 68(2):68– ??, June 2017. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

                                        //academic.oup.com/bjps/
                                        article/68/2/597/3044189/
                                        The-Quantum-Doomsday-
                                        Argument.

Williams:2019:SRM

[Wil19] Porter Williams. Scientific realism made effective. British Journal for the Philosophy of Science, 70(1):209–237, March 2019. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://academic.
                                        oup.com/bjps/article/70/1/
                                        209/4098116.

Williams:2022:EIB

[Wil22] Daniel Williams. Epistemic irrationality in the Bayesian brain. British Journal for the Philosophy of Science, 72(4):913–938, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Williamson:2023:BAE

[Wil23a] Jon Williamson. A Bayesian account of establishing. British Journal for the Philosophy of Science, 73 (4):903–925, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wills:2023:CPI

[Wil23b] James Wills. Classical particle indistinguishability, precisely. British Journal for the Philosophy of Science, 74(2):335–358, ???? 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Winnie:1967:IDT

[Win67] John A. Winnie. The implicit definition of theoretical terms. British Journal for the Philosophy of Science, 18(3): 223–229, November 1967. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/3/223.full.pdf+html;
http://www.jstor.org/
stable/686592.

Winning:2021:CMC

[Win21a] Jason Winning. Corrigendum to:

Mechanistic causation and constraints: Perspectival parts and powers, non-perspectival modal patterns. British Journal for the Philosophy of Science, 72(1): ??, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic). See [Win21b].

Winning:2021:MCC

[Win21b] Jason Winning. Mechanistic causation and constraints: Perspectival parts and powers, non-perspectival modal patterns. British Journal for the Philosophy of Science, 71(4):1385–1409, ???? 2021. CODEN ????

ISSN 0007-0882 (print), 1464-3537 (electronic). See corrigendum [Win21a].

Wisdom:1950:R

[Wis50] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, I(3):248–249, November 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/3/248.full.pdf+
html;   http://www.jstor.
org/stable/685371.

Wisdom:1951:HC

[Wis51] J. O. Wisdom. The hypothesis of cybernetics. British Journal for the Philosophy of Science, II(5):1–24, May 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/5/1.full.pdf+html;
http://www.jstor.org/
stable/685605.

Wisdom:1952:NMM

[Wis52a] J. O. Wisdom. A new model for the mind-body relationship. British Journal for the Philosophy of Science, II(8):295– 301, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/8/295.full.pdf+html;
http://www.jstor.org/
stable/686018.

Wisdom:1952:R

[Wis52b] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, III(9):97–99, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/9/97.full.pdf+html;
http://www.jstor.org/
stable/685482.

Wisdom:1953:BCI

[Wis53a]  J. O. Wisdom.  Berkeley’s

criticism of the infinitesimal. British Journal for the Philosophy of Science, IV(13):22–25, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IV/13/22.full.pdf+
html;   http://www.jstor.
org/stable/686042.

Wisdom:1953:OBL

[Wis53b] J. O. Wisdom. An outline of Berkeley’s life. British Journal for the Philosophy of Science, IV(13):78–87, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/13/78.full.pdf+html;
http://www.jstor.org/
stable/686047.

Wisdom:1953:Ra

[Wis53c]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, IV(13):88–91, May 1953. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/IV/13/88.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686048.

Wisdom:1953:Rb

[Wis53d] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, IV(13):91, May 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/13/91.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686049.

Wisdom:1953:Rc

[Wis53e]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, IV(15):260– 262, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        IV/15/260.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685909.

Wisdom:1953:Rd

[Wis53f] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, IV(16):353, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        IV/16/353.full.pdf+html.

Wisdom:1953:RM

[Wis53g] J. O. Wisdom. Reviews: Metabiology. British Journal for the Philosophy of Science, IV(16): 339–344, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/16/339.full.pdf+html.

Wisdom:1954:RM

[Wis54a] J. O. Wisdom. Review: Metabiology. British Journal for the Philosophy of Science, IV(16): 339–344, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686085.

Wisdom:1954:RU

[Wis54b] J. O. Wisdom. Review: [untitled]. British Journal for the Philosophy of Science, IV (16):353, February 1954. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686091.

Wisdom:1954:Ra

[Wis54c] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, V(17):87, May 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/17/87.full.pdf+html;
http://www.jstor.org/
stable/685960.

Wisdom:1954:Rb

[Wis54d] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, V(18):171–172, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/18/171.full.pdf+
html;   http://www.jstor.
org/stable/685181.

Wisdom:1954:Rc

[Wis54e] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, V(18):174, August 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
V/18/174.full.pdf+html;
http://www.jstor.org/
stable/685184.

Wisdom:1954:Rd

[Wis54f] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, V(19):263–266, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/19/263.full.pdf+
html;   http://www.jstor.
org/stable/685723.

Wisdom:1954:Re

[Wis54g] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, V(19):269–270, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/V/19/269.full.pdf+
html;   http://www.jstor.
org/stable/685726.

Wisdom:1955:Ra

[Wis55a] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, VI(22):171–172, August 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/22/171.full.pdf+
html;   http://www.jstor.
org/stable/685527.

Wisdom:1955:Rb

[Wis55b] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, VI(23):257, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VI/23/257.full.pdf+
html;   http://www.jstor.
org/stable/685690.

Wisdom:1956:PAT

[Wis56a] J. O. Wisdom. Psycho-analytic technology. British Journal for the Philosophy of Science, VII (25):13–28, May 1956. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/25/13.full.pdf+html;
http://www.jstor.org/
stable/685933.

Wisdom:1956:Ra

[Wis56b] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, VI(24):351, February 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VI/24/351.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/685423.

Wisdom:1956:Rb

[Wis56c]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, VII(25):105–110, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/25/105.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685941.

Wisdom:1956:Rc

[Wis56d] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, VII(25):113–114, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/25/113.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685943.

Wisdom:1956:Rd

[Wis56e]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, VII(26):175, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VII/26/175.full.
pdf+html;   http://www.
jstor.org/stable/685702.

Wisdom:1957:RU

[Wis57a] J. O. Wisdom. Review: [untitled]. British Journal for the Philosophy of Science, VII (28):363, February 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/685126.

Wisdom:1957:Ra

[Wis57b] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, VII (28):363b–363, February 1957. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VII/28/363-b.full.pdf+html.

Wisdom:1957:Rb

[Wis57c]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, VIII(31):256– 257, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/31/256.full.pdf+html;
http://www.jstor.org/
stable/685765.

Wisdom:1958:RDD

[Wis58a] J. O. Wisdom. A reply to Dr Das’s criticisms. British Journal for the Philosophy of Science, VIII(32):325–328, February 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/VIII/32/325.full.
pdf+html;   http://www.
jstor.org/stable/685860.

Wisdom:1958:Ra

[Wis58b] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, IX(33):81–84, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/81.full.pdf+
html;   http://www.jstor.
org/stable/686113.

Wisdom:1958:Rb

[Wis58c]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, IX(33):84–85, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/84.full.pdf+
html;   http://www.jstor.
org/stable/686114.

Wisdom:1958:Rc

[Wis58d] J. O. Wisdom. Reviews. British Journal for the Philosophy of Science, IX(34):177, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/177.full.pdf+

html;   http://www.jstor.
org/stable/685163.

Wisdom:1958:Rd

[Wis58e]  J. O. Wisdom.  Reviews.

British Journal for the Philosophy of Science, IX(35):258– 261, November 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/35/258.full.pdf+html;
http://www.jstor.org/
stable/685662.

Wisdom:1958:SB

[Wis58f] J. O. Wisdom. The sceptic at bay. British Journal for the Philosophy of Science, IX(34): 159–163, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/34/159.full.pdf+html;
http://www.jstor.org/
stable/685152.

Wisdom:1962:BCP

[Wis62] J. O. Wisdom. On Broad’s contribution to philosophy. British Journal for the Philosophy of Science, XIII(49):70– 75, May 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/49/70.full.
pdf+html;   http://www.
jstor.org/stable/685231.

Wisdom:1963:MCD

[Wis63a] J. O. Wisdom. Mid-century developments within psychoanalytic theory. British Journal for the Philosophy of Science, XIV(53):54–63, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIV/53/54.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685353.

Wisdom:1963:RIL

[Wis63b] J. O. Wisdom. The refutability of ‘ irrefutable ’ laws. British Journal for the Philosophy of Science, XIII(52):303– 306, February 1963. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XIII/52/303.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685327.

Wisdom:1965:WC

[Wis65] J. O. Wisdom. A word to colleagues. British Journal for the Philosophy of Science, XV (60):368b–368, February 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        XV/60/368-b.full.pdf+html.

Withers:1950:R

[Wit50] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, I(2):166–171, August 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/2/166.full.pdf+
html;   http://www.jstor.
org/stable/685809.

Withers:1951:R

[Wit51] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, II(7):265– 266, November 1951. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/265.full.pdf+html;
http://www.jstor.org/
stable/685303.

Withers:1952:R

[Wit52] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, II(8):334– 337, February 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/8/334.full.pdf+html;
http://www.jstor.org/
stable/686032.

Withers:1953:R

[Wit53] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, III(12):386– 388, February 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/12/386.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/685465.

Withers:1955:R

[Wit55] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, V(20):357– 358, February 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        V/20/357.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685751.

Withers:1956:R

[Wit56] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, VII(25):110–113, May 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/VII/25/110.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685942.

Withers:1957:Ra

[Wit57a] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, VIII(30): 171–172, August 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VIII/30/171.full.pdf+html;
                                         http://www.jstor.org/
                                        stable/685799.

Withers:1957:Rb

[Wit57b] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, VIII(31):252– 253, November 1957. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/31/252.full.pdf+html;
http://www.jstor.org/
stable/685762.

Withers:1958:R

[Wit58a] R. J. F. Withers. Reviews.

British Journal for the Philosophy of Science, IX(33):67–68, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/67.full.pdf+
html;   http://www.jstor.
org/stable/686106.

Wittgenstein:1958:BBB

[Wit58b] Ludwig Wittgenstein. The Blue and Brown Books. Basil Blackwell, Oxford, UK, 1958. xiv + 185 pp.

Wittgenstein:1958:PSP

[Wit58c] Ludwig Wittgenstein. Preliminary studies for the “Philosophical investigations”, generally known as the blue and brown books. Harper, New York, NY, USA, 1958. xiv + 185 pp.

Withers:1959:ESS

[Wit59]  R. F. J. Withers.  Episte-

mology and scientific strategy. British Journal for the Philosophy of Science, X(38):89–102, August 1959. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/89.full.pdf+
html;   http://www.jstor.
org/stable/686001.

Withers:1961:MPR

[Wit61a] R. F. J. Withers. A methodological problem in rheology: III: Explanations and models in scientific theory construction. British Journal for the Philosophy of Science, XI(44): 280–288, February 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/44/280.full.pdf+html.

Withers:1961:Ra

[Wit61b] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, XII(45):83–84, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/45/83.full.pdf+
html;   http://www.jstor.
org/stable/685549.

Withers:1961:Rb

[Wit61c] R. F. J. Withers. Reviews. British Journal for the Philosophy of Science, XII(46):167–170, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XII/46/167.full.

pdf+html;   http://www.
jstor.org/stable/685499.

Withers:1961:Rc

[Wit61d] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, XII(47):258– 259, November 1961. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/47/258.full.pdf+html;
http://www.jstor.org/
stable/685214.

Withers:1962:Ra

[Wit62a] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, XII(48):346– 347, February 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/48/346.full.pdf+html;
http://www.jstor.org/
stable/685405.

Withers:1962:Rb

[Wit62b] R. F. J. Withers. Reviews.

British Journal for the Philosophy of Science, XIII(50):183, August 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XIII/50/183.full.
pdf+html;   http://www.
jstor.org/stable/685976.

Witmer:2000:LOP

[Wit00]  D. G. Witmer.  Locating

the overdetermination problem.

British Journal for the Philosophy of Science, 51(2):273–286, June 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/51/2/273.full.pdf+
html;   http://www.jstor.
org/stable/3541805.

Wilhelm:2022:TTT

[WMR^{+}22] Isaac Wilhelm, Lukas J. Meier, Mark Rubin, Bendik Hellem Aaby, Grant Ramsey, and Isaac Wilhelm. Typical: a theory of typicality and typicality explanation. British Journal for the Philosophy of Science, 73(2): 561–581, ???? 2022. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Weatherall:2021:HBS

[WOB21] James Owen Weatherall, Cailin O’Connor, and Justin P. Bruner. How to beat science and influence people: Policymakers and propaganda in epistemic networks. British Journal for the Philosophy of Science, 71(4): 1157–1186, ???? 2021. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wojcicki:1978:R

[Wój78] Ryszard Wójcicki. Reviews. British Journal for the Philosophy of Science, 29(2):197–201, June 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/2/197.full.pdf+

html;   http://www.jstor.
org/stable/687130.

Wollaston:1959:R

[Wol59] Arthur Wollaston. Reviews.

British Journal for the Philosophy of Science, IX(36):339– 340, February 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IX/36/339.full.pdf+html;
http://www.jstor.org/
stable/685200.

Wolff:2012:DOD

[Wol12] Johanna Wolff. Do objects depend on structures? British Journal for the Philosophy of Science, 63(3):607–625, September 2012. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
63/3/607.full.pdf+html.

Woodger:1951:Ra

[Woo51a] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, I(4):332–333, February 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/I/4/332.full.pdf+
html;   http://www.jstor.
org/stable/685100.

Woodger:1951:Rb

[Woo51b] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, II(6):168–169, October 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/6/168.full.pdf+
html;   http://www.jstor.
org/stable/685512.

Woodger:1951:SP

[Woo51c] J. H. Woodger. Science without properties. British Journal for the Philosophy of Science, II(7):193–216, November 1951. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
II/7/193.full.pdf+html;
http://www.jstor.org/
stable/685292.

Woodger:1952:BM

[Woo52a] J. H. Woodger. From biology to mathematics. British Journal for the Philosophy of Science, III(9):1–21, May 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/9/1.full.pdf+html;
http://www.jstor.org/
stable/685471.

Woodger:1952:Ra

[Woo52b] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, II(8):337, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/

content/II/8/337.full.pdf+
html;   http://www.jstor.
org/stable/686033.

Woodger:1952:Rb

[Woo52c] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, II(8):338–342, February 1952. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/II/8/338.full.pdf+
html;   http://www.jstor.
org/stable/686034.

Woodger:1952:Rc

[Woo52d] J. H. Woodger. Reviews.

British Journal for the Philosophy of Science, III(11):293– 294, November 1952. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
III/11/293.full.pdf+html;
http://www.jstor.org/
stable/685286.

Woodger:1953:R

[Woo53a] J. H. Woodger. Reviews.

British Journal for the Philosophy of Science, IV(15):251– 252, November 1953. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
IV/15/251.full.pdf+html;
http://www.jstor.org/
stable/685904.

Woodger:1953:WDW

[Woo53b] J. H. Woodger. What do we mean by ‘inborn’? British Journal for the Philosophy of Science, III(12):319–326, February 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/III/12/319.full.
                                        pdf+html;   http://www.
                                        jstor.org/stable/685446.

Woodger:1956:RPH

[Woo56] J. H. Woodger. A reply to Professor Haldane. British Journal for the Philosophy of Science, VII(26):149–155, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        VII/26/149.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/685698.

Woodger:1958:RU

[Woo58a] J. H. Woodger. Review: [untitled]. British Journal for the Philosophy of Science, IX (34):172, August 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/685159.

Woodger:1958:Ra

[Woo58b] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, VIII(32):348– 349, February 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
VIII/32/348.full.pdf+html;
http://www.jstor.org/
stable/685866.

Woodger:1958:Rb

[Woo58c] J. H. Woodger. Reviews.

British Journal for the Philosophy of Science, IX(33):70–72, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/33/70.full.pdf+
html;   http://www.jstor.
org/stable/686108.

Woodger:1958:Rc

[Woo58d] J. H. Woodger. Reviews. British Journal for the Philosophy of Science, IX(34):172a–172, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/
IX/34/172-a.full.pdf+html.

Woodger:1959:R

[Woo59]  J. H. Woodger.  Reviews.

British Journal for the Philosophy of Science, X(38):157–159, August 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/X/38/157.full.pdf+
html;   http://www.jstor.
org/stable/686007.

Woodger:1960:BP

[Woo60a] J. H. Woodger. Biology and physics. British Journal for the Philosophy of Science, XI (42):89–100, August 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XI/42/89.full.pdf+html;
http://www.jstor.org/
stable/685581.

Woodger:1960:SP

[Woo60b] J. H. Woodger. Science and persons. British Journal for the Philosophy of Science, XI(41):65–71, May 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/XI/41/65.full.pdf+
html;   http://www.jstor.
org/stable/685823.

Woodward:1979:SE

[Woo79] James Woodward. Scientific explanation. British Journal for the Philosophy of Science, 30 (1):41–67, ???? 1979. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
30/1/41.full.pdf+html;
http://www.jstor.org/
stable/687416.

Woodward:1990:LC

[Woo90] J. Woodward. Laws and causes. British Journal for the Philosophy of Science, 41(4):553– 573, December 1990. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
41/4/553.full.pdf+html;
http://www.jstor.org/
stable/687865.

Woodward:1994:R

[Woo94]  James Woodward.  Review.

British Journal for the Philosophy of Science, 45(1):353–374, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/45/1/353.full.pdf+
html;   http://www.jstor.
org/stable/687978.

Woodfield:1998:R

[Woo98] Andrew Woodfield. Review.

British Journal for the Philosophy of Science, 49(1):168–174, March 1998. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/49/1/168.full.pdf+
html;   http://www.jstor.
org/stable/688149.

Woodward:2000:EIS

[Woo00a] J. Woodward. Explanation and invariance in the special sciences. British Journal for the Philosophy of Science, 51 (2):197–254, June 2000. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
51/2/197.full.pdf+html.

Woolf:2000:BRB

[Woo00b] R. Woolf. Book review: Cause and Explanation in Ancient Greek Thought. R. J. Hankinson. British Journal for the Philosophy of Science, 51(3):545–547, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        51/3/545.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/3541786.

Woodward:2003:MTH

[Woo03] James Woodward. Making things happen: a theory of causal explanation. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-515527-0. vi + 410 pp. LCCN BD541 .W64 2003.

Woodward:2014:SBS

[Woo14] James Woodward. Simplicity in the best systems account of laws of nature. British Journal for the Philosophy of Science, 65 (1):91–123, January 2014. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        65/1/91.full.pdf+html.

Worrall:1984:UI

[Wor84] John Worrall. An unreal image. British Journal for the Philosophy of Science, 35(1):65–80, March 1984. CODEN BJPIA5.

ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/35/1/65.full.pdf+
html;   http://www.jstor.
org/stable/687561.

Worrall:1985:R

[Wor85] John Worrall. Reviews. British Journal for the Philosophy of Science, 36(1):81–85, March 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/36/1/81.full.pdf+
html;   http://www.jstor.
org/stable/687160.

Worrall:1988:VFM

[Wor88] John Worrall. The value of a fixed methodology. British Journal for the Philosophy of Science, 39(2):263–275, June 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
39/2/263.full.pdf+html;
http://www.jstor.org/
stable/687271.

Worrall:1989:FID

[Wor89]  John Worrall.  Fix it and

be damned: a reply to Laudan. British Journal for the Philosophy of Science, 40(3):376– 388, September 1989. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/3/376.full.pdf+html;

                                        http://www.jstor.org/
                                        stable/687785.

Worrall:1995:R

[Wor95] John Worrall. Review. British Journal for the Philosophy of Science, 46(4):621–623, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/46/4/621.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687904.

Worrall:1999:OJW

[Wor99] J. Worrall. Obituary. John Watkins (1924–1999). British Journal for the Philosophy of Science, 50(4):787–789, December 1999. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/50/4/787.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/40072270.

Worrall:2000:SLD

[Wor00] J. Worrall. The scope, limits, and distinctiveness of the method of ’deduction from the phenomena’: some lessons from Newton’s ’demonstrations’ in optics. British Journal for the Philosophy of Science, 51(1):45–80, March 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/51/1/45.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/3541748.

Worrall:2007:WTN

[Wor07] John Worrall. Why there’s no cause to randomize. British Journal for the Philosophy of Science, 58(3):451–488, September 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/58/3/451.full.pdf+
html;   http://www.jstor.
org/stable/30115184.

Wray:2011:KES

[Wra11] K. Brad Wray. Kuhn’s evolutionary social epistemology. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-01223-6 (hardcover). xiii + 229 pp. LCCN Q175 .W78

2011. URL http://assets.
cambridge.org/97811070/
12233/cover/9781107012233.
jpg.

Wright:1968:CAT

[Wri68] Larry Wright. The case against teleological reductionism. British Journal for the Philosophy of Science, 19(3):211–223, November 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/19/3/211.full.pdf+
html;   http://www.jstor.
org/stable/686241.

Wright:1983:WLR

[Wri83] Crispin Wright. What the liar really says. British Journal for the Philosophy of Science, 34(3):277–288, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
34/3/277.full.pdf+html;
http://www.jstor.org/
stable/687325.

Wright:1986:BZN

[Wri86] Edmond Wright. Ben-Zeev on the non-epistemic. British Journal for the Philosophy of Science, 37(3):351–359, September 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687293.

Wright:2018:ADE

[Wri18] Jessey Wright. The analysis of data and the evidential scope of neuroimaging results. British Journal for the Philosophy of Science, 69(4):1179–1203, December 1, 2018. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL https:

//academic.oup.com/bjps/
article/69/4/1179/4079370.

Williamson:2023:LAC

[WS23] Timothy L. Williamson and Alexander Sandgren. Lawabiding causal decision theory. British Journal for the Philosophy of Science, 74(4):899–920, December 2023. CODEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Wallace:2010:QMS

[WT10] David Wallace and Christopher G. Timpson. Quantum mechanics on spacetime I: Spacetime state realism. British Journal for the Philosophy of Science, 61(4):697–727, December 2010. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
61/4/697.full.pdf+html.

Woodward:1972:IMP

[WY72] James F. Woodward and Wolfgang Yourgrau. The incompatibility of Mach’s Principle and the Principle of Equivalence in current gravitation theory. British Journal for the Philosophy of Science, 23(2):111–116, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/23/2/111.full.pdf+
html;   http://www.
jstor.org/stable/686434. See

comments [New73] and rejoinder [WY73].

Woodward:1973:MPE

[WY73] James F. Woodward and Wolfgang Yourgrau. Mach’s Principle, the Equivalence Principle and Gravitation: a Rejoinder to Newburgh. British Journal for the Philosophy of Science, 24(3):264–270, September 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

                                        24/3/264.full.pdf+html;
                                        http://www.
                                        jstor.org/stable/686231. See

[New73].

Woodward:1975:MPM

[WY75] James F. Woodward and Wolfgang Yourgrau. Mach’s Principle: Micro- or macrophysical? British Journal for the Philosophy of Science, 26(2):137–141, June 1975. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/26/2/137.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/686513.

Wykstra:1976:ESP

[Wyk76] Steve Wykstra. On Einstein’s Second Postulate. British Journal for the Philosophy of Science, 27(3):259–261, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        27/3/259.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/686123.

Young:1968:RU

[YHF68] Robert M. Young, Mary Hesse, and Arthur I. Fine. Review: [untitled]. British Journal for the Philosophy of Science, 18(4): 325–339, February 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/686363.

Yolton:1959:E

[Yol59] John W. Yolton. Explanation. British Journal for the Philosophy of Science, X(39): 194–208, November 1959. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
X/39/194.full.pdf+html;
http://www.jstor.org/
stable/686053.

Yoshida:1978:R

[Yos78] Ron Yoshida. Reviews. British Journal for the Philosophy of Science, 29(3):285–289, September 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/29/3/285.full.pdf+
html;   http://www.jstor.
org/stable/687020.

Yoshida:1981:RR

[Yos81] Ron Yoshida. Reduction as replacement. British Journal for the Philosophy of Science, 32(4):400–410, December 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
32/4/400.full.pdf+html;
http://www.jstor.org/
stable/687311.

Yoshimi:2012:SDS

[Yos12] Jeffrey Yoshimi. Supervenience, dynamical systems theory, and non-reductive physicalism. British Journal for the Philosophy of Science, 63 (2):373–398, June 2012. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
63/2/373.full.pdf+html.

Yourgrau:1958:R

[You58] Wolfgang Yourgrau. Reviews. British Journal for the Philosophy of Science, IX(34):164–165, August 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/IX/34/164.full.pdf+
html;   http://www.jstor.
org/stable/685153.

Yourgrau:1961:CD

[You61] Wolfgang Yourgrau. Challenge to dualism. British Journal for the Philosophy of Science, XII(46):158–166, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
XII/46/158.full.pdf+html;
http://www.jstor.org/
stable/685498.

Young:1966:RU

[You66] Robert M. Young. Review: [untitled]. British Journal for the Philosophy of Science, 17 (1):73–76, May 1966. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/686407.

Young:1968:R

[You68] Robert M. Young. Reviews.

British Journal for the Philosophy of Science, 18(4):325– 330, February 1968. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
18/4/325.full.pdf+html.

Yourgrau:1986:TAD

[You86] Palle Yourgrau. On time and actuality: The dilemma of privileged position. British Journal for the Philosophy of Science, 37(4):405–417, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687367.

Yudell:2013:LCA

[Yud13] Zanja Yudell. Lange’s challenge: Accounting for metalaws. British Journal for the Philosophy of Science, 64(2): 347–369, June 2013. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/347.full.pdf+html.

Yudell:2016:BRM

[Yud16] Zanja Yudell. Book review: Matteo Morganti, Combining Science and Metaphysics: Contemporary Physics, Conceptual Revision, and Common Sense. British Journal for the Philosophy of Science, 67(2):657–661, June 2016. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/67/2/657.

Yuxin:1990:LMD

[Yux90] Zheng Yuxin. From the logic of mathematical discovery to the methodology of scientific research programmes. British Journal for the Philosophy of Science, 41(3):377–399, September 1990. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/41/3/377.full.pdf+
html;   http://www.jstor.
org/stable/687826.

Zahar:1973:WDEa

[Zah73a] Elie Zahar. Why did Einstein’s programme supersede Lorentz’s? (I). British Journal for the Philosophy of Science, 24(2): 95–123, June 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
24/2/95.full.pdf+html.

Zahar:1973:WDEb

[Zah73b] Elie Zahar. Why did Einstein’s programme supersede Lorentz’s? (II). British Journal for the Philosophy of Science, 24(3):223– 262, September 1973. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/

24/3/223.full.pdf+html;
http://links.jstor.org/
sici?sici=0007-0882%
28197309%2924%3A3%3C223%
3AWDEPSL%3E2.0.CO%3B2-Z;
http://www.jstor.org/
stable/686229.

Zahar:1977:MER

[Zah77] Elie Zahar. Mach, Einstein, and the rise of modern science. British Journal for the Philosophy of Science, 28(3):195– 213, September 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/3/195.full.pdf+html;
http://www.jstor.org/
stable/686806.

Zahar:1978:EDL

[Zah78] Elie Zahar. Einstein’s debt to Lorentz: a reply to Feyerabend and Miller. British Journal for the Philosophy of Science, 29(1):49–60, March 1978. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
29/1/49.full.pdf+html;
http://www.jstor.org/
stable/686392.

Zahar:1980:EMR

[Zah80] Elie Zahar. Einstein, Meyerson and the role of mathematics in physical discovery. British Journal for the Philosophy of Science, 31(1):1–43, March 1980. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        31/1/1.full.pdf+html;
                                        http://links.jstor.org/
                                        sici?sici=0007-0882%
                                        28198003%2931%3A1%3C1%
                                        3AEMATRO%3E2.0.CO%3B2-B;
                                        http://www.jstor.org/
                                        stable/687250.

Zahar:1981:STA

[Zah81] Elie Zahar. Second thoughts about Machian positivism: a reply to Feyerabend. British Journal for the Philosophy of Science, 32(3):267–276, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        32/3/267.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687590.

Zahar:1982:FOE

[Zah82a] Elie Zahar. Feyerabend on observation and empirical content. British Journal for the Philosophy of Science, 33(4): 397–408, December 1982. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        33/4/397.full.pdf+html.

Zahar:1982:RFO

[Zah82b] Elie Zahar. Review: Feyerabend on observation and empirical content. British Journal for the Philosophy of Science, 33(4):397–409, December 1982. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://www.
jstor.org/stable/687171.

Zahar:1983:PLC

[Zah83a] E. G. Zahar. The Popper– Lakatos controversy in the light of ’Die Beiden Grundprobleme Der Erkenntnistheorie’ (German [The two basic problems of the theory of knowledge]). British Journal for the Philosophy of Science, 34(2):149–171, June 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/2/149.full.pdf+
html;   http://www.jstor.
org/stable/687447.

Zahar:1983:LDP

[Zah83b] Elie Zahar. Logic of discovery or psychology of invention? British Journal for the Philosophy of Science, 34(3):243–261, September 1983. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/34/3/243.full.pdf+
html;   http://www.jstor.
org/stable/687322.

Zahar:1987:MRD

[Zah87] Elie Zahar. Meyerson’s ‘relativistic deduction’: Einstein versus Hegel. British Journal for the Philosophy of Science, 38 (1):93–106, March 1987. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.

                                        oxfordjournals.org/content/
                                        38/1/93.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687084.

Zahar:1991:NAR

[Zah91] Elie Zahar. Natural axiomatization: a revision of ‘Wajsberg’s requirement’. British Journal for the Philosophy of Science, 42(3):391–396, September 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        42/3/391.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687915.

Zahar:1997:R

[Zah97]  Elie Zahar.  Review.

British Journal for the Philosophy of Science, 48(1):132–139, March 1997. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

                                        /bjps.oxfordjournals.org/
                                        content/48/1/132.full.pdf+
                                        html;   http://www.jstor.
                                        org/stable/687726.

Zaret:1979:ASC

[Zar79] David Zaret. Absolute space and conventionalism. British Journal for the Philosophy of Science, 30(3):211–226, September 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

                                        tronic). URL http://bjps.
                                        oxfordjournals.org/content/
                                        30/3/211.full.pdf+html;
                                        http://www.jstor.org/
                                        stable/687136.

Zemach:1987:TLL

[Zem87] E. M. Zemach. Are there logical limits for science? British Journal for the Philosophy of Science, 38(4):527–532, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/38/4/527.full.pdf+
html;   http://www.jstor.
org/stable/687357.

Zenzen:1977:PGF

[Zen77] Michael J. Zenzen. Popper, Grünbaum and de facto irreversibility. British Journal for the Philosophy of Science, 28(4): 313–324, December 1977. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
28/4/313.full.pdf+html;
http://www.jstor.org/
stable/686876.

Zhang:2021:UCI

[ZG21] Kun Zhang and Madelyn R. K.

Glymour. Unmixing for causal inference: Thoughts on McCaffrey and Danks. British Journal for the Philosophy of Science, 71 (4):1319–1330, ???? 2021. CO-DEN ???? ISSN 0007-0882 (print), 1464-3537 (electronic).

Zhang:2013:CTO

[Zha13] Jiji Zhang. A comparison of three Occam’s razors for Markovian causal models. British Journal for the Philosophy of Science, 64(2):423–448, June 2013.

CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
64/2/423.full.pdf+html.

Zimba:2008:ID

[Zim08] Jason Zimba. Inertia and determinism. British Journal for the Philosophy of Science, 59(3): 417–428, September 2008. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
59/3/417.full.pdf+html;
http://www.jstor.org/
stable/40072293.

Zucker:1988:DRL

[Zuc88] Arthur Zucker. Davy refuted Lavoisier not Lakatos. British Journal for the Philosophy of Science, 39(4):537–540, November 1988. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/
content/39/4/537.full.pdf+
html;   http://www.jstor.
org/stable/687439.

Zycinski:1989:TMS

[Zyc89] Joseph Zycinski. Taking mathematics seriously? British Journal for the Philosophy of Science, 40(1):77–82, March 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.
oxfordjournals.org/content/
40/1/77.full.pdf+html;

REFERENCES 640

http://www.jstor.org/
stable/687465.